A Complete Bibliography of Publications in the Journal of Cell Biology: 1995–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

1 [BZB^{+}96, BGR^{+}95, DNB^{+}96, GHP^{+}96, SMH98, SYO95]. 2 [GRS^{+}99, ROG^{+}96, TBR^{+}96, WBG^{+}97]. 3 [AJS96, BDE^{+}95, BGN^{+}97, CGSY^{+}99, CMC^{+}98, Dav95, GPRS^{+}95, RWO95, SOE^{+}95, SMFC95, SDTE95, SCB^{+}97, TCM96, VSGF96]. 4 [SEC^{+}95]. 5 [LCS96, SMA96]. 6 [CBM^{+}98, KDvFP96, KKG^{+}98, PHB96, dPF95, RSJK95, SKR96]. < 100 [KE98]. ^{+} [BG97, FWG^{+}96, GHS98, MCS^{+}97b, PBCS98, THS^{+}98, WSG^{+}98]. −/− [TYT^{+}99]. ^{138/303} [FIC^{+}97]. ^{2}+ [PT97, AMY^{+}98, ABM^{+}99, BHG^{+}97, BCC^{+}95, BML^{+}97, BRD99, BCWA97, CT97b, CBTZ98, EKF^{+}97, EAOS^{+}98, FCTPA^{+}99, GL95, HFP98a, IMK99, KJBM^{+}96, LCD^{+}98, LRG^{+}98, LYH^{+}99, LZC^{+}95, MAB98, MP98, MAS^{+}97, NYN^{+}95, RL96, RROA95, RWOA97, SLW^{+}96, SMH98, TT96, TMM97, TBD^{+}99, WWM^{+}98, YHMS^{+}98]. ^{3} [VB98]. ^{65} [GCKC^{+}97]. _{2}^{B} [XMMW97]. ^{BBP} [SPV^{+}99]. ^{c−−Y es} [MKB^{+}99]. c−neu [CWMD97]. ^{CAS} [YTY^{+}99, CHP^{+}98]. ^{cdc2} [NNSN97]. ^{cip1} [BZBA99]. ctn [ANRT99, YNG98]. ^{FAK} [CLRR99]. ^{Fyn} [HR99, vtHR97]. ^{KIP1} [SSR^{+}98]. lck [BINRM97, BM99, LWSL^{+}98]. ^{RB} [KLO^{+}98]. ^{v−src} [ZKN99]. ^{Val12} [CCMG97]. _{0} [LBWF^{+}97]. _{1} [MHC^{+}97, MNG^{+}98]. _{12} [KBS^{+}99]. _{13} [KBS^{+}99]. _{2} [AAB^{+}97, ALM97b, BPA^{+}97, FKNM97, GSKL99, GTC^{+}98, vtHR97, HPL97, sHWM^{+}99, KD97a, MBA97, MSH^{+}97, OWD^{+}99, SAR^{+}99, SWK^{+}99, SKC97, TNFK^{+}98, WRG98]. _{3} [PFAF97]. _{7} [CSF^{+}97]. _{B} [SZBB98]. _{CRAC} [HFP98a]. _{G} [WS98, ZB98, ZLM^{+}98]. _{i} [VLM^{+}98]. _{L} [NNK^{+}97]. _{o} [FAZ^{+}99]. α [ALVMM98, BMB97, BF97a, BGH^{+}98, CHLÅ97, CBS^{+}97, CAH^{+}99, CIK^{+}99, DCVF98, DHDJ^{+}97, DSM^{+}98b, DOP^{+}98, EKSC98, FNDS99, GMGGD98, GFN^{+}99, GBvdN^{+}98, GZC^{+}99, GGK97, HOS^{+}97, HTWC98, HZR98, HDDKH98, HHRdB^{+}98, IIM^{+}99, INMT97, IRW^{+}99, KHA^{+}99, KGL^{+}99, KRCP99, LNG^{+}98, LHB^{+}99, MGB^{+}97, MZL^{+}99, MZNR98, MMN^{+}99, MYC^{+}98, MPL^{+}97, MCS98, MLC99, MKW^{+}99, OSM98, PSPG^{+}98, PHN^{+}98, PLB97, PH97b, PWGG98, RM97, RTM99, RDS^{+}99, RYB^{+}99, RKB^{+}98, SAC^{+}98, SSZ^{+}97, SGR97, SRA^{+}99, SMS^{+}98, SDDV97, SB99c, TCYS99, TDR^{+}98, VSRP97, WLF99, WUUI^{+}98, WMR^{+}98, WKR^{+}98, YMAC^{+}98, ZDK^{+}99, vHGB^{+}97]. α6β4 [BRM^{+}99]. α_{2} [LPD^{+}99, XC97, XZSC98]. α_{5}β_{1} [BSW^{+}99, WMP^{+}98, HSML98]. α_{6}β_{4} [RdPRW98, VNTM^{+}99]. α_{E}β_{7} [HMS^{+}98]. α_{IIb} [GXE^{+}99, HPS98, PWD^{+}99]. α_{IIb}β_{3} [KSE^{+}97]. α_{V} [BP98]. α_{v}β_{3} [BSW^{+}99, JCR97, WMP^{+}98]. αvβ3 [TSC^{+}98]. β [BPD99, BPA^{+}97, BGH^{+}97, BRP^{+}97, BFL^{+}97, BKN^{+}97, BLRS97, BGH^{+}98, CBM^{+}98, CHLÅ97, CSN99, CAH^{+}99, CGC^{+}97a, DHDJ^{+}97, DDWM97, EKSC98, EPMB^{+}99, FCMN97, GHW99, GFN^{+}99, GWLH98, GvdFvD^{+}99, GZC^{+}99, HC99, HRD98, HDDKH98, IRW^{+}99, KRH98, KcZS97, KKLA97, LNG^{+}98, LTR^{+}97, LMW99, LBO^{+}98, MCS^{+}97a, MN98, MMAK97, MZNR98, MM97b, MBCS98, MGD97, NMK^{+}97, NG99, NGMR97, OSM98, OBW98, OOB99, PBJ^{+}98, PBA^{+}97, PH97b, PWGG98, RM97, RTM99, SZFM98, SLN^{+}97, SSR^{+}97, SPV^{+}99, SAC^{+}98, SBG^{+}98, SBPR^{+}98, SJF^{+}97, SKS99, SRA^{+}99, SSS^{+}98, SDL98, SDDV97, SB99c, TKY^{+}97, TGW99, TCYS99, TOSF^{+}98, VSRP97, WLF99, WMR^{+}98, WAV98, WRG98, WS98, YMAC^{+}98, ZSB99, ZLL^{+}98, ZOKS99]. β_{1} [CWSL97, LCB^{+}99]. β_{3} [BP98, GXE^{+}99, HPS98, KSE^{+}97, PWD^{+}99]. β_{4} [RdPRW98]. β_{II} [MDC^{+}99]. βγ [NA99, RP97]. δ [ALVMM98, LPM^{+}99]. ϵ [DSG^{+}97, SHB99, SKY^{+}98, ZLPL99]. γ [ALVMM98, ALC^{+}99, CRM^{+}97, DZS^{+}98, DKN^{+}97, FGA^{+}98, GSB97, HLJ^{+}98, JDK97, KR99, KOA^{+}99, MGIZ98, MZAO98, MUS98, NMH^{+}98, OWM^{+}99, PSLR99, SSI^{+}97, Sal99, SBGR98, SKY^{+}98, TCMB98, ZJB^{+}97]. κ [FWBSO97, HDV^{+}99, LDH^{+}98, SAC^{+}98, XZSC98]. λ [ÜHK^{+}99]. μ [BKMN^{+}98, BGBK99]. σ [ALM^{+}97a]. θ [TGW99]. ζ [HR99, KNT^{+}99, MN98, SLPE^{+}98, ÜHK^{+}99, XC97].

-/- [SGGS^{+}99, SZBB98]. -10 [SHK^{+}99]. -2 [FFH^{+}98, FSFT98, SVF^{+}98]. -3 [SHK^{+}99]. -4 [DVL^{+}98]. -6 [SHK^{+}99]. -7 [SHK^{+}99]. -8 [SHK^{+}99]. -Actin [KcZS97, ZSB99]. -Actinin [PLB97, RKB^{+}98]. -activated [SM95a]. -adaptin-coated [FGA^{+}98]. -Amyloid [CGC^{+}97a]. -anchored [LMC^{+}98].

-And [CIK^{+}99]. -ATPase [GHS98]. -based [RdPRW98, KTD^{+}97]. -binding [YTT99, SLW^{+}96, YCR^{+}98]. -Bungarotoxin [RDS^{+}99]. -C-kine [CBM^{+}98]. -Catenin [BPA^{+}97, vHGB^{+}97, LPM^{+}99, BLRS97, EPMB^{+}99, IIM^{+}99, LTR^{+}97, MM97b, NG99, OOB99, PBA^{+}97, SSR^{+}97, SGR97, SSS^{+}98, WKR^{+}98, WAV98, WRG98]. -Catenin-Vinculin [WUUI^{+}98]. -Cells [KKLA97]. -chain [DNB^{+}96]. -complexed [LRG^{+}98]. -containing [TBR^{+}96]. -converting [MMAK97]. -COP [DSG^{+}97]. -coupled [VLM^{+}98]. -deficient [BBJ98]. -dependent [GHP^{+}96, GRS^{+}99, LCS96, FCTPA^{+}99, TGW99, CT97b, CNP^{+}95, MAB98, NMK^{+}97, PT97, SKS99, SMH98, TMM97]. -Dystrobrevin [PSPG^{+}98]. -Dystroglycan [MLC99]. -endonexin [SOE^{+}95, KSE^{+}97]. -Factor [HZR98]. -Filagenin [LBO^{+}98]. -Glucan [KHA^{+}99]. -glycoprotease [CFS98]. -Golgi [CSA^{+}97, JMS97, KHHS97, MM97a, WBG^{+}97]. -Granules [HOS^{+}97]. -Helical [SDDV97]. ’-Iminodipropionitrile [ZLL^{+}98]. -independent [SKS99, VvZ99, WLG^{+}96, MLW^{+}95]. -induced [DOP^{+}98, ABM^{+}99]. -kinase [AJS96, BDE^{+}95, CMC^{+}98, Dav95, GPRS^{+}95, SMFC95, SDTE95, TCM96]. -kinases [BGN^{+}97]. -mediated [HSML98, SEC^{+}95, MSW^{+}98, GSvR^{+}96, SMH98]. -methylene-diphosphonate [HCAW95]. -methyltetrahydrofolate [SMA96]. -micron [PHB96]. -O-benzyl [HHRdB^{+}98]. -phosphate [BGR^{+}95, KDvFP96, dPF95, RWO95, RSJK95, SKR96]. -RANTES [RFVCM^{+}99]. -regulated [HIW^{+}98, BML^{+}97, RWOA97]. -Regulatory [LNH^{+}98]. -Sarcoglycan [DSM^{+}98b, HLJ^{+}98]. -Secretase [ALC^{+}99, CGC^{+}97a]. -signaling [vdBCH^{+}95]. -SNAP [BMB97]. -specific [ZB98]. -Spectrin [WS98]. -Subspecies [SSI^{+}97, SKY^{+}98]. -subunit [OBW98, BPD99]. -Synergin [PSLR99]. -syntrophin [SYO95]. -tagged [OTQM98]. -Terminal [sHWM^{+}99, vtHR97, BPA^{+}97, FKNM97, HPL97, KD97a, MBA97, MSH^{+}97, OWD^{+}99, SAR^{+}99, SWK^{+}99, SKC97, TNFK^{+}98, WRG98]. -terminal-processed [AAB^{+}97]. -Tubulin [MMN^{+}99, OWM^{+}99, DZS^{+}98, FNDS99, KR99, MZAO98, MUS98, NMH^{+}98, TCMB98].

/c [YTY^{+}99]. /CD9 [NIM95]. /Cyclin [NNSN97]. /eIF6 [SPV^{+}99]. /G [KBS^{+}99]. /H [FWG^{+}96, WSG^{+}98]. /PTH [SSA^{+}96]. /PTH-related [SSA^{+}96]. /Ribosome [YHS^{+}99]. /Ribosome-Associated [YHS^{+}99]. /RPTP [MN98]. /third [DLB96].

000-dalton [RFNF96]. 000-M [KPKWW95]. 029 [CSC^{+}98].

1 [AKC^{+}95, AFFS96, ABT^{+}98, ALVMM98, AKM^{+}98, BNW^{+}99, BSRG99, BHHF97, CWSL97, CC98, CMC^{+}98, CMS98, CDH^{+}99, DPZ^{+}97, DAP^{+}95, DKN^{+}97, DHT96, FCMN97, FFH^{+}98, FSFT98, GMGGD98, GvdFvD^{+}99, GGD^{+}97, HGW^{+}98, IMW^{+}99, ISN^{+}99, INMT97, IFM^{+}99, JMM^{+}98, JMWG98, JLID98, KGL^{+}99, KRS^{+}98b, KKG^{+}98, KJR^{+}98, KSG^{+}98, LPM^{+}98, LR96, LGM^{+}97, LVN^{+}98, MCS^{+}97a, MGM98, MSH^{+}98, MSS^{+}99, MFBK99, MWB98, MAB98, MSF^{+}99a, MMR99b, MBDS^{+}98, NMK^{+}97, OST^{+}95, OTL^{+}97, PHU^{+}95, PDS^{+}97b, PFP97, PH97b, RM97, REdC95, RKR^{+}97, SKA^{+}96, SVF^{+}98, SBPR^{+}98, SWS97b, SBGR98, SOPM^{+}97, SCH^{+}98, SWVR98, SGJP99, SJW97, TCYS99, TYSP^{+}96, TPM^{+}99, TTAH^{+}99, TRA^{+}99, VLM^{+}98, VZNR98, WLF99, WG99, YHK^{+}97, YMAC^{+}98, YKC95, ZZYG96, ZOKS99, WXS96]. 1- [PH97b, SYO95]. 1-5 [MPL^{+}97]. 1-Deficient [GvdFvD^{+}99]. 1-Dependent [WLF99]. 1-mediated [PH97b]. 1/CD31 [FSDA97]. 1/Fas [GSvR^{+}96]. 1/Semaphorin [MSF^{+}99a]. 10 [HSML96, SGGW99]. 10T1 [HRD98]. 10T1/2 [HRD98]. 11/ [MSF^{+}99b]. 110 [FSHD96]. 110-kD [FSHD96]. 115 [HIW^{+}98, MK95a, WMR^{+}96]. 115-kD [WMR^{+}96]. 12 [WG99]. 120 [CRCH95]. 13 [WGF^{+}98]. 13-kD [WGF^{+}98]. 138 [HS97]. 138-kD [HS97]. 14 [HSB98]. 14-3-3 [CRMB95, GXE^{+}99, LO96]. 16 [SCKG95]. 16-kD [SCKG95]. 17 [HSB98]. 170 [DPG^{+}99, DWM^{+}98, FNDS99, LM98]. 180 [BBAN^{+}99, WSSM95]. 180-kD [HBJ95, BBAN^{+}99, WSSM95]. 186 [FBM99]. 186-Kilodalton [FBM99]. 19 [GTC^{+}98]. 19-kD [GTC^{+}98]. 193 [KSK^{+}99a]. 193-Kd [KSK^{+}99a]. 19K [PSBB96, PW98]. 1A [BZB^{+}96, SZFM98, SWR^{+}96]. 1alpha [LTE^{+}96]. 1B [TKH^{+}97]. 1D [BRP^{+}97].

2 [ABT^{+}98, ANT^{+}97, BGB^{+}95a, BGGR^{+}96, BTNZ^{+}95, CG95, CWM^{+}98, CDH^{+}99, DKN^{+}97, DDM98, EY99, FCMN97, FEF^{+}97, GK99, GF97, GBBSO96, HMB95a, HLMD97, HRD98, HR99, HYK^{+}98, HHRdB^{+}98, IFM^{+}99, JM97a, JLID98, KGL^{+}99, KAS^{+}99, LLN^{+}95, LEB^{+}97, LBCB96, Mah96, MWGE97, MBCS98, OSK^{+}96, PKBHK97, SRR96a, Sal99, SGS95, SGD98, SPS97b, SPR^{+}98, SVH^{+}98, TKM^{+}96, WMC^{+}95a, YKC95, YHD^{+}98, ZRHO99, ZOKS99, vDSH^{+}97]. 2-3 [HHRdB^{+}98]. 2-mannosyltransferase [LSKB95]. 2/ [BLB^{+}98]. 2/Bcl [NNK^{+}97]. 20 [BBW^{+}98, CHC97, CSEM99]. 210 [IRMF^{+}99]. 24 [POR^{+}97]. 2’5 [DCJ97, GMC^{+}95, SLM^{+}95, WSBE^{+}95]. 250 [OGS^{+}96]. 250-kD [OGS^{+}96]. 26S [TYT^{+}97]. 270-kD [ZB98]. 280 [CLF^{+}97b, LTW^{+}97]. 29 [HHRdB^{+}98]. 2a [MKF96, RTP^{+}98, SNCBM95, TFF^{+}95, WPS^{+}96]. 2b [MKF96, SCB^{+}97]. 2c [CLF^{+}97b, HFSHK96]. 2H3 [PHB96].

3’ [SMS^{+}98, APG^{+}96, BMC^{+}99, BKK^{+}97, CBM^{+}98, CDH^{+}99, DBE98, DSvdV^{+}96, DFC^{+}96, ELL^{+}99, EPMB^{+}99, GGPG99, HGW^{+}98, HOS^{+}97, vtHR97, HHRdB^{+}98, IFM^{+}99, IHT^{+}98, JH97, KAS^{+}99, MNR^{+}98, MLN^{+}98, MS96, MSP97, ODB98, PWZS97, PMKR97, PBM^{+}97, SALdP^{+}97, SJHC96, SPCR97, SKC97, SHL97, SB99d, TFYH99, VMT^{+}99, WDV^{+}97, YMAC^{+}98]. 3-Dependent [BRM^{+}99]. 3-Kinase [HTK^{+}99, SvDtK^{+}98, GGPG99, PMKR97, PBM^{+}97, TFYH99]. 3-Kinase/Protein [HTK^{+}99]. 3’-Kinases [SMS^{+}98]. 3-Oh [EPMB^{+}99].

3-P-glycoprotein [DSvdV^{+}96]. 3/ICE [JWR96, MWGE97]. 33 [SBD^{+}99]. 34 [KKG95]. 34-kD [KKG95]. 35 [SMW^{+}99, XJM^{+}95]. 39 [HTS^{+}95]. 39-kD [HTS^{+}95]. 3A [MST^{+}96]. 3F [KAT^{+}95]. 3F3 [CG95, SVH^{+}98]. 3F3/2 [CG95, SVH^{+}98]. 3T3 [MAMJ95, NJB^{+}98, BL99]. 3T3-L1 [MAMJ95, BL99].

4 [AKC^{+}95, AC95b, HYE^{+}98, LHK^{+}96, MAMJ95, VC97]. 4.1-actin [SMB^{+}95]. 4.1R [MHH^{+}99]. 43/beta [SL95]. 43/beta-galactosidase [SL95]. 440 [Kun95, SZBB98]. 440-kD [Kun95, SZBB98]. 45 [BBOE98]. 480/270 [ZB98]. 480/270-kD [ZB98].

5 [RZZ96, GWB^{+}98, GSJ98, KCV^{+}98, KLS^{+}99, PFP97, RAH^{+}95, TAK95]. 5-Bisphosphate [KLS^{+}99]. 5-bisphosphate-sensitive [GHF^{+}95]. 5-Kinase [GWB^{+}98]. 5-trisphosphate [TT96, WS95a, YHMS^{+}98]. 5-Trisphosphate- [PFP97]. 5’-untranslated [RZZ96]. 5/ [GKS^{+}98]. 50 [EPVV96]. 50-kD [EPVV96]. 53 [ISH95, VKIH98]. 550 [WKL^{+}99]. 550-Kd [WKL^{+}99]. 5A [TYSP^{+}96]. 5S [DKLS^{+}99].

6 [AMY^{+}98, BGK^{+}95, BSM^{+}99, CC98, KKG^{+}98, LH97, MNM^{+}98, PFHWN98, TYK^{+}98, YHK^{+}97]. 6-bisphosphatase [CC98, CC98]. 6-glucan [LMB^{+}95]. 6-Phosphate [NCV^{+}98, LH97]. 6-Phosphate/Insulin-like [NCV^{+}98]. 60B [OBB99]. 60S [SPV^{+}99]. 61hck [LFB96]. 65 [DBOM98]. 65-kD [DBOM98]. 67C [MMN^{+}99]. 6A [GBvdN^{+}98].

7 [ARR^{+}98, AYP^{+}99, DBJDT99]. 70 [ANS^{+}97, HIWL97, SLPE^{+}98, SWVR98]. 70-kD [ANS^{+}97]. 70S [VSGF96]. 76 [KNT^{+}99]. 76p [FRMLT^{+}99]. 78 [IMS^{+}98, KPKWW95]. 79 [CFS98]. 7B2 [ZL95c, MZL97].

8 [MSL^{+}98a]. 8-kD [BRLM98]. 8-Positive [MSL^{+}98a]. 80 [XGC96]. 80-kD [XGC96]. 80s [VSGF96].

9 [LGZ^{+}99, MCTMK98]. 95 [GCML98, TBP^{+}99]. 95-kD [GCML98, TBP^{+}99]. 95/SAP90 [KNH^{+}97]. 95F [MM95a]. 97 [KIT^{+}97]. 97-kD [KIT^{+}97].

a-factor [BGN^{+}96, RD96b, FKNM97]. A-inducible [SRRB95]. A-kinase [YDFB98]. A-treated [LCK95]. A1 [IJB^{+}97, SD95]. A2 [KBC99]. A20 [HDV^{+}99]. A431 [BKJL^{+}98, FLW^{+}95, GJ95a, MSH^{+}98]. abalone [SFSV95]. ABC [EM96]. Aberrant [STN^{+}96]. Aberrantly [WM96]. Abilities [ZGYS99]. Ability [FIC^{+}97, NDBR95, OWM^{+}99]. ABIN [HDV^{+}99]. Ablation [RPS^{+}97, WGP98]. abLIM [RHA^{+}97]. Abnormal [KPZ^{+}97, NKK^{+}99, PJE^{+}99, CLT^{+}95, SdCAH^{+}97, DBH^{+}97, KRS^{+}98b, WHB^{+}99]. abnormalities [ADP^{+}95, AYF96, CGSY^{+}99, DPT^{+}97, KZS^{+}98, RLMC99]. Abolish [QPH^{+}98]. Abolishes [GRTB97]. ABP [CRCH95, LTW^{+}97]. ABP-120 [CRCH95]. ABP-280 [LTW^{+}97]. abrogation [PHB96]. abrupt [ZKSB96]. Absence [AC95a, EFB^{+}98, SVM^{+}99, DSGF95, DPH95, KTC95, LYC^{+}95, MKB^{+}95, SSY^{+}96, SSL^{+}95, CML^{+}98, KDBW97, KGHG98, SG99a, ULMT99]. absent [RACHV96]. abundant [CSDR96, KSY^{+}98]. Ac [CGJ^{+}95]. Acanthamoeba [BCMK95, JOPM98, KAP95, MSP97, OP96b]. Accelerated [GBW98]. Accelerating [The97]. acceleration [ZPS96]. acceptor [KKG95]. access [MGS96]. accessory [WMR^{+}96]. accompanies [VIF^{+}96, JLM99]. accompany [NMHS95]. accompanying [HCAW95]. Accomplished [CH97]. accumulate [NMJL96, RvSL^{+}95]. Accumulates [SDL98]. accumulating [TSS^{+}95b]. Accumulation [BAC97, ISFB^{+}96, HCA^{+}95, MM95b, SS95a, WMC^{+}95b, EPMB^{+}99, LBWF^{+}97, MM97a, MSL^{+}98a, PFHWN98, RPS^{+}97, TFYH99]. accurate [WGG96, PMBM^{+}99]. acetyl [STN^{+}96]. acetylated [KSL^{+}95]. Acetylation [GCB^{+}95]. Acetylcholine [GDR95, ACHG95, FDH96, MPW95, Wal95, WHH96, BGH^{+}98, DP98, LE97, MMM^{+}98, OMS97, SGYH97]. Acetylcholinesterase [PXRR99, FKM^{+}99, RRA97]. acetylgalactosaminylphosphotransferase [BEZ^{+}95]. Acetyltransferase [RPY^{+}99]. Achiasmate [MMN^{+}99]. achieved [SSN96]. AChR [TCM96]. AChRs [KBSW95]. Acid [CCMC97, CHH97, BAC^{+}96, EB95, HWPC96, KSL96, MLLT95, MMMT95, NNMW96, OSK^{+}96, OABD^{+}96, STN^{+}96, WPS^{+}96, XGR95, YFST^{+}96, CBS^{+}97, CNG98, DSG98, ESR97, GF97, TSD^{+}97, YHD^{+}98]. Acid-induced [CHH97, EB95, OSK^{+}96]. acid-neural [WPS^{+}96]. acid-responsive [OABD^{+}96]. Acidic [DBE98, PJE^{+}99, CMHT96, GCGE^{+}96, LDCB95, FSH^{+}99, HTY^{+}99, SVB^{+}97, TSB99]. acidification [HP95]. Acids [LHS97a, SKY^{+}98, SBS^{+}98]. acinar [RRH^{+}95, IMK99]. Acquire [LLG^{+}99]. Acquisition [JDK97]. Acrosomal [GRB^{+}97, AZF96, WSG^{+}95]. acrosome [SOH^{+}95, hYGYF97]. acrosomes [WS95a]. across [DFD^{+}97, NMHS95, PCL^{+}96, VJHR96, WHM^{+}96]. Acrp30 [BL99]. act [KA95, LO96, AvdG99, BS97, KLS^{+}99]. ActA [LZK^{+}97, STP96]. ACTH [BBW^{+}98]. Actin [BMC^{+}99, BZL^{+}98, CLS^{+}97, Cun95, ELL^{+}99, EGKC^{+}99, GCST97, HCW96, HCM96a, KTC95, KcZS97, LLH^{+}99, gLSE^{+}99, MEFH97, MM99a, MMP^{+}95, MKXY95, MSP97, OBB99, PPWL97, RAA^{+}97, SSP95, SHA95, SB99d, The97, VMH97, WDV^{+}97, ZC99, AOMB96, ANS96, ASY96, BWKH96, BMLM96, COML95, CGB^{+}95, CGSP95, CGJ^{+}95, CLL95, CRM^{+}96, CRCH95, DTB96, DWS^{+}95, EAL^{+}95, FWHW96, FUB^{+}96, GW95, GAW^{+}96, GPRS^{+}95, KAP95, LTVC96, LZD95, LTE^{+}96, MMMW96, MH95b, MELC96, PAV^{+}96, PMCT96, RFNF96, RJP96b, RAE96, SHC95, SJC96, SMB^{+}95, SH95, STP96, SRRB95, TET^{+}96, TTG95, TCSG96, TTG96, ÜHK^{+}99, VSB95, WKWC96, WSG^{+}95, YWW^{+}96, ZPS96, vABSP95, ABE^{+}98, ASP^{+}97, BALL98, BD98, BVKP98, BGN^{+}97, CYSD99, CCC^{+}99, CSM97, CLF^{+}97b, DS98b, DSM^{+}97, EGM^{+}99, EHC97, FCTPA^{+}99, GFN^{+}99, GDL98, GWB^{+}99, HMD97, HC99]. Actin [HKS97, HB99, HYE^{+}98, HHR^{+}98, INMT97, JL98, JOPM98, KMMC98, KHA^{+}99, KMHK97, KSK^{+}99b, KWS^{+}99, LZK^{+}97, LL97, LSZ^{+}99a, Li97, LL98b, MCJK98, MH97, MI99, MNS^{+}97, MLB98, MPCW97, MSL98b, MMR99a, MHM^{+}97, NOS^{+}97, NHA^{+}97, NLM^{+}98, ONI^{+}98, OC98, OFM^{+}98, PBJ^{+}98, PTJ^{+}98, PSB98, RM97, RTM99, RKB^{+}98, RC97, RTL^{+}99, RHA^{+}97, SWM^{+}98, SWS97a, SMS^{+}98, SGA97, SMTN99, SVMB97, TIBL99, TCV^{+}98, TWP98, TLBK98, VMH97, YAD97, ZSB99, ZJB^{+}97, ZJY^{+}98]. Actin-associated [MHM^{+}97]. Actin-Based [ELL^{+}99, CLS^{+}97, LLH^{+}99, MMP^{+}95, STP96, ZPS96, LZK^{+}97]. Actin-Binding [EGKC^{+}99, MSP97, VMH97, CRM^{+}96, PMCT96, RFNF96, CLF^{+}97b, OFM^{+}98, RKB^{+}98, RHA^{+}97]. Actin-bundling [BZL^{+}98, TET^{+}96, HYE^{+}98]. actin-capping [CLL95]. Actin-containing [RM97]. actin-cross-linking [WSG^{+}95]. Actin-dependent [SSP95, EAL^{+}95]. actin-depolymerizing [AOMB96]. Actin-encoding [KMHK97]. actin-related [KAP95, MMMW96]. acting [MC95, REdC95, CH97, KBC99, SW98]. Actinin [XWK^{+}97, KSJW95, PWG^{+}95, HYE^{+}98, PLB97, RKB^{+}98]. Actinin-4 [HYE^{+}98]. Actinin-associated [XWK^{+}97]. action [DPHW96, GEJ^{+}95, OWA95, AMY^{+}98, FEF^{+}97, MBDS^{+}98, SHMC99, ZLM^{+}98]. Actions [FCMN97, MBF^{+}97, VLM^{+}98]. activate [WM96, BKK^{+}97]. Activated [PMKR97, SBC99, YTT99, GPRS^{+}95, HTK^{+}99, HG95a, LS95a, LYR96, SK96, STT^{+}96, VTG^{+}96, Wal95, WMXE96, DSM^{+}97, GCM^{+}97, KDO^{+}99, PWD^{+}99, PHN^{+}98, SMTN99, TCYS99, TLF^{+}97, UNPW98, CMS98, DCJ97, EKSC98, GTY98, KCG^{+}97, SM95a, WZF97, ZTH^{+}97, ZOS^{+}97]. activates [HWFA96, RK95, BxQK^{+}99, FWBSO97, FFLP99, SSH^{+}98, XZSC98]. activating [CKS^{+}95, ZBG99, JK97, MCB96, MWB98]. Activation [AGS^{+}97, AKM^{+}98, BHD^{+}97, BP98, ELL^{+}99, GXE^{+}99, HPD^{+}97, HSML98, KD97a, KBS^{+}99, KAS^{+}99, LPM^{+}98, LKSW98, MCS^{+}97a, MSH^{+}97, NNSN97, PSBB96, SVH^{+}98, SHK^{+}99, SP96, TYY^{+}99, TMN97, TSC^{+}98, AZF96, BFP^{+}96, BMLM96, CRG^{+}95, FLW^{+}95, GJ95a, GLS96, KRMMA96, LLN^{+}95, LYR96, MBF^{+}95a, MTGY96, RD96a, SKA^{+}96, SKS^{+}95, SJA96, TCM96, TMM^{+}95, ZOB^{+}96, BAC97, BATP99, BOM^{+}98, CAS^{+}98, CWMD97, EKF^{+}97, EPMB^{+}99, GvdFvD^{+}99, IRW^{+}99, JP98, JWFS99, LPM^{+}98, LBBP98, LWSL^{+}98, LAS^{+}97, MSH^{+}98, MGP97, MZL97, NGMR97, NMNL98, OWD^{+}99, OF98, OWW^{+}99, PH97b, RPS99, SBGR98, TYT^{+}97, TSB99, VGVF97, VLAHV98, vHGB^{+}97]. Activator [NCW^{+}99, SPK^{+}97, DCW^{+}96, SM95c, CMD^{+}98, EvdESvdH^{+}97, GSJ98, JLID98, OTL^{+}97]. Activator-Stimulated [NCW^{+}99]. Activator/Plasmin [SPK^{+}97]. Active [DSVL^{+}98, FTH96, HTW97, KLN^{+}96, RMY97, SVH^{+}98, ZCE^{+}98, EDB^{+}96, GDR95, JMTCF96, NAP^{+}96, OMW^{+}96, RACHV96, SFSV95, VSB95, VBZ96, DPM^{+}97, GCKC^{+}97, HSB98, JVRNM99, WKL^{+}99]. Actively [CL99, RV98]. activin [YtDH^{+}95]. activin-like [YtDH^{+}95]. Activities [TYSP^{+}96, YF99, DRR96, GSB^{+}96, KSG^{+}96, TRJ^{+}95, WXS96, LSH^{+}99, MKS^{+}97, MM97b, OWAlW99, WAV98]. Activity [MSP97, MBBT98, PBB^{+}96, SCR98, SKC97, WLF98, XC97, AHM^{+}96, BYG96, BM96, FFGG96, FaAR^{+}96, FWG^{+}96, GDR95, GMFN95, GBBSO96, HIG^{+}95, KWK^{+}96, KJBM^{+}96, KvEdV^{+}96, KSVZ95, MAS^{+}95, OCS^{+}95, PMH96, RL96, RR96, ROG^{+}96, SWB^{+}96, SLL96, SCG^{+}96, STT^{+}96, SMFC95, SSP95, SIAS96, TLF^{+}95, TMM^{+}95, ZL95c, BMB97, BBAN^{+}99, BZBA99, BJR^{+}99, CR99, CS97b, EKSC98, FSA^{+}99, GWB^{+}98, GLMG98, HCLG99, HLK^{+}98, KTYT99, LDH^{+}98, LJK97, iMOeM^{+}97, MBDBB98, MLB98, MBCS98, MGD97, MSH^{+}99, NRP98, OK98, PDS^{+}97b, QPH^{+}98, RME98, StKvD^{+}99, SHMC99, SEH^{+}97a, SGGW99, qSmLY97, VWG^{+}97, ZH97, ZS98]. Activity-dependent [MBBT98]. Activity-induced [PBB^{+}96, TLF^{+}95]. Actomyosin [WSS97, BML^{+}98, LL98a]. Actomyosin-based [WSS97]. acts [KM95, KCBR96, OMW^{+}96, ANRT99, CWSG99, CS98, SHLD97]. actuating [CMUW96]. Acute [PHA^{+}98]. Acyl [SBS^{+}99]. Acylated [BM99, SLS^{+}96]. Acylation [HR99, LHS97a]. Aczonin [WKL^{+}99]. ADAM [CAH^{+}99, WPMW95, YPM97]. Adaptation [ACHG95]. adaptin [FGA^{+}98]. adaptor [BGB^{+}95a, DHT96, PR95, SBW^{+}96, CRYC99, DKN^{+}97, KKG^{+}98, ODB98, SPCR97, SHL97, WE98, WBR97, vDSH^{+}97]. adaptor-related [SBW^{+}96, SPCR97]. adaptors [WB96]. ADAR1 [EJ99]. added [LCP95]. Addition [GDC97]. Adducin [MLB98, FOK^{+}99]. Adenine [BSRG99]. Adenomatous [BPA^{+}97, VWG^{+}97, NAP^{+}96]. Adenosine [DVL^{+}98, MBM95, NH95]. Adenovirus [SBR^{+}96, BRJP96, CSO^{+}95, PSBB96, PPW^{+}95, LNM^{+}98, SNK^{+}99]. Adenovirus-mediated [SBR^{+}96]. Adenylate [WLF99]. adenylyl [LD95, SKS^{+}95]. Adequate [VUC^{+}95]. ADF [CLS^{+}97, OBB99, The97, GBW^{+}95]. ADF/Cofilin [CLS^{+}97, OBB99, The97]. Adherence [FIC^{+}97, MBD^{+}96]. Adherens [MNS^{+}99, TNM^{+}99, CKP96, DZC^{+}96, FSIG^{+}95, FSR^{+}96, MW96, HMC^{+}99, LWS^{+}97b, MNS^{+}97, ONI^{+}98]. adherent [AFFS96]. adhering [HHM^{+}95]. Adhesion [ACSN98, BPA^{+}97, CBM^{+}98, CHP^{+}98, CT97b, FSDA97, HSML98, KMI^{+}97, KBB^{+}97b, LCN^{+}97, MEFH97, RdPRW98, SvDtK^{+}98, SMH98, TNM^{+}99, ZOB^{+}96, ANS96, AC95a, ABN96, BEZ^{+}95, BLE^{+}96, BNA^{+}96, BOLS95, BGK^{+}95, BPT96, BKB^{+}96, BHP^{+}95, CQB^{+}96, DLB96, DZC^{+}96, DNC^{+}95, DCW^{+}96, DYF96, DSC^{+}95, DMD^{+}96, DFC^{+}96, FSIG^{+}95, FPM^{+}95, FBH^{+}95, FWHW96, FVRCH96, GPRS^{+}95, HG95b, HH95b, ICW^{+}96, KSJW95, KZKS96, LCS96, LXK^{+}95, LFB96, MAFJ^{+}95, MBS^{+}96b, OP96a, PLR^{+}95, PHU^{+}95, RWM^{+}96, RD96a, RAH^{+}95, SOHP95, SEC^{+}95, SPB^{+}95, TNY^{+}95, UHJ^{+}96, VDA^{+}96, WPS^{+}96, WvdVV^{+}95, WF96, YWW^{+}96, ZZvdB^{+}96, dPSMNSM95, ANS^{+}97, ALCC^{+}99, ABL98, BO99, BR98, CCS^{+}98, CCB^{+}98, DCR98, ETP^{+}99, FHSM^{+}97, GRTB97, GCM^{+}97, GGD^{+}97, GB98, GSP^{+}98, HSL^{+}99, HHM^{+}98, HSSW99, IAS^{+}98, IIM^{+}99, INMT97, JLID98, KEJ97, KSG^{+}98, LKKL97, LBW^{+}97, LVN^{+}98].

Adhesion [MH97, MPLS^{+}98, MZL^{+}99, MSB^{+}97, MGD97, NWM^{+}97, NB97, NLM^{+}98, NH99, OWD^{+}99, OFM^{+}98, OK98, PHM^{+}98, RPS99, SLW^{+}99, SZBB98, SGR97, SALdP^{+}97, SWS97b, SSEM98, SLS^{+}99, SEBF98, TSK^{+}97, TCYS99, WYL^{+}99, WAV98, WFS97, WSWR99, cXP97, YPM97, ZRG98, vHGB^{+}97, GSvR^{+}96]. Adhesion-dependent [ZOB^{+}96, FVRCH96, LFB96, PLR^{+}95]. Adhesion-triggering [CBM^{+}98, CQB^{+}96]. adhesions [AFFS96, CWB96, CH95, FAP^{+}96, KCDB95, CLRR99, KRS98a, NJB^{+}98]. Adhesive [CT98, DAP^{+}95, GSC^{+}96, PCL^{+}96, RS95, BRP^{+}97, HPS98, LPM^{+}99, ODJ99, YNG98, iMOeM^{+}97]. adhesiveness [OSPJ^{+}96]. Adipocyte [CJH^{+}98]. adipocytes [MAMJ95, BL99, CMC^{+}98]. adipogenic [PNL^{+}95]. adipose [KSP95]. adjacent [TKVC95, STF98]. Adp [ALK^{+}99, KSK^{+}99a, SAM^{+}99, DHT96, FHK97, JM97b, KBW^{+}96, MCS^{+}97b, ODB98, RD97, WBR97]. ADP-dependent [JM97b]. Adp-Ribosyl [SAM^{+}99]. Adp-Ribosylation [ALK^{+}99, DHT96, WBR97, ODB98, RD97]. Adrenal [SN97]. Adult [MAM99, AC95b, FHV^{+}96, ASES^{+}97, CAS97, PMCS97, RV98, SGGS^{+}99, WS98, YDFB98]. Ae1 [AKCC99]. Aequorin [ABM^{+}99]. Aerolysin [AvdG99]. AF [BSM^{+}99, TYK^{+}98, YHK^{+}97]. AF-6 [BSM^{+}99, TYK^{+}98, YHK^{+}97]. Afadin [INM^{+}99, MNS^{+}97, MNS^{+}99, TNM^{+}99]. affect [CPL^{+}95, WGG96]. affected [HCM96a, SSSKR95]. affects [FK95, FPM^{+}95, MC95, MMMT95, SLL96, WLML95, ASACF98, HBS^{+}97, HMC^{+}99, LMB98, PURB98, vLKvdK^{+}97]. Affinity [KSE^{+}97, CLL95, HMB95a, HGH96, LHDC95, WS96]. after [BMC^{+}99, BKJL^{+}98, FJT^{+}95, GMS^{+}98, HE96, KD97a, KZKS96, LWM^{+}95, MRM^{+}98, MW97, MCPB99, OS96a, PTBC96, SVM^{+}99, SDDV97, WTCV95, cXP97, vGvM95, JHM98]. against [KKD^{+}95, LW96, VWK^{+}95, KTAX98, MZNR98]. Age [EFML99, BOLS95]. Age-Related [EFML99, BOLS95]. Agent [ZRHO99]. Agent-induced [ZRHO99]. agents [LLN^{+}95, MH95a, PMP^{+}98]. agglutinin [LMB^{+}95]. Aggregates [MSL^{+}98a, MPSC98, SHC^{+}97]. aggregating [GDR95]. Aggregation [JLM99, WBKB97, MTGY96, MMM^{+}98, PPP^{+}98]. Aggresome [GMBSS99]. Aggresomes [JWK98]. aging [HAS^{+}96]. Agonist [CBM^{+}98, LCD^{+}98, VB98, SKWY95, LYH^{+}99]. Agonist-induced [VB98, SKWY95]. Agonist-sensitive [LCD^{+}98]. Agonists [CMS98]. Agrin [DBG^{+}97, FDH96, MMM^{+}98, DGSR95, GDR95, MPW95, Wal95, CJY^{+}97]. agrin-activated [Wal95]. Agrin-induced [FDH96, MPW95, CJY^{+}97]. AGT [MLO^{+}95]. Aip1p [RTL^{+}99]. AIR [SGD98]. AIR-2 [SGD98]. Airway [LGZ^{+}99, MKB^{+}99]. AKAP1 [sHWM^{+}99]. AKAP100 [YDFB98]. Akap95 [CLT99]. Akt [BRM^{+}99, HTK^{+}99, PMKR97, VMT^{+}99]. Akt/PKB [BRM^{+}99]. alanine [LOD96, MLO^{+}95]. alanine/glyoxylate [MLO^{+}95]. Alcohol [RWM^{+}96, WRdVC97]. Aldrich [Li97]. Alf1p [FNDS99]. alga [SG95]. alignment [EPVV96, ETF95, DWM^{+}98, SCM^{+}97]. Alkaline [PBS97, FSH^{+}99, SHL97]. allele [GDC95]. allow [WSGPS95].

Allows [VG99]. along [BG98, HW96b, KDvFP96, KCBR96, MH95b, MKI^{+}96, SWO^{+}99, YvdEH^{+}95, ZCOP99, BSB^{+}97, WBS^{+}98]. alpha [AM95b, BLB^{+}96, BLLB95, BGGR^{+}96, CGJ^{+}95, CJG^{+}95, CNP^{+}95, DNB^{+}96, DMP^{+}96, DLIB^{+}95, ES96, FWHW96, GCB^{+}95, GHJJ95, GLD^{+}96, GBG^{+}96, GGC96, GHF^{+}95, GPRS^{+}95, HSL^{+}95, HBJ95, HHM^{+}95, HCAW95, ISFB^{+}96, KKD^{+}95, KSJW95, LCC^{+}95, LRM^{+}95, LCS96, LMB^{+}95, LSKB95, MBD^{+}96, MTLW95, MS96, MBS^{+}96b, MELC96, NIM95, PWG^{+}95, PHU^{+}95, PPF96, RHF^{+}96, RJP^{+}96a, RJP96b, REdC95, SWR^{+}96, SLT^{+}96, SHS96, SCK^{+}95, SEC^{+}95, SYO95, TKM^{+}96, WHH96, WAMS96, XZSC96, YRM^{+}96, ZMM^{+}95, CRMB95, KRMMA96, LDCB95, SSSKR95]. alpha-2 [CNP^{+}95]. alpha-actin [FWHW96]. alpha-actinin [KSJW95, PWG^{+}95]. alpha-agglutinin [LMB^{+}95]. alpha-catenin [KSJW95, LCC^{+}95]. alpha-dystroglycan [CJG^{+}95, DLIB^{+}95]. alpha-satellite [SHS96]. alpha-smooth [CGJ^{+}95, RJP96b]. alpha-Tubulin [GGC96, GCB^{+}95]. alpha/beta [ES96]. alpha1 [BFP^{+}96]. alpha2 [VKR^{+}96]. alpha6beta4 [MPY^{+}96, XGC96]. alphaV [WKL^{+}96]. alphaV-related [WKL^{+}96]. alphavirus [LLN^{+}95, LDH^{+}98]. alphavirus-induced [LLN^{+}95, LDH^{+}98]. Alport [MS96]. ALSV [HPD^{+}97]. alter [ROG^{+}96, GHW99]. alteration [DNB^{+}96]. Alterations [KSM96, PLZ^{+}95, DSvdV^{+}96, NMHS95, LGM^{+}97]. Altered [BPA^{+}97, FWHW96, NTAN98, STS96, VCJA99, FFZ^{+}95, TFF^{+}95, KD97b]. Alternate [SHL97]. Alternative [TFS^{+}99, SYO95, ZL95a, BRP^{+}97, CMS^{+}97, SGYH97]. alternatively [AK95, JBD^{+}95, LBCB96, PJL^{+}95, FPF^{+}99, iMOeM^{+}97, ZBW^{+}97]. alters [BCSG96, CRCH95, KGW^{+}95, MHLB96, RWO95, SHC95, XC96, CSEM99, MQR^{+}97, PKQ^{+}98, PH97b, SMG^{+}97]. Aluminum [RKD96]. alveolar [WPLK95, NBS^{+}98]. Alzheimer [EGS^{+}98, JM97a, SOJM99, VRD96, YF99]. ameloblast [MLLT95]. Amino [KSL96, BLL95, DGSR95, HAS95, NNMW96, GF97, KBB^{+}97a, LHS97a, YHD^{+}98, ZOKS99]. amino- [HAS95]. amino-terminal [BLL95, DGSR95, KBB^{+}97a]. aminoacyl [LTE^{+}96]. aminoacyl-tRNA [LTE^{+}96]. Aminopeptidase [SBOK97, OSHG^{+}96]. Aminophospholipid [CIRG99]. aminopterin [HSWC96]. aminotransferase [LOD96, MLO^{+}95]. amnion [CLR^{+}96]. Amoeboid [ISR99]. among [BLL^{+}96]. Amphiphysin [BDO^{+}97]. Amphiphysin/Rvs [BDO^{+}97]. Amplified [SIUW98]. amplifiers [Val96]. AMPPNP [SRR^{+}97]. Amylase [CHC97, CFB^{+}98]. Amyloid [ALC^{+}99, CGC^{+}97a, AHM^{+}95, AM95b, HKC^{+}95, YSK95, MZNR98, SDL98]. Amyloid- [TOSF^{+}98, MZNR98]. analogs [HP95]. analogue [HSWC96, HCAW95, RW95]. analogues [vGvM95, MSM99]. analyses [BTNZ^{+}95, KGGK99]. Analysis [CT97a, EGM^{+}99, FKS95, GXE^{+}99, KS97, MJMS96, MM97b, OPH^{+}99, OMO95, PvDA^{+}98, SBS^{+}99, SVMB97, TYT^{+}99, WJH^{+}98, ZBG99, ANS96, AHM^{+}96, ABN96, BFL95, BHP^{+}95, CGB^{+}95, DKD^{+}96, FTH96, HAS^{+}96, HG95b, MGA^{+}96, PSW95, VMN95, WMO^{+}95, YN95, YRM^{+}96, BBP^{+}98, BMS^{+}97, BWWT97, DFMG97, FLA99, FKM^{+}99, FY97, GSK^{+}99, HME^{+}98, JMWG98, JM97b, LMF98, OAB^{+}98, PGP^{+}99, PAN97, RDP^{+}99, SWM^{+}98, SGA97, TWiK^{+}99, WLK97]. Anaphase [DMMS98, KWD^{+}98, LARH97, CG95, KCBR96, PBT^{+}95, RCKS95, SKE^{+}95, STT^{+}96, TMM^{+}95, WFBC96, YGK96a, YGK96b, ZKB95, GCM98, MHWW98, MHC^{+}97, SSM98, WHG^{+}97, EAM^{+}97, YYSB97]. Anaphase-promoting [KWD^{+}98]. Anastral [MMGT96]. anatomy [SM95b]. ance [KB99a]. anchor [HER95, GSB97, WS97]. Anchorage [Ass97, KGF^{+}97, XGC96, BCSG96, OOB99, Sch97]. Anchorage-dependent [Ass97]. Anchorage-regulated [KGF^{+}97]. anchored [GH95, HE96, HIG^{+}95, KE98, MWC95, MVvdBD96, NIM95, vdBCH^{+}95, BRS99, Sal99, AFG^{+}98, BGK^{+}95, DTB96, KSY^{+}98, LMC^{+}98, MBOC97, WYY95]. Anchoring [vtHR97, CLT99, LBBP98, QGE^{+}99, TFS^{+}99, WSA^{+}99, YDFB98, sHWM^{+}99, WWO^{+}98]. Anchors [KSF^{+}99, BLC95, CGB^{+}95, MTLW95]. Androgen [RPS^{+}97]. angiogenesis [MA95, BFL^{+}97, EKSC98, SPR^{+}98]. Angiogenic [BAS^{+}97]. Angiotensin [LYR96, LYLR98]. Anillin [FA95]. Animal [SRCC97, GSL98]. animals [FWHW96, LKD^{+}95]. Anion [AKCC99, SSS99a, vHMV^{+}97]. Anion-selective [SSS99a]. anisotropy [DFGL96]. Ank3 [PJL^{+}95]. AnkG119 [DSM^{+}96]. Ankyrin [HPL97, TBD^{+}99, DLB96, DSM^{+}96, Kun95, OFY^{+}95, PJL^{+}95, BAC97, GRTB97, SZBB98, TBP^{+}99, WS98, ZB98, ZBW^{+}97, ZLM^{+}98]. Ankyrin-3 [HPL97]. Ankyrin-B [TBD^{+}99]. ankyrin-binding [DLB96]. ankyrin-related [OFY^{+}95]. ankyrinB [Kun95]. Annexin [CGVS^{+}96, KsYA98, LLVS98, CSDR96, FLPS95, FLPS95]. annulate [MMF95, LBWF^{+}97]. anoikis [FVKS96, OOB99]. Antagonism [SGR97]. Antagonistic [KTAX98]. Antagonistically [CH97, SJ99]. antagonists [LW96, SCG^{+}96]. antagonized [TYSP^{+}96]. antagonizes [FFGG96]. Anterior [SCEB99]. Anterograde [SNCH95, NH95, NSYK^{+}95, GE97]. anti [CG95, GSC^{+}96, ASACF98]. anti-adhesive [GSC^{+}96]. Anti-coilin [ASACF98]. anti-phosphoepitope [CG95]. antibioticus [MHM99]. Antibodies [KKD^{+}95, LNY^{+}95, LW96, NKI^{+}95, VWK^{+}95, ZKH^{+}95, ASACF98, SSS^{+}99b, WPW^{+}97]. Antibody [LN96, CG95, ZZvdB^{+}96, GCM98, HTY^{+}99, ZBG99]. Antigen [AWC99, KMI^{+}97, SBG^{+}98, BGK^{+}95, CLT^{+}95, HBJ95, LGdC95, RS95, SMFC95, WLR96, ASB^{+}99, BKN^{+}97, CWT^{+}98, CCB^{+}98, CFS98, GAZ^{+}98, GSP^{+}98, JLM99, PH97b, SPS97b]. Antigen-1 [PH97b]. antigen/BP180 [HBJ95]. Antigenic [FKWP97, GAZ^{+}98]. antiglycolipid [ZZvdB^{+}96]. Antioxidant [FEF^{+}97]. Antiparallel [MCBE98]. antisense [OST^{+}95]. Antiviral [DCJ97]. anucleate [FT95]. AOP [RFVCM^{+}99]. Aorta [DBC^{+}99]. Aortic [GCM^{+}97]. AP [BGB^{+}95a, BLB^{+}98, DBE98, DHT96, KKG^{+}98, MSH^{+}98, ODB98, SKA^{+}96, SPCR97, SHL97, TBR^{+}96, GK99, TRA^{+}99]. AP- [TBR^{+}96]. AP-1 [DHT96, KKG^{+}98, MSH^{+}98, SKA^{+}96, TRA^{+}99, WXS96]. AP-2 [BGB^{+}95a, BLB^{+}98, GK99]. AP-2/ [BLB^{+}98]. AP-3 [ODB98, SPCR97, SHL97]. AP-3-dependent [DBE98]. APC [BHHF97, CJS^{+}99, KTYT99, PBA^{+}97, YGK96a]. APC-interacting [BHHF97]. Apc2 [MDK^{+}99]. apCAM [SEBF98]. Apg8 [KBI^{+}99]. Apg8/Aut7p [KBI^{+}99]. aphidicolin [DKD^{+}96]. aphidicolin-treated [DKD^{+}96]. Apical [DMY^{+}97, HHRdB^{+}98, IMS^{+}98, MGB^{+}98, NDBR95, OPD^{+}96, YHMS^{+}98, YLB^{+}97, ZT99, vIZKH97, vIH98, BWF^{+}96, DvD95, FLPS95, KE98, MVvdBD96, ALK^{+}99, BRS99, BFRB99, CS98, GFM^{+}98, LLVS98, LMC^{+}98, LCW^{+}98, MKB^{+}99, PMBM^{+}99, SRV^{+}97, Sal99, TFYH99, VTHAA99, WUUI^{+}98, WC97a, ZSP^{+}98, ZH97, vHMV^{+}97]. apical-basolateral [BWF^{+}96]. Apicobasal [ZT99]. Apicomplexan [KBG^{+}99]. APO [GSvR^{+}96]. APO-1 [GSvR^{+}96, AKM^{+}98]. APO-1/Fas [GSvR^{+}96]. Apolipoprotein [DSM^{+}98a]. Apoptosis [AKM^{+}98, BBL^{+}99a, CCMC97, CCMG97, KD97a, KTD^{+}97, MFL^{+}99, MMAK97, PW98, SKC97, AHBW96, DH96, GBBSO96, GSvR^{+}96, HCM^{+}96b, LKD^{+}95, LK96, MBF^{+}95a, MG96, OSK^{+}96, PLZ^{+}95, RPW96, REA95, SKA^{+}96, SSL^{+}95, SBR^{+}96, WMC^{+}95a, BMP^{+}98, BSRG99, BLRS97, CSO97, CIK^{+}99, DOSN^{+}99, FLO^{+}99, FCMN97, GKS^{+}98, GM99, GSL98, GDM^{+}99, HLJ^{+}98, HBPJ98, HHR^{+}98, IAS^{+}98, JM97a, KA98, LNM^{+}98, LDH^{+}98, MCS^{+}97a, MFF97, MAC^{+}97, MDE^{+}99, MZNR98, MSB^{+}97, MSP99, NTAN98, NRP98, SMZ^{+}98, SSSS97, WHS^{+}97, cXP97, ZRHO99]. Apoptotic [MSEP98, SKC97, TMN97, YF99, AHM^{+}96, FHV^{+}96, LLN^{+}95, GDM^{+}99, KEP^{+}99, MNR^{+}98, RPS^{+}97, STK^{+}98, VCJA99, ZRHO99]. Apparatus [FNH99, LAW^{+}98, MAT^{+}98, ZWT^{+}97, DSM^{+}96, HM96, MKLS96, SLL96, SL96, KCFS98, LRS^{+}98, MHH^{+}99, SHP^{+}97, SCPPW98, VCS^{+}99, WBG^{+}97]. apparently [BGK^{+}95]. Appearance [DFL^{+}99]. appears [LKD^{+}95, RLVM98]. appraisal [Bor95]. Approaches [QBvD^{+}98]. apsA [FT95]. Arabidopsis [LWS^{+}97a, YNM^{+}99]. Arborization [KTAX98, PURB98]. Architecture [FAO^{+}96, FSIG^{+}95, LKWG96, MWT^{+}96, SPT^{+}95, EKS97, FHAP98, KGGK99]. area [DFC^{+}96]. ARF [HW98, PAL^{+}98, TBP^{+}99]. ARF1 [ODB98, NJB^{+}98, PHO^{+}95]. ARF6 [DSvDH^{+}98, PHO^{+}95, RKD96]. ARG118 [TSD^{+}97]. Arginase [GM99]. arginine [KMOO95]. Arginine-rich [YTO^{+}99]. ARIA [GYS^{+}95, LF95, TCM96]. ARIA/HRG [TCM96]. Arise [FGA^{+}98]. Arising [JM97a]. Ark1p [CYSD99]. arm [KPKWW95, WKK^{+}95, HS97, HOCK98, KMHK97, KD97b, KPW^{+}97, MKW^{+}99, PKBHK97]. Armadillo [CKP96, FFMG95, MW96, OP96a, SL96, WCW98]. arms [PMH96, ROG^{+}96]. ARP1 [HTKH96, EGM^{+}99]. Arp2 [BMC^{+}99, ELL^{+}99, KAP95, MMMW96, MSP97, SB99d, WDV^{+}97]. Arp2/3 [BMC^{+}99, ELL^{+}99, MSP97, SB99d, WDV^{+}97]. Arp3 [KAP95]. arranged [YvdEH^{+}95]. arrangements [XJM^{+}95]. Array [SB99d, AN99, SVS97, XCLM98]. Arrays [FWL^{+}98, GSC95]. arrest [BS95a, JGM96, TALM96, FK99, JCC^{+}98, OOB99]. Arrested [NNSN97, BBZ^{+}95]. arterial [BGR^{+}95]. Arteriolar [WMP^{+}98]. Articular [MSWL^{+}99]. articulins [HGP^{+}95]. Ascaris [RSS98]. asexual [FT95]. asialoglycoprotein [WGR^{+}95]. ASIP [IHT^{+}98]. Asp [MBS96a]. aspartate [WPS^{+}96]. aspects [BCC^{+}95, VMV^{+}95, GPKH99, PAN97]. Aspergillus [FT95, BRLM98, MGM95, WOM98]. Aspiration [HMCL97]. Assay [VG99, ABN96, MMF95, HHS^{+}99, NG99, RDP^{+}99, UPT98, WLK97]. assemble [SJ95, TBR^{+}96, EZC^{+}97]. assembled [RACHV96, TCSG96, ILSH99]. assembles [KGHMT95, LWM^{+}95, GJW^{+}97]. assemblies [VSB95]. Assembly [BBOE98, BGH^{+}97, CSN99, CDH^{+}98, DKLS^{+}99, DVL^{+}98, GHS98, HSML98, KMMC98, gLSE^{+}99, MF96a, MEFH97, MM99a, MB96, OBB99, RP97, SPV^{+}99, SLSW96, SVS97, WDV^{+}97, BTTB96, BCSG96, CKP96, DNB^{+}96, DSGF95, DNH^{+}96, DMD^{+}96, GF95, HM95, HSML96, HW96b, HH95b, HGG96, KSS^{+}95, LHDC95, LZD95, MMP^{+}95, MMF95, NCZ^{+}95, ORW^{+}95, RHHRB96, SHC95, SJC96, STS96, SJ95, SGBD96, SR95, WSGPS95, WM96, WKK^{+}95, WRW^{+}95, YSB96, ZN95a, ZN95b, CML^{+}98, CYH98, CHH97, DPG^{+}99, EHC97, FH97, FKM^{+}99, FRHS98, GKB^{+}98, GHW99, GCAHW97, GWB^{+}99, HTH^{+}97, HDDKH98, KGL^{+}99, KRS^{+}98b, LL97, Li97, LXC^{+}98, LWS99, LHWH97, LL98b, LFP98, MCJK98, MPSC98, MGIZ98, MRM^{+}98, MLC99, MS97b, MKW^{+}99, PDW99, PYM^{+}98, PHM^{+}98, PFAA98, RDP^{+}99, SWM^{+}98, SGA97, TGN97, WVM97, WZF97, WRdVC97, WHB^{+}99, ZCWB^{+}98]. Assembly [dS98a]. assisted [WJH^{+}98]. Associate [HPL97, LO96, CYSD99, FTB97, KDBW97, ZGYS99]. Associated [BdPE^{+}97, MKB^{+}99, RMY97, SGD98, WLF99, YHS^{+}99, ZWT^{+}97, BGB^{+}95a, BLLB95, CDRB^{+}96, CSO^{+}95, CM95a, CWSL97, CLF^{+}97b, FKNM97, GLD^{+}96, HCA^{+}95, HFSHK96, IRMF^{+}99, KR96, KOA^{+}99, KJBM^{+}96, KBR95, LGRB96, MN98, MTL^{+}96, MLLT95, NIM95, NNK^{+}97, OHB^{+}95, OS96b, PBT^{+}95, PPWL97, PT97, RHF^{+}96, RR96, SWS^{+}96, SCR98, SNCBM95, SGGS96, SVD96, SJW97, SJA96, TSD^{+}97, WSW95, WWP^{+}96, WM95, XGC96, XWK^{+}97, BRS99, EGM^{+}99, EJ97, FSC^{+}97, FNDS99, GZC^{+}99, HAB^{+}99, HYE^{+}98, IFM^{+}99, KZS^{+}98, LM98, LRAB99, MBW^{+}99, MGIZ98, MDK^{+}99, TMPM99, WSS97, ZR98, ZBW^{+}97, dBCKS98, BSM^{+}99, CSC^{+}98, FFB^{+}97, FKO^{+}98, GCD97, GF97, HIW^{+}98, IKR^{+}98, IFSV^{+}98, JM97a, JMD^{+}98, KOM^{+}98, KAS^{+}99, LPM^{+}99, MHWW98, MB96, MS99, MBB^{+}97, MHM^{+}97, PH97b, RSS98, SdCAH^{+}97, SYF^{+}97, SDC^{+}97, SV98, TKH^{+}97, TWP98, WBA^{+}98]. associated [YAC97, YAiN^{+}98]. Associates [MBA97, BKRT95, CLR^{+}96, DSM^{+}96, FK95, GHF^{+}95, HY95, HTKH96, TET^{+}96, WM95, vdBCH^{+}95, EJ99, IHT^{+}98, KLO^{+}98, hKsKC99, LLVS98, PWZS97, RBB97, RKB^{+}98, SdCAH^{+}97, SSW^{+}98, YPDM98]. Association [AARS^{+}97, CSC^{+}98, FKO^{+}98, HM96, KWD^{+}98, KSR96, LSZ^{+}99b, MSB99, MMD^{+}98, MZN97, SHT^{+}95, SLPE^{+}98, WBSNV97, vDSH^{+}97, BAC^{+}96, BSMH95, HSK^{+}96, LCC^{+}95, LvSC^{+}96, NAV96, PHB96, SH95, ZMG96b, BBP^{+}98, BM98, EPMB^{+}99, GFN^{+}99, HKJ^{+}99, HZC^{+}97, HVL^{+}99, HR99, JVRNM99, MGM98, MKB^{+}99, PDS^{+}97a, PAF97, RTM99, RKR^{+}99, SHB99, SMW^{+}99, SL97, SO97, SO98, TCMB98, WSSB99, WWD^{+}97, WFL^{+}99, WJWM^{+}98, ZCE^{+}98]. Associations [BLL^{+}96, LCS96, HFP^{+}98b, PSPG^{+}98]. AST1 [CF95]. aster [GSC95]. aster-like [GSC95]. asterless [BGG98]. asters [DFGL96]. Astral [SYM^{+}97]. astrocytes [GCGE^{+}96, BRD99]. asymmetric [REA95]. Asymmetries [LTR^{+}97]. Asymmetry [TYT^{+}99]. asynchronous [ASSS^{+}95]. AT-rich [dBCKS98]. ATDC5 [SSA^{+}96]. ATP [BGH^{+}97, FWBSO97, FNH99, JMD^{+}98, JGCS97, LHWH97, MCL^{+}96, MB96, OMS97, OSR^{+}98, RL99, REC97, SVT96, ZPL^{+}98]. ATP- [JGCS97]. ATP-binding [SVT96]. ATP-Dependent [RL99, FNH99, LHWH97]. ATP-independent [ZPL^{+}98]. ATP-induced [REC97]. ATPase [BMB97, CF95, CS97b, GBG^{+}96, GHS98, jLC97, OP96b, RCE^{+}99, TLBK98, CIRG99, MSF^{+}95]. Atrial [SGM^{+}97]. atrium [CFC^{+}95]. Atrophy [CAC^{+}99, EFML99]. AtT [BBW^{+}98, CHC97, CSEM99]. AtT-20 [BBW^{+}98, CHC97, CSEM99]. attached [CRRF97]. attachment [CLR^{+}96, HER95, MHK^{+}95, NHB^{+}95, WSGPS95, JB99, RWS^{+}99, VZNR98, WCMS98]. attachments [WGG96]. Atypical [EJ99, IHT^{+}98, ÜHK^{+}99]. AUG [VTG^{+}96]. AUG-initiated [VTG^{+}96]. Augment [KcZS97]. augmenting [SGYL^{+}96]. Augments [CCMG97]. Aurora [SGD98]. Aurora/Ipl1 [SGD98]. Aurora/Ipl1-related [SGD98]. Aut7p [KBI^{+}99]. Autoantibodies [MSA98]. autoantigen [HBJ95]. Autocrine [AGS^{+}97, BSN95, DCVF98, HKM^{+}96, SPR^{+}98]. Automatic [CS99]. Autophagic [DRE97, LGGS97]. Autophagosome [KBI^{+}99]. autophagy [AGB^{+}96]. autotypic [FSIG^{+}95]. auxilin [HUU96]. avian [TRJ^{+}95, WDL^{+}98]. avidity [LXK^{+}95, WAMS96]. Award [Sim99]. away [KR96]. axes [Gol95]. axial [CMM^{+}95, SH96, ALM97b, SWS97a]. Axin [FhJZ^{+}99]. axis [FFGG96, FFMG95, LTR^{+}97, MM97b, MRL^{+}99, VWG^{+}97]. Axl2 [SGTH99]. Axl2/Bud10p [SGTH99]. Axl2p [PB98]. Axo [DGP99]. Axo-Glial [DGP99]. Axon [BDO^{+}97, ZB98, BKEC96, BSJJ95, DLB96, GMC^{+}95, MGRS96, YSB96, AEVB98, LHB^{+}99, LDP^{+}99, WNB97, ZLM^{+}98, EHM^{+}95]. axon/Schwann [EHM^{+}95]. Axonal [BR99, MH95b, VGVF97, dCVCV95, BOLS95, IKT^{+}96, MWL^{+}96, OFY^{+}95, THK95, VLZR96, WMC^{+}95b, XML^{+}96, DGP99, EZC^{+}97, EFB^{+}98, GDB^{+}98, HSL^{+}99, KB99b, KNT^{+}99, RHW^{+}98, RPY^{+}99, ZP98, ZLL^{+}98]. Axoneme [LXC^{+}98]. axonemes [Sch96, MS97b]. axonin [BKB^{+}96, KZKS96, SPB^{+}95, KSG^{+}98, VZNR98]. axonin-1 [BKB^{+}96, KZKS96, SPB^{+}95, KSG^{+}98, VZNR98]. Axonopathy [ZLL^{+}98]. axons [BOLS95, CLS96, FTH96, HFSHK96, Kun95, Sch96, Bri99, EFK^{+}98, EFML99, HSL^{+}99, NTH98, RHW^{+}98, SZBB98, ULMT99, ZhTA^{+}97]. axotomy [TLF^{+}95]. axotomy- [TLF^{+}95].

B [HDV^{+}99, KLO^{+}98, TBD^{+}99, YNOH95, AWC99, CCP96, DRAT95, FWBSO97, GMC^{+}95, GMS^{+}96, GTL^{+}96, HBS^{+}97, HTK^{+}99, HDV^{+}99, HFE^{+}98, IMS^{+}98, KSV^{+}99, KKFN^{+}95, KMG^{+}97, KRMMA96, KGHG98, LLN^{+}95, LDH^{+}98, LSM^{+}98, MAT^{+}98, MAM95, MAE^{+}97, MEYPR95, NNSN97, NSYK^{+}95, OHB^{+}95, OK96, PCD^{+}98, RK95, SAC^{+}98, TYT^{+}97, WOM98, XZSC98, YNM^{+}99, YYH^{+}96]. B-alpha [KRMMA96]. B-cadherin [MEYPR95]. B-dependent [HDV^{+}99]. B-fragment [MAT^{+}98]. B-inhibiting [HDV^{+}99]. B-type [KKFN^{+}95]. B1 [JHM98, MYTK99]. B100 [DSM^{+}98a]. B16 [BBOB96]. B2 [LW96]. B3 [FKS95]. b5 [BAC^{+}96]. b559 [MML96]. B78 [DFL^{+}99]. bacteria [STP96]. Bacterial [ELL^{+}99, ES96, MGP97]. bacterium [MMP^{+}95]. baculovirus [PZ96, MPSC98]. Bak [MWGE97, GDM^{+}99]. Balance [HTK^{+}99, WXS96, PH97a]. Balbiani [SAEV^{+}98, VIF^{+}96]. Bamacan [WC97b]. Band [LRC^{+}99, FC95, OGS^{+}96, SBSG96, FSH^{+}99, TSK98, vHMV^{+}97, FPF^{+}99]. BAP31 [ABK^{+}97, NNK^{+}97]. BAPTA [MF96a]. Barbed [BMC^{+}99, BWKH96, MMP^{+}95, SJC96]. Bard1 [IFSV^{+}98]. barrier [BWF^{+}96, hCMPG97, DFD^{+}97, WG97]. Barriers [SK95a]. Basal [CSN99, GYS^{+}95, LYC^{+}95, BFRB99, DBG^{+}97, GCML98, HPKL97, PC98]. Basal-lateral [CSN99]. base [DH96]. Based [ELL^{+}99, SBS^{+}99, BD99, CLS^{+}97, DAF^{+}97, DTB96, GLS96, IIM^{+}99, INMT97, KTD^{+}97, LZK^{+}97, LLH^{+}99, LKSW98, MNS^{+}97, MMP^{+}95, MWOB96, MSF^{+}99b, NAV96, NSYK^{+}95, RdPRW98, RPV98, SYM^{+}97, STP96, TNM^{+}99, TNY^{+}95, WSS97, WMXE96, WBG^{+}97, ZPS96, DSC^{+}99]. Basement [KOA^{+}99, WC97b, DNB^{+}96, NKI^{+}95, CGA^{+}99, DHDJ^{+}97, GJW^{+}97, SVM^{+}99]. Bases [BD99, dBCKS98]. Basic [NNMW96, BKP^{+}95, HTS^{+}95, AAD^{+}97, KBC99]. basis [LKD^{+}95, HCM^{+}97, ZKN99]. basket [HUU96]. Basolateral [CWM^{+}98, vIZKH97, BWF^{+}96, HH95a, NDBR95, OPD^{+}96, ZZvdB^{+}96, DMY^{+}97, FGA^{+}98, KFM99, MGB^{+}98, OT97, vHMV^{+}97]. Basonuclin [hYGYF97]. Bassoon [DSVL^{+}98, WKL^{+}99]. Bax [EAOS^{+}98, MMK^{+}97, GGL^{+}98, DOSN^{+}99, KEP^{+}99, WHS^{+}97]. Bax-induced [EAOS^{+}98]. bazooka [MW96]. BB [KAAFZ95]. BC [PFAF97]. BC1 [MBBT98]. bcl [FVKS96, LLRTP95, ANT^{+}97, DDM98, FCMN97, FEF^{+}97, GBBSO96, HLMD97, KAS^{+}99, LLN^{+}95, LEB^{+}97, MWGE97, NNK^{+}97, OSK^{+}96, SPS97b, WMC^{+}95a, YKC95, ZRHO99]. bcl-2 [FVKS96, LLRTP95, ANT^{+}97, DDM98, FCMN97, FEF^{+}97, GBBSO96, HLMD97, KAS^{+}99, LLN^{+}95, LEB^{+}97, MWGE97, NNK^{+}97, OSK^{+}96, SPS97b, WMC^{+}95a, YKC95, ZRHO99]. Bcl-2/Bcl [NNK^{+}97]. Be [DFD^{+}99, FWL^{+}98, BKN^{+}97, LKD^{+}95, LF95]. bead [LGRB96]. beaded [GGR^{+}96]. beads [PHB96]. Bearing [DSG^{+}97, SFD^{+}98, YHMS^{+}98]. Becker [CRM^{+}96]. becomes [vdBCH^{+}95]. Bed [AEWG^{+}97]. BED1 [MR96]. Bee1 [Li97]. before [BMS^{+}97, SAEV^{+}98, VLAHV98]. Begin [HT98]. Behave [RWOA97]. behavior [BEG^{+}95, PPW^{+}95, RD96a, SHS96, RKRG99, WWD^{+}97]. Behaviors [HKBM99]. Behmel [CGSY^{+}99, GKS^{+}98]. belonging [CBB^{+}96, YPM97]. Belongs [RHPW97]. bent [ASSS^{+}95]. benzyl [HHRdB^{+}98]. besides [OSH^{+}96]. Bet1p [ZWT^{+}97]. Beta [BZB^{+}96, SOE^{+}95, WLG^{+}96, AGPG96, AM95b, BLE^{+}96, BLLB95, BKEC96, DMS^{+}95, DMP^{+}96, DSM^{+}96, EHM^{+}95, ES96, FFGG96, FPM^{+}95, FBD^{+}95, FSAP95, FFMG95, GLD^{+}96, GBG^{+}96, GKR95, HKC^{+}95, HSL^{+}95, HGW95, HMB95b, HWC^{+}96, HHM^{+}95, HCAW95, KLMR95, LCC^{+}95, LRM^{+}95, LCS96, LMB^{+}95, MRvD^{+}95, MTLW95, MBS^{+}96b, NIM95, OP96a, PHU^{+}95, RZZ96, SOHP95, SMB^{+}95, SMFC95, SEC^{+}95, SR96, SYO95, TNY^{+}95, TET^{+}96, WAMS96, WVVD95, YSK95, ZMM^{+}95, DMS^{+}95, ED96, LLZ^{+}96]. beta-actin [SMB^{+}95]. beta-amyloid [AM95b, HKC^{+}95, YSK95]. beta-Catenin [TET^{+}96, BLE^{+}96, FFGG96, FFMG95, LCC^{+}95, OP96a]. beta-dependent [FSAP95]. beta-galactosidase [SL95]. beta-receptors [SR96]. beta-tectorin [KLMR95]. beta1 [SR96]. Beta4 [DYF96]. Between [LCD^{+}98, AFFS96, ABRB95a, BMC^{+}99, BSY97, BFL95, BVKP98, CT97b, CSY^{+}95, CS96b, DSC^{+}99, DW97, FST99, GCAHW97, GCML98, GW95, HCAC99, HBJ95, HHTT95, HPY^{+}98, HKKH99, KKM^{+}97, KMMC98, KD97a, KSR96, LAW^{+}98, LWS^{+}97b, LFP98, LS95b, MCB96, MNM^{+}98, MDS95, MGD97, NDBR95, NAsG^{+}98, NB97, POR^{+}97, RdPRW98, SBG^{+}98, SZBB98, SDTE95, SNCH95, SPB^{+}95, TH97, TWS97, VV95, VvdKMvD99, WKR^{+}98, XWK^{+}97, ZMG96b, CSC^{+}98, GB98, IYM98, SF98, XCLM98]. Beyond [KC99, JCC^{+}98, DFMG97, HPL99]. bFGF [KvEdV^{+}96, YFST^{+}96]. BHK [YKRS96]. BHK21 [NNSN97]. Bi [CCMG97, SWR^{+}96, BSB^{+}97]. Bi-directional [BSB^{+}97]. Bi-transgenic [CCMG97]. biased [KKG95]. Bid [DOSN^{+}99, KEP^{+}99]. Bid-induced [DOSN^{+}99]. bidirectional [LZC^{+}95]. bifunctional [ZSMV95]. Bifurcation [CRYC99]. Biglycan [LHR^{+}97]. bilayer [MWC95]. Bim1p [TOB^{+}99]. BIN1 [BDO^{+}97]. Bind [VB98, SOHP95, SYO95, DDF^{+}98, HVL^{+}99, YHD^{+}98]. Binding [BGH^{+}97, EGKC^{+}99, FFGG96, FSDA97, GFN^{+}99, KSE^{+}97, KGL^{+}99, MK95a, NCV^{+}98, SPB^{+}95, WH97, BLE^{+}96, BLG95, BJS^{+}95, BMG^{+}95, BYG96, BSJJ95, BPT96, BS95b, CBB^{+}96, CMHT96, CLT^{+}96, CRM^{+}96, DMS^{+}95, DvD95, DLB96, DGSR95, DWS^{+}95, FBH^{+}95, GLS96, HIG^{+}95, HSML96, HGH96, IRMF^{+}99, JSR96, KPKWW95, KBM^{+}95, LJ95, LGRB96, LF95, MN98, MNS^{+}99, MCL^{+}96, MFMW95, MMATe96, MMMT95, MSP97, MEYPR95, NIM95, NH95, NHB^{+}95, PPWL97, PMCT96, RRRA95, RLCM96, RFNF96, RS95, SRH^{+}95, SRH^{+}96, SOR^{+}96, SZB^{+}95, SSF^{+}95, TGG95, TOSF^{+}98, VMH97, VIF^{+}96, WXS96, WHH96, YGG97b, YTT99, ASP99, BJR^{+}99, CSY98, DKN^{+}97, GRTB97, GFN^{+}99, GDB^{+}97, HMD97, HSL^{+}99, HPD^{+}97, INMT97, IFM^{+}99, KSV^{+}99, KFO^{+}99, LMW99, MAW98, MSFA99, NMK^{+}97, NGMR97, OSR^{+}98, STG^{+}99, STF98, SZNS98, qSmLY97, SAEV^{+}98, TSD^{+}97, TGTL97, TWP98]. Binding [TRA^{+}99, WGF^{+}98, BFHB97, BH98, CLF^{+}97b, DBOM98, DDRC99, EJ97, FKO^{+}98, GKP^{+}97, HTS^{+}95, HBS^{+}97, IKR^{+}98, JMTCF96, KSK^{+}99b, LM98, LVD96, LLNH^{+}98, MCJK98, MNS^{+}97, NOS^{+}97, NIK^{+}97, OWAlW99, ONI^{+}98, OFM^{+}98, PKR95, PRB97, PRB99, PCQH98, RKB^{+}98, RHA^{+}97, RPBB98, SLW^{+}96, SLEB95, SCG^{+}95, SOJM99, SVT96, WA95, WSH98, YCR^{+}98, ZR98, dBCKS98]. Binds [CJS^{+}99, CWSL97, CWM^{+}98, HDV^{+}99, RL99, AK95, DSM^{+}96, DHT96, GEK95, GEW^{+}95, VV95, VIF^{+}96, YtDH^{+}95, APH^{+}97, ABE^{+}98, BALL98, BFHB97, CGM^{+}99, DBG^{+}97, KBB^{+}97a, MTW^{+}99, PFKvD99, RKR^{+}97, SSH^{+}98, TPMB97, TBP^{+}99, WKL^{+}99, ZOKS99, ZLB^{+}97]. Biochemical [CBTZ98, LHS96, OP96b, SWR^{+}96, AGB^{+}96, CGM^{+}95, MML96, MMF95, KGGK99, qXLCH97]. biochemically [MF96a, HBH98]. Biogenesis [AAB^{+}97, CSCM97, DCOK95, PAL^{+}98, RN99, DEG^{+}96, GMS^{+}96, FBM99, GPSF99, GAPG97]. biologic [MBD^{+}96]. biological [MW95a, TRJ^{+}95]. Biologically [DPM^{+}97]. Biology [CQB^{+}96, Sim99]. Listeria [LLH^{+}99]. biosynthesis [BLC95, BvdBvM96]. biosynthetic [CMM96, VOP^{+}95, YKRS96, DRE97, OT97, RPE^{+}97]. BiP [CS97b, HVL^{+}99, KA95, LS95b, SHR^{+}95, SFNRH95]. BiP/ [SFNRH95]. BiP/Kar2p [SHR^{+}95]. bipartite [GBMS95]. Bipolar [WGG96, VH96, JL98, SMB^{+}99, YAD97]. Bipotential [SATW98]. bisphosphatase [CC98, CC98]. bisphosphate [GHF^{+}95, KLS^{+}99]. bladder [RVD^{+}95]. Blastocyst [JDK97]. blastocysts [FPM^{+}95]. blastoderm [MM95a, MW96]. blebbing [Cun95, HHR^{+}98, MSEP98]. Bleeding [KZS^{+}98]. blister [MZL^{+}99]. Block [BBL^{+}99a, OST^{+}95]. blocked [JCPK95, TMM^{+}95, MFBK99]. Blocking [NRP98, WPW^{+}97]. Blocks [HHRdB^{+}98, BRHD95, NH95, KBF^{+}98, NBJ^{+}99, YSEI^{+}99]. Blood [ILSH99, MWC95, DFD^{+}97, WSR^{+}97]. blown [DFMG97]. BM28 [TAK95]. BMP [BSK96, CJH^{+}98, ZOKS99]. BMP-2 [ZOKS99]. BMP-6 [BSK96]. Bni1p [LKE^{+}99]. bodies [BFL95, CSO^{+}95, HFSHK96, KP96, ASACF98, CRGW98, FFLP99, GCML98, IYM98, JWW^{+}99, ZRHO99]. Body [ISR99, OBB99, SGD98, WH97, AWS96, DK96, FSHD96, HY95, MCA96, SCG^{+}95, SGGS96, SGBD96, WW96a, AK99, CYH98, HPKL97, HFE^{+}98, MB97, MUS98, SGSW97, SCEB99, SMM99, SVMB97]. body-wall [MCA96]. Bond [CS99]. Bonded [RDS^{+}99, KGHG98]. Bonds [LG99]. Bone [EIIM^{+}98, WSH98, BSK96, CCR96, AMY^{+}98, CJH^{+}98, CR99, IKR^{+}98, NMT97, SKS99, WSH98]. Border [BZL^{+}98, JM97b]. bordering [UHJ^{+}96]. Borders [FCTPA^{+}99]. boss [KP96]. Both [ASSS^{+}95, BGH^{+}97, TB99, ETF95, GDC95, HH95b, MWC95, OP96a, RRRA95, SJ95, WCW98, ALVMM98, BKK^{+}97, CLF^{+}97b, DDF^{+}98, HTY^{+}99, IW97, KQB^{+}98, LBBP98, SFR^{+}98, SZNS98, WS98]. Botulinum [NBJ^{+}99, IKT^{+}96]. bound [OCS^{+}95, dBCKS98, FHWV97, ZBW^{+}97]. boundaries [TK95]. Bovine [PAN97, SN97, WSH98]. box [BBOB96, KKAN^{+}96, RGS99]. BP180 [BKN^{+}97, HBJ95]. BPAG1 [LSL99]. BPAG1-n [LSL99]. Brain [BDO^{+}97, PT97, GCM^{+}95, HTS^{+}95, Kun95, SWR^{+}96, TLF^{+}95, WYY95, ASP99, BSM^{+}99, DFD^{+}97, LYLR98, MSF^{+}99b, SVB^{+}97, TNCM97]. brain-derived [TLF^{+}95]. Braking [CS99]. Branch [Nei98]. branching [AC95a, KKD^{+}95, AGP^{+}98, DCR98, NBS^{+}98]. Bravo [SPB^{+}95]. Brca1 [IFSV^{+}98]. Brca1-associated [IFSV^{+}98]. Breakage [WSS97]. Breakdown [KTD^{+}97, STRM95, ZKSB96]. Breaking [BCSG96]. breaks [DH96, RBRB99]. breast [KCDB95, ARR^{+}98, KRCP99, NPJW99, WPW^{+}97, WBG^{+}97]. brefeldin [LCK95, KBF^{+}98, LLG^{+}99, SPL^{+}97, SPS^{+}97a]. bridge [NCZ^{+}95, SCG^{+}95]. Brief [Dov99]. bristle [HCM96a]. bristles [TTG95, TCSG96]. Bronchial [SSZ^{+}97]. Bronchitis [MM97a]. brown [VXBH95]. brucei [MGB^{+}97, RFK^{+}97, SSF^{+}95, WCV^{+}97]. Brush [BZL^{+}98, JM97b, FCTPA^{+}99]. bub1 [BBL^{+}99a, BHJ98, THM98]. Bub1-related [THM98]. Bub2 [FFLP99]. Bub3 [THM98]. Bubr1 [CJS^{+}99]. bud [BMLM96, CP95, KM95, MMATe96, MR96, TN95, MBTW97, OPB98, SYM^{+}97, YAD97, SH96]. bud-site [CP95]. Bud10p [SGTH99]. Bud3p [CMM^{+}95]. BUD4 [SH96]. Budding [FFLP99, CMM^{+}95, HM95, KM95, LR95, LZD95, SH96, SOR^{+}95, SOG96, SHG96, TMDD96, VH96, WM96, YGK96a, YSC^{+}95, AWB^{+}98, ASP^{+}97, BSY97, BVKP98, CYSD99, FHK97, GDL98, KsYA98, LL97, LL98b, MCM^{+}99, OMS98, PSB98, SIUW98, SSM98, YYSB97]. Building [VLAHV98]. Bullous [SBG^{+}98, HBJ95, BKN^{+}97]. Bundle [TCV^{+}98]. bundles [BCSG96, TTG95, TCSG96, TTG96, VSB95, WW96b, CSM97, SMB^{+}99]. Bundling [MELC96, ASY96, BZL^{+}98, TET^{+}96, TTG95, HYE^{+}98]. Bungarotoxin [RDS^{+}99]. burst [MBD^{+}96]. Butyrylcholinesterase [FKM^{+}99]. BY-2 [SGS95]. bypasses [dPF95]. byr4 [SMC^{+}96]. bZIP [KWN^{+}99].

C [CBM^{+}98, FHY99, GSC^{+}96, CGSP95, DOSN^{+}99, MDE^{+}99, NRP98, NBS^{+}98, SHK^{+}99, VDA^{+}96, VCJA99, YTY^{+}99, AARS^{+}97, BH98, BYG96, CRM^{+}97, DRR96, EAOS^{+}98, FMM^{+}98, IMW^{+}99, ISFB^{+}96, JL98, JCR97, KHA^{+}99, KNT^{+}99, LCS96, MGP97, MLB98, MSH^{+}97, MYC^{+}98, MDC^{+}99, ND96, OTQM98, OIT^{+}95, PHN^{+}98, PMCT96, PPBJ96, RTM99, RL96, RTP^{+}98, SSI^{+}97, SBGR98, SKY^{+}98, SIAS96, SZB^{+}95, qSmLY97, TOT^{+}96, TWiK^{+}99, ÜHK^{+}99, WCV^{+}97, XZSC96, XC97, ZBG99, vHGB^{+}97, ESH^{+}99, FSR^{+}96, NBS^{+}98, PHA^{+}98, RME^{+}95, SKC97]. C- [ÜHK^{+}99, XC97]. C-cadherin [BYG96, ZBG99]. c-erbB2 [NBS^{+}98]. c-initiated [SHK^{+}99]. C-Jun [IMW^{+}99, SKC97]. c-kit [VDA^{+}96, SBGR98]. C-like [SIAS96]. c-met [NBS^{+}98]. C-Nap1 [FMM^{+}98]. c-Src [CGSP95, AARS^{+}97]. C-src/Microtubule [AARS^{+}97]. C-Type [FHY99, BH98]. C-zeta [XZSC96]. C. [GKR95, MCA96]. C1 [IKT^{+}96]. C2 [GSKL99, MPW95]. C2a [SEB99]. C2b [SEB99]. Ca [SLW^{+}96, TT96, AMY^{+}98, ABM^{+}99, BHG^{+}97, BCC^{+}95, BML^{+}97, BRD99, BCWA97, CT97b, CBTZ98, EKF^{+}97, FCTPA^{+}99, GL95, HFP98a, IMK99, KJBM^{+}96, LCD^{+}98, LYH^{+}99, LZC^{+}95, MLW^{+}95, MAB98, MP98, MAS^{+}97, NYN^{+}95, PT97, RL96, RROA95, RWOA97, SM95a, SMH98, TT96, TMM97, TBD^{+}99, WWM^{+}98, YHMS^{+}98, vdBCH^{+}95]. CAAX [TNFK^{+}98]. Cab45 [SLW^{+}96]. Cables [GCST97, KMMC98, PSB98]. Caco [Sal99]. Caco-2 [Sal99]. Cadherin [CSN99, HMS^{+}98, NPJW99, RFCR^{+}97, SSR^{+}98, TNM^{+}99, ANS96, AYF96, ABN96, ABRB^{+}95b, BEZ^{+}95, BLE^{+}96, FSIG^{+}95, HG95b, KSJW95, KSR96, LCC^{+}95, MAFJ^{+}95, MEYPR95, SSS^{+}95, TNY^{+}95, TYSP^{+}96, UHJ^{+}96, ANRT99, BALL98, BMHH97, BGBK99, CMR^{+}97, CRA^{+}98, DBJDT99, GvdFvD^{+}99, HLK^{+}98, IIM^{+}99, INMT97, KBB^{+}97a, KBB^{+}97b, LYS99, MNS^{+}97, MNM^{+}98, NRD98, NPJW99, NGT99, OK98, SNT^{+}98, WKR^{+}98, YNG98, BLE^{+}96, BYG96, CT98, ICW^{+}96, MGD97, MEYPR95, RS95, SGYL^{+}96, ZBG99]. Cadherin-6 [MNM^{+}98]. Cadherin-7 [DBJDT99]. Cadherin-based [TNM^{+}99, TNY^{+}95, IIM^{+}99, INMT97, MNS^{+}97]. cadherin-binding [MEYPR95]. cadherin-catenin [KSR96]. cadherin-catenin-actin [ANS96]. Cadherin-dependent [BMHH97, BGBK99]. cadherin-mediated [BEZ^{+}95, BLE^{+}96]. Cadherin/ [CSN99]. cadherin/catenin [KSJW95, SHT^{+}95, SSS^{+}95]. Cadherin/Green [ACSN98]. Cadherins [CT97b, RNK97, BKRT95, CLT^{+}96, FFGG96, ZMG96b, BKMN^{+}98, LBW^{+}97, NRD98]. Caenorhabditis [BBOE98, BTTB96, CDH^{+}98, CRA^{+}98, FR98, Gol95, GPKH99, GSK^{+}99, GJW^{+}97, GGK97, HMC^{+}99, HW96b, HSSW99, IMFS96, IHT^{+}98, KNT^{+}99, LHK^{+}96, LBO^{+}98, MCTMK98, MAB98, MMR99b, MGA^{+}96, OBB99, OFY^{+}95, SGD98, SCEB99, SWO^{+}99, SLP^{+}96, TWiK^{+}99, WG99]. Caffeine [KGE^{+}98]. Caffeine-sensitive [KGE^{+}98]. CAG [DSVL^{+}98, ZL95a]. CAG/Glutamine [DSVL^{+}98]. CAG/Glutamine-repeat [DSVL^{+}98]. Calcineurin [SM95a]. Calcitonin [CKM95]. Calcium [BAS95, BSW^{+}99, CRM^{+}97, HCPW^{+}99, JL98, JGCS97, MKI^{+}96, TT96, VB98, WPS^{+}98, WMP^{+}98, BSJJ95, CM95a, CMHT96, FHCL95, GCM^{+}95, GG95a, LW96, LK95, PLR^{+}95, SWR^{+}96, SH95, VBZ96, WWP^{+}96, ZMG^{+}96a, BR98, CT98, DLC98, EKF^{+}97, FRHS98, GCE^{+}97, HLMD97, HFL97, KBF^{+}98, mLZB99, LLNH^{+}98, OMS97, PFP97, PFAA98, REC97, SOJM99, WMSM97, ZOS^{+}97]. Calcium- [VB98]. Calcium-activated [ZOS^{+}97]. calcium-binding [BSJJ95, CMHT96, LLNH^{+}98, SOJM99]. calcium-independent [ZMG^{+}96a]. Calcium-regulated [BAS95]. Calcium/Calmodulin [BSW^{+}99]. Calcium/Calmodulin-dependent [BSW^{+}99]. caldesmon [COML95]. CALEB [SVB^{+}97]. Calibers [EFB^{+}98, EFK^{+}98]. calmodulin [CBB^{+}96, CRMB95, SGBD96]. calmodulin-binding [CBB^{+}96]. Calmodulin-dependent [BSW^{+}99]. Calnexin [KA95, MB96, BCAS98, HZC^{+}97, TH97]. CALNUC [LLNH^{+}98, LYH^{+}99]. Calpain [PTJ^{+}98, SMH98, SGGW99]. Calreticulin [MNZ^{+}99, OSPJ^{+}96, BCC^{+}95, SAPLHD95, BCAS98, HZC^{+}97, JLC98, WMSM97]. Calsensin [BSJJ95]. calsequestrin [RPC^{+}95]. CAM [ETP^{+}99, LBW^{+}97, BHP^{+}95, CCB^{+}98, MGB^{+}98, SLN^{+}97, OIT^{+}95]. CAM/ [SPB^{+}95]. cAMP [BBOB96, BAH^{+}98, GLMG98, HG95a, ISFB^{+}96, OSM98, OMS97, RKB^{+}98, dPSMNSM95]. cAMP-dependent [GLMG98, ISFB^{+}96]. cAMP-regulated [RKB^{+}98]. camptothecin [MG96]. Can [DFD^{+}99, LKWG96, LF95, MTGY96, ND96, RSC96, RFNF96, SHR^{+}95, SJ95, BdPE^{+}97, BSRG99, GBS99, HVL^{+}99, KCM^{+}97, MMM^{+}98, SFD^{+}98, YPB^{+}98]. CAN/Nup214 [BdPE^{+}97]. Canal [gLSE^{+}99, RC97]. canalicular [MKB^{+}95]. canals [TTG96]. Cancer [ARR^{+}98, HYE^{+}98, NPJW99, WBG^{+}97, vHGB^{+}97]. Candidate [FMM^{+}98]. canine [ZZvdB^{+}96, vGvM95, ALK^{+}99, BRS99, DMY^{+}97, OT97, PMBM^{+}99]. cap [GTC^{+}98, VIF^{+}96, KBC99, MB97]. cap-binding [VIF^{+}96]. Capacitative [HFP98a, FHCL95]. capacity [BCC^{+}95, DSE^{+}95, MR95, SRH^{+}96, JDK97]. CapG [SKWY95]. Capillaries [SPR^{+}98]. capillary [CCB^{+}95]. Capping [EHC97, BWKH96, CLL95, HCM96a, KTC95, SJC96, HC99]. Capsids [LHWH97, SEH97b]. Capture [KRS98a]. CapZ [SHC95]. cAR1 [VBG96]. carbon [AGB^{+}96]. carboxy [CPL^{+}95, HAS95, OGS^{+}96, DW97]. carboxy-terminal [HAS95, OGS^{+}96]. carboxyl [NCZ^{+}95]. carboxyl-terminal [NCZ^{+}95]. carboxylase [STN^{+}96]. carboxypeptidase [SFNRH95]. Carcinoembryonic [SPS97b]. Carcinogenesis [MDC^{+}99]. Carcinoma [BMC^{+}99, CCMC97, DSC^{+}95, GSvR^{+}96, ICW^{+}96, RVD^{+}95, ANRT99, GKO99, MAC^{+}97, OSM98, RM97, RTM99, SPBCM97]. Cardiac [BGH^{+}97, MAM99, CLL95, GF95, KAT^{+}95, LK96, SPT^{+}95, CMS98, KFL^{+}98, LRC^{+}99, MNZ^{+}99, SNK^{+}97, ZTH^{+}97]. Cardiac-specific [ZTH^{+}97]. cardiomyocyte [vABSP95]. Cardiomyocytes [DLC98, PRV98, SGM^{+}97]. Cardiomyopathy [BGH^{+}97]. Cargo [AWB^{+}98, GE97, LPD^{+}99]. Cargo-selective [GE97]. carried [NSYK^{+}95]. Cartilage [CHLÅ97, HBH98, AKP95, TKM^{+}96, CAR^{+}97, CGA^{+}99, KPZ^{+}97]. CAS [KLM^{+}98]. CAS/ [KLM^{+}98]. Cascade [SHK^{+}99, SP96, TDW95, FRHS98, LGM^{+}97]. case [GJ95b, RFNF96]. casein [GHF^{+}95]. CASK [CWM^{+}98, HYK^{+}98]. CASK/LIN [CWM^{+}98, HYK^{+}98]. CASK/LIN-2 [CWM^{+}98, HYK^{+}98]. Caspase [BRM^{+}99, CSO97, KOM^{+}98, SHK^{+}99, BLRS97, MNR^{+}98, MMK^{+}97, STK^{+}98]. Caspase-3 [MNR^{+}98]. Caspase-9-dependent [SHK^{+}99]. Caspase-dependent [BLRS97, STK^{+}98]. Caspase-independent [KOM^{+}98, MMK^{+}97, STK^{+}98]. Caspases [MMAK97, SHK^{+}99, IJR98, SMZ^{+}98, ZRHO99]. Caspases-2 [SHK^{+}99]. Caspr [EZC^{+}97]. cassette [SVT96]. castellanii [BCMK95, JOPM98]. catalase [PL96]. Catalytic [ZLPL99, ISFB^{+}96, OCS^{+}95, MKS^{+}97]. catalyzed [AKP95, CFL^{+}98, MW97, SSS99a]. cataract [DSvdV^{+}96, LKD^{+}95]. Cataracts [WGP98]. catecholamine [CGVS^{+}96]. Catenin [BPA^{+}97, CSN99, TET^{+}96, WUUI^{+}98, vHGB^{+}97, ANS96, BLE^{+}96, FFGG96, FFMG95, KSJW95, KSR96, LCC^{+}95, LPM^{+}99, OP96a, SHT^{+}95, SSS^{+}95, TNY^{+}95, UHJ^{+}96, BLRS97, CRA^{+}98, EPMB^{+}99, GvdFvD^{+}99, IIM^{+}99, INMT97, LTR^{+}97, MM97b, NG99, OOB99, PBA^{+}97, SSR^{+}97, SGR97, SSS^{+}98, WKR^{+}98, WAV98, WRG98]. catenin/cadherin [UHJ^{+}96]. Catenins [RHHRB96, BKRT95, RS95, CT98, MDV98]. CATGTG [BBOB96]. Cathepsin [DBLD^{+}99, GTL^{+}96, OCS^{+}95, PFHWN98]. cation [AZF96, RSJK95, SKR96, PFHWN98, ZOS^{+}97]. cation-dependent [RSJK95, SKR96]. Cation-independent [PFHWN98]. Cause [BBL^{+}99a, HTW97, KEP^{+}99]. Caused [MAB98, WMC^{+}95a, LMS97, PHA^{+}98]. Causes [KMI^{+}97, ADP^{+}95, Dav95, FSR^{+}96, GDC95, JGD^{+}96, SSY^{+}96, SKA^{+}96, AFG^{+}98, KKT^{+}98, NMH^{+}98, PFHWN98, SYT97]. causing [KKA96]. caveolae [CSY^{+}95, PGS95, SRR96b, SYA95, SM95c, HKOM98, LMC^{+}98, MYC^{+}98, OMS98, OF98, WJWM^{+}98]. Caveolae-like [OF98, WJWM^{+}98]. Caveolar [SGJP99]. Caveolin [CSY^{+}95, LMC^{+}98, LSZ^{+}99b, GH95, KSP95, MBS95, PWZS97, SVF^{+}98, WYL^{+}99]. Caveolin-1 [SVF^{+}98]. Caveolin-3 [PWZS97]. CBF3 [RGS99]. CC [WAMS96]. CCR7 [CBM^{+}98]. CD1 [DAP^{+}95]. CD11b [DAP^{+}95, LXK^{+}95, LRG^{+}98]. CD11b/CD1 [DAP^{+}95]. CD11b/CD18 [LXK^{+}95]. CD11c [LXK^{+}95]. CD11c/CD18 [LXK^{+}95]. CD151 [YMAC^{+}98]. CD151/PETA [YMAC^{+}98]. CD151/PETA-3 [YMAC^{+}98]. CD18 [LXK^{+}95]. CD2 [DFC^{+}96, SMFC95]. CD23 [HAB^{+}99]. CD3 [DNH^{+}96, DKN^{+}97]. CD31 [FSDA97, FBH^{+}95, LCN^{+}97, PHU^{+}95]. CD31/ [PHU^{+}95]. CD34 [PFGS95, SCM^{+}96]. CD36 [DPZ^{+}97, KJE^{+}99]. CD36/LIMP [KJE^{+}99]. Cd38 [SAM^{+}99]. Cd38/ [SAM^{+}99]. CD4 [BINRM97]. CD43 [YHD^{+}98]. CD437 [SKA^{+}96]. CD44 [BNW^{+}99, BNA^{+}96, BJS^{+}95, BMG^{+}95, CAS96, DSE^{+}95, GSvR^{+}96, JBD^{+}95, SZNS98, SRH^{+}96, SSV98, WSR^{+}97, YHD^{+}98, ZKH^{+}95, OPH^{+}99]. Cd44-Containing [OPH^{+}99]. CD44-mediated [BNA^{+}96]. CD45 [RR96]. Cd47 [GZC^{+}99, HAB^{+}99, PCL^{+}96]. CD59 [vdBCH^{+}95]. CD62L [LKKL97]. CD73 [ANS^{+}97]. CD81 [YMAC^{+}98, TH99]. CD81/TAPA [YMAC^{+}98]. CD81/TAPA-1 [YMAC^{+}98]. CD9 [HIG^{+}95, TH99, NIM95]. CD9/DRAP [HIG^{+}95]. CD95 [AKM^{+}98, GSvR^{+}96]. CD95L [AKM^{+}98]. cdc12p [CDN97]. cdc15 [LL98a]. cdc15/PSTPIP [LL98a]. cdc2 [OSH^{+}96, JGM96, MYTK99]. Cdc2/Cyclin [MYTK99]. CDC20 [DRAT95]. Cdc24p [NA99]. Cdc25B [NNSN97, KKH^{+}99]. Cdc25c [KKH^{+}99]. Cdc31p [SCG^{+}95, SBR98]. Cdc31p-binding [SCG^{+}95]. CDC37 [SGSW97]. cdc4 [MBP^{+}95]. CDC42 [BOM^{+}98, EAL^{+}95, AZRJ98, BFHB97, EWS96, ELL^{+}99, MCJK98, MM96, SMTN99, ZJB^{+}97, ZJY^{+}98]. Cdc42-Activated [SMTN99]. Cdc42-binding [BFHB97]. Cdc42-induced [ZJY^{+}98]. Cdc42p [OC98]. cdc6 [HN98]. cdk [DVL^{+}98, MCC^{+}97, MYTK99]. Cdk-dependent [MCC^{+}97]. Cdk/Cyclin [MYTK99]. CDK1 [WHG^{+}97]. CDK2 [FLW^{+}95, HN98, YN95, BBJ98, HYN97, MYTK99, ZOB^{+}96]. Cdk2/Cyclin [MYTK99]. CDMP [TTAH^{+}99]. CDMP-1 [TTAH^{+}99]. cDNA [AASH^{+}96, WKL^{+}96, WC97b]. CDO [KGF^{+}97, KMM^{+}98]. Cdx2 [LDD^{+}97]. Ced [JWR96, MWGE97]. Ced-3 [JWR96, MWGE97]. Ced-3/ICE-family [JWR96]. Ced-3/ICE-related [MWGE97]. Cek9 [SHPP96]. Cell [ACSN98, AWC99, BPA^{+}97, BMPP99, BGN^{+}96, BKB^{+}96, CHP^{+}98, CLK99, CSF^{+}97, CT97b, CCMG97, DH99, EY99, FSDA97, GR97, GZK^{+}95, Gol95, GMFN95, GDM^{+}99, GXE^{+}99, HW98, HB98, HRD98, HSML98, ISR99, JB99, JCR97, KMI^{+}97, KBB^{+}97b, LSB98, LMP98, LCN^{+}97, MCC^{+}97, MNS^{+}99, MWGE97, MKS^{+}97, MSF^{+}99a, MHM99, MM96, NAV96, OCS^{+}95, PBCS98, RYB^{+}99, RdPRW98, RBM^{+}99, SvDtK^{+}98, SS95a, SVH^{+}98, SW99, SKC97, SLPE^{+}98, SWVR98, TNM^{+}99, VG99, VNTM^{+}99, YMAC^{+}98, ANS96, ASY96, AHBW96, AWS96, AW96, ABN96, BEZ^{+}95, BNA^{+}96, BOLS95, BAS95, BKP^{+}95, BSD^{+}96, BCSG96, BGK^{+}95, BHP^{+}95, CDRB^{+}96, CCB^{+}95, CMUW96, CKM95, CKP96, CM95b, Cun95, DLB96, DRAT95, DNB^{+}96, DZC^{+}96, DSGF95, DCW^{+}96, DNH^{+}96, DYF96, DMD^{+}96, EHM^{+}95, EM96, ETF95, FLW^{+}95, FHCL95, FSIG^{+}95, FK95, FPM^{+}95, FBH^{+}95]. cell [FWHW96, FMJ^{+}95, FSR^{+}96, FA95, FBD^{+}95, FUB^{+}96, FSAP95, FHV^{+}96, FSHD96, FVRCH96, GW95, GTL^{+}96, GKC^{+}96, GJ95b, HP95, HB95a, HAS^{+}96, HG95b, HCW96, HFSHK96, HGW95, HGH96, ICW^{+}96, JWR96, KSJW95, KJBM^{+}96, KZKS96, KF95, LCC^{+}95, LG95, LR96, LR95, LKD^{+}95, LZD95, LO96, LDF95b, LK95, LLRTP95, LMB^{+}95, MAFJ^{+}95, MAS^{+}95, MAM95, MFM^{+}95, MK95a, MW95a, MWC95, MW95b, MLLT95, MMMT95, MBS^{+}96b, MG96, MXSRB96, NIM95, NAP^{+}96, NWG95, NHB^{+}95, NMJL96, OWA95, OSPJ^{+}96, OK96, OP96a, OABD^{+}96, PNL^{+}95, PPF96, RMWW95, RME^{+}95, RWM^{+}96, RD96a, RHF^{+}96, RFNF96, REA95, RSP^{+}95, RVD^{+}95, RSR^{+}96, RRH^{+}95, SRR96a, SGS95, SH96, SPMB^{+}96, SRR96b, SM95a, SSA^{+}96, SSP95, SS95b, SNCBM95, SEC^{+}95, SLP^{+}96, SVD96, SRRB95, SPB^{+}95, TKVC95, TIJ^{+}95, TNY^{+}95, TALM96, TMM^{+}95, TFF^{+}95, VBG96]. cell [VKP^{+}95, VMN95, VKR^{+}96, VOP^{+}95, WPS^{+}96, WJC^{+}96, WMXE96, WvdVV^{+}95, WPMW95, WHP^{+}96, WFBO^{+}96, YSK95, YSC^{+}95, ZL95a, ZOB^{+}96, ZZvdB^{+}96, ARR^{+}98, AEWG^{+}97, Ass97, ASP^{+}97, BSW^{+}98, BGBS98, BHHF97, BBL^{+}99b, BOM^{+}98, BMHH97, BLRS97, BGR98, BR98, BHSAJ98, BSM^{+}99, CCS^{+}98, CAR^{+}97, CS97a, CSO97, CDN97, CCB^{+}98, CRYC99, CLF^{+}97b, DAF^{+}97, DCR98, DH99, DAB99, DDRC99, DDWM97, DBJDT99, EM98, ETP^{+}99, FK99, FHSM^{+}97, FR98, FHWV97, FMM^{+}98, FOK^{+}99, GR97, GRTB97, GvdFvD^{+}99, GPKH99, GSK^{+}99, GKS^{+}98, GB98, GTP^{+}99, HSL^{+}99, HOS^{+}97, HCAC99, HCRS98, HR99, HB99, HYE^{+}98, HHM^{+}98, HSSW99, HYN97, HDES97, HCW^{+}98, HIWL97, IKR^{+}98, IIM^{+}99, INMT97, JLM99, JLID98, JE98, JN98, KEJ97, KMM^{+}98, KBG^{+}99, KHA^{+}99, KOM^{+}98, KR99, KDO^{+}99, KcZS97, KCV^{+}98, KCG^{+}97, KLM^{+}98, KFL^{+}98, KLL^{+}97, KGE^{+}98, KSG^{+}98, KSY^{+}98]. Cell [LPM^{+}98, LMM^{+}99, LGZ^{+}99, LMS97, LHR^{+}97, LHWH97, LBW^{+}97, LDD^{+}97, LVN^{+}98, LPM^{+}99, iMOeM^{+}97, MAC^{+}97, MNS^{+}97, MZL^{+}99, MLN^{+}98, MOY^{+}98, MNM^{+}98, MPR^{+}97, MP98, MVM^{+}99, MMK^{+}97, MAE^{+}97, MSFA99, MLC99, MNG^{+}98, NG99, NIK^{+}97, NB97, NLM^{+}98, NH99, NT97, NMNL98, ONI^{+}98, OWD^{+}99, OKE^{+}99, OOB99, OFM^{+}98, PH97a, PH98, PFP97, PCQH98, PAN97, PDO^{+}99, PHA^{+}98, PTJ^{+}98, QBvD^{+}98, RM97, RTM99, RHP99, RC98, RKR^{+}99, SNT^{+}98, SLW^{+}99, SWK^{+}99, SN98, SAC^{+}98, SZBB98, SGR97, SBC99, SALdP^{+}97, SEH^{+}97a, SWS97b, SSS^{+}99b, Sim99, SZNS98, SATW98, SBS^{+}98, SSSS97, SB99c, SBR98, SEBF98, SVMB97, TH99, TSK^{+}97, TGN97, TSL^{+}98, TGGB97, TWSD98, TGW99, TCYS99, TIBL99, TRG^{+}99, TTAH^{+}99, ULMT99, UPT98, VCJA99, WSR^{+}97, WSH98, WWO^{+}98, WFS97, YGG97b, YHO^{+}98, YMD99, ZTH^{+}97, ZH97, ZRG98, ZJB^{+}97, vHGB^{+}97, GMS^{+}96]. Cell-adhesive [iMOeM^{+}97]. cell-associated [CDRB^{+}96]. Cell-binding [WSH98]. cell-cycle [LK95]. Cell-derived [TWSD98]. Cell-Free [SW99, VG99, HW98, MW95b, RMWW95, SSP95, KGE^{+}98, LHWH97, UPT98, ZJB^{+}97]. Cell-induced [HRD98]. Cell-intrinsic [GR97]. Cell-like [BMPP99]. cell-matrix [WPMW95]. cell-specific [OABD^{+}96, NMNL98]. cell-to-cell [DZC^{+}96]. Cells [ABM^{+}99, AGS^{+}97, BMC^{+}99, BHD^{+}97, CSN99, CLK99, CCMC97, CGC^{+}97a, CWM^{+}98, DCVF98, FNH99, GRB^{+}97, GCST97, HTK^{+}99, HRD98, HHRdB^{+}98, JGCS97, KD97a, LCD^{+}98, LKSW98, MMAK97, MAS^{+}97, NCW^{+}99, NNSN97, RWOA97, SvDtK^{+}98, SGTH99, SKC97, SSL^{+}95, TSC^{+}98, WLD^{+}98, YLB^{+}97, ZLPL99, vIZKH97, vIH98, AAP^{+}95, AGB^{+}96, AS96, BBZ^{+}95, BEG^{+}95, BZB^{+}96, BBOB96, BGR^{+}95, BFS^{+}95, BDE^{+}95, CG95, CQB^{+}96, CGVS^{+}96, CMM96, CGM^{+}95, DBvdBS95, DvD95, DC96a, DNC^{+}95, DH96, DHT96, DPH95, DSC^{+}95, DSE^{+}95, EELF^{+}95, FB96, FPM^{+}95, FFZ^{+}95, FGA^{+}96, GJ95a, GHP^{+}96, GM95, GH95, GGR^{+}96, GSvR^{+}96, HTS^{+}95, HKC^{+}95, HE96, HMB95a, HM96, HH95a, HMB95b, ICW^{+}96, JCWP95, JGM96, KSP95, KE98, KCBR96, KLW^{+}96, KCN^{+}96, KvEdV^{+}96, KGW^{+}95, KSVZ95, KGHMT95, KKE^{+}95, LCK95, LR96, LMMO96, LLL^{+}96, LDCB95, LLRTP95, MBS96a]. cells [MLW^{+}95, MTL^{+}96, MBRN95, MSF^{+}95, MBF^{+}95a, MVvdBD96, MKXY95, MXSRB96, MGA^{+}96, NDBR95, NYN^{+}95, NSC96, NCS95, NMJL96, OPD^{+}96, OSK^{+}96, OMO95, OK96, PGS95, PWE95, PLR^{+}95, PHB96, RKD96, RME^{+}95, RFNF96, RWO95, RROA95, REdC95, SRR96a, SGS95, SKA^{+}96, SLS^{+}96, SSM^{+}95, SMB^{+}95, SHS96, SM95a, SJ95, SMA96, SCM^{+}96, SL95, TMS^{+}95, TT96, TALM96, VTG^{+}96, VGM96, VPH95, VOP^{+}95, WXS96, WPS^{+}96, WW96b, WNB^{+}95, YKRS96, ZN95b, ZSMV95, ZL95c, vGvM95, vdBCH^{+}95, vdBLCvM96, vdLST^{+}96, ANS^{+}97, ALK^{+}99, ALVMM98, ANRT99, BBW^{+}98, BM97, BMG^{+}99a, BD97, BPS99, BRS99, BHC99, BKM^{+}98, BBJ98, BCWA97, BKJL^{+}98, BP98, CHC97, CJH^{+}98, CWT^{+}98, CS98, CF99, CSEM99, CEZA97, CCB^{+}98, CJY^{+}97, CGC^{+}97b, DSG^{+}97, DPZ^{+}97, DMY^{+}97, DSM^{+}97, DDM98, DBLD^{+}99, DOP^{+}98, EAM^{+}97, ESM^{+}97, EKF^{+}97, FSC^{+}97]. Cells [FHK97, FGA^{+}98, GMGGD98, GR97, GCAHW97, GMS^{+}98, GFM^{+}98, GvdFvD^{+}99, GBW98, GCM98, GM99, GLMG98, GCE^{+}97, GBFL^{+}98, HSLB99, HTK^{+}98, HKJ^{+}99, HGI^{+}99, HTY^{+}99, HLG^{+}98, HIW^{+}98, HME^{+}98, HHB^{+}97, IFSV^{+}98, IMK99, JP98, JR98, wJTSSL98, JL98, KFM99, KLH98, KRCP99, KCV^{+}98, KFL^{+}98, KA98, KHHS97, KLL^{+}97, KKLA97, LHR^{+}97, LNRR98, LMC^{+}98, LE97, LHS97a, LGM^{+}97, LCW^{+}98, MSD^{+}98, MCS^{+}97a, MSH^{+}98, MKC99, MRM^{+}98, MOY^{+}98, MLB98, MNM^{+}98, MHC^{+}97, MMK^{+}97, MGD97, MQR^{+}97, MSF^{+}99b, NRD98, NHA^{+}97, NPJW99, NWM^{+}97, OWAlW99, OSM98, OTQM98, OT97, PC98, PGP^{+}99, PHN^{+}98, PH97b, PCD^{+}98, PHM^{+}98, PFAF97, PMBM^{+}99, QWS^{+}98, RTM99, RLVM98, RPS^{+}97, RYB^{+}99, RC98, RLMC99, SGGS^{+}99, SPBCM97, SFD^{+}98, SSI^{+}97, SRV^{+}97, Sal99, SFA^{+}98, SWM^{+}98, SVF^{+}98, SHH97, SPS^{+}97a, SPR^{+}98, SHP^{+}97, SN97, SW98, SM97, SG99b]. Cells [TKY^{+}97, TSK^{+}97, TWSD98, TCYS99, TFYH99, VGVF97, VTHAA99, VZNR98, VKIH98, WZF97, WUUI^{+}98, WSS97, WPW^{+}97, WBS^{+}98, WJM^{+}99, WWD^{+}97, WC97a, WSWR99, XZMD97, qXLCH97, YHK^{+}97, YMPG98, YPDM98, YSK^{+}97, YKO97, ZSP^{+}98, ZH97, ZOS^{+}97, dBCKS98]. Cellubrevin [CGM^{+}95, ABK^{+}97]. Cellular [ALM^{+}97a, BJR^{+}99, DCVF98, LNG^{+}98, SLPE^{+}98, SJZW97, WDV^{+}97, KAP95, LLZ^{+}96, LKD^{+}95, MF96b, RAH^{+}95, SS95a, WXS96, WPZ^{+}96, WKRB96, dPSMNSM95, HSW^{+}98, JWK98, LBB97, MPCW97, NTAN98, PBM^{+}97, SPS97b, SB99b, dPCM^{+}97]. Cellularization [BDS97]. CENP [YAC97, CSY98, HBS^{+}97, LWM^{+}95, LNY^{+}95, SCM^{+}97, SVS97, ZCE^{+}98, CJS^{+}99]. CENP-E [SCM^{+}97, CJS^{+}99]. CENP-F [LWM^{+}95]. center [BIR96, MJMS96, WSW95]. Centractin [HTKH96]. Central [PJE^{+}99, SL96, WS96, WRW^{+}95, BPS99, MS99, PURB98]. Centrin [SBR98]. Centriolar [KSYK^{+}99]. Centriole [BKM^{+}98, LG95]. Centromere [HFE^{+}98, SWS^{+}96, SWL^{+}97, RHF^{+}96, SKK95, WGG96, WM95, AKLME98, BHJ98, EAM^{+}97, HBS^{+}97, wJTSSL98, MAEE98, MHWW98, MMR99b, YAC97, ZKS^{+}99]. centromere-associated [WM95]. Centromere/Kinetochore [SWL^{+}97, RHF^{+}96, WGG96]. centromeres [WM96, WM95, MPT^{+}98, YAC97]. centromeric [MC95, SVS97]. centrosomal [OWA95, FMM^{+}98, NMH^{+}98]. centrosome [BBZ^{+}95, GBW^{+}95, BKM^{+}98, DZS^{+}98, GPKH99, HCRS98, KR99, MTLGC99, MGIZ98, MZAO98, VWO99, WFL^{+}99, hYGYF97]. Centrosomes [VSRP97, DKD^{+}96, LN96, MBF^{+}95b, ZN95b, GDC97, HTH^{+}97, QGE^{+}99, RWS^{+}99, dS98a]. Centrosomin [LXC^{+}98]. CEP3 [SKK95]. CEPU [MSS^{+}99]. CEPU-1 [MSS^{+}99]. Ceramide [FNH99, MM97a, FCMN97]. cerebellar [Kun95, MMK^{+}97]. cerebellum [NA96]. cerebral [HWPC96]. cerevisiae [CRGW98, DK96, DPH95, HCG^{+}98, NWPW95, SKE^{+}95, AK99, BLC95, BML^{+}98, BGR98, CP95, CSCM97, CCBK96, CS97b, CH97, CGMH99, DRAT95, DAF^{+}97, DEPR97, EKK^{+}96, EB95, FGP96, FY99, FSHD96, GLF^{+}96, HCA^{+}95, HCW96, HB99, HRA^{+}98, HKKH99, IFHR^{+}96, JBRA98, JCC^{+}98, KKG95, hKsKC99, LKE^{+}99, LMB^{+}95, LSKB95, LS95b, MCM^{+}99, MR96, MMMW96, OAB^{+}98, PBT^{+}95, PH97a, PH98, PCYS95, RRK97, RSB^{+}99, SHLD97, SFD^{+}96, SCR98, SRG^{+}96, SCK^{+}95, SS95b, SGGS96, SD98, SKK95, SVT96, THP96, TN95, TIBL99, TDR99, VS98, WW96a, WMO^{+}95, XW96, YGK96b, YAD97, YSC^{+}95, ZL95b]. certain [YtDH^{+}95]. cervical [ZSMV95, PMKR97]. cessation [GDC95]. cGMP [DRR96, LMS97]. Chain [BGH^{+}97, KBS^{+}99, KOA^{+}99, MZN97, NCW^{+}99, SBS^{+}99, VLA^{+}98, AM95a, BKEC96, DNB^{+}96, DRR96, GW95, KSG^{+}96, KAT^{+}95, KPKWW95, KFE96, LW96, MRvD^{+}95, NSYK^{+}95, ROG^{+}96, VKR^{+}96, WLR96, WKK^{+}95, ASES^{+}97, BRLM98, BRAM97, DSG98, FKWP97, HS97, HCW99, HOCK98, HR99, JK97, KSY97, KAFB99, MTLGC99, MZL^{+}99, MSEP98, MCS98, MKW^{+}99, PKBHK97, PWW98, PTVD99, RKRG99, RYB^{+}99, SSZ^{+}97, SCEB99, SD98, SDDV97, TWP98, YWMH99, NMHS95]. chain-deficient [VKR^{+}96]. Chain-dependent [VLA^{+}98]. chain-specific [DRR96]. chains [HSL^{+}95, VGM96, VV95, WSW95, GDB^{+}98, MPL^{+}97, SLAS^{+}98, SSRP99, SCB^{+}97]. Change [BGH^{+}97, KBS^{+}99, KJBM^{+}96, THT^{+}95, DOSN^{+}99, HMD97, HPS98]. changed [KZKS96]. Changes [BEG^{+}95, HTK^{+}99, JE98, PCQH98, SLPE^{+}98, VB98, DRR96, FGA^{+}96, HAS^{+}96, HCAW95, PLM^{+}96, SR95, GDM^{+}99, IFSV^{+}98, JM97b, LHR^{+}97, MPCW97, PBCS98, SLW^{+}99, SWK^{+}99, TIBL99, TSH^{+}98]. Channel [KNH^{+}97, GCM^{+}95, LHK^{+}96, LW96, LBL95, SLL96, SMRT96, TT96, LJK97, SSS99a, SG97]. channels [AZF96, EELF^{+}95, FHCL95, PBB^{+}96, SWR^{+}96, LWSL^{+}98, PBCS98, WWM^{+}98, WS98, ZLM^{+}98, ZKN99]. chaperone [SNL96, SFNRH95]. chaperones [HD96, KA95, BCAS98, TH97]. Chaperonin [LTVC96]. Chaperonin-mediated [LTVC96]. Characteristics [KFL^{+}98]. Characterization [BO99, CSY98, FHCL95, GMBSS99, GKR95, GGK97, HCPW^{+}99, JMWG98, KLH98, KOA^{+}99, KKE^{+}95, LSZ^{+}99b, MS99, NRW^{+}95, NHB^{+}95, OWM^{+}99, OABD^{+}96, OS96b, SPCR97, TCMB98, AGB^{+}96, BMLM96, BPT96, CAA95, EPVV96, FAO^{+}96, HMSK95, HGG96, HGP^{+}95, KKG95, KGMF95, MML96, MBF^{+}95b, OP96b, dCBR95, SOL^{+}95, WSSM95, WM95, ZLC^{+}96, GYRWR97, HME^{+}98, KIH^{+}97, MGM98, RHA^{+}97, TRG^{+}99, WSH98, WFS97, qXLCH97]. characterized [FHV^{+}96, SK95a]. Charged [YHD^{+}98]. Charges [MS98]. Checkpoint [BBL^{+}99a, CJS^{+}99, KWD^{+}98, LARH97, HM95, LR95, NWG95, RCKS95, WW96a, WM96, YGK96a, ALM97b, BHJ98, CSMM98, FFLP99, KGE^{+}98, MSL98b, RC98, TGN97, WZF97, YYSB97, YMD99]. Chediak [FRT^{+}98]. Chemical [ANK97]. chemically [BLB^{+}96]. chemistry [NWG95, MSM99]. Chemoattractant [DRR96, LD95, CQB^{+}96, KGHMT95, VBG96, EHC97, XZMD97]. Chemoattractant-induced [EHC97]. Chemoattractant-mediated [DRR96]. chemoattractants [WVVD95, LCB^{+}99]. Chemoattraction [TWSD98]. Chemokine [CBM^{+}98, CQB^{+}96, ASP99, SOPM^{+}97]. Chemokines [dPSMNSM95, WAMS96, dPCM^{+}97]. Chemosensory [SWO^{+}99]. chemotactic [BGR^{+}95, CQB^{+}96, KAAFZ95, MHK^{+}95, FKB99, OSM98, RTM99]. chemotaxis [BGR^{+}95, WAMS96, AZRJ98, CF99, FCB97, PBJ^{+}98, RFVCM^{+}99, WLF99, XZMD97]. Chemotherapeutic [ZRHO99]. chick [BKEC96, BMLM96, DGSR95, GYS^{+}95, GF95, KSS^{+}95, KLMR95]. Chicken [SVB^{+}97, CLL95, SK96, AKCC99]. Chimera [GMBSS99, CRLM96, LDF^{+}95a]. chimeras [OMO95, MSFA99]. Chimeric [MM99b, CLT^{+}95, LNR^{+}96, WLML95, GMS^{+}98, TDR^{+}98]. Chinese [BBZ^{+}95, FNH99, HP95, ZOS^{+}97]. chitin [CS96a, DAF^{+}97, SS97, TDR99]. Chk1 [KGE^{+}98]. Chlamydia [SFH96]. Chlamydomonas [KMHK97, BPKB^{+}99, BM98, CDH^{+}98, GJ95b, KPKWW95, KD97b, KPW^{+}97, KBR95, KGS98, MS99, MKW^{+}99, PKBHK97, PSW95, PMH96, ROG^{+}96, SL96, WKK^{+}95, WFS97, ZR98]. Chloride [MBD^{+}96, JMD^{+}98, LWSL^{+}98]. chloroplast [MKLS96, RRRA95, ST96, HCM^{+}97, KCFS98, RB98, RKB97, YCR^{+}98, ZR98]. Chloroplast-encoded [RKB97]. Chloroplastic [ANK97, LHS97b, MS98]. Chloroplasts [RL99, KS97, WRC^{+}99]. Chloroquine [WSG^{+}98, BJR^{+}99]. CHO [DSG^{+}97, GMS^{+}98, LE97, RWO95, WBS^{+}98, YKRS96]. CHO1 [KRK^{+}98]. CHO1/MKLP1 [KRK^{+}98]. cholecystokinin [RRP^{+}95]. cholera [LCM^{+}95, MBS96a, MSW^{+}98, OF98, WJWM^{+}98]. Cholesterol [KS98, TSS^{+}95a, KSL^{+}95, GZC^{+}99, SHB99, VPM^{+}98]. Choline [RPY^{+}99]. Chondroadherin [CHLÅ97]. Chondrocyte [ANT^{+}97, SJF^{+}97, TTAH^{+}99, YSEI^{+}99]. chondrocytes [EMB^{+}96, LGdC95, REA95, TIJ^{+}95, EIIM^{+}98, KNSP97, MSWL^{+}99, NMNL98]. Chondrogenesis [CBS^{+}97]. Chondrogenic [SSA^{+}96, PNL^{+}95, FSA^{+}99]. Chondroitin [WC97b]. choreographer [GOT96]. Choreography [LSB98]. Chromaffin [ABM^{+}99, CGVS^{+}96, PAN97, SN97]. chromatid [GACM^{+}95]. Chromatids [MPT^{+}98]. Chromatin [ACF^{+}97, BM97, BHDW95, FPRL97, RMY97, CCP96, DFGL96, MC95, RSL^{+}96, TGG95, TAK95, YvdEH^{+}95, CLT99, MRM^{+}98, MBB^{+}97, RBRB99, SVS97, TSB99, YLD97, dBCKS98]. Chromatin-associated [MBB^{+}97]. Chromogranin [KGHG98]. Chromokinesin [WA95]. Chromosomal [EAM^{+}97, HSB98, WHB^{+}99, CH98, HSLB99, MSA98, MAEE98, MPL^{+}97, PCQH98]. Chromosome [BBL^{+}99a, FPRL97, FMD^{+}98, KCM^{+}97, LKWG96, SO97, SO98, ASH95, BLG95, CMWL96, EPVV96, EDB^{+}96, KCBR96, KR96, KLN^{+}96, RCKS95, SKE^{+}95, SWS^{+}96, WGG96, ABAAS98, ALM97b, CLL^{+}98, DMMS98, DWM^{+}98, GCD97, HBS^{+}97, JP98, hKsKC99, MSB99, MKC99, MHWW98, MHC^{+}97, MCPB99, MBB^{+}97, MMR99b, RC98, SAR^{+}99, SCM^{+}97, TMPM99, VvdKMvD99, YHC^{+}97]. Chromosomes [ALM97b, SGD98, ZN95a, EDB^{+}96, GACM^{+}95, LNY^{+}95, UPAS95, ZN95b, AN99, CBB^{+}99, EJ99, FSC^{+}97, HSB98, HMCL97, HD99, MCBE98, MMR99b, SR97]. Chronic [KMOO95]. Chs2p [CS96a]. Chs3p [CS96a]. Chs5p [SS97]. Chs7p [TDR99]. chymase [FGA^{+}96]. Cik1p [MBW^{+}99]. Cilia [SWO^{+}99]. Ciliary [CDH^{+}98, MS97b]. ciliates [FC95, HGP^{+}95]. Cingulin [CDH^{+}99]. Cip1 [YN95]. circadian [GJ95b]. cis [DDF^{+}98, IRMF^{+}99, LNH^{+}98, LSZ^{+}99b, MM97a, MC95, NRW^{+}95, REdC95, TKM^{+}96, KBC99, UvMJ^{+}99]. cis- [DDF^{+}98]. cis-acting [MC95, REdC95]. cis-Golgi [IRMF^{+}99, LSZ^{+}99b, NRW^{+}95, ISH95]. cis-regulatory [TKM^{+}96]. cisterna [MR95]. Cisternae [SW99, AMJM95, HW98, MWL97]. city [EGA97]. clams [TMM^{+}95]. claret [MMGT96]. Class [SKC97, AM95a, GMS^{+}96, GBN95, MRvD^{+}95, RvSL^{+}95, SOG96, THP96, TYSP^{+}96, WLR96, WFBO^{+}96, BRAM97, DBLD^{+}99, FRT^{+}98, FKWP97, GDB^{+}97, KMG^{+}97, LM98, LMB98, SEH^{+}97a, SSRP99, SWO^{+}99]. classes [BLC95, ZPS96]. classic [JWH96, LBW^{+}97]. classical [CQB^{+}96]. Clathrin [DBvdBS95, EGKC^{+}99, GK99, HUU96, SK96, SOG96, TMDD96, TBR^{+}96, WB96, FMP^{+}97, FGA^{+}98, KRGA98, KKG^{+}98, LH97, LMB98, SSEM98, SFR^{+}98, vDSH^{+}97]. Clathrin- [FGA^{+}98]. Clathrin-Coated [EGKC^{+}99, GK99, SK96, SOG96, FMP^{+}97, LH97, SSEM98]. Clathrin-dependent [SFR^{+}98]. Clathrin-independent [DBvdBS95, SFR^{+}98]. Claudin [FFH^{+}98, FSFT98, FST99, MSF^{+}99b, MSFT99]. Claudin-1 [FFH^{+}98, FSFT98]. Claudin-11 [MSF^{+}99b]. Claudin-11/ [MSF^{+}99b]. Claudins [IFM^{+}99, SFS^{+}99]. Clb1 [KM95]. Clb5 [SCR98]. Clb5-associated [SCR98]. Clear [FSH^{+}99]. Cleavage [BRM^{+}99, CGC^{+}97a, BHP^{+}95, CDRB^{+}96, LWB^{+}95, LF95, MAM95, MKI^{+}96, ZLC^{+}96, BLRS97, CLRR99, CSO97, DMY^{+}97, EAM^{+}97, GDTBD98, SDC^{+}97]. cleavage-stage [MKI^{+}96]. CLIP [DPG^{+}99, DWM^{+}98, FNDS99]. CLIP-170 [DPG^{+}99, DWM^{+}98, FNDS99]. clock [GJ95b]. clonal [BLB^{+}96, SSA^{+}96, DDM98, PGP^{+}99]. Cloning [HCPW^{+}99, WKL^{+}96, WC97b, AASH^{+}96, CLL95, HGP^{+}95, KLMR95, GYRWR97, KIH^{+}97, MPL^{+}97, SGJP99]. close [CFK95]. Closely [HFP98a]. Clostridium [SFS^{+}99, KIH^{+}97]. Closure [EGA97]. cluster [NNMW96, BMS^{+}97, DP98, YHD^{+}98]. Clustered [SMA96]. Clustering [KNH^{+}97, FDH96, GDC95, GLF^{+}96, HCA^{+}95, KZKS96, SMRT96, SRH^{+}96, BGH^{+}98, CJY^{+}97, HPS98, wJTSSL98, KRGA98, LBWF^{+}97, PXRR99, SGYH97, WSWR99, YNG98, ZLM^{+}98]. clusters [vdBCH^{+}95, BSM^{+}99, JP98, KBB^{+}97a]. CMF [VBG96]. CNS [BQF^{+}95, CBB^{+}96, FJT^{+}95]. CO [CSC^{+}98]. CO-029 [CSC^{+}98]. CoA [LW95, STN^{+}96, TOSF^{+}98, TMM^{+}95]. coagulation [SRR96b]. Coagulopathy [CSC^{+}98]. coassemble [GGR^{+}96]. Coassembling [LBO^{+}98]. Coat [SME98, EGH^{+}95]. Coated [EGKC^{+}99, GK99, KBW^{+}96, LS95a, LNR^{+}96, NDBR95, PR95, SK96, SOG96, KsYA98, FMP^{+}97, FGA^{+}98, LH97, SSEM98, SOR^{+}96, SWC^{+}96]. Coatomer [PAL^{+}98, FAO^{+}96, SOR^{+}96, DSG^{+}97, DDF^{+}98, SPL^{+}97]. coats [ABRB95a, TBR^{+}96]. Cochaperone [SHMC99]. coded [LHS97b]. Coexistence [MAS^{+}97]. Coexpress [DBC^{+}99]. coexpressed [CCR96]. coexpression [FHV^{+}96]. cofactor [LO96]. Cofactors [TLF^{+}97]. Cofilin [CLS^{+}97, MPCW97, OBB99, RAA^{+}97, SMTN99, SB99d, The97, AOMB96, ASY96, RTL^{+}99, GBW^{+}95]. cofilin/ADF [GBW^{+}95]. Cognate [KBC99]. Cohesion [MPT^{+}98]. Coil [WC97b, FT95, HW96b, CDH^{+}99, FMM^{+}98, QPH^{+}98]. Coiled [WC97b, BFL95, FT95, HW96b, ASACF98, CDH^{+}99, FMM^{+}98, IYM98, JWW^{+}99, MCBE98, QPH^{+}98]. coiled-coil [FT95, HW96b]. coilin [BFL95, ASACF98]. Coils [MCBE98]. Coincides [MC98, MRL^{+}99]. COL2A1 [CLT^{+}95, LNH^{+}98]. COL2A1-directed [CLT^{+}95]. COL4A5 [NKI^{+}95]. COL4A6 [NKI^{+}95]. cold [SFD^{+}96, NIK^{+}97]. Cold-inducible [NIK^{+}97]. Coli [BPA^{+}97, NAP^{+}96, VWG^{+}97, LCM^{+}95, VSGF96]. collaborate [MTGY96]. Collaboration [YTT99]. Collagen [ACH^{+}98, LRM^{+}95, XC97, BGGR^{+}96, BFP^{+}96, BTNZ^{+}95, EMB^{+}96, HG95a, KBM^{+}95, LGdC95, LWB^{+}95, MHLB96, MS96, NKI^{+}95, REdC95, SHY^{+}96, TKM^{+}96, CLRR99, CML^{+}98, DBH^{+}97, FKM^{+}99, GJW^{+}97, GGK97, HBH98, IRW^{+}99, KPZ^{+}97, KZS^{+}98, PDS^{+}97b, PWGG98, RRA97, RKR^{+}97, XZSC98]. Collagen-dependent [PWGG98]. collagen-producing [REdC95]. collagen-specific [SHY^{+}96]. Collagen-tailed [RRA97]. Collagenase [XC97, LRM^{+}95, LWB^{+}95, TDW95, PDS^{+}97b]. Collagenase-1 [PDS^{+}97b]. collapse [IKT^{+}96]. Collapsin [MSF^{+}99a]. Collapsin-1 [MSF^{+}99a]. Collapsin-1/Semaphorin [MSF^{+}99a]. Colocalization [BPA^{+}97, GLF^{+}96, vDSH^{+}97]. colocalize [BRJP96, GMC^{+}95]. Colocalized [HPKL97, WS98]. colocalizes [JMTCF96, IHT^{+}98]. colon [GSvR^{+}96, ANRT99, MAC^{+}97, MDC^{+}99, vHGB^{+}97]. Colonic [MDC^{+}99]. Colony [MHM99, SRR96a, PHN^{+}98, WDL^{+}98]. Colony-stimulating [PHN^{+}98, WDL^{+}98]. Columnarization [VTHAA99]. Combination [BS97]. combinations [JWH96]. Combinatorial [FCB97]. Combined [KGGK99]. Comet [SWS97a]. Coming [Col99]. Commitment [LMA^{+}97]. common [LKD^{+}95, PWK^{+}96, Val96, JMM^{+}98, KLS^{+}99]. communication [SL95, GCE^{+}97, HCW^{+}98, LNG^{+}98, NB97, PHA^{+}98]. Compaction [ACSN98, BHDW95]. Comparative [FLA99]. Comparison [CJY^{+}97, HST^{+}98, SFA^{+}98, ZCOP99]. Compartment [BPWS98, CGC^{+}97a, VS98, ZWT^{+}97, GMS^{+}96, JGD^{+}96, LSKB95, MLJM95, MFMW95, PCYS95, PPBJ96, RvSL^{+}95, SNL96, VYB95, CTB^{+}97, GMS^{+}98, HW98, HDES97, HDES98, IMS^{+}98, JMS97, KHHS97, MKB^{+}99, PFHWN98, SHH97, WWD^{+}97]. compartment-specific [SNL96]. compartmentalization [KSP95, KPZ^{+}97]. Compartmentalized [SM97, KLS^{+}99]. compartmented [CLS96, SC97]. Compartments [ALC^{+}99, NAsG^{+}98, WBKB97, vIH98, CS96b, DMM^{+}96, MRvD^{+}95, PHO^{+}95, dPF95, WLR96, BL99, FKWP97, KMG^{+}97, PvDA^{+}98, RKB^{+}98, YDFB98]. compenents [SGBD96]. competence [RRRA95, CMR^{+}97]. competent [vdBCH^{+}95, KNSP97]. Competes [NRD98]. Competing [SNK^{+}99]. Complementary [HPS98]. complementation [ARM^{+}95, BLC95]. complete [BL96]. completed [HER95]. completion [AJS96, FT95, LS95b, ATH^{+}97]. Complex [BMC^{+}99, BdPE^{+}97, CSN99, ELL^{+}99, GHS98, KWD^{+}98, LAW^{+}98, LSZ^{+}99b, MSP97, PT97, RMY97, SBG^{+}98, SME98, SKC97, SB99d, WDV^{+}97, YTY^{+}99, AM95a, BLEB96, BS95b, CPL^{+}95, CSY^{+}95, DWS^{+}95, DHT96, EPVV96, FLW^{+}95, FAO^{+}96, FaAR^{+}96, GMC^{+}95, HCA^{+}95, HGG96, KLW^{+}96, KSJW95, KSR96, LvSC^{+}96, MCB96, MGS^{+}95, MDS95, MMF95, MGY^{+}95, NIM95, NRM^{+}96, NW96, OS96b, PG95, RHHRB96, RvSL^{+}95, RDM^{+}96, SLSW96, SMB^{+}95, SCK^{+}95, SBW^{+}96, SSS^{+}95, SSF^{+}95, SNCH95, SR95, TET^{+}96, TN95, VIF^{+}96, WMR^{+}96, ATH^{+}99, BBP^{+}98, BMB97, BHC99, BRAM97, Bri99, BENV97, BKJL^{+}98, mCBLK98, CBTZ98, CRRF97, DZS^{+}98, DTMG99, DFMG97, FHAP98, FKWP97, GZC^{+}99, HOCK98, HN98, HFP^{+}98b, KFM99, KRH98, KSV^{+}99, KDK^{+}99, KHS^{+}97, KD97b, KMG^{+}97, KIT^{+}97, KPW^{+}97, LMB98, MAW98, MGIZ98, MWB98, MZAO98, MUS98, NA99, OTS98]. Complex [ODB98, PKBHK97, PAF97, PFKvD99, RN99, RGS99, SSW^{+}98, SPCR97, SHL97, dCBR99, SAEV^{+}98, UNPW98, UvMJ^{+}99, VCS^{+}99, VZNR98, WUUI^{+}98, WKR^{+}98, YPB^{+}98, ZDK^{+}99]. complexed [LRG^{+}98]. Complexes [MEFH97, MYTK99, YMAC^{+}98, AM95a, BRB96, FPHH96, GH95, GDC95, HH95b, PR95, SHT^{+}95, VBZ96, WSGPS95, ANK97, BO99, BHBB99, BW97, CGM^{+}99, CRGW98, CT98, HCW99, KBB^{+}97a, KSG^{+}98, LBWF^{+}97, MLB98, OWM^{+}99, RHPW97, SB99c, TMPM99]. Component [EGKC^{+}99, MSMC99, WH97, DK96, FSHD96, GHJJ95, HM95, OSH^{+}96, SCG^{+}95, STVR95, WRW^{+}95, WGG96, BSRG99, BSY97, CDN97, CYH98, CRGW98, CRRF97, DSG^{+}97, DS98b, DTMG99, EZC^{+}97, HRA^{+}98, JMM^{+}98, KIT^{+}97, LJK97, MCTMK98, MTW^{+}99, PFDL97, RN99, RHPW97, YHK^{+}97, YAC97]. Components [RMY97, TDW95, BGB95b, KSL96, MKLS96, MWOB96, SVB96, TN95, WvdVV^{+}95, AK99, HSVS98, KCFS98, KPZ^{+}97, MC97, MJBR98, MUS98]. Composed [GCST97, WDV^{+}97, BTTB96, TCSG96]. Composes [KSY97]. Composition [SBS^{+}99, DLB96, XML^{+}96]. Compromises [PHM^{+}98, SEAB^{+}99, YSK^{+}97]. Con [SRRB95]. Concentrated [BDO^{+}97, GYS^{+}95]. Concentrates [CGM^{+}99]. concentration [NYN^{+}95, AvdG99, mLZB99]. concerted [DPHW96]. Condensation [LSB98, CMHT96, CLT99, CLL^{+}98, SAR^{+}99]. Condensation/Decondensation [LSB98]. Condensed [MPT^{+}98]. conditional [GDC95, NPRT95]. Conditions [FFDP98, FAP^{+}96, RME^{+}95, SWC^{+}96]. conductance [LBL95, JMD^{+}98, LJK97]. Conducting [PBCS98]. cone [IKT^{+}96, SSW^{+}96, THK95, TK95, MQR^{+}97, PKQ^{+}98, SEBF98, ZS98]. cones [AC95a, SWD^{+}98, ZSB99]. confer [KAT^{+}95]. conferred [HLSM95]. confers [RS95, SNL96, SCG^{+}96]. Conflicting [FKB99]. conformation [BHDW95, TFF^{+}95, CLF^{+}97a, KRGA98]. conformational [LBCB96, DOSN^{+}99, GDM^{+}99, HMD97, HPS98, JM97b, TLF^{+}97]. conformations [ES96, RDS^{+}99]. conforms [TT96]. congenital [KKA96, LKD^{+}95]. Congress [KCM^{+}97]. Congression [MAEE98, MHC^{+}97]. conjugate [MKB^{+}95]. Conjugated [SSRP99]. Conjugating [KSF^{+}99, HBPJ98]. conjugation [EM96, PNEH98]. Connecting [dCBR99]. connection [BCSG96, HCAC99]. connective [HKS^{+}96, KZS^{+}98]. Connects [DSC^{+}99]. Connexin [MSFA99, EELF^{+}95, SL95, VUC^{+}95, FZZN98, ZKN99]. connexin-mediated [VUC^{+}95]. Connexin-Occludin [MSFA99]. connexin-transfected [EELF^{+}95]. Connexin26 [GJB^{+}98]. Connexin26-deficient [GJB^{+}98]. Connexin32 [CFB^{+}98]. Connexin32-deficient [CFB^{+}98]. connexin43 [KGW^{+}95, LCK95, qXLCH97]. connexin45 [KGW^{+}95]. Connexin46 [KHHS97]. Connexin50 [WGP98]. consequences [SS95a, GSJ98, SB99b].

Conservation [KBG^{+}99, SWL^{+}97]. Conserved [CWSL97, WLF98, WDV^{+}97, ES96, KMOO95, NSR95, SHR^{+}95, SCKG95, AKLME98, AL97, BSY97, BG97, BF97b, EGA97, FBM99, GRTB97, HMC^{+}99, KQB^{+}98, MC97]. constant [DFC^{+}96]. Constitutive [AGS^{+}97, SAPLHD95, VC97, WJC^{+}96, CMM96, HKL96, MKF96, SRH^{+}95, CRRF97, FSH^{+}99, JH97, MCRB97]. constitutively [BGB^{+}95a, HIWL97, YSEI^{+}99]. constructs [RCC^{+}96]. Consumption [CSC^{+}98]. contact [DNC^{+}95, DMD^{+}96, DFC^{+}96, GZK^{+}95, NIM95, BSM^{+}99, HLK^{+}98, NB97, OOB99, SSZ^{+}97, SVS97]. contact-inhibition [DNC^{+}95]. Contactin [SLN^{+}97, ZDK^{+}99]. contacts [DSGF95, FBH^{+}95, Gol95, SRRB95, BMHH97, BLRS97, KKS99, MSFA99]. Contain [CDH^{+}98, ASSS^{+}95, BSMH95, ND96, OPD^{+}96, FHWV97, RDS^{+}99, VSRP97]. contained [DMG96]. containing [BLL^{+}96, BJS^{+}95, BKEC96, FPHH96, GDC95, JBD^{+}95, JGD^{+}96, dPF95, PSB98, SLS^{+}96, TNM^{+}99, TBR^{+}96, WKRB96, WPMW95, ZL95a, MAW98, MSFA99, OPH^{+}99, Sal99, SVB^{+}97, ZOKS99, CC98, EKK^{+}96, FNDS99, KFM99, PH98, RM97, RBB97, RKB97]. Contains [NAsG^{+}98, CLR^{+}96, GSC^{+}96, MFMW95, CDH^{+}99, MUS98, SDHM99]. content [CMHT96, MVvdBD96, PPW^{+}95, EFB^{+}98, ZLL^{+}98]. context [MLO^{+}95]. Continual [DSGF95]. Continuous [SHH97]. continuously [WW96b]. Contractile [RL96, FA95, BML^{+}98, HHB^{+}97, KSY97, SWS97b]. contractility [CWB96, ML97, ZCWB^{+}98]. Contraction [CLK99, GW95, HG95a, MCA96, SPT^{+}95, TPM^{+}99]. Contractions [PvDA^{+}98]. Contrasting [YMD99]. contribute [TKM^{+}96, DBC^{+}99, TCV^{+}98]. Contributes [CT97b, FJT^{+}95, HFSHK96, SMB^{+}95, JP98, MLN^{+}98, PTVD99]. Contributing [SPR^{+}98]. Control [AK97a, BMG^{+}99b, FJ98, FVRCH96, LARH97, LKE^{+}99, VNTM^{+}99, AGB^{+}96, BLE^{+}96, BFP^{+}96, EAL^{+}95, FWHW96, FUB^{+}96, GEJ^{+}95, RSP^{+}95, RK95, STRM95, VBZ96, WHP^{+}96, YFST^{+}96, BR99, BBJ98, BR98, CCS^{+}98, DBJDT99, EIIM^{+}98, FLO^{+}99, FSM^{+}99, FCB97, GR97, GKP^{+}97, IMK99, KGE^{+}98, LMF98, LWS99, MGB^{+}97, MBOC97, MGD97, NH99, PWD^{+}99, TDR99, WG99, WWM^{+}98]. Controlled [PNL^{+}95, ABK^{+}97, CCBL98, JMD^{+}98, KGA^{+}97, ZH97]. Controlling [MBCS98]. Controls [ALC^{+}99, MM99a, PCYS95, SS95a, STP96, BRAM97, BP98, DBLD^{+}99, JHM98, OKB^{+}98, RHLG98, SSRvdB98, VPM^{+}98]. Conventional [KMG^{+}97]. Converge [LGGS97, VMT^{+}99]. conversion [PNL^{+}95, LGM^{+}97]. convertase [DMM^{+}96]. Convertases [CR99]. converting [MMAK97, FCMN97, MCS^{+}97a, TPM^{+}99]. COOH [BL96, hCMPG97, GF97, HLSM95, HTY^{+}99, LCM^{+}95, LDF95b, MSW^{+}98, MAS^{+}95, MMD^{+}98, PL96, SYO95, TNFK^{+}98, TSS^{+}95a, WHM^{+}96, IFM^{+}99, SHPB99]. COOH-terminal [BL96, GF97, HLSM95, HTY^{+}99, LCM^{+}95, MAS^{+}95, PL96, TNFK^{+}98, TSS^{+}95a, WHM^{+}96, LDF95b, MMD^{+}98]. cooperate [BKB^{+}96]. Cooperates [XMMW97, SPS97b]. Cooperative [HHM^{+}95, KvEdV^{+}96, SVS97]. coordinate [KSP95, OWA95]. Coordinated [BWKH96, PFAF97, WB99, DSGF95, ZL95a]. Coordinates [WE98, WJM^{+}99]. Coordination [AWS96, XCLM98, HCRS98, NTAN98]. COP [AGPG96, DSG^{+}97, DDF^{+}98, FAO^{+}96, GAPG97, MW95b, SSSKR95, SWC^{+}96]. COPI [DSG^{+}97, ABRB95a, GE97, LPD^{+}99, MSW^{+}98, RAM^{+}96, SOR^{+}96, SLL^{+}98, TPMB97, SBD^{+}99]. COPI-coated [SOR^{+}96]. COPI-independent [GE97]. COPII [ABRB95a, AWB^{+}98, KSL96, RAM^{+}96]. Coracle [WLF98]. Core [WC97b, CMHT96, GACM^{+}95, NRM^{+}96, TGG95, WBG^{+}97, AK99, ATH^{+}99, KSV^{+}99, RN99, RGS99]. core- [WBG^{+}97]. Corepressor [QWS^{+}98]. Corequirement [MCJK98]. Cores [ALM97b, HD99, SR97, TMPM99]. corneal [MHLB96, CML^{+}98]. corneal-specific [MHLB96]. cornichon [PB98]. cornified [RHS^{+}96, RHPW97]. Corona [YAC97]. coronavirus [ORW^{+}95]. Coronin [GWB^{+}99]. Corralling [SNT^{+}98]. Correct [SCR98, BHJ98]. correlate [GHP^{+}96]. Correlated [ABAAS98]. Correlates [ISR99, MAM95, PTBC96, CLRR99, LPM^{+}98, RSB^{+}99]. Correlation [KD97a, ZH97]. correlations [RCC^{+}96, LRG^{+}98]. Correlative [SGA97]. Corresponding [WG97]. cortex [FK95, FA95, HWPC96, CS97a, HCAC99, HIWL97, PAN97]. Cortical [BDO^{+}97, BGB95b, LZD95, MST^{+}96, WKWC96, ZLC^{+}96, BRD99, CBTZ98, CYSD99, DSM^{+}97, GDTBD98, GDL98, LL97, Li97, MR98, MMR99a, MRL^{+}99, OWAlW99]. cotranscriptionally [VIF^{+}96]. Cotranslational [KRH98, MJBR98]. Coupled [Bar97, RRH^{+}95, TOT^{+}96, VLM^{+}98, BLRS97, DPG^{+}99, KsYA98, CMS98, MKS^{+}97, PHA^{+}98]. couplers [DM95]. Coupling [CSN99, ABRB95a, MCL^{+}96, SLP^{+}96, SPT^{+}95, VUC^{+}95, FJ98, KLM^{+}98, SEBF98]. Couplings [PFAF97]. Course [KS97]. Covalent [SJW97]. covalently [CLR^{+}96]. covers [WSW95]. CP190 [OWA95]. CPP32 [TMN97]. CPP32-like [TMN97]. CPRF2 [KWN^{+}99]. CRAC [FHCL95, LD95]. cRac1B [AGP^{+}98]. Crawling [HB98]. Cre [GBvdN^{+}98]. Cre- [GBvdN^{+}98]. Crest [HCW^{+}98, MGD97]. Crithidia [JE98]. Critical [SHB99, MW96, WPS^{+}96, YGK96a, BBJ98, ISN^{+}99, Li97, LRC^{+}99, PHA^{+}98, SEBW^{+}98]. Crk [KLM^{+}98]. CRM1 [KDG98, KDK^{+}99, PIB^{+}99]. CRM1-mediated [PIB^{+}99]. crmA [FVKS96]. CRO1 [BBOE98]. CRO1/She4p [BBOE98]. CRO1/She4p-like [BBOE98]. Cross [LWS^{+}97b, MGD97, AKP95, LMB^{+}95, NCZ^{+}95, TTG95, WSG^{+}95, ANK97, BFHB97, GWB^{+}99, LSZ^{+}99a, MSH^{+}99, NGMR97, PH97b, SMB^{+}99, TCV^{+}98, YPB^{+}98]. cross-bridge [NCZ^{+}95]. Cross-link [TCV^{+}98]. Cross-linkers [TCV^{+}98]. cross-linking [AKP95, LMB^{+}95, GWB^{+}99]. cross-links [TTG95, BFHB97, SMB^{+}99]. Cross-striated [YPB^{+}98]. Cross-Talk [LWS^{+}97b]. Crosslinking [EKS97]. Crosstalk [BSW^{+}99]. CRP1 [PLB97]. Crucial [FOK^{+}99, SBPR^{+}98]. Cruising [BG98]. cruzi [BCWA97]. cryoelectron [BHDW95, MML96, JM97b, JOPM98].

Cryofixation [PAN97]. Cryp [LHB^{+}99]. cryptic [MGS96, ZCWB^{+}98]. Crystal [SFSV95, SEB99, LRG^{+}98]. crystals [KSL^{+}95, MML96]. Ctf13 [RGS99]. Ctf19p [HKKH99]. Cues [SWD^{+}98]. CUG [VTG^{+}96]. CUG- [VTG^{+}96]. culture [KSS^{+}95, LDCB95, VKR^{+}96, dHvPL^{+}95, mCBLK98, CMR^{+}97, MBBT98, SDL98, WPW^{+}97]. cultured [GGR^{+}96, HMB95b, HCM^{+}96b, KKE^{+}95, LLL^{+}96, NYN^{+}95, PBB^{+}96, RL96, WW96b, YSK95, KB99b, MOY^{+}98, OWAlW99, PKQ^{+}98, SPBCM97, WNB97]. Cultures [CSF^{+}97, BLB^{+}96, CLS96, RNK97, SC97, SSSS97]. Current [WMP^{+}98]. Currents [ZOS^{+}97]. Cushion [NMK^{+}97]. Cutaneous [APG^{+}96]. CXCR4 [SOPM^{+}97]. cyclase [LD95, SKS^{+}95, SAM^{+}99, WLF99]. Cycle [CCMG97, LSB98, AW96, DRAT95, DSGF95, FSHD96, GMFN95, GJ95b, KJBM^{+}96, KF95, LG95, LR95, LO96, LK95, MK95a, MG96, NWG95, NAV96, NMJL96, OWA95, RHF^{+}96, RSP^{+}95, SH96, SNCBM95, TKVC95, TMDD96, TALM96, TMM^{+}95, TFF^{+}95, VMN95, YSC^{+}95, ZOB^{+}96, Ass97, BOM^{+}98, DDRC99, FK99, FMM^{+}98, HCRS98, HB99, HYN97, HHS^{+}99, JE98, JB99, KR99, KCV^{+}98, LHR^{+}97, LMP98, MCC^{+}97, MKS^{+}97, OWD^{+}99, OOB99, PBCS98, PCQH98, RC98, RKR^{+}99, SLW^{+}99, TGN97, TGGB97, TIBL99, VWO99, YGG97b, YMD99, ZRG98]. cycle- [RHF^{+}96]. cycle-associated [KJBM^{+}96]. cycle-dependent [FSHD96, KF95, MK95a, OWA95, SH96, TKVC95]. cycle-independent [MG96]. Cycles [NAsG^{+}98, BBZ^{+}95, CSY^{+}95, CS96b, FA95, SO97, SO98]. Cyclic [DVL^{+}98, TPM^{+}99, WWM^{+}98]. Cyclin [DVL^{+}98, MYTK99, NNSN97, OHB^{+}95, PMP^{+}98, SSR^{+}98, JCPK95, KSVZ95, NWPW95, ZOB^{+}96, FdEP99, JHM98, LSM^{+}98, MYTK99, TYT^{+}97, WOM98]. Cyclin-dependent [DVL^{+}98, PMP^{+}98, SSR^{+}98, KSVZ95]. cycling [DG96, BF97b, MM97a, PKCS98]. cyclins [DRAT95, KKFN^{+}95]. Cycloheximide [FEF^{+}97]. Cyclosome [CJS^{+}99, KWD^{+}98]. Cyclosome/Anaphase [KWD^{+}98]. Cyclosome/Anaphase-promoting [KWD^{+}98]. Cyclosome/APC [CJS^{+}99]. Cyk2 [LL98a]. Cylindrotheca [WFBC96]. Cyritestin [YPM97]. cysteine [AHM^{+}96, GLL^{+}99, vtHR97, OTQM98, ZOKS99]. Cysteine-3 [vtHR97]. Cysteine-rich [OTQM98, ZOKS99]. Cysteine34 [SKR96]. cysteinyl [SLS^{+}96]. cysteinyl-containing [SLS^{+}96]. Cystic [JMD^{+}98]. Cyto [SATW98]. cytoarchitecture [PFF^{+}95, vABSP95]. cytochemistry [JCWP95]. Cytochrome [SHK^{+}99, MML96, DOSN^{+}99, EAOS^{+}98, MDE^{+}99, NRP98, VCJA99, BAC^{+}96]. cytokeratin [PLM^{+}96]. Cytokine [LCN^{+}97, SSM^{+}95, HAB^{+}99]. Cytokine-independent [LCN^{+}97]. Cytokinesis [SGS95, SGD98, TXB^{+}99, CM95a, GBW^{+}95, LVD96, MBP^{+}95, RSP^{+}95, SMC^{+}96, WLML95, WW96b, WRW^{+}95, YYH^{+}96, ATH^{+}97, ALCC^{+}99, BML^{+}98, BGG98, CRM^{+}98, CDN97, CF99, EAM^{+}97, KSY97, KWS^{+}99, LWS^{+}97a, LL98b, MAEE98, NWM^{+}97, SCEB99, SB99a, SMM99, TCM^{+}97, WHG^{+}97, YAiN^{+}98]. Cytokinesis-specific [LWS^{+}97a]. Cytokinetic [JCC^{+}98, YAiN^{+}98]. Cytological [CLF^{+}97b]. Cytomatrix [BDO^{+}97]. Cytomegalovirus [ISM97]. Cytoplasm [BOS^{+}97, HMSK95, OSHG^{+}96, SK95b, SMA96, CL99, KSOK97, PIB^{+}99, SBV97, WCW98]. Cytoplasmic [AEVB98, BdPE^{+}97, CWSL97, FSDA97, GPKH99, HZR98, HH95a, KSE^{+}97, KBB^{+}97b, NB97, RWS^{+}99, SNT^{+}98, TMN97, VV95, All95, BLLB95, CAA95, DSM^{+}96, FBD^{+}95, GBG^{+}96, GKM^{+}96, HLSM95, LNR^{+}96, MRvD^{+}95, MGS^{+}95, NNMW96, OABD^{+}96, PWG^{+}95, RSJK95, RSRK96, SLT^{+}96, SSN96, SPMB^{+}96, SMRT96, SOHP95, SLEB95, SKR96, SOE^{+}95, STRM95, TFF^{+}95, VGM96, WGR^{+}95, WKRB96, WF96, ZKSB96, vABSP95, BALL98, BSS^{+}97, BRLM98, BKN^{+}97, CS98, EPMB^{+}99, GCML98, HTK^{+}98, HOCK98, HMB^{+}98, HFP^{+}98b, IRW^{+}99, KLH98, KRCP99, LM98, LG99, MR98, OK98, PDW99, PTVD99, SZFM98, SHLD97, SALdP^{+}97, SSEM98, SMZ^{+}98, YWMH99, YNG98, YHD^{+}98, YPDM98]. Cytoskeletal [AAP^{+}95, LTW^{+}97, LKSW98, RK95, SL97, SMH98, BGB95b, CK96, CH95, GKC^{+}96, HFSHK96, HGH96, LCS96, MTC^{+}95, PWG^{+}95, vdLST^{+}96, DOP^{+}98, EGA97, GKB^{+}98, GB98, HPKL97, HCW99, HDDKH98, KSK^{+}99b, MH97, MSH^{+}98, SV98, SMTN99, SEBF98, TBP^{+}99, WGF^{+}98]. Cytoskeletal-associated [SV98]. Cytoskeletally [MDK^{+}99]. Cytoskeleton [ISR99, MM99a, MB99, RdPRW98, The97, ZC99, CGB^{+}95, CGSP95, DTB96, FUB^{+}96, GAW^{+}96, GPRS^{+}95, LZD95, MCL^{+}96, MMMW96, SVB96, Wal95, YWW^{+}96, BRP^{+}97, BVKP98, BGN^{+}97, CF99, CWY99, CYSD99, GDL98, HKBM99, HKS97, JL98, KMMC98, KHA^{+}99, Li97, MSL98b, OC98, PBJ^{+}98, SRV^{+}97, SMS^{+}98, ÜHK^{+}99, VMH97, VTHAA99, YAD97]. Cytoskeletons [CCC^{+}99, GWB^{+}99]. cytosol [HP95, MCB96, MB96, PGS95, HC97, KEP^{+}99, SPBCM97, TSSR98, TSH^{+}98, WHS^{+}97]. Cytosolic [GMBSS99, CM95a, PG95, RLCM96, Bar97, BM99, BCWA97, DSM^{+}98a, GSKL99, HRK^{+}98, MNR^{+}98, SDDV97, SEB99]. Cytotoxicity [DCVF98]. cytotoxin [LHM^{+}96, CHH97].

D [sHWM^{+}99, NGT99, WPS^{+}96, BZB^{+}96, CSA^{+}97, Dav95, HG95a, KBW^{+}96, PFHWN98, RME98, VSGF96, WBR97]. D1 [MML96, ZOB^{+}96]. D2 [DSE^{+}95]. D3 [BAS^{+}97, DVL^{+}98]. Daip1 [KWS^{+}99]. dalton [RFNF96]. Dam1p [HCG^{+}98]. Damage [MWGE97, GDM^{+}99, JHM98, RC98]. Damage-induced [GDM^{+}99]. Damaged [MTW^{+}99]. damaging [PMP^{+}98]. Dap [CIK^{+}99]. Dap-Kinase [CIK^{+}99]. Darby [ALK^{+}99, BRS99, DMY^{+}97, OT97, PMBM^{+}99, ZZvdB^{+}96, vGvM95]. Daughter [SGTH99]. Daxx [ISN^{+}99]. DBP [SS95a]. DdCAD [SWS97b]. DdCAD-1 [SWS97b]. DdPIK1 [BGN^{+}97]. DdPIK2 [BGN^{+}97]. De-adhesion [GSvR^{+}96]. Death [ABP^{+}98, MWGE97, MHM99, SMZ^{+}98, SKC97, FMJ^{+}95, HG95b, JWR96, LLRTP95, MFM^{+}95, RME^{+}95, SM95a, WJC^{+}96, BMP^{+}98, BNJ99, CIK^{+}99, KOM^{+}98, LNM^{+}98, LMS97, LEB^{+}97, MMK^{+}97, NMNL98, PMP^{+}98, RHP99, TSL^{+}98, TNCM97, VCJA99, YHO^{+}98]. Death-effector [SMZ^{+}98]. Decatenation [ALM97b]. Deciphering [LBB97, SK96]. Decondensation [LSB98]. Deconvolutes [GCE^{+}97]. Decorated [HST^{+}98, JOPM98]. Decorin [DBH^{+}97, HBH98]. Decreased [GJB^{+}98]. Decreases [EFB^{+}98]. Decreasing [SGGW99]. Dedifferentiated [SW98]. Deep [FHWV97]. default [DPH95]. Defect [YHO^{+}98, GDF^{+}96, SSY^{+}96, DSG^{+}97]. Defective [RKRG99, HP95, LFB96, PMMF95, STN^{+}96, LBWF^{+}97, NWM^{+}97, SJF^{+}97, jLC97]. defects [FHCL95, GBW^{+}95, MS96, NPRT95, NMJL96, RFNF96, ASES^{+}97, BRS98, BGN^{+}97, GMS97, HLJ^{+}98, KAFB99, LEB^{+}97, PSH^{+}98, SZBB98]. Defense [DCJ97]. Deficiency [LFB96, THT^{+}95, HLJ^{+}98]. Deficient [CGSY^{+}99, FRT^{+}98, BBJ98, MKB^{+}95, SFD^{+}98, VWK^{+}95, VKR^{+}96, EFML99, GvdFvD^{+}99, CFB^{+}98, CLF^{+}97b, DPT^{+}97, DSM^{+}98b, GJB^{+}98, GMS97, RFCR^{+}97, YHO^{+}98]. define [GLS96]. defined [BLB^{+}96, OGS^{+}96, BL99, FCMN97, SSS^{+}99b]. defines [BM96, HKL96, NSR95, SJSJ95, THP96, MPR^{+}97, Nei98]. Defining [SF98, LYC^{+}95, PTBC96]. Definition [ISM97, MC98]. DEG [AAM^{+}98]. DEG/ENaC [AAM^{+}98]. degeneracy [MLO^{+}95]. degeneration [MWT^{+}96, LYD^{+}97, MAE^{+}97]. Degradation [JBRA98, LW95, AHBW96, DRAT95, HD96, KRMMA96, MB96, PBB^{+}96, ATD97, ASB^{+}99, BM97, BSH99, CC98, FKWP97, FJ98, GLL^{+}99, HC97, MGP97, MDV98, RSD98, SKR^{+}98, SSRP99, SDDV97, TYT^{+}97]. Degraded [CLRR99, LWM^{+}95]. Degree [GBFL^{+}98]. Dehydrogenase [TOSF^{+}98, RWO95]. Dehydrogenase/Amyloid [TOSF^{+}98]. Dehydrogenase/Amyloid- [TOSF^{+}98]. Delay [MAM95]. Delayed [LWKK97, TKH^{+}97]. delaying [RCKS95]. delays [CG95, LCP95, LEB^{+}97, PC98]. Deletion [BPA^{+}97, MHM99, CRM^{+}96, CRCH95, RCC^{+}96, STS96, FH97, KGHG98, MMK^{+}97]. deliver [SMA96]. Delivered [PBS97, GMS^{+}98]. Delivery [CLS96, WSCN97, OS96a, KsYA98, LLVS98, PSB98, SHL97]. Delta [MSMC99, HCM^{+}97, FAO^{+}96]. demonstrate [BTNZ^{+}95]. demonstrated [LCH96, NKI^{+}95]. Demonstration [BJB^{+}96, BMLM96, KRS^{+}98b]. denatured [SHC^{+}97]. Dendrite [YSK^{+}97]. Dendrite-like [YSK^{+}97]. dendrites [CBB^{+}96, HFSHK96, MHM^{+}97, SWO^{+}99]. Dendritic [SB99d, SYB95, DBLD^{+}99, EY99, KTAX98, MWF^{+}99, MLB98, MBBT98, PURB98, SYF^{+}97, TRG^{+}99, WSR^{+}97]. denervation [ACHG95]. dense [CMHT96, FCMN97, KQB^{+}98]. dense-core [CMHT96]. density [FFZ^{+}95, GLP^{+}95, VBG96, YJB^{+}95, KZS^{+}98, MBF^{+}98, ZR98]. density-sensing [YJB^{+}95]. Dental [HKJ^{+}99]. Deoxyribonuclease [RPS^{+}97]. dependence [MB96, VPM^{+}98]. Dependent [BRM^{+}99, HCPW^{+}99, MSMC99, RL99, VG99, VS98, AGPG96, AZF96, Ass97, BJB^{+}96, BHC99, BSW^{+}99, BHD^{+}97, BMHH97, BLRS97, BDvdZW95, BGBK99, BENV97, CS97a, CT97b, CAS96, CDH^{+}98, CNP^{+}95, CSY^{+}95, DBE98, DVL^{+}98, EAL^{+}95, EGS^{+}98, FSH^{+}99, FSR^{+}96, FSAP95, FIC^{+}97, FSHD96, FVRCH96, FNH99, GAZ^{+}98, GHP^{+}96, GACM^{+}95, GRS^{+}99, GLMG98, GPRS^{+}95, HDV^{+}99, HSK^{+}96, HS96, HHR^{+}98, ISFB^{+}96, JSR96, JM97b, JGCS97, KSV^{+}99, KSVZ95, KBW^{+}96, KF95, LCS96, LHWH97, LTE^{+}96, LNY^{+}95, LFB96, LKSW98, LMB^{+}95, MEFH97, MCC^{+}97, MK95a, MAB98, MBF^{+}95a, MSH^{+}97, MBBT98, NMK^{+}97, NGMR97, OWA95, OF98, PWD^{+}99, PMP^{+}98, PW98, PLR^{+}95, PBS97, PWGG98, PYM^{+}98, PT97, PSB98, PL96, RD96a, RHW^{+}98, RSJK95, RSRK96, STK^{+}98, SvDtK^{+}98, SH96, SSN96, SSS99a, SKR96, SKS99, SRR96b, SFR^{+}98, SSA^{+}96, SSP95]. dependent [SHK^{+}99, SWVR98, SMC^{+}96, SSR^{+}98, SMH98, SNK^{+}99, TKVC95, TMM97, TALM96, TSC^{+}98, VLA^{+}98, VvZ99, VJHR96, WLG^{+}96, YNM^{+}99, YAC97, ZOB^{+}96, BF97a, BJR^{+}99, EAOS^{+}98, FCTPA^{+}99, FSM^{+}99, HVL^{+}99, HN98, KBC99, MMR99a, MRL^{+}99, MBF^{+}98, OSR^{+}98, RTM99, SSRP99, TGW99, TRA^{+}99, TLBK98, WLF99, WSWR99]. depending [BSK96]. depends [SSSKR95, vWDG^{+}95, BM98, CLF^{+}97a, LWS^{+}97b, PB98, WCMS98]. Dephosphorylated [CGB^{+}95]. dephosphorylation [BLE^{+}96]. depleted [PMMF95, AC95a]. Depletion [GG95a, BQF^{+}95, TSS^{+}95a, LMS97, ZOS^{+}97]. Depolarization [VMT^{+}99, MLW^{+}95, MAC^{+}97]. Depolarizations [DLC98]. Depolarizes [DH99]. Depolymerization [ISR99, LNY^{+}95, RK95]. depolymerization-dependent [LNY^{+}95]. Depolymerizing [CLS^{+}97, RAA^{+}97, SB99d, AOMB96, DM95]. Deposited [ZOKS99]. deposition [IM96, DAF^{+}97]. deprivation [AGB^{+}96, MFM^{+}95, WMC^{+}95a, HE96]. deprived [DVD^{+}96, MDE^{+}99]. Deregulation [SB99b]. derived [KSL^{+}95, LXK^{+}95, NYN^{+}95, RAM^{+}96, TLF^{+}95, TDW95, DPM^{+}97, HSLB99, SGGS^{+}99, SATW98, TRG^{+}99, GM99, KFL^{+}98, LLSO98, LHS96, MMM^{+}98, SOR^{+}96, TWSD98]. Desmin [LMA^{+}97, MWT^{+}96]. Desmocollin [CT97b]. Desmoglein [CT97b, KMI^{+}97, RS95]. Desmoplakin [GKB^{+}98, BCSG96, RHS^{+}96, KBB^{+}97a]. Desmosomal [CT97b, AYF96, BCSG96, CLT^{+}96, DSGF95, MKF96, KBB^{+}97a]. desmosome [OS96b, SYT97]. desmosome-intermediate [OS96b]. desmosomes [AYF96, RBL^{+}96, GKB^{+}98, LWS^{+}97b, RHPW97, SF98, vHGB^{+}97]. Desorption [WJH^{+}98]. Desorption/Ionization [WJH^{+}98]. despite [VWK^{+}95]. destabilization [FSR^{+}96]. destined [AGPG96]. Detachment [KD97a, NYN^{+}95, GSvR^{+}96]. Detectable [GJW^{+}97]. Detection [CWT^{+}98, STT^{+}96, SSS^{+}99b, SDL98, LDF^{+}95a]. detergent [DvD95, GH95, HE96, SHB99, SBS^{+}98]. detergent-insoluble [DvD95, SBS^{+}98]. detergent-resistant [GH95, SHB99]. Determinant [WBKB97, BPT96, OSHG^{+}96, RSJK95]. determinants [HH95a, CAS97, HCM^{+}97, WS97]. Determination [HTWC98, TYT^{+}99, DFC^{+}96, MLLT95, LMP98, SEBW^{+}98]. Determine [HTK^{+}99, HZC^{+}97, LHS97a, MS98]. determined [OWA95, GF97, JBRA98, JOPM98]. Determines [WSG^{+}98, WPS^{+}98]. determining [PGC^{+}95, APF^{+}97, BSW^{+}98]. Detrimental [SB99b]. detyrosinated [GG95b]. Deubiquitinating [TYK^{+}98]. develop [RFNF96, ULMT99]. developing [AC95b, GYS^{+}95, HWPC96, MLLT95, MMMT95, TTG96, WYY95, PWZS97, PMCS97, SSZ^{+}97, TNCM97, hYGYF97]. Development [AAM^{+}98, BSAJ97, MHM99, QBvD^{+}98, AOMB96, ANS96, BSMH95, CLT^{+}95, DLGB95, DLIB^{+}95, FT95, GEJ^{+}95, HKM^{+}96, IMFS96, KFE96, LKD^{+}95, NA96, OST^{+}95, TYSP^{+}96, VMV^{+}95, WKL^{+}96, WKJH^{+}95, WD95, YRM^{+}96, ZLC^{+}96, ANT^{+}97, BR99, BRS98, BHSAJ98, CL99, CCB^{+}98, DHDJ^{+}97, ESH^{+}99, EFK^{+}98, GKB^{+}98, GBvdN^{+}98, JDK97, JWFS99, JN98, LGM98, MZL^{+}99, MC98, MNZ^{+}99, NBS^{+}98, RFCR^{+}97, SSZ^{+}97, TKH^{+}97, VMH97, VLAHV98, WBG^{+}97]. Developmental [ABP^{+}98, AHCC95, MGA^{+}96, SHPP96, ZHR95, GDF^{+}96, CGSY^{+}99, LNH^{+}98, LSM^{+}98, MPL^{+}97, PAT99, SWK^{+}99, YSEI^{+}99]. developmentally [FGP96, DCR98, Nei98, RKB^{+}98]. Devoid [SHH97]. dexamethasone [PLZ^{+}95]. dexamethasone-induced [PLZ^{+}95]. DHC1b [PDW99, SWO^{+}99]. DHC2 [PDW99]. DHFR [HSWC96]. Di-leucine [DBE98]. di-leucine-based [MWOB96]. di-lysine [SSSKR95]. Diacylglycerol [OTQM98]. Diagnostic [MFF97]. diameter [MWL^{+}96, XML^{+}96]. diaphragm [AHCC95]. Diathesis [KZS^{+}98]. diatom [WFBC96]. dictate [LKWG96]. dictates [JH96, WJWM^{+}98]. Dictyostelium [KJE^{+}99, ATH^{+}97, ASY96, BRS98, BGN^{+}97, BRHD95, CF99, CRCH95, DRR96, EHC97, GDTBD98, JWH96, KFE96, KWS^{+}99, KGHMT95, LWS99, LD95, MTLGC99, NHA^{+}97, NWM^{+}97, NPRT95, PdHTV99, RFNF96, RKB^{+}98, SKS^{+}95, SWS97b, TCM^{+}97, VBG96, WVVD95, YJB^{+}95]. differ [EDB^{+}96, NDBR95, KKH^{+}99, OWM^{+}99]. Differences [CBB^{+}99, DFL^{+}99, HPY^{+}98, RFVCM^{+}99]. Different [BHBB99, GF97, LNH^{+}98, MBRN95, OS96a, YKRS96, BKP^{+}95, BFP^{+}96, DMM^{+}96, EAL^{+}95, FGA^{+}96, KKD^{+}95, KSG^{+}96, MAFJ^{+}95, NDBR95, PPW^{+}95, REdC95, VGM96, WXS96, YFST^{+}96, FFLP99, MWF^{+}99, MC97, RDS^{+}99, SDHM99, SRCC97, TCV^{+}98, VvZ99]. Differential [ANS^{+}97, ALVMM98, BGH^{+}97, BKP^{+}95, CJH^{+}98, CS96a, DMP^{+}96, DBJDT99, EMB^{+}96, HJJJ97, JLC98, KKLA97, LE97, LCB^{+}99, MBW^{+}99, NRD98, NKI^{+}95, OIT^{+}95, PAF97, PSPG^{+}98, PXR^{+}95, RLVM98, RPC^{+}95, SSS^{+}98, TALM96, TFF^{+}95, WSG^{+}98, vGvM95, CQB^{+}96, KKAN^{+}96, KBM^{+}95, LCC^{+}95, HBH98, HHB^{+}97, SL97]. Differentially [DCVF98, HKM^{+}96, WLD^{+}98, GCM^{+}95, LRM^{+}95, RGS^{+}96, SFH96, BKK^{+}97, GTP^{+}99, LMB98, WBS^{+}98]. Differentiate [SFD^{+}98]. Differentiated [HHRdB^{+}98, MAM99, CRG^{+}95, HKS^{+}96, RPC^{+}95, EIIM^{+}98, ILSH99, PGP^{+}99, PHM^{+}98]. differentiating [AKP95, GTL^{+}96, SSL^{+}95, BM97, SWD^{+}98]. Differentiation [AGS^{+}97, CCMC97, HRD98, TNCM97, WXS96, AYF96, AW96, BLB^{+}96, BS95a, DNB^{+}96, DPC^{+}96, EBR96, HG95b, HGW95, JSR96, JH96, KSVZ95, MMMT95, MMS96a, OSK^{+}96, REA95, SLT^{+}96, SSA^{+}96, SBS^{+}96, VKP^{+}95, YSM^{+}95, BAH^{+}98, CJH^{+}98, CEZA97, ELMS98, HPKL97, IJR98, KMM^{+}98, KLO^{+}98, KA98, LMA^{+}97, LMW99, LDD^{+}97, MAC^{+}97, MNM^{+}98, OWW^{+}99, PWZS97, PHN^{+}98, PLB97, PBM^{+}97, QWS^{+}98, RNK97, RLMC99, SGGS^{+}99, SVB^{+}97, SJF^{+}97, SYF^{+}97, SVM^{+}99, TTAH^{+}99, VTHAA99, WDL^{+}98]. differentiation-dependent [SSA^{+}96]. Differing [MSM99]. Differs [TWS97]. diffusion [BWF^{+}96, GG95a, SK95a, PÖS^{+}98, SBV97, TSK98]. dihydrocytochalasin [MAM95]. dihydropyridine [CFC^{+}95, PPF96, PFAF97]. Dilation [BSD^{+}96, MBDS^{+}98]. Dilute [Bri99, WBR^{+}98]. dilysine [LHDC95]. dim1 [BG97]. Dimensional [STG^{+}99, WSSB99, BHDW95, CFK95, CRCH95, DFC^{+}96, EDB^{+}96, JOPM98, LRM^{+}95, MML96, MBF^{+}95b, NHA^{+}97, RME^{+}95, SBSG96, UPAS95, WMO^{+}95, XC97, WPW^{+}97, BMS^{+}97, RNK97, SRR^{+}97, XZSC98]. Dimensions [LMM^{+}99]. dimer [SFSV95]. Dimeric [OZF^{+}99, PdHTV99, HST^{+}98]. Dimerization [SCH^{+}98, BYG96, LOD96, SCG^{+}96, BCAS98, OK98]. dimers [SFSV95, CWT^{+}98]. diphosphonate [HCAW95]. diphtheria [NIM95]. Direct [AKM^{+}98, CT97b, HKBM99, HMS^{+}98, HYK^{+}98, IFM^{+}99, MAT^{+}98, MKC99, PTVD99, RdPRW98, SSI^{+}97, SBG^{+}98, SSS99a, SVH^{+}98, WTCV95, WFBO^{+}96, YTT99, HH95a, MBB96, REdC95, VV95, sHWM^{+}99, INMT97, mLZB99, NG99, PSB98, SHPB99, WS97]. Directed [PC98, HCW96, KKE^{+}95, YNOH95, PDS^{+}97a, CLT^{+}95, SNK^{+}99]. directing [BPT96]. directional [BSB^{+}97]. Directionality [GTP^{+}99]. Directly [HDV^{+}99, FPHH96, BFHB97, BxQK^{+}99, IHT^{+}98, ZLB^{+}97]. directs [MRvD^{+}95, LTW^{+}97, MBTW97, RSD98]. Disassemble [RTL^{+}99]. Disassembled [UNPW98]. disassembly [MKXY95, ALCC^{+}99, BMB97, BKM^{+}98, CLRR99, NLM^{+}98, SCPPW98]. disc [EAL^{+}95, MAM95, GTC^{+}98]. discoideum [BGN^{+}97, CF99, RFNF96, RKB^{+}98]. Discrete [BMPP99, BDvdZW95, DFL^{+}99]. discs [BSMH95, MCL^{+}96, ACH^{+}98]. discs-large [MCL^{+}96]. Disease [YF99, GMS^{+}96, KKA96, MMS^{+}96b, VRD96, CAC^{+}99, EGS^{+}98, FRT^{+}98, GSL98, JM97a, SOJM99]. Disease-Related [YF99]. Dishevelled [MRL^{+}99]. disintegrin [WKRB96, WPMW95, YPM97]. disjunctional [MMGT96]. Dislocation [TSH^{+}98]. Dismantling [BLRS97]. Dismutase [LEB^{+}97]. Disorder [FH97]. Disordered [KZS^{+}98]. disorders [WNB^{+}95]. disorganization [DZC^{+}96]. Dispensable [TGN97]. Dispersal [DP98]. Dispersed [YGG97a]. Dispersion [HTK^{+}98, TZL^{+}98]. displaced [TAK95]. displacement [BCSG96]. displaces [BZB^{+}96]. Display [BGH^{+}97, CRM^{+}96, NMJL96, PKR95, wJTSSL98, KZS^{+}98, VCJA99]. disrupt [NWPW95]. Disrupted [GCGE^{+}96, GSL98, ICW^{+}96, TKH^{+}97]. Disruption [KMI^{+}97, MWT^{+}96, PHM^{+}98, SMG^{+}97, ZLL^{+}98, BWF^{+}96, DSE^{+}95, MM95b, DBH^{+}97, FCTPA^{+}99, RLMC99, SRCC97, WWO^{+}98]. Disruptions [TMM97]. Disrupts [BGH^{+}97, BCSG96, CRG^{+}95, HMB95b, SEC^{+}95, BENV97, MAEE98, MS99].

Dissecting [AHM^{+}96]. Dissection [GSK^{+}99, VZNR98, ZKN99, HAS95, UPAS95, DAB99, GAPG97, HIW^{+}98]. dissociate [SFSV95]. Dissociation [BBZ^{+}95, SPL^{+}97, YTY^{+}99, BWF^{+}96, DFC^{+}96, HUU96, SMB^{+}95, CHH97, KKS99, KBF^{+}98, SYT97]. dissolution [ZN95a]. distal [CLS96, PMH96, MMS96a]. Distal-less [MMS96a]. distance [MC95]. Distinct [BOS^{+}97, BFP^{+}96, CRMB95, CLK99, FKWP97, HSML98, HFP^{+}98b, MM99b, MF96b, MYTK99, NAsG^{+}98, PBS97, PSH^{+}98, SBS^{+}99, SKY^{+}98, SMS^{+}98, SKC97, VYB95, VvZ99, ZGYS99, BSN95, CCR96, DBBB95, GSC^{+}96, HSL^{+}95, HAS95, JWH96, KI96, LMMO96, MF96a, MWOB96, MBRN95, NCZ^{+}95, NSYK^{+}95, PHO^{+}95, SG95, ST96, SG96, SSW^{+}96, VGM96, ABT^{+}98, CAH^{+}99, CFS98, CLF^{+}97b, DEPR97, GPKH99, HW98, HC99, HKO^{+}98, KCV^{+}98, LMB98, PRB97, PvDA^{+}98, SLAS^{+}98, SDHM99, TSD^{+}97, WNB97]. distinctive [SHPP96]. distinguish [NDBR95]. Distorter [PKBHK97]. distributed [PJL^{+}95, VBZ96]. Distributes [MPLS^{+}98]. Distribution [CJG^{+}95, HYK^{+}98, KE98, MB99, PMCS97, vIH98, ABR95, DMP^{+}96, DG96, LLK^{+}95, LvSC^{+}96, SWR^{+}96, YKC95, BD97, BHC99, BHSAJ98, BENV97, GGK97, HBH98, LL98a, MNR^{+}98, MQR^{+}97, PvDA^{+}98, SSRvdB98, SMG^{+}97, XZMD97]. Distributions [GF97]. Disulfide [FZZN98, GLL^{+}99, HTW97, KGHG98, LG99, RDS^{+}99]. Disulfide-bonded [KGHG98]. Dithiothreitol [HTW97]. divergent [SS95b]. diverse [SLS^{+}96, SCKG95, TBD^{+}99]. Diversity [Bor95, AASH^{+}96, MAG^{+}96, PC99]. Dividing [KLH98]. Division [SJ99, BBZ^{+}95, Gol95, GJ95b, HCW96, KJBM^{+}96, MDQG96, MK95b, BSW^{+}98, BHSAJ98, CDN97, GSK^{+}99, KLL^{+}97, MCM^{+}99, OWAlW99]. divisions [REA95]. dj2 [TKBM97]. Djp1p [HRK^{+}98]. Dlg [WHP^{+}96]. DM [DC96b, MRvD^{+}95, LBWF^{+}97]. Dmc1 [TMPM99]. DNA [ALM97b, BBZ^{+}95, CJR^{+}97, CL99, DC96a, EKS97, FB96, GHW99, HTS^{+}95, HBS^{+}97, HYN97, HN98, IM96, JCPK95, JHM98, JE98, LKWG96, LNH^{+}98, LK96, MSD^{+}98, MKC99, MC95, MWGE97, MTW^{+}99, MG96, OPB98, PMP^{+}98, RSL^{+}96, RBRB99, REdC95, SFD^{+}96, SHS96, SVS97, SIUW98, SSS^{+}99b, SMG^{+}97, TAK95, TRA^{+}99, WA95, WXS96, WSA^{+}99, YN95, ZSMV95]. DNA-binding [HTS^{+}95, HBS^{+}97, WA95]. DNA-damaging [PMP^{+}98]. DNA-immortalized [ZSMV95]. DnaJ [HRK^{+}98, HUU96, SHR^{+}95, SSSG98, TKBM97]. DnaJ-like [HRK^{+}98]. DNM1 [GKVR95]. Dnm1p [OKB^{+}98, SJ99]. do [KR96, GJW^{+}97, KR99]. Docking [MW97, EKK^{+}96, GKM^{+}96, CWSG99, FZZN98, KPW^{+}97, SLL^{+}98, UNPW98, VCL^{+}98]. Does [KBB^{+}97b, MWGE97, PWG^{+}95, SEC^{+}95, EHC97, GFN^{+}99, LB98, RHW^{+}98, TCV^{+}98]. dog [AASH^{+}96]. Domain [BBOE98, CWSL97, FSDA97, KNH^{+}97, LSZ^{+}99b, MBA97, MSH^{+}97, TNM^{+}99, WLF98, WBKB97, WC97b, YLB^{+}97, ZLPL99, ASH95, BLG95, BLL^{+}96, BLL95, BRJP96, BTNZ^{+}95, BS95b, BSMH95, CLL95, CRM^{+}96, DLB96, DWS^{+}95, EKK^{+}96, FFMG95, GLS96, HKS^{+}96, IM96, KKD^{+}95, KGMF95, LW96, LGRB96, LF95, LNY^{+}95, LBCB96, MCL^{+}96, MAS^{+}95, MMATe96, NCZ^{+}95, PWG^{+}95, PJL^{+}95, PXR^{+}95, RS95, SSN96, SPMB^{+}96, SMRT96, SHR^{+}95, SD95, STP96, SRRB95, TGG95, WGR^{+}95, WHH96, WSG^{+}95, WPMW95, WF96, APF^{+}97, ABE^{+}98, AL97, BALL98, BINRM97, BKN^{+}97, CSY98, CWMD97, CIK^{+}99, CS97b, DSC^{+}99, DW97, FNDS99, FHSM^{+}97, GSKL99, HSVS98, HMB^{+}98, HTWC98, HR99, IFSV^{+}98, KOM^{+}98, KRGA98, KBB^{+}97a, KJR^{+}98, LDP^{+}99, LRG^{+}98, LNRR98, MNS^{+}97, MLB98, MS98, MAE^{+}97, MSFA99, OBW98, OK98, PNEH98, PH98, PLB97, RKR^{+}97, SZFM98, SSW^{+}98]. Domain [SAR^{+}99, SHPB99, SVB^{+}97, SBPR^{+}98, SSEM98, SDDV97, TSB99, WS97, WWO^{+}98, WG97, XWK^{+}97, YHD^{+}98, YPM97, ZH97]. domain- [DLB96]. Domain-2 [HR99]. Domain-associated [CWSL97]. domain-binding [MCL^{+}96, MMATe96]. domain-containing [BLL^{+}96, FNDS99, PH98]. Domains [DFD^{+}99, FhJZ^{+}99, FFB^{+}97, FPRL97, HSML98, MKW^{+}99, NAsG^{+}98, VB98, ZB98, ASSS^{+}95, BTTB96, BDvdZW95, DTB96, GBMS95, HAS95, HUU96, HMB95b, LBCB96, MCL^{+}96, NHB^{+}95, OGS^{+}96, OWA95, RR96, SLT^{+}96, SOHP95, SHS96, WKRB96, XJM^{+}95, vdBLCvM96, BBP^{+}98, CMS^{+}97, CR99, CDH^{+}99, FCMN97, FGA^{+}98, HIWL97, IIM^{+}99, KRCP99, KSK^{+}99b, LMF98, LDP^{+}99, NGMR97, OTQM98, OF98, PFKvD99, RBRB99, SMW^{+}99, SBS^{+}98, SEB99, TRS^{+}99, VHF97, WJWM^{+}98, XCLM98, ZGYS99]. dome [APG^{+}96]. Dominant [WF96, vABSP95, LCH96, MHLB96, TYSP^{+}96, ABL98, LMB98, MAEE98, RDP^{+}99, HDDKH98]. Dominant-negative [WF96, MHLB96, ABL98, LMB98]. Dominantly [BSRG99]. Dopamine [VvZ99]. Dormancy [GKO99, YKO97]. Dorsal [DBC^{+}99, EGA97, GS99, MRL^{+}99, RV98]. Dorso [LTR^{+}97]. Dorso-ventral [LTR^{+}97]. Dorsoanterior [VWG^{+}97]. dosage [SMC^{+}96]. dosage-dependent [SMC^{+}96]. Dot [SJW97, SJZW97]. Dot-associated [SJW97]. double [DH96, NPRT95, CTB^{+}97, DCJ97, EJ97, QPH^{+}98, RBRB99, RHA^{+}97]. double-strand [DH96]. Double-stranded [CTB^{+}97, EJ97]. down [GBBSO96, SOPM^{+}97]. down-regulates [GBBSO96]. Downregulates [GM99, StKvD^{+}99]. Downregulation [AHM^{+}95, GKO99, GvdFvD^{+}99, SHMC99, SEH^{+}97a]. Downstream [NCW^{+}99, LYR96, PHB96, ANT^{+}97, CWSG99, KHA^{+}99]. DR [CWT^{+}98]. Drab6 [PAT99]. dragging [SK95a]. DRAP [HIG^{+}95]. DRAP27 [NIM95]. Drive [HOCK98, RTM99]. driven [All95, BML^{+}97, OMS98]. drives [CWB96]. Drosophila [BPKB^{+}99, CRM^{+}98, DLGB95, EGA97, LXC^{+}98, MAG^{+}96, MDK^{+}99, RC97, AAM^{+}98, ASH95, ASSS^{+}95, BBL^{+}99a, BGG98, BSAJ97, BHSAJ98, BDS97, CKP96, CH98, DRAT95, DKD^{+}96, EAL^{+}95, EWS96, EK97, FaAR^{+}96, FAZ^{+}99, FMD^{+}98, GCA^{+}98, GDB^{+}98, GPSF99, GB98, GCST97, GBW^{+}95, HCPW^{+}99, HCM96a, HHM^{+}98, HFP^{+}98b, JK97, KP96, KSM96, KAFB99, KF95, LMF98, LM98, LBWF^{+}97, gLSE^{+}99, MZL^{+}99, MMN^{+}99, MF96b, MGS^{+}95, MM95a, MAE^{+}97, MBB^{+}97, MBF^{+}95b, MZAO98, MW96, MM96, NGT99, OWM^{+}99, PGD^{+}97, PDS^{+}97a, PCQH98, PB98, PURB98, RPY^{+}99, RWS^{+}99, RV98, SdCAH^{+}97, SWK^{+}99, SMB^{+}99, SSB96, SO97, SO98, TTG95, TCSG96, TTG96, TSS^{+}95b, TOSF^{+}98, UPAS95, VCJA99, WLF98, WAV98, WRW^{+}95, WGG96, WBSNV97, WHP^{+}96, YF99, ZT99, ZLC^{+}96]. Drs2p [CIRG99]. DSK1 [WFBC96]. Dual [DMS^{+}95, HR99, LL98a, NNSN97, RRP^{+}95, TNFK^{+}98, VLM^{+}98, MKF96, Bri99, FKO^{+}98, sHWM^{+}99, LAS^{+}97]. Dual-Specificity [NNSN97]. Ductal [YHMS^{+}98]. Due [NMT97, PBJ^{+}98, SVM^{+}99]. Dumping [GCST97]. Duo1p [HCG^{+}98]. Duplication [TRS^{+}99, BIR96, DK96, WW96a, AK99, MB97, SGSW97, VWO99]. during [AAP^{+}95, ANS96, AK99, ANT^{+}97, ALVMM98, AHCC95, AWB^{+}98, AZF96, AWC99, BMS^{+}97, BAC97, BMS95, BLRS97, BGN^{+}96, BSAJ97, CL99, CMD^{+}98, CSO97, CF99, CEZA97, CJG^{+}95, CCR96, CKP96, DRAT95, DDWG96, DOSN^{+}99, DK96, DFL^{+}99, EPVV96, EZC^{+}97, EM96, EM98, EKSC98, ETF95, EK97, EKF^{+}97, FFGG96, FK95, FSA^{+}99, FWG^{+}96, GCAHW97, GCST97, GSvR^{+}96, HPKL97, HG95a, HAS^{+}96, HCW96, HGW95, HYN97, HHR^{+}98, JE98, JDK97, JWFS99, KRH98, KsYA98, KCBR96, KA95, KSY97, KFE96, KJBM^{+}96, KLL^{+}97, KZKS96, LGM98, LHDC95, LZD95, gLSE^{+}99, LO96, MKLS96, MCS^{+}97a, MAC^{+}97, MGP97, MZL^{+}99, MDE^{+}99, MOY^{+}98, MC98, MHM99, MM97b, MGRS96, NMK^{+}97, NA99, NGT99, NH99, OWAlW99, OTL^{+}97, OSK^{+}96, PWZS97, PDS^{+}97a, PH97a, PCQH98, PTJ^{+}98, RHHRB96, RPW96, RS99, RC98, RWS^{+}99, RACHV96, SAR^{+}99, SHC95]. during [SWS^{+}96, SALdP^{+}97, SVH^{+}98, SCEB99, SHP^{+}97, SIUW98, SRG^{+}96, SFNRH95, SNCBM95, SSM98, TSL^{+}98, THS^{+}98, TIBL99, TAK95, TFF^{+}95, UPT98, VKP^{+}95, VMV^{+}95, VMN95, VLAHV98, VIF^{+}96, VJHR96, WS95b, WBSNV97, WHS^{+}97, WM95, WD95, YHS^{+}99, YTY^{+}99, YGG97a, YGG97b, YSB96, ZBG99, dPSMNSM95, vGvM95, BGG98, BHSAJ98, CH98, LHR^{+}97, PNEH98, RLVM98, RME98, SEBW^{+}98, XZMD97]. DxxxLL [DKN^{+}97]. dy [VKR^{+}96]. dy/dy [VKR^{+}96]. Dynactin [QGE^{+}99, EPVV96, HTKH96, VV95, AEVB98, BENV97, EGM^{+}99, KLH98, MTLGC99, SWHG98]. Dynamic [BKMN^{+}98, BSAJ97, DPG^{+}99, JCWP95, KLW^{+}96, KWCS96, PFP97, SHS96, SLPE^{+}98, The97, VB98, WDV^{+}97, WFL^{+}99, XZMD97, BJB^{+}96, FHWV97, GBS99, GGD^{+}97, HDES97, KR99, KLL^{+}97, LHR^{+}97, PCD^{+}98, SFA^{+}98, TWS97, WSS97, WWD^{+}97]. Dynamics [AWC99, BB97, BMPP99, BD97, GMBSS99, LGM98, LSB98, LLK^{+}95, MAM99, PBA^{+}97, REC97, SJC96, ZP98, OHB^{+}95, RSP^{+}95, SFSV95, THK95, VE95, WSGPS95, YSC^{+}95, ZKB95, ZKSB96, ACF^{+}97, BD98, BW97, CHW^{+}99, ESM^{+}97, EK97, GLMG98, HSLB99, HKO^{+}98, KSY^{+}98, LGZ^{+}99, LL98a, MSD^{+}98, MI99, MCM^{+}99, MKC99, MPCW97, NTH98, SFA^{+}98, SYM^{+}97, SGA97, SMTN99, TOB^{+}99, WHG^{+}97, WBR^{+}98]. Dynamin [HKOM98, OMS98, DBvdBS95, GKVR95, HM96, TMDD96, ABT^{+}98, LRS^{+}98, OKB^{+}98, SY99, SSRvdB98, VvZ99, YPDM98]. Dynamin-1 [ABT^{+}98]. Dynamin-2 [ABT^{+}98]. Dynamin-dependent [VvZ99]. dynamin-like [HM96, SY99, YPDM98]. Dynamin-mediated [HKOM98]. dynamin-related [GKVR95, OKB^{+}98, SSRvdB98]. Dynamitin [BENV97]. Dynein [MTLGC99, All95, KPKWW95, MGS^{+}95, NAV96, PMH96, ROG^{+}96, SKE^{+}95, VGM96, VV95, WKK^{+}95, YSC^{+}95, AEVB98, BRLM98, BENV97, CS97a, GPKH99, HS97, HTK^{+}98, HOCK98, KLH98, KMHK97, KD97b, KPW^{+}97, MTLGC99, MKW^{+}99, PKBHK97, PWW98, PDW99, PTVD99, RWS^{+}99, SWO^{+}99, SWHG98, Ste97, WBA^{+}98, YWMH99]. Dynein-dependent [BENV97, CS97a]. dynein-driven [All95]. dynein-related [SKE^{+}95]. Dyneins [HOCK98]. Dysfunction [MAB98]. dysgenic [PPF96]. Dystonin [LSL99]. Dystrobrevin [PSPG^{+}98]. Dystroglycan [MLC99, CJG^{+}95, DLIB^{+}95, CJY^{+}97, PXRR99]. Dystrophic [TDR^{+}98, WB99]. dystrophin [AK95, CRM^{+}96, PWE95, RCC^{+}96, RAE96, SYO95, VWK^{+}95, BHBB99, GB98, HLJ^{+}98, KFM99, PAF97]. dystrophin-deficient [VWK^{+}95]. Dystrophin-like [BHBB99]. dystrophy [TMS^{+}95, DSM^{+}98b, SRCC97, SEAB^{+}99, SSH^{+}98].

E- [LVN^{+}98]. E-cadherin [ABRB^{+}95b, ACSN98, CT98, RS95, CSN99, HMS^{+}98, SHT^{+}95]. E-Cadherin-dependent [SSR^{+}98]. E-cadherin-mediated [ABN96, FSIG^{+}95]. E-Cadherin/ [CSN99]. E-cadherin/catenin [SHT^{+}95]. E-Cadherin/Green [ACSN98]. E-cdk2 [ZOB^{+}96]. E-selectin [GGD^{+}97, KCN^{+}96, YWW^{+}96]. E. [VSGF96]. E1A [CSO^{+}95, LK96]. E1B [PW98]. E1B19K [MFM^{+}95]. E2 [SRA^{+}99]. E3 [KKD^{+}95, PSBB96]. E3/19K [PSBB96]. E4 [CSO^{+}95]. E4-ORF3 [CSO^{+}95]. E48 [BGK^{+}95]. E4orf4 [LNM^{+}98]. E587 [BOLS95]. Each [TCV^{+}98]. eag [PBCS98]. ear [KLMR95, HGG^{+}97]. Earlier [NAsG^{+}98]. Early [AAM^{+}98, GSvR^{+}96, JCPK95, KTD^{+}97, MAT^{+}98, RDP^{+}99, SCR98, GM95, HW96a, MGY^{+}95, MBF^{+}95b, MW96, PLZ^{+}95, RHHRB96, SWS^{+}96, TYSP^{+}96, VMV^{+}95, VJHR96, WPLK95, WWP^{+}96, WKJH^{+}95, ZLC^{+}96, CL99, hCMPG97, CLF^{+}97a, DTMG99, DDM98, GKB^{+}98, HMB^{+}98, HHR^{+}98, ISM97, KRS98a, LTR^{+}97, LTMR97, LGM98, MFF97, NG99, NZ97, PFHWN98, RC98, SDHM99, TPMB97, VLAHV98]. Early/Recycling [MAT^{+}98]. eat [SLP^{+}96]. eat-5 [SLP^{+}96]. EB [BHHF97]. EB-1 [BHHF97]. EBP50 [RBB97, MKB^{+}99]. Ect2 [TXB^{+}99]. ectodomain [CDRB^{+}96, DMY^{+}97]. Ectopic [NMH^{+}98, VWG^{+}97]. ED [SBPR^{+}98]. EDA [iMOeM^{+}97]. Edg [VLM^{+}98]. Edg-1 [VLM^{+}98]. Edge [BMC^{+}99, KGHMT95, PTBC96, BKJL^{+}98]. editing [EJ99]. EEED [CGJ^{+}95]. EF [MBP^{+}95]. EF-hand [MBP^{+}95]. Effect [AGS^{+}97, DSE^{+}95, KvEdV^{+}96, SGYL^{+}96, WF96, vABSP95, BKM^{+}98, vHMV^{+}97]. effectively [LK95]. Effector [ELL^{+}99, SMZ^{+}98, CGC^{+}97b, DSP97, FLA99]. Effects [BGH^{+}97, JN98, SKWY95, WRG98, CRMB95, EMB^{+}96, WXS96, YtDH^{+}95, DPZ^{+}97, GGK97, OC98, SPS^{+}97a, SKS99, SKY^{+}98, SZNS98, SSSS97, VHF97]. Efficiency [GHW99]. Efficient [AM95a, CSN99, LGRB96, MR96, VH96, JP98, SHL97, YWMH99]. efficiently [SLS^{+}96]. efflux [MBD^{+}96]. Eg5 [MSH^{+}99]. EGF [NIM95, BLB^{+}96, DMY^{+}97, FPHH96, HIG^{+}95, MPY^{+}96, MSH^{+}98, MLW^{+}95, MBF^{+}95a, NIM95, SVB^{+}97]. EGF-induced [MSH^{+}98]. EGF-like [HIG^{+}95, NIM95, SVB^{+}97]. Egg [GRB^{+}97, LKWG96, MMS^{+}96b, SK96, SFSV95, WSGPS95, CLL^{+}98, HTH^{+}97, MCJK98, SBGR98, TGN97, TGTL97, TYT^{+}97, WVM97, YPM97]. eggs [MST^{+}96, WTCV95, CRM^{+}97]. Eh [DSC^{+}99]. Eif [KGL^{+}99]. Eif-2 [KGL^{+}99]. eIF6 [SPV^{+}99]. Either [SvDtK^{+}98, JMTCF96, sHWM^{+}99]. Elaboration [RB98]. elastase [OCS^{+}95]. elastic [SHS96]. Elasticity [HMCL97, HD99, TGL^{+}98]. electrical [BWF^{+}96]. electron [FTH96, KKM^{+}97, KGGK99]. Electrospray [SBS^{+}99]. elegans [BBOE98, BTTB96, CDH^{+}98, CRA^{+}98, FR98, GKR95, Gol95, GPKH99, GSK^{+}99, GJW^{+}97, GGK97, HMC^{+}99, HW96b, HSSW99, IMFS96, IHT^{+}98, KNT^{+}99, LHK^{+}96, LBO^{+}98, MCA96, MCTMK98, MAB98, MMR99b, MGA^{+}96, OBB99, OFY^{+}95, SGD98, SCEB99, SWO^{+}99, SLP^{+}96, TWiK^{+}99, WG99]. element [MC95, KEJ97, LRC^{+}99]. elements [REdC95, TKM^{+}96, AAD^{+}97, LNH^{+}98]. Elevated [WBG^{+}97]. elevation [CM95a, mLZB99]. Elicited [MGD97]. Elicits [VHF97]. Elimination [MZNR98, RPS^{+}97]. elongation [LTE^{+}96, MELC96, NCZ^{+}95, WFBC96, DSG98, GMGGD98, RYB^{+}99, RSS98]. Embryo [BSAJ97, CM95a, Gol95, CRA^{+}98, GCAHW97, GPKH99, GSK^{+}99, NG99]. embryogenesis [CCR96, CKP96, FSA^{+}99, MCS98, SCEB99]. Embryonal [CCMC97, OSK^{+}96]. Embryonic [BHD^{+}97, FFMG95, JDK97, BEG^{+}95, BMLM96, CJG^{+}95, FPM^{+}95, GF95, KSS^{+}95, MBB96, RD96a, RBL^{+}96, SSA^{+}96, TKM^{+}96, ZLC^{+}96, CJY^{+}97, DBC^{+}99, DBJDT99, GJB^{+}98, GSP^{+}98, MZL^{+}99, MDK^{+}99, NMK^{+}97, PHM^{+}98, RHP99, RYB^{+}99, SFD^{+}98, SMB^{+}99, SVM^{+}99]. Embryos [BDS97, SGD98, DKD^{+}96, FHV^{+}96, KJBM^{+}96, MBF^{+}95b, MKI^{+}96, WWP^{+}96, BHSAJ98, hCMPG97, LM98, LTR^{+}97, MRL^{+}99, MS97b, SGR97, VWG^{+}97, dS98a]. emergence [MR96]. emerging [WSW95]. EMMPRIN [MGB^{+}98]. emo [IMFS96]. emo-1 [IMFS96]. emp24 [DDF^{+}98]. emp24/p24 [DDF^{+}98]. Emp47p [SSSKR95]. employing [LCH96]. Employment [Car99]. Ena [LLH^{+}99]. Ena/VASP [LLH^{+}99]. enables [WHH96, BPS99]. ENaC [AAM^{+}98]. encoded [MKB^{+}95, OABD^{+}96, RRRA95, RZZ96, GKS^{+}98, KSY97, HOCK98, RKB97, YCR^{+}98]. encodes [BTTB96, EGH^{+}95, FT95, GEK95, HWPC96, MCA96, MBP^{+}95, MGM95, SCKG95, SL96, SKK95, TWVC95, TSS^{+}95b, ZL95b, BBR97, BG97, BF97b, EGA97, GB98, KPW^{+}97, PGD^{+}97, SdCAH^{+}97]. encoding [GBW^{+}95, KBSW95, MR96, BGN^{+}97, ESR97, PAT99, KMHK97]. end [IRMF^{+}99, KKE^{+}95, LJ95, MGRS96, YNOH95, EGM^{+}99, PDS^{+}97a, SNK^{+}99]. end-directed [KKE^{+}95, YNOH95]. end-specific [LJ95]. Endcompartments [BAC97]. Endocardial [NMK^{+}97]. Endochondral [SKS99, YSEI^{+}99]. endocrine [AS96]. Endocytic [BQF^{+}95, MSM99, FPWN95, GLP^{+}95, HP95, MLJM95, PCYS95, SJHC96, GMS^{+}98, KMG^{+}97, LGGS97, LTW^{+}97, OT97, PFHWN98, VvZ99, WBS^{+}98]. Endocytosed [GMS^{+}98, LRS^{+}98]. Endocytosis [DTB96, GK96, HSL^{+}95, LNR^{+}96, RJP^{+}96a, SK96, BMS95, DSGF95, HH95a, KP96, RWO95, SRH^{+}95, THP96, VE95, WPLK95, WMXE96, ALK^{+}99, BLB^{+}98, BGN^{+}97, DSC^{+}99, KWS^{+}99, MBF^{+}97, NBJ^{+}99, RS99, RSD98, SOPM^{+}97, SN97, SDDV97, TFYH99, WE98, ZPL^{+}98]. Endoderm [SVM^{+}99]. endogenous [BEZ^{+}95, KGW^{+}95, TEL^{+}95, BL99, KGHG98, SMW^{+}99]. Endoglin [LLZ^{+}96]. Endolyn [IMS^{+}98]. Endolyn-78 [IMS^{+}98]. Endonexin [KSE^{+}97, SOE^{+}95]. Endopeptidase [BCWA97]. Endoplasmic [BB97, CSN99, CGC^{+}97a, FFDP98, FNH99, GHS98, LAW^{+}98, MB99, MAS^{+}97, MZN97, NNK^{+}97, SSN96, YHS^{+}99, ZWT^{+}97, All95, ABRB95a, BRB96, GEK95, GG95a, KA95, KKA96, LCP95, LHDC95, MBS96a, MBM95, MTLW95, NBW96, RAM^{+}96, RW95, SHR^{+}95, SOR^{+}95, SNCH95, VJHR96, AFG^{+}98, ABK^{+}97, DSM^{+}98a, EGS^{+}98, GLL^{+}99, HLMD97, HTW97, sHWM^{+}99, JBRA98, LPD^{+}99, MJBR98, NZ97, RCE^{+}99, RSB^{+}99, SHC^{+}97, SPL^{+}97, SSSG98, TH97, TDR99, VKIH98, YGG97a, YPDM98]. Endoreplication [SO98]. Endosomal [BPWS98, MM99b, AGPG96, AM95a, GKVR95, PPW^{+}95, BF97b, FRT^{+}98, FGA^{+}98, jLC97, WC97a]. Endosome [SMCE97, SME98, URU^{+}96, WPLK95, DSG^{+}97, DSP97, MTK^{+}98]. Endosome-to-Golgi [SME98]. Endosomes [MAT^{+}98, AGPG96, FPHH96, GM95, MTL^{+}96, OPD^{+}96, SOG96, vWDG^{+}95, AYP^{+}99, GAPG97, MBF^{+}98, PKCS98, SDHM99, WC97a, ZSP^{+}98]. Endothelial [FSDA97, HRD98, LCN^{+}97, LVN^{+}98, MSF^{+}99a, MSFT99, MA95, YMAC^{+}98, CCB^{+}95, DZC^{+}96, GHP^{+}96, GW95, HMB95a, KvEdV^{+}96, LCC^{+}95, PLR^{+}95, PFGS95, ANS^{+}97, AEWG^{+}97, BBL^{+}99b, CAR^{+}97, CFS98, DPZ^{+}97, DBC^{+}99, EvdESvdH^{+}97, EWW^{+}98, KDO^{+}99, KA98, LHS97a, MDV98, NMK^{+}97, NRD98, SAC^{+}98, SPR^{+}98, SGJP99, TGW99, TCYS99, WSWR99]. Endothelin [CMS98, TPM^{+}99]. Endothelin-1 [CMS98]. Endothelin-converting [TPM^{+}99]. Endothelium [CBM^{+}98, PHU^{+}95, YWW^{+}96, dPSMNSM95, GCM^{+}97, OMS98]. Ends [BMC^{+}99, BWKH96, CFK95, MMP^{+}95, MELC96, SJC96, WSGPS95, DPG^{+}99, ILSH99, TWS97]. energetics [ST96]. energy [KE98, BF97a]. Engaged [TDR99]. Engagement [AWC99, PWD^{+}99]. Enhance [KBC99, SSEM98]. Enhanced [CFB^{+}98, SCB^{+}97, AGP^{+}98, IKR^{+}98, KEP^{+}99, RFK^{+}97]. Enhancement [TGW99, NHB^{+}95, NTAN98]. enhancer [BGGR^{+}96, OMW^{+}96]. Enhances [CLS^{+}97, APG^{+}96, SSM^{+}95, KSK^{+}99b]. Enlarged [ESR97]. Enriched [RMY97, RR96, SS95a]. Enrichment [DMY^{+}97, MRL^{+}99]. enriettii [PXR^{+}95]. entactin [AHBW96]. Enterocytes [CCMG97]. Enterotoxin [KIH^{+}97, SFS^{+}99]. Entrapment [SIUW98]. Entropic [TGL^{+}98]. Entry [HFP98a, LR98, RC98, AAP^{+}95, FHCL95, WWP^{+}96, CHC97, JHM98, TGN97, VPM^{+}98, WWM^{+}98]. Envelope [ORW^{+}95, RST^{+}99, MKLS96, RHS^{+}96, SPMB^{+}96, ST96, STRM95, ZKSB96, ZN95a, ANK97, DFL^{+}99, FHWV97, GHW99, HPD^{+}97, KCFS98, LHS97b, MGM98, MS98, MBDS^{+}98, RDP^{+}99, SEAB^{+}99, TGTL97, YGG97a]. envelopes [SFSV95, dCBR95, LBWF^{+}97, RHPW97, THS^{+}98]. environment [BEG^{+}95, WSW95, ISM97]. Envoplakin [RHS^{+}96, RHPW97]. Enzymatic [RRRA95, KSG^{+}96, ZL95c]. Enzyme [KSF^{+}99, MMAK97, SSY^{+}96, SKR96, EJ99, FCMN97, HBPJ98, HCLG99, MCS^{+}97a, MKS^{+}97, TYK^{+}98, TPM^{+}99, VKIH98]. Enzyme-1 [TPM^{+}99]. Enzyme-like [MCS^{+}97a]. enzymes [BDE^{+}95, KDvFP96, KKLA97, LLG^{+}99, SPL^{+}97]. Ep [LBW^{+}97, CCB^{+}98]. Eph [BSM^{+}99, SHPP96]. Eph-related [SHPP96]. Epidermal [BMC^{+}99, JCR97, SSM^{+}95, ADP^{+}95, AYF96, CGSP95, CMUW96, FLW^{+}95, GJ95a, HBJ95, LS95a, MMS96a, Car99, DHDJ^{+}97, EPMB^{+}99, PC99, SGGW99, SVMB97, WSR^{+}97, WWO^{+}98]. epidermis [BSK96, PLM^{+}96, MDK^{+}99]. Epididymis [KKT^{+}98]. epiligrin [XGC96]. Epimorphin [HLG^{+}98]. epiplasmic [HGP^{+}95]. Episialin [WvdVV^{+}95]. epithelia [KCDB95, LCM^{+}95, LYC^{+}95, MHK^{+}95, PCL^{+}96, SSB^{+}95, WHP^{+}96, GSP^{+}98, HGG^{+}97, HPY^{+}98, MKB^{+}99, WJWM^{+}98]. Epithelial [ACSN98, CWM^{+}98, GSP^{+}98, KD97a, LBW^{+}97, RWOA97, SvDtK^{+}98, TSC^{+}98, ZT99, AHBW96, BFS^{+}95, CLR^{+}96, DvD95, DNC^{+}95, DSC^{+}95, DLIB^{+}95, FSR^{+}96, FLPS95, FSAP95, HMB95b, HWC^{+}96, KKD^{+}95, LKD^{+}95, LLK^{+}95, LO96, LDCB95, LLRTP95, MAFJ^{+}95, MSF^{+}95, MBF^{+}95a, MVvdBD96, MLLT95, MMMT95, PJL^{+}95, RME^{+}95, SWB^{+}96, SSM^{+}95, WW96b, ZSMV95, vdBLCvM96, BMG^{+}99a, CSO97, CCB^{+}98, CGC^{+}97b, DOP^{+}98, FK99, GvdFvD^{+}99, HLG^{+}98, IFSV^{+}98, IHT^{+}98, JN98, KFM99, KRCP99, LGZ^{+}99, LNRR98, LMC^{+}98, LGM^{+}97, MLN^{+}98, MC98, PBA^{+}97, RPS^{+}97, RLMC99, SFD^{+}98, SRV^{+}97, Sal99, SYT97, SSZ^{+}97, SW98, TWSD98, VTHAA99, WUUI^{+}98, WSS97, WWO^{+}98, WRG98, YHK^{+}97, ZSP^{+}98]. epithelial-mesenchymal [MLLT95, MMMT95]. epithelial-stromal [CLR^{+}96]. epithelialization [PTBC96]. epithelium [EAL^{+}95, HG95b, MW96, BFRB99, DCR98, FLO^{+}99, FAZ^{+}99, HKJ^{+}99, MGB^{+}98, YHMS^{+}98]. epitope [NSR95, WHM^{+}96, SG97, SSS^{+}99b]. Epitope-defined [SSS^{+}99b]. Eps15 [BLB^{+}98, BGB^{+}95a, WMXE96, WE98, DSC^{+}99, vDSH^{+}97]. epsilon [PHB96, PMCT96, TCM96]. Equator [KCM^{+}97]. ER-associated [MB96]. ER-derived [LHS96]. ER/ [LLG^{+}99]. ER/ERGIC/cis [ISH95]. ER/ERGIC/cis-Golgi [ISH95]. ErbB [VC97]. ErbB-4 [VC97]. erbB2 [NBS^{+}98]. Erbb4 [JWFS99]. Erd2p [MSW^{+}98, MSW^{+}98]. Erg30 [SBD^{+}99]. ERGIC [ISH95, ISH95, VKIH98]. ERGIC-53 [ISH95, VKIH98]. ERK [BZBA99, HTK^{+}99, NCW^{+}99, SVH^{+}98]. ERK/p38MAPK [HTK^{+}99]. ERM [YHD^{+}98, HSK^{+}96, KTD^{+}97, MMD^{+}98, YTT99]. ERS [POR^{+}97]. ERS-24 [POR^{+}97]. Erv14p [PB98]. Erythrocyte [TSK98, WGF^{+}98]. Erythrocytes [TLBK98]. erythroid [ABRB^{+}95b]. escaped [MR95]. Escherichia [LCM^{+}95]. Esi [SBS^{+}99]. Espin [BZL^{+}98]. Essential [ABP^{+}98, BBL^{+}99a, COML95, DBE98, FFDP98, KSF^{+}99, KGHG98, MEFH97, SPV^{+}99, SME98, TOSF^{+}98, WH97, ZT99, AOMB96, CRLM96, DPC^{+}96, DK96, DBBB95, GCB^{+}95, GKR95, HY95, HG95b, HB95b, IWW95, IFHR^{+}96, JRLS95, LN96, MBP^{+}95, MGM95, NNMW96, RLCM96, RBL^{+}96, SCK^{+}95, SCKG95, SS95b, SDTE95, WVVD95, ZL95b, ZLC^{+}96, AYMB99, AKLME98, ACH^{+}98, BMP^{+}98, BSD97, BHJ98, BG97, BFL^{+}97, BBJ98, DRE97, EGA97, FRHS98, GWB^{+}98, GCA^{+}98, GDB^{+}98, KHA^{+}99, LMA^{+}97, MGIZ98, MVM^{+}99, MNZ^{+}99, MAE^{+}97, PGD^{+}97, PWW98, PFDL97, PYL98, SCM^{+}97, SYF^{+}97, SHL97, SV98, UvMJ^{+}99, WSR^{+}97, WE98]. Established [KRK^{+}98]. establishes [SYB95]. Establishment [LTR^{+}97, MRL^{+}99, ASP^{+}97, BMHH97, DWM^{+}98]. Esters [SOPM^{+}97]. estrogen [FSR^{+}96, ARR^{+}98]. estrogen-dependent [FSR^{+}96]. Ethylene [SRR^{+}97]. Eukaryote [HMCL97]. eukaryotic [SCKG95, VSGF96, SHMC99]. EVA [GSP^{+}98]. event [ZMG^{+}96a, MDE^{+}99]. Events [KWD^{+}98, AW96, BBZ^{+}95, BSD^{+}96, JCPK95, PCL^{+}96, PLZ^{+}95, RPW96, AK97a, HCRS98, NG99, PHN^{+}98, TMM97]. Evidence [AN99, CRLM96, CAH^{+}99, DWM^{+}98, FWL^{+}98, FNH99, GM95, HMD97, KBW^{+}96, MHM99, PT97, SJW97, TFYH99, ÜHK^{+}99, YvdEH^{+}95, CMWL96, CPL^{+}95, DG96, FHCL95, FFMG95, GGC96, HBJ95, JMTCF96, KAAFZ95, LMMO96, MWOB96, MG96, RZZ96, ALM97b, BR98, Dav95, GBvdN^{+}98, JP98, JR98, KBF^{+}98, KKT^{+}98, LAS^{+}97, MLN^{+}98, OTL^{+}97, OF98, SB99c, WNB97]. Evoked [PMP^{+}98]. Evolution [RPBB98]. Evolutionarily [WDV^{+}97]. Evolutionary [HCM^{+}97]. examined [KE98]. Exchange [TXB^{+}99, WSCN97, FWG^{+}96, JLRS95, GKP^{+}97, KR99, vLKvdK^{+}97, vHMV^{+}97]. Exchanger [HCPW^{+}99, WSG^{+}98]. Exchangers [AKCC99]. exchanging [BCC^{+}95]. excitation [SPT^{+}95]. excitation-contraction [SPT^{+}95]. Exclusion [RR96, PHB96]. execute [DPH95]. execution [YGK96b, MCS^{+}97a, STK^{+}98, SGSW97]. Exencephaly [MCS98]. exert [KR96, BRD99]. exhibit [AYF96, FGA^{+}96, NPRT95, RFNF96, CGSY^{+}99]. exhibits [FaAR^{+}96]. Exist [TLF^{+}97]. Existence [MHM99, HD99]. Exit [BMG^{+}99b, CSN99, LCP95, CHC97, JMS97, LPD^{+}99, TGN97, VPM^{+}98]. exocrine [DMP^{+}96, CFB^{+}98]. Exocytic [RWOA97, BM99, SVF^{+}98]. Exocytosis [BML^{+}97, CT97a, BAS95, CRMB95, CGVS^{+}96, MST^{+}96, MM95b, MKXY95, VBZ96, WTCV95, BMB97, BL99, EKF^{+}97, FN98, LRAB99, NBJ^{+}99, PAN97, ZCOP99]. Exocytosis-sensitive [CT97a]. exocytotic [LBL95, TKY^{+}97]. Exogenous [OOB99, vdBCH^{+}95, OMS97]. Exon [FSDA97, AK95, BJS^{+}95, JBD^{+}95]. exons [BMG^{+}95, BRJP96]. Exoribonuclease [BSS^{+}97]. Exosomes [TRG^{+}99]. exotoxin [FFZ^{+}95]. exotoxin-mediated [FFZ^{+}95]. expanded [RvSL^{+}95]. expansion [BLB^{+}96, DDM98, RS99]. experimental [ZN95a, SEH^{+}97a]. Explanation [SWR^{+}98]. Exploits [RBM^{+}99]. Export [ABK^{+}97, GLL^{+}99, BRB96, BLEB96, CCBK96, GDC95, HB95b, MKB^{+}95, ND96, RLCM96, VIF^{+}96, AWB^{+}98, BBP^{+}98, BF97a, HHS^{+}99, IW97, IJB^{+}97, KDG98, KDK^{+}99, PFDL97, PFKvD99, RCE^{+}99, TDR99, YLD97]. Exposed [DSM^{+}98a]. Exposes [ZCWB^{+}98]. express [VGM96, GJW^{+}97, ZhTA^{+}97]. Expressed [AAM^{+}98, CBM^{+}98, OZF^{+}99, WLD^{+}98, AC95b, BSJJ95, FAP^{+}96, GCM^{+}95, HKL96, HKM^{+}96, HWPC96, PJL^{+}95, RGS^{+}96, SLM^{+}95, SCKG95, VDA^{+}96, WYY95, WKRB96, BHBB99, GSP^{+}98, PDS^{+}97a, WDL^{+}98, YSEI^{+}99]. Expressing [CSF^{+}97, AYF96, CRM^{+}96, DNC^{+}95, FWHW96, KGW^{+}95, KMOO95, ABL98, GBW98, SATW98, SCB^{+}97]. Expression [ADP^{+}95, BBW^{+}98, BBAN^{+}99, BRHD95, FSA^{+}99, GRB^{+}97, GXE^{+}99, HDV^{+}99, HMB95b, ICW^{+}96, KSP95, KOA^{+}99, LMMO96, LRS^{+}98, MKB^{+}95, MLLT95, OTL^{+}97, SJF^{+}97, SL95, VKP^{+}95, XC97, ZOS^{+}97, ACHG95, AW96, BWF^{+}96, BEG^{+}95, BBOB96, BSK96, BGGR^{+}96, CLT^{+}95, CLL95, DNB^{+}96, DSvdV^{+}96, EMB^{+}96, ED96, FFZ^{+}95, FHV^{+}96, GZK^{+}95, GEJ^{+}95, HHM^{+}95, HGP^{+}95, KKAN^{+}96, KAT^{+}95, KBM^{+}95, LRM^{+}95, LGdC95, LDCB95, MAFJ^{+}95, MF96b, MVvdBD96, NKI^{+}95, NMJL96, OSPJ^{+}96, PLM^{+}96, RCC^{+}96, RL96, RME^{+}95, RZZ96, RK95, REdC95, SRR96a, SPMB^{+}96, SMRT96, SSM^{+}95, SSA^{+}96, SHPP96, SAPLHD95, SSB^{+}95, TCM96, TLF^{+}95, TDW95, TRJ^{+}95, TKM^{+}96, WPS^{+}96, WLML95, WJC^{+}96, WMC^{+}95b, WD95, XZSC96, YKC95, ZMM^{+}95, ZHR95, ZL95c, ZOB^{+}96, AEWG^{+}97, BZBA99, CSEM99, DCJ97, DDM98, GvdFvD^{+}99, GYRWR97, HHB^{+}97, JM97a, JLID98, JDK97, JN98, KEJ97, KCV^{+}98, KFL^{+}98]. Expression [KGHG98, LNH^{+}98, LMP98, MBTW97, MFBK99, ML97, MC98, MMM^{+}98, MPL^{+}97, MNG^{+}98, NTAN98, NPJW99, NMNL98, OKE^{+}99, OOB99, PC98, PGD^{+}97, RYB^{+}99, RB98, RFK^{+}97, SPR^{+}98, SW98, SJZW97, SEAB^{+}99, TDR^{+}98, TWiK^{+}99, TPM^{+}99, VHF97, WMSM97, WRG98, cXP97, XZSC98, ZTH^{+}97]. extend [PNM95, VLZR96]. extends [MR95]. Extensibility [TGL^{+}98]. Extension [MM99a, DC96b, DGSR95, MAS^{+}95, HJJJ97, MQR^{+}97]. Extensive [WB99, MMS^{+}96b, ZL95a]. extent [VBZ96]. external [MBS96a]. Externalization [NMNL98]. Extracellular [FLO^{+}99, FWBSO97, IAS^{+}98, LDCB95, NMK^{+}97, XC96, AC95b, DMS^{+}95, DGSR95, HMB95a, HH95b, KLMR95, LF95, RSC96, RS95, WHH96, WvdVV^{+}95, ZHR95, dPSMNSM95, APF^{+}97, AMY^{+}98, BZBA99, FZZN98, GCAHW97, GLMG98, LGZ^{+}99, LDD^{+}97, MH97, MLC99, SLN^{+}97, SSSS97, VLM^{+}98, WLF99, WG97, ZOKS99]. Extrachromosomally [SIUW98]. Extract [DMMS98]. extracts [LKWG96, WSGPS95, CLL^{+}98, GHW99, HTH^{+}97, KGE^{+}98, MCJK98, TGN97, WVM97, ZJY^{+}98]. Extranuclear [MSP99]. Extrasynaptic [CAS97]. Extreme [FFDP98]. Extrusion [SGD98]. Eye [BM97]. Eyes [RCD99]. Ezrin [BGB95b, BFRB99, CGC^{+}97b, KTD^{+}97, MEFH97, MAS^{+}95, MMD^{+}98, YTT99, HSK^{+}96, SH95, RBB97, YHD^{+}98]. Ezrin/ [YHD^{+}98]. Ezrin/Radixin/Moesin [KTD^{+}97, MEFH97, MMD^{+}98, YTT99, HSK^{+}96].

F [LWM^{+}95, CSY98, FCTPA^{+}99, GFN^{+}99, HHR^{+}98, LTE^{+}96, MPCW97, MH95b, PPWL97, RGS99, RAE96, SH95, SRRB95, TTG95, TCSG96, TLBK98, ÜHK^{+}99]. F-actin [LTE^{+}96, MH95b, RAE96, SH95, TCSG96, ÜHK^{+}99]. F-Actin-binding [PPWL97]. F-actin-rich [SRRB95]. F-box [RGS99]. F11 [NHB^{+}95, VLZR96, VZNR98]. F11-mediated [NHB^{+}95]. F9 [CCMC97]. Fab1p [GWB^{+}98]. Face [BdPE^{+}97]. faces [ZM95]. Facilitate [AvdG99]. facilitated [GK96, BCAS98]. Facilitates [YHS^{+}99, RPW96, ZL95c, KRH98, TKBM97]. Facilitation [MZL97]. Factor [AGS^{+}97, BMC^{+}99, BHD^{+}97, CLS^{+}97, DCVF98, FKNM97, JCR97, NAsG^{+}98, NCV^{+}98, RAA^{+}97, SB99d, TXB^{+}99, TB99, TSC^{+}98, AOMB96, AM95b, BSN95, BJS^{+}95, BKP^{+}95, BFS^{+}95, BGN^{+}96, CGSP95, CMUW96, CAA95, DMS^{+}95, DPC^{+}96, DHT96, EHM^{+}95, EB96, FLW^{+}95, GJ95a, GHP^{+}96, GKM^{+}96, GBBSO96, HIG^{+}95, HWFA96, IWW95, JSR96, JLRS95, KBW^{+}96, LDF^{+}95a, LS95a, LLN^{+}95, LTE^{+}96, Mah96, MBD^{+}96, MMMT95, MELC96, NIM95, OSH^{+}96, PSBB96, PG95, RVD^{+}95, RROA95, RD96b, SRR96a, SLT^{+}96, SS95a, SCG^{+}96, SSM^{+}95, SRR96b, SVD96, TIJ^{+}95, TLF^{+}95, TRJ^{+}95, VTG^{+}96, VBG96, WMC^{+}95a, WKJH^{+}95, WD95, YN95, YSM^{+}95, YJB^{+}95, ZMM^{+}95, ZM95, ZSMV95, APRB98, ALK^{+}99, AK97b, BKJL^{+}98, Car99, CSCM97, CEZA97, CGC^{+}97b, DPM^{+}97, EWW^{+}98, FHK97, FWBSO97, FJ98, GMGGD98, GKP^{+}97, GWLH98, GGPG99, GTY98, HTWC98, HZR98, IW97, JLID98]. Factor [KWN^{+}99, LNM^{+}98, LBBP98, LSZ^{+}99a, LMP98, MGB^{+}97, MBTW97, MSH^{+}98, MFBK99, MRM^{+}98, NGMR97, ODB98, OFM^{+}98, PHN^{+}98, PFKvD99, RD97, RPS99, RSD98, SYT97, SPR^{+}98, SBPR^{+}98, SKS99, SRA^{+}99, SGGW99, SW98, SV98, TNFK^{+}98, TWSD98, VHF97, WSSB99, WDL^{+}98, WBR97, WWO^{+}98, vLKvdK^{+}97]. Factor- [MGB^{+}97, NGMR97, SBPR^{+}98]. Factor-1 [GMGGD98]. Factor-2 [JLID98, SPR^{+}98]. factor-alpha [AM95b, MBD^{+}96]. factor-beta [DMS^{+}95, EHM^{+}95, ZMM^{+}95]. Factor-dependent [BHD^{+}97, TSC^{+}98, JSR96, KBW^{+}96, SRR96b]. Factor-II [NCV^{+}98]. factor-immunoglobulin [LDF^{+}95a]. factor-induced [RVD^{+}95]. factor-stimulated [DPC^{+}96]. Factor/Cofilin [RAA^{+}97, SB99d, AOMB96]. Factor/Scatter [AGS^{+}97, BFS^{+}95, TIJ^{+}95, WKJH^{+}95]. Factors [WLD^{+}98, YTO^{+}99, BJB^{+}96, DEG^{+}96, HKM^{+}96, HS96, KWCS96, MTGY96, NSR95, SS95a, TYSP^{+}96, BZBA99, BM98, JVRNM99, KDG98, KB99a, KDBW97, KCV^{+}98, KTYT99, Ped98, SVB^{+}97, SMG^{+}97]. FADD [PW98]. FADD-dependent [PW98]. Fail [BBL^{+}99a]. Failure [SVM^{+}99]. faithful [YGK96b]. FAK [LAS^{+}97, YTY^{+}99, GTP^{+}99]. FAK/p130 [YTY^{+}99]. falciparum [BJR^{+}99, GDF^{+}96, WSG^{+}98]. Fam [TYK^{+}98]. Families [DDWM97]. Family [BDO^{+}97, EGKC^{+}99, HB99, KGF^{+}97, KNH^{+}97, KND^{+}97, LLH^{+}99, OBB99, OZF^{+}99, PdHTV99, BOLS95, CBB^{+}96, CK96, DVD^{+}96, ES96, FUB^{+}96, FSAP95, GKR95, JWR96, KKFN^{+}95, LVD96, LFB96, NSR95, PJL^{+}95, SSL^{+}95, SLP^{+}96, TAK95, WPMW95, ZL95b, AGP^{+}98, CCS^{+}98, CF99, CKB^{+}98, DDF^{+}98, FCMN97, GRTB97, GvdFvD^{+}99, HGW^{+}98, HHM^{+}98, LL98a, NLM^{+}98, PRV98, RBB97, RHPW97, SVB^{+}97, SEBW^{+}98, SSH^{+}98, YPM97]. far [BGGR^{+}96]. far-upstream [BGGR^{+}96]. FAR1 [VPH95]. Far1p [NA99]. Fas [FHV^{+}96, GSvR^{+}96, TMN97, CIK^{+}99, RHP99, AKM^{+}98, FHV^{+}96, KOM^{+}98, PW98, SKC97]. Fas-associated [KOM^{+}98]. Fas-induced [SKC97]. Fas-mediated [PW98]. Fas/APO [AKM^{+}98]. Fas/APO-1 [AKM^{+}98]. Fascicle [HJJJ97]. fasciculata [JE98]. fasciculating [BSJJ95]. fasciculation [BOLS95, KZKS96, KSG^{+}98, LDP^{+}99]. fascin [TET^{+}96]. fatal [WNB^{+}95]. Fate [HRD98, WBKB97, HSL^{+}95, MLLT95]. fates [OS96a]. fatty [STN^{+}96, DSG98, HR99, LHS97a, SKY^{+}98, SBS^{+}98]. FBR [JH96]. Fc [GSB97, PHB96, SHB99]. Fcgamma [FWG^{+}96]. Fe [GYRWR97]. feature [LNR^{+}96]. features [GLS96, SW95, KKM^{+}97, LRG^{+}98, NKK^{+}99, TWP98]. Feedback [BRD99]. Female [GCA^{+}98, WHB^{+}99]. Females [LARH97]. fence [SK95a]. Fenestrations [EWW^{+}98]. Ferriprotoporphyrin [BJR^{+}99]. fertilization [AZF96, CRM^{+}97, WFS97]. fetal [DvD95, ISFB^{+}96, KBSW95, PBB^{+}96]. FGF [Mah96, OST^{+}95, SPR^{+}98, GHP^{+}96, HMB95a, KJR^{+}98, LPM^{+}98, Mah96, HKJ^{+}99]. FGF- [GHP^{+}96]. FGF-1 [LPM^{+}98]. FGF-2 [SPR^{+}98, HMB95a, Mah96]. FGF-4 [OST^{+}95]. fgr [LFB96]. FH3 [PNEH98]. Fiber [BM97, IJR98, MHC^{+}97, NJB^{+}98]. Fibers [BMG^{+}99b, CWB96, FCTS99, KMK^{+}98, PvDA^{+}98, SCB^{+}97, YAC97, YPB^{+}98]. fibril [MHLB96, CML^{+}98, DBH^{+}97]. Fibrillar [NMK^{+}97]. Fibrillary [PJE^{+}99, GCGE^{+}96]. fibrillin [ZHR95, KRS^{+}98b, WPS^{+}98]. Fibrillin-1 [KRS^{+}98b]. Fibrillin-rich [WPS^{+}98]. fibrillogenesis [MHLB96, KZS^{+}98]. Fibrils [HBH98]. Fibrin [CCB^{+}95, KvEdV^{+}96, LCB^{+}99]. Fibroblast [MBTW97, SPR^{+}98, BKP^{+}95, CH95, GHP^{+}96, HKM^{+}96, HG95a, Mah96, SVD96, SOL^{+}95, VTG^{+}96, WLML95, XC96]. fibroblastic [ICW^{+}96]. Fibroblasts [MEFH97, RWOA97, GG95b, HKS^{+}96, HHM^{+}95, HCM^{+}96b, PAV^{+}96, RJP96b, SR96, VSB95, ALCC^{+}99, ABL98, BKJL^{+}98, CSM97, FSFT98, HKBM99, LAS^{+}97, MSB^{+}97, NJB^{+}98, SBC99, XZSC98]. Fibronectin [BOM^{+}98, HSML98, KGF^{+}97, BEG^{+}95, HSML96, HHM^{+}95, RD96a, STS96, WLG^{+}96, iMOeM^{+}97, SBPR^{+}98, ZCWB^{+}98]. fibronectin-specific [BEG^{+}95]. Fibrosis [JMD^{+}98]. Filagenin [LBO^{+}98]. Filament [CLS^{+}97, FWL^{+}98, GCST97, MAM99, RdPRW98, SB99d, The97, WDV^{+}97, BWKH96, BCSG96, COML95, GF95, MMP^{+}95, MKXY95, MELC96, OS96b, SHA95, TTG95, TTG96, VSB95, XHUC96, ASP^{+}97, Bri99, CSM97, FRHS98, GFN^{+}99, LSL99, LWS99, LRC^{+}99, MNS^{+}97, MPCW97, NOS^{+}97, NLM^{+}98, ONI^{+}98, PYM^{+}98, RSS98, SWS97a, STF98, SGA97, TWP98, YMPG98]. Filament-binding [ONI^{+}98]. filamentous [RJP96b, LL98b]. Filaments [BMC^{+}99, BGH^{+}97, MEFH97, RAA^{+}97, ASY96, CK96, CRCH95, FaAR^{+}96, GKC^{+}96, GGR^{+}96, GPRS^{+}95, GG95b, KTC95, MELC96, NCZ^{+}95, PTBC96, SHC95, SVB96, TCSG96, CSO97, CRRF97, GCML98, GWB^{+}99, INMT97, JOPM98, MSL^{+}98a, RTL^{+}99, Sal99, SMZ^{+}98, SF98, YAiN^{+}98]. Filamin [gLSE^{+}99]. Filensin [GGR^{+}96]. Filipin [OF98]. Filipin-dependent [OF98]. Filling [HFP98a]. Filopodia [mLZB99, LWKK97]. filopodial [SSW^{+}96]. Filopodium [MM99a]. fimbrin [KTC95, HMD97]. final [MKXY95]. Finger [HDV^{+}99, DSVL^{+}98, RHA^{+}97, SYT97]. Finish [Ste97]. Firing [ZLM^{+}98]. First [BMG^{+}99b, FKNM97, HB98, JRLS95, KJBM^{+}96, STS96, LHS97a, MCM^{+}99, VL99]. Fish [BOLS95, SHA95, SVMB97]. Fission [BHJ98, BG97, KHA^{+}99, VMN95, HY95, LW95, STN^{+}96, SMC^{+}96, BSW^{+}98, BHHF97, BSY97, CDN97, DS98b, HBS^{+}97, OMS98, PNEH98, SN98, THS^{+}98, YWMH99]. Five [WC97b]. fizzy [DRAT95]. FK506 [DBOM98, WGF^{+}98]. FK506-binding [DBOM98]. FKBP13 [WGF^{+}98]. FKBP65 [DBOM98]. FLA10 [PMH96, KBR95]. flagella [PMH96, SL96, HPKL97]. flagellar [KBR95, SM95b, HS97, HOCK98, LXC^{+}98, MS99, MKW^{+}99, PDW99]. Flagellate [GCML98]. Flagellum [WWM^{+}98]. Flexibility [BWWT97, BSAJ97]. Flexible [MCBE98]. flexneri [AAP^{+}95, BGL^{+}99, ELL^{+}99]. Flexural [FFB^{+}97]. FLICE [PW98]. flight [KSM96, RV98, SRR^{+}97]. flow [AKC^{+}95, AFFS96, Cun95, FAP^{+}96, SSEM98, WSS97]. Fluctuations [TLBK98]. Fluid [JDK97, LKKL97]. Fluorescence [SLPE^{+}98, GJ95a, KE98, WMXE96, CWT^{+}98]. fluorescence-activated [WMXE96]. Fluorescent [ACSN98, CRLM96, HP95, JCWP95, OMO95, GGPG99, KGA^{+}97, KFL^{+}98, LHS97a, OTQM98, PÖS^{+}98, SSI^{+}97]. Fluorescent-labeled [KGA^{+}97]. fluoride [RKD96]. Flux [DMMS98]. Fly [MK95b, GCD97, RFK^{+}97]. fmk [LNM^{+}98]. FNIII [DLB96]. Focal [CHP^{+}98, MEFH97, RPS99, SOHP95, BPT96, CWB96, FVRCH96, GPRS^{+}95, LCS96, ALCC^{+}99, CLRR99, DLC98, IAS^{+}98, KRS98a, NB97, NJB^{+}98, OWD^{+}99, PHM^{+}98, SLW^{+}99, cXP97, ZRG98]. Foci [TMS^{+}95]. focus [OCS^{+}95]. folate [HSWC96, RJP^{+}96a, SMA96]. folate-analogue [HSWC96]. Folded [MCBE98, TGL^{+}98]. Folding [BCAS98, KA95, LTVC96, SSY^{+}96, SFNRH95, HCW99, HZC^{+}97, HVL^{+}99, HTW97, MZL97, SSSG98]. follicle [APG^{+}96]. follicles [TTG96]. following [CGSP95]. force [KCBR96, KR96, CSM97]. Forced [RCC^{+}96, WLML95, WRG98]. forces [NWG95, GS99, YKO97]. Forest [BWG95, KKGD96, VPM^{+}98]. Forked [BZL^{+}98]. form [BK96, BKB^{+}96, CK96, GH95, GGR^{+}96, LZC^{+}95, NIM95, NWPW95, RHHRB96, WHH96, BHHF97, DZS^{+}98, GWLH98, GBW98, OMS98, SSW^{+}98, SBGR98, TMPM99, UNPW98, YPB^{+}98, ZDK^{+}99]. Formation [AS96, BMG^{+}99b, GMBSS99, KBI^{+}99, KRK^{+}98, MBA97, PPF96, SPK^{+}97, SBG^{+}98, TTG96, AMJM95, All95, AGPG96, BSK96, CCB^{+}95, CWB96, DLGB95, DYF96, EELF^{+}95, ETF95, FFGG96, FGP96, FBH^{+}95, GACM^{+}95, GEK95, HY95, KvEdV^{+}96, KBW^{+}96, LCH96, MMATe96, MMF95, MW96, NCZ^{+}95, RRRA95, RROA95, SGS95, SK96, SGGS96, TSS^{+}95a, VSB95, ZDT^{+}96, ZLC^{+}96, ZZvdB^{+}96, ACH^{+}98, CFL^{+}98, CT98, DP98, EAM^{+}97, FMP^{+}97, FAZ^{+}99, GHW99, GZC^{+}99, HPKL97, HJJJ97, IKR^{+}98, ISN^{+}99, JH97, KSY^{+}98, LPD^{+}99, LH97, LMC^{+}98, MSL^{+}98a, MKC99, MC97, NMK^{+}97, NOS^{+}97, Nei98, NJB^{+}98, OSM98, PKQ^{+}98, RM97, RME98, SFS^{+}97, SVB^{+}97, SKS99, TGW99, TCV^{+}98, VZNR98, WKR^{+}98, YSK^{+}97, ZRHO99, ZS98]. Formed [SFR^{+}98]. Formin [HCAC99, LKE^{+}99, PNEH98]. forming [PHB96, WSW95, SPR^{+}98, AFG^{+}98]. Formins [PNEH98]. Forms [MM99b, BKP^{+}95, BSMH95, FaAR^{+}96, JBD^{+}95, PJL^{+}95, PZ96, RDM^{+}96, TFF^{+}95, VTG^{+}96, GCML98, IMK99, RFK^{+}97, RHPW97, RGS99, SN97, ZBW^{+}97]. Forward [ISR99]. fos [JH96]. Found [HGW^{+}98, PNEH98]. Four [MCS^{+}97a]. fpr [BTTB96]. fraction [GEW^{+}95, FCMN97]. fractionation [BMS95]. Fractions [HFP^{+}98b]. fragility [KMOO95, PLM^{+}96, CML^{+}98, DBH^{+}97]. fragment [CDRB^{+}96, MAT^{+}98, SFS^{+}99, YGG97b]. Fragmentation [DFD^{+}99, MW95b]. fragments [EMB^{+}96, RMWW95, CLRR99, KCM^{+}97]. Frame [RKB97]. Free [SW99, VG99, CM95a, DKD^{+}96, HW98, MW95b, RMWW95, SSP95, GLL^{+}99, NG99, SPBCM97, SNT^{+}98, KGE^{+}98, LHWH97, UPT98, ZJB^{+}97]. Freezing [KKM^{+}97]. Frog [RRA97]. Fructose [CC98]. Fructose-1 [CC98]. frustrated [DSGF95]. FSP1 [SOL^{+}95]. FtsZ [ES96]. fucosylated [FAP^{+}96]. fucosyltransferase [KCN^{+}96]. FucT [KCN^{+}96]. Full [MBOC97]. Fully [MAM99]. Function [BSW^{+}99, FWL^{+}98, IMW^{+}99, NAsG^{+}98, SGTH99, SB99c, VS98, BCC^{+}95, BLEB96, Bor95, BRHD95, CRLM96, CS96b, DK96, EPVV96, GLD^{+}96, GKC^{+}96, GMFN95, GG95b, HG95b, HCW96, IFHR^{+}96, KSM96, LYC^{+}95, MAG^{+}96, MJMS96, MBB96, MGS^{+}95, MBF^{+}95a, MTC^{+}95, OMO95, OP96a, PWE95, PLZ^{+}95, PMCT96, RCC^{+}96, RRRA95, RS95, RJP96b, SHR^{+}95, SJHC96, VJHR96, vABSP95, ANS^{+}97, ANRT99, BRP^{+}97, BBL^{+}99b, BKM^{+}98, BF97b, CBS^{+}97, hCMPG97, CIRG99, CIK^{+}99, DSG^{+}97, DFD^{+}97, DKN^{+}97, DAB99, DBJDT99, FKM^{+}99, FY97, GvdFvD^{+}99, GZC^{+}99, HT98, HMC^{+}99, HDDKH98, HCG^{+}98, KMMC98, KHA^{+}99, KPZ^{+}97, LLVS98, LRAB99, LL98a, LBB97, LFP98, LVN^{+}98, MGIZ98, MCTMK98, MPCW97, MP98, MNG^{+}98, OSR^{+}98, OBW98, PMCS97, PH97b, SHC^{+}97, SCM^{+}97, SMCE97, TRA^{+}99, VMH97, WSR^{+}97, WBR^{+}98, vHMV^{+}97]. Function-associated [PH97b]. Functional [BWF^{+}96, BGBS98, BBP^{+}98, BHP^{+}95, DBE98, GAPG97, GYRWR97, HPY^{+}98, IIM^{+}99, KKG95, KFL^{+}98, KMK^{+}98, LCD^{+}98, PCD^{+}98, WPS^{+}96, WSSM95, WLF98, AAP^{+}95, ARM^{+}95, ADP^{+}95, ASSS^{+}95, BFL95, CSDR96, GBG^{+}96, HMB95b, HGG96, LDF^{+}95a, LN96, MS96, NHB^{+}95, PKR95, RPC^{+}95, SSL^{+}95, BGH^{+}98, CBTZ98, DW97, GBvdN^{+}98, GSJ98, HTWC98, JSN98, KIH^{+}97, LMF98, LMM^{+}99, LRG^{+}98, LHR^{+}97, LLSO98, OBW98, PC99, SZBB98, TPM^{+}99]. Functionality [MS98]. Functionally [CWSL97, ABRB^{+}95b, HBJ95, SG95, ZDT^{+}96, BG97, DFL^{+}99, FNDS99, SLAS^{+}98]. Functions [CJS^{+}99, CCMC97, ESH^{+}99, NCW^{+}99, DBBB95, EKK^{+}96, FK95, HAS95, HUU96, JWH96, KM95, LFB96, MPY^{+}96, MF96b, NCZ^{+}95, OP96c, RBL^{+}96, SOR^{+}95, SM95a, WPMW95, YRM^{+}96, ABT^{+}98, FLA99, HC99, HPS98, HLG^{+}98, IYM98, RM97, SZFM98, SDHM99, SY99, SBD^{+}99, WRC^{+}99, WCW98, WUUI^{+}98, WE98, ZKS^{+}99]. Furin [MM99b, WBKB97, LTW^{+}97]. Furrow [EAM^{+}97, GDTBD98, SDC^{+}97]. furrows [MKI^{+}96, ZLC^{+}96]. Further [MMK^{+}97]. Fus1 [PNEH98]. Fus2 [ETF95]. Fus2p [BGR98]. fuse [DFMG97, FPHH96]. fusiformis [WFBC96]. Fusion [CMR^{+}97, CSF^{+}97, MBF^{+}98, RST^{+}99, AW96, BL96, BSD^{+}96, BGN^{+}96, DPHW96, EM96, ETF95, GTL^{+}96, KKGD96, LCH96, LBL95, LHS96, NW96, PZ96, SL95, VBZ96, ZMG^{+}96a, BBR97, BGR98, CLF^{+}97a, CFL^{+}98, CBTZ98, DFMG97, EM98, EGA97, HTM^{+}98, HPD^{+}97, KKM^{+}97, LMA^{+}97, MPSC98, MBOC97, MBDS^{+}98, PGD^{+}97, PH97a, PH98, QPH^{+}98, SSI^{+}97, SJ99, TH99, THS^{+}98, TMM97, UNPW98, UvMJ^{+}99, UPT98, WLK97, WFS97, YNM^{+}99, YPM97]. Fusogenic [DFD^{+}99]. Fyn [OWW^{+}99, YYH^{+}96, ZDK^{+}99]. Fzo1p [HTM^{+}98, SJ99].

G [KBS^{+}99, NA99, PBJ^{+}98, ALM97b, CMS98, DMP^{+}96, FUB^{+}96, FAZ^{+}99, GHJJ95, GZC^{+}99, HPL99, HMB95b, KBS^{+}99, MXSRB96, OCS^{+}95, PHA^{+}98, RLVM98, RP97, RRH^{+}95, TSK^{+}97, Val96, VLM^{+}98, WVVD95]. G-protein [GHJJ95]. G-spectrin [HMB95b]. G0 [YvdEH^{+}95]. G0/G1 [YvdEH^{+}95]. G1 [BZBA99, DTMG99, OWD^{+}99, TOB^{+}99, TALM96, YvdEH^{+}95]. G1-Phase [DTMG99]. G1-specific [TOB^{+}99]. G2 [FK95, JGM96, LWM^{+}95, NMJL96, TXB^{+}99, ZKSB96]. G2/M [TXB^{+}99, ZKSB96]. G4 [BSS^{+}97]. G93A [ZhTA^{+}97]. Gaa1p [HER95]. gag [SHG96]. gaining [MGS96]. galactosidase [SL95]. galactosphingolipids [vdBLCvM96]. galactosylceramide [vGvM95]. galactosyltransferase [MR96]. galactosyltransferases [BvdBvM96]. GalNAc [HHRdB^{+}98]. GalNAc- [HHRdB^{+}98]. GAM [IKR^{+}98, MN98]. Gamete [TGTL97, KGS98]. Gamete-specific [KGS98]. Gametes [WFS97]. gametogenesis [GDF^{+}96, GCA^{+}98]. Gamma [SJ95, DMP^{+}96, DNH^{+}96, ES96, HSL^{+}95, IMFS96, LJ95, LD95, MF96a, MJMS96, ROG^{+}96, SS95b, SGGS96]. gamma-dynein [ROG^{+}96]. Gamma-tubulin [SJ95, LJ95, SS95b]. gamma-Tubulin-like [SGGS96, MJMS96]. ganglion [BOLS95, ULMT99]. Ganglioside [WJWM^{+}98]. Gap [HCW^{+}98, JGCS97, LCK95, EELF^{+}95, KGW^{+}95, LLL^{+}96, SL95, AC95a, TBP^{+}99, BGBS98, FZZN98, GCE^{+}97, LNG^{+}98, MBTW97, qXLCH97, ZKN99]. GAP-43-depleted [AC95a]. gastrulation [RD96a]. gated [GJ95b, WWM^{+}98, WS98, ZLM^{+}98]. Gating [OSM98, ZKN99]. Gbf1 [CZK^{+}99]. GDNF [TWSD98, TRJ^{+}95]. GDP [GKP^{+}97]. GDP/GTP [GKP^{+}97]. Gelatinase [OTL^{+}97]. Gelsolin [qSmLY97, BWKH96, CMUW96, LWKK97, PPWL97]. Gemin3 [CPP^{+}99]. Gene [BBL^{+}99a, GTL^{+}96, HDV^{+}99, HT98, KMI^{+}97, MMN^{+}99, MAB98, PLM^{+}96, TOSF^{+}98, ACHG95, BTTB96, BBOB96, BGGR^{+}96, CF95, CKM95, DRAT95, EMB^{+}96, EGH^{+}95, FPM^{+}95, GKVR95, GEJ^{+}95, GAW^{+}96, GBW^{+}95, HTS^{+}95, HY95, HCA^{+}95, HWPC96, HWC^{+}96, HHM^{+}95, IFHR^{+}96, KP96, KBSW95, LRM^{+}95, LGdC95, LDCB95, MAG^{+}96, MKB^{+}95, MBP^{+}95, MR96, MMS96a, NMJL96, OFY^{+}95, OABD^{+}96, PJL^{+}95, PGC^{+}95, RK95, SSA^{+}96, SCKG95, SBR^{+}96, SS95b, SSB^{+}95, SVT96, TWVC95, TCM96, TSS^{+}95b, TDW95, TKM^{+}96, VKP^{+}95, WS95b, WW96a, WD95, XJM^{+}95, XZSC96, ZLC^{+}96, AEWG^{+}97, BAH^{+}98, CAC^{+}99, DDWM97, ESR97, FH97, FSFT98, GCA^{+}98, GPSF99, GKS^{+}98, GB98, HKS97, IFSV^{+}98, IRW^{+}99, KMHK97, KSY97, KPW^{+}97, LBWF^{+}97, LNH^{+}98, LDD^{+}97, ML97, MMM^{+}98, NTAN98, PB98, PHM^{+}98, RB98, RKB97, RLMC99, SdCAH^{+}97, SGSW97, SMCE97, SEH^{+}97a, SCEB99]. Gene [SMW^{+}99, SVM^{+}99, TKH^{+}97, THS^{+}98, TWiK^{+}99, WRC^{+}99, WMSM97, XZSC98, ZTH^{+}97, ZLL^{+}98]. gene-1 [HWPC96]. gene-encoded [MKB^{+}95]. gene-related [CKM95]. general [LCH96]. generated [NH95, TLF^{+}97]. generates [CDRB^{+}96, ZL95a, FBM99]. generating [MSF^{+}95]. generation [FJT^{+}95, KJBM^{+}96, SKWY95, BCWA97, CSM97]. genes [BIR96, BGGR^{+}96, DG96, GCM^{+}95, KLN^{+}96, NKI^{+}95, OMW^{+}96, REdC95, VMN95, ZHR95, ASES^{+}97, BGN^{+}97, ELMS98, JVRNM99, jLC97, SW98, TRA^{+}99]. Genetic [CRCH95, DVD^{+}96, DFMG97, FKM^{+}99, FNH99, YRM^{+}96, FHCL95, HBJ95, MGA^{+}96, FRT^{+}98]. genetically [NRM^{+}96, CMR^{+}97]. genome [MW95a, LGM98]. Genomes [SLF^{+}99]. Genomic [TWiK^{+}99, dBCKS98]. germ [OABD^{+}96, GCAHW97]. GFP [OTQM98, OMO95, YNM^{+}99, GMBSS99, RST^{+}99]. Gfp-Fusion [RST^{+}99]. GI [WLF99]. Giant [EB95, BTTB96, LRAB99, HBPJ98]. Giantin [SLL^{+}98]. Gin4 [AK97a, CAS^{+}98]. Gin4p [LFP98]. Gives [PHN^{+}98]. GKAP [KNH^{+}97]. Gla [YSEI^{+}99]. gland [FMJ^{+}95, KKD^{+}95, YSM^{+}95, JWFS99, NBS^{+}98, RFCR^{+}97, YHMS^{+}98]. GLAST [WS96]. GLAST-1 [WS96]. GLC [LCP95, FFDP98]. GlcNAc [DNC^{+}95]. GlcNAc-transferase [DNC^{+}95]. GLEBS [PFKvD99]. GLEBS-like [PFKvD99]. GLFG [IWW95, IW97, PMMF95, PFDL97]. glia [BHP^{+}95, SPB^{+}95, BHBB99]. Glial [DGP99, PJE^{+}99, FJT^{+}95, GCGE^{+}96, WPS^{+}96, GCE^{+}97, SLN^{+}97, TWSD98, YAiN^{+}98]. Gliding [KBG^{+}99]. Glimpse [VL99]. Glioma [BPS99, LHR^{+}97]. Globular [CDH^{+}99]. GLP [WG99]. GLP-1 [WG99]. Glu [MBS96a]. glucan [LMB^{+}95, KHA^{+}99]. Glucocorticoid [FMJ^{+}95, CRLM96, SHMC99, TRA^{+}99, YLD97]. glucocorticoids [COML95]. Glucose [SGM^{+}97, HLSM95, LCP95, MTL^{+}96, PXR^{+}95, VYB95, WHM^{+}96, GJB^{+}98, HOS^{+}97, SL97]. Glucosylceramide [vIZKH97, vGvM95]. Glucosyltransferase [FFDP98, LCP95]. Glued [MGS^{+}95]. GLUT [MTL^{+}96, SGM^{+}97, HOS^{+}97, MAMJ95]. GLUT-3 [HOS^{+}97]. GLUT-4 [MTL^{+}96, SGM^{+}97, MAMJ95]. GLUT4 [HLSM95, KSP95, PvDA^{+}98, VYB95, WBS^{+}98, BL99, WHM^{+}96]. glutamate [WS96]. Glutamine-repeat [DSVL^{+}98]. glutaminyl [AKP95]. glutathione [KRMMA96, LMS97]. glycan [PGS95, GSB97]. Glycans [BRS99, HZC^{+}97, WBG^{+}97]. glyceraldehyde [RWO95]. Glycerol [RHLG98]. Glycine [KWK^{+}96, NIK^{+}97]. Glycine-rich [NIK^{+}97]. Glyco [vIH98]. Glycol [SRR^{+}97]. Glycolipid [GH95, DTB96, RR96]. Glycolipid-anchored [GH95]. glycolipid-based [DTB96]. glycolipid-enriched [RR96]. Glycolytic [HCLG99]. glycoprotease [CFS98]. Glycoprotein [FFDP98, GXE^{+}99, BOLS95, LCP95, MPB^{+}95, ORW^{+}95, SCM^{+}96, BNW^{+}99, GGD^{+}97, HPD^{+}97, MBDS^{+}98, RDP^{+}99, SCH^{+}98, TSD^{+}97, DSvdV^{+}96]. Glycoproteins [HHRdB^{+}98, HGG96, LCP95, dPF95, SPMB^{+}96, SFH96, VOP^{+}95, GCAHW97, JBRA98, TH97, TGTL97]. Glycosaminoglycans [RSC96]. glycosomes [HSWC96]. glycosphingolipid [BvdBvM96]. Glycosyl [LMB^{+}95, vdBCH^{+}95, ANS^{+}97]. glycosylated [YLB^{+}97]. Glycosylation [BNA^{+}96, GBS99, HHRdB^{+}98, SGTH99, SZNS98, YHS^{+}99, BMG^{+}95, HJJJ97]. glycosylphosphatidylinositol [BLC95, HE96, KE98, LMC^{+}98, MBOC97]. glycosylphosphatidylinositol-anchored [HE96, KE98, MBOC97]. Glycosyltransferase [RHLG98]. Glycosyltransferases [SWR^{+}98]. glyoxylate [LOD96, MLO^{+}95]. Glypican [LHR^{+}97, MVvdBD96, SVD96, WYY95, CGSY^{+}99, GKS^{+}98, LHR^{+}97]. Glypican- [CGSY^{+}99]. GM130 [NRW^{+}95]. GMAP [IRMF^{+}99]. GMAP-210 [IRMF^{+}99]. Going [Col99]. goiter [KKA96]. Golabi [CGSY^{+}99, GKS^{+}98]. goldfish [BOLS95]. Golgi [AMJM95, AFHS99, ALC^{+}99, ABRB95a, AS96, BB97, Bar97, BS97, BvdBvM96, BKJL^{+}98, CSA^{+}97, CIRG99, CES^{+}98, CSY^{+}95, CS96b, DMP^{+}96, DSM^{+}96, DDF^{+}98, DFD^{+}99, FTB97, FNH99, GE97, HW98, HTK^{+}98, HW96a, HKO^{+}98, HM96, HME^{+}98, IRMF^{+}99, ISH95, JMS97, KLW^{+}96, KBF^{+}98, KHHS97, KBW^{+}96, LMM^{+}99, LAW^{+}98, LLNH^{+}98, LLG^{+}99, LYH^{+}99, LSCM^{+}95, LHS97a, LRS^{+}98, LLSO98, LSZ^{+}99b, LHS96, LSKB95, MM97a, MSW^{+}98, MAT^{+}98, MM99b, MR95, MCS^{+}97b, MW95b, NOR^{+}96, NRW^{+}95, NAsG^{+}98, NCS95, POR^{+}97, PPBJ96, RMWW95, RvSL^{+}95, RHLG98, RSB^{+}99, SLL96, SLSW96, SPL^{+}97, SLW^{+}96, SSSKR95, SPS^{+}97a, SMCE97, SME98, SHP^{+}97, SCPPW98, SW99, SIAS96, SOR^{+}96, SLL^{+}98, SS98, SNCH95, SWR^{+}98, VCL^{+}98, VCS^{+}99, VS98, WGR^{+}95, WSCN97, WJM^{+}99, WBG^{+}97, WBKB97, ZWT^{+}97, vIH98]. Golgi-derived [LLSO98, SOR^{+}96]. Golgi-localization [SSSKR95]. Golgi-localized [CES^{+}98]. Golgi-Related [vIH98]. Golgi-resident [SWR^{+}98]. Gonad [GCAHW97]. gondii [KQB^{+}98]. Govern [VWO99]. governs [DCW^{+}96]. gp25L [DDF^{+}98]. gp25L/ [DDF^{+}98]. gp41 [MBDS^{+}98]. gp64 [MPSC98, PZ96]. gp75 [RDM^{+}96, VXBH95]. GPC3 [GKS^{+}98]. GPI [LMC^{+}98, AFG^{+}98, BGK^{+}95, DvD95, DTB96, HER95, MWC95, MVvdBD96, RJP^{+}96a, SM95c, WYY95, WCV^{+}97, BRS99]. GPI-anchored [AFG^{+}98, BGK^{+}95, DTB96, WYY95, MWC95, MVvdBD96]. GPI-linked [DvD95, RJP^{+}96a, SM95c]. Graded [CSM97]. gradients [DPH95, MHK^{+}95]. Granin [KGHG98]. Granular [HMB^{+}98]. Granule [CMHT96, FGA^{+}96, LW96, MST^{+}96, EvdESvdH^{+}97, MMK^{+}97]. granules [CMHT96, DHT96, LBL95, CHC97, HOS^{+}97, KQB^{+}98, KGL^{+}99, KKG^{+}98, KGHG98, SGM^{+}97, UPT98]. Granulocyte [SRR96a, WDL^{+}98]. GRASP [DC96b]. grasshopper [ZN95a]. Grd19p [VS98]. Green [ACSN98, OTQM98, CRLM96, OMO95, SG95, GCML98, GGPG99, KFL^{+}98, LHS97a, PÖS^{+}98, SSI^{+}97]. Grl1p [CMHT96]. group [KKAN^{+}96, BHSAJ98, FSC^{+}97, SSW^{+}98]. grow [RFNF96]. Growing [FSH^{+}99, CLS96, CFK95, FTH96, RPC^{+}95, ZP98, ZCOP99]. grown [ESR97]. Growth [ASES^{+}97, AGS^{+}97, BMC^{+}99, IKT^{+}96, JCR97, LBBP98, MN98, NCV^{+}98, Ped98, SEH^{+}97a, SSR^{+}98, TSC^{+}98, AC95a, AM95b, BJS^{+}95, BKP^{+}95, BLB^{+}96, BS95a, BKEC96, BFS^{+}95, CGSP95, CMUW96, DMS^{+}95, DPC^{+}96, DSE^{+}95, EHM^{+}95, FLW^{+}95, GJ95a, GHP^{+}96, GTL^{+}96, HKM^{+}96, HIG^{+}95, HWFA96, JSR96, KM95, KFE96, LDF^{+}95a, LS95a, LZD95, Mah96, MGM95, MMMT95, MTGY96, MR96, NIM95, PMMF95, RME^{+}95, RVD^{+}95, SLT^{+}96, SKA^{+}96, SCG^{+}96, SSM^{+}95, SCK^{+}95, SS95b, SVD96, SSW^{+}96, TIJ^{+}95, THK95, TK95, VTG^{+}96, VH96, VKP^{+}95, WMC^{+}95a, WMC^{+}95b, WKJH^{+}95, WD95, YSM^{+}95, YSB96, ZMM^{+}95, ZM95, ZL95a, ZSMV95, dCVCV95, BFL^{+}97, BZBA99, BKJL^{+}98, CAS97, Car99, CEZA97, CGC^{+}97b, DBC^{+}99, DH99, DPM^{+}97, EWW^{+}98, FSH^{+}99, GWLH98, GGPG99, GTY98, IKR^{+}98, KHA^{+}99, KB99a, KLS^{+}99, LMF98, LHB^{+}99, MGB^{+}97, MBTW97, MFBK99]. Growth [MQR^{+}97, NIK^{+}97, NGMR97, OKE^{+}99, OOB99, PKQ^{+}98, RHW^{+}98, RPS99, SS97, SYT97, SPR^{+}98, SBPR^{+}98, SKS99, SWD^{+}98, SGGW99, SEBF98, TCM^{+}97, WWO^{+}98, ZTH^{+}97, ZSB99, ZS98]. Growth-associated [MN98, IKR^{+}98]. Growth-dependent [FSH^{+}99]. Growth-inhibitory [SEH^{+}97a]. growth/differentiation [MMMT95]. gruberi [HPKL97]. GTP [DDRC99, HCAW95, LVD96, LD95, MCJK98, MF96a, MGY^{+}95, OWAlW99, OMS98, SRH^{+}95, ZJB^{+}97, GKP^{+}97]. GTP-binding [DDRC99, LVD96, MCJK98, OWAlW99]. GTP-driven [OMS98]. GTPase [MCB96, AGP^{+}98, CRLM96, CKS^{+}95, FaAR^{+}96, HTM^{+}98, JRLS95, JLRS95, MWB98, MW97, MTLW95, OBW98, OKB^{+}98, RKD96, SWK^{+}99, SG96, BPD99]. GTPase-activating [MWB98]. GTPases [BLL^{+}96, BOM^{+}98, BMHH97, CKB^{+}98, HH95b, JMS97, JN98, JSN98, MSH^{+}98, NH99, TXB^{+}99, vLKvdK^{+}97, GvdFvD^{+}99]. Guanine [vLKvdK^{+}97]. Guanosine [CNG98]. Guanylate [KNH^{+}97]. guanylyl [HCAW95]. guanylyl- [HCAW95]. Guarded [RC98]. guidance [OFY^{+}95, LDP^{+}99].

H [CF95, EPMB^{+}99, FWG^{+}96, RME^{+}95, RHW^{+}98, VB98, WSG^{+}98, GHS98, KTAX98, MWL^{+}96, SEH^{+}97a, WMC^{+}95b, ZLL^{+}98]. H-Ras [EPMB^{+}99]. H-rev107 [SEH^{+}97a]. H1 [PFAF97, WGR^{+}95]. H3 [SAR^{+}99]. Haemanthus [KCBR96]. hair [APG^{+}96, EWS96]. half [DSGF95, SCG^{+}95]. half-desmosomal [DSGF95]. Hamster [FNH99, NNSN97, BBZ^{+}95, HP95, ZOS^{+}97]. hand [MBP^{+}95]. Hansenula [TWVC95]. HB [IKR^{+}98, NIM95, MN98]. HB-GAM [MN98]. hBUBR1 [CSY98]. hcf106 [WRC^{+}99]. Hcp [MMR99b]. Hcp-1 [MMR99b]. HD1 [SBG^{+}98]. HD1/Plectin [SBG^{+}98]. hDlg [LBCB96]. Head [MMD^{+}98, WTM^{+}99]. Head-to-Tail [MMD^{+}98]. Heads [BWWT97, HH98]. Healing [BMLM96, ZMM^{+}95, CMD^{+}98, OTL^{+}97]. heart [BSMH95, RBL^{+}96, CSM97, ESH^{+}99, FAZ^{+}99, GBvdN^{+}98, NMK^{+}97, YDFB98]. Heat [BSAJ97, BM96, LCM^{+}95, SFD^{+}96, SRG^{+}96, SHC^{+}97, SSH^{+}98]. Heat-denatured [SHC^{+}97]. Heavy [BGH^{+}97, DRR96, KSG^{+}96, KFE96, NSYK^{+}95, ROG^{+}96, VGM96, ASES^{+}97, BRLM98, EFK^{+}98, KSY97, KAFB99, MKW^{+}99, RKRG99, RHW^{+}98, SLAS^{+}98, SSRP99, SCB^{+}97, TWP98, YWMH99]. hedgehog [DDWM97, FK99]. HeLa [AAP^{+}95, EELF^{+}95, MAM95, MAS^{+}97, SHP^{+}97, VKIH98]. Helical [SDDV97]. Helicases [SSS^{+}99b]. Helicobacter [CHH97, LHM^{+}96]. Helps [SWHG98]. hemagglutinin [BSD^{+}96, DPHW96, LNR^{+}96, MWC95, NDBR95, CLF^{+}97a, CFL^{+}98, HZC^{+}97, KS98, LNRR98, MBOC97, PMBM^{+}99, QPH^{+}98, RLVM98]. Hemagglutinin-induced [MBOC97]. hematopoiesis [ZM95]. hematopoietic [SBS^{+}96, SCM^{+}96, TSS^{+}95b, MPR^{+}97, PHN^{+}98, WDL^{+}98]. HEMCAM [VDA^{+}96]. hemichannels [LLL^{+}96]. Hemidesmosome [SBG^{+}98, DYF96]. hemidesmosomes [SEC^{+}95, BKN^{+}97, RTM99]. hemifusion [BL96, MWC95, CFL^{+}98, MBOC97]. hemopoietic [VDA^{+}96, MNM^{+}98, SG99b]. Hensin [VTHAA99]. HEp2 [GM95]. Heparan [HYK^{+}98, MVvdBD96, WYY95, EY99, FSA^{+}99, GCM^{+}97]. Heparin [DAP^{+}95, SCG^{+}96, BJS^{+}95, HIG^{+}95, LF95, MMMT95, MKI^{+}96, NIM95, IKR^{+}98]. Heparin-binding [MN98, BJS^{+}95, HIG^{+}95, LF95, MMMT95, NIM95, IKR^{+}98]. Hepatic [AAB^{+}97, SATW98, TFYH99]. hepatitis [KMOO95, KSV^{+}99]. hepatocarcinoma [VOP^{+}95]. Hepatocellular [SPBCM97]. Hepatocyte [AGS^{+}97, BFS^{+}95, SW98, TIJ^{+}95, KMOO95, SCG^{+}96, WKJH^{+}95, YSM^{+}95, ZM95, CGC^{+}97b]. hepatocytes [BLB^{+}96, MKB^{+}95, IMS^{+}98]. Hepatoma [SW98, ZH97]. hepatopoiesis [ZM95]. Hepatotoxic [KMS^{+}98]. HEPG2 [SPBCM97, VOP^{+}95, ZH97, vIZKH97, vIH98]. Heptahelical [HPL99]. HER2 [VGVF97]. HER3 [VGVF97]. HER4 [VGVF97]. Herpes [SEH97b]. Heterochromatic [LSB98]. Heterochromatin [AKLME98, HFP^{+}98b, PCQH98, SSW^{+}98, TRS^{+}99]. Heterochromatin-binding [PCQH98]. Heterodimer [XMMW97, ZMM^{+}95]. heterodimeric [WHH96, YNOH95]. Heterodimers [CCMC97, BKB^{+}96, GKR95]. heterogeneity [ZHR95, MP98]. Heterogeneous [EJ97, FST99, HBH98, SEK^{+}97]. Heterokaryons [EAM^{+}97]. Heteromeric [XCLM98]. Heterophilic [CT97b, DC96b, HSL^{+}99, VLZR96]. Heterotrimeric [LAW^{+}98, MS97b, RP97, TZL^{+}98, DMP^{+}96, WMR^{+}96, FAZ^{+}99, GZC^{+}99, MPL^{+}97]. Heterotypic [HRD98, MNM^{+}98]. HGF [BLB^{+}96]. HGF/SF [BLB^{+}96]. HGM [BLB^{+}96]. Hierarchical [SHK^{+}99]. hierarchies [MTC^{+}95]. Hierarchy [MSF^{+}95, DAF^{+}97]. Higashi [FRT^{+}98]. High [ACSN98, ASP^{+}97, CBM^{+}98, CSF^{+}97, FSC^{+}97, HMB95a, MAS^{+}97, BKP^{+}95, BGN^{+}96, CLL95, KKAN^{+}96, LHDC95, PFGS95, WS96, BBL^{+}99b, CFS98, GBFL^{+}98, RHLG98]. high-affinity [WS96]. High-resolution [ACSN98]. higher [AGB^{+}96, BvdBvM96, SGS95, SG95]. highly [BGK^{+}95, SCKG95, SS95b, WYY95, EAOS^{+}98, GRTB97]. Highways [BG98]. Hinge [MCBE98]. Hippocampal [ALC^{+}99, dHvPL^{+}95, EY99, MBBT98, WNB97]. Histocompatability [SKC97]. histocompatibility [AM95a, RvSL^{+}95, BRAM97, FKWP97, KMG^{+}97, LMB98]. histocompatible [VWK^{+}95]. Histogenesis [GSP^{+}98]. Histone [MCPB99, SJA96, SAR^{+}99]. histones [TGG95]. HIV [JMWG98, ZZYG96]. HIV-1 [JMWG98, ZZYG96]. HLA [CWT^{+}98, MRvD^{+}95]. HMG [HSB98, LW95, TMM^{+}95]. HMG-14 [HSB98]. HMG-17 [HSB98]. Hnrnp [KBC99, BSAJ97, IJB^{+}97, MFMW95, ND96, SD95]. HOG [RHLG98]. Homeobox [LDD^{+}97]. homeodomain [MMS96a, KGS98]. Homeostasis [LCD^{+}98, MAS^{+}97, FJ98, GWB^{+}98, HLMD97, KBF^{+}98, TBD^{+}99, WRG98]. homing [JBD^{+}95]. Homo [XCLM98]. Homo- [XCLM98]. homodimer [LCH96]. homodimerization [MDQG96]. Homodimers [IMW^{+}99, GWLH98]. Homogeneously [LSB98]. Homolog [LZK^{+}97]. Homologous [FMD^{+}98, TOSF^{+}98, ZWT^{+}97, BGK^{+}95, BSAJ97, SCK^{+}95, BH98, MNG^{+}98, VvZ99]. Homologue [KJE^{+}99, MWGE97, THM98, ARM^{+}95, GBW^{+}95, HKL96, HUU96, IMFS96, KGHMT95, MR96, MBS^{+}96b, OP96a, PMH96, SHR^{+}95, WMXE96, BNW^{+}99, BHHF97, CSC^{+}98, DRE97, EZC^{+}97, IHT^{+}98, KWS^{+}99, KNT^{+}99, LM98, OC98, PKBHK97, RCE^{+}99, RB98, SHC^{+}97, SSSG98, TCMB98, TKBM97, WGF^{+}98, YMD99]. homologues [AASH^{+}96, MF96b, KLS^{+}99, LRAB99, MUS98]. Homology [MSH^{+}97, VB98, DRAT95, GLS96, MMATe96, MMS^{+}96b, RHS^{+}96, VSGF96, ABE^{+}98, DSC^{+}99, FCMN97, HBS^{+}97, HR99, Li97, WKL^{+}99]. homophilic [BYG96]. Homotypic [UPT98, WLK97, FBH^{+}95, BBR97, GSP^{+}98, UvMJ^{+}99]. Homozygous [TKH^{+}97]. hook [KP96]. Hormonal [SYA95]. hormone [EMB^{+}96, SSA^{+}96, ANT^{+}97, SKS99, SSSS97]. Host [RBM^{+}99, RSR^{+}96, DSC^{+}95, RROA95, BGL^{+}99, ISM97, PDO^{+}99]. Hox [BAS^{+}97]. HP1 [AKLME98]. HPV16 [ZSMV95]. Hrb [DSC^{+}99]. HRG [TCM96]. hrp23 [SAEV^{+}98]. Hset [MSH^{+}99]. HSP47 [SHY^{+}96]. hsp70 [ATD97, SHC^{+}97]. Hsp70s [SHR^{+}95]. Hsp78 [SNL96]. Hsp90 [OSR^{+}98]. Hst [OST^{+}95]. Hst-1 [OST^{+}95]. HSY [TT96]. HT [HHRdB^{+}98]. HT-29 [HHRdB^{+}98]. hTom20p [MGS96]. Human [BGH^{+}97, CLR^{+}96, CJS^{+}99, CLT^{+}95, CWM^{+}98, DSvdV^{+}96, FRMLT^{+}99, FIC^{+}97, FdEP99, ISM97, LEB^{+}97, MSA98, MBS^{+}96b, RMY97, SOR^{+}95, SKC97, TXB^{+}99, THM98, VNTM^{+}99, WDV^{+}97, WNB^{+}95, AASH^{+}96, BGR^{+}95, BGK^{+}95, DTB96, DDWG96, EDB^{+}96, FWG^{+}96, GHP^{+}96, GMS^{+}96, GMFN95, HSL^{+}95, HBJ95, ICW^{+}96, JGM96, KCN^{+}96, KvEdV^{+}96, KSL^{+}95, LN96, LR96, LHDC95, LLK^{+}95, LLRTP95, MCL^{+}96, MGS96, MPB^{+}95, MMS^{+}96b, OCS^{+}95, PLR^{+}95, PL96, PFGS95, RWM^{+}96, SHS96, TAK95, TEL^{+}95, VTG^{+}96, VXBH95, VOP^{+}95, WLML95, YvdEH^{+}95, ZMM^{+}95, AN99, ASP99, BHHF97, CWT^{+}98, CFS98, CBB^{+}99, DMY^{+}97, FRT^{+}98, FH97, FHSM^{+}97, GKO99, HBS^{+}97, HOS^{+}97, HAB^{+}99, JP98, KMG^{+}97, KLL^{+}97, LNH^{+}98, LHWH97, MCS^{+}97a, MAC^{+}97, MRM^{+}98, MSWL^{+}99, MBDS^{+}98, NRD98, NPJW99, OT97, PWD^{+}99, PGP^{+}99, PH97b, SPBCM97, SSW^{+}98, SPS97b, SSRvdB98, TCMB98]. Human [TKBM97, TLBK98, VHF97, WPW^{+}97, WSWR99, qXLCH97, XZSC98, ZH97, ZhTA^{+}97, ZLB^{+}97, vHGB^{+}97]. humans [LKD^{+}95]. Huntingtin [EGKC^{+}99, HSW^{+}98]. hVPS34 [SMS^{+}98]. hyaluronan [BNA^{+}96, BMG^{+}95, CAS96, ZKH^{+}95, BNW^{+}99, SZNS98]. hyaluronan-dependent [CAS96]. hyaluronate [GSvR^{+}96, SRH^{+}96]. hybrid [HSWC96, SZB^{+}95, MBF^{+}98]. hydrolases [CS96b]. hydrolysis [HCAW95, MGY^{+}95, OSR^{+}98]. hydrolyzable [HCAW95]. hydrophilic [DMG96, ZL95b, SG97]. Hydrophobic [MSM99, WS97]. Hydrophobicity [HW96b]. Hydrostatic [RSS98]. hydroxyacyl [TOSF^{+}98]. Hydroxyurea [NNSN97, BBZ^{+}95]. hydroxyurea-arrested [BBZ^{+}95]. Hyperactivation [DH99]. hyperoxaluria [LOD96, MLO^{+}95]. Hyperphosphorylated [KDBW97]. hyperphosphorylation [SP96]. Hyperproliferation [MDC^{+}99]. Hypertrophic [BGH^{+}97, REA95, KNSP97, NMNL98]. hypertrophy [FSAP95, CMS98, EIIM^{+}98, SNK^{+}97]. Hyphal [MHM99]. Hypodermal [HSSW99]. hypothyroid [NA96]. hypothyroidism [KKA96]. hypotrichous [FC95].

I-Band [LRC^{+}99]. I-cell [GMS^{+}96]. I-coated [SWC^{+}96]. I-independent [MW95b]. I1 [HOCK98]. IA [CJH^{+}98]. IAP [HBPJ98]. IAP75 [MKLS96]. IAP86 [MKLS96]. IB [CJH^{+}98, JOPM98, GXE^{+}99]. ICAM [YHD^{+}98]. ICAM-2 [YHD^{+}98]. ICAMs [dPCM^{+}97]. ICAP [CWSL97]. ICAP-1 [CWSL97]. ICE [DVD^{+}96]. ICE-family [JWR96]. ICE-related [MWGE97]. Icsa [ELL^{+}99]. ida5 [KMHK97]. identical [YKC95]. Identification [ASH95, ANK97, BLC95, BPT96, CHP^{+}98, CRRF97, CSM97, CSDR96, DBOM98, DSM^{+}96, DSM^{+}98a, EB96, FGP96, GKP^{+}97, GLP^{+}95, HOCK98, HN98, HC97, HHB^{+}97, KKAN^{+}96, KJE^{+}99, LSZ^{+}99b, MYTK99, MZN97, OSHG^{+}96, OK96, PG95, PH98, PMCT96, RBB97, SMRT96, SYF^{+}97, SATW98, SOL^{+}95, TLF^{+}95, WM95, YYSB97, AKP95, DLB96, HBJ95, SSP95, VMN95, VXBH95, AKLME98, BBP^{+}98, MPL^{+}97, PvDA^{+}98, SKR^{+}98, TGTL97, XWK^{+}97]. Identified [KFL^{+}98, LBO^{+}98, MWF^{+}99]. Identifies [RST^{+}99, GAW^{+}96, WMXE96, DDM98]. identify [LLN^{+}95, PWK^{+}96]. Identity [DDWM97]. IFT [PWW98, PSH^{+}98, CDH^{+}98, SWO^{+}99]. Ig [BTTB96, HJJJ97, KGF^{+}97, LDP^{+}99, MSS^{+}99, SEBF98]. Ig-like [MSS^{+}99]. Ig/Fibronectin [KGF^{+}97]. IgE [SM97]. IGF [EBR96]. Ii [BRAM97, Col99, AM95a, ACH^{+}98, BSW^{+}99, BRAM97, BDvdZW95, BRHD95, BDO^{+}97, CCB^{+}95, CGVS^{+}96, CDH^{+}98, DRR96, DDF^{+}98, DBLD^{+}99, DW97, FRT^{+}98, FKWP97, GWLH98, GACM^{+}95, GMS^{+}96, GM99, KJE^{+}99, KDBW97, KSY97, KMG^{+}97, LAW^{+}98, LWS99, LL98b, LMB98, LYR96, LYLR98, MRvD^{+}95, MML96, MFBK99, MOY^{+}98, MKS^{+}97, MS97b, MCRB97, NHA^{+}97, NCV^{+}98, OIT^{+}95, OZF^{+}99, RHF^{+}96, RPY^{+}99, Sch96, SJF^{+}97, SEH^{+}97a, SVMB97, TOSF^{+}98, TZL^{+}98, VSB95, WLR96, WMR^{+}96, WFBO^{+}96, YtDH^{+}95]. II-dependent [GACM^{+}95]. II-invariant [AM95a]. IIA [TB99, ZOKS99, SLAS^{+}98]. iib [SRA^{+}99]. IId [SLAS^{+}98]. III [BGL^{+}99, HSML96, KGF^{+}97, MFMW95, MSF^{+}99a, STS96, SEB99]. III-10 [HSML96]. IL-1-induced [GL95]. Image [HST^{+}98, JM97b]. Imaginal [MZL^{+}99]. Imaging [SLPE^{+}98, GJ95a, KE98, CWT^{+}98, KMK^{+}98, MKC99]. Iminodipropionitrile [ZLL^{+}98]. immature [DHT96, KKG^{+}98, UPT98]. Immediate [ISM97]. Immobilization [MPW95]. Immortalization [SSS^{+}99b]. immortalized [LGdC95, PNL^{+}95, ZSMV95]. immune [FHV^{+}96, GLS96]. immune-privileged [FHV^{+}96]. immune-receptor [GLS96]. immunodeficiency [SPMB^{+}96, FRT^{+}98, LHWH97, MBDS^{+}98]. immunoelectron [OGS^{+}96]. immunofluorescence [NKI^{+}95]. Immunoglobulin [TNM^{+}99, CSM^{+}96, LDF^{+}95a, LHDC95, LGRB96, dHvPL^{+}95, FHSM^{+}97, GSP^{+}98, MPLS^{+}98]. Immunoglobulin-like [TNM^{+}99, FHSM^{+}97]. Immunohistochemical [BTNZ^{+}95]. Immunolocalization [CFC^{+}95]. immunoreactive [HMB95a]. Immunoreceptor [LKSW98]. imp2 [DS98b]. impact [ZN95b, BHSAJ98]. Impair [MMN^{+}99, QPH^{+}98, YAiN^{+}98]. Impaired [ABL98, GHP^{+}96, BCAS98, NWM^{+}97, PBJ^{+}98]. Impairment [TTT^{+}99]. Impairs [FIC^{+}97, LWB^{+}95, RRRA95, MVM^{+}99, NT97, VKIH98]. Imperatoxin [STG^{+}99]. Implicate [BBOE98]. implicated [BNA^{+}96, FLPS95, GMFN95, NOR^{+}96, KSY^{+}98, POR^{+}97, PLB97, SVB^{+}97]. Implication [CLS^{+}97, BBOB96, CAR^{+}97]. Implications [GRB^{+}97, TSC^{+}98, CSDR96, HMB95a, MJMS96, MS96, MLO^{+}95, ST96, WAMS96, ZMM^{+}95, ZKSB96, ACF^{+}97, CSM97, EGS^{+}98, HST^{+}98, KPZ^{+}97, MNR^{+}98, MHC^{+}97, MPCW97, RFVCM^{+}99]. imply [MF96b]. Import [HSWC96, KB99a, KSF^{+}99, MYTK99, CRLM96, CAA95, CAVA96, CKS^{+}95, DG96, EKK^{+}96, GLB^{+}95, IWW95, KWCS96, MKLS96, MGS96, MGY^{+}95, PWK^{+}96, PG95, PMMF95, ST96, SG96, TKVC95, WNB^{+}95, ZL95b, APRB98, CWSG99, EEHW^{+}98, GRS^{+}99, HTWC98, HRK^{+}98, HTY^{+}99, IW97, JMWG98, KBD99, KWN^{+}99, KS97, KCFS98, LJK97, LHS97b, MVM^{+}99, PRB97, PRB99, RPB97, RPBB98, STF98, SEK^{+}97, TKBM97, WRdVC97]. Important [CWSL97, FPWN95, HBJ95, SMRT96, LMF98, MHWW98, SL97]. imported [MLO^{+}95, HSB98]. Importin [HTWC98, TSS^{+}95b]. Importin-like [TSS^{+}95b]. Imports [TB99]. imprints [MHK^{+}95]. Improve [QBvD^{+}98]. Improves [SCB^{+}97]. inability [GHP^{+}96]. Inactivation [BGN^{+}97, HWC^{+}96, CCBL98, GBvdN^{+}98, MCPB99, RKB97, WHG^{+}97]. inactive [CMWL96, EDB^{+}96, JMTCF96, KLN^{+}96, RACHV96, MCPB99]. Inappropriate [PBJ^{+}98]. INCENP [MAEE98, AKLME98]. incisional [BMLM96]. Including [WC97b, PJL^{+}95, GPKH99, KDG98, MLB98]. inclusion [SFH96]. Incoming [SEH97b]. Incorporation [PFAF97]. Increase [CS99, YGG97b]. Increased [ED96, GL95, JM97a, KMOO95, MH95a, BBAN^{+}99, JM97a, KZS^{+}98, MLO^{+}95, MDC^{+}99]. increases [APG^{+}96, BCC^{+}95, DRR96, SRH^{+}96, WHM^{+}96, GRTB97, LYH^{+}99, OKE^{+}99, RS99, ZLL^{+}98, ZS98]. Increasing [WMC^{+}95b]. Independence [Sch97]. Independent [FMD^{+}98, TMN97, CRLM96, CMD^{+}98, CSY^{+}95, CCMG97, DBvdBS95, GE97, HKC^{+}95, JVRNM99, KOM^{+}98, LNM^{+}98, LCN^{+}97, LBCB96, MBF^{+}95a, MG96, SKS99, SR96, VvZ99, WLG^{+}96, ZMG^{+}96a, BJR^{+}99, Car99, EAOS^{+}98, HLJ^{+}98, HW98, HN98, LGM98, LAS^{+}97, OOB99, TRA^{+}99, WRC^{+}99, MLW^{+}95, MMK^{+}97, MW95b, PFHWN98, STK^{+}98, SSN96, SFR^{+}98, ZPL^{+}98]. Independently [AKM^{+}98, ZLPL99, BINRM97, BS97, HIWL97]. indicates [BFL95, MBRN95]. Indicators [OTQM98]. Indirect [SH95]. individual [BSD^{+}96]. induce [DRR96, BSRG99, SOPM^{+}97, ZGYS99]. Induced [MWGE97, VHF97, ABM^{+}99, BHDW95, BLB^{+}96, COML95, CJY^{+}97, CHH97, DBvdBS95, DOSN^{+}99, DOP^{+}98, EB95, EAOS^{+}98, FLW^{+}95, FDH96, GBBSO96, HDV^{+}99, HRD98, JWR96, JGM96, LLN^{+}95, LW95, MMAK97, MFM^{+}95, MPW95, NMT97, NYN^{+}95, OSK^{+}96, PBB^{+}96, PLZ^{+}95, PZ96, RME^{+}95, RVD^{+}95, SPK^{+}97, SAR^{+}99, SRR96b, SKC97, SGYH97, SKWY95, SR95, TLF^{+}95, VB98, ALM97b, BAH^{+}98, CIK^{+}99, CNG98, GKO99, HMD97, OOB99, SBGR98, SGGW99, TSB99, ZKN99, BGR^{+}95, CJY^{+}97, EHC97, GL95, GDM^{+}99, HPD^{+}97, HHR^{+}98, KBF^{+}98, LBBP98, LWSL^{+}98, LDH^{+}98, LYLR98, MH97, MSH^{+}98, MBOC97, QWS^{+}98, REC97, SYT97, SAC^{+}98, SPL^{+}97, SWR^{+}98, ZRHO99, ZJY^{+}98]. Induces [AARS^{+}97, AKM^{+}98, MBA97, SKC97, TMN97, XC97, ZLPL99, BFS^{+}95, CCB^{+}95, DMD^{+}96, DSvdV^{+}96, ICW^{+}96, MWC95, RROA95, SRR96a, SKA^{+}96, SCG^{+}96, SBR^{+}96, SR96, VLZR96, YtDH^{+}95, YWW^{+}96, ZMM^{+}95, AGP^{+}98, BSH99, BBAN^{+}99, CWY99, EWW^{+}98, EY99, GKS^{+}98, GCM98, HCLG99, IFSV^{+}98, KA98, LNM^{+}98, MDC^{+}99, NTAN98, RYB^{+}99, SPR^{+}98, SSSS97, VTHAA99, ZPL^{+}98, ZCWB^{+}98]. Inducible [ALM^{+}97a, SRRB95, BBR97, NIK^{+}97]. Inducing [MCJK98, GCE^{+}97]. Induction [BAS^{+}97, CSEM99, GvdFvD^{+}99, KD97a, mLZB99, LCN^{+}97, LK96, MG96, SLN^{+}97, TYT^{+}99, cXP97, ZS98, FLW^{+}95, FJT^{+}95, FFMG95, XZSC96, FEF^{+}97, KLM^{+}98, LPM^{+}98, SBPR^{+}98, VWG^{+}97, vHGB^{+}97]. Inductive [NG99]. Inefficiently [MVM^{+}99]. infected [FGA^{+}96]. infection [KKGD96, KGMF95, MGP97]. Infectious [MM97a]. Infertility [KKT^{+}98]. inflammation [HWC^{+}96]. Inflammatory [HMB^{+}98]. Influence [DFGL96, ZKH^{+}95, GBS99, HSW^{+}98, WHB^{+}99]. influences [CMHT96, WSS97, WCV^{+}97]. influenza [BSD^{+}96, DPHW96, LNR^{+}96, MWC95, SLL96, CLF^{+}97a, CFL^{+}98, HZC^{+}97, KKM^{+}97, KS98, LNRR98, MBOC97, PMBM^{+}99, QPH^{+}98, RLVM98]. Influx [EKF^{+}97]. Infoldings [BFRB99]. information [BHDW95, HCAW95]. inherent [Bor95]. Inheritance [XMMW97, NWPW95, XW96, BSY97, FY97, FY99, HKS97, SY99, SCPPW98, WCW98]. inhibit [FB96, FMJ^{+}95, KP96, LNY^{+}95, TSS^{+}95a, SBS^{+}98]. inhibited [AS96, HSWC96, LNM^{+}98, WTM^{+}99]. inhibiting [HDV^{+}99, KKGD96]. Inhibition [DSG^{+}97, DNB^{+}96, HCM^{+}96b, KSVZ95, KRMMA96, LOD96, MWGE97, MSF^{+}99a, SHC95, YSK^{+}97, ZRHO99, BEZ^{+}95, BSK96, DNC^{+}95, FLW^{+}95, GHJJ95, IKT^{+}96, MKI^{+}96, SKA^{+}96, VKP^{+}95, WS95b, FhJZ^{+}99, HHS^{+}99, HLK^{+}98, OOB99, OF98, SSR^{+}97, TSD^{+}97, WLF99]. Inhibitor [SSR^{+}98, AHBW96, DCW^{+}96, DVD^{+}96, MG96, SJHC96, SMC^{+}96, YGK96a, ASP^{+}97, DMY^{+}97, HDDKH98, MCC^{+}97, MFBK99]. inhibitor-1 [DCW^{+}96]. inhibitors [AS96, TMM^{+}95, HBPJ98, TFYH99]. Inhibitory [BHD^{+}97, JGM96, RZZ96, RCKS95, SEH^{+}97a, ANRT99, DPZ^{+}97, MBDS^{+}98, SWD^{+}98, SZNS98, TTT^{+}99]. Inhibits [BRM^{+}99, HDV^{+}99, HHRdB^{+}98, PW98, AHM^{+}95, BGR^{+}95, GG95a, LLRTP95, MAS^{+}95, MELC96, OSK^{+}96, RWM^{+}96, RR96, SBS^{+}96, SAPLHD95, SL95, WvdVV^{+}95, WMC^{+}95b, ZZvdB^{+}96, BBW^{+}98, DOP^{+}98, EGS^{+}98, GTY98, HTY^{+}99, LDH^{+}98, LRS^{+}98, MLB98, MQR^{+}97, NRP98, QWS^{+}98, SGGW99, SMG^{+}97, WBG^{+}97, YGG97b, ZRHO99]. Initial [BDO^{+}97, OSH^{+}96, ABN96, PZ96, SYT97, ZLM^{+}98]. initiate [ZN95a, EHC97]. Initiated [SBG^{+}98, SHK^{+}99, VTG^{+}96]. Initiates [RL99]. Initiation [TYT^{+}97, LKWG96, WRW^{+}95, BMS^{+}97, MCPB99, YHMS^{+}98, ZRHO99]. Initiations [FMD^{+}98]. Initiators [KKH^{+}99]. Injury [PJE^{+}99, FJT^{+}95, PTBC96, KMS^{+}98]. Inner [MBOC97, PMH96, FPM^{+}95, KLMR95, SW95, EAM^{+}97, ESM^{+}97, HS97, HOCK98, HGG^{+}97, KMHK97, KHS^{+}97, KD97b, KCFS98, LJK97, LHS97b, MAEE98, MKW^{+}99]. innervation [APG^{+}96]. iNOS [GL95]. Inositol [WS95a, TT96, FSM^{+}99, IMK99, PFP97, YHMS^{+}98, VB98]. inositol-labeled [VB98]. Insect [SRR^{+}97]. insensitive [PFP97]. Insertion [KHS^{+}97, SSS99a, ZP98]. Insertional [LBWF^{+}97, MS99]. Insertions [SG97]. Inserts [BGL^{+}99]. Insights [TYT^{+}99, Blo98, DSvDH^{+}98, LMM^{+}99]. Insoluble [Sal99, DvD95, SDL98, SBS^{+}98]. Instability [HSLB99, TWS97, WSS97]. Insulation [RRH^{+}95]. Insulin [HLSM95, JSR96, MFBK99, MBS95, WHM^{+}96, MTL^{+}96, MBF^{+}95a, PGS95, SLM^{+}95, VYB95, VUC^{+}95, BCAS98, BL99, CMC^{+}98, FLO^{+}99, PvDA^{+}98]. Insulin-like [JSR96, MFBK99, NCV^{+}98]. insulin-responsive [VYB95]. Insulin-sensitive [HLSM95, MTL^{+}96]. insulin-stimulated [PGS95]. Intact [LCD^{+}98, MAS^{+}97, DC96a, FB96, ALM97b, GLMG98, NRP98, NJB^{+}98, SFA^{+}98]. Integral [JR98, YGG97a, CCP96, DMG96, SW95, TWVC95, vWDG^{+}95, BVKP98, CSEM99, FFH^{+}98, YAC97]. Integrating [CCC^{+}99, FKB99]. Integration [KA98, DWS^{+}95, FPM^{+}95]. Integrin [BRM^{+}99, BLLB95, BSW^{+}99, CHLÅ97, CWSL97, CKB^{+}98, EKSC98, GXE^{+}99, GPRS^{+}95, HMS^{+}98, HLK^{+}98, IRW^{+}99, JCR97, KSE^{+}97, LAS^{+}97, LGRB96, MSWL^{+}99, MTC^{+}95, NCW^{+}99, OWD^{+}99, PWD^{+}99, PH97b, PWGG98, SPV^{+}99, SLT^{+}96, SBG^{+}98, TSC^{+}98, WLF99, WMP^{+}98, XC97, YMAC^{+}98, AKC^{+}95, BEG^{+}95, BZB^{+}96, DAP^{+}95, DYF96, FPM^{+}95, FBD^{+}95, GLD^{+}96, GKR95, HSML96, HBJ95, HH95b, HGW95, HWC^{+}96, HGH96, KKD^{+}95, MPY^{+}96, MBB96, MTGY96, MBS^{+}96b, NIM95, PHU^{+}95, SOHP95, SOE^{+}95, SMFC95, SEC^{+}95, TDW95, WKL^{+}96, WAMS96, WLG^{+}96, WvdVV^{+}95, XGC96, XZSC96, YRM^{+}96, ZMM^{+}95, ABL98, BRP^{+}97, BO99, BFL^{+}97, BKN^{+}97, BGH^{+}98, CAH^{+}99, DHDJ^{+}97, GFN^{+}99, GKO99, GBvdN^{+}98, GZC^{+}99, GTY98, HPS98, HDDKH98, KRCP99, LNG^{+}98, LMW99, LAS^{+}97, MBDBB98, MZNR98, OSM98, PWD^{+}99, RM97, RTM99, RPS99, RdPRW98, SZFM98, SLW^{+}99, SAC^{+}98, SRA^{+}99]. Integrin [SLS^{+}99, TGW99, TCYS99, TDR^{+}98, WPW^{+}97, WMR^{+}98, WYL^{+}99, XZSC98]. Integrin- [GTY98]. Integrin-Associated [WLF99, LGRB96, BLLB95, GLD^{+}96]. integrin-binding [HSML96]. Integrin-dependent [GPRS^{+}95, PWD^{+}99, WLG^{+}96]. integrin-derived [TDW95]. Integrin-induced [SAC^{+}98]. Integrin-mediated [CKB^{+}98, LAS^{+}97, OWD^{+}99, HGH96, WvdVV^{+}95, ABL98, SLS^{+}99, WYL^{+}99]. Integrin-regulated [MSWL^{+}99]. Integrins [MTGY96, Sch97, VNTM^{+}99, HWC^{+}96, HHM^{+}95, LRM^{+}95, LXK^{+}95, PKR95, SR96, XC96, APF^{+}97, FHSM^{+}97, GvdFvD^{+}99, LCB^{+}99, MGD97]. integrity [GKC^{+}96, BHHF97, CGA^{+}99, GTC^{+}98, HMC^{+}99, LMA^{+}97, MS97a, SBR98, SEAB^{+}99]. interact [KKFN^{+}95, MKLS96, OK96, CAH^{+}99, RV98]. Interacting [EGKC^{+}99, KNT^{+}99, ISN^{+}99, KWS^{+}99, SBR98, BHHF97]. Interaction [ARR^{+}98, BVKP98, BLL95, CT97b, GMGGD98, HMS^{+}98, HYK^{+}98, IMW^{+}99, KSJW95, LS95b, NCV^{+}98, PT97, RdPRW98, SBG^{+}98, TNM^{+}99, VLA^{+}98, WLD^{+}98, YTO^{+}99, BEZ^{+}95, BFL95, CAVA96, DFC^{+}96, GBG^{+}96, HBJ95, HHTT95, LTE^{+}96, MDS95, OHB^{+}95, PWG^{+}95, RAE96, SHY^{+}96, SGBD96, SPB^{+}95, SVB96, VV95, Wal95, WB96, dPSMNSM95, BLB^{+}98, BKMN^{+}98, DSC^{+}99, DW97, FST99, GLL^{+}99, GTC^{+}98, HAB^{+}99, IW97, ISM97, LNG^{+}98, LSL99, MTLGC99, OFM^{+}98, PDO^{+}99, PTVD99, TCYS99, TWP98, WUUI^{+}98, XWK^{+}97, YNG98]. Interactions [AFFS96, GK99, GCAHW97, GRB^{+}97, HRD98, KBB^{+}97b, LCD^{+}98, SSEM98, SVH^{+}98, TH97, ARM^{+}95, BS95b, CAS96, DC96b, DWS^{+}95, EHM^{+}95, FAO^{+}96, GPRS^{+}95, KWCS96, MLLT95, MMMT95, NMHS95, ORW^{+}95, PR95, SHR^{+}95, SDTE95, VLZR96, WPMW95, AMY^{+}98, CS97a, CSY98, DGP99, FhJZ^{+}99, HKO^{+}98, HR99, KS97, LBW^{+}97, MNM^{+}98, PBA^{+}97, SF98, TGTL97, VZNR98, WE98, WTM^{+}99, XCLM98, vMNS97]. Interactor [SPV^{+}99]. Interacts [JCR97, KNH^{+}97, EGH^{+}95, IWW95, NRM^{+}96, PG95, PWG^{+}95, SOE^{+}95, TGG95, ZDT^{+}96, AFG^{+}98, BGR98, BSM^{+}99, CDN97, CWSG99, CYH98, CDH^{+}99, HGW^{+}98, KRH98, KRGA98, MSS^{+}99, MHH^{+}99, PLB97, RTL^{+}99, SGSW97, SALdP^{+}97, SOJM99, TSD^{+}97, WGF^{+}98, YHK^{+}97]. Interaptin [RKB^{+}98]. Intercalated [VTHAA99]. Intercellular [JGCS97, BCSG96, LXK^{+}95, MPLS^{+}98, NGT99, SALdP^{+}97, vHGB^{+}97]. interconnected [OPD^{+}96]. interdependent [HAS95]. interendothelial [FBH^{+}95]. Interface [BB97, FZZN98]. interfaces [BCSG96]. Interfere [GFN^{+}99]. Interference [SBS^{+}96]. interferes [ZZvdB^{+}96]. Interfering [HDV^{+}99]. interferon [KGMF95, DOP^{+}98, JLID98]. Interior [dCBR99]. Interleukin [MMAK97, HSL^{+}95, AMY^{+}98, FCMN97, MCS^{+}97a, MSWL^{+}99, SDDV97]. Interleukin-1 [FCMN97, MCS^{+}97a]. Interleukin-6 [AMY^{+}98]. Intermediate [CGC^{+}97a, RdPRW98, ZWT^{+}97, BCSG96, GKC^{+}96, GGR^{+}96, GG95b, KPKWW95, OS96b, RvSL^{+}95, ST96, SVB96, VV95, WKK^{+}95, CSO97, GFN^{+}99, HS97, LSL99, LHS97b, MTLGC99, MBF^{+}98, PYM^{+}98, PTVD99, Sal99, SSRP99, SF98, TWS97, YMPG98]. intermediates [HWFA96, HME^{+}98, LLG^{+}99, SKR^{+}98]. Intermolecular [PSPG^{+}98]. Internal [HFP98a, LCD^{+}98, STF98]. internalization [CNP^{+}95, NDBR95, PBB^{+}96, RRP^{+}95, SPMB^{+}96, SYA95, VYB95, HKOM98, HZR98, OF98]. internalized [FPHH96]. Interphase [FHWV97, LSB98, CMWL96, EDB^{+}96, MAM95, SWC^{+}96, WSGPS95, YvdEH^{+}95, ACF^{+}97, ALVMM98, CH98, ESM^{+}97, MTLGC99, SFA^{+}98]. Interpolar [SMB^{+}99]. Interspecies [AASH^{+}96]. Intervertebral [ACH^{+}98]. Intestinal [CCMG97, DvD95, HG95b, MHK^{+}95, PCL^{+}96, RME^{+}95, BMG^{+}99a, FCTPA^{+}99, LDD^{+}97, WC97a, WRG98]. Intestine [SSSS97]. Intoxication [AvdG99]. intra [NOR^{+}96, ZL95b]. intra-Golgi [NOR^{+}96]. intra-peroxisomal [ZL95b]. Intracellular [AKCC99, ASB^{+}99, CMC^{+}98, DDWG96, GGPG99, HHRdB^{+}98, KZKS96, MBB96, MB99, RP97, RS95, SHY^{+}96, VXBH95, dHvPL^{+}95, vIH98, BKP^{+}95, Cun95, FFMG95, GL95, HSL^{+}95, HTKH96, HH95b, KDvFP96, LCK95, MPY^{+}96, NYN^{+}95, RROA95, TOT^{+}96, WPLK95, YKC95, BHG^{+}97, BRD99, BR98, FhJZ^{+}99, FKWP97, mLZB99, LTW^{+}97, MGP97, MWL97, MGD97, NMNL98, OKE^{+}99, PRV98, RKB^{+}98, TBD^{+}99, VLM^{+}98, VHF97, ZDK^{+}99, ZS98]. Intracrine [WWO^{+}98]. Intraflagellar [CDH^{+}98, RCD99, PWW98, PSH^{+}98]. Intraluminal [HFP98a]. Intralysosomal [ATD97]. intramembrane [BWF^{+}96]. Intramembranous [YSEI^{+}99]. Intraneuronal [SDL98]. Intranuclear [WSA^{+}99, CRRF97, PRB99, RPS^{+}97]. intraperoxisomal [STVR95, ZL96, ESR97]. Intraretinal [LHB^{+}99]. Intrinsic [BINRM97, CAS97, GR97]. Intron [HS96, JVRNM99, RKB97]. Intron-dependent [HS96]. Intron-independent [JVRNM99]. Invaginations [FHWV97]. invariant [AM95a, WLR96, FKWP97, AM95a]. Invasion [TSC^{+}98, RSR^{+}96, CAR^{+}97, HYE^{+}98, KBG^{+}99, KRCP99, MSH^{+}98, SB99c]. Invasive [BPS99, OSM98]. inversely [MVvdBD96]. Investigation [DSvDH^{+}98]. involution [FMJ^{+}95]. Involved [BMG^{+}99b, KWD^{+}98, MMAK97, SPV^{+}99, TXB^{+}99, WSCN97, WLD^{+}98, WBSNV97, ZWT^{+}97, AGPG96, ABRB^{+}95b, BLC95, BIR96, BGK^{+}95, CMWL96, CM95b, DC96b, DLIB^{+}95, HH95a, IKT^{+}96, KBR95, MCA96, MMATe96, MSF^{+}95, MMMW96, MA95, NAP^{+}96, PHU^{+}95, SLM^{+}95, SOR^{+}96, SW95, SLP^{+}96, SPT^{+}95, SPB^{+}95, SG96, TWVC95, TTG95, WFBC96, ATH^{+}97, DCR98, FhJZ^{+}99, FKWP97, HRK^{+}98, HCG^{+}98, HSSW99, KLO^{+}98, KWS^{+}99, KNT^{+}99, LL98b, LCW^{+}98, jLC97, MGB^{+}97, MLC99, MMR99b, MCRB97, NOS^{+}97, NGMR97, PH98, RBRB99, RPB97, SHC^{+}97, TDR99]. Involvement [BML^{+}98, DSG98, FSAP95, GRB^{+}97, GRS^{+}99, INMT97, KSS^{+}95, MN98, MST^{+}96, PAL^{+}98, RPE^{+}97, SRH^{+}95, SBGR98, YYH^{+}96, YTT99, BGB95b, Dav95, FGP96, HSK^{+}96, RL96, AKLME98, CMC^{+}98, DAB99, HOCK98, KBF^{+}98, MCC^{+}97, MZL97, SFS^{+}97, dPSMNSM95]. Involves [SBG^{+}98, AWS96, BBOB96, GEJ^{+}95, GPRS^{+}95, HW96a, REA95, SSN96, DMY^{+}97, FEF^{+}97, GKO99, MSW^{+}98, MPSC98, PXRR99, SSRP99]. involving [FSR^{+}96, GGC96, TMDD96, TT96, GZC^{+}99, IRW^{+}99, SWS97b]. ion [LHK^{+}96, LBL95, MBD^{+}96, SLL96, mLZB99]. Ionization [SBS^{+}99, WJH^{+}98]. Ions [EAOS^{+}98]. Ip [SGGW99]. Ip-10 [SGGW99]. IP3 [RROA95]. Ipab [BGL^{+}99]. Ipac [BGL^{+}99]. Ipl1 [hKsKC99]. Ipl1-related [SGD98]. Iqg1p [OC98]. IQGAP [ATH^{+}97, LL98b]. IQGAP-like [LL98b]. IQGAP-related [ATH^{+}97]. IQGAP1 [BFHB97]. IQGAPs [OC98]. iron [DvD95]. iron-binding [DvD95]. Irradiation [MRM^{+}98]. irreversibly [SMB^{+}95]. isa [WKK^{+}95]. Islet [CCB^{+}98]. islets [SLM^{+}95, ETP^{+}99, MBCS98]. isoform [BZB^{+}96, Kun95, MAG^{+}96, PMCT96, RL96, SH95, TSS^{+}95a, IRW^{+}99, MHH^{+}99, MPL^{+}97, MTK^{+}98, OTS98, PDW99]. Isoforms [HPL97, WLR96, ASSS^{+}95, BCMK95, BJS^{+}95, DMM^{+}96, GDR95, JWH96, KKAN^{+}96, KSG^{+}96, OP96c, RGS^{+}96, SWR^{+}96, ZL95a, ABT^{+}98, ALVMM98, HC99, JLC98, PSPG^{+}98, SL97, TFS^{+}99, WSR^{+}97]. Isolated [GGD^{+}97, PKR95, SZB^{+}95, BRS98, SWS97a, UNPW98, UvMJ^{+}99, VSRP97]. Isolation [HMSK95, LLSO98, PGS95, SGJP99, dCBR95, ZLC^{+}96, HP95, WSH98, WFS97]. Isomerase [GLL^{+}99, HTW97]. isometric [GW95]. isoprenylation [FKS95]. isoproteins [vABSP95]. ITAM [GLS96]. itself [ZL96]. IV [GGK97, MS96, EZC^{+}97, GJW^{+}97, GGK97]. IX [GXE^{+}99, BJR^{+}99, HBH98].

J [MVM^{+}99]. J-related [MVM^{+}99]. jejunum [FGA^{+}96]. jelly [MMS^{+}96b]. Jif-1 [IMW^{+}99]. JNK/SAPK [TMN97]. Jun [MSH^{+}97, ESH^{+}99, FVKS96, IMW^{+}99, KD97a, OWD^{+}99, SWK^{+}99, SKC97, TKVC95, WXS96]. Jun/AP [WXS96]. Jun/AP-1 [WXS96]. Junction [JGCS97, WLF98, BWF^{+}96, BCSG96, CKP96, EELF^{+}95, GYS^{+}95, ISFB^{+}96, KSK^{+}96, KGW^{+}95, LCK95, WHP^{+}96, ZZvdB^{+}96, BSM^{+}99, CAS97, hCMPG97, FSFT98, FST99, GCE^{+}97, HGW^{+}98, HMC^{+}99, HCW^{+}98, IFM^{+}99, IHT^{+}98, LPM^{+}99, MBTW97, MNS^{+}97, NZ97, ONI^{+}98, PXRR99, SFS^{+}97, SFS^{+}99, SV98, WG97, WS98, qXLCH97, ZKN99]. Junction-associated [BSM^{+}99]. Junction-dependent [JGCS97]. Junctional [MPLS^{+}98, BZB^{+}96, LLL^{+}96, SL95, LNG^{+}98, NRD98, WUUI^{+}98, WKR^{+}98]. Junctions [MNS^{+}99, TNM^{+}99, YMAC^{+}98, DZC^{+}96, FSIG^{+}95, FSR^{+}96, LCC^{+}95, NNMW96, RHHRB96, SM95b, UHJ^{+}96, BGBS98, DPT^{+}97, EZC^{+}97, FZZN98, FFH^{+}98, JSN98, LWS^{+}97b, LBB97, MPLS^{+}98, MSFA99, MSF^{+}99b, SFD^{+}98, YHK^{+}97]. JunER [FSR^{+}96]. Jurkat [dBCKS98]. just [WSW95, SAEV^{+}98]. Juxta [YHD^{+}98]. juxtacrine [HIG^{+}95]. Juxtamembrane [YNG98].

K-glypican [WYY95]. K-homologous [BSAJ97]. K-ras [CCMG97]. K14 [LYC^{+}95]. K15 [LYC^{+}95]. Kakapo [SV98, GB98, PURB98]. kangaroo [AASH^{+}96]. Kap114p [PRB99]. Kap121p [MAW98]. Kap122p [TB99]. Kap122p/Pdr6p [TB99]. Kap95p [IWW95, IW97]. kappa [KRMMA96, KRMMA96, LLN^{+}95, RK95]. kappaB [GBBSO96, PSBB96]. Kar1p [SCG^{+}95]. Kar2p [SFNRH95, SHR^{+}95]. Kar3p [BSD97, MBW^{+}99, SHLD97]. KAR5 [BBR97]. Kar9p [MR98, MMR99a]. Karyogamy [THS^{+}98]. Karyomeres [LGM98]. Karyopherin [TB99, MAW98, PRB99]. karyoplasm [MKF96]. Katanin [AYMB99]. KCl [MLW^{+}95]. kD [BRLM98, HBJ95, KSK^{+}99a, WKL^{+}99, ANS^{+}97, BBAN^{+}99, DBOM98, EPVV96, FSHD96, GCML98, GTC^{+}98, HTS^{+}95, HS97, KKG95, KIT^{+}97, Kun95, OGS^{+}96, SZBB98, SCKG95, TBP^{+}99, WGF^{+}98, WSSM95, WMR^{+}96, XGC96, ZB98]. KDEL [MBS96a, LCM^{+}95, MSW^{+}98]. Kel1p [PH98]. Kelch [PH98, RC97]. Keratan [WSH98]. keratin [ADP^{+}95, BK96, KMOO95, LLK^{+}95, LYC^{+}95, PTBC96, CSO97, KMS^{+}98, MSL^{+}98a, MC98, PC98, PC99]. Keratinocyte [KMI^{+}97, ZSMV95, BGK^{+}95, HGW95, LDF^{+}95a, CCS^{+}98, PDS^{+}97b]. keratinocytes [PTBC96, ZMM^{+}95, EPMB^{+}99, HDDKH98]. keratins [LLK^{+}95, LO96, HPY^{+}98, PC99]. keratocyte [AWS96, SHA95]. Keratocytes [GS99, ODJ99, SVMB97]. Kex2p [BF97b]. Key [LDD^{+}97, CLT99, HLG^{+}98, LRG^{+}98, MP98, MBCS98]. KHP1 [PMH96]. Kic1p [SBR98]. kidney [DSM^{+}96, MMS^{+}96b, PJL^{+}95, WYY95, ZZvdB^{+}96, vGvM95, AKCC99, ALK^{+}99, BRS99, DCR98, DMY^{+}97, LMM^{+}99, LTMR97, LYD^{+}97, OT97, PMBM^{+}99]. kidneys [MS96]. KIF1A [YHO^{+}98, PdHTV99]. KIF2 [MQR^{+}97, NSYK^{+}95]. KIF21A [MWF^{+}99]. KIF21B [MWF^{+}99]. KIF3A [NSYK^{+}95, TYT^{+}99, YNOH95]. KIF3A/ [YNOH95]. KIF3A/B [NSYK^{+}95]. Killing [KOM^{+}98]. Kilodalton [FBM99]. Kinase [AGS^{+}97, AWC99, BSW^{+}99, CHP^{+}98, CJS^{+}99, CIK^{+}99, CLT99, DVL^{+}98, HTK^{+}99, sHWM^{+}99, KD97a, KNH^{+}97, KBS^{+}99, MMD^{+}98, MSH^{+}97, NCW^{+}99, SvDtK^{+}98, SGD98, SCR98, SKC97, SWVR98, SSR^{+}98, THM98, VMT^{+}99, WLD^{+}98, XC97, ZC99, AJS96, AS96, ASSS^{+}95, BGB^{+}95a, BDE^{+}95, BPT96, CMC^{+}98, Dav95, DRR96, FUB^{+}96, FVRCH96, FVKS96, GJ95a, GOT96, GW95, GMFN95, GHF^{+}95, GPRS^{+}95, ISFB^{+}96, KFE96, KSVZ95, LS95a, LN96, LCS96, LYR96, MTGY96, NWPW95, OIT^{+}95, OSH^{+}96, OHB^{+}95, PMCT96, PPBJ96, RL96, RDM^{+}96, SOHP95, SKS^{+}95, STT^{+}96, SHT^{+}95, SMFC95, SJHC96, SIAS96, SHPP96, SDTE95, SZB^{+}95, SP96, TOT^{+}96, TNY^{+}95, TCM96, Wal95, XZSC96, YDFB98, YYH^{+}96, AK97a, BSW^{+}98, CRM^{+}98, CAS^{+}98, CWMD97, CMS98, DCJ97, DSM^{+}97, EKSC98, EPMB^{+}99, FMM^{+}98, FKO^{+}98, FOK^{+}99, GWB^{+}98, GGPG99, GCD97, GLMG98, GTY98, HIW^{+}98, HYN97]. Kinase [HIWL97, IAS^{+}98, JL98, JH97, KHA^{+}99, KFO^{+}99, hKsKC99, KDO^{+}99, KCG^{+}97, KJR^{+}98, KGE^{+}98, KNT^{+}99, LPM^{+}98, LWSL^{+}98, LAS^{+}97, LFP98, MLB98, MYC^{+}98, MDC^{+}99, OTQM98, OWD^{+}99, OKE^{+}99, OWW^{+}99, PMKR97, PHN^{+}98, PBM^{+}97, PRV98, RTM99, RTP^{+}98, RPS99, RHLG98, SSI^{+}97, SWK^{+}99, SBC99, SJF^{+}97, SBGR98, SKY^{+}98, SMM99, SLS^{+}99, SBR98, SMTN99, SSH^{+}98, TGN97, TFS^{+}99, TFYH99, ÜHK^{+}99, VC97, WZF97, cXP97, YAiN^{+}98, ZCE^{+}98, ZTH^{+}97, ZDK^{+}99, ZRG98, vHGB^{+}97]. Kinase-induced [SKC97]. Kinase-like [KNH^{+}97]. Kinase-promoted [CHP^{+}98]. kinase-regulated [GW95]. Kinase/Protein [HTK^{+}99]. Kinases [HTK^{+}99, KND^{+}97, BGN^{+}97, DTB96, FWG^{+}96, LFB96, MTGY96, MA95, SWB^{+}96, vdBCH^{+}95, BATP99, BSM^{+}99, CCS^{+}98, CYSD99, HB99, PMP^{+}98, SMS^{+}98, WLF99]. kine [CBM^{+}98]. Kinesin [BML^{+}97, CDH^{+}98, Col99, DEPR97, GDB^{+}98, LAW^{+}98, LSCM^{+}95, PdHTV99, RPY^{+}99, Sch96, TYT^{+}99, VLA^{+}98, ASH95, KBR95, KKE^{+}95, LNY^{+}95, MMGT96, NH95, NSYK^{+}95, PMH96, SKE^{+}95, WA95, WFBC96, WRW^{+}95, WM95, YNOH95, APH^{+}97, Blo98, EGS^{+}98, HH98, LB98, MHWW98, MBW^{+}99, MWF^{+}99, MBB^{+}97, MS97b, MSH^{+}99, PYM^{+}98, RKRG99, RPV98, SHLD97, SMB^{+}99, Ste97, TZL^{+}98, WVM97, WBA^{+}98, WMR^{+}96]. Kinesin- [BML^{+}97, SKE^{+}95]. Kinesin-dependent [EGS^{+}98, PYM^{+}98]. Kinesin-Ii [Col99, RPY^{+}99, Sch96, WMR^{+}96]. Kinesin-II-dependent [CDH^{+}98]. kinesin-like [ASH95, KBR95, KKE^{+}95, MMGT96, WA95, WRW^{+}95, WBA^{+}98]. kinesin-mediated [RSR^{+}96]. Kinesin-related [DEPR97, WFBC96, WM95, APH^{+}97, LB98, MBW^{+}99, MBB^{+}97, SHLD97, WVM97]. Kinesins [HST^{+}98, MWF^{+}99, SSM98]. Kinetic [HME^{+}98, IMK99, ABN96, OP96b]. kinetics [BSD^{+}96, SRG^{+}96, ACP^{+}97]. Kinetochore [BMG^{+}99b, MHC^{+}97, NWG95, SWL^{+}97, THM98, ZKB95, KCBR96, RHF^{+}96, WGG96, CSY98, EKS97, HKKH99, KCM^{+}97, RGS99, SWHG98, YAC97, YMD99]. Kinetochores [CJS^{+}99, KR96, SVH^{+}98, LWM^{+}95, RCKS95, CSMM98, SCM^{+}97, WCMS98, ZCE^{+}98]. Kinetoplast [JE98]. KIP3 [DEPR97]. kit [VDA^{+}96, SBGR98]. KLP38B [APH^{+}97]. KLP3A [WRW^{+}95]. KLP61F [SMB^{+}99]. knock [RLC^{+}96]. knock-out [RLC^{+}96]. knockout [GCGE^{+}96, MSL^{+}98a]. KNOLLE [LWS^{+}97a]. known [LWB^{+}95]. Kre2p [LSKB95]. Kre2p/Mnt1p [LSKB95]. KRP85 [WMR^{+}96]. KRP85/95 [WMR^{+}96]. KSA [CCB^{+}98]. Kv2.1 [SMRT96].

L [CFS98, EFB^{+}98, ZLB^{+}97, MNS^{+}99, TOSF^{+}98, AFFS96, ACP^{+}97, BBL^{+}99b, BQF^{+}95, CTB^{+}97, FAP^{+}96, GCM^{+}97, KTAX98, PWG^{+}95, PFGS95, SCM^{+}96, WS96]. l-3-hydroxyacyl-CoA [TOSF^{+}98]. l-Afadin- [MNS^{+}99]. L-glutamate [WS96]. L-Selectin [CFS98, ZLB^{+}97, AFFS96, ACP^{+}97, BBL^{+}99b, FAP^{+}96, PWG^{+}95, PFGS95, SCM^{+}96]. L1 [BOLS95, BL99, DCR98, FHSM^{+}97, GRTB97, HSL^{+}99, HHM^{+}98, KEJ97, MAMJ95, MBS^{+}96b, RWM^{+}96, RAH^{+}95, SZBB98]. L6 [HLSM95, RLVM98]. L929 [DCVF98]. labeled [VB98, KGA^{+}97]. labile [LCM^{+}95]. lac [RSL^{+}96]. Lack [FPM^{+}95, HPL97, KD97a, LARH97, ZMG96b, LNR^{+}96, PJL^{+}95, KZS^{+}98]. Lacking [PJE^{+}99, EB95, JWH96, MWT^{+}96, MS96, NMJL96, RFNF96, SFD^{+}96, ASES^{+}97, HTK^{+}98, KPZ^{+}97, LYD^{+}97, LEB^{+}97, MCS98, MNG^{+}98, NKK^{+}99, TTT^{+}99]. Lactation [JWFS99]. laevis [AOMB96, FB96, GHW99, LGM98, NG99, SRH^{+}95, TGTL97, VLAHV98, VWG^{+}97, WVM97]. LAMA3 [RLMC99]. LAMC1 [SVM^{+}99]. lamella [VSB95, WSS97]. lamellae [MMF95, LBWF^{+}97, OSM98]. Lamellipodia [SB99d]. Lamellipodial [RSS98]. lamellipodium [SHA95]. Lamin [RPW96, YGG97b, CCP96, FKS95, LBWF^{+}97, SMG^{+}97, SEAB^{+}99]. Lamin-binding [YGG97b]. lamina [GYS^{+}95, DBG^{+}97, GHW99]. Laminin [CJY^{+}97, CWY99, KOA^{+}99, RKR^{+}97, SSZ^{+}97, SSB^{+}95, SGYH97, VNTM^{+}99, BKEC96, CLR^{+}96, DNB^{+}96, KKD^{+}95, LW96, VLX^{+}96, VKR^{+}96, XGC96, ZZvdB^{+}96, CAH^{+}99, DBG^{+}97, GSJ98, LNG^{+}98, MZL^{+}99, MPL^{+}97, MCS98, MLC99, OTS98, RM97, RYB^{+}99]. laminin-1 [KKD^{+}95, RM97]. laminin-4 [BKEC96]. Laminin-5 [GSJ98]. Laminin-induced [CJY^{+}97, SGYH97]. laminins [SEC^{+}95, MPL^{+}97, PMCS97]. lamins [TGG95]. lamp [HH95a, GF97, MSS^{+}99]. LAMP-2 [GF97]. Lamp1 [RSRK96]. lampbrush [BLG95, EJ99]. Langerhans [MBCS98, SLM^{+}95, WSR^{+}97]. LAP2 [GHW99, YGG97b]. Lapse [SSM98]. LAR [ALM^{+}97a, OTS98, ZL95a]. Large [CH98, FPRL97, KBM^{+}95, TMM97, TSB99, BDvdZW95, LGdC95, MCL^{+}96, MAG^{+}96, PZ96, RSL^{+}96, WTCV95, EFK^{+}98, GB98, KSOK97, RSD98]. Large-Scale [FPRL97, TSB99, CH98, RSL^{+}96]. Largely [GF97]. larger [NSR95]. larvae [LRAB99]. Larval [RV98]. Laser [WJH^{+}98, SK95a]. latches [Val96]. Late [BNJ99, KWD^{+}98, LSB98, VS98, AGPG96, BRJP96, CS96b, FPWN95, LWM^{+}95, VMV^{+}95, vWDG^{+}95, BRS98, BS97, CIRG99, MFF97, MBF^{+}98, NMT97, RC98, RRK97, RLMC99]. Late-Golgi [VS98, CS96b]. Late-Replicating [LSB98]. Latent [NGMR97, DMS^{+}95, NMK^{+}97, NGMR97]. Lateral [BYG96, YMAC^{+}98, CSN99, SK95a, CT98, GRTB97, TSK98, YNG98]. Laterality [TYT^{+}99]. Latrunculin [ASP^{+}97]. Lattice [XC97, CMHT96, SM95b, DZS^{+}98, XZSC98]. lattices [LRM^{+}95]. Layers [GB98]. Layilin [BH98]. LBR [DFL^{+}99]. LBR-like [DFL^{+}99]. Lc7 [BPKB^{+}99]. LCAM [MEYPR95]. LD4 [TBP^{+}99]. LDL [DFD^{+}97, HE96, KsYA98, KRGA98, KSL^{+}95, CNP^{+}95]. LDL-deprivation [HE96]. Lead [SB99b, TFYH99]. Leading [BMC^{+}99, Blo98, DSvdV^{+}96, KGHMT95, BKJL^{+}98, SEAB^{+}99]. leads [BL96, DNB^{+}96, KJBM^{+}96, SP96, DBH^{+}97, HLJ^{+}98, LGM^{+}97, RPS^{+}97, WWO^{+}98]. Leaflets [MBOC97]. least [DPHW96, MCS^{+}97a]. Lectin [VKIH98]. Lectins [BH98, FHY99]. leech [BSJJ95]. LeechCAM [HJJJ97].

Left [TYT^{+}99]. Left-Right [TYT^{+}99]. Leishmania [PXR^{+}95, SL97]. length [KCBR96, PNM95, HSLB99, SHLD97]. Lens [LKD^{+}95, DSvdV^{+}96, BM97, IJR98]. lesioned [SRR96a]. lesions [BSK96]. less [SEC^{+}95, MMS96a]. Lessons [Lan97, MSL^{+}98a, FN98]. lethal [LRAB99, Bri99, GJB^{+}98]. lethality [GAW^{+}96, SVM^{+}99]. Leu [MBS96a]. leucine [DBE98, MWOB96, SVB^{+}97, WBA^{+}98]. Leucine-rich [SVB^{+}97]. Leukemia [BHD^{+}97, MMAK97, SHG96]. Leukocyte [YWW^{+}96, DZC^{+}96, DAP^{+}95, LXK^{+}95, APF^{+}97, CS99, CWT^{+}98, FCB97, LKKL97, SSEM98]. leukocytes [AFFS96, FAP^{+}96, PHU^{+}95]. level [TEL^{+}95, WPZ^{+}96]. Levels [CSF^{+}97, BGN^{+}96, HGW95, EFB^{+}98, HFP^{+}98b, LKKL97, SSR^{+}97]. Lewis [FHY99]. LFA [DFC^{+}96, SWVR98, SMH98]. LFA- [SMH98]. LFA-1-dependent [SWVR98]. LFA-3 [DFC^{+}96]. lgp120 [HH95a]. Lhs1p [SHC^{+}97]. LI [KBB^{+}97b]. LI-Cadherin-mediated [KBB^{+}97b]. libraries [PKR95]. Library [RST^{+}99]. licensing [YN95, MCC^{+}97]. Lies [ANT^{+}97]. Life [Ols99]. Ligand [AKM^{+}98, FSDA97, GRB^{+}97, MN98, SRR96b, BEZ^{+}95, DAP^{+}95, FLW^{+}95, FHV^{+}96, HGH96, KCN^{+}96, KBM^{+}95, KP96, MPB^{+}95, PHU^{+}95, PFGS95, RAH^{+}95, SCM^{+}96, WHH96, WKRB96, BxQK^{+}99, DBOM98, GGD^{+}97, GTC^{+}98, KBC99, OTS98, SCH^{+}98]. ligand-1 [MPB^{+}95, GGD^{+}97, SCH^{+}98]. ligand-binding [HGH96, WHH96]. Ligand-induced [SRR96b, FLW^{+}95]. Ligands [WMP^{+}98, FAP^{+}96, MBS^{+}96b, BBL^{+}99b, CFS98]. Ligase [CJR^{+}97]. Ligation [SKC97]. Light [AKM^{+}98, KBS^{+}99, NCW^{+}99, VLA^{+}98, YNM^{+}99, BRHD95, GW95, KAT^{+}95, BRLM98, EFB^{+}98, GDB^{+}98, HOCK98, JK97, MAE^{+}97, MSEP98, PKBHK97, PWW98, PTVD99, RKRG99, SCEB99, SD98, TWP98]. Light-dependent [YNM^{+}99]. like [ASH95, BLE^{+}96, BBOE98, BMPP99, BIR96, BHBB99, CLR^{+}96, CCP96, DvD95, DFL^{+}99, EZC^{+}97, ED96, FHSM^{+}97, GSC95, GSP^{+}98, HM96, HRK^{+}98, HIG^{+}95, JSR96, KNH^{+}97, KBR95, KKE^{+}95, LN96, LL98b, MCS^{+}97a, MN98, MSH^{+}98, MSS^{+}99, MJMS96, MFBK99, MMGT96, MCB96, MW97, NIM95, NCV^{+}98, OF98, PFKvD99, RSD98, SHH97, SVB^{+}97, SY99, SIAS96, SGGS96, SV98, TNM^{+}99, TSS^{+}95b, TMN97, WA95, WRW^{+}95, WBSNV97, WBA^{+}98, WJWM^{+}98, XWK^{+}97, YtDH^{+}95, YPDM98, YSK^{+}97, ZLPL99]. LIM [HMC^{+}99, PLB97, SSB96, XWK^{+}97, SMTN99]. LIM3 [BPT96]. limb [OST^{+}95, CBS^{+}97, MBTW97, YSEI^{+}99]. Limit [SHLD97]. Limited [KEP^{+}99, KGS98]. limits [GGC96]. LIMP [KJE^{+}99]. Limulus [WSG^{+}95]. LIN [WG99]. LIN-12 [WG99]. LIN-2 [CWM^{+}98, HYK^{+}98]. line [RVD^{+}95, SSA^{+}96, CLF^{+}97b, GKS^{+}98, MAC^{+}97, MNG^{+}98, BTTB96]. lineage [ABRB^{+}95b, MNM^{+}98, VGVF97, WDL^{+}98]. Lineage-specific [WDL^{+}98]. Lineages [CJH^{+}98, MC98]. lines [GTL^{+}96, PPF96, PFP97, SEH^{+}97a, SATW98]. link [HTKH96, NWG95, BRP^{+}97, GWB^{+}99, HKKH99, IYM98, KGL^{+}99, YAC97, TCV^{+}98]. linkage [YWW^{+}96, BALL98, GKB^{+}98]. linkages [HGH96, FZZN98]. Linked [HFP98a, ZOB^{+}96, ZDK^{+}99, ANS^{+}97, DvD95, RJP^{+}96a, SM95c]. linker [SJA96, GCML98, GB98, LM98]. linkers [TCV^{+}98]. Linking [RdPRW98, AKP95, LMB^{+}95, WSG^{+}95, ANK97, LSZ^{+}99a, NGMR97, GWB^{+}99]. links [TTG95, MSH^{+}99, BFHB97, SMB^{+}99]. Lipid [HSVS98, SBS^{+}99, SLS^{+}96, GCE^{+}97, JLM99, LLVS98, MSM99, OPH^{+}99]. Lipid-modified [SLS^{+}96]. Lipidic [CFL^{+}98]. Lipids [CLF^{+}97a, CFL^{+}98, vHMV^{+}97]. lipolytica [ESR97, STVR95, TSSR98]. Lipopolysaccharide [BMG^{+}99a]. lipoprotein [FFZ^{+}95, GLP^{+}95]. Liposome [HPD^{+}97, KKM^{+}97]. Listeria [MMP^{+}95, MGP97, RBM^{+}99, RAA^{+}97, SWS97a, STP96]. Live [DKD^{+}96, WJM^{+}99]. lived [MLJM95]. liver [ADP^{+}95, MKB^{+}95, SS95a, ESH^{+}99, MSL^{+}98a, SATW98]. liver-enriched [SS95a]. Liver-specific [MSL^{+}98a]. living [JCWP95, KLW^{+}96, MH95b, OMO95, SHS96, ZN95b, CWT^{+}98, ESM^{+}97, HKBM99, HTY^{+}99, HME^{+}98, MKC99, NTH98, OTQM98, SSI^{+}97, SWM^{+}98, SPS^{+}97a, SHP^{+}97, SWO^{+}99, SM97, WWD^{+}97, XZMD97, YMPG98]. loaded [QPH^{+}98]. Loading [FRT^{+}98, FKWP97]. Lobuloalveolar [JWFS99]. Local [KBSW95, WWP^{+}96, MMP^{+}95, BAC97, EKF^{+}97, mLZB99]. Localization [AK99, DSM^{+}97, HKJ^{+}99, HW96a, HKO^{+}98, JWW^{+}99, LSKB95, MSH^{+}97, NAsG^{+}98, SGTH99, THM98, WLD^{+}98, WLF98, WCMS98, WBA^{+}98, ZZYG96, ASH95, AM95a, BCMK95, BZB^{+}96, BPT96, CGM^{+}95, CS96a, CPL^{+}95, CSDR96, DC96a, GLS96, GMFN95, HSL^{+}95, HFSHK96, ISFB^{+}96, JCWP95, KAP95, KWCS96, KFE96, KF95, MAG^{+}96, MML96, MBB96, OGS^{+}96, OWA95, PGC^{+}95, PMCT96, RPC^{+}95, RSC96, RLCM96, RSL^{+}96, SSN96, SD95, STP96, TKVC95, WMR^{+}96, APF^{+}97, AAD^{+}97, ALVMM98, ASB^{+}99, AL97, BBP^{+}98, BRLM98, BF97a, BKN^{+}97, BF97b, BKJL^{+}98, CMS^{+}97, CMC^{+}98, CCBL98, CBB^{+}99, DCJ97, DPG^{+}99, DGP99, FhJZ^{+}99, FHAP98, GRS^{+}99, HSW^{+}98, JHM98, KLH98, KcZS97, LMF98, LHR^{+}97, LE97, MOY^{+}98, MM97b, MMR99a, NRD98, OAB^{+}98, PPP^{+}98, PSPG^{+}98, PCQH98, RHLG98, SSR^{+}97, SGJP99, SJZW97, SV98, TFS^{+}99, WOM98, qXLCH97, hYGYF97]. Localization [YMD99, ZCE^{+}98, ZSB99, SSSKR95]. localizations [KSG^{+}96]. localize [GG95b, KLN^{+}96, BS97, KHA^{+}99, MSFA99]. Localized [DSVL^{+}98, ISR99, MNS^{+}99, WDV^{+}97, YMAC^{+}98, CBB^{+}96, CM95a, FC95, SH96, SLW^{+}96, SLEB95, SM95c, UHJ^{+}96, VGM96, VH96, WS95a, AK97b, BH98, CC98, DAF^{+}97, EJ99, FCTS99, FFLP99, MNS^{+}97, MMR99b, ONI^{+}98, YDFB98, ZS98, CES^{+}98]. Localizes [CWM^{+}98, ETF95, NAP^{+}96, SL96, SRRB95, TN95, MKB^{+}99, MPT^{+}98, SR97, WBG^{+}97]. Localizing [FFH^{+}98]. Locally [CAS97]. located [PPBJ96, SOR^{+}95, GCML98]. Location [PGP^{+}99, MKF96, HZC^{+}97, STG^{+}99]. locations [FGA^{+}96, MPL^{+}97]. locomote [RFNF96]. Locomoting [CSM97, ODJ99]. Locomotion [MOY^{+}98]. locus [ASSS^{+}95, PLM^{+}96, VXBH95]. long [ACHG95, MLJM95, DSG98, MCBE98, SWS97a]. long-lived [MLJM95]. long-term [ACHG95]. longitudinal [NCZ^{+}95, RV98]. longlasting [FJT^{+}95]. Loop [AGS^{+}97, DMG96, EJ99, FZZN98, KGHG98]. loops [BLG95, YvdEH^{+}95, dBCKS98]. Loss [KMI^{+}97, MSB^{+}97, MS97a, SEAB^{+}99, FSR^{+}96, MHLB96, PHA^{+}98, SFD^{+}96, KsYA98, KKG95, MSL^{+}98a, NLM^{+}98]. Low [MAS^{+}97, ZR98, FFZ^{+}95, GLP^{+}95, CLF^{+}97a]. low-density [FFZ^{+}95]. Lower [HFE^{+}98]. loxP [GBvdN^{+}98]. LRP [GLP^{+}95, CNP^{+}95]. LST1 [RCE^{+}99]. LT [DCVF98]. LT- [DCVF98]. Luciferase [TKBM97]. lumen [GG95a, SLW^{+}96]. lumenal [BCC^{+}95, MLJM95, CS97b]. Lumican [CML^{+}98, CML^{+}98]. Luminal [WBKB97, TT96]. lung [HWC^{+}96, SSZ^{+}97]. Ly [BGK^{+}95]. Ly-6 [BGK^{+}95]. Lymph [BNW^{+}99]. Lymph-specific [BNW^{+}99]. lymphoblastoid [LR96]. lymphoblasts [GMS^{+}96, KMG^{+}97]. Lymphocyte [CBM^{+}98, PLR^{+}95, JBD^{+}95, YYH^{+}96, dPSMNSM95, CES97, GBvdN^{+}98, PH97b]. lymphocytes [CAS96, DTB96, HSL^{+}95, ANS^{+}97, CFS98, HFL97, LWSL^{+}98, SALdP^{+}97, dPCM^{+}97]. lymphoid [CQB^{+}96, SLS^{+}99]. lymphosarcoma [DSE^{+}95]. Lymphotoxin [DCVF98]. Lyn [SHB99]. Lyn-mediated [SHB99]. Lys [MBS96a]. lysates [BK96]. lysin [SFSV95]. lysine [GCB^{+}95, SSSKR95]. Lysophosphatidic [YFST^{+}96, HWPC96, CNG98]. Lysosomal [WBKB97, BDE^{+}95, GTL^{+}96, HH95a, KDvFP96, LCH96, MRvD^{+}95, dPF95, SKR96, WFBO^{+}96, ATD97, KKLA97, TFYH99, VKIH98]. Lysosome [BOS^{+}97, VG99, NH95, ZMG^{+}96a, BGN^{+}97, GF97, HTK^{+}98, WLK97]. Lysosome-associated [GF97]. Lysosomes [HPL97, NCV^{+}98, RWOA97, Dav95, FPHH96, OS96a, RSR^{+}96, RSJK95, RSRK96, TBR^{+}96, vWDG^{+}95, AYP^{+}99, BRAM97, KsYA98, MBF^{+}98, SPBCM97]. LYVE [BNW^{+}99]. LYVE-1 [BNW^{+}99].

M [KPKWW95, BTTB96, BMG^{+}99b, DFGL96, KTAX98, MM97a, NCZ^{+}95, OGS^{+}96, OSH^{+}96, OHB^{+}95, TGN97, TXB^{+}99, WKK^{+}95, WMC^{+}95b, ZKSB96]. M-line [BTTB96]. M-phase [DFGL96, OSH^{+}96, OHB^{+}95]. M-protein [OGS^{+}96]. m/Odz [OZF^{+}99]. M2 [SLL96]. M9 [IJB^{+}97]. MA104 [SMA96]. MAC [SRR96a, DAP^{+}95]. Mac-1 [DAP^{+}95]. MAC-2 [SRR96a]. Macf [LSZ^{+}99a]. machineries [MBRN95, ST96]. machinery [MGS96, RACHV96, STVR95, WJC^{+}96, AWB^{+}98, DSC^{+}99, HRA^{+}98, KBG^{+}99, KS97, LCW^{+}98, WFL^{+}99]. Macroglobulin [CNP^{+}95]. Macroh2a1.2 [MCPB99]. Macromolecular [BS95b]. Macromolecule [SBV97]. Macromolecule-sized [SBV97]. Macrophage [CSF^{+}97, HPL97, OS96a, SRR96a, AZRJ98, GM99, PHN^{+}98, SSV98, WDL^{+}98]. Macrophage-derived [GM99]. macrophages [AJS96, BMS95, KSL^{+}95, SRR96a, WPLK95, ZMG^{+}96a, WLK97]. macropinocytosis [AJS96]. mad [VKP^{+}95]. Mad1p [HM95]. MAD2 [YMD99, FFLP99, GCM98, KWD^{+}98, WCMS98]. Mad3 [THM98]. Mad3/Bub1 [THM98]. Mad3/Bub1-related [THM98]. Madin [ALK^{+}99, BRS99, DMY^{+}97, OT97, PMBM^{+}99, ZZvdB^{+}96, vGvM95]. MAG [BQF^{+}95]. Magnified [FPRL97]. MAGUK [HGW^{+}98]. Maguks [IFM^{+}99]. Maintained [TLF^{+}97]. Maintaining [LRC^{+}99]. Maintains [CGA^{+}99].

Maintenance [CRG^{+}95, HLMD97, MAM99, HG95b, MMP^{+}95, MW96, PFF^{+}95, ASP^{+}97, BHHF97, BENV97, EIIM^{+}98, GWB^{+}98, MS97a, MCBE98, MCS^{+}97b, SO97, SO98, TH99, TRS^{+}99]. Maize [WRC^{+}99, YHC^{+}97, YMD99]. Major [CT97a, HGP^{+}95, ISR99, KMG^{+}97, STF98, SKC97, AKP95, AM95a, BK96, BGB95b, PJL^{+}95, PFGS95, RvSL^{+}95, SOR^{+}96, BRAM97, FKWP97, HGI^{+}99, KMS^{+}98, LMB98, MC98, RSS98]. making [OMW^{+}96]. MAL [PMBM^{+}99]. Mal3 [BHHF97]. MALDI [WJH^{+}98]. male [GDF^{+}96, BGG98, GCA^{+}98, LXC^{+}98]. Malignancy [MLN^{+}98]. malignant [JH96, WPW^{+}97, YKO97]. Mammalian [GXE^{+}99, KWD^{+}98, KNT^{+}99, LARH97, MLO^{+}95, SYO95, TOSF^{+}98, VGM96, WLD^{+}98, WDL^{+}98, ZWT^{+}97, AZF96, BDE^{+}95, DC96a, EKPK95, FDH96, GSB^{+}96, HM96, KKE^{+}95, LG95, LvSC^{+}96, RHF^{+}96, RAM^{+}96, SLS^{+}96, SBSG96, SM95a, SJ95, SCK^{+}95, SRRB95, TRJ^{+}95, WS95a, WMXE96, ZZYG96, ALVMM98, BGN^{+}97, BCWA97, DTMG99, FCTS99, FHWV97, GCD97, GCM98, HRR^{+}99, IHT^{+}98, KGL^{+}99, KGGK99, LLNH^{+}98, MSD^{+}98, MSH^{+}99, MUS98, NZ97, NIK^{+}97, OC98, POR^{+}97, PFP97, SLF^{+}99, SBC99, WAV98, WWD^{+}97, YPDM98, ZBW^{+}97]. Mammals [HBH98]. mammary [AHBW96, FMJ^{+}95, FSR^{+}96, LCK95, LDCB95, LLRTP95, MBF^{+}95a, SSB^{+}95, YSM^{+}95, FLO^{+}99, HLG^{+}98, IFSV^{+}98, JWFS99, LGM^{+}97, NBS^{+}98, RFCR^{+}97, WWO^{+}98]. man [SWS^{+}96]. Manganese [LK95]. Manner [FST99, NCW^{+}99, SHK^{+}99, FSHD96, RHF^{+}96, SH96, GKS^{+}98, MLN^{+}98]. Mannose [KKG^{+}98, LH97, NCV^{+}98, KDvFP96, dPF95, RSJK95, SKR96, PFHWN98]. mannosylphosphoryldolichol [RW95]. mannosyltransferase [HW96a, LSKB95]. Many [OZF^{+}99, ZM95, FHWV97, MLB98]. MAP [GTY98, MK95a, LPM^{+}98, LAS^{+}97, MTGY96, OMO95, RHLG98, SKS^{+}95, SP96, TGN97, ZCE^{+}98, BATP99, RPS99]. MAP1B [TKH^{+}97, TWP98]. MAP4 [OHB^{+}95, WPZ^{+}96]. MAPK [GKO99, LBBP98, TCM96]. Mapping [CJR^{+}97, FBH^{+}95, CLL95, GDR95]. MARCKS [LYLR98, MLB98]. MARCKS-related [MLB98]. Markedly [KKH^{+}99]. marker [BSN95, LG95, SOL^{+}95, HHB^{+}97, MM97a, SPS97b, SW98]. Markers [DBC^{+}99, MFF97]. Marrow [NMT97]. Mas37p [GLB^{+}95]. masking [LHDC95]. Mass [SBS^{+}99, WJH^{+}98, FPM^{+}95]. Massive [RME^{+}95]. Mast [FGA^{+}96, PHB96]. mate [DPH95]. Maternal [WBSNV97, BLG95]. mating [BGN^{+}96, VPH95, VH96, CSCM97, ELMS98, MCM^{+}99, NA99, PH97a, WFS97]. matricellular [Bor95]. matrices [HG95a]. Matrix [CHLÅ97, HSML98, KSF^{+}99, LTMR97, LGM^{+}97, MNS^{+}99, RMY97, SvDtK^{+}98, SPV^{+}99, SLS^{+}99, WJH^{+}98, YSEI^{+}99, AKP95, AC95b, CSO^{+}95, DMS^{+}95, DGSR95, HMB95a, HSML96, HH95b, KLMR95, KvEdV^{+}96, LWM^{+}95, LDCB95, LF95, LvSC^{+}96, MGS96, MHK^{+}95, NRW^{+}95, SSY^{+}96, STS96, STVR95, WKL^{+}96, WvdVV^{+}95, WPMW95, XC96, dPSMNSM95, BPS99, BRP^{+}97, BZBA99, BOM^{+}98, CMD^{+}98, CWSG99, EJ99, FLO^{+}99, FTB97, GCAHW97, IAS^{+}98, IKR^{+}98, KLO^{+}98, KNSP97, KPZ^{+}97, LGZ^{+}99, LDD^{+}97, MSB99, MH97, MSB^{+}97, MBCS98, MLC99, ONI^{+}98, OTL^{+}97, PÖS^{+}98, PDS^{+}97b, SSSS97, SB99c, WSH98, ZCWB^{+}98, ZOKS99, dBCKS98]. Matrix-assisted [WJH^{+}98]. matrix-associated [CSO^{+}95]. Matrix-dependent [SvDtK^{+}98]. Matter [BPS99]. maturation [ABRB^{+}95b, GHJJ95, GBG^{+}96, LN96, LG95, MAG^{+}96, NDBR95, OSH^{+}96, ZL95c, ANT^{+}97, DBLD^{+}99, EY99, GR97, MHC^{+}97, MZL97, PC98, RPB97, TNFK^{+}98, UPT98]. maturation-promoting [OSH^{+}96]. mature [FPHH96, HFSHK96, TBR^{+}96, AMY^{+}98, BxQK^{+}99]. Maturing [CHC97, MWL97]. May [FWL^{+}98, CK96, HW96b, NWG95, GWB^{+}99]. Mbs [KFO^{+}99]. MCC [LJK97]. MCF [ARR^{+}98]. MCF-7 [ARR^{+}98]. Mcl [YKC95]. Mcl-1 [YKC95]. MCM2 [TAK95, DTMG99]. MCM2-3-5 [TAK95]. MDAY [DSE^{+}95]. MDC9 [WKRB96]. MDCK [JN98, BPA^{+}97, CSN99, CS98, FGA^{+}98, GFM^{+}98, HKC^{+}95, HE96, HH95a, KE98, LNRR98, LCW^{+}98, MSF^{+}95, NDBR95, OPD^{+}96, SVF^{+}98, TSK^{+}97, YLB^{+}97, MSFA99]. Mdm10p [BVKP98]. Mdm12p [BSY97]. Mdm1p [FY97]. MDM20 [HKS97]. mdx [CRM^{+}96, VWK^{+}95, WB99]. Measured [LCD^{+}98]. measurement [KLW^{+}96]. Measurements [HFP98a, HD99]. MEC [LHK^{+}96]. MEC-4 [LHK^{+}96]. MECA [CFS98]. MECA-79 [CFS98]. Mechanical [DFD^{+}99, PWE95, TLBK98, HKBM99]. Mechanics [ACP^{+}97]. Mechanism [ABN96, BSW^{+}99, HUU96, KTD^{+}97, LARH97, LKSW98, MZL97, ZJY^{+}98, AM95b, GGC96, HSK^{+}96, HTKH96, JSR96, LYR96, MCL^{+}96, MDQG96, OPD^{+}96, OCS^{+}95, SS95a, SRR96b, TT96, WMC^{+}95a, WS95b, BATP99, IRW^{+}99, NMT97, NB97, PRB99, SBOK97, SPR^{+}98, SYM^{+}97, SVMB97, TSK98, VC97, dPCM^{+}97]. Mechanisms [ACSN98, GF95, KKGD96, MYTK99, HKM^{+}96, MSF^{+}95, SOH^{+}95, SSN96, VRD96, CCBL98, GFM^{+}98, KQB^{+}98, MHC^{+}97, MC97, STK^{+}98, SFA^{+}98, VvZ99]. mechanistic [ST96]. Mechanochemistry [DWS^{+}95]. Mechanosensory [AAM^{+}98, HMC^{+}99]. mechanotransduction [LHK^{+}96, MSWL^{+}99]. media [LSKB95]. media-Golgi [LSKB95]. medial [WBG^{+}97, OPB98]. medial/trans [WBG^{+}97]. Mediate [HRD98, FAP^{+}96, HW96b, PNM95, SVB96, BRS99, KDG98, LNH^{+}98, MMM^{+}98, SY99, ZGYS99]. Mediated [MSH^{+}97, SSR^{+}98, BEZ^{+}95, BLE^{+}96, BLL95, CMUW96, CKB^{+}98, CCP96, DPHW96, DRR96, DCW^{+}96, DG96, DMD^{+}96, FUB^{+}96, FFZ^{+}95, FMP^{+}97, GBvdN^{+}98, GG95a, GSvR^{+}96, HKOM98, HDV^{+}99, HSML98, HGH96, IAS^{+}98, JLRS95, KBS^{+}99, KBB^{+}97b, LTVC96, LAS^{+}97, LSCM^{+}95, MXSRB96, OWD^{+}99, PIB^{+}99, PW98, PXR^{+}95, PH97b, PPW^{+}95, RWM^{+}96, RCKS95, RSR^{+}96, RW95, SKS^{+}95, SEK^{+}97, SBR^{+}96, SEH97b, SEC^{+}95, SDTE95, SMH98, VUC^{+}95, WvdVV^{+}95, XGC96, XZSC96, YHC^{+}97, ZPS96, ZCWB^{+}98, BATP99, BKN^{+}97, CRM^{+}97, CLF^{+}97a, CLH^{+}99, GAZ^{+}98, GBW98, GXE^{+}99, KSG^{+}98, LBW^{+}97, MTLGC99, MPSC98, OFM^{+}98, SGGW99, SBD^{+}99, WLF99, ACP^{+}97, ABN96, ABL98, BNA^{+}96, BLB^{+}98, CGC^{+}97b, DOP^{+}98, FSIG^{+}95, FJ98, GM99, GTY98, HCW^{+}98, KsYA98, LD95, LDD^{+}97, MSW^{+}98, NHB^{+}95, PPP^{+}98, PHB96, SHB99, SM97, SLS^{+}99, ÜHK^{+}99, WYL^{+}99]. Mediates [KWD^{+}98, MSF^{+}99a, WSCN97, CRLM96, FLW^{+}95, KBW^{+}96, LR96, MBM95, MPB^{+}95, PCL^{+}96, SSB^{+}95, AYP^{+}99, BMP^{+}98, CCB^{+}98, CNG98, CS97b, CRA^{+}98, DPZ^{+}97, HSL^{+}99, HAB^{+}99, IRW^{+}99, KMM^{+}98, KHS^{+}97, KGE^{+}98, LSH^{+}99, LWSL^{+}98, MZNR98, MNM^{+}98, NJB^{+}98, OMS98, OC98, PWGG98, PKCS98, RPS99, RME98, SAC^{+}98, SDDV97, SEBF98, TCYS99, TPM^{+}99, TRA^{+}99, vMNS97]. Mediating [WLD^{+}98, TLF^{+}95, BM98, NIK^{+}97, RM97]. Mediation [PRB99]. Mediator [CHP^{+}98]. medium [BLB^{+}96]. Megabase [RBRB99, YvdEH^{+}95]. Megakaryocytes [ILSH99, NMT97, SRA^{+}99]. MEI [MPT^{+}98]. Meiosis [BSD97, BGG98, EK97, LXC^{+}98, RME98, WHB^{+}99]. Meiotic [BMG^{+}99b, LARH97, MMGT96, SWS^{+}96, TMM^{+}95, BMS^{+}97, HOCK98, wJTSSL98, MCTMK98, MPT^{+}98, SR97, TMPM99, WHB^{+}99, YWMH99, YMD99]. Meiotically [HFE^{+}98]. Melanocytes [BAH^{+}98, WBR^{+}98]. melanogaster [BGG98, FMD^{+}98, MMN^{+}99, YF99]. melanoma [BBOB96, FBD^{+}95, SKA^{+}96, CLF^{+}97b]. Melanophores [RTP^{+}98, TZL^{+}98]. melanosomal [VXBH95]. Melanosome [RKR^{+}99, WBR^{+}98]. Member [BDO^{+}97, KGF^{+}97, BOLS95, LVD96, PJL^{+}95, TAK95, ZL95b, CF99, GSP^{+}98, HGW^{+}98, HHM^{+}98, MPLS^{+}98, NLM^{+}98, SVB^{+}97, SSH^{+}98]. Members [KKFN^{+}95, CKB^{+}98, DDF^{+}98, DDWM97, GvdFvD^{+}99, HJJJ97, RBB97]. Membrane [BB97, BPS99, CSN99, CLK99, CWM^{+}98, CLH^{+}99, DPHW96, FKNM97, GK99, GRB^{+}97, vtHR97, KOA^{+}99, KSF^{+}99, LAW^{+}98, MBA97, MM99b, MPSC98, MSH^{+}97, MZN97, NIM95, NAsG^{+}98, PPWL97, PBS97, PdHTV99, RS99, RST^{+}99, SBS^{+}99, SME98, VS98, WS96, WHH96, WLF98, WB99, WC97b, YHS^{+}99, YTT99, ARM^{+}95, ASY96, AHCC95, BWF^{+}96, BWG95, BGB^{+}95a, BAS95, BLL95, BKB^{+}96, CF95, CCP96, CSY^{+}95, DNB^{+}96, DWS^{+}95, DMD^{+}96, DSvdV^{+}96, DMG96, EKK^{+}96, EB95, EB96, FPWN95, FLPS95, GHJJ95, GLS96, GKM^{+}96, HIG^{+}95, HSK^{+}96, KBR95, KSL96, LSCM^{+}95, MBS96a, MCL^{+}96, MTL^{+}96, MKB^{+}95, MTLW95, MM95b, NH95, NAP^{+}96, NW96, NBW96, NMHS95, NKI^{+}95, NCS95, PNM95, PAV^{+}96, PHO^{+}95, dPF95, RJP^{+}96a, RR96, RSRK96, SK95a, Sch96, SLS^{+}96, SW95, TWVC95, VE95, VXBH95, VJHR96, WGR^{+}95, WTCV95, WPMW95, XGC96]. membrane [YNOH95, YKRS96, vWDG^{+}95, vdBCH^{+}95, APF^{+}97, AvdG99, BBAN^{+}99, BINRM97, BM99, BVKP98, BH98, BF97b, BS97, BENV97, BKJL^{+}98, CHW^{+}99, CLF^{+}97a, CFL^{+}98, CSEM99, CMC^{+}98, DSvDH^{+}98, DHDJ^{+}97, DSM^{+}98a, EZC^{+}97, ESR97, ESM^{+}97, FCMN97, FHWV97, FFH^{+}98, GWB^{+}98, GFM^{+}98, GWLH98, GBS99, GF97, HLJ^{+}98, HSVS98, HPD^{+}97, HME^{+}98, HHR^{+}98, IMS^{+}98, JR98, KFM99, KKM^{+}97, KGA^{+}97, KHS^{+}97, KEP^{+}99, KBF^{+}98, KCFS98, KAS^{+}99, KSY^{+}98, LRAB99, LJK97, LCW^{+}98, LHS97b, jLC97, MAC^{+}97, MS97a, MBOC97, MSEP98, MQR^{+}97, MJBR98, NTH98, Nei98, NZ97, OMS98, OTL^{+}97, OK98, PSPG^{+}98, PDO^{+}99, PKCS98, QPH^{+}98, RD97, RKB^{+}98, RRK97, RCE^{+}99, RSS98, RB98, RME98, SNT^{+}98, SSS99a, SHH97, SY99, SBS^{+}98, TMM97, TSK98, TSH^{+}98, TFYH99, TLBK98, VCL^{+}98, VL99, WGF^{+}98, WSGPS95, WWO^{+}98, WJWM^{+}98, YGG97a, YHD^{+}98, YPM97]. Membrane [ZSP^{+}98, ZP98, ZH97, ZBW^{+}97]. Membrane-anchored [NIM95, HIG^{+}95]. Membrane-anchoring [WWO^{+}98]. Membrane-associated [FKNM97, KOA^{+}99, PPWL97, XGC96, KAS^{+}99, RSS98]. Membrane-bound [FHWV97, ZBW^{+}97]. Membrane-organizing [WLF98]. Membrane-proximal [OK98]. Membrane-type [BPS99]. Membrane/ [WSGPS95]. Membranes [DFD^{+}99, MBA97, SBS^{+}99, AMJM95, BAC^{+}96, CCP96, DMP^{+}96, GHJJ95, LLL^{+}96, LD95, LW95, LHS96, MWC95, NAV96, ST96, SPT^{+}95, TOT^{+}96, WSGPS95, ANK97, BGL^{+}99, BG98, CGA^{+}99, FTB97, GJW^{+}97, HCM^{+}97, ODB98, SPL^{+}97, SHB99, SVM^{+}99, SLL^{+}98, TSK98, VvZ99, ZR98]. membranous [BSN95, NSYK^{+}95]. merlin [MF96b, LMF98]. meromyosin [BRHD95]. Merosin [VLX^{+}96]. Merzbacher [GSL98]. mesenchymal [MLLT95, MMMT95, CJH^{+}98, LGM^{+}97, NMK^{+}97, RYB^{+}99, SYT97, SSZ^{+}97, TTAH^{+}99]. Mesenchyme [MBTW97]. Mesoderm [TYT^{+}99]. mesodermal [PNL^{+}95, PGD^{+}97]. messenger [KKAN^{+}96, SKWY95, KcZS97]. metabolic [DVD^{+}96]. Metabolism [HCLG99, KBF^{+}98]. Metalloprotease [FKNM97, BPS99, DMY^{+}97, WPMW95]. metalloproteinase [HHM^{+}95, CMD^{+}98, LGM^{+}97, MFBK99, MLN^{+}98, MBCS98, OTL^{+}97, SB99c]. metalloproteinases [AHBW96, LTMR97, OTL^{+}97]. metalloproteinases-1 [AHBW96]. metanephric [WKL^{+}96]. metanephros [WKJH^{+}95]. Metaphase [LARH97, GACM^{+}95, KCBR96, STT^{+}96, TMM^{+}95, ZKB95, DWM^{+}98]. metaphase-anaphase [STT^{+}96, ZKB95]. Metaphase/Anaphase [LARH97, TMM^{+}95]. Metastable [TWS97]. Metastasis [CSC^{+}98]. Metastasis-associated [CSC^{+}98]. metastatic [DSE^{+}95]. Metazoan [OMW^{+}96]. method [LCH96]. methyl [WPS^{+}96]. methylation [TFF^{+}95, CEZA97]. methylene [HCAW95]. methyltetrahydrofolate [SMA96]. Methyltransferase [CL99]. Mg [EAOS^{+}98]. Mgm101p [MTW^{+}99]. Mgm1p [SY99]. MHC [FKWP97, LMB98, FRT^{+}98, GMS^{+}96, MRvD^{+}95, SKC97, WLR96, WFBO^{+}96, DBLD^{+}99, SSRP99]. MHP1 [IFHR^{+}96]. MIB [JOPM98]. Mibp [LMW99]. Mice [AYF96, CCMG97, KMI^{+}97, KZS^{+}98, PJE^{+}99, TYT^{+}99, WB99, ADP^{+}95, BSK96, BQF^{+}95, BGGR^{+}96, BFP^{+}96, CLT^{+}95, CRM^{+}96, DSvdV^{+}96, FHV^{+}96, FGA^{+}96, GCGE^{+}96, KAT^{+}95, KMOO95, LMMO96, LDCB95, MWT^{+}96, MS96, REdC95, RBL^{+}96, RLC^{+}96, TLF^{+}95, TEL^{+}95, ASES^{+}97, BNJ99, CGSY^{+}99, CMD^{+}98, CFB^{+}98, DPT^{+}97, DSM^{+}98b, EFML99, FH97, GJB^{+}98, GMS97, HSLB99, HFE^{+}98, HCLG99, KKT^{+}98, KMS^{+}98, KMK^{+}98, KPZ^{+}97, Lan97, LNH^{+}98, LYD^{+}97, LWKK97, MSL^{+}98a, MCS98, NKK^{+}99, PMCS97, RFCR^{+}97, RLMC99, SZBB98, SATW98, TKH^{+}97, TDR^{+}98, TTT^{+}99, WCV^{+}97, WGP98, WHB^{+}99, YHO^{+}98, ZhTA^{+}97]. Microautophagy [SKR^{+}98]. Microdomain [EKF^{+}97]. microdomains [DvD95, AvdG99, LLVS98]. microenvironment [BCC^{+}95, AEWG^{+}97]. microfibrils [ZHR95, KRS^{+}98b, WPS^{+}98]. microfilament [RROA95]. Microfilaments [BFHB97, CLF^{+}97b]. Microinjection [ASACF98, CG95, GCM98, LN96]. Micromanipulation [HMCL97]. micron [PHB96]. Micronucleation [SIUW98]. Micronuclei [HRR^{+}99]. Microphthalmia [BAH^{+}98, BBOB96, WGP98]. Micropipette [HMCL97]. microscopic [FTH96, KGGK99]. microscopy [BHDW95, GJ95a, MML96, OGS^{+}96, SOH^{+}95, JM97b, JOPM98, KKM^{+}97, SSM98]. Microsomal [AAB^{+}97]. microsomes [RAM^{+}96]. Microtubular [KRK^{+}98, FK95]. Microtubule [AARS^{+}97, CLF^{+}97b, KKS99, LSZ^{+}99a, MCM^{+}99, RSP^{+}95, SNK^{+}97, SEH97b, SNK^{+}99, WH97, YSB96, ZKSB96, BIR96, CFK95, CSY^{+}95, DFGL96, HFSHK96, IRMF^{+}99, IFHR^{+}96, KPKWW95, LJ95, LSCM^{+}95, LNY^{+}95, MJMS96, NAV96, NSYK^{+}95, OHB^{+}95, PBT^{+}95, RSR^{+}96, SJ95, SL96, SS95b, THK95, WSGPS95, WW96b, YNOH95, ZKB95, BHHF97, BG98, CYH98, CH97, DMMS98, DPG^{+}99, FFB^{+}97, GWB^{+}99, GLMG98, GBFL^{+}98, HPKL97, HTH^{+}97, LM98, LSH^{+}99, MTLGC99, MCTMK98, MR98, MMR99a, MS99, NMH^{+}98, PDS^{+}97a, PYM^{+}98, QGE^{+}99, RPV98, SFA^{+}98, SdCAH^{+}97, SYF^{+}97, SMB^{+}99, SYM^{+}97, TKH^{+}97, TWP98, TWS97, TZL^{+}98, WSS97, WCMS98, WBA^{+}98, YAC97, ZLL^{+}98]. Microtubule-associated [CLF^{+}97b, HFSHK96, OHB^{+}95, PBT^{+}95, FFB^{+}97, MS99, SdCAH^{+}97, SYF^{+}97, TKH^{+}97, TWP98, WBA^{+}98]. microtubule-based [NAV96, NSYK^{+}95, RPV98, SYM^{+}97]. Microtubule-Binding [WH97, IRMF^{+}99, KPKWW95, LM98]. Microtubule-dependent [SNK^{+}99, CSY^{+}95, YAC97]. microtubule-independent [CSY^{+}95]. Microtubule-mediated [SEH97b, LSCM^{+}95]. microtubule-organizing [MJMS96]. microtubule/kinesin [RSR^{+}96]. microtubule/kinesin-mediated [RSR^{+}96]. Microtubules [CS97a, SGD98, VLA^{+}98, DM95, GSC95, GG95b, HCAW95, MK95a, MH95a, MH95b, MGRS96, OHB^{+}95, PAV^{+}96, RK95, SLEB95, SYB95, SM95b, SNCBM95, SGGS96, SVB96, TK95, WPZ^{+}96, ZKSB96, AEVB98, AYMB99, BSB^{+}97, CNG98, CLF^{+}97b, FFB^{+}97, GCML98, HPKL97, HCAC99, HST^{+}98, KRS98a, KR99, LB98, MSH^{+}99, OWM^{+}99, PCD^{+}98, SHLD97, SN98, SMB^{+}99, TOB^{+}99, WSS97, WWD^{+}97, YAC97, ZP98]. microvascular [KvEdV^{+}96]. Microvesicles [SHH97]. Microvessels [ASP99]. Microvillar [KTD^{+}97]. Microvilli [BZL^{+}98, YTT99, BGB95b, PWG^{+}95, APF^{+}97, BFRB99]. Mid [EAM^{+}97, EFB^{+}98, EFML99, FdEP99, YYSB97]. Mid-anaphase [EAM^{+}97, YYSB97]. Mid-sized [EFB^{+}98]. Mid1p [BSW^{+}98]. Midbody [SGD98, WRW^{+}95, KLH98]. Middle [SPK^{+}97, LNRR98]. Midkine [MMMT95]. Midzone [WW96b]. Migrating [MGD97, WSS97]. Migration [CHP^{+}98, MN98, NCW^{+}99, SvDtK^{+}98, SWVR98, FPM^{+}95, FBD^{+}95, GBW^{+}95, HGH96, NAP^{+}96, PCL^{+}96, WPS^{+}96, ALCC^{+}99, BRLM98, BPS99, CAR^{+}97, CGC^{+}97b, DEPR97, FH97, GCAHW97, GBvdN^{+}98, HCW^{+}98, HLK^{+}98, KRCP99, KDO^{+}99, KLM^{+}98, KIT^{+}97, LPM^{+}98, LGZ^{+}99, LCB^{+}99, MZL^{+}99, NT97, OSM98, OFM^{+}98, PDS^{+}97b, RM97, RTM99, RWS^{+}99, SYM^{+}97, TWSD98, TGW99, TCYS99]. Migratory [CLK99, BEG^{+}95, GHP^{+}96, RD96a, CRYC99]. mild [CRM^{+}96]. mimic [PKR95]. mimicking [SOHP95]. Mimicks [PHN^{+}98]. mineralization [SAPLHD95, KNSP97, YSEI^{+}99]. Mineralization-competent [KNSP97]. mini [CDGO99]. Minichromosome [SO97, SO98]. Minifilament [EGM^{+}99]. minifly [GPSF99]. miniparamyosin [MAG^{+}96]. Minority [BMPP99]. minus [IRMF^{+}99, KJE^{+}99, KKE^{+}95, LJ95, MGRS96, SNK^{+}99, TWS97, WFS97]. minus-end [MGRS96]. MIP [CBM^{+}98]. MIP-3 [CBM^{+}98]. Misfolded [RvSL^{+}95, HDK96, GLL^{+}99, JBRA98, JWK98]. mislocalization [CF95]. Missegregation [BBL^{+}99a]. Mistargeting [VKIH98, Dav95, LOD96, MLO^{+}95]. Mitochondria [AAB^{+}97, BRD99, LCD^{+}98, AGB^{+}96, GLB^{+}95, MH95b, SNL96, BVKP98, EGS^{+}98, EAOS^{+}98, GGL^{+}98, sHWM^{+}99, KMK^{+}98, MDE^{+}99, RN99, SSRvdB98, WHS^{+}97, hYGYF97]. Mitochondrial [AAB^{+}97, BHG^{+}97, HTM^{+}98, HFL97, MAC^{+}97, PDS^{+}97a, TOSF^{+}98, BAC^{+}96, DC96a, MGS96, PLZ^{+}95, SFD^{+}96, SSP95, BSY97, BVKP98, DOSN^{+}99, DLC98, FY97, FY99, HKS97, KHS^{+}97, KEP^{+}99, LJK97, MNR^{+}98, MTW^{+}99, MVM^{+}99, OPB98, OKB^{+}98, PÖS^{+}98, SSS99a, SJ99, SY99, TKBM97, VL99]. mitochondrion [LOD96, MLO^{+}95]. Mitogen [HTK^{+}99, LYR96, CMS98, EKSC98, GTY98, KCG^{+}97, WZF97, ZTH^{+}97]. Mitogen-activated [HTK^{+}99, LYR96, CMS98, EKSC98, GTY98, KCG^{+}97, WZF97, ZTH^{+}97]. Mitogenesis [SMS^{+}98]. Mitosis [SVH^{+}98, TOT^{+}96, YTY^{+}99, DRAT95, EPVV96, LWM^{+}95, RACHV96, STN^{+}96, TSS^{+}95b, WS95b, WWP^{+}96, WM95, ALVMM98, BHJ98, BG97, CAS^{+}98, CLT99, ESM^{+}97, FdEP99, GCM98, KKH^{+}99, KR99, MOY^{+}98, NMT97, PDS^{+}97a, RS99, RC98, RWS^{+}99, SHP^{+}97, TYY^{+}99, qXLCH97, YGG97a, ZCE^{+}98, dS98a]. mitosis-entry [WWP^{+}96]. Mitosis-specific [TOT^{+}96, YTY^{+}99, CAS^{+}98, TYY^{+}99, qXLCH97]. Mitotic [CJS^{+}99, CH97, GDC97, KWD^{+}98, MHWW98, SVH^{+}98, SW99, VWO99, VRD96, BBZ^{+}95, CG95, GSC95, GSB^{+}96, GOT96, GMFN95, KM95, KJBM^{+}96, LN96, MW95b, NWG95, RMWW95, SSP95, SWC^{+}96, SGGS96, WW96a, WMO^{+}95, WM95, AN99, APH^{+}97, AK97a, AK97b, BHJ98, CS97a, CGMH99, CLL^{+}98, EAM^{+}97, FFLP99, GCD97, HSB98, HCG^{+}98, HD99, HKKH99, JHM98, LKE^{+}99, MHH^{+}99, MCTMK98, MBB^{+}97, MPT^{+}98, MMR99b, MSH^{+}99, NHA^{+}97, PTVD99, SdCAH^{+}97, SAR^{+}99, SMB^{+}99, SCPPW98, Ste97, SO97, WVM97, YSK^{+}97, YMD99, ZKS^{+}99]. Mitotically [HFE^{+}98]. Mitsugumin29 [NKK^{+}99]. Mixed [TMPM99]. MK [MMMT95].

MKBP [SSH^{+}98]. MKK [SVH^{+}98]. MKK/ERK [SVH^{+}98]. MKK6 [TMN97]. MKK6-p38 [TMN97]. MKK7 [TMN97]. MKK7-JNK [TMN97]. MKK7-JNK/SAPK [TMN97]. MKLP1 [KRK^{+}98]. mlc [SCEB99]. mlc-4 [SCEB99]. Mlc1p [SD98]. Mlh1 [WHB^{+}99]. Mmm1p [BVKP98]. MMP [BPS99, SB99c, LGZ^{+}99, MBCS98]. Mmp-2 [SB99c, MBCS98]. MMP-9 [LGZ^{+}99]. MMP2 [LTMR97]. MMP9 [LTMR97]. Mn [LRG^{+}98]. Mnt1p [LSKB95]. mobile [EKK^{+}96, EEHW^{+}98]. mobility [KKAN^{+}96, FSC^{+}97, GRTB97]. mobilization [RROA95, RTM99]. Mode [AMY^{+}98]. Model [GXE^{+}99, MAM99, GJ95a, RAE96, SGS95, FH97, FZZN98, GSL98, LWS99, MNM^{+}98]. models [KAP95, BHSAJ98, SRCC97]. modes [KKD^{+}95, YFST^{+}96]. Modification [SJW97, MCB96, ROG^{+}96, SMB^{+}95, TSS^{+}95a, MGM98, MWB98, SBS^{+}98]. modified [FSHD96, SLS^{+}96, ISN^{+}99]. Modifier [PAT99]. modifies [TEL^{+}95]. modify [BEG^{+}95]. Modular [MCL^{+}96, BTTB96, MMS^{+}96b, CMS^{+}97]. Modulate [DCVF98, RSC96]. Modulated [GTP^{+}99, KEJ97, LTR^{+}97]. modulates [GLD^{+}96, JSR96, LLZ^{+}96, MWL^{+}96, MCB96, NAV96, OSPJ^{+}96, SPMB^{+}96, TIJ^{+}95, WXS96, ZMM^{+}95, HCW^{+}98, LMW99, LBW^{+}97, LTW^{+}97, MPLS^{+}98, qSmLY97]. Modulating [BF97b, TGW99]. Modulation [HGH96, KSE^{+}97, iMOeM^{+}97, MBDBB98, SZFM98, WMP^{+}98, BKP^{+}95, BGR^{+}95, HMB95a, MAFJ^{+}95, BRP^{+}97, JLC98, MKS^{+}97, SOPM^{+}97]. Modulator [CBS^{+}97]. module [HSML96]. Modules [GCST97, TCSG96]. Moesin [KTD^{+}97, MEFH97, MMD^{+}98, SALdP^{+}97, YHD^{+}98, YTT99, HSK^{+}96, MF96b, FKO^{+}98, PKQ^{+}98, RBB97]. Mok1 [KHA^{+}99]. Molecular [BMS95, BSB^{+}97, BSW^{+}99, BD99, mCBLK98, CLL95, DLB96, DAB99, EPVV96, FHAP98, FTB97, HAS95, HIW^{+}98, HBJ95, HGG96, KIH^{+}97, KLMR95, KTD^{+}97, KGMF95, LSZ^{+}99b, MGM98, MAMJ95, MS96, RdPRW98, RHA^{+}97, SBS^{+}99, SPT^{+}95, TRG^{+}99, YTO^{+}99, BKP^{+}95, DCW^{+}96, FAO^{+}96, HD96, KA95, LCC^{+}95, LG95, MTC^{+}95, SNL96, SFNRH95, Val96, BCAS98, DSC^{+}99, IYM98, KLM^{+}98, LRC^{+}99, LHS97a, LGM^{+}97, PSH^{+}98, SWM^{+}98, TGL^{+}98, VZNR98, ZKN99]. Molecule [FSDA97, LCN^{+}97, LSZ^{+}99b, MN98, TNM^{+}99, BGK^{+}95, BHP^{+}95, DSC^{+}95, FWHW96, HBJ95, KLMR95, LXK^{+}95, MBS^{+}96b, RAH^{+}95, SIAS96, VDA^{+}96, WPS^{+}96, ANS^{+}97, BR98, CSC^{+}98, DCR98, ETP^{+}99, FHSM^{+}97, GSP^{+}98, HAB^{+}99, IKR^{+}98, JLID98, KEJ97, KSG^{+}98, LBW^{+}97, LVN^{+}98, MPLS^{+}98, SZBB98, SALdP^{+}97, SWS97b, SEBF98, TCYS99]. Molecule-1 [FSDA97, JLID98, LVN^{+}98, TCYS99]. molecule-2 [LXK^{+}95]. Molecule-3 [SALdP^{+}97]. Molecules [KNH^{+}97, LCN^{+}97, SKC97, YMAC^{+}98, BOLS95, BKB^{+}96, DLB96, FTH96, KZKS96, MAFJ^{+}95, MTC^{+}95, RvSL^{+}95, SPB^{+}95, WLR96, WFBO^{+}96, BRAM97, DBLD^{+}99, FKWP97, GRTB97, HHM^{+}98, LMB98, MGD97, RRA97]. Moloney [SHG96]. Monitors [CJS^{+}99, LR95, MSL98b]. Monoamine [LE97]. Monoclonal [ZKH^{+}95, NKI^{+}95, HTY^{+}99, SSS^{+}99b]. Monocyte [GCM^{+}97, WSWR99, KSL^{+}95, MPLS^{+}98]. monocyte-derived [KSL^{+}95]. monocytes [SCM^{+}96, WAMS96, HAB^{+}99, WMR^{+}98]. Monocytic [MCS^{+}97a]. monocytogenes [LLH^{+}99, MMP^{+}95, MGP97, RBM^{+}99, RAA^{+}97, SWS97a, STP96]. monohexosyl [BvdBvM96]. monomer [SFSV95, GDL98]. Monomer-sequestering [GDL98]. Monomeric [HST^{+}98]. Monomers [BCAS98, KHHS97]. monoorientation [RCKS95]. Monophosphate [DVL^{+}98]. Monophosphate-dependent [DVL^{+}98]. Morphogen [SW98]. Morphogenesis [JCC^{+}98, ZT99, BGB95b, CKP96, KKD^{+}95, LR95, LLRTP95, MMGT96, MMMT95, VMN95, VSB95, YSM^{+}95, ZKSB96, BAC97, BFRB99, CRA^{+}98, CGC^{+}97b, DCR98, KHA^{+}99, LTMR97, MBDBB98, MSL98b, MBCS98, NBS^{+}98, TTAH^{+}99, VMH97, ZBG99]. morphogenetic [BSK96, CCR96, CJH^{+}98, CR99, EIIM^{+}98]. morphogenic [BFS^{+}95, SWB^{+}96, EM98]. Morphological [FSM^{+}99, MHM99, OWW^{+}99, VSGF96, PSH^{+}98, SKR^{+}98]. morphology [FPM^{+}95, HCM96a, HMB95b, MHLB96, CBB^{+}99, DBH^{+}97, GKP^{+}97, OKB^{+}98, PKQ^{+}98, PH98, SCEB99, vLKvdK^{+}97]. Morphometry [PAN97]. Morphoregulator [HLG^{+}98]. Morphoregulatory [CCB^{+}98]. mortality [TALM96]. Mosaic [MB99]. most [MR95, GJW^{+}97]. mother [FK95, SH96]. mother/BUD [SH96]. Motif [CWSL97, DBE98, LKSW98, BPT96, FBD^{+}95, GLS96, JGD^{+}96, LHDC95, MLO^{+}95, NH95, NNMW96, PMCT96, RSRK96, SSSKR95, DKN^{+}97, KRGA98, PFKvD99, TBP^{+}99, FPF^{+}99]. Motif-dependent [LKSW98]. Motifs [WC97b, BBOB96, LDF95b, AKLME98, sHWM^{+}99, LMB98, XWK^{+}97]. Motile [YMPG98, CSM97, HLK^{+}98, MS97b]. Motilities [SNK^{+}99]. Motility [CLS^{+}97, ELL^{+}99, FIC^{+}97, LLH^{+}99, MSF^{+}99a, YMAC^{+}98, BSN95, CMUW96, KBR95, RJP96b, SSP95, STP96, SKWY95, SSW^{+}96, TIJ^{+}95, THK95, ZPS96, DAB99, DBJDT99, FOK^{+}99, GTP^{+}99, HST^{+}98, HYE^{+}98, KBG^{+}99, KcZS97, KCG^{+}97, KWS^{+}99, LZK^{+}97, MSH^{+}98, MKW^{+}99, NPJW99, NGT99, PKQ^{+}98, RM97, RSS98, RPV98, SBC99, SGGW99, TCM^{+}97]. motion [KCBR96, LNY^{+}95]. Motogenic [SWB^{+}96]. Motoneurons [RHP99]. Motor [HB98, KRK^{+}98, PdHTV99, PT97, BKEC96, DM95, GSB^{+}96, KWK^{+}96, KKE^{+}95, LSCM^{+}95, LNY^{+}95, NSYK^{+}95, Sch96, SSP95, WMC^{+}95b, YNOH95, BSD97, BNJ99, Bri99, CH97, EFML99, GDC97, HH98, KTAX98, PDS^{+}97a, RHW^{+}98, SHLD97, SYF^{+}97, TZL^{+}98, YAC97, YHO^{+}98, YSK^{+}97, ZCE^{+}98]. Motoring [Ste97]. Motorneurons [PURB98]. motors [MGRS96, Val96, Blo98, CGMH99, FTB97, SWO^{+}99]. Mouse [BMG^{+}99b, HPL97, KBS^{+}99, OZF^{+}99, AASH^{+}96, AHCC95, BGGR^{+}96, FMJ^{+}95, HTS^{+}95, HG95b, KKAN^{+}96, KJBM^{+}96, KSVZ95, LGdC95, LDCB95, LWB^{+}95, MST^{+}96, MPW95, MLLT95, MMMT95, PLM^{+}96, REdC95, SWS^{+}96, SSA^{+}96, TKM^{+}96, VMV^{+}95, VWK^{+}95, VKR^{+}96, WKL^{+}96, YRM^{+}96, BSS^{+}97, mCBLK98, FSA^{+}99, GCAHW97, HDDKH98, HCW^{+}98, JDK97, KRS^{+}98b, KMG^{+}97, LSL99, NMK^{+}97, PC98, RKRG99, RYB^{+}99, REC97, SLAS^{+}98, SSZ^{+}97, SJF^{+}97, SBGR98, TMPM99, WRG98, YPM97, ZLB^{+}97]. Move [BG98]. Movement [ISR99, MB99, WKWC96, WHS^{+}97, ASY96, BvdBvM96, HCW96, MMP^{+}95, PBT^{+}95, RFNF96, RW95, BSB^{+}97, DMMS98, GGPG99, NH99, PWW98, RTP^{+}98, RKR^{+}99, SNT^{+}98, YWMH99, YHC^{+}97]. movements [SWS^{+}96, Bri99, CH98, PYM^{+}98]. moving [KR96, STP96]. MPS1 [SGSW97, WW96a]. MRF4 [ACHG95]. mRNA [AAD^{+}97, ABR95, BBP^{+}98, BLEB96, BJB^{+}96, DW97, GL95, GDC95, HS96, IJB^{+}97, KDBW97, KBC99, NSR95, PRB97, PFKvD99, RZZ96, VIF^{+}96, WLD^{+}98, XJM^{+}95, XC97, YTO^{+}99, YCR^{+}98, ZSB99]. mRNA-binding [PRB97]. mRNAs [HPKL97, SMW^{+}99]. mRNP [SAEV^{+}98]. MRP [JCWP95, MKB^{+}95]. Ms [SBS^{+}99]. Ms/Ms [SBS^{+}99]. MT1 [BPS99]. MT1-MMP [BPS99]. mtHsp70 [MVM^{+}99]. Mtoc [GPKH99]. mu [GZK^{+}95, PPBJ96, ZMG96b]. MUC1 [WvdVV^{+}95]. mucin [DLB96]. Mucus [HHRdB^{+}98]. MukB [MCBE98]. Multicellular [MHM99, KA98]. multicomponent [CPL^{+}95]. multidomain [EGH^{+}95]. multidrug [DSvdV^{+}96]. Multifunctional [RPBB98]. multigene [SLP^{+}96]. multimerization [AM95a]. Multinucleation [SSV98]. Multiple [DG96, FMD^{+}98, GLS96, RdPRW98, SN97, TKY^{+}97, WS97, ZB98, CS96b, JWH96, WVVD95, BKMN^{+}98, FHSM^{+}97, IMK99, JMM^{+}98, KSK^{+}99b, LSH^{+}99, LJK97, PFDL97, PFKvD99, YDFB98]. Multipotent [MPR^{+}97]. multipotential [WPMW95]. Multiprotein [MPSC98]. Multispecificity [DRE97]. Multistep [FCB97, LHWH97]. Multistranded [YPB^{+}98]. multisubunit [TN95]. Multivalent [CLT99, WE98]. Multivesicular [FPHH96, KP96, GAPG97]. multiview [SOH^{+}95]. murine [BGK^{+}95, GCM^{+}95, OST^{+}95, SLEB95, SHG96, THT^{+}95, GM99, MNM^{+}98, PKBHK97]. Muscle [BRP^{+}97, BDO^{+}97, HTK^{+}99, HRD98, JCR97, MAB98, OBB99, WMP^{+}98, ACHG95, AHCC95, ASSS^{+}95, BTTB96, BGR^{+}95, BSMH95, CRG^{+}95, CGJ^{+}95, CJG^{+}95, DSM^{+}96, DLIB^{+}95, FWHW96, FDH96, HKS^{+}96, HKM^{+}96, HW96b, ISFB^{+}96, JSR96, KAT^{+}95, KWK^{+}96, KSM96, LMMO96, MAG^{+}96, MCA96, MWT^{+}96, MGA^{+}96, NMJL96, PWE95, PPF96, RJP96b, SLT^{+}96, SBSG96, SSZ^{+}97, SBS^{+}96, SSB96, SPT^{+}95, VKR^{+}96, WHM^{+}96, XGR95, XHUC96, YRM^{+}96, vdLST^{+}96, ASES^{+}97, BAC97, CLRR99, CJY^{+}97, DBC^{+}99, DBG^{+}97, DDM98, DDWM97, FCTS99, HLJ^{+}98, HSSW99, HHB^{+}97, JLID98, KCV^{+}98, KMK^{+}98, LMA^{+}97, LMW99, LMP98, LRC^{+}99, MMM^{+}98, MLC99, NKK^{+}99, PWZS97, PSPG^{+}98, PvDA^{+}98, PLB97, PFAF97, PFAA98, RLVM98, RNK97, RV98, SGGS^{+}99, SRR^{+}97, SCB^{+}97, SV98, TH99, TDR^{+}98, WTM^{+}99, ZBW^{+}97]. Muscle-derived [MMM^{+}98]. muscle-specific [BSMH95, SSB96, XGR95]. muscles [BZB^{+}96, MBB96, HCLG99, KAFB99, RV98]. Muscular [CAC^{+}99, DSM^{+}98b, HMC^{+}99, SRCC97, SEAB^{+}99]. Musk [ZGYS99]. must [PNM95]. Mutant [BPA^{+}97, PFHWN98, YF99, AYF96, AHM^{+}96, BWF^{+}96, CF95, DBvdBS95, HP95, KMOO95, MKB^{+}95, NDBR95, NCS95, PHO^{+}95, RME^{+}95, SFD^{+}96, ABL98, BD97, BD98, GBW98, JN98, KMHK97, KRS^{+}98b, KD97b, KGHG98, LMB98, LRS^{+}98, jLC97, MAEE98, MFF97, MVM^{+}99, MS99, PMCS97, SGR97, WHB^{+}99]. mutants [FHCL95, HMSK95, JWH96, KSM96, NPRT95, PWK^{+}96, STN^{+}96, WLML95, WMXE96, BGG98, GE97, HHS^{+}99, MM97b, PBJ^{+}98, PSH^{+}98, RKRG99]. mutated [dHvPL^{+}95, GKS^{+}98]. Mutation [KMS^{+}98, MAB98, BTNZ^{+}95, HCA^{+}95, LCH96, LWB^{+}95, NH95, RWO95, RBL^{+}96, SSSKR95, JM97a, KMHK97, LBWF^{+}97, PURB98, VPM^{+}98, ZhTA^{+}97]. Mutational [BFL95, FY97, OAB^{+}98, PSW95, GSK^{+}99]. Mutations [BBOE98, BBL^{+}99a, BGH^{+}97, GBW^{+}95, KP96, LNRR98, MMN^{+}99, HCM96a, KKGD96, MGA^{+}96, ROG^{+}96, WGR^{+}95, GGK97, HTW97, KAFB99, YAiN^{+}98]. Mx [DCJ97]. mXRN1p [BSS^{+}97]. Myc [SPS97b]. Myelin [TSD^{+}97, BQF^{+}95, HTS^{+}95, WF96, AAD^{+}97, BGBS98, KBC99, MSF^{+}99b]. Myelin-associated [TSD^{+}97]. Myelination [EZC^{+}97]. Myeloma [MMAK97]. Myelomonocytic [FIC^{+}97]. Myf [KCV^{+}98]. Myf-5 [KCV^{+}98]. Myf5 [LMP98]. Myo [VB98]. Myo- [VB98]. myo2 [KSY97, GBN95]. Myo2p [PSB98, SS97, SHPB99, SD98]. Myo2p-dependent [PSB98]. MYO5 [GAW^{+}96]. Myo5p [ABE^{+}98]. myoA [MGM95]. Myoblast [AGS^{+}97, BMPP99, EGA97, GTL^{+}96, VWK^{+}95, DFMG97, EGA97, PGD^{+}97, QBvD^{+}98, SLW^{+}99]. myoblasts [HLSM95, KSS^{+}95, RPC^{+}95, YFST^{+}96, CMR^{+}97, LMP98, RV98]. myocardial [MWT^{+}96, ZTH^{+}97]. Myocyte [CMS98]. Myocytes [MAM99, GF95, LK96, CMS98, FRHS98]. MyoD [KCV^{+}98, LCH96, SGGS^{+}99]. Myofibers [WB99]. myofibril [KSS^{+}95]. myofibrillogenesis [SHC95]. Myofibrils [OBB99, LMA^{+}97]. Myofibroblastic [SBPR^{+}98]. Myofibroblasts [ML97]. myogenesis [AW96, EBR96, FWHW96, VLX^{+}96, VMV^{+}95, DDM98, KJR^{+}98, RLVM98]. Myogenic [AGS^{+}97, BMPP99, Ord99, DBC^{+}99, KMM^{+}98, LMA^{+}97, LMW99, MNG^{+}98, SGGS^{+}99]. Myogenin [AW96, HCLG99, VMV^{+}95, ACHG95]. myomesin [OGS^{+}96]. Myosin [BBOE98, BGH^{+}97, BML^{+}97, Bri99, CM95b, GW95, JK97, KAT^{+}95, KBS^{+}99, MCRB97, NCW^{+}99, NT97, PT97, SLAS^{+}98, VSB95, XHUC96, BCMK95, BRHD95, DRR96, GAW^{+}96, GBN95, HCW96, HW96b, HHTT95, JWH96, KSG^{+}96, KWK^{+}96, KFE96, MGM95, MM95a, NPRT95, SRRB95, ASES^{+}97, ABE^{+}98, BWWT97, BKJL^{+}98, CDH^{+}99, FRHS98, FKO^{+}98, JM97b, JOPM98, KFO^{+}99, KSY97, KAFB99, LM98, LWS99, LL98b, LBO^{+}98, MOY^{+}98, MSEP98, NHA^{+}97, RKR^{+}99, SHPB99, SCEB99, SCB^{+}97, SD98, SVMB97, TYY^{+}99, WTM^{+}99, WBR^{+}98, OP96b, OP96c]. Myosin-binding [FKO^{+}98]. Myosin-driven [BML^{+}97]. Myosins [HGG^{+}97]. Myotactin [HSSW99]. myotonic [TMS^{+}95, SSH^{+}98]. myotube [VLX^{+}96, BAC97, TH99]. myotubes [CRG^{+}95, MPW95, RL96, mCBLK98, MLC99, TGGB97]. myristic [BAC^{+}96]. Myristoylated [SOJM99]. Myristoylation [vtHR97, BAC^{+}96]. Myristylation [JH96].

N [BRS99, ETP^{+}99, FVKS96, NPJW99, BEZ^{+}95, BLE^{+}96, BALL98, BAC^{+}96, BGBK99, CMR^{+}97, DBJDT99, ELL^{+}99, ICW^{+}96, MGB^{+}98, MGD97, NRD98, SGYL^{+}96, WPS^{+}96, LSL99]. N-acetylgalactosaminylphosphotransferase [BEZ^{+}95]. N-Cadherin [NPJW99, BLE^{+}96, ICW^{+}96, MGD97, SGYL^{+}96]. N-CAM [ETP^{+}99, MGB^{+}98]. N-Glycans [BRS99]. N-myristoylation [BAC^{+}96]. N-Wasp [ELL^{+}99]. N135I [BBW^{+}98]. N141I [JM97a]. N1E [HIW^{+}98]. N1E-115 [HIW^{+}98]. N375 [HMD97]. Na/K [MSF^{+}95]. NAB2 [QWS^{+}98]. NAC [WSW95]. NAD [MCS^{+}97b]. NADH [BAC^{+}96]. NADH-cytochrome [BAC^{+}96]. NADPH [RRRA95]. Naegleria [HPKL97]. NAP [KKFN^{+}95]. NAP/SET [KKFN^{+}95]. NAP1 [KM95, AK97a, FMM^{+}98]. Nascent [HCW99, MZN97, AS96, NMHS95, WSW95, CSA^{+}97, EJ99, GBS99, LGGS97]. Native [BKEC96, VSGF96]. Natural [AGS^{+}97]. Naturally [BMP^{+}98]. Navigation [FCB97]. NC [BTNZ^{+}95, RKR^{+}97]. NC-1 [RKR^{+}97]. NC-2 [BTNZ^{+}95]. NCAM [BSMH95, RL96]. ncl [FR98]. ncl-1 [FR98]. ND10 [IM96, ISN^{+}99]. Ndc1p [CRGW98]. NDP52 [KGMF95]. near [ETF95]. Necessary [KRK^{+}98, BLLB95, LGRB96, OFY^{+}95, SKS^{+}95, AL97, DOP^{+}98, HKS97, PMBM^{+}99, SYT97, TCV^{+}98]. neck [SH96, OMS98, RPV98]. Necrosis [DCVF98, DH96, MBD^{+}96, MGB^{+}97, VHF97]. Nectin [TNM^{+}99]. Nectin/PRR [TNM^{+}99]. Nef [FMP^{+}97]. Nef-mediated [FMP^{+}97]. negative [LCH96, MHLB96, TYSP^{+}96, WXS96, WF96, vABSP95, BRD99, BWWT97, DDWM97, PBM^{+}97, RDP^{+}99, ABL98, LMB98, vHGB^{+}97]. Negatively [HYN97, OK98]. Neither [KDvFP96]. Nek2 [FMM^{+}98]. NEM [MF96a]. Nematode [ISR99, CSDR96]. Neocentromere [YHC^{+}97]. Neocentromere-mediated [YHC^{+}97]. Neonatal [CMS98]. Neoplastic [MFBK99]. Neosynthesis [LBBP98]. Neovascularization [CAR^{+}97]. NEP [DFL^{+}99]. Nerve [CEZA97, DSVL^{+}98, CJG^{+}95, DPC^{+}96, FSIG^{+}95, WD95, DBG^{+}97, DPM^{+}97, GBW98, LMS97, NBJ^{+}99, SCB^{+}97, ULMT99, ZCOP99]. nerve-muscle [CJG^{+}95]. nerves [SRR96a]. Nervous [PJE^{+}99, SZBB98, AC95b, DSC^{+}95, EKPK95, THT^{+}95, WS96, ZL95a, BPS99, TKH^{+}97]. nestin [FJT^{+}95]. Network [MM99b, AS96, DLGB95, DWS^{+}95, GCGE^{+}96, IRMF^{+}99, LYC^{+}95, NNMW96, BHG^{+}97, CSA^{+}97, CWY99, DDF^{+}98, MM97a, Sal99, TH97]. network-associated [IRMF^{+}99]. networks [All95, MW95b, WSGPS95, PYM^{+}98, YMPG98]. NeuAc [HHRdB^{+}98]. Neurabin [NOS^{+}97]. Neural [ETP^{+}99, BLL95, FWHW96, KZKS96, MBS^{+}96b, OSK^{+}96, SSM^{+}95, SPB^{+}95, WPS^{+}96, AGP^{+}98, BR98, FHSM^{+}97, HCW^{+}98, KEJ97, MGD97, NOS^{+}97, SVB^{+}97]. Neural-specific [AGP^{+}98]. neuregulin [YSM^{+}95, SV98, VGVF97]. Neuregulin-like [SV98]. Neuregulins [MMM^{+}98]. Neurexin [EZC^{+}97]. Neurexins [LBB97]. Neurite [BGBK99, KSG^{+}98, ZLPL99, AHM^{+}95, AC95b, BKB^{+}96, DC96b, GSC^{+}96, IKT^{+}96, KZKS96, LW96, MLW^{+}95, NHB^{+}95, AGP^{+}98, HIW^{+}98, HJJJ97, MSS^{+}99, MQR^{+}97, NOS^{+}97, SLN^{+}97, SVB^{+}97, TSD^{+}97, VZNR98].

Neurite-like [ZLPL99]. neurites [LZC^{+}95, VLZR96, ZP98]. Neuritic [HGI^{+}99]. Neuritogenesis [AGP^{+}98, BKK^{+}97]. Neuroblastoma [ZLPL99, GH95, KSVZ95, HIW^{+}98, YSK^{+}97]. neuroectoderm [SSL^{+}95]. Neuroendocrine [SFR^{+}98, ZL95c, BHC99, SHH97]. Neurofascin [VLZR96, DLB96, GRTB97, VZNR98]. neurofibromatosis [MF96b]. Neurofilament [MWL^{+}96, RHW^{+}98, NCZ^{+}95, WMC^{+}95b, EFB^{+}98, EFK^{+}98, EFML99, HT98, KB99b, KTAX98, RHW^{+}98, ZLL^{+}98]. Neurofilament-dependent [RHW^{+}98]. neurofilamentous [WMC^{+}95b]. Neurofilaments [ZhTA^{+}97, XML^{+}96, RHW^{+}98]. neurogenic [HWPC96]. Neuroglian [DMD^{+}96, HHM^{+}98]. Neuroglian-mediated [DMD^{+}96]. Neurolin [LDP^{+}99]. neuromodulation [LYR96, LYLR98]. neuromuscular [GYS^{+}95, ISFB^{+}96, KBSW95, TCM96, CAS97, DPT^{+}97, GMS97, PMCS97, PXRR99, RRA97, WS98]. Neuron [ABP^{+}98, APG^{+}96, BHP^{+}95, Kun95, SPB^{+}95, AYMB99, BMP^{+}98]. neuron-glia [BHP^{+}95, SPB^{+}95]. neuron-specific [Kun95]. Neuronal [HYK^{+}98, MN98, BKB^{+}96, KSVZ95, MG96, SSW^{+}96, TLF^{+}95, BMP^{+}98, CAS97, CEZA97, FH97, FFB^{+}97, FJ98, GKP^{+}97, GBFL^{+}98, KLO^{+}98, SWD^{+}98, SEBW^{+}98, VMT^{+}99, YHO^{+}98, ZDK^{+}99, vLKvdK^{+}97]. Neurons [AAM^{+}98, ALC^{+}99, CDH^{+}98, BSJJ95, CLS96, CBB^{+}96, DVD^{+}96, GSC^{+}96, HFSHK96, Kun95, LW96, MFM^{+}95, MJ96, MH95b, NA96, PBB^{+}96, SBR^{+}96, TRJ^{+}95, WMC^{+}95b, WD95, YSK95, dCVCV95, dHvPL^{+}95, AGP^{+}98, BNJ99, BKK^{+}97, BHBB99, EY99, KB99b, KTAX98, KSY^{+}98, LM98, LSL99, LHR^{+}97, LYLR98, MDE^{+}99, MLB98, MBBT98, NRP98, PKQ^{+}98, PMP^{+}98, PMKR97, SC97, SWO^{+}99, TSD^{+}97, WNB97]. Neuropilin [MSF^{+}99a]. Neuropilin-1 [MSF^{+}99a]. neuropoietic [SSM^{+}95]. Neuroprotective [FEF^{+}97]. neurotoxin [IKT^{+}96, NBJ^{+}99]. Neurotractin [MSS^{+}99]. Neurotransmitter [ZCOP99]. neurotrophic [KBSW95, TLF^{+}95, TRJ^{+}95, TWSD98]. Neurotrophin [ABP^{+}98, YLB^{+}97, ZSB99, RDM^{+}96, BMP^{+}98, BKK^{+}97, SCB^{+}97]. Neurotrophin- [SCB^{+}97]. Neurotrophin-3 [BKK^{+}97]. Neurotrophins [NA96, VMT^{+}99]. neutrophil [LFB96, NYN^{+}95, PCL^{+}96, LCB^{+}99, MDV98, TCYS99, ZJY^{+}98]. neutrophil-derived [NYN^{+}95]. neutrophils [DDWG96, FWG^{+}96, MHK^{+}95, MBD^{+}96, MPB^{+}95, OCS^{+}95, PNM95, SCM^{+}96, ZLB^{+}97]. Newly [vtHR97, BDE^{+}95, CLS96, SHY^{+}96, vdBLCvM96, BM99, HVL^{+}99, LBO^{+}98, TH97]. Newt [TGGB97]. Nexilin [ONI^{+}98]. NF [EFB^{+}98, RHW^{+}98, FWBSO97, GBBSO96, HDV^{+}99, KTAX98, KRMMA96, LLN^{+}95, LDH^{+}98, MWL^{+}96, NCZ^{+}95, PSBB96, RK95, SAC^{+}98, SS95a, WMC^{+}95b, XZSC98, ZLL^{+}98, SRA^{+}99]. NF- [FWBSO97, HDV^{+}99, LDH^{+}98, SAC^{+}98, XZSC98, HDV^{+}99]. NF-kappa [KRMMA96, LLN^{+}95, RK95]. NF-kappaB [GBBSO96, PSBB96]. NFAT [KDG98]. NFH [TEL^{+}95]. NFL [TEL^{+}95]. NFM [TEL^{+}95]. Ng [BHP^{+}95, KSG^{+}98]. NgCAM [BKB^{+}96, DC96b, KZKS96]. NGF [BKK^{+}97, DVD^{+}96, EKPK95, MFM^{+}95, MDE^{+}99, QWS^{+}98].

NGF-deprived [DVD^{+}96, MDE^{+}99]. NGF-induced [QWS^{+}98]. NGF-stimulated [EKPK95]. NH [KD97a, MSH^{+}97, SKC97, AAB^{+}97, BPA^{+}97, FKNM97, GTC^{+}98, vtHR97, HPL97, sHWM^{+}99, MBA97, MSH^{+}97, OWD^{+}99, SAR^{+}99, SWK^{+}99, TNFK^{+}98, WRG98]. NH2 [CGJ^{+}95, GBG^{+}96, MAS^{+}95, PXR^{+}95, ZL96]. NH2-terminal [CGJ^{+}95, MAS^{+}95, PXR^{+}95, ZL96]. NH2Cl [NYN^{+}95]. NHE [BJR^{+}99]. Nic96p [ZDT^{+}96]. nicotinic [ACHG95, MPW95, Wal95]. Nidogen [OTS98]. nidulans [BRLM98, FT95, MGM95, WOM98]. NIH3T3 [SL95]. NILE [MBS^{+}96b]. NIMA [WOM98]. Nitric [GM99, LHS97a]. NK1 [SCG^{+}96]. NK2 [SCG^{+}96]. NO [GM99, DSE^{+}95, WPZ^{+}96, GL95, GM99, FFH^{+}98]. NO-mediated [GM99]. Nocodazole [SWR^{+}98]. Nocodazole-induced [SWR^{+}98]. nod [ASH95]. nodal [DLB96]. node [DLB96]. Nodes [BDO^{+}97]. nomenclature [DEG^{+}96]. Non [DLIB^{+}95, KOA^{+}99, vIH98, HP95, LKD^{+}95, MMGT96, OCS^{+}95]. Non-Basement [KOA^{+}99]. non-claret [MMGT96]. non-congenital [LKD^{+}95]. non-endocytic [HP95]. Non-muscle [DLIB^{+}95]. non-oxidative [OCS^{+}95]. noncadherin [TET^{+}96]. noncaveolar [DvD95]. Nonclassical [SBOK97, SWS97b]. Noncontractile [HHB^{+}97]. Nondisjunction [LARH97]. Nondividing [wJTSSL98]. Nonendothelial [KKT^{+}98]. Nonerythroid [MHH^{+}99]. Nonessential [WCV^{+}97]. nonexcitable [MKXY95]. nonmuscle [BZB^{+}96, KSG^{+}96, MP98, SCEB99]. nonneuronal [RGS^{+}96]. nonoverlapping [PHO^{+}95]. nonpolarized [MXSRB96, WC97a]. Nonrandom [XJM^{+}95]. Nonreceptor [BALL98]. nonstriated [MGA^{+}96]. Nonsynaptic [MQR^{+}97]. nontetrameric [BK96]. Nontranscriptional [BR98]. nontransformed [TALM96, WXS96]. Nonuniform [KRK^{+}98, SYB95]. nonvectorial [GM95]. Nopp140 [IYM98]. Normal [BGH^{+}97, RBM^{+}99, TSC^{+}98, BRHD95, DRAT95, DKD^{+}96, KF95, PHB96, SBSG96, SKR96, ZMM^{+}95, dCVCV95, Bri99, DHDJ^{+}97, ETP^{+}99, HFE^{+}98, LMM^{+}99, LJK97, MBB^{+}97, PMBM^{+}99, RB98, TGN97, TCM^{+}97, ZLM^{+}98]. normally [SJA96]. NORs [RACHV96]. Notch [HKJ^{+}99, MLLT95, PAT99]. Notochord [ACH^{+}98]. Novel [AAM^{+}98, AAB^{+}97, BdPE^{+}97, CWSL97, CGC^{+}97a, CYSD99, CGMH99, DSVL^{+}98, EGKC^{+}99, FSIG^{+}95, FKNM97, HDV^{+}99, HDDKH98, KNH^{+}97, KOA^{+}99, KND^{+}97, LCN^{+}97, jLC97, MWF^{+}99, MZN97, NCV^{+}98, PPWL97, PdHTV99, SKC97, TWP98, ZC99, ARM^{+}95, BSJJ95, BPT96, CBB^{+}96, CLR^{+}96, CF95, CMWL96, CCR96, FLPS95, GAW^{+}96, GDC95, GGR^{+}96, GEW^{+}95, GLB^{+}95, HKS^{+}96, HSML96, KSK^{+}96, KGMF95, LVD96, MC95, MCB96, MBP^{+}95, OSH^{+}96, OABD^{+}96, OS96b, OCS^{+}95, PL96, RHS^{+}96, SLW^{+}96, SOE^{+}95, SJ95, SBW^{+}96, SMC^{+}96, SNCBM95, SJSJ95, SRRB95, SOG96, TWVC95, WYY95, WFBC96, WMXE96, WPMW95, ZDT^{+}96, ZL95a, vdLST^{+}96, AMY^{+}98, APRB98, BBR97, BATP99, BH98, BF97b, CC98, CS98, CSM97, DSP97, DZS^{+}98, EGM^{+}99, FBM99, FMM^{+}98, FFH^{+}98, GKP^{+}97, GCML98, GCD97, GDB^{+}97, GTC^{+}98, GSP^{+}98, HGW^{+}98, HCG^{+}98, HYE^{+}98, HSSW99].

Novel [HJJJ97, HC97, HRA^{+}98, HHB^{+}97, HHS^{+}99, HKKH99, KSK^{+}99a, KLO^{+}98, KSK^{+}99b, LNM^{+}98, LMW99, MSL^{+}98a, MNS^{+}97, MSS^{+}99, MPLS^{+}98, MPR^{+}97, MTW^{+}99, MR98, MSWL^{+}99, MPL^{+}97, NMH^{+}98, NOS^{+}97, ONI^{+}98, PDO^{+}99, PYL98, RD97, RHA^{+}97, RKB97, RHPW97, SHC^{+}97, SdCAH^{+}97, SSW^{+}98, SHH97, SMZ^{+}98, SDL98, SGJP99, SWR^{+}98, SV98, SSH^{+}98, TGW99, THS^{+}98, TBP^{+}99, VCL^{+}98, WGF^{+}98, WSH98, WJM^{+}99, YPDM98, ZDK^{+}99]. novo [GGR^{+}96, ZMM^{+}95, BMS^{+}97]. NPC [ABR95]. NPFXD [THP96]. NPXY [CWSL97, FBD^{+}95]. Nr [SLN^{+}97, SPB^{+}95]. Nr-CAM [SPB^{+}95]. Nr-CAM/ [SPB^{+}95]. NrCAM [DLB96, VLZR96]. Nrk [MSFA99]. NRP [KLO^{+}98]. NRP/ [KLO^{+}98]. NSF [MW97, AMJM95, BMB97, MBF^{+}98, XMMW97]. Nsp1p [FHAP98, GEW^{+}95]. NT [APG^{+}96]. NT-3 [APG^{+}96]. NT2N [CGC^{+}97a]. NTF2 [PG95]. Nuclear [BdPE^{+}97, BSAJ97, ESM^{+}97, FB96, GBMS95, HCA^{+}95, JHM98, KWN^{+}99, LSB98, Mah96, MMF95, MYTK99, NB97, PRB99, PGC^{+}95, RMY97, RST^{+}99, RPBB98, SLF^{+}99, SPV^{+}99, SCR98, STRM95, SJW97, ARM^{+}95, BLEB96, BLG95, BS95a, BDvdZW95, BRJP96, BS95b, CRLM96, CSO^{+}95, CAA95, CAVA96, CMWL96, CKS^{+}95, ETF95, FKS95, FT95, GDC95, GEW^{+}95, HGG96, IWW95, IM96, KWCS96, KGMF95, KF95, LWM^{+}95, LvSC^{+}96, MF96a, MCB96, MDS95, MGY^{+}95, ND96, NRM^{+}96, NW96, NNMW96, PG95, PMMF95, RPW96, RRRA95, RLCM96, SRG^{+}96, SD95, SW95, dCBR95, SG96, TKVC95, TGG95, TDW95, TFF^{+}95, VIF^{+}96, WA95, ZDT^{+}96, ZKSB96, ZN95a, ACF^{+}97, APRB98, BBP^{+}98, BRLM98, BBR97, BD97, BATP99, BF97a, BW97, CRGW98, CRRF97, DEPR97, DFL^{+}99, EJ97, ESM^{+}97, FHAP98, FY97, FHWV97, GHW99, HSW^{+}98, HTWC98, HTY^{+}99, HCRS98]. Nuclear [HFP^{+}98b, HHS^{+}99, IW97, ISM97, ISN^{+}99, IJB^{+}97, JP98, JMWG98, KC99, KSV^{+}99, KBD99, KDG98, KDK^{+}99, KLO^{+}98, KIT^{+}97, LBBP98, LBWF^{+}97, LHR^{+}97, LHS97b, MSB99, MGM98, MAW98, MHH^{+}99, MWB98, PRB97, PPP^{+}98, PFKvD99, RWS^{+}99, RPB97, SSW^{+}98, SHMC99, STF98, SYM^{+}97, SIUW98, SSS^{+}98, SEK^{+}97, SMG^{+}97, SW98, SJZW97, dCBR99, SEAB^{+}99, SAEV^{+}98, SNK^{+}99, THS^{+}98, WSSB99, WOM98, YWMH99, YGG97a, YGG97b, YLD97, YCR^{+}98, dBCKS98]. Nuclear-coded [LHS97b]. nuclear-encoded [RRRA95, YCR^{+}98]. nuclear/chromosome [CMWL96]. nucleate [AFFS96, SJ95, OWM^{+}99]. Nucleation [ELL^{+}99, NMH^{+}98, RSS98]. nuclei [BJB^{+}96, FB96, PMMF95, TMS^{+}95, ZZYG96, ACF^{+}97, ABAAS98, FHWV97, wJTSSL98, LHR^{+}97]. Nucleobindin [LYH^{+}99, LLNH^{+}98]. Nucleocytoplasmic [BF97a, KDG98]. Nucleocytosolic [DSC^{+}99]. Nucleoid [MTW^{+}99]. Nucleoli [VLAHV98]. nucleolus [BFL95, JCWP95, IYM98, JWW^{+}99, OAB^{+}98, Ped98]. Nucleoplasm [SSS^{+}99b]. Nucleoporin [BdPE^{+}97, ABR95, BS95b, GDC95, HB95b, IWW95, NRM^{+}96, ZDT^{+}96, BD97, FBM99, GCA^{+}98, HHS^{+}99, IW97, PFDL97, STF98]. nucleoporins [ARM^{+}95, FBM99]. Nucleosomal [SVS97]. Nucleotide [WSCN97, JLRS95, BSRG99, WWM^{+}98, vLKvdK^{+}97]. Nucleotide-gated [WWM^{+}98]. Nucleus [FPRL97, TB99, FA95, FKS95, GG95a, LDF95b, PWK^{+}96, SJA96, TSS^{+}95b, YvdEH^{+}95, CBB^{+}99, GCKC^{+}97, HSB98, HDES97, KB99a, KLL^{+}97, NB97, SBV97, SEH97b, YNM^{+}99]. Null [NWM^{+}97, CMR^{+}97, HSLB99, HFE^{+}98, LWKK97, PC99, PBJ^{+}98]. Num1p [FK95]. NuMA [MHH^{+}99, GSC95]. number [APG^{+}96, CS99, HZC^{+}97, SHLD97]. Nup100p [IW97]. Nup116p [IWW95, IW97]. NUP120 [HCA^{+}95]. Nup120p [ABR95]. Nup153 [BLEB96, STF98]. Nup154 [GCA^{+}98]. Nup155 [GCA^{+}98]. Nup155p [ARM^{+}95]. Nup157p [ARM^{+}95]. NUP159 [GDC95]. Nup170p [ARM^{+}95, MAW98]. Nup188p [NRM^{+}96, ZDT^{+}96]. Nup214 [BdPE^{+}97]. Nup53p [MAW98]. Nup59p [MAW98]. NUP82 [HB95b]. Nup82p [GEW^{+}95]. Nup84 [BdPE^{+}97]. Nup96 [FBM99]. NUP98 [PFKvD99, FBM99, PFDL97]. Nurse [GCST97].

O [HHRdB^{+}98, CFS98, SGTH99, WPS^{+}96, WBG^{+}97, YLB^{+}97]. O-2A [WPS^{+}96]. O-glycosylated [YLB^{+}97]. O-Glycosylation [SGTH99]. observable [WPZ^{+}96]. Observations [HKBM99]. observed [JCWP95, ZN95b, GBFL^{+}98]. occludens [RHHRB96]. occludens-1 [RHHRB96]. Occludin [MSFA99, SFD^{+}98, AASH^{+}96, hCMPG97, FFH^{+}98, FSFT98, HGW^{+}98, MSFA99, SFS^{+}97, WG97]. Occludin-deficient [SFD^{+}98]. occupancy [MTGY96, SVD96, SSV98]. Occur [DMMS98]. Occurring [BMP^{+}98]. Occurs [BPWS98, GHS98, vIZKH97, MGY^{+}95, MCS^{+}97a, MLB98]. Och1p [HW96a]. OCI [GKS^{+}98]. OCI-5 [GKS^{+}98]. OCI-5/ [GKS^{+}98]. Oct1 [SS95a]. Ocular [PGP^{+}99]. ODA3 [KPW^{+}97]. Odz [OZF^{+}99]. Off [GLMG98]. Oh [EPMB^{+}99]. Oleamide [GCE^{+}97]. oleic [EB95, ESR97]. Oligoadenylate [DCJ97]. Oligodendrocyte [BR99, GR97, GSL98, VGVF97]. Oligodendrocytes [OWW^{+}99, SWD^{+}98, TNCM97, ULMT99]. oligodendrocytic [HTS^{+}95]. Oligomeric [GWLH98, LHM^{+}96, KSOK97, RC97]. Oligomerization [GJ95a, GBG^{+}96, BWG95, TSSR98]. Oligomerized [MLJM95]. oligomers [BGB95b]. oligonucleotides [OST^{+}95]. oligoproline [ZPS96]. Oligosaccharide [JBRA98]. Oligosaccharides [FHY99, SPBCM97]. oligosaccharyltransferase [KKG95, SCK^{+}95, SCKG95]. Oncogene [KGF^{+}97, RME^{+}95]. Oncogene- [KGF^{+}97]. Oncogenes [MWGE97, Sch97]. oncoprotein [LK96, GLMG98]. One [GCST97, GPKH99, GSK^{+}99, KCM^{+}97, MNS^{+}97]. only [KCBR96, KCM^{+}97, LTMR97]. Onset [KWD^{+}98, MC98, PTBC96, CG95, SWS^{+}96, TSS^{+}95b, BNJ99, GDM^{+}99, KR99, MHC^{+}97, STK^{+}98, SMM99]. onto [HER95, LWM^{+}95, SCG^{+}96, RRA97]. Oocyte [TMM^{+}95, GHJJ95, IMFS96, MBF^{+}97]. Oocytes [BMG^{+}99b, GRB^{+}97, BK96, FKS95, OSH^{+}96, SRH^{+}95, EK97, VSRP97]. Oogenesis [gLSE^{+}99, WBSNV97, MM96, JK97, NGT99]. Op18 [LSH^{+}99]. Op18/ [LSH^{+}99]. Opacity [CML^{+}98]. Open [RKB97]. operated [HFL97]. operator [RSL^{+}96]. operator/repressor [RSL^{+}96]. OPGL [BxQK^{+}99]. Opposes [FK99, MSH^{+}99]. Opposing [GSB^{+}96, vLKvdK^{+}97]. Optic [ULMT99]. Optical [SNT^{+}98]. Optimal [SCH^{+}98]. order [BOLS95]. Ordered [DFL^{+}99, SCPPW98]. Ordering [SHK^{+}99]. Orders [MMK^{+}97]. ORF3 [CSO^{+}95]. organelle [BSN95, NAV96, RSP^{+}95, WPLK95, YNOH95, BENV97, LHS97b, MBF^{+}98, RTP^{+}98, WFS97]. organelles [FC95, HTKH96, NSYK^{+}95, NSC96, VGM96, WSBE^{+}95]. organisms [SCKG95]. Organization [FPRL97, gLSE^{+}99, SB99d, YTT99, ZC99, AHCC95, BRB96, BJB^{+}96, CMHT96, CRCH95, EPVV96, GSC95, GSB^{+}96, HCM96a, LCC^{+}95, MCL^{+}96, MM95a, RSL^{+}96, SHA95, SS95b, WSG^{+}95, XJM^{+}95, YvdEH^{+}95, BGN^{+}97, mCBLK98, EGA97, HTH^{+}97, HKS97, HTWC98, HDES97, JP98, JE98, KSK^{+}99b, MTLGC99, MCTMK98, PAN97, PTVD99, RHW^{+}98, RSB^{+}99, SLF^{+}99, SRV^{+}97, SMG^{+}97, TWiK^{+}99, WPS^{+}98, WWO^{+}98, ZTH^{+}97, BGG98]. organize [SGGS96, WUUI^{+}98]. Organized [FWL^{+}98, DZS^{+}98, GDC97]. Organizer [RC97]. Organizers [RV98]. organizing [BIR96, MJMS96, WLF98]. Organogenesis [LTMR97]. organs [TSS^{+}95b]. organs-31 [TSS^{+}95b]. orient [Gol95, CS97a, Ste97]. orientation [SYB95, MR98, MMR99a, NA99, SYM^{+}97, TIBL99, WS97]. oriented [LZC^{+}95, VPH95]. Origin [HN98, HFP^{+}98b]. Originate [SHH97]. Originating [DBC^{+}99]. Orthograde [ZhTA^{+}97]. Ortholog [BZL^{+}98]. Oscillatory [YWMH99]. Osmolarity [RHLG98]. Osmotic [PH97a]. OSP [MSF^{+}99b]. OSP-based [MSF^{+}99b]. Ossification [YSEI^{+}99]. OST2 [SCKG95]. Ost3p [KKG95]. Osteoadherin [WSH98]. Osteoarthritis [SJF^{+}97]. Osteoblast [IKR^{+}98, CJH^{+}98]. Osteoblastic [JGCS97, KHHS97]. osteoblasts [ED96, SAPLHD95]. Osteoclast [AARS^{+}97, MNM^{+}98]. Osteoclastic [AMY^{+}98]. Osteoclasts [AMY^{+}98, BxQK^{+}99]. Osteogenic [REA95, YtDH^{+}95, PNL^{+}95]. osteonectin [AKP95]. osteoporosis [ED96]. osteoporosis-like [ED96]. Osteoprotegerin [BxQK^{+}99]. osteosarcoma [JH96]. other [SVB96, BF97b, SC97]. outer [BAC^{+}96, KPKWW95, MKLS96, PMH96, ROG^{+}96, WKK^{+}95, BVKP98, KEP^{+}99, KPW^{+}97, MS98, MBOC97, PKBHK97, SY99, VL99]. Outgrowth [BGBK99, AHM^{+}95, AC95b, BKB^{+}96, EWS96, GSC^{+}96, LW96, MLW^{+}95, NHB^{+}95, KNT^{+}99, MSS^{+}99, SLN^{+}97]. Outgrowth-promoting [MSS^{+}99]. Ovary [FNH99, BBZ^{+}95, HP95, ZOS^{+}97]. Overexpressed [NMH^{+}98]. overexpressing [DBvdBS95, RKD96, BNJ99, ESR97]. Overexpression [ASY96, AGP^{+}98, BCC^{+}95, BSK96, BENV97, EGS^{+}98, KKT^{+}98, LYH^{+}99, MDC^{+}99, PHO^{+}95, TEL^{+}95, APG^{+}96, KRMMA96, LLRTP95, SKWY95, WvdVV^{+}95, NT97, SSR^{+}97]. overgrown [TSS^{+}95b]. Overgrowth [CGSY^{+}99, GKS^{+}98]. overhangs [DH96]. overlap [YKC95, SZBB98]. Overlapping [HYK^{+}98, OP96c, OP96a]. Overload [SNK^{+}97]. override [MGRS96, ND96, ALM97b]. ovulation [IMFS96]. own [GGC96]. oxidant [NYN^{+}95]. Oxidase [WRdVC97]. oxidative [OCS^{+}95, HCLG99]. Oxidatively [MTW^{+}99]. Oxide [GM99, LHS97a]. oxidoreductase [RRRA95]. Oxygen [MFL^{+}99, HWFA96, TSL^{+}98].

P [DSvdV^{+}96, LKKL97, PNM95, SCM^{+}96, BHC99, CIRG99, GWB^{+}98, GRB^{+}97, GGD^{+}97, HMB^{+}98, JWW^{+}99, LVN^{+}98, MPB^{+}95, NSC96, RFCR^{+}97, SSEM98, SCH^{+}98]. P- [GGD^{+}97]. P-selectin [PNM95, SCM^{+}96, GGD^{+}97, NSC96, GRB^{+}97, MPB^{+}95]. p100 [SSS^{+}95]. p110 [SMS^{+}98, KLO^{+}98]. p115 [NAsG^{+}98, SW99]. p120 [ANRT99, SHT^{+}95, SSS^{+}95, YNG98]. p120-related [SSS^{+}95]. p120RasGAP [CGSP95]. p130 [CHP^{+}98, YTY^{+}99]. p150Glued [VV95]. p160ROCK [HIW^{+}98]. p185 [CWMD97]. p190RhoGAP [CGSP95]. p205 [PPWL97]. p21 [BBJ98, SBC99, BZBA99, KDO^{+}99, DSM^{+}97]. p21-Activated [SBC99, KDO^{+}99, DSM^{+}97]. p21CIP1 [JCPK95, FLW^{+}95]. p21Cip1/WAF1/CDK2/PCNA [FLW^{+}95]. p22 [KLH98]. p23 [MSW^{+}98, RPE^{+}97]. p24 [DDF^{+}98, LPD^{+}99, WG99]. p27 [SPV^{+}99, SSR^{+}98]. p270 [CRRF97]. p270/Tpr [CRRF97]. p28 [NNK^{+}97]. p29 [SJSJ95]. P2X [CSF^{+}97]. P2Z [CSF^{+}97, FWBSO97]. P2Z/P2X [CSF^{+}97]. p34 [NNSN97]. p34cdc2 [OHB^{+}95]. p35 [MFM^{+}95]. p36 [CGVS^{+}96]. p38 [CMS98, HHR^{+}98, IRW^{+}99, ZTH^{+}97, TMN97]. p38-dependent [HHR^{+}98]. p38MAPK [HTK^{+}99]. p42 [LBBP98]. p42/p44 [LBBP98]. p44 [LBBP98]. P450MT2 [AAB^{+}97]. p50 [BENV97]. p53 [CCMG97, IAS^{+}98, LNM^{+}98, MBF^{+}95a, SBR^{+}96, SBS^{+}96, TPMB97, ABP^{+}98, BRM^{+}99, FLA99]. p53-dependent [MBF^{+}95a]. p53-independent [CCMG97, LNM^{+}98, MBF^{+}95a]. p53-mediated [IAS^{+}98]. p53/58 [TPMB97]. p54 [BS95b]. p55 [VHF97]. p55CDC [KWD^{+}98]. p56 [BINRM97, BM99, LWSL^{+}98]. p56lck [RR96]. p58 [RGS99]. p58-Ctf13 [RGS99]. p58c [LFB96]. p58c-fgr [LFB96]. p59 [vtHR97, HR99, LFB96]. p59/61hck [LFB96]. p59fyn [GLS96]. p62 [BS95b, HGG96, PG95, MKB^{+}99]. p65 [FWBSO97]. P73 [FLA99]. p75 [ABP^{+}98, BMP^{+}98]. p80 [BFL95]. p85 [GPRS^{+}95]. p97 [CAA95, CAVA96, PMMF95]. packaging [KSL96]. painted [EDB^{+}96]. paints [CMWL96]. pair [YvdEH^{+}95, MS99, PYL98]. pair-sized [YvdEH^{+}95]. Pairing [FMD^{+}98, SWS^{+}96]. Pairs [PAF97]. PAK [TBP^{+}99, CF99, HB99]. Pak-Family [HB99]. PAK1 [DSM^{+}97, SBC99]. Paka [CF99]. PAL1 [SVT96]. Palmitoyl [KSY^{+}98]. palmitoylated [SKR96]. Palmitoylation [vtHR97]. PAM [CSEM99]. Pan1p [WE98]. pancreas [DMP^{+}96, CFB^{+}98]. pancreatic [RRH^{+}95, CCB^{+}98, ETP^{+}99, IMK99, KKLA97, MBCS98, TKY^{+}97]. Pantophysin [HKL96]. PAR-3 [IHT^{+}98]. paracellular [BWF^{+}96]. paracrine [HKM^{+}96, MLN^{+}98]. Paradigm [HPL99, WBR^{+}98]. Paradoxical [MAC^{+}97]. Paralemmin [KSY^{+}98]. Parallel [HB95a, PRB97]. Paralleling [RPS^{+}97]. Paramecium [EKF^{+}97]. Parameters [SB99a]. paramyosin [MAG^{+}96, LBO^{+}98]. paramyosin/miniparamyosin [MAG^{+}96]. paramyxovirus [BL96]. Paranodal [DGP99, EZC^{+}97]. Parasite [KQB^{+}98]. Parasitemia [WCV^{+}97]. Parasites [KBG^{+}99]. Parathyroid [SKS99, EMB^{+}96, SSA^{+}96, ANT^{+}97]. Parsley [KWN^{+}99]. part [PMH96, WW96a, WKR^{+}98]. Parthenogenetic [dS98a]. Partial [VKR^{+}96, RFNF96, WM95, PC99]. Partially [SRR^{+}97]. participate [WSBE^{+}95, PMP^{+}98, ÜHK^{+}99]. participates [BOLS95, GKVR95, GEK95, PMCT96, CIK^{+}99, LDP^{+}99, SRV^{+}97]. Participating [KND^{+}97]. Participation [BHG^{+}97]. particle [MDS95, MTLW95, VIF^{+}96, BPD99, BWWT97, CWT^{+}98, NWM^{+}97, PWW98, SNT^{+}98]. Particles [CDH^{+}98, OS96a, ARR^{+}98, KSV^{+}99, SWO^{+}99, SAEV^{+}98]. Partitioning [SHP^{+}97, MCB96]. Partner [KSE^{+}97, LSL99]. PAS7 [ZL95b]. Pas7p [ZL96]. passage [TMM^{+}95]. Passive [CHC97, GG95a]. Passively [GCST97]. Past [GCST97]. pastoris [EEHW^{+}98, RSB^{+}99, SKR^{+}98, WRdVC97, KSF^{+}99]. pat [GKR95, TWiK^{+}99]. pat-10 [TWiK^{+}99]. pat-3 [GKR95]. patches [WKWC96, KMMC98]. Patching [HSVS98]. path [YvdEH^{+}95]. Pathway [AAB^{+}97, LR98, MAT^{+}98, PBS97, SCR98, SVH^{+}98, SKC97, CMM96, CCBK96, DPH95, HMSK95, HG95a, HSK^{+}96, HDK96, ISH95, JRLS95, LLN^{+}95, MW95b, MLO^{+}95, PSW95, SRR96b, SJHC96, TCM96, VOP^{+}95, YGK96a, dPSMNSM95, ATD97, APRB98, ANT^{+}97, BM99, BG98, CFL^{+}98, CMS98, DBOM98, DKLS^{+}99, DOP^{+}98, FhJZ^{+}99, FBM99, FFLP99, GFM^{+}98, HC97, KSOK97, KLS^{+}99, KKLA97, KGE^{+}98, LPM^{+}98, LTR^{+}97, LNM^{+}98, LHWH97, LTW^{+}97, LMB98, MM97a, MMK^{+}97, MSWL^{+}99, Nei98, PRB97, PWGG98, PYL98, RD97, RHLG98, RRK97, RPB97, SWK^{+}99, SVF^{+}98, SGR97, SWS97b, SSRP99, SDHM99, SHL97, SWR^{+}98, SGYH97, TWSD98, TPMB97, VMT^{+}99, WBS^{+}98, WJM^{+}99, ZTH^{+}97]. Pathways [BOS^{+}97, HSML98, MM99b, MC97, MSMC99, TMN97, HB95a, JH96, KI96, PNL^{+}95, RRP^{+}95, Car99, FEF^{+}97, GTY98, JMWG98, JR98, LAS^{+}97, LGGS97, MGP97, OT97, PFDL97, SFR^{+}98, TKY^{+}97, vMNS97]. patients [WNB^{+}95]. pattern [BEG^{+}95, BSK96, CMM^{+}95, MAG^{+}96, YKC95, CHH97, FZZN98, HHB^{+}97, RV98]. Patterning [RV98]. Patterns [CP95, BLB^{+}96, CLT^{+}96, MF96b, EAM^{+}97, NHA^{+}97, SRCC97, YMD99]. Paxillin [TBP^{+}99, BPT96, SOHP95, CLRR99, NJB^{+}98]. Pay32p [STVR95]. PC1 [CSEM99]. PC12 [BHC99, CGM^{+}95, CEZA97, DHT96, FHK97, HGI^{+}99, LE97, LZC^{+}95, MLW^{+}95, MQR^{+}97, NSC96, QWS^{+}98]. PC2 [MZL97]. PC5 [DMM^{+}96]. PCNA [FLW^{+}95]. PDGF [BGR^{+}95, GEJ^{+}95, HRD98, KAAFZ95, SR96, XC96, XZSC96]. PDGF-induced [BGR^{+}95]. PDGFs [GEJ^{+}95]. PDMP [KBF^{+}98]. PDR1 [KGA^{+}97]. PDR3 [KGA^{+}97]. Pdr6p [TB99]. Pds1p [YGK96a, YGK96b]. PDZ [FPF^{+}99, KFM99, MNS^{+}97, MCL^{+}96, RBB97, TNM^{+}99, XWK^{+}97, XCLM98]. PDZ-containing [KFM99, RBB97]. PDZ-motif [FPF^{+}99]. PDZ1 [LBCB96]. PDZ1-2 [LBCB96]. Pea2 [VH96]. PEB1 [ZL95b]. Peb1p [ZL96]. PECAM [FSDA97, PHU^{+}95]. PECAM-1 [FSDA97, PHU^{+}95]. PECAM-1/CD31 [FSDA97]. pEg7 [CLL^{+}98]. Pelizaeus [GSL98]. Pellucida [GRB^{+}97]. Pemphigoid [SBG^{+}98, HBJ95, BKN^{+}97]. pemphigus [RS95, KMI^{+}97]. Pendulin [KF95]. Pentameric [UvMJ^{+}99]. Pep12p [vMNS97]. Peptidase [RL99, KRH98]. Peptide [TOSF^{+}98, CKM95, LXK^{+}95, NKI^{+}95, PKR95, RZZ96, SSA^{+}96, ZL96, ANT^{+}97, FRT^{+}98, FKWP97, GAZ^{+}98, MZNR98, MBDS^{+}98, SKS99, WG97]. Peptide-binding [TOSF^{+}98]. peptide-specific [NKI^{+}95]. Peptides [RL99, SOHP95, SLS^{+}96, DPM^{+}97]. peptidyl [WSW95]. PER8 [TWVC95]. percentage [STP96]. perform [DBBB95, KM95]. perfringens [KIH^{+}97, SFS^{+}99]. Pericellular [CMD^{+}98]. Pericentrin [DZS^{+}98, PTVD99]. pericentriolar [URU^{+}96]. Pericentromeric [SSW^{+}98]. perinuclear [SZB^{+}95, WWP^{+}96]. perinucleolar [MFMW95, HDES97, HDES98]. Peripheral [TRJ^{+}95, BSJJ95, FSIG^{+}95, GCM^{+}95, MTLW95, SRR96a, ESR97, PFAF97, VCL^{+}98, YHK^{+}97]. Peripherin [BNJ99, LSL99]. periphery [FUB^{+}96, IM96, KLN^{+}96]. Periplakin [RHPW97]. Perlecan [CGA^{+}99, FSA^{+}99, Ols99, PXRR99]. Permanently [TGL^{+}98]. permeability [BWF^{+}96, EELF^{+}95, KGW^{+}95, BSRG99, EAOS^{+}98, WG97]. Permeabilization [KEP^{+}99]. Permeabilized [MEFH97, FB96, NJB^{+}98, SFA^{+}98, KMK^{+}98]. Permease [BSH99]. permeases [KSL96]. peroxidase [KRMMA96]. Peroxin [ESR97, HRA^{+}98]. Peroxins [PYL98]. Peroxisomal [KSF^{+}99, WRdVC97, DMG96, EKK^{+}96, EB95, EB96, MKL^{+}95, MDQG96, PL96, SSY^{+}96, SVT96, TWVC95, WNB^{+}95, ZL95b, ZL96, CWSG99, EEHW^{+}98, GRS^{+}99, HRK^{+}98, HRA^{+}98, PYL98, WWD^{+}97]. Peroxisome [PAL^{+}98, SKR^{+}98, SG99a, DEG^{+}96, GKM^{+}96, LOD96, MLO^{+}95, FH97, TSSR98]. peroxisome-to-mitochondrion [LOD96, MLO^{+}95]. peroxisomes [EB95, MLO^{+}95, PL96, ZL95b, ZL96, ESR97, GRS^{+}99, SG99a]. Persist [GWLH98]. persistent [AC95a, XZMD97]. perspective [MW95a]. Pertinent [PAN97]. perturb [KKD^{+}95]. perturbations [PHO^{+}95]. Perturbs [WG97]. PEST [ALCC^{+}99, RSD98, SDC^{+}97]. PEST-like [RSD98]. PETA-3 [YMAC^{+}98]. PEX [DG96]. Pex10 [CWSG99]. Pex11p [MDQG96]. Pex12 [CWSG99]. Pex13p [EKK^{+}96, EB96, GRS^{+}99, GKM^{+}96]. Pex14p [GRS^{+}99]. PEX16 [ESR97]. Pex17p [HRA^{+}98]. Pex18p [PYL98]. PEX2 [FH97]. Pex20p [TSSR98]. Pex21p [PYL98]. Pex22p [KSF^{+}99]. Pex4p [KSF^{+}99]. Pex5 [CWSG99]. Pex5p [DG96, EKK^{+}96]. pf [ROG^{+}96]. PF16 [SL96]. PH [BLL^{+}96, MBA97, MSMC99, AGPG96, CLF^{+}97a, DDWG96, HCM^{+}97, KLW^{+}96, LTE^{+}96, PRV98]. Ph-Dependent [MSMC99, AGPG96, LTE^{+}96]. phage [PKR95]. phagocytosed [OS96a]. phagocytosis [AJS96, BLLB95, FWG^{+}96, PBJ^{+}98]. Phagosome [ZMG^{+}96a, PHB96]. Phagosome-lysosome [ZMG^{+}96a]. Phagosomes [BSB^{+}97]. phakinin [GGR^{+}96]. Phase [BMG^{+}99b, FK95, TALM96, BZBA99, DTMG99, HYN97, JP98, MCS^{+}97a, OWD^{+}99, SIUW98, TGN97, YGG97b, DFGL96, OSH^{+}96, OHB^{+}95]. Phases [TXB^{+}99, NMJL96, VJHR96, FCMN97]. phenomenon [MLW^{+}95]. Phenotype [HTK^{+}99, KJE^{+}99, KMI^{+}97, BKP^{+}95, CRM^{+}96, ED96, FGA^{+}96, ICW^{+}96, MAFJ^{+}95, SSM^{+}95, WPZ^{+}96, BAS^{+}97, EIIM^{+}98, LGM^{+}97, PC99, SBPR^{+}98, WPW^{+}97]. phenotypic [AW96, IFSV^{+}98]. Phenylephrine [CMS98]. Pheromone [ELMS98, BGN^{+}96, DPH95, BBR97, CSCM97]. Pheromone-inducible [BBR97]. Pheromone-regulated [ELMS98]. pheromones [VPH95]. Pho80 [NWPW95]. Pho85 [NWPW95]. Phorbol [SOPM^{+}97]. Phosphatase [MN98, NNSN97, PBS97, All95, AS96, BLE^{+}96, BKRT95, GZK^{+}95, KSR96, LDF95b, SNCBM95, TFF^{+}95, ZL95a, ZMG96b, ALVMM98, ALCC^{+}99, BALL98, BATP99, DP98, FKO^{+}98, KFO^{+}99, LHB^{+}99, OTS98, RTP^{+}98, SLN^{+}97, SDC^{+}97, SHL97, TYY^{+}99, ZDK^{+}99]. Phosphatase-1 [ALVMM98]. Phosphatases [ALM^{+}97a]. phosphate [BGR^{+}95, KDvFP96, dPF95, RWO95, RSJK95, SKR96, KKG^{+}98, LH97, OKE^{+}99, PFHWN98, VLM^{+}98]. Phosphate/Insulin [NCV^{+}98]. phosphatidylcholine [dCVCV95]. Phosphatidylethanolamine [KGA^{+}97]. Phosphatidylinositide [BGN^{+}97, CMC^{+}98]. Phosphatidylinositol [ANS^{+}97, JH97, PBM^{+}97, SvDtK^{+}98, BDE^{+}95, Dav95, GHF^{+}95, HSK^{+}96, LMB^{+}95, SMFC95, TCM96, vdBCH^{+}95, GGPG99, KLS^{+}99, MAE^{+}97, PMKR97, SMS^{+}98, VMT^{+}99]. phosphatidylinositol-anchored [vdBCH^{+}95]. phosphatidylinositol-dependent [LMB^{+}95]. phosphatidylserine [HP95]. Phosphoantigen [SVH^{+}98]. phosphoepitope [CG95]. phosphoglycokeratins [KMOO95]. Phosphoinositide [HTK^{+}99, VB98, AJS96, GPRS^{+}95, EPMB^{+}99, GK99, TFYH99]. Phosphoinositides [VB98, MCJK98]. phosphoinositol [PGS95, GSB97]. Phospholipase [CSA^{+}97, HG95a, KBW^{+}96, CRM^{+}97, GSKL99, MGP97, RME98, SBGR98, qSmLY97, WCV^{+}97, WBR97]. Phospholipid [qSmLY97]. phosphooligosaccharide [PGS95]. phosphooligosaccharide/phosphoinositol [PGS95]. Phosphoprotein [CT97a, DPC^{+}96, FSHD96, HM95, STP96, KSY^{+}98, RBB97]. Phosphoregulation [KD97b]. Phosphorylated [TXB^{+}99, BCMK95, CDRB^{+}96, DK96, LLK^{+}95, LO96, MRM^{+}98, MOY^{+}98, SDC^{+}97]. Phosphorylates [MMD^{+}98, GHF^{+}95]. phosphorylating [SIAS96]. Phosphorylation [FIC^{+}97, FOK^{+}99, JCR97, KSV^{+}99, KFO^{+}99, KBS^{+}99, MBA97, SAR^{+}99, SJA96, YTY^{+}99, YTO^{+}99, AOMB96, CH95, FDH96, GHP^{+}96, GW95, JGM96, KCDB95, KRMMA96, LCK95, LCS96, LLK^{+}95, MK95a, MBS95, MPW95, MTGY96, SHPP96, SR96, TKVC95, TOT^{+}96, TEL^{+}95, WB96, XGC96, ZOB^{+}96, BKJL^{+}98, CCS^{+}98, FKO^{+}98, GRTB97, GSKL99, HS97, HZR98, HFP^{+}98b, JK97, KGL^{+}99, KTYT99, KMS^{+}98, LWS99, MLB98, MSEP98, MCC^{+}98, RHW^{+}98, SFS^{+}97, SC97, SHB99, SMTN99, TGGB97, TYY^{+}99, qXLCH97, YAiN^{+}98]. Phosphorylation-dependent [KSV^{+}99]. Phosphorylation-induced [SAR^{+}99]. phosphorylation-regulated [AOMB96]. phosphotyrosine [NYN^{+}95, HR99]. photoactivated [FTH96]. photoreceptor [PFF^{+}95, MAE^{+}97]. Photoreceptors [KWN^{+}99]. photosystem [MML96, RKB97]. Phototactic [KD97b]. phototransduction [PSW95]. Phragmoplast [SG95]. physically [NRM^{+}96, FNDS99]. physiologic [FAP^{+}96]. Physiological [RRK97, FHCL95]. Phytochrome [KWN^{+}99, YNM^{+}99]. PI [SJHC96, SKC97]. PI-3 [SJHC96, SKC97]. PIC1 [SJW97]. PIC1/SUMO [SJW97]. PIC1/SUMO-1 [SJW97]. Pichia [EEHW^{+}98, KSF^{+}99, RSB^{+}99, SKR^{+}98, WRdVC97]. PICK1 [SZB^{+}95]. Pickpocket [AAM^{+}98]. Pigment [BFRB99, MGB^{+}98, TZL^{+}98]. pinin [OS96b]. Pinocytic [DSM^{+}97]. pinocytosis [DBvdBS95, NPRT95]. pit [SK96, FMP^{+}97, KsYA98]. Pits [EGKC^{+}99, GK99, LS95a, LNR^{+}96, NDBR95, SSEM98]. pivotal [KWK^{+}96]. PIX [TBP^{+}99]. PKA [RTP^{+}98, ZH97, BATP99]. PKB [BRM^{+}99]. PKC [RTP^{+}98, IHT^{+}98, ZH97, ZLPL99, TGW99]. PKD1 [MMS^{+}96b]. place [LK95]. placental [BGB95b]. Placentopathy [MCS98]. Plakin [RHPW97]. plakoglobin [CLT^{+}96, LCC^{+}95, RBL^{+}96, SGYL^{+}96, KBB^{+}97a, LWS^{+}97b, SSR^{+}97, SSS^{+}98, WAV98]. Plakophilins [MKF96]. planar [EWS96, MWC95]. plant [AGB^{+}96, SG95, JR98]. plants [SGS95]. plaque [BCSG96, KSK^{+}96, MKF96]. Plasma [AvdG99, GK99, vtHR97, KGA^{+}97, MBA97, MM99b, NMNL98, SBS^{+}99, BGB^{+}95a, BLL95, CF95, CSY^{+}95, FLPS95, GHJJ95, GLS96, HSK^{+}96, LLL^{+}96, MBS96a, MM95b, NAP^{+}96, NCS95, PNM95, SK95a, SLS^{+}96, WGR^{+}95, YKRS96, BINRM97, BM99, DSvDH^{+}98, GWLH98, HSVS98, HME^{+}98, IMS^{+}98, KSY^{+}98, LRAB99, LCW^{+}98, jLC97, MS97a, NTH98, OMS98, PKCS98, RD97, RCE^{+}99, SHH97, TMM97, TFYH99, YPM97, ZH97]. plasmalemmal [RRH^{+}95]. Plasmin [SPK^{+}97, CMD^{+}98, GSJ98]. Plasminogen [NCW^{+}99, SPK^{+}97, DCW^{+}96, SM95c, CMD^{+}98, EvdESvdH^{+}97, GSJ98]. Plasmodium [BJR^{+}99, GDF^{+}96, WSG^{+}98]. Plasticity [MAFJ^{+}95]. plastid [RRRA95]. plastin [AAP^{+}95]. plate [SGS95]. Platelet [FSDA97, GXE^{+}99, KSE^{+}97, LCN^{+}97, BWKH96, GHJJ95, FHSM^{+}97]. Platelet/Endothelial [LCN^{+}97]. Platelets [KBS^{+}99, GPRS^{+}95, HOS^{+}97, ILSH99, PWD^{+}99]. Platforms [AvdG99]. play [SSW^{+}96]. plays [CGJ^{+}95, SWW^{+}95, YGK96a, FOK^{+}99, KHA^{+}99, MBCS98]. Pleckstrin [MBA97, MSH^{+}97, VB98]. Plectin [RdPRW98, SBG^{+}98, SVB96, NNMW96, GFN^{+}99, GB98]. plectin-vimentin [NNMW96]. Pleiotrophin [MN98]. Pleiotrophin/Heparin [MN98]. Pleiotrophin/Heparin-binding [MN98]. Plk1 [LN96, GMFN95]. Ploidy [BHJ98]. Plus [MGRS96, WSGPS95, YNOH95, DPG^{+}99, PDS^{+}97a, SNK^{+}99, TWS97, WFS97]. Plus- [SNK^{+}99]. Plus-end [MGRS96]. Pml [ISN^{+}99, CSO^{+}95, SJW97, ISN^{+}99]. Pmp27 [MKL^{+}95]. Pmp27p [EB95]. PMP47 [DMG96, SSY^{+}96]. Pmt4p [SGTH99]. Pocket [AAM^{+}98]. Podocytes [KRK^{+}98, MHM^{+}97]. Point [BGH^{+}97, NH95, VPM^{+}98]. Pointed [EGM^{+}99]. Poisson [VBZ96]. Poisson-distributed [VBZ96]. Polar [SGD98, KM95, KLS^{+}99, SdCAH^{+}97]. Polarity [KRK^{+}98, ZT99, CKP96, FSR^{+}96, HE96, MSF^{+}95, RSP^{+}95, SYB95, SM95b, WHP^{+}96, ASP^{+}97, CSM97, IHT^{+}98, JN98, LNRR98, MGB^{+}98, NH99, SN98, SCEB99, VMH97, ZH97]. Polarization [LKSW98, EWS96, GAW^{+}96, VPH95, ZZvdB^{+}96, dPSMNSM95, ABE^{+}98, DDRC99, HB99, MMR99a, SALdP^{+}97]. Polarized [CSN99, HKC^{+}95, WSCN97, YLB^{+}97, vIH98, HFSHK96, LCM^{+}95, LZD95, LLK^{+}95, MGM95, MVvdBD96, MR96, MW96, MXSRB96, OPD^{+}96, PCL^{+}96, PAV^{+}96, SMRT96, VH96, AL97, BMG^{+}99a, CS98, FGA^{+}98, GFM^{+}98, IMS^{+}98, PSB98, SFD^{+}98, Sal99, SS97, TFYH99, WJWM^{+}98, ZSP^{+}98]. Pole [WH97, WJH^{+}98, DK96, FSHD96, GSB^{+}96, HY95, KR96, SCG^{+}95, SGGS96, SGBD96, WW96a, AK99, CYH98, CRGW98, FFLP99, LM98, MB97, MC97, MUS98, SRV^{+}97, SGSW97, SMM99]. Poles [GDC97, MCTMK98, SHLD97, WBA^{+}98]. poleward [KCBR96, DMMS98]. Pollen [FSH^{+}99, KLS^{+}99]. Polo [GOT96, LN96, BSW^{+}98, CRM^{+}98]. polo-like [LN96]. Poly [KSK^{+}99a, YCR^{+}98, HCA^{+}95, HB95b]. Polycomb [BHSAJ98, SSW^{+}98]. polycomplex [SR95]. polycystic [MMS^{+}96b]. Polyglutamine [PPP^{+}98]. Polyglutamine-mediated [PPP^{+}98]. Polymannose [SPBCM97]. Polymannose-type [SPBCM97]. polymerase [BDvdZW95, JMTCF96, MFMW95, DW97, KSK^{+}99a, KDBW97, VLAHV98]. polymeric [CSM^{+}96, dHvPL^{+}95]. Polymerization [FWL^{+}98, CGJ^{+}95, Cun95, GW95, MELC96, WLG^{+}96, CWY99, MH97, ZJB^{+}97, ZJY^{+}98]. polymorpha [TWVC95]. Polymorphonuclear [DZC^{+}96]. Polyoma [SPK^{+}97]. polypeptide [NHB^{+}95, SOE^{+}95, BG97]. polypeptides [BSMH95, MML96, GBS99]. polyphosphoinositides [SJC96]. Polyploidization [NMT97]. Polyposis [BPA^{+}97, NAP^{+}96, VWG^{+}97]. polypyrimidine [MFMW95]. polysialic [WPS^{+}96]. polysialylation [RL96, BR98]. polytene [BJB^{+}96, UPAS95]. Polyunsaturated [SBS^{+}98]. Pom152p [ARM^{+}95]. pombe [MC95, DS98b, FFDP98, KSY97, LW95, MBP^{+}95]. POMC [CSEM99]. Ponsin [MNS^{+}99]. Ponsin/SH3P12 [MNS^{+}99]. Ponticulin [SWW^{+}95]. pool [SNCBM95, BR98, PFP97, SDL98]. Pools [VB98]. Population [BLB^{+}96]. Populations [JMM^{+}98, SWS97a, SDHM99]. Porcine [GRB^{+}97]. Pore [BdPE^{+}97, ARM^{+}95, BLEB96, BSD^{+}96, GDC95, GEW^{+}95, HCA^{+}95, HGG96, LBL95, MF96a, MCB96, MDS95, MMF95, MGY^{+}95, NRM^{+}96, PG95, PZ96, PMMF95, ZDT^{+}96, AFG^{+}98, BBP^{+}98, BSRG99, BD97, BW97, CFL^{+}98, CRGW98, CRRF97, EAOS^{+}98, FHAP98, KSV^{+}99, KDK^{+}99, KIT^{+}97, LBWF^{+}97, MAW98, MWB98, MBDS^{+}98, PFKvD99, dCBR99, SAEV^{+}98]. Pore-forming [AFG^{+}98]. pore-membrane [ARM^{+}95]. pores [BS95b]. Portion [HTY^{+}99]. position [RD96a, LKE^{+}99]. position-specific [RD96a]. positional [SWW^{+}95]. Positioning [LSB98, SCR98, FT95, PWG^{+}95, RSP^{+}95, CH97, EAM^{+}97, GPKH99]. Positions [BMS^{+}97]. Positive [DDWM97, MS98, KKG^{+}98, RVD^{+}95, WXS96, MSL^{+}98a]. Positively [YHD^{+}98]. Possess [FSH^{+}99]. Possessing [KCM^{+}97]. Possible [SFS^{+}97, BKP^{+}95, HSK^{+}96, MAG^{+}96, OSH^{+}96, WS95b, LLSO98, PRB99, SSZ^{+}97].

Post [SW99, WSCN97, WBKB97, PCL^{+}96, SIAS96]. post-adhesive [PCL^{+}96]. Post-Golgi [WSCN97, WBKB97, SIAS96]. Post-Mitotic [SW99]. Posterior [LM98, SCEB99]. Postlysosome [BGN^{+}97]. postmitotic [CM95b, MAM95, SBR^{+}96]. Postnatal [DBC^{+}99]. postprophase [GACM^{+}95]. Postsynaptic [DPT^{+}97, ZGYS99]. posttranslational [BHP^{+}95, SMB^{+}95]. Potassium [HCPW^{+}99, SG97]. Potassium-Dependent [HCPW^{+}99]. potent [BGGR^{+}96]. Potential [BZL^{+}98, SJHC96, BZB^{+}96, CLT^{+}96, HTKH96, OHB^{+}95, PTBC96, WXS96, CIRG99, CMC^{+}98, GRB^{+}97, GSP^{+}98, HKKH99, NB97, PGD^{+}97, SGGS^{+}99, SSS^{+}98, TSH^{+}98, ZLM^{+}98]. potentiates [HIG^{+}95, NJB^{+}98]. POU [GBW98]. Powered [PdHTV99]. PP1 [APH^{+}97]. Pp125 [CLRR99]. pp125FAK [HCM^{+}96b]. PP2A [RTP^{+}98, MTK^{+}98]. pp60 [ZKN99]. pp60c [RVD^{+}95]. pp60c-src [RVD^{+}95]. pRB [FSAP95]. Pre [ALC^{+}99, WLD^{+}98, YTO^{+}99, ZWT^{+}97, DKLS^{+}99, DW97, KDBW97, NSR95, SMW^{+}99, SAEV^{+}98, VLAHV98, BJB^{+}96, HS96, OK96, VIF^{+}96, XJM^{+}95]. Pre-Golgi [ALC^{+}99, ZWT^{+}97]. Pre-mRNA [WLD^{+}98, YTO^{+}99, DW97, KDBW97, NSR95, BJB^{+}96, HS96, VIF^{+}96, XJM^{+}95]. Pre-mRNAs [SMW^{+}99]. Pre-mRNP [SAEV^{+}98]. Pre-rRNA [DKLS^{+}99]. Pre-rRNAs [VLAHV98]. Preanaphase [SCR98, SHLD97]. precede [AW96, GDM^{+}99]. precedes [EBR96, FKS95, VCJA99]. Prechondrogenic [ZOKS99]. Precise [LSB98]. Precocious [RFCR^{+}97]. Precursor [ALC^{+}99, CGC^{+}97a, FKNM97, AHM^{+}95, AM95b, HKC^{+}95, MKLS96, RRRA95, RHS^{+}96, SHG96, WPS^{+}96, YSK95, CJH^{+}98, DMY^{+}97, DPM^{+}97, FBM99, GR97, KFL^{+}98, MNR^{+}98, SATW98, YHO^{+}98]. Precursors [AARS^{+}97, BMPP99]. Predisposes [KMS^{+}98]. Predominant [UNPW98]. predominantly [GKM^{+}96]. Preexisting [SG99a]. Preference [BSS^{+}97]. Preferential [LCH96, ZL95a, DMY^{+}97, OPB98]. preferentially [KLN^{+}96, SOJM99]. Prefoldin [HCW99]. Prefusion [DFMG97]. Preinitiation [HN98]. Premalignant [LGM^{+}97]. Premature [GCM98]. premeiotic [OABD^{+}96]. Premyelinated [SZBB98]. Premyelinating [TNCM97]. Prenyl [KSY^{+}98]. prenylation [PFF^{+}95]. Preproteins [KS97]. Prereplication [DTMG99]. prerequisite [LOD96]. Presaged [LTR^{+}97]. Presecretory [NZ97]. Presence [DH96, VLAHV98, MBRN95, HVL^{+}99, LHR^{+}97]. Presenilin [ALC^{+}99, IMW^{+}99, JM97a, SOJM99]. Presenilin-2 [JM97a]. Presenilins [YF99]. present [WSBE^{+}95, ESR97, SBGR98, UvMJ^{+}99]. Presentation [APF^{+}97, BJS^{+}95]. Presented [SWD^{+}98]. Preservation [BHJ98]. preserve [OCS^{+}95]. Pressure [RSS98, SNK^{+}97]. Presumptive [LSM^{+}98]. Presynaptic [DSVL^{+}98, THT^{+}95, WKL^{+}99]. Prevacuolar [BPWS98, VS98, CS96b, PCYS95]. Prevacuolar/Endosomal [BPWS98]. Prevent [MWGE97]. prevents [CF95, MGS96, RSJK95, GFN^{+}99, GM99, KA98, OK98, SSV98]. Primarily [ZWT^{+}97]. Primary [SRA^{+}99, BLB^{+}96, FT95, LOD96, MLO^{+}95, VKR^{+}96, AGP^{+}98, BAC97, PKQ^{+}98, SGGS^{+}99, SSSS97, WDL^{+}98]. priming [CRMB95, MW97]. Primordial [HHB^{+}97]. principal [BPT96]. Principally [ILSH99]. prion [TSS^{+}95a]. privileged [FHV^{+}96]. Prk1p [CYSD99, ZC99]. Pro [FIC^{+}97, BGGR^{+}96, REdC95, CR99, GDM^{+}99, KEP^{+}99, OTL^{+}97]. pro-alpha [REdC95]. Probed [SG97]. Probing [EKS97]. Problems [Ols99]. Procaspase [NNK^{+}97, KAS^{+}99]. Procaspase-3 [KAS^{+}99]. Procaspase-8 [NNK^{+}97]. procathepsin [Dav95]. Proceeds [FMD^{+}98]. Process [KBI^{+}99, KRK^{+}98, MHM99, WSG^{+}95, KSY^{+}98, MWF^{+}99, PKQ^{+}98, VLAHV98]. processed [AAB^{+}97, BTNZ^{+}95]. Processes [RBM^{+}99, ZLPL99, EAL^{+}95, GSK^{+}99, ILSH99, SNK^{+}99, VTHAA99, YSK^{+}97, ZCOP99]. Processing [ALM^{+}97a, ALC^{+}99, BM98, FKNM97, GSJ98, RL99, SMW^{+}99, AHM^{+}96, WLR96, DKLS^{+}99, FBM99, GAZ^{+}98, LTW^{+}97, MCS^{+}97a, TNFK^{+}98]. Processing/ [MCS^{+}97a]. Procession [Blo98]. Processive [RPV98]. Processivity [HH98]. procollagen [BTNZ^{+}95, SHY^{+}96, ZOKS99]. Procyclin [RFK^{+}97]. Produce [NMNL98]. produced [RCKS95, SRR96a, ILSH99, LTMR97]. produces [WPZ^{+}96, MBF^{+}98]. producing [CMM^{+}95, REdC95]. product [ASSS^{+}95, BBOB96, HY95, PGC^{+}95, RRRA95, SVT96, BAH^{+}98, CAC^{+}99, FSFT98, PB98]. production [GL95, SIAS96, VWK^{+}95, VUC^{+}95, GM99, KRS^{+}98b, KNSP97, MCRB97, SB99c, TSL^{+}98]. Products [WBSNV97, BHP^{+}95, DBBB95, HKC^{+}95, MAG^{+}96, OK96, SMCE97]. Proenkephalin [BS95a]. Profilin [KJE^{+}99, STP96, ZPS96, CDN97, WKL^{+}99]. profilin-mediated [ZPS96]. Profilin-minus [KJE^{+}99]. progenitor [PNL^{+}95, SCM^{+}96, YYH^{+}96, PC98, PHN^{+}98]. progenitors [VDA^{+}96, DBC^{+}99, MPR^{+}97, WDL^{+}98]. Progeny [PGP^{+}99]. progesterone [FMJ^{+}95]. Program [GR97]. Programmed [RHP99, FMJ^{+}95, HG95b, JWR96, RME^{+}95, WJC^{+}96, AEWG^{+}97, TSL^{+}98]. programs [BFS^{+}95, LMMO96]. Progress [VWO99]. progression [LO96, LK95, TFF^{+}95, YYH^{+}96, ZOB^{+}96, Ass97, BOM^{+}98, CH98, DFMG97, DDRC99, GDTBD98, OWD^{+}99, YGG97b]. Progressive [DSM^{+}98b, LYD^{+}97]. Proinflammatory [HAB^{+}99]. projectin [ASSS^{+}95]. projection [BLL95, PNEH98]. Projections [MBA97]. Prokaryote [MHM99]. proliferating [YFST^{+}96, GMGGD98, GR97, LMP98]. Proliferation [CCMC97, DVL^{+}98, EBR96, VNTM^{+}99, AYF96, BSK96, CKM95, KF95, MKL^{+}95, MMMT95, SLT^{+}96, TIJ^{+}95, TWVC95, WHP^{+}96, BBAN^{+}99, BBJ98, EIIM^{+}98, KJR^{+}98, LPM^{+}98, MAC^{+}97, MFBK99, PWGG98, VLM^{+}98, WDL^{+}98]. proliferative [GHP^{+}96, SCG^{+}96, WXS96, CRYC99]. Proline [CHC97, QPH^{+}98]. Proline-rich [CHC97]. Prolonged [FLW^{+}95]. Prometaphase [MAEE98]. Prominent [wJTSSL98]. Promote [FIC^{+}97, JCR97, NCW^{+}99, XMMW97, CLR^{+}96, NA96, RSC96, BGR98, CLRR99, KRCP99, hKsKC99, LKKL97, PMKR97, RNK97, VZNR98]. promoted [CHP^{+}98]. promoter [BFP^{+}96, FWHW96, LMMO96, REdC95, TLF^{+}95]. Promotes [ELL^{+}99, SvDtK^{+}98, The97, AC95b, BKEC96, CKM95, LLRTP95, MKL^{+}95, PAV^{+}96, SSM^{+}95, WMC^{+}95a, BFRB99, BF97b, CAR^{+}97, DDM98, EPMB^{+}99, GWB^{+}99, KKS99, LNG^{+}98, LHB^{+}99, LPM^{+}99, NPJW99, NLM^{+}98, OKE^{+}99, SJF^{+}97, TOB^{+}99, WVM97]. Promoting [BRM^{+}99, GSC^{+}96, KWD^{+}98, OSH^{+}96, MLB98, AK97b, MSS^{+}99]. promotion [BKB^{+}96, TTAH^{+}99]. prone [SYO95]. Propagation [BRD99, JP98, MCPB99]. ProPC2 [MZL97, ZL95c]. Propeptide [ZOKS99]. Proper [OBB99, DG96, MM95a, OFY^{+}95, RHF^{+}96, SS95b, SGBD96, VUC^{+}95, CYH98, hKsKC99, SGSW97]. Properties [BMPP99, LLL^{+}96, CGM^{+}95, SWR^{+}96, KKH^{+}99, PBCS98, PFP97, YMPG98]. prophase [SWS^{+}96, BMS^{+}97, FdEP99, wJTSSL98, RC98, YWMH99]. Proplatelet [ILSH99]. proposal [BGB95b]. proposed [MDQG96, HCM^{+}97]. proprotein [DMM^{+}96, CR99, LTW^{+}97]. Propulsive [ODJ99]. Prospore [Nei98]. Prostate [RPS^{+}97]. Protease [SMH98, AHM^{+}96, CCR96, GTL^{+}96, LCH96, SRR96b, EM98, MDV98]. Proteases [MMAK97, MWGE97, TMN97, DVD^{+}96, JWR96, FCMN97, MCS^{+}97a]. Proteases/Caspases [MMAK97]. Proteasomal [ASB^{+}99, WFL^{+}99]. Proteasome [MB97, SSR^{+}97, TYT^{+}97]. protect [MFM^{+}95]. protective [WSW95]. Protects [MZNR98]. Protein [ACSN98, AAB^{+}97, ALM^{+}97a, All95, ALCC^{+}99, ALC^{+}99, AWC99, BPA^{+}97, BZL^{+}98, BSW^{+}99, BSAJ97, BDS97, CHLÅ97, CWSL97, CGC^{+}97a, CWM^{+}98, DSVL^{+}98, DSC^{+}95, ELL^{+}99, EGKC^{+}99, HTK^{+}99, HDV^{+}99, ISR99, KWD^{+}98, KNH^{+}97, KRK^{+}98, KSF^{+}99, LCS96, MN98, MNS^{+}99, MWOB96, MB99, MAM99, MSMC99, NNK^{+}97, NNSN97, OBB99, PJE^{+}99, PPWL97, PBS97, PT97, PPBJ96, RTM99, RST^{+}99, SPV^{+}99, SGD98, SW99, SWL^{+}97, TNM^{+}99, TYT^{+}99, THM98, TOSF^{+}98, WH97, WLD^{+}98, WBKB97, WC97b, XC97, YHS^{+}99, ZDK^{+}99, ZWT^{+}97, vHGB^{+}97, AOMB96, ASH95, ARM^{+}95, ADP^{+}95, AHM^{+}95, AC95b, AM95b, BL96, BWF^{+}96, BWKH96, BWG95, BLEB96, BLG95, BTTB96, BSK96, BLLB95, BAC^{+}96, BCSG96, BS95a, BKRT95, BSJJ95, BM96, CRLM96, CBB^{+}96, CAA95, CAVA96, CMHT96, CCBK96, CCP96, CKS^{+}95, DMS^{+}95, DvD95, DTB96, DRR96, DWS^{+}95, DG96, DMG96, EKK^{+}96]. protein [EB95, EGH^{+}95, FAO^{+}96, FSR^{+}96, FA95, FT95, FFZ^{+}95, GCGE^{+}96, GLD^{+}96, GZK^{+}95, GEK95, GLP^{+}95, GMFN95, GKM^{+}96, GEW^{+}95, GLB^{+}95, HTS^{+}95, HKC^{+}95, HKS^{+}96, HE96, HMSK95, HSWC96, HD96, HM96, HIG^{+}95, HSK^{+}96, HFSHK96, HUU96, HCM96a, ISFB^{+}96, IRMF^{+}99, ISH95, JSR96, JLRS95, KTC95, KE98, KSK^{+}96, KPKWW95, KGMF95, KBR95, KSL96, KKE^{+}95, KF95, KSR96, LJ95, LWM^{+}95, LDCB95, LGRB96, LDF95b, LYR96, LS95b, MKLS96, MBS96a, MCL^{+}96, MJMS96, MTL^{+}96, MFMW95, MMGT96, MCB96, MBM95, MGS96, MBP^{+}95, MB96, MPW95, MGY^{+}95, MMMW96, MMS^{+}96b, MXSRB96, NRW^{+}95, NIM95, NAP^{+}96, NW96, NMJL96, OWA95, OIT^{+}95, OMO95, OHB^{+}95, OP96a, OABD^{+}96, OS96b, PSBB96, PWK^{+}96, PG95, PMH96, PFF^{+}95, PMMF95, PMCT96, RL96, RSC96, RRRA95, RLCM96, RFNF96, RRH^{+}95, RW95, SLL96, SH96, SHY^{+}96, SPMB^{+}96]. protein [SJC96, SLW^{+}96, SFD^{+}96, SRH^{+}95, SLEB95, SK95b, ST96, STT^{+}96, SMB^{+}95, SRG^{+}96, SCKG95, SIAS96, SBW^{+}96, SD95, SL96, STP96, SBS^{+}96, SOR^{+}96, SNCBM95, SW95, SCG^{+}95, SDTE95, SSS^{+}95, SM95c, SZB^{+}95, SJSJ95, SSF^{+}95, SKK95, SL95, SG96, STVR95, TOT^{+}96, TWVC95, TET^{+}96, TSS^{+}95a, TSS^{+}95b, TFF^{+}95, VH96, VXBH95, VJHR96, WGR^{+}95, Wal95, WA95, WSG^{+}95, WFBC96, WNB^{+}95, WKK^{+}95, WRW^{+}95, WF96, WHP^{+}96, WM95, WVVD95, XGC96, XZSC96, YtDH^{+}95, YNOH95, YSK95, ZL95b, ZOB^{+}96, ZMG96b, vdLST^{+}96, AFG^{+}98, ATH^{+}97, ATD97, AAD^{+}97, AKLME98, ANK97, APH^{+}97, ALVMM98, BALL98, BSD97, BFHB97, BAC97, BRS98, BBR97, BHHF97, BRS99, BHJ98, BM99, BH98, BF97b, BS97, BSM^{+}99, CHC97, CDN97, CWSG99, CJH^{+}98, CSMM98, CC98, CEZA97, CMS98, CLT99, CRYC99, CR99, CYSD99, CRRF97, CLL^{+}98, CLF^{+}97b]. Protein [DRE97, DBOM98, DCJ97, DZS^{+}98, DKN^{+}97, DDF^{+}98, DDRC99, DFL^{+}99, EGS^{+}98, EAM^{+}97, EGM^{+}99, EJ97, EHC97, EZC^{+}97, EEHW^{+}98, EKSC98, ESM^{+}97, EIIM^{+}98, EGA97, EKS97, FhJZ^{+}99, FSC^{+}97, FPF^{+}99, FNDS99, FJ98, FAZ^{+}99, FMM^{+}98, GE97, GKP^{+}97, GCML98, GMS^{+}98, GFM^{+}98, GLL^{+}99, GGPG99, GRS^{+}99, GBW98, GDL98, GCM98, GF97, GSL98, GLMG98, GZC^{+}99, GB98, GDM^{+}99, GTY98, HRR^{+}99, HSW^{+}98, HPL99, HBS^{+}97, HH98, HC99, HGW^{+}98, HKO^{+}98, HCM^{+}97, HPD^{+}97, HGI^{+}99, HRK^{+}98, HTY^{+}99, HIW^{+}98, HME^{+}98, HMC^{+}99, HTW97, HYE^{+}98, HSSW99, HC97, HFP^{+}98b, sHWM^{+}99, HIWL97, HFE^{+}98, HRA^{+}98, HKKH99, ISN^{+}99, IHT^{+}98, JWW^{+}99, JR98, JL98, JB99, KRH98, KMM^{+}98, KHA^{+}99, KOM^{+}98, KSK^{+}99a, KSOK97, KLO^{+}98, hKsKC99, KSK^{+}99b, KcZS97, KCG^{+}97, KKG^{+}98, KB99b, KFL^{+}98, KWS^{+}99, KS97, KCFS98, KPW^{+}97, KLL^{+}97, KGE^{+}98, KNT^{+}99]. Protein [KGS98, KBC99, LM98, LWS^{+}97a, LSL99, Li97, LXC^{+}98, LMW99, LB98, LLNH^{+}98, LL98a, LL98b, LTW^{+}97, LHS97a, LBO^{+}98, LJK97, LFP98, LPM^{+}99, LHS97b, MCJK98, MM97a, MSA98, MAEE98, MNS^{+}97, MBW^{+}99, MSS^{+}99, MGIZ98, MLB98, MHH^{+}99, MCTMK98, MWB98, MS98, MDK^{+}99, MPR^{+}97, MVM^{+}99, MR98, MMR99a, MAE^{+}97, MYC^{+}98, MKB^{+}99, MBB^{+}97, MPT^{+}98, MNG^{+}98, MMR99b, MJBR98, MSH^{+}99, MHM^{+}97, MDC^{+}99, NMK^{+}97, NMH^{+}98, NOS^{+}97, NTAN98, NA99, NZ97, NIK^{+}97, NB97, NGMR97, OWAlW99, OTS98, OTQM98, ONI^{+}98, OFM^{+}98, PÖS^{+}98, PGD^{+}97, PRB97, PRB99, PBJ^{+}98, PH98, PHN^{+}98, PBA^{+}97, PLB97, PHA^{+}98, PCD^{+}98, PB98, RLVM98, RTP^{+}98, RKB^{+}98, RRK97, RSS98, RPE^{+}97, RHA^{+}97, RPB97, RPBB98, RGS99, SSI^{+}97, SHC^{+}97, SdCAH^{+}97, SYT97, SSS99a, SVB^{+}97, SMCE97, STF98, SYF^{+}97, SY99, SKY^{+}98, SSSG98, SPCR97, SDL98, SEBW^{+}98]. Protein [SSRvdB98, SR97, SBD^{+}99, SDC^{+}97, SOJM99, SGJP99, SHL97, SJZW97, SV98, SB99c, SBR98, SAEV^{+}98, SSH^{+}98, TKH^{+}97, TGGB97, TGW99, TKBM97, TLF^{+}97, TWP98, TDR99, TFS^{+}99, TZL^{+}98, TBP^{+}99, ÜHK^{+}99, VMH97, VCL^{+}98, VCS^{+}99, VWG^{+}97, WVM97, WGF^{+}98, WRC^{+}99, WZF97, WCW98, WKL^{+}99, WLF99, WE98, WNB97, WC97a, WBA^{+}98, XWK^{+}97, qXLCH97, YSEI^{+}99, YNM^{+}99, YDFB98, YAC97, YCR^{+}98, YHO^{+}98, YPDM98, YMD99, YPM97, YPB^{+}98, ZCE^{+}98, ZTH^{+}97, ZKS^{+}99, ZSB99, dBCKS98, vDSH^{+}97, vHMV^{+}97, CNP^{+}95, GHJJ95, JMTCF96, OGS^{+}96]. protein-1 [YtDH^{+}95, DKN^{+}97, NMK^{+}97]. protein-2 [MTL^{+}96, DKN^{+}97, vDSH^{+}97]. Protein-280 [CLF^{+}97b]. protein-5 [JSR96]. protein-6 [BSK96]. protein-coupled [RRH^{+}95]. protein-glucocorticoid [CRLM96]. Protein-specific [WLD^{+}98]. protein/plasma [HSK^{+}96]. protein/SL1 [JMTCF96]. proteinases [OCS^{+}95]. Proteins [BOS^{+}97, CDH^{+}98, LLH^{+}99, MEFH97, MMD^{+}98, OZF^{+}99, RP97, SJW97, dCBR99, VS98, YHS^{+}99, YTT99, BLL^{+}96, Bor95, BM96, CSO^{+}95, CRMB95, CS96a, CCR96, DTB96, DMP^{+}96, DM95, EKK^{+}96, FB96, FUB^{+}96, GAW^{+}96, GH95, GLF^{+}96, GHF^{+}95, HER95, HFSHK96, HDK96, HGP^{+}95, KAP95, KKFN^{+}95, LVD96, LNR^{+}96, LO96, MKLS96, MCL^{+}96, MFM^{+}95, MMATe96, MKF96, MR95, NDBR95, ND96, NSR95, NCS95, PWG^{+}95, PBT^{+}95, PHB96, SKE^{+}95, SSL^{+}95, SRH^{+}96, SSB96, Val96, VXBH95, WPMW95, vWDG^{+}95, Bar97, BVKP98, BF97b, BHSAJ98, BGN^{+}97, CMS^{+}97, CSY98, CES^{+}98, CH97, DAF^{+}97, DCJ97, DGP99, EAM^{+}97, FFB^{+}97, FFH^{+}98, GDC97, GHW99, GBS99, GDB^{+}97, GZC^{+}99, GBFL^{+}98, HPKL97, HCW99, HVL^{+}99, HT98, HSB98, HCG^{+}98, HFP^{+}98b, IMS^{+}98, JE98, JWK98, KQB^{+}98, KBD99, KGL^{+}99, KHS^{+}97, KEP^{+}99, LBBP98]. Proteins [LMC^{+}98, LG99, MBW^{+}99, MCBE98, MMR99a, MCC^{+}98, NTH98, OPB98, PWD^{+}99, PCQH98, PKCS98, SSI^{+}97, SC97, SEK^{+}97, SO97, SO98, TH99, TSK^{+}97, TSH^{+}98, TFYH99, TBD^{+}99, WS97, WG99, WSH98, XCLM98, YGG97a, YHD^{+}98, ZR98]. Proteoglycan [HYK^{+}98, JBD^{+}95, WC97b, BEZ^{+}95, MVvdBD96, TIJ^{+}95, WYY95, EY99, FSA^{+}99, WSH98]. Proteoglycans [GCM^{+}97]. Proteolipid [GSL98, PMBM^{+}99]. proteolysis [RPW96, SRR96b, CMD^{+}98, LZK^{+}97, TNFK^{+}98, VC97]. Proteolytic [ALM^{+}97a, MGP97, HKC^{+}95, BLRS97, FBM99, ZS98]. Prothrombin [BP98]. protochlorophyllide [RRRA95]. protocol [HP95]. Protofilaments [ES96]. proton [vWDG^{+}95]. Prototypic [WFS97]. Protozoan [KQB^{+}98]. Protracted [YKO97]. protrusion [AWS96, NH99]. protrusions [RKD96, RTM99]. Provide [Bri99]. provides [LMMO96, FH97, SZNS98, SWR^{+}98]. Provokes [SW98]. Proximal [ZB98, OK98]. PRR [TNM^{+}99]. psbA [BM98, YCR^{+}98]. PSD [KNH^{+}97]. PSD-95 [KNH^{+}97]. PSD-95/SAP90 [KNH^{+}97]. Pseudomonas [FFZ^{+}95]. Pseudopodium [SWS97a]. pseudopods [CRCH95, SWW^{+}95]. PSGL [SCH^{+}98]. PSGL-1 [SCH^{+}98]. psoriatic [BSK96]. PSTPIP [SDC^{+}97, LL98a]. PtdIns [GWB^{+}98, SDTE95]. PTEN [GTY98, GTP^{+}99]. PTH [SSA^{+}96]. PTH-related [SSA^{+}96]. PTHrP [SKS99]. PtK [MHC^{+}97]. Ptp [ZDK^{+}99, ALM^{+}97a, BKMN^{+}98, BGBK99, BATP99]. PTP1B [ABL98, BLE^{+}96, BALL98]. PTP1B-like [BLE^{+}96]. PTPmu [BKRT95]. PTS1 [DG96, EKK^{+}96, EB96, MLO^{+}95, GKM^{+}96].

PTS1-containing [EKK^{+}96]. PTS2 [EEHW^{+}98, PYL98]. pull [WSGPS95, GS99]. Pulverulent [WGP98]. pump [MKB^{+}95, vWDG^{+}95, JDK97]. Purified [FWL^{+}98, PdHTV99, FaAR^{+}96, LHS96, Bar97]. Purinoreceptor [FWBSO97]. purse [BMLM96]. purse-string [BMLM96]. push [WSGPS95]. pushing [KR96]. putative [SSY^{+}96, CF99, CRA^{+}98, GKP^{+}97, HKJ^{+}99, JDK97, PKBHK97, WKL^{+}99]. PV [SGJP99]. PV-1 [SGJP99]. PwA33 [BLG95]. PYK2 [cXP97]. pylori [CHH97, LHM^{+}96].

Q [FKM^{+}99]. Qm [IMW^{+}99]. Qm/Jif [IMW^{+}99]. Qm/Jif-1 [IMW^{+}99]. quail [CRG^{+}95, RRA97]. quaking [BQF^{+}95]. Quality [WG99]. Quantification [BCMK95]. Quantitative [ANS96, SLW^{+}99, ABN96]. quantity [MAG^{+}96]. Quick [KKM^{+}97]. quickly [PZ96]. quiescent [FB96].

R [KRCP99, WMC^{+}95a, KPKWW95, PBCS98, WKK^{+}95, MNG^{+}98, NHB^{+}95, VZNR98]. r-eag [PBCS98]. R-Ras [KRCP99, WMC^{+}95a]. Rab [FPWN95]. Rab11 [URU^{+}96]. Rab17 [ZSP^{+}98]. Rab3D [BBW^{+}98]. Rab5 [MBF^{+}97]. Rab6 [WJM^{+}99]. Rab7 [MBF^{+}97, PFHWN98]. Rab8 [PAV^{+}96]. Rab9 [DSP97]. rabbit [BK96, CFC^{+}95]. Rabphilin [MST^{+}96]. Rabphilin-3A [MST^{+}96]. Rac [KLS^{+}99, StKvD^{+}99, HH95b, BFHB97, BMHH97, MH97, MEFH97, MSH^{+}97, MM96, SvDtK^{+}98, TSK^{+}97, vLKvdK^{+}97]. Rac- [BFHB97, MH97]. Rac-dependent [MEFH97, MSH^{+}97]. Rac1 [EAL^{+}95, EWS96, JN98, JSN98]. RacE [GDTBD98]. Rad51 [HRR^{+}99, TMPM99]. Rad51-Recombination [HRR^{+}99]. radial [WMC^{+}95b, RHW^{+}98]. radiation [JGM96, OOB99]. radiation-induced [JGM96]. Radixin [KTD^{+}97, MEFH97, MMD^{+}98, RBB97, YTT99, HAS95, HSK^{+}96, PKQ^{+}98, YHD^{+}98]. Radixin/ [YHD^{+}98]. RAE1 [PFKvD99]. Raft [BRS99, SWO^{+}99]. Rafts [JLM99, OPH^{+}99]. Raji [LR96]. Ral [SWK^{+}99]. Ramp4 [YHS^{+}99]. Ran [KIT^{+}97, PWK^{+}96, CRLM96, CAVA96, CKS^{+}95, HTY^{+}99, HHS^{+}99, KDG98, MCB96, MWB98, MGY^{+}95, RLCM96, SG96]. Ran-unassisted [KIT^{+}97]. RAN/ [PWK^{+}96]. Ran/TC4 [CKS^{+}95, CRLM96]. RanBP1 [CAVA96, KDK^{+}99, RLCM96]. RanBPM [NMH^{+}98]. random [YvdEH^{+}95, AN99, SWS97a]. RanGAP1 [MGM98, MCB96, MWB98]. RanGTP [GDB^{+}97]. RANTES [RFVCM^{+}99, RFVCM^{+}99]. RANTES- [RFVCM^{+}99]. Ranvier [BDO^{+}97, DLB96]. Rap1 [CPL^{+}95, GLF^{+}96]. Rap46 [SHMC99]. Rapid [FJT^{+}95, vtHR97, PÖS^{+}98, PYM^{+}98, SC97, The97, GDC95, PBB^{+}96, BCAS98, CLRR99, GWB^{+}99, MSB^{+}97, SOPM^{+}97]. rapidly [BCC^{+}95, CLS96, KGHMT95, LWM^{+}95, TMM97]. RAR [RZZ96]. ras [KCDB95, EPMB^{+}99, KRCP99, KAAFZ95, MW97, NCW^{+}99, TCM96, TYK^{+}98, ÜHK^{+}99, WMC^{+}95a, YHK^{+}97, CCMG97, RME^{+}95]. Ras-like [MW97]. Ras-mediated [ÜHK^{+}99]. ras-transformed [KCDB95].

Ras/ERK [NCW^{+}99]. Ras/MAPK [TCM96]. ras1 [SMC^{+}96]. RasG [TCM^{+}97]. Rat [ULMT99, AASH^{+}96, BS95b, HH95a, KBSW95, Kun95, LCK95, MBS^{+}96b, RVD^{+}95, SRRB95, dCVCV95, BRD99, CSC^{+}98, CMS98, DCJ97, DLC98, EY99, LMM^{+}99, OTL^{+}97, RPS^{+}97, SC97, SGM^{+}97, SGJP99, WS98, YDFB98]. rat-kangaroo [AASH^{+}96]. RAT2 [HCA^{+}95]. RAT2/NUP120 [HCA^{+}95]. RAT7 [GDC95]. RAT7/NUP159 [GDC95]. Rate [CLS^{+}97, STS96, STP96, VBZ96, HVL^{+}99, HTW97, RS99]. rates [CFK95, NDBR95, ASP^{+}97, DMMS98]. Ratio [VSRP97]. ratios [SLT^{+}96]. rats [NA96]. RAW [GM99]. Ray [HST^{+}98]. Rb [BBJ98]. RB47 [YCR^{+}98]. rbA [SWR^{+}96]. rbet1 [ZWT^{+}97]. RBL [PHB96]. RBL-2H3 [PHB96]. rbSec1A [GMC^{+}95]. RCC1 [BF97a]. rDNA [OMW^{+}96, RACHV96]. re [PTBC96]. re-epithelialization [PTBC96]. reach [NCS95, PMH96]. Reaction [PJE^{+}99, BWG95, LTE^{+}96, SOH^{+}95]. reactive [HWFA96, TSL^{+}98]. Reading [RKB97]. rearrangement [CGSP95, RROA95, SAR^{+}99]. rearrangements [AAP^{+}95, DOP^{+}98, KLL^{+}97, MSH^{+}98]. Reassembling [SW99]. Reassembly [RMWW95, DFL^{+}99, ESM^{+}97]. Recapitulates [NG99]. Receptor [AWC99, BKRT95, CBM^{+}98, CHLÅ97, CMD^{+}98, CCMC97, CSF^{+}97, FIC^{+}97, JCR97, MN98, NCV^{+}98, SLPE^{+}98, ACHG95, BSN95, CDRB^{+}96, CSM^{+}96, CRLM96, CFC^{+}95, CMUW96, CCP96, CNP^{+}95, CS96b, DCW^{+}96, DNH^{+}96, DG96, EKK^{+}96, EB96, FLW^{+}95, FDH96, FFZ^{+}95, FPHH96, GJ95a, GHP^{+}96, GLS96, GZK^{+}95, GDR95, GCM^{+}95, GLP^{+}95, GKM^{+}96, GLB^{+}95, HWPC96, HSL^{+}95, JBD^{+}95, JGD^{+}96, KDvFP96, KLW^{+}96, KWCS96, LDF^{+}95a, LHDC95, LZC^{+}95, MCA96, MGS96, MTLW95, MTGY96, MMS^{+}96b, MA95, NIM95, PWG^{+}95, PPF96, RRRA95, RJP^{+}96a, RRP^{+}95, RRH^{+}95, RSJK95, RDM^{+}96, RD96b, SWB^{+}96, SK95a, SKR96, SKS^{+}95, SRR96b, SSA^{+}96, SHPP96, SM95c, SVD96, SSF^{+}95, TT96, VBG96, WGR^{+}95, WHH96, WSSM95, WNB^{+}95, WKJH^{+}95, WD95, ZL95a, ZL96, ZMG96b, dHvPL^{+}95, dPSMNSM95, ARR^{+}98, BPD99, BMP^{+}98, BNW^{+}99, BCAS98, BRS98, BBAN^{+}99, BLB^{+}98, BF97a]. Receptor [BSM^{+}99, BGH^{+}98, Car99, CBS^{+}97, CWSG99, CWMD97, CJH^{+}98, CMC^{+}98, CMS98, CWY99, DP98, DFD^{+}97, EEHW^{+}98, GKO99, GGPG99, GSB97, HPL99, HPS98, HAB^{+}99, HPD^{+}97, HZR98, HDDKH98, HR99, JLM99, KIH^{+}97, KBD99, KRGA98, KJR^{+}98, LHB^{+}99, MSW^{+}98, MMM^{+}98, MLC99, MNG^{+}98, MQR^{+}97, OT97, OBW98, PFHWN98, RHP99, RSD98, SLN^{+}97, STG^{+}99, SHH97, SHMC99, SMCE97, SJF^{+}97, STF98, SDHM99, SGGW99, SOPM^{+}97, SM97, SDDV97, SGYH97, TRA^{+}99, VLM^{+}98, VHF97, VC97, WBS^{+}98, WYL^{+}99, WDL^{+}98, WSWR99, YKO97, ZDK^{+}99, CNP^{+}95]. receptor-1 [GHP^{+}96, SVD96]. receptor-aggregating [GDR95]. receptor-alpha [RJP^{+}96a]. receptor-associated [NIM95]. receptor-binding [RRRA95]. Receptor-dependent [FIC^{+}97]. Receptor-independent [CMD^{+}98]. Receptor-induced [HPD^{+}97]. Receptor-like [MN98, CCP96]. receptor-mediated [CMUW96, DCW^{+}96, SKS^{+}95, BLB^{+}98]. receptor-oriented [LZC^{+}95]. receptor-related [CNP^{+}95, FFZ^{+}95, GLP^{+}95]. receptor/calcium [GCM^{+}95]. receptor/Ldl [CNP^{+}95]. Receptor/Retinoic [CCMC97]. Receptors [ABP^{+}98, YLB^{+}97, CQB^{+}96, DFC^{+}96, FWG^{+}96, GJ95a, GKR95, GEJ^{+}95, HMB95a, HSL^{+}95, HWFA96, LS95a, Mah96, MLJM95, MPW95, MTGY96, dPF95, SK96, SMA96, SR96, WKL^{+}96, WS95a, Wal95, WPS^{+}96, YtDH^{+}95, ZMM^{+}95, FCTS99, GWLH98, KKG^{+}98, LH97, LVN^{+}98, OMS97, PHA^{+}98, PFAF97, PFAA98, RDS^{+}99, VvZ99, XZMD97, YHMS^{+}98, YLD97]. Recognition [AARS^{+}97, ZPS96, EB96, LNR^{+}96, MKLS96, MTLW95, RSL^{+}96, ZKH^{+}95, BPD99, BD99, BP98, HN98, HFP^{+}98b, MGB^{+}98, MJBR98, SCH^{+}98]. Recombinant [HGI^{+}99, HGP^{+}95, SEC^{+}95]. Recombination [HRR^{+}99, YWMH99]. Reconstituted [GXE^{+}99, PMMF95, Bar97]. Reconstitutes [FSFT98]. Reconstitution [AMJM95, NBS^{+}98, PdHTV99, SS98, VG99, SK95b, BBL^{+}99b, LL97, MNG^{+}98, SFA^{+}98]. reconstruction [EDB^{+}96, SRR^{+}97]. Reconstructions [HST^{+}98]. recovery [MAM95, MAE^{+}97, SCB^{+}97]. recruit [RAM^{+}96, SMZ^{+}98, SWHG98]. Recruited [TNM^{+}99, PvDA^{+}98]. Recruitment [BML^{+}97, HRD98, LS95a, MRM^{+}98, MZAO98, PPP^{+}98, HS96, BKJL^{+}98, CS97b, DFL^{+}99, IKR^{+}98, KR99, MCC^{+}98, NJB^{+}98, ODB98, TTAH^{+}99, WBR97, dPCM^{+}97]. Recruits [CSMM98, FSFT98, ISN^{+}99, MVM^{+}99]. Recycling [LYS99, MAT^{+}98, SWR^{+}98, ISH95, JGD^{+}96, KI96, MLJM95, MTL^{+}96, RLC^{+}96, URU^{+}96, VOP^{+}95, WSBE^{+}95, BF97a, DSvDH^{+}98, GMS^{+}98, HTY^{+}99, IW97, PIB^{+}99, PKCS98, RD97, SDHM99, YLD97, ZSP^{+}98]. red [MWC95]. Red1 [SR97]. redirects [NWPW95]. redistributed [KGMF95, SALdP^{+}97, dPCM^{+}97]. Redistribution [ALM^{+}97a, BSAJ97, BDvdZW95, dPSMNSM95, NHA^{+}97]. redistributions [WGG96]. Redox [MDQG96, TSH^{+}98]. Redox-sensitive [MDQG96]. Reduced [DNC^{+}95, SGGS^{+}99, HTW97, wJTSSL98, ZhTA^{+}97]. reduces [WMC^{+}95b]. reducing [GL95]. reductase [BAC^{+}96, LW95, TMM^{+}95]. reductase-induced [LW95]. Reduction [MHLB96, YKO97]. redundancy [RFNF96]. Redundant [ABT^{+}98, CCMC97]. Reenter [CCMG97, TGGB97]. reentry [VOP^{+}95]. Reflect [DFL^{+}99, DLC98]. refolding [BM96, SHC^{+}97, TKBM97]. Regardless [NPJW99]. Regeneration [DBC^{+}99, GBW98, SCB^{+}97]. Region [LSB98, BL96, CGVS^{+}96, IWW95, KKAN^{+}96, MMATe96, OGS^{+}96, RZZ96, SYO95, BM98, LSH^{+}99, LHS97a, OK98, QPH^{+}98, RPV98, SALdP^{+}97, YNG98]. regionalized [KAT^{+}95]. Regions [DFD^{+}99, WDV^{+}97, BFP^{+}96, HWPC96, OP96a, TLF^{+}95, WSW95, DSM^{+}98a, FZZN98, IW97, SLN^{+}97, SdCAH^{+}97, WKL^{+}99]. regulate [CH95, HKM^{+}96, MDQG96, SLT^{+}96, TDW95, WXS96, WLR96, dPSMNSM95, BKK^{+}97, CAS97, CYSD99, DGP99, GTP^{+}99, HB99, JL98, KLS^{+}99, LH97, MBF^{+}97, PBA^{+}97, SLW^{+}99, SJ99, VMT^{+}99, VLM^{+}98, ZP98]. Regulated [ABP^{+}98, BLE^{+}96, GGL^{+}98, HMS^{+}98, KNSP97, LCN^{+}97, LKSW98, MM99a, MSH^{+}97, SvDtK^{+}98, SRH^{+}96, SMH98, SG99b, AOMB96, BAS95, BML^{+}97, CRMB95, CMM96, DHT96, ELMS98, FGP96, GZK^{+}95, GW95, HKC^{+}95, KGF^{+}97, LRM^{+}95, MPY^{+}96, MK95a, MSWL^{+}99, MLLT95, MMMT95, NSC96, RWOA97, SNCBM95, TKVC95, BBW^{+}98, BKN^{+}97, CKB^{+}98, DCR98, JB99, JN98, KWN^{+}99, KAS^{+}99, KKLA97, MSEP98, Nei98, SMTN99, WLF99, FMM^{+}98, HIW^{+}98, LMP98, RKB^{+}98, WSWR99]. Regulates [BSW^{+}99, BGBK99, FSDA97, MMD^{+}98, YTO^{+}99, All95, AZF96, BGGR^{+}96, CGSP95, EHM^{+}95, GBBSO96, HHM^{+}95, KCDB95, KCN^{+}96, KFE96, LCS96, LXK^{+}95, LDCB95, MBD^{+}96, RL96, TCM96, URU^{+}96, VBG96, WB96, XGC96, YJB^{+}95, ANT^{+}97, ALCC^{+}99, BPD99, BALL98, BOM^{+}98, CML^{+}98, CC98, FLO^{+}99, HJJJ97, HHR^{+}98, HLK^{+}98, JLID98, KJR^{+}98, LCB^{+}99, MDK^{+}99, OWAlW99, ODB98, OOB99, OK98, PWD^{+}99, PH97a, PTJ^{+}98, PRV98, RD97, SWK^{+}99, SBC99, SMM99, SLS^{+}99, TFS^{+}99, WHG^{+}97, ZSP^{+}98]. Regulating [KWD^{+}98, HWC^{+}96, MCA96, SRR96b, WPS^{+}96, GFM^{+}98, HYN97, LSH^{+}99, LL98a]. Regulation [AWC99, BMG^{+}95, BBOB96, BZBA99, BR98, CLK99, CR99, DKN^{+}97, GDL98, GLMG98, HS97, HSK^{+}96, HGW95, JCR97, KCG^{+}97, KTYT99, LZD95, LYLR98, MGS^{+}95, MTK^{+}98, MMS96a, NZ97, RTP^{+}98, RZZ96, RKR^{+}99, SSR^{+}97, SN98, SVH^{+}98, TSK^{+}97, TSK98, VLA^{+}98, Wal95, WMSM97, WD95, XGR95, YMAC^{+}98, ZC99, ZRG98, ZJB^{+}97, AHCC95, BWKH96, BLLB95, CSM^{+}96, DDWG96, GTL^{+}96, GMFN95, HG95b, KSP95, LLL^{+}96, LYR96, MHLB96, MAMJ95, NAV96, OSPJ^{+}96, PMCT96, RFNF96, SRR96a, SS95a, SMFC95, SYA95, TT96, WAMS96, XC96, YN95, YSC^{+}95, AMY^{+}98, ANS^{+}97, CF99, CEZA97, DCJ97, DDWM97, FLA99, FN98, FR98, FKO^{+}98, GSKL99, HFL97, JSN98, LCB^{+}99, MCC^{+}97, MBW^{+}99, MBBT98, PBJ^{+}98, RRK97, SNT^{+}98, SMS^{+}98, TSL^{+}98, TWP98, TDR99, TPM^{+}99, WDL^{+}98, ZSB99, ZBG99]. Regulator [MFL^{+}99, FPWN95, OHB^{+}95, RVD^{+}95, ZSMV95, ANRT99, BBJ98, HDDKH98, JMD^{+}98, PBM^{+}97, YSEI^{+}99]. Regulators [GCD97, KGA^{+}97, VWO99]. Regulatory [SOH^{+}95, ZLPL99, ISFB^{+}96, KAT^{+}95, LMMO96, SDTE95, TKM^{+}96, BATP99, JLID98, KCV^{+}98, KTYT99, LNH^{+}98, NMT97, SCEB99]. Reinforces [BRP^{+}97]. reinhardtii [BM98, GJ95b, KD97b, KPW^{+}97, PSW95, ZR98]. RelA [FWBSO97]. Related [EFML99, MMAK97, YF99, vIH98, ARM^{+}95, BOLS95, CKM95, CNP^{+}95, DEPR97, FFZ^{+}95, GKVR95, GLP^{+}95, KAP95, MWGE97, MVvdBD96, MMMW96, OFY^{+}95, SKE^{+}95, SGD98, SSA^{+}96, SBW^{+}96, SSS^{+}95, THM98, TALM96, WFBC96, WM95, BGN^{+}97, EGM^{+}99, GCA^{+}98, GCD97, GB98, HBPJ98, JB99, KFM99, MWF^{+}99, MSH^{+}99, OWM^{+}99, PB98, PYL98, SBD^{+}99, WRC^{+}99, cXP97, ZOS^{+}97, ATH^{+}97, APH^{+}97, ANT^{+}97, BRS98, KMM^{+}98, LB98, MBW^{+}99, MLB98, MVM^{+}99, MBB^{+}97, OKB^{+}98, PBCS98, RKB97, SHLD97, SKS99, SPCR97, SSRvdB98, SHPP96, WKL^{+}96, WVM97].

Relationship [BMC^{+}99, VB98, GW95, SK96, LFP98, VvdKMvD99]. Relationships [KMMC98]. relative [RSRK96]. Relatively [AN99]. Relaxation [KKS99]. Relaxed [SRR^{+}97]. Releasable [LYH^{+}99]. Release [ABM^{+}99, OSM98, CMHT96, CH95, DCW^{+}96, PPW^{+}95, SLM^{+}95, TT96, UHJ^{+}96, CRM^{+}97, CSA^{+}97, DOSN^{+}99, DLC98, EAOS^{+}98, JMD^{+}98, KDK^{+}99, MDE^{+}99, PFAA98, ZRHO99]. Released [DFD^{+}99, LF95, HSB98, SAEV^{+}98]. releasing [KJBM^{+}96]. Relief [ZLL^{+}98]. Relocalization [RME98]. Relocates [MCPB99]. Relocation [MKS^{+}97]. Remains [MVM^{+}99]. Remodeling [SBS^{+}99, LWB^{+}95, HIW^{+}98, LGZ^{+}99, PTJ^{+}98, TSB99, TBP^{+}99, XZSC98]. Remodels [VTHAA99]. Removal [WPZ^{+}96, WWO^{+}98, BTNZ^{+}95, ACH^{+}98]. Removes [SFS^{+}99]. Renal [DCR98, LTMR97, MHM^{+}97]. Rendered [CMR^{+}97]. Reorganization [WB99, ANS96, COML95, JGD^{+}96, PTBC96, PAV^{+}96, TOT^{+}96, CSO97, HHR^{+}98, MH97, ÜHK^{+}99]. Reorganizes [CLF^{+}97b]. Reorientation [GBS99]. Repair [MTW^{+}99]. Repeat [KGF^{+}97, DSVL^{+}98, FFMG95, GDC95, PJL^{+}95, STP96, TMS^{+}95, ZL95b, BAC97, IW97, TBP^{+}99, XWK^{+}97, CBB^{+}96, WKK^{+}95]. repeat-containing [GDC95]. Repeats [HPL97, STS96, SL96, ZL95a, WCW98]. repetitive [IWW95, WSA^{+}99]. Replicating [LSB98]. Replication [MB99, BBZ^{+}95, DC96a, FC95, FB96, JCPK95, LKWG96, PMMF95, TAK95, CJR^{+}97, CTB^{+}97, GHW99, HYN97, HN98, JE98, LGM98, MSD^{+}98, MTLGC99, MCC^{+}97, SMG^{+}97]. replicative [HAS^{+}96]. Replicon [JP98]. reporter [BGGR^{+}96, REdC95]. represent [SLP^{+}96, KMG^{+}97]. repression [GGC96, KBSW95, IFSV^{+}98, TRA^{+}99]. repressor [RSL^{+}96]. Reproduction [HCRS98]. Require [KBB^{+}97b, JCPK95, KSL96, PWG^{+}95, PMH96, CT98, KR99, LB98, RHW^{+}98]. Required [BDS97, CDH^{+}98, FWL^{+}98, FKNM97, GK99, gLSE^{+}99, MSH^{+}97, NAsG^{+}98, OBB99, SGTH99, SGD98, SCR98, SWVR98, THM98, WLF98, YLB^{+}97, ABR95, BLG95, BTTB96, BAS95, BYG96, BvdBvM96, CPL^{+}95, CKS^{+}95, CKP96, DRAT95, DGSR95, DG96, DYF96, ETF95, FT95, GSC95, GSB^{+}96, GAW^{+}96, HER95, HB95b, IFHR^{+}96, IMFS96, KF95, LHK^{+}96, LVD96, LS95b, MBM95, MGM95, MM95a, MR96, NW96, OP96a, PCL^{+}96, PBT^{+}95, RHF^{+}96, SH96, SKE^{+}95, SKR96, SGGS96, TNY^{+}95, VPH95, VH96, VMV^{+}95, VUC^{+}95, WW96b, WKK^{+}95, WRW^{+}95, WGG96, WHP^{+}96, XW96, YGK96b, ZDT^{+}96, ZL95c, dCVCV95, ATD97, AEVB98, ABE^{+}98, ALM97b, BMB97, BRLM98, BBR97, BHHF97, BLB^{+}98, BSY97, BMHH97, BCWA97, CRM^{+}98, CAS^{+}98, CDN97, hCMPG97, CYH98, CF99, CLL^{+}98, DAF^{+}97, DEPR97, DKLS^{+}99, DHDJ^{+}97, DSP97, DFMG97, DDRC99, EIIM^{+}98]. Required [ELMS98, ETP^{+}99, FHK97, FCTPA^{+}99, FR98, FAZ^{+}99, FdEP99, GKB^{+}98, GPSF99, GPKH99, GTC^{+}98, GB98, HSL^{+}99, HZR98, HR99, IIM^{+}99, IW97, JMS97, JDK97, JH97, JWFS99, JK97, KS98, LTMR97, LXC^{+}98, LJK97, MTLGC99, MSH^{+}98, MZL^{+}99, MCTMK98, MB97, MTW^{+}99, MR98, MMR99a, MCC^{+}98, MBB^{+}97, MS97b, MKW^{+}99, NA99, NGT99, OBW98, OWD^{+}99, PDW99, PDS^{+}97b, QGE^{+}99, RPY^{+}99, RWS^{+}99, RB98, SPL^{+}97, SGSW97, SCEB99, SCH^{+}98, SBR98, TGN97, THS^{+}98, TPMB97, TSSR98, TBD^{+}99, TCM^{+}97, VMH97, VCL^{+}98, VCS^{+}99, YWMH99, ZLM^{+}98]. Requirement [DVL^{+}98, EFK^{+}98, FBD^{+}95, vtHR97, JLRS95, LMS97, AMJM95, BMLM96, PWK^{+}96, PFF^{+}95, VLX^{+}96, YSM^{+}95, EKSC98, NBS^{+}98, RPS99, SSZ^{+}97, SEBW^{+}98]. Requirements [MBA97, CQB^{+}96, LHS96, BSB^{+}97, CWMD97, HHM^{+}98, OT97]. Requires [GHS98, SBG^{+}98, ZB98, AM95a, BAC^{+}96, BGN^{+}96, CGVS^{+}96, DPHW96, DLGB95, FDH96, GF95, HH95b, LS95a, LD95, LLN^{+}95, PSBB96, RSR^{+}96, SLSW96, SSN96, SIAS96, VLZR96, BVKP98, CIK^{+}99, GLL^{+}99, GvdFvD^{+}99, HH98, HTM^{+}98, KHA^{+}99, LBBP98, OWW^{+}99, RSS98, SS97, SMCE97, SVS97, WBA^{+}98, YCR^{+}98]. Rer1p [SSN96]. Rer1p-dependent [SSN96]. Rer1p-independent [SSN96]. Rescue [AHBW96, AK97b, PC99]. Rescue-promoting [AK97b]. rescued [LW96]. Resealed [TMM97]. resealing [BAS95]. research [MW95a]. Resemble [VTHAA99]. reside [FGA^{+}96]. Resident [VS98, KSL96, LLNH^{+}98, TH97, TFYH99, SWR^{+}98]. Residue [FSDA97, NNMW96, GF97]. Residues [SZFM98]. resistance [BWF^{+}96, DSvdV^{+}96]. resistant [GH95, CFS98, SHB99]. resolution [ACSN98, KE98]. resolved [GJ95a]. resonance [KE98]. Resorption [AMY^{+}98]. Respective [HTH^{+}97]. respiratory [AGB^{+}96, MBD^{+}96]. response [KGHMT95, KAAFZ95, Mah96, RCKS95, TDW95, VPH95, GGPG99, JWK98, RHLG98]. Responses [TMN97, CQB^{+}96, GHP^{+}96, LLZ^{+}96, WVVD95, EM98, HMB^{+}98, NTAN98, SWD^{+}98]. Responsible [RAA^{+}97, BJS^{+}95, DOSN^{+}99, GBFL^{+}98, TZL^{+}98]. responsive [BS95a, OABD^{+}96, VYB95]. restoration [VKR^{+}96, MS97a]. restores [WNB^{+}95]. Restraint [SMH98]. Restricted [SV98]. restrictin [NHB^{+}95]. Restriction [DBE98, ZB98, CFL^{+}98]. Restrictive [KEJ97]. result [GBW^{+}95, WGR^{+}95]. Resulting [MGB^{+}98, VWG^{+}97, ZLL^{+}98]. Results [BPA^{+}97, KBS^{+}99, BEZ^{+}95, CLT^{+}95, ED96, HCM^{+}96b, KKG95, LFB96, MMP^{+}95, MHLB96, PHB96, THT^{+}95, WMC^{+}95b, BGN^{+}97, GSL98, LBWF^{+}97, LMC^{+}98, SVM^{+}99, WGP98, ZOS^{+}97]. RET [TWSD98]. ret1 [SSSKR95]. ret1-1 [SSSKR95]. retained [MLJM95, CL99, KHHS97]. retardation [RJP96b]. Retention [LCP95, LHDC95, MAM95, ND96, PSBB96, SSN96, WGR^{+}95, BS97, LBBP98]. reticular [CFC^{+}95, DLC98]. reticulocyte [BK96]. Reticulum [BB97, CSN99, CGC^{+}97a, FFDP98, FNH99, GHS98, LAW^{+}98, MB99, MAS^{+}97, MZN97, NNK^{+}97, YHS^{+}99, ZWT^{+}97, All95, ABRB95a, BRB96, GEK95, GG95a, KA95, KKA96, LCP95, LHDC95, MBS96a, MBM95, MTLW95, NBW96, RAM^{+}96, RW95, SSN96, SHR^{+}95, SOR^{+}95, SNCH95, VJHR96, AFG^{+}98, ABK^{+}97, DSM^{+}98a, EGS^{+}98, GLL^{+}99, HLMD97, HTW97, sHWM^{+}99, JBRA98, LPD^{+}99, MJBR98, NZ97, RCE^{+}99, RSB^{+}99, SHC^{+}97, SPL^{+}97, SSSG98, TH97, TDR99, VKIH98, YGG97a, YPDM98, ZBW^{+}97]. Reticulum-Golgi [BB97]. Reticulum-to-Golgi [FNH99]. Reticulum/Intermediate [CGC^{+}97a]. retina [BEZ^{+}95, BOLS95]. retinal [PFF^{+}95, BFRB99, JL98, LDP^{+}99, MGB^{+}98, MAE^{+}97, ULMT99]. retinoblastoma [NMJL96, SSL^{+}95, ZOB^{+}96, SEBW^{+}98, TGGB97]. Retinoic [CBS^{+}97, CCMC97, MLLT95, MMMT95, OSK^{+}96, OABD^{+}96, XGR95]. Retinoid [CCMC97, SKA^{+}96]. Retraction [MM99a, LWKK97]. Retrieval [VS98, SSN96, WTCV95, BS97, SMCE97]. retrieve [MR95]. Retrograde [BPWS98, CES^{+}98, LR98, SME98, EKPK95, GM95, HW96a, KLW^{+}96, SNCH95, YSK95, KBF^{+}98, LLG^{+}99, LLSO98, MSW^{+}98, PWW98, PSH^{+}98, SC97, SWO^{+}99, SS98, WSS97, WJM^{+}99]. Retroviral [HPD^{+}97, RDP^{+}99]. rev107 [SEH^{+}97a]. Reveal [BMPP99, CCMG97, JWH96, HKO^{+}98, HT98, KGGK99, MSA98, SRA^{+}99]. Revealed [ACSN98, MAT^{+}98, BSD^{+}96, FTH96, MF96a, SOH^{+}95, SJHC96, ASP^{+}97, BWWT97, HSVS98, JM97b, KKM^{+}97, KGHG98, jLC97, MBDS^{+}98, PC99]. revealing [HTKH96]. Reveals [SBS^{+}99, BCC^{+}95, EPVV96, HWC^{+}96, LN96, MBF^{+}95a, OP96a, RCC^{+}96, RBL^{+}96, ST96, BAC97, BD98, BZBA99, CSY98, CHH97, DAB99, EGM^{+}99, GBS99, HHS^{+}99, KLH98, LMF98, MKC99, RKB97, RLMC99, SCPPW98, SWR^{+}98, SSM98, YMD99]. Reversal [OMS97]. Reverse [KRGA98, LNRR98]. Reverses [LEB^{+}97, RC98]. reversible [FSR^{+}96, GDC95, MDE^{+}99, RKR^{+}99]. reversibly [GG95a, SKR96]. Reversion [WPW^{+}97, DSC^{+}95]. RGD [PKR95, RD96a]. RGD-binding [PKR95]. RGD/synergy [RD96a]. rhinovirus [PPW^{+}95]. Rho [CWB96, CNG98, KBS^{+}99, MEFH97, MMD^{+}98, NH99, TXB^{+}99, ZCWB^{+}98, HH95b, LVD96, BOM^{+}98, AGP^{+}98, BLL^{+}96, BMHH97, BMLM96, CKB^{+}98, FUB^{+}96, FKO^{+}98, FOK^{+}99, GKP^{+}97, GvdFvD^{+}99, HSK^{+}96, HIW^{+}98, KFO^{+}99, MH97, MSH^{+}98, NLM^{+}98, NJB^{+}98, OWAlW99, StKvD^{+}99, SRH^{+}95, SMTN99, TSK^{+}97, WSWR99, YAiN^{+}98, vLKvdK^{+}97]. Rho- [MEFH97, SMTN99]. Rho-associated [FKO^{+}98, HIW^{+}98, YAiN^{+}98]. Rho-dependent [HSK^{+}96]. Rho-induced [MH97]. Rho-Kinase [MMD^{+}98]. Rho-Kinase-mediated [KBS^{+}99]. Rho-like [MSH^{+}98]. Rho-mediated [ZCWB^{+}98]. Rho-regulated [WSWR99]. Rho-specific [GKP^{+}97]. Rho-stimulated [CWB96, NJB^{+}98]. Rho-type [BLL^{+}96]. rho/rac [HH95b]. Rho/Rho [KBS^{+}99]. Rho/Rho-Kinase-mediated [KBS^{+}99]. Rho1 [DH99]. Rho1p [DDRC99]. RhoA [GKP^{+}97, JN98, JSN98]. RhoA-binding [GKP^{+}97]. Rhodopsin [CS98]. RhoL [MM96]. RI [ISFB^{+}96, PHB96, SHB99]. RI-mediated [PHB96]. Ribonuclease [JWW^{+}99]. ribonucleoprotein [VIF^{+}96, EJ99, JWW^{+}99]. ribophorin [SCK^{+}95]. Ribose [KSK^{+}99a]. Ribosomal [SPV^{+}99, GBMS95, WS95b, DKLS^{+}99, FR98, HHS^{+}99, NTAN98]. Ribosome [MZN97, NMHS95, VSGF96, WSW95, WSSM95, BPD99, BBAN^{+}99, BM98, GPSF99, NZ97]. Ribosome-Associated [YHS^{+}99, WSW95]. ribosome-specific [NMHS95]. Ribosome/Nascent [MZN97]. Ribosomes [EJ97]. Ribosyl [SAM^{+}99]. Ribosylation [ALK^{+}99, ODB98, RD97, KBW^{+}96, FHK97, MCS^{+}97b, DHT96, WBR97]. rich [SRRB95, YTO^{+}99, RTM99, WSSB99, CHC97, NIK^{+}97, OTQM98, SVB^{+}97, WPS^{+}98, ZOKS99, dBCKS98]. Ricin [LRS^{+}98]. Right [TYT^{+}99]. Rigid [HD99]. Rigidity [MH95a, FFB^{+}97]. rigor [NH95]. Rim [WKL^{+}99]. Ring [gLSE^{+}99, FA95, TTG96, VIF^{+}96, IFSV^{+}98, AN99, BML^{+}98, CDN97, DS98b, KSY97, LL98a, LL98b, MGIZ98, MZAO98, RC97, SAEV^{+}98]. rings [ES96, VSRP97]. RIP [HDV^{+}99]. RIP- [HDV^{+}99]. Ripped [AAM^{+}98]. RNA [BDvdZW95, CTB^{+}97, CMWL96, CCBL98, DKLS^{+}99, DCJ97, DW97, EJ97, EJ99, FC95, FR98, GBMS95, HCA^{+}95, HB95b, JCWP95, JMTCF96, KKAN^{+}96, KGL^{+}99, KDBW97, KcZS97, KBC99, MB99, MFMW95, MBBT98, NIK^{+}97, PFDL97, RPBB98, SS95a, SLEB95, WS95b, ZR98, ZZYG96]. RNA-activated [DCJ97]. RNA-binding [EJ97, NIK^{+}97, RPBB98, SLEB95, ZR98, KGL^{+}99]. RNA-editing [EJ99]. RNA1p [CKS^{+}95]. RNAs [BRJP96]. RNase [JCWP95]. Rnd1 [NLM^{+}98]. RNP [MDS95, SEK^{+}97]. RNPs [EJ97]. roadblock [BPKB^{+}99]. Robo [KMM^{+}98]. Robo-related [KMM^{+}98]. rod [BS95b, HW96b]. Rodent [TNCM97]. Role [BSW^{+}98, BDE^{+}95, CMS^{+}97, CMM^{+}95, CIRG99, DNH^{+}96, EM96, GSKL99, GSB97, GZC^{+}99, HMB^{+}98, HG95a, JWR96, JGM96, LLH^{+}99, LAW^{+}98, LFP98, MH97, MEFH97, MB99, MCS^{+}97b, PFAA98, RPV98, SPK^{+}97, SW99, SWO^{+}99, SGBD96, TH99, TTAH^{+}99, WBKB97, XC97, YTY^{+}99, vIH98, AAP^{+}95, AJS96, BAC^{+}96, BM96, CSM^{+}96, COML95, CGJ^{+}95, CMUW96, DMS^{+}95, DTB96, DM95, FBH^{+}95, FVKS96, FWG^{+}96, GBN95, HSML96, HWC^{+}96, KAAFZ95, LN96, LCM^{+}95, MAG^{+}96, MMGT96, MW95b, NYN^{+}95, PTBC96, RPC^{+}95, SMFC95, SRG^{+}96, SWW^{+}95, THK95, TK95, VMN95, VSB95, WKL^{+}96, WPS^{+}96, YGK96a, APF^{+}97, AYMB99, AZRJ98, BBP^{+}98, BZBA99, BGH^{+}98, CMD^{+}98, CMS98, CLT99, DP98, DWM^{+}98, EM98, FY99, FOK^{+}99, GDTBD98, GBvdN^{+}98, GCM^{+}97, GSJ98, GSP^{+}98, HYN97, IJR98, IJB^{+}97, KHA^{+}99, KDK^{+}99, KDO^{+}99, KvEdV^{+}96, KGHG98, LGZ^{+}99, LG99]. Role [LDD^{+}97, LLSO98, LSM^{+}98, MGM98, MGIZ98, MWB98, MP98, MBCS98, NMK^{+}97, OFM^{+}98, PPP^{+}98, PHA^{+}98, PAT99, SHPB99, SSZ^{+}97, SHB99, SCPPW98, SRA^{+}99, SSSG98, SLL^{+}98, SB99c, TIBL99, TBP^{+}99, WS97, WZF97, WMR^{+}98, WYL^{+}99, WSR^{+}97, WBR97, WOM98, XZSC98, YAD97, YPM97, ZTH^{+}97]. Roles [EWS96, GHW99, GXE^{+}99, HD96, LPD^{+}99, MCS98, WKJH^{+}95, YAiN^{+}98, FHCL95, HG95b, MTGY96, MM96, SSW^{+}96, ASP^{+}97, BD98, CJH^{+}98, CCB^{+}98, CGMH99, FKM^{+}99, HPS98, HKO^{+}98, HTH^{+}97, HDDKH98, KTAX98, SMS^{+}98, SSM98, TNFK^{+}98, vLKvdK^{+}97]. rolling [AKC^{+}95, AFFS96, AWS96, CAS96, FAP^{+}96, MPB^{+}95, PNM95, PGD^{+}97, ACP^{+}97, CS99, CES97, CFS98]. root [SGS95, ZhTA^{+}97]. rost [PGD^{+}97]. Rotation [MRL^{+}99]. rough [SNCH95]. Route [PBS97, SBS^{+}99, vIZKH97, NBW96]. routes [KDvFP96, YKRS96]. routing [dHvPL^{+}95]. Rpa [DTMG99]. RPTP [GZK^{+}95, MN98, ZMG96b]. rRNA [DKLS^{+}99, JMTCF96]. rRNAs [VLAHV98]. Ruffles [BH98, BKJL^{+}98].

Ruffling [CLK99, MSH^{+}97, ASY96]. Rvs [BDO^{+}97]. Rvs161p [BGR98]. Ry [MNG^{+}98]. ryanodine [CFC^{+}95, GCM^{+}95, MCA96, FCTS99, MNG^{+}98, PFAF97, PFAA98, STG^{+}99].

s [BKEC96, YGK96a, DK96, DBLD^{+}99, FK95, GHJJ95, HYN97, JP98, LD95, MF96a, MC95, NWPW95, NMJL96, SLS^{+}96, SIUW98, YGG97b, ZJB^{+}97]. S-acylated [SLS^{+}96]. s-laminin [BKEC96]. S-mediated [LD95]. S/G2 [FK95]. S180 [BEG^{+}95]. S332 [MPT^{+}98]. S3a [NTAN98]. Sac1p [MBM95]. Saccharomyces [CRGW98, DPH95, HCG^{+}98, SKE^{+}95, AK99, BLC95, BML^{+}98, BGR98, CP95, CSCM97, CCBK96, CS97b, CH97, CGMH99, DRAT95, DAF^{+}97, DEPR97, EKK^{+}96, EB95, FGP96, FY99, FSHD96, GLF^{+}96, HCA^{+}95, HCW96, HB99, HRA^{+}98, HKKH99, IFHR^{+}96, JBRA98, JCC^{+}98, KKG95, hKsKC99, LKE^{+}99, LMB^{+}95, LSKB95, LS95b, MCM^{+}99, MR96, MMMW96, OAB^{+}98, PBT^{+}95, PH97a, PH98, PCYS95, RRK97, RSB^{+}99, SHLD97, SFD^{+}96, SCR98, SRG^{+}96, SCK^{+}95, SS95b, SGGS96, SD98, dCBR95, SKK95, SVT96, THP96, TN95, TIBL99, TDR99, VS98, WW96a, WJH^{+}98, WMO^{+}95, XW96, YGK96b, YAD97, YSC^{+}95, ZL95b]. sad1 [HY95]. Salivary [CHC97]. Salmonella [MHK^{+}95, dPF95]. salt [BHDW95, MZAO98]. salt-induced [BHDW95]. Salt-stripped [MZAO98]. Same [FCTS99]. SAP90 [KNH^{+}97]. SAPK [TMN97]. SAPK2 [HHR^{+}98]. SAPK2/ [HHR^{+}98]. Saponin [KMK^{+}98]. Saponin-permeabilized [KMK^{+}98]. Sarcoglycan [DSM^{+}98b, HLJ^{+}98, mCBLK98, CLH^{+}99]. sarcolemmal [CFC^{+}95, SRCC97]. sarcoma [BEG^{+}95]. Sarcomere [MAM99]. sarcomeres [CRG^{+}95]. Sarcomeric [BGH^{+}97, ML97, OGS^{+}96, GTC^{+}98, ZTH^{+}97]. sarcoplasmic [CFC^{+}95, DLC98, ZBW^{+}97]. Sarcospan [CLH^{+}99]. SATB1 [dBCKS98]. satellite [SHS96]. Satellites [KSYK^{+}99]. saturable [MWOB96]. saved [DVD^{+}96]. SC-35 [SMW^{+}99, XJM^{+}95]. Scaffold [KFM99]. Scaffolding [WKL^{+}99]. Scaffolds [MZAO98]. Scale [FPRL97, TSB99, CH98, RSL^{+}96]. scar [FJT^{+}95, BRS98]. Scatter [AGS^{+}97, BFS^{+}95, TIJ^{+}95, WKJH^{+}95]. scattered [ICW^{+}96]. scattering [RVD^{+}95, GvdFvD^{+}99, LPM^{+}99, SWR^{+}98]. Schizosaccharomyces [DS98b, FFDP98, KSY97, LW95, MBP^{+}95]. Schwann [BGBS98, CKM95, EHM^{+}95, FSIG^{+}95, GBW98, HSL^{+}99, SRR96a, VGVF97]. Sciara [dS98a]. SCIP [GBW98]. Scj1p [SHR^{+}95, SSSG98]. SCP3 [YPB^{+}98]. scrapie [TSS^{+}95a]. Screen [RST^{+}99, GAW^{+}96, WMXE96]. scruin [WSG^{+}95]. scully [TOSF^{+}98]. SDF [SOPM^{+}97]. SDF-1 [SOPM^{+}97]. SDYQRL [JGD^{+}96]. sea [CCP96, MMS^{+}96b, STRM95, WTCV95, WWP^{+}96, HCRS98, LSM^{+}98, MS97b]. seam [SM95b]. Searching [SYM^{+}97]. Sec [MSMC99, HCM^{+}97, WRC^{+}99]. Sec12p [SSN96]. SEC13Rp [SOR^{+}95]. Sec15 [TN95]. SEC16 [EGH^{+}95]. Sec16p [GEK95]. Sec18p [MW97, XMMW97]. Sec1p [CGM^{+}99]. Sec23p [EGH^{+}95]. SEC24 [RCE^{+}99]. Sec2p [WSCN97]. Sec34p [VCS^{+}99]. Sec35p [VCL^{+}98, VCS^{+}99]. Sec4p [WSCN97]. Sec6 [TN95]. Sec61 [KRH98].

Sec61p [IMFS96, NMHS95]. Sec63p [CS97b, LS95b]. Sec8 [TN95]. Sec9 [LRAB99]. second [SKWY95]. secondary [FWHW96, MSL^{+}98a]. Secretase [ALC^{+}99, CGC^{+}97a]. secreted [HKS^{+}96, TYSP^{+}96]. secretion [AZF96, CGVS^{+}96, DGSR95, MGM95, BBW^{+}98, CGM^{+}99, CFB^{+}98, GAZ^{+}98, MSWL^{+}99, SG99b, VKIH98, ZCOP99]. Secreton [BGL^{+}99]. Secretory [LR98, NAsG^{+}98, AM95b, AS96, CMM96, CMHT96, DHT96, HB95a, JRLS95, NSC96, BBAN^{+}99, BG98, CSA^{+}97, DSG98, DBOM98, GLL^{+}99, HME^{+}98, KKG^{+}98, KGHG98, KKLA97, Nei98, PSB98, RRK97, SHPB99, SGM^{+}97, TPMB97, UPT98]. SecY [RB98]. SED4 [GEK95]. Sed5p [vMNS97]. Seems [BKN^{+}97]. Seen [ABM^{+}99]. Segment [iMOeM^{+}97, MVM^{+}99]. Segments [BDO^{+}97, ZB98, DLB96, TGL^{+}98, ZLM^{+}98]. segregated [MTL^{+}96]. segregating [WM96]. Segregation [MMN^{+}99, vIZKH97, KSP95, RSP^{+}95, SKE^{+}95, WGG96, ETP^{+}99, GCD97, HBS^{+}97, hKsKC99, MHWW98, MBB^{+}97, MMR99b, YAiN^{+}98]. select [DPH95]. Selectin [CFS98, ZLB^{+}97, GRB^{+}97, MPB^{+}95, PNM95, SCM^{+}96, BHC99, CS99, GCM^{+}97, HMB^{+}98, LVN^{+}98, SSEM98, SCH^{+}98, ZLB^{+}97, AFFS96, ACP^{+}97, BBL^{+}99b, FAP^{+}96, GGD^{+}97, KCN^{+}96, NSC96, PWG^{+}95, PFGS95, SCM^{+}96, YWW^{+}96]. Selectin-mediated [ACP^{+}97]. selectins [AFFS96, LKKL97]. selection [CP95, FFZ^{+}95, AWB^{+}98, YAD97]. Selective [HFSHK96, vtHR97, KSE^{+}97, NCW^{+}99, SIUW98, VOP^{+}95, EELF^{+}95, MKB^{+}95, ATD97, CNG98, MSB^{+}97, RCE^{+}99, TPMB97, GE97, SSS99a]. Selectively [DSVL^{+}98, GL95, MLJM95, MWL^{+}96, NH95, WS95a, AGP^{+}98, FWBSO97, SCB^{+}97]. self [SJ95, BGG98, HTH^{+}97, TWP98]. self-assemble [SJ95]. Self-organization [BGG98]. Semaphorin [MSF^{+}99a]. Seminiferous [DCJ97, TPM^{+}99]. Semliki [BWG95, KKGD96, VPM^{+}98]. senescent [GHP^{+}96]. sensing [YJB^{+}95, AMY^{+}98]. sensitive [CT97a, DBvdBS95, GHF^{+}95, HLSM95, LCD^{+}98, LGdC95, MDQG96, MTL^{+}96, NWG95, DMY^{+}97, LLG^{+}99, TSH^{+}98, DSG^{+}97, KGE^{+}98]. Sensitivity [DCVF98, MBRN95, SFD^{+}96, SH95, BR98, CBTZ98, HTW97, MDC^{+}99]. Sensor [JMD^{+}98]. Sensory [CDH^{+}98, APG^{+}96, HSL^{+}99, HGG^{+}97]. Separable [TKM^{+}96, AW96, OWA95]. Separate [REdC95, HUU96, OP96a, BS97, CCBL98, MPT^{+}98]. Separation [ODJ99, GACM^{+}95, GPKH99, MTLGC99]. Septate [WLF98, EZC^{+}97]. Septate-like [EZC^{+}97]. Septin [FWL^{+}98, FGP96, FaAR^{+}96, DAF^{+}97, LL98a, LFP98]. Septin-based [DAF^{+}97]. Septins [CK96, FWL^{+}98, FGP96, CAS^{+}98, JB99]. Sequence [CWSL97, CAA95, LHK^{+}96, WSG^{+}95, WMR^{+}96, AASH^{+}96, AM95a, CGJ^{+}95, DMG96, LDF95b, MKLS96, MGS96, ND96, NNMW96, PL96, THP96, TSS^{+}95a, VJHR96, ZL96, CJR^{+}97, CWMD97, FFH^{+}98, MJBR98, RSD98, dBCKS98]. Sequence-binding [dBCKS98]. sequence-dependent [VJHR96]. Sequences [KAP95, HBJ95, KAT^{+}95, NBW96, RSL^{+}96, ZPS96, HVL^{+}99, WSA^{+}99, dBCKS98]. Sequential [ABRB95a, LL98b, MBF^{+}97, WAMS96, YSM^{+}95, FGA^{+}96, KA95, HPKL97, NTAN98, VC97]. Sequestered [DFD^{+}99]. sequestering [GDL98]. sequesters [LTE^{+}96]. Sequestration [HRR^{+}99, KSL^{+}95, PW98]. Ser [FIC^{+}97]. SERCA2 [JLC98]. Serine [KND^{+}97, MCC^{+}98, YTY^{+}99, YTO^{+}99, ZC99, OCS^{+}95, TKVC95, BCWA97, MOY^{+}98]. Serine/Arginine [YTO^{+}99]. Serine/Arginine-rich [YTO^{+}99]. Serine/Threonine [KND^{+}97, ZC99]. serotype [PPW^{+}95]. Serp1 [YHS^{+}99]. Sertoli [MSF^{+}99b]. Serum [KGF^{+}97]. Serum- [KGF^{+}97]. serves [SFNRH95, KLM^{+}98, YHK^{+}97]. SET [KKFN^{+}95]. several [FHV^{+}96, MFMW95]. severe [PLM^{+}96]. Severing [AYMB99]. Sex [KGS98]. Sex-limited [KGS98]. SFT [GYRWR97]. SH2 [HIWL97]. SH3 [ABE^{+}98, BLL^{+}96, EKK^{+}96, GKM^{+}96, KSK^{+}99b, SRRB95, WKRB96]. SH3P12 [MNS^{+}99]. SH3P9 [BDO^{+}97]. Shaker [SG97]. Shape [KBS^{+}99, CMHT96, EDB^{+}96, GKC^{+}96, Ord99, SWK^{+}99, SJ99]. Shaping [KTAX98]. share [AM95b, PSH^{+}98]. shared [JWH96, CRGW98]. Shares [WKL^{+}99]. Shc [GTP^{+}99, CRYC99]. She4p-like [BBOE98]. shear [AFFS96, CS99, LKKL97]. Sheath [BGBS98]. Sheaths [MSF^{+}99b]. shedding [GSvR^{+}96]. sheets [CFK95]. shift [TNY^{+}95, HCLG99]. Shifters [Ord99]. shifts [TNY^{+}95]. Shiga [MAT^{+}98]. Shigella [AAP^{+}95, BMG^{+}99a, BGL^{+}99, DOP^{+}98, ELL^{+}99, LZK^{+}97, ZPS96]. Shock [BSAJ97, BM96, SFD^{+}96, SRG^{+}96, SSH^{+}98]. short [DMG96, TCSG96, SWS97a]. show [MAG^{+}96, HD99, SRCC97]. Showing [MFF97]. Shows [RSS98]. Shr3p [KSL96]. Shuttling [KDG98, NB97, PFKvD99]. Sialic [TSD^{+}97]. Sialomucin [PFGS95]. Sialyl [FHY99]. Sialylated [FAP^{+}96]. Sialyltransferase [WBG^{+}97]. sickled [SMB^{+}95]. Sid2p [SMM99]. Side [DSM^{+}98a]. sidearms [SVB96]. sigma [DSM^{+}96]. Signal [CSM^{+}96, HDV^{+}99, MJBR98, NBW96, VJHR96, BTTB96, BGR^{+}95, GHP^{+}96, GG95a, GK96, HG95a, MRvD^{+}95, MLJM95, MTLW95, MXSRB96, RSC96, RLCM96, RCKS95, SPMB^{+}96, TKVC95, THP96, WNB^{+}95, YJB^{+}95, BPD99, BAH^{+}98, BHC99, BF97a, CS98, GRS^{+}99, IW97, IJB^{+}97, KRH98, KBC99, MGB^{+}98, PHN^{+}98, SM97, SBS^{+}98, SDDV97, WS97, WJWM^{+}98]. signal-1 [WNB^{+}95]. Signal-dependent [BHC99]. signal-mediated [GG95a]. Signaling [BRM^{+}99, HSML98, JGCS97, MI99, SvDtK^{+}98, WAV98, BZB^{+}96, BKP^{+}95, FFGG96, FFMG95, HSK^{+}96, HHM^{+}95, KZKS96, LZC^{+}95, LYR96, MTC^{+}95, OP96a, PLR^{+}95, PHB96, SVD96, TYSP^{+}96, TDW95, dPSMNSM95, ANT^{+}97, ABL98, BCWA97, Car99, CRYC99, EIIM^{+}98, FLO^{+}99, FRHS98, GKO99, GGPG99, GZC^{+}99, GTY98, HPL99, HKJ^{+}99, HPS98, HHM^{+}98, HFL97, JLM99, JWFS99, LTR^{+}97, LAS^{+}97, LVN^{+}98, MNR^{+}98, MFBK99, MDK^{+}99, MP98, MM97b, NG99, NBS^{+}98, OTQM98, OWD^{+}99, PHN^{+}98, RPS99, REC97, SLW^{+}99, SRA^{+}99, VWG^{+}97, WYL^{+}99, WWO^{+}98, XCLM98, SGR97, vdBCH^{+}95]. signalings [OIT^{+}95]. Signals [CLK99, SW95, BS95a, HKC^{+}95, MWOB96, MBB96, MBRN95, NDBR95, ND96, PR95, TDW95, VYB95, VXBH95, WWP^{+}96, BINRM97, BD99, BF97b, BS97, CKB^{+}98, FSM^{+}99, FKB99, GLMG98, IAS^{+}98, JMWG98, KRCP99, LHR^{+}97, MGD97, RFVCM^{+}99, VGVF97]. significant [KR96]. Significantly [TWS97]. Silencer [KEJ97]. silencing [CPL^{+}95, SP96, BHSAJ98]. silent [SJA96]. simian [LGdC95, SPMB^{+}96]. Similar [KMI^{+}97, CGM^{+}95, EDB^{+}96, LBL95, DMMS98, GS99, NMH^{+}98, PSH^{+}98]. Similarities [RFVCM^{+}99]. Similarity [FFH^{+}98]. simple [LLK^{+}95, LO96, EKS97, HPY^{+}98]. Simplex [SEH97b]. Simply [HVL^{+}99]. Simpson [CGSY^{+}99, GKS^{+}98]. Simultaneous [HFP98a, CGSP95]. Single [FPRL97, BSJJ95, DH96, SOH^{+}95, SL96, TMDD96, BWWT97, CWT^{+}98, DLC98, FHSM^{+}97, FSFT98, QPH^{+}98, REC97, SNT^{+}98, VPM^{+}98]. single-base [DH96]. Sir3 [CPL^{+}95, GLF^{+}96]. Sir3p [SP96]. Sir4 [CPL^{+}95, GLF^{+}96]. Sister [MPT^{+}98]. Sisyphus [DSGF95]. Site [BGH^{+}97, CP95, EKK^{+}96, ETF95, GDR95, IM96, LKWG96, LWB^{+}95, NHB^{+}95, PKR95, RD96a, SHG96, TGG95, TKM^{+}96, WHH96, BSW^{+}98, DKN^{+}97, GRTB97, HTW97, KMS^{+}98, LPD^{+}99, MVM^{+}99, STG^{+}99, SV98, TSD^{+}97, YHMS^{+}98, YAD97, ZCWB^{+}98]. site-dependent [RD96a]. site-specific [LKWG96, TKM^{+}96]. Sites [BSAJ97, RdPRW98, vIH98, CLT^{+}96, CCR96, DMD^{+}96, FBH^{+}95, GSC^{+}96, HS96, MBB96, MM95b, NIM95, NAP^{+}96, SJHC96, VH96, ABAAS98, ABE^{+}98, ASP99, BSM^{+}99, CJR^{+}97, CGM^{+}99, GCKC^{+}97, GSKL99, JK97, KHA^{+}99, LL97, LHS97a, MSD^{+}98, MCC^{+}98, NB97, SS97, STF98, SVS97, VvdKMvD99, WSSB99, YAiN^{+}98]. Situ [LCD^{+}98, DC96a, HAS^{+}96, SHS96, TGL^{+}98]. six [BLC95]. size [BMS95, CMHT96, HAS^{+}96, SS95a, FR98, FJ98, GHW99, GR97, GWB^{+}98, HSW^{+}98]. size-fractionation [BMS95]. sized [YvdEH^{+}95, EFML99, EFB^{+}98, SBV97]. Skeletal [BDO^{+}97, DBC^{+}99, NMJL96, CRG^{+}95, CLT^{+}95, FWHW96, GTL^{+}96, HKM^{+}96, KAT^{+}95, KWK^{+}96, PPF96, SBSG96, WHM^{+}96, YRM^{+}96, ANT^{+}97, FCTS99, LMA^{+}97, NKK^{+}99, PSPG^{+}98, PvDA^{+}98, PFAF97, PFAA98, RNK97, SGGS^{+}99, SJF^{+}97, TTAH^{+}99, ZBW^{+}97]. Skeletogenesis [CBS^{+}97]. Skeleton [WB99, DMD^{+}96, CMC^{+}98, SNT^{+}98, TSK98, TLBK98]. skin [HWC^{+}96, PTBC96, PLM^{+}96, CML^{+}98, DBH^{+}97, KRS^{+}98b, MC98, OTL^{+}97, PC98, TRA^{+}99]. SL1 [JMTCF96]. Sla2 [EGKC^{+}99]. Sla2/Huntingtin [EGKC^{+}99]. SLC [CBM^{+}98]. Sleep [GCE^{+}97]. Sleep-inducing [GCE^{+}97]. Sli15 [hKsKC99]. Slide [GCST97]. Slightly [NWM^{+}97]. Slk19p [ZKS^{+}99]. Slow [KB99b, SMB^{+}95]. slowing [MWL^{+}96]. slowly [HCAW95]. Slug [SYT97]. SM22alpha [LMMO96]. Small [BZL^{+}98, CCMG97, ZBW^{+}97, DSM^{+}96, FUB^{+}96, KBM^{+}95, LVD96, LBL95, TN95, AL97, BMHH97, JN98, JSN98, MCJK98, OWAlW99, SSH^{+}98, TSK^{+}97, vLKvdK^{+}97]. SMC [MCBE98]. Smn [CAC^{+}99]. Smooth [HTK^{+}99, HRD98, JCR97, WMP^{+}98, BGR^{+}95, CGJ^{+}95, LMMO96, RJP96b, XHUC96, vdLST^{+}96, CLRR99, HHB^{+}97, SSZ^{+}97, WTM^{+}99]. Smoothelin [vdLST^{+}96]. Smy1p [LB98]. SNAP [BMB97, CRMB95, GMC^{+}95, SLM^{+}95, WSBE^{+}95]. SNAP-25 [GMC^{+}95, SLM^{+}95, WSBE^{+}95]. SNARE [BMB97, CGM^{+}99, CHW^{+}99, CBTZ98, DBE98, IKT^{+}96, LCW^{+}98, NOR^{+}96, POR^{+}97, SLSW96, UNPW98, UvMJ^{+}99, vMNS97, LRAB99]. SNAREs [HKO^{+}98, UvMJ^{+}99, WSBE^{+}95, vMNS97]. SNF [RMY97]. snRNP [BRJP96, KGGK99, PWK^{+}96]. Snurportin [PIB^{+}99]. SOD1 [ZhTA^{+}97]. Sodium [HCPW^{+}99, PBB^{+}96, JDK97, WS98]. Sodium/Calcium [HCPW^{+}99]. SOI1 [BF97b]. Solely [MSM99]. solidissima [VSRP97]. solubility [HE96, LO96]. soluble [BK96, HDK96, KMOO95, RROA95, RW95, HAB^{+}99, HPD^{+}97, LHS97b, SZNS98]. Solutes [SBV97]. solvent [Cun95]. Somatic [SVH^{+}98, FSC^{+}97, RC98]. some [Gol95, CGSY^{+}99, CGA^{+}99]. Sonic [FK99]. Sorted [vIH98, DMM^{+}96, KKG^{+}98, WBS^{+}98]. sorter [WMXE96]. sorter-based [WMXE96]. Sorting [DBE98, GFM^{+}98, SBS^{+}99, SWC^{+}96, YLB^{+}97, vdBLCvM96, BDE^{+}95, CMHT96, CS96b, DMG96, HKC^{+}95, HH95a, MLJM95, MBRN95, MVvdBD96, OPD^{+}96, RSRK96, SKR96, SDTE95, VXBH95, WPLK95, vWDG^{+}95, BRS99, BD99, CHC97, CS98, FRT^{+}98, JR98, KGHG98, KKLA97, LMC^{+}98, LMB98, LHS97b, MWF^{+}99, MTK^{+}98, MSM99, OT97, OPB98, PMBM^{+}99, RRK97, SMCE97, SDHM99, SDDV97, TBD^{+}99, WNB97]. Sorting/Remodeling [SBS^{+}99]. Source [BMPP99]. Sox17 [KKAN^{+}96]. Sp100 [SJW97]. Spa2p [AL97]. Space [KQB^{+}98]. spacing [RSRK96]. Spans [GTC^{+}98]. Spatial [FPRL97, FN98, MSD^{+}98, VvdKMvD99, BGR^{+}95, BMS^{+}97]. spatiotemporal [GEJ^{+}95]. SPB [AK99, AK99, DK96]. Spc110p [SGBD96]. SPC4 [CCR96]. Spc42p [DK96]. SPC6 [CCR96]. SPC7 [CCR96]. Spc72p [CYH98]. Spc97p [MUS98]. Spc98p [MUS98, TCMB98]. SPD [DK96]. Special [dBCKS98]. Specialization [PCD^{+}98]. Specializations [ZGYS99]. Specialized [DFD^{+}99, BSM^{+}99, CTB^{+}97]. Species [SBS^{+}99, BK96, RHF^{+}96, SFSV95, FST99, TSL^{+}98]. species-specific [RHF^{+}96]. Specific [CCMC97, EELF^{+}95, KAFB99, LDF^{+}95a, MGB^{+}97, MAW98, MOY^{+}98, QPH^{+}98, BEG^{+}95, BLB^{+}96, BRJP96, BFS^{+}95, BSMH95, CGJ^{+}95, DRR96, FLPS95, GPRS^{+}95, KEJ97, Kun95, LKWG96, LJ95, MHLB96, MDS95, MM96, NMHS95, NKI^{+}95, OABD^{+}96, RD96a, RHF^{+}96, SSY^{+}96, SHY^{+}96, SNL96, SSB^{+}95, SSB96, TOT^{+}96, TLF^{+}95, TALM96, TKM^{+}96, VLX^{+}96, WLD^{+}98, XGR95, XJM^{+}95, YTY^{+}99, ZB98, vdLST^{+}96, BSM^{+}99, BGH^{+}98, DKLS^{+}99, GBvdN^{+}98, HW98, IRW^{+}99, JBRA98, KRCP99, LSH^{+}99, LSL99, LMW99, MCJK98, MTK^{+}98, OTS98, RFK^{+}97, SKR^{+}98, SMW^{+}99, SFS^{+}99, AEWG^{+}97, AGP^{+}98, BNW^{+}99, CAS^{+}98, CEZA97, GKP^{+}97, GKS^{+}98, KCV^{+}98, KFL^{+}98, KGS98, LWS^{+}97a, LNH^{+}98, MSL^{+}98a, NOS^{+}97, NMNL98, TOB^{+}99, TYY^{+}99, WDL^{+}98, qXLCH97, ZTH^{+}97]. specific-related [TALM96]. Specifically [MMN^{+}99, GEW^{+}95, KKFN^{+}95, LVD96, STT^{+}96, SOE^{+}95, ATH^{+}97, EJ99, HPKL97, HRK^{+}98, TDR99]. Specification [CJH^{+}98, MM97b]. specificities [MEYPR95, BBL^{+}99b]. Specificity [DBE98, FSDA97, MSMC99, NNSN97, MLO^{+}95, SH95, CMS^{+}97, sHWM^{+}99, SSSS97]. specifies [SHR^{+}95, XML^{+}96]. specify [NBW96, SB99a]. Speckles [WSSB99]. Spectrin [WS98, ZT99, DLGB95, DSM^{+}96, DWS^{+}95, HTKH96, SMB^{+}95, FTB97, KsYA98, MLB98, XWK^{+}97, HMB95b]. spectrin-actin-binding [DWS^{+}95]. Spectrin-like [XWK^{+}97]. spectrin-protein [SMB^{+}95]. Spectrometry [SBS^{+}99, WJH^{+}98]. Sperm [GRB^{+}97, ISR99, AZF96, CCP96, MMS^{+}96b, SOH^{+}95, SFSV95, WS95a, WSG^{+}95, RSS98, TGTL97, WWM^{+}98, YPM97]. Spermatids [hYGYF97]. spermatocytes [ZN95a]. spermatogenesis [KKAN^{+}96]. Spermatozoa [SBGR98]. Spermiogenesis [KND^{+}97]. spherical [FC95]. Spheroids [KA98]. Sphingolipid [ZH97, BvdBvM96, KBF^{+}98]. Sphingolipids [SFH96, BvdBvM96, vIH98]. Sphingomyelin [FNH99, vIZKH97]. Sphingomyelinase [ZPL^{+}98]. Sphingosine [BGR^{+}95, OKE^{+}99, OKE^{+}99, VLM^{+}98]. Sphingosine- [BGR^{+}95]. Sphingosine-1-Phosphate [OKE^{+}99, VLM^{+}98]. Spikes [IMK99]. Spinal [CAC^{+}99, KTAX98]. Spindle [BBL^{+}99a, BGG98, BMG^{+}99b, CSMM98, EK97, HTH^{+}97, KWD^{+}98, SCR98, WH97, WJH^{+}98, YSC^{+}95, dS98a, DK96, EPVV96, FSHD96, GSB^{+}96, GMFN95, HY95, HM95, MAM95, MMGT96, PBT^{+}95, SdCAH^{+}97, SCG^{+}95, SGGS96, SGBD96, WFBC96, WW96a, WM96, WRW^{+}95, WGG96, WMO^{+}95, ZKSB96, ZN95a, ZN95b, AK99, AKLME98, AK97b, BHJ98, CS97a, CYH98, CRGW98, CH97, CGMH99, DMMS98, FFLP99, GDC97, GCD97, HCG^{+}98, HKKH99, KCM^{+}97, LKE^{+}99, MCTMK98, MB97, MC97, MSH^{+}99, MUS98, PTVD99, SdCAH^{+}97, SHLD97, SGSW97, SYM^{+}97, SCPPW98, SMM99, Ste97, TGN97, TIBL99, WVM97, WZF97, WHG^{+}97, WBA^{+}98, WHB^{+}99, YAC97]. spindles [CDGO99, Sch96, DMMS98, SMB^{+}99, ZKS^{+}99]. Spines [EY99, MLB98]. spiralis [FGA^{+}96]. Spirogyra [SG95]. Spisula [VSRP97]. splice [KBM^{+}95, SYO95, GBvdN^{+}98, GF97, LSM^{+}98, OTS98]. splice-prone [SYO95]. Spliced [BRJP96, AK95, BMG^{+}95, JBD^{+}95, LBCB96, PJL^{+}95, BGH^{+}98, FPF^{+}99, iMOeM^{+}97, ZBW^{+}97]. Splicing [KDBW97, WLD^{+}98, YTO^{+}99, BJB^{+}96, HS96, NSR95, XJM^{+}95, ZL95a, BRP^{+}97, CMS^{+}97, DW97, GCKC^{+}97, JVRNM99, TFS^{+}99, WSSB99]. Spnr [SLEB95]. Sponge [WBSNV97]. Sponge-like [WBSNV97]. Spontaneous [CSF^{+}97]. spore [FGP96]. Spread [RBM^{+}99]. spreading [AC95a, CM95b, DDWG96, LR96, MBD^{+}96, PTJ^{+}98, WSWR99]. spreads [GBMS95]. Spring [QPH^{+}98, LRC^{+}99]. Spring-loaded [QPH^{+}98]. Sprouting [PURB98]. squamous [ICW^{+}96, LYC^{+}95]. squid [MGRS96]. SR [WLD^{+}98, CMS^{+}97, KBD99, MCC^{+}98, NSR95]. src [GLS96, MMATe96, RVD^{+}95, TNY^{+}95, ABE^{+}98, CCS^{+}98, CRG^{+}95, CGSP95, DOP^{+}98, FCMN97, FUB^{+}96, HR99, LPM^{+}98, LFB96, PHA^{+}98, PRV98, YTY^{+}99]. Src-mediated [DOP^{+}98]. src/Microtubule [AARS^{+}97]. SRP [OBW98]. SRPK2 [WLD^{+}98]. SRY [PGC^{+}95]. Ssa1 [BM96]. Ssa2 [BM96]. Stability [SGTH99, DG96, GL95, VLX^{+}96, HVL^{+}99]. Stabilization [CCBL98, OMS97, HFSHK96, SWW^{+}95, CNG98, CLH^{+}99, GBFL^{+}98, KRS98a, RM97, SHC^{+}97, SNK^{+}97, TWP98]. Stabilize [ZKS^{+}99]. stabilized [FSIG^{+}95]. Stabilizes [YHS^{+}99, CAVA96, CS99]. stabilizing [MH95a]. Stable [BPA^{+}97, GG95b, ABN96, CLR^{+}96, GMC^{+}95, JP98, LGM^{+}97, MPSC98, PCD^{+}98]. Stably [FSC^{+}97]. Stack [DFD^{+}99, AMJM95, MR95]. Stacking [SW99]. stacks [AMJM95, MW95b, RMWW95]. Stage [KTD^{+}97, MZN97, NMHS95, GOT96, MKI^{+}96, CLF^{+}97a, DDM98, GPKH99, GSK^{+}99, LNH^{+}98, RLMC99]. Stage- [NMHS95, LNH^{+}98]. stages [MW96, RHHRB96, LTMR97, NZ97]. stain [VE95, BWWT97]. Staining [LSB98, NKI^{+}95]. Stalk [YLB^{+}97]. stardust [MW96]. starfish [OSH^{+}96, SOH^{+}95, CRM^{+}97]. Starvation [BSH99]. Stat5 [JWFS99]. State [HFP98a, CRG^{+}95, TNY^{+}95, TT96, TEL^{+}95, GF97, IIM^{+}99, TLF^{+}97, TWS97, WTM^{+}99, YKO97, LDH^{+}98]. States [CAH^{+}99, LRG^{+}98]. Stathmin [LSH^{+}99, DPC^{+}96]. status [DVD^{+}96]. staurosporine [JWR96]. staurosporine-induced [JWR96]. Ste24p [FKNM97, TNFK^{+}98]. Ste6 [EM96]. steady [TT96, GF97, LDH^{+}98]. steady-state [TT96, LDH^{+}98]. Steering [SEBF98]. Stem [BMPP99, BHD^{+}97, FPM^{+}95, OSK^{+}96, HKJ^{+}99, PGP^{+}99, PHM^{+}98, SFD^{+}98]. Step [FKNM97, GACM^{+}95, MGY^{+}95, TMDD96, DEPR97, DKLS^{+}99, HN98, KDK^{+}99, RDP^{+}99, SYT97]. Steps [BML^{+}97, CSY^{+}95, JRLS95, MF96a, SKR^{+}98, VWO99, WNB97]. Stepwise [SFA^{+}98]. stereochemical [GJ95a]. Steric [LHDC95]. sterigmata [FT95]. Stimulate [KJR^{+}98]. Stimulated [NCW^{+}99, CWB96, DPC^{+}96, EKPK95, MBD^{+}96, PGS95, STP96, BL99, DSM^{+}97, EBR96, NJB^{+}98, PHN^{+}98]. stimulates [ASY96, HTS^{+}95, MBS95, RKD96, CSA^{+}97, CS97b, HOS^{+}97, OSM98, TWSD98, WLF99]. stimulating [SRR96a, PHN^{+}98, WDL^{+}98]. Stimulation [BMC^{+}99, BMB97, CMS98, GHJJ95, MLW^{+}95, SLPE^{+}98, SVD96, SR96, WSG^{+}98, BSK96, CGSP95, LD95, BKJL^{+}98, MAE^{+}97, XZMD97]. Stimulator [GYRWR97]. Stimulatory [SZNS98]. Stimulus [LDH^{+}98]. Stimulus-induced [LDH^{+}98]. Stoichiometric [VSRP97]. stomatitis [MXSRB96, RLVM98]. stone [PGD^{+}97]. STOP [GBFL^{+}98]. storage [DMS^{+}95, KKA96, CSEM99, EvdESvdH^{+}97, LYH^{+}99, PvDA^{+}98]. Store [LCD^{+}98, EKF^{+}97, HFL97, ZOS^{+}97]. Store-operated [HFL97]. Stores [HFP98a, BCC^{+}95, MP98]. Stra8 [OABD^{+}96]. Strains [KD97b]. strand [DH96, RBRB99]. Stranded [DCJ97, CTB^{+}97, EJ97]. Strands [FSFT98, FST99, SFS^{+}99]. strategy [MHLB96, SCPPW98]. stratified [LYC^{+}95, HPY^{+}98]. Strength [LMA^{+}97]. strengthening [DWS^{+}95, YNG98]. strengthens [SP96]. Streptomyces [MHM99]. Stress [FFDP98, MFL^{+}99, YHS^{+}99, CWB96, SHY^{+}96, DH99, HHR^{+}98, KGL^{+}99, NJB^{+}98]. Stress-Associated [YHS^{+}99]. Stress-induced [HHR^{+}98]. stressed [VTG^{+}96]. Stresses [ODJ99]. striated [BZB^{+}96, HW96b, MGA^{+}96, GCML98, YPB^{+}98]. string [BMLM96]. stripped [MZAO98]. stroma [CAS96]. Stromal [RL99, CLR^{+}96, OK96, MNM^{+}98, OTL^{+}97]. Stromelysin [LGM^{+}97, MLN^{+}98]. Stromelysin-1 [LGM^{+}97]. Stromelysin-3 [MLN^{+}98]. strong [TNY^{+}95, IIM^{+}99]. Strongly [NWM^{+}97]. Structural [HHM^{+}98, HCAW95, JSN98, KKM^{+}97, KLL^{+}97, LMF98, MDS95, OT97, ADP^{+}95, BHDW95, DMS^{+}95, DWS^{+}95, HGP^{+}95, KSP95, KAP95, MBF^{+}95b, PKR95, SW95, THT^{+}95, UPAS95, XJM^{+}95, GDC97, LWS99, MCBE98, RGS99]. Structurally [TBD^{+}99, VvZ99, cXP97]. Structure [BGH^{+}97, CTB^{+}97, CFK95, LvSC^{+}96, MSP97, SJSJ95, WBSNV97, WC97b, CMWL96, EDB^{+}96, HKS^{+}96, LHM^{+}96, MC95, OGS^{+}96, OP96a, PLZ^{+}95, PFF^{+}95, RCC^{+}96, RHF^{+}96, RFNF96, SBSG96, SFSV95, SR95, VSGF96, WHP^{+}96, XHUC96, ACF^{+}97, ASACF98, ALM97b, BKM^{+}98, FKM^{+}99, GWLH98, HSVS98, HST^{+}98, HMCL97, ISN^{+}99, JP98, JBRA98, JM97b, JOPM98, LMM^{+}99, LRC^{+}99, LL98b, MCBE98, MCS^{+}97b, NMK^{+}97, OAB^{+}98, RSB^{+}99, SGA97, SJZW97, SEB99, VL99, WJWM^{+}98]. structure-function [OP96a, RCC^{+}96]. structures [BZB^{+}96, DSGF95, KvEdV^{+}96, NRM^{+}96, SLS^{+}96, SJ95, SRRB95, WFBO^{+}96, CLF^{+}97b, DSM^{+}97, ISM97, LRG^{+}98, MS99, NLM^{+}98, RM97, Sal99, SMZ^{+}98]. Stu2p [CYH98, WH97]. studied [GJ95a, MMF95, HMCL97, SNT^{+}98]. Studies [HHTT95, HBJ95, LDCB95, MBS96a, CBTZ98, HT98, MNG^{+}98]. Study [HFP98a, MAT^{+}98, OP96a, LHS97a]. Sub [vIH98]. Subapical [IMS^{+}98]. Subcellular [SBS^{+}99, CGM^{+}95, DMM^{+}96, DG96, GMFN95, KSG^{+}96, LBCB96, MF96b, SWR^{+}96, ALVMM98, GF97, LMF98, TFS^{+}99, YDFB98]. Subcompartments [MAS^{+}97]. Subcomplex [CLH^{+}99]. Subcomplexes [EGM^{+}99, FHAP98]. subcutaneous [DSE^{+}95]. subepithelial [MHK^{+}95]. subjected [AGB^{+}96]. submandibular [KKD^{+}95, YHMS^{+}98]. Submembrane [SRV^{+}97]. submolecular [WMR^{+}96]. Subnuclear [YLD97, FC95, LvSC^{+}96, CMS^{+}97]. subplasmalemmal [CGVS^{+}96, EKF^{+}97]. Subsequent [YHS^{+}99, KRMMA96]. subset [BSJJ95, CBB^{+}96, GHF^{+}95, NSR95]. Subspecies [SSI^{+}97, SKY^{+}98]. Substitutions [QPH^{+}98]. Substrate [AARS^{+}97, AKP95, BGB^{+}95a, SHT^{+}95, SZB^{+}95, TK95, CMC^{+}98, FMM^{+}98, KKS99, MLB98, NWM^{+}97, SDC^{+}97, SEBF98, TYK^{+}98]. Substrate-1 [CMC^{+}98]. substrates [AGB^{+}96, GHP^{+}96, KKG95, BSS^{+}97]. substratum [DNC^{+}95, NYN^{+}95]. Substructure [YMD99]. substructures [SGGS96]. Subtle [GMS97]. Subunit [LAW^{+}98, MSP97, RdPRW98, SPV^{+}99, SBG^{+}98, XML^{+}96, BDvdZW95, EPVV96, FBD^{+}95, GHJJ95, GBG^{+}96, GLB^{+}95, HWC^{+}96, KKD^{+}95, KKG95, LHK^{+}96, LHM^{+}96, MWL^{+}96, MTLW95, OBW98, PR95, SLT^{+}96, SOE^{+}95, SFSV95, SCK^{+}95, SCKG95, SEC^{+}95, TCM96, TEL^{+}95, WGR^{+}95, WHH96, WMR^{+}96, WMC^{+}95b, WVVD95, BPD99, BF97a, BKN^{+}97, BENV97, EFB^{+}98, EFK^{+}98, EFML99, FKO^{+}98, JDK97, KRH98, KLH98, KFO^{+}99, MSW^{+}98, MAW98, MB97, PBJ^{+}98, RHW^{+}98, TYY^{+}99]. Subunits [AAM^{+}98, RP97, TB99, WDV^{+}97, BWG95, DMP^{+}96, ISFB^{+}96, PPF96, SWR^{+}96, TEL^{+}95, GAPG97, JWW^{+}99, KTAX98, RDS^{+}99, TLF^{+}97]. Successful [VWK^{+}95]. Sudden [KR99]. sufficient [BLLB95, KDvFP96, LBCB96, MKXY95, PNM95, AL97, CFS98, CCBL98, KJR^{+}98]. Suggest [SZBB98]. suggests [CSM^{+}96, MCL^{+}96, ZHR95, FZZN98, MMK^{+}97, WBR^{+}98]. Sulfate [HYK^{+}98, WC97b, MVvdBD96, WYY95, EY99, FSA^{+}99, GCM^{+}97, WSH98]. Sulfotransferases [BBL^{+}99b]. SUMO [MGM98, MWB98, ISN^{+}99, JB99]. SUMO-1 [MGM98, MWB98, SJW97, ISN^{+}99]. sup [ROG^{+}96]. sup-pf- [ROG^{+}96]. Superfamily [PPWL97, TYT^{+}99, YNOH95, GSP^{+}98, HJJJ97, MPLS^{+}98, RKB^{+}98, SEBF98, TH99]. Superior [PMKR97]. Superoxide [LEB^{+}97]. Superposition [EKF^{+}97]. Supervillin [PPWL97]. supply [AGB^{+}96]. Support [Bri99]. supports [AKC^{+}95, LK95, CES97, FHSM^{+}97, HOCK98, YNG98]. suppress [KM95, SEC^{+}95, IAS^{+}98, KAFB99]. suppressed [SFD^{+}96]. Suppresses [IMW^{+}99, GL95]. suppressible [WMC^{+}95a]. Suppression [BSW^{+}99, LDH^{+}98, MQR^{+}97, PKQ^{+}98, SGYL^{+}96, SSR^{+}98, FVKS96, LDCB95, RME^{+}95, jLC97, MGB^{+}98, NTAN98, NIK^{+}97]. Suppressor [KJE^{+}99, MF96b, NAP^{+}96, SBR^{+}96, SMC^{+}96, TSS^{+}95b, BRS98, GTY98, VWG^{+}97, SEH^{+}97a]. Supramolecular [WPS^{+}98]. Surface [MHK^{+}95, BEZ^{+}95, Cun95, DvD95, EDB^{+}96, GZK^{+}95, HB95a, HW96b, HGW95, KE98, MMP^{+}95, MXSRB96, OCS^{+}95, RKD96, RRH^{+}95, SRR96a, SPMB^{+}96, SRR96b, SSP95, SVD96, VLZR96, VOP^{+}95, WFBO^{+}96, YSK95, vdBLCvM96, ALK^{+}99, CWT^{+}98, EY99, HOS^{+}97, KMM^{+}98, KS98, MPR^{+}97, PGP^{+}99, SNT^{+}98, SWS97b, SZNS98, YKO97, YPM97]. surface-associated [SVD96]. surface-bound [OCS^{+}95]. surfaces [MBD^{+}96, GS99, NWM^{+}97]. Surrogate [GAZ^{+}98]. surrounding [BBOB96]. Survival [BRM^{+}99, JCR97, RFK^{+}97, DYF96, FVRCH96, MBF^{+}95a, NA96, VLX^{+}96, BKK^{+}97, HSL^{+}99, IAS^{+}98, LEB^{+}97, MAE^{+}97, OKE^{+}99, PMKR97, QBvD^{+}98, RLMC99, SAC^{+}98, VMT^{+}99, VLM^{+}98]. Sustained [EKSC98]. SV40 [CLT^{+}95]. SV5 [BL96]. Swelling [LWSL^{+}98]. Swelling-induced [LWSL^{+}98]. SWI [RMY97]. SWI/SNF [RMY97]. Swiss [NJB^{+}98]. switch [DCW^{+}96, KLM^{+}98]. Switches [Val96, GLMG98]. Syk [SLS^{+}99]. Symmetric [MW97]. Symmetrical [MCBE98]. sympathetic [APG^{+}96, CLS96, DVD^{+}96, MFM^{+}95, WD95, dCVCV95, BMP^{+}98, BKK^{+}97, MDE^{+}99, SC97]. Symplekin [KSK^{+}96]. synapse [TCM96, RRA97]. Synapses [HYK^{+}98, KBSW95]. Synapsin [THT^{+}95, CGB^{+}95, RLC^{+}96, TTT^{+}99]. Synapsis [BMS^{+}97]. Synaptic [BDS97, KNH^{+}97, KI96, PT97, RLC^{+}96, SHH97, CGB^{+}95, DCOK95, GYS^{+}95, GHF^{+}95, GK96, PMCT96, SJSJ95, TMDD96, UHJ^{+}96, WSBE^{+}95, CAS97, FHK97, FKM^{+}99, JMM^{+}98, JL98, NTH98, NBJ^{+}99, SFR^{+}98, TTT^{+}99, WNB97, YHO^{+}98, ZCOP99]. Synaptic-like [SHH97]. Synaptobrevin [PT97, RGS^{+}96, WNB97]. synaptobrevins [CGM^{+}95]. synaptogenesis [AHCC95, CJG^{+}95]. synaptogyrin [SJSJ95]. synaptonemal [SR95, TMPM99, YPB^{+}98]. Synaptophysin [PT97, HKL96]. Synaptopodin [MHM^{+}97]. Synaptotagmin [SEB99]. Synchronized [WPLK95]. synchronous [ASSS^{+}95, MAM95]. syncytial [MM95a]. Syndactyly [MCS98]. Syndecan [HYK^{+}98, CWM^{+}98, EY99, LR96]. Syndecan-1 [LR96]. Syndecan-2 [CWM^{+}98, EY99]. syndecans [SVD96]. syndrome [MS96, CGSY^{+}99, FRT^{+}98, FH97, GKS^{+}98, Li97]. Synergin [PSLR99]. synergistic [NYN^{+}95]. Synergistically [VMT^{+}99]. synergy [RD96a, SLT^{+}96, GSB97, HLK^{+}98]. Syntaxin [PKCS98, GMC^{+}95, WSBE^{+}95, DRE97, KKG^{+}98, LWS^{+}97a]. Syntaxin1 [BDS97]. synthase [CS96a, KHA^{+}99, LHS97a, SS97]. Synthesis [FNH99, BBZ^{+}95, GGC96, JMTCF96, KCN^{+}96, LK96, TIJ^{+}95, YN95, dCVCV95, vGvM95, FR98, FJ98, HAB^{+}99, KcZS97, KGHG98, PWD^{+}99, SSZ^{+}97, SG99a, SMG^{+}97, TDR99]. Synthesized [vtHR97, BDE^{+}95, CLS96, SHY^{+}96, vdBLCvM96, BM99, HVL^{+}99, TH97]. synthetase [STN^{+}96, DCJ97]. Synthetic [GAW^{+}96, SKA^{+}96, MBDS^{+}98, WG97]. Syntrophin [AK95, SYO95, PAF97]. System [PJE^{+}99, SPK^{+}97, SW99, AC95b, DSC^{+}95, EKPK95, MW95b, RMWW95, SZB^{+}95, THT^{+}95, UHJ^{+}96, WS96, ZL95a, BPS99, CS99, CRA^{+}98, DCJ97, LHWH97, MCC^{+}97, NG99, PMCS97, SSR^{+}97, SZBB98, SVMB97, TKH^{+}97]. systems [SSP95, HCM^{+}97, ZJB^{+}97].

T [GTC^{+}98, HOCK98, PKBHK97, UNPW98, AAP^{+}95, BDO^{+}97, CLT^{+}95, DNH^{+}96, FHCL95, HSL^{+}95, HR99, HFL97, HIWL97, JLM99, KCN^{+}96, LDF95b, LKSW98, MAB98, MGA^{+}96, PWZS97, PH97b, SPK^{+}97, SALdP^{+}97, SKC97, SLPE^{+}98, SWVR98, SBS^{+}98, VNTM^{+}99, WHM^{+}96, dBCKS98, dPCM^{+}97, DBE98, IKT^{+}96, UvMJ^{+}99, WSBE^{+}95, vMNS97]. T-1 [MAB98]. T-cap [GTC^{+}98]. T-induced [SPK^{+}97]. T-plastin [AAP^{+}95]. t-SNARE [DBE98]. T-tubules [PWZS97, WHM^{+}96]. TACs [WSGPS95]. Tadpole [SSSS97]. Tagged [GGPG99, OTQM98]. Tail [KSE^{+}97, MMD^{+}98, BLLB95, FBD^{+}95, HLSM95, MRvD^{+}95, NCZ^{+}95, RSJK95, RSRK96, SKR96, SOE^{+}95, SEC^{+}95, TGG95, CS98, HZR98, IRW^{+}99, SWS97a, YNG98]. tail-less [SEC^{+}95]. tailed [RRA97]. Tails [RAA^{+}97, MSM99]. Talin [NWM^{+}97, SSW^{+}96, KGHMT95, BH98, CLRR99, PHM^{+}98]. Talin-binding [BH98]. Talin-Null [NWM^{+}97]. Talk [LWS^{+}97b, MGD97, PH97b]. Tandem [SBS^{+}99, STP96]. TAP [GAZ^{+}98, NAsG^{+}98]. TAP-dependent [GAZ^{+}98]. TAP/p115 [NAsG^{+}98]. Tap20 [TGW99]. TAPA-1 [YMAC^{+}98]. target [LDF95b, BINRM97, TYK^{+}98, VvZ99, XCLM98, YHK^{+}97]. Targeted [ABM^{+}99, DBH^{+}97, DSE^{+}95, FH97, KMI^{+}97, RKB97, RBL^{+}96, RLMC99, WGP98, JGD^{+}96, Kun95, LWB^{+}95, ZL96, HGI^{+}99, HIWL97, SBOK97, SWS97b, SGM^{+}97]. Targeting [AAB^{+}97, BOS^{+}97, BLEB96, CSO^{+}95, CF95, CSN99, CCP96, GK99, GCKC^{+}97, HCM^{+}97, ISH95, KRS98a, LCM^{+}95, LSZ^{+}99b, MYC^{+}98, NSC96, PR95, dPF95, PL96, RP97, SS97, SHG96, WNB97, WC97a, WBKB97, YTO^{+}99, AM95a, BLG95, BAC^{+}96, GK96, HLSM95, HMSK95, HDK96, HH95a, KDvFP96, LBCB96, LHS96, LSKB95, MRvD^{+}95, MWOB96, MAMJ95, MGS96, NBW96, NNMW96, OSHG^{+}96, OIT^{+}95, PXR^{+}95, PLM^{+}96, RSRK96, SK95b, SRG^{+}96, SW95, TSS^{+}95a, VYB95, VXBH95, WNB^{+}95, ZL96, vGvM95, AKLME98, ABE^{+}98, BSH99, CC98, CLH^{+}99, ESM^{+}97, FWBSO97, GRS^{+}99, GGL^{+}98, HMB^{+}98, HT98, sHWM^{+}99, KKS99, KSOK97, LHS97a, jLC97, MWF^{+}99, MWB98, PRB99, PYL98, RLVM98, SHPB99, SVM^{+}99, SL97, SNK^{+}99, TSSR98, TYY^{+}99, WCW98, WRdVC97, KIT^{+}97]. Targeting-defective [jLC97]. Targets [NCV^{+}98, OHB^{+}95, CJR^{+}97, DSvDH^{+}98, KQB^{+}98, KKT^{+}98, PNEH98]. TATA [BBOB96, JMTCF96, PRB99, WRC^{+}99]. TATA-binding [JMTCF96, PRB99]. tatB [WRC^{+}99]. tau [BLL95, HFSHK96, EGS^{+}98]. Taxol [TALM96]. Taxol-dependent [TALM96]. TC4 [CKS^{+}95, PWK^{+}96, CRLM96]. TCR [SLPE^{+}98]. Tctex [PKBHK97]. Tctex-2 [PKBHK97]. Tctex1 [HOCK98]. Tectal [VZNR98]. tectorin [KLMR95]. Telencephalic [MHM^{+}97]. Telomerase [FC95, HSLB99, Lan97]. Telomere [HSLB99, HAS^{+}96, SWS^{+}96, BMS^{+}97]. Telomeres [BMS^{+}97, GLF^{+}96, SB99b]. telomeric [CPL^{+}95, LvSC^{+}96]. telophase [MAM95]. temperature [DBvdBS95, LGdC95, DSG^{+}97]. temperature-sensitive [DBvdBS95, LGdC95, DSG^{+}97]. Template [ALM97b]. Templates [RV98]. Temporal [DFL^{+}99, FCMN97, BGR^{+}95, MSD^{+}98, SEBW^{+}98]. temporally [AW96]. Ten [OZF^{+}99]. Ten-m [OZF^{+}99]. Ten-m/Odz [OZF^{+}99]. Tenascin [CES97, GSC^{+}96, HKS^{+}96, JCR97, VZNR98, NHB^{+}95]. Tenascin-C [GSC^{+}96, JCR97]. Tenascin-Y [HKS^{+}96]. Tendon [SV98]. Tensile [LMA^{+}97]. tensin [CLL95, LYD^{+}97]. tension [CH95, GW95, NWG95, GDTBD98, WCMS98]. Teratoma [BFL^{+}97]. term [ACHG95]. Terminal [vtHR97, AAB^{+}97, BPA^{+}97, BLL95, DGSR95, FKNM97, HAS95, HPL97, HGW95, KD97a, LDF95b, MBA97, MMD^{+}98, MSH^{+}97, NCZ^{+}95, OGS^{+}96, OWD^{+}99, SAR^{+}99, SWK^{+}99, SKC97, TNFK^{+}98, WRG98, DW97, sHWM^{+}99, JL98, KDK^{+}99, KRGA98, MAC^{+}97, NBJ^{+}99, PURB98, SHPB99, SJF^{+}97, VTHAA99, BL96, CGJ^{+}95, FVKS96, GF97, HLSM95, HTY^{+}99, KBB^{+}97a, LCM^{+}95, MAS^{+}95, PXR^{+}95, PL96, TNFK^{+}98, TSS^{+}95a, WHM^{+}96, ZL96]. Terminals [DSVL^{+}98, THT^{+}95]. Terminates [EHC97]. termini [CPL^{+}95, IFM^{+}99]. terminus [GBG^{+}96, SYO95, hCMPG97, GTC^{+}98, MSW^{+}98]. ternary [DWS^{+}95]. Territories [FPRL97, EDB^{+}96, KLN^{+}96, ABAAS98, MSB99, VvdKMvD99]. Testicular [GEJ^{+}95, DCJ97]. testis [PGC^{+}95, HFE^{+}98, KKT^{+}98, MSF^{+}99b]. tether [SK95a]. Tethering [SW99, AKC^{+}95, FAP^{+}96, WHH96, CFS98, SNT^{+}98, VCS^{+}99]. Tethers [ACP^{+}97]. Tetrahymena [CMHT96, CT97a, GCB^{+}95]. Tetraplex [BSS^{+}97]. Tetraspan [YMAC^{+}98]. Tetraspanin [BO99, CSC^{+}98, SB99c].

TFIIS [APRB98]. TGF [SKS99, BLB^{+}96, DMS^{+}95, DDWM97, ED96, FSAP95, HRD98, LLZ^{+}96, LDCB95, MBCS98, NMK^{+}97, SJF^{+}97, ZOKS99]. TGF- [SKS99, DDWM97, MBCS98, SJF^{+}97, ZOKS99, HRD98]. TGF-alpha [LDCB95]. TGF-beta [DMS^{+}95, ED96, LLZ^{+}96]. TGN [BF97b, BPWS98, DSP97, GMS^{+}98, JGD^{+}96, JH97, LH97, MCRB97]. TGN38 [GMS^{+}98, JGD^{+}96, MM99b]. tha4 [WRC^{+}99]. Thapsigargin [LYH^{+}99, PFP97]. Thapsigargin-insensitive [PFP97]. ThB [BGK^{+}95]. Their [GRB^{+}97, HRD98, HYK^{+}98, MM99a, MMD^{+}98, RL99, BCC^{+}95, BGB95b, BM96, DSGF95, FGA^{+}96, KKAN^{+}96, KR96, LS95a, NDBR95, SYB95, SRH^{+}96, VMN95, VSB95, WB96, ZKH^{+}95, CT98, GS99, HKJ^{+}99, HFP^{+}98b, KKH^{+}99, KKS99, MCC^{+}98, MSM99, NPJW99, OWM^{+}99, OMS98, PGP^{+}99, RPS^{+}97, SHB99, SSS^{+}99b, WSSB99]. Theme [MWL97]. themes [Val96]. Therapy [QBvD^{+}98]. thereby [WSW95]. thermophila [CMHT96, GCB^{+}95]. thermotolerance [SNL96]. Thick [BGH^{+}97, FRHS98, LWS99]. Thin [BGH^{+}97, MAM99, GF95, HD99, LRC^{+}99]. Thiol [LLN^{+}95]. thiolase [ZL95b, ZL96, TSSR98]. Thioredoxin [XW96, XMMW97]. Third [BZL^{+}98, DLB96]. those [NSYK^{+}95, YKC95]. Three [EDB^{+}96, JOPM98, MBF^{+}95b, NHA^{+}97, STG^{+}99, SBSG96, UvMJ^{+}99, WSSB99, WMO^{+}95, XC97, BHDW95, CRCH95, DPHW96, LRM^{+}95, RME^{+}95, UPAS95, VGM96, BMS^{+}97, GF97, IFM^{+}99, LMM^{+}99, LRC^{+}99, RNK97, SRR^{+}97, WPW^{+}97, XZSC98]. Three-Dimensional [STG^{+}99, WSSB99, EDB^{+}96, JOPM98, MBF^{+}95b, NHA^{+}97, SBSG96, WMO^{+}95, XC97, BHDW95, CRCH95, LRM^{+}95, RME^{+}95, UPAS95, BMS^{+}97, RNK97, SRR^{+}97, XZSC98]. Threonine [KND^{+}97, ZC99]. Threonines [MMD^{+}98]. Threshold [LKKL97]. Thrombin [HOS^{+}97, GPRS^{+}95]. thrombin-activated [GPRS^{+}95]. Thrombocytes [MPR^{+}97]. Thrombomucin [MPR^{+}97]. Thrombopoietin [NMT97]. Thrombopoietin-induced [NMT97]. Thrombospondin [GLD^{+}96, Bor95, GLP^{+}95, AC95b, DPZ^{+}97, KZS^{+}98]. thrombospondin-1 [GLP^{+}95, DPZ^{+}97]. Thrombospondin-4 [AC95b]. throughout [GMC^{+}95, YGG97a, HB99, KR99]. tht1 [THS^{+}98]. Thy [JMM^{+}98]. Thy-1 [JMM^{+}98]. Thylakoid [HCM^{+}97, RB98]. Thylakoidal [MSMC99]. Thylakoids [ZR98]. thymic [SSM^{+}95]. thymocyte [PLZ^{+}95]. Thymocytes [REC97]. Thymus [VNTM^{+}99, GSP^{+}98]. Thyrocytes [DVL^{+}98]. thyroglobulin [KA95]. Thyroid [SSSS97]. Tia [KGL^{+}99]. Tia-1 [KGL^{+}99]. Tiam1 [MSH^{+}97, SvDtK^{+}98, vLKvdK^{+}97]. Tiam1/Rac [SvDtK^{+}98]. Tiar [KGL^{+}99]. Tic20 [KCFS98]. Tic22 [KCFS98]. tight [BWF^{+}96, KSK^{+}96, RHHRB96, ZZvdB^{+}96, hCMPG97, FFH^{+}98, FSFT98, FST99, HGW^{+}98, IFM^{+}99, IHT^{+}98, JSN98, KRS^{+}98b, MSFA99, MSF^{+}99b, SFD^{+}98, SFS^{+}97, SFS^{+}99, WG97, YHK^{+}97]. Tightly [dBCKS98]. Tim22p [KHS^{+}97]. Tim23 [LJK97]. Tim44 [MVM^{+}99]. Tim54p [KHS^{+}97]. Time [SSM98, GJ95a, SDL98]. Time-Lapse [SSM98]. time-resolved [GJ95a]. timed [CS96a]. Timing [SB99a]. Timp [MFBK99]. Timp-1 [MFBK99]. tip [SGS95, WSGPS95, FSH^{+}99, PNEH98]. Tips [MM99a, TN95, HGI^{+}99]. Tissue [KEJ97, AHBW96, BFS^{+}95, FJT^{+}95, HKS^{+}96, LWB^{+}95, SRR96b, SSB^{+}95, TLF^{+}95, TKM^{+}96, AEWG^{+}97, EvdESvdH^{+}97, GSJ98, KFL^{+}98, KZS^{+}98, LNH^{+}98, NMK^{+}97, NOS^{+}97, OFM^{+}98, SJF^{+}97, XZSC98, ZBG99, ZOKS99]. tissue- [TKM^{+}96]. Tissue-specific [KEJ97, BFS^{+}95, SSB^{+}95, TLF^{+}95, KFL^{+}98, LNH^{+}98]. Tissue-type [GSJ98]. Tissues [OZF^{+}99, BFP^{+}96, FHV^{+}96, GCM^{+}95, LLK^{+}95, RGS^{+}96, TMS^{+}95, GJW^{+}97]. Titin [TGL^{+}98, HHTT95, OGS^{+}96, GTC^{+}98, LRC^{+}99, MSA98]. Tm4sf [TH99]. TNF [MBD^{+}96, VHF97, CIK^{+}99, DCVF98, HDV^{+}99, VHF97]. TNF- [CIK^{+}99]. TNF-induced [HDV^{+}99]. TNF/LT [DCVF98]. TNF/LT- [DCVF98]. TNFalpha [KvEdV^{+}96]. Tobacco [MB99, MBRN95, SGS95, RKB97]. Toc34 [MS98]. Together [Ste97]. Tom [ATH^{+}99, RN99, VL99]. Tom20 [SSS99a]. Tom40 [RN99]. Tomographic [SRR^{+}97]. Tomosyn [LRAB99]. tonsil [PFGS95]. tonsillar [CAS96]. tooth [MLLT95, MMMT95]. Topogenesis [GBS99]. topogenic [CLT^{+}96, WS97]. Topoisomerase [RHF^{+}96, GACM^{+}95, MG96]. Topoisomerases [MKS^{+}97]. Topology [BvdBvM96, MSP97, SG97, LHK^{+}96, WS96, MS98, ZP98]. touch [APG^{+}96]. toxicity [MG96, HSW^{+}98, LRS^{+}98]. Toxin [LR98, MAT^{+}98, LCM^{+}95, MBS96a, NIM95, AFG^{+}98, MSW^{+}98, OF98, WJWM^{+}98, ZLL^{+}98]. Toxoplasma [KQB^{+}98]. TPM1 [DBBB95]. TPM2 [DBBB95]. Tpr [BBP^{+}98, CRRF97, STF98]. Traced [KBI^{+}99]. trachomatis [SFH96]. Tracking [ACSN98, SNT^{+}98]. Tracks [PYM^{+}98]. tract [BSJJ95, MFMW95]. tract-binding [MFMW95]. Tractin [HJJJ97]. Traction [ODJ99]. TRAF2 [HDV^{+}99]. TRAF2-mediated [HDV^{+}99]. Traffic [BPWS98, HE96, AHCC95, FPWN95, FLPS95, LSCM^{+}95, PCYS95, Sch96, SNCH95, GFM^{+}98, HME^{+}98, jLC97, POR^{+}97, PCD^{+}98, SPS^{+}97a]. trafficked [SFH96]. Trafficking [BMG^{+}99a, BM99, FNH99, YSK95, CS96a, GKVR95, GM95, LCK95, RSJK95, SKR96, AKCC99, BHC99, CHW^{+}99, CSEM99, DSG98, DBLD^{+}99, EGS^{+}98, GMS^{+}98, GSL98, KBC99, LTW^{+}97, LMB98, LCW^{+}98, MWL97, SMS^{+}98, WG99, YLD97, ZSP^{+}98]. TRAIL [MMAK97]. TRAIL-induced [MMAK97]. TRAM [VJHR96]. Trans [MM99b, AS96, CSA^{+}97, JMS97, KHHS97, MR95, WGR^{+}95, WBG^{+}97, HDDKH98, KBC99]. Trans-dominant [HDDKH98]. Trans-Golgi [MM99b, AS96, WGR^{+}95]. trans-most [MR95]. Transactivation [HDV^{+}99, SSS^{+}98]. transbilayer [BvdBvM96]. Transcript [BAC97, ISM97]. Transcription [ABAAS98, BDvdZW95, BSAJ97, TB99, BJB^{+}96, GBBSO96, HTS^{+}95, HS96, LLN^{+}95, PSBB96, RACHV96, SS95a, WS95b, XJM^{+}95, APRB98, CTB^{+}97, FWBSO97, HDES98, IRW^{+}99, KGA^{+}97, KWN^{+}99, MSH^{+}98, MCC^{+}98, SRA^{+}99, VLAHV98, VvdKMvD99, WSSB99]. Transcription-dependent [BDvdZW95, BJB^{+}96]. Transcriptional [BHD^{+}97, BFP^{+}96, LMMO96, SP96, SJA96, JLID98, QWS^{+}98]. transcriptionally [OMW^{+}96]. transcripts [KBSW95, MFMW95, TMS^{+}95]. transcytosis [CSM^{+}96, vGvM95, DFD^{+}97]. Transcytotic [vIZKH97, YSK95, GFM^{+}98]. Transduced [IAS^{+}98]. Transducer [BAH^{+}98]. transduction [BTTB96, BGR^{+}95, CSM^{+}96, HG95a, RSC96, YJB^{+}95, SM97, SBS^{+}98, WJWM^{+}98]. transendothelial [WAMS96, TCYS99]. transepithelial [BWF^{+}96]. Transfected [KGW^{+}95, EELF^{+}95, LR96, WBG^{+}97]. Transfection [LMC^{+}98]. Transfer [SPBCM97, KE98, SBR^{+}96, MAE^{+}97, QBvD^{+}98, SPL^{+}97]. transferase [WSW95, DNC^{+}95]. Transferrin [CAR^{+}97, JGD^{+}96, SSF^{+}95, DvD95, GM95, MLJM95, SK95a, OT97, SHH97, WBS^{+}98]. Transferrin-binding [SSF^{+}95]. transferrin-like [DvD95]. transformation [HAS^{+}96, MSB^{+}97, NMK^{+}97, SPS97b, SSS^{+}99b]. transformed [KCDB95, TALM96, VTG^{+}96, WXS96]. Transforming [AM95b, EHM^{+}95, ZMM^{+}95, DMS^{+}95, GWLH98, NGMR97, SBPR^{+}98, SKS99]. transgene [AHBW96, BSK96, XGR95]. Transgenic [CRM^{+}96, ADP^{+}95, BSK96, BGGR^{+}96, BFP^{+}96, CLT^{+}95, CCMG97, DSvdV^{+}96, FWHW96, KAT^{+}95, KMOO95, LMMO96, LDCB95, REdC95, TLF^{+}95, TEL^{+}95, HCLG99, KKT^{+}98, KMS^{+}98, LNH^{+}98, MFBK99, MNG^{+}98, PC98, SATW98, YNM^{+}99, ZhTA^{+}97]. Transglutaminase [AKP95, KSS^{+}95, NGMR97, NMNL98]. Transglutaminase-catalyzed [AKP95]. Transglutaminase-dependent [NGMR97]. Transient [DLC98, WPLK95, BZBA99, ZS98]. transients [KJBM^{+}96]. Transit [RL99, MKLS96]. Transition [LARH97, STK^{+}98, ABN96, STT^{+}96, TMM^{+}95, ZKB95, ZKSB96, BSRG99, EAOS^{+}98, MBOC97, SYT97]. Transitional [RSB^{+}99]. Transitions [MPL^{+}97]. Translation [VTG^{+}96, YCR^{+}98, GMGGD98, KBC99, PWD^{+}99]. Translational [SBV97, GGC96]. Translocase [BSRG99, CIRG99, VL99]. Translocase-1 [BSRG99]. translocating [MDS95]. Translocation [FUB^{+}96, AWS96, BM96, CGVS^{+}96, GPRS^{+}95, LD95, LS95b, MKLS96, MBM95, MGS96, MLO^{+}95, NMHS95, RRRA95, ST96, SRR96b, SRG^{+}96, STVR95, BATP99, HOS^{+}97, HRA^{+}98, KRH98, KGA^{+}97, LBBP98, Mah96, MVM^{+}99, MJBR98, NZ97, SSI^{+}97, SKY^{+}98, SSS^{+}98, SVMB97, YNM^{+}99]. Translocon [CS97b, GBS99]. Transmembrane [CWMD97, OZF^{+}99, RW95, KP96, KSR96, LHK^{+}96, MTLW95, SSN96, SPMB^{+}96, SOR^{+}96, BH98, HTM^{+}98, JMD^{+}98, LNRR98, MS98, OTS98, OBW98, PGD^{+}97, RPE^{+}97, TH99]. Transmigration [MPLS^{+}98, WMR^{+}98]. Transmission [HB98, GCE^{+}97, OPB98, TTT^{+}99]. transmitter [UHJ^{+}96]. Transplacental [GJB^{+}98]. Transplantation [BMPP99, RRA97, VWK^{+}95, VKR^{+}96]. Transport [AAD^{+}97, CDH^{+}98, HHRdB^{+}98, KSOK97, LAW^{+}98, LR98, MBS96a, MAT^{+}98, MSMC99, MXSRB96, NAsG^{+}98, PdHTV99, PB98, RCD99, SME98, SYB95, VG99, WBSNV97, ZWT^{+}97, vWDG^{+}95, ABR95, AGPG96, ABRB95a, BDE^{+}95, Dav95, EKPK95, FTH96, GBN95, GG95a, HKL96, HW96a, JLRS95, KLW^{+}96, KWCS96, KSL96, LHS96, MWL^{+}96, MKB^{+}95, MBM95, MJ96, MH95b, MGRS96, NOR^{+}96, NH95, NAV96, PAV^{+}96, RSR^{+}96, SLL96, SSY^{+}96, SOR^{+}95, SRG^{+}96, SDTE95, VJHR96, WHM^{+}96, WLR96, WFBO^{+}96, YNOH95, YSK95, YKRS96, YSB96, AYP^{+}99, AEVB98, ANK97, AFHS99, Bar97, BRAM97, BGN^{+}97, CES^{+}98, DRE97, DSP97, DSC^{+}99, GE97, GDB^{+}98, GYRWR97, HW98, HOS^{+}97, HCM^{+}97, HME^{+}98, HSB98, IJB^{+}97, JDK97, JH97, KC99, KRH98, KS98, KB99b, KBF^{+}98, LNRR98, LLG^{+}99, LRS^{+}98, LLSO98, MSW^{+}98, MWL97, MZL97, MCRB97, OMS98, PWW98, PSH^{+}98, PMBM^{+}99]. Transport [RPY^{+}99, RPE^{+}97, SWS97b, SWO^{+}99, SEH97b, SBD^{+}99, SS98, TPMB97, WRC^{+}99, WJM^{+}99, YHO^{+}98, ZH97, ZhTA^{+}97, ZLL^{+}98, vMNS97]. transport-deficient [MKB^{+}95]. Transported [MM99b]. transporter [EM96, HLSM95, MTL^{+}96, SSY^{+}96, SVT96, VYB95, WS96, HOS^{+}97, SGM^{+}97, SL97]. transporters [PXR^{+}95, LE97]. Transportin [KBD99, SEK^{+}97]. Transportin-mediated [SEK^{+}97]. Transportin-SR [KBD99]. transports [NSYK^{+}95]. trapping [GTL^{+}96]. Travel [IMS^{+}98]. Travels [BM99]. Treadmill [The97]. Treadmilling [SB99d, WSS97]. treated [DKD^{+}96, LCK95]. Treatment [SKA^{+}96, KGMF95, ZPL^{+}98]. triadin [CFC^{+}95]. triads [PPF96, FCTS99]. Trichinella [FGA^{+}96]. trigger [MKXY95, SMZ^{+}98]. Triggered [RHP99, RC98, RFVCM^{+}99]. triggering [CQB^{+}96, CBM^{+}98, CRMB95, OSH^{+}96]. Triggers [NNSN97, DZC^{+}96, DSC^{+}95, KSVZ95, LGM^{+}97, MSB^{+}97]. Trimerized [VHF97]. trimers [DPHW96, MPSC98]. trinucleotide [TMS^{+}95]. Triorganellar [BHC99]. Tripartite [BGL^{+}99]. Triphosphatase [CNG98]. triphosphate [MBM95, NH95]. triphosphate-binding [NH95]. Trisphosphate [FSM^{+}99, TT96, WS95a, YHMS^{+}98, IMK99, PFP97]. TrkA [ABP^{+}98, BKK^{+}97, CDRB^{+}96, RDM^{+}96, EKPK95, SC97]. tRNA [LTE^{+}96, RPB97]. Trophic [FJ98]. Tropoelastin [DBOM98]. tropomodulin [GF95]. Tropomyosin [DBBB95, PSB98, GF95, KSM96, WLML95]. Tropomyosin-containing [PSB98]. Troponin [MAB98, MGA^{+}96, KAFB99, TWiK^{+}99]. TRPC3 [ZOS^{+}97]. truncated [NWPW95, WF96, GHW99, SJF^{+}97, SBGR98, WRG98]. Truncation [BL96]. Trypanosoma [BCWA97, MGB^{+}97, RFK^{+}97, SSF^{+}95, WCV^{+}97]. trypanosome [RSP^{+}95, RROA95]. trypanosome-soluble [RROA95]. trypanosomes [RSR^{+}96]. Tryptophan [BSH99]. Tsetse [RFK^{+}97]. Tub4p [MJMS96, SGGS96]. tube [CCB^{+}95, KLS^{+}99, TGW99]. tubes [CFK95, FSH^{+}99]. tubular [BSN95, KvEdV^{+}96, MW95b, SJ95, FHWV97]. Tubule [DCJ97, SPS^{+}97a, TPM^{+}99, WFS97]. Tubules [BDO^{+}97, FGA^{+}98, MWL97, YPDM98, PWZS97, WHM^{+}96]. Tubulin [GGC96, MMN^{+}99, MJ96, Sal99, SGGS96, TLF^{+}97, CLS96, ES96, FTH96, GCB^{+}95, LTVC96, LJ95, MJMS96, OWM^{+}99, SJ95, SS95b, DZS^{+}98, FNDS99, KR99, LSH^{+}99, MGIZ98, MZAO98, MUS98, NMH^{+}98, TCMB98, VSRP97]. Tubulogenesis [PBA^{+}97]. tubulovesicular [WSGPS95, PKCS98]. Tumor [DCVF98, GKO99, GTY98, SPK^{+}97, TSC^{+}98, LCK95, LGdC95, MF96b, MBD^{+}96, NAP^{+}96, SBR^{+}96, TSS^{+}95b, MCS^{+}97a, MGB^{+}97, SPS97b, SEH^{+}97a, SB99c, VHF97, VWG^{+}97]. Tumor-suppressor [SEH^{+}97a]. tumorigenesis [LLRTP95]. tumorigenicity [SGYL^{+}96]. Tumors [SEH^{+}97a]. TuRC [MGIZ98]. Turn [KRGA98]. turning [TK95]. Turnover [CLS^{+}97, RL99, RAA^{+}97, The97, FHV^{+}96, HSK^{+}96, LCK95, ASP^{+}97, WSS97]. tweezers [SK95a, SNT^{+}98]. Twinfilin [GDL98]. twinstar [GBW^{+}95]. Twist [MPCW97]. Two [ARM^{+}95, BOS^{+}97, BL99, BS97, DPM^{+}97, JMS97, LRG^{+}98, LBCB96, MKLS96, MML96, NCZ^{+}95, PBT^{+}95, SSB96, BBOB96, CFK95, CS96a, DFC^{+}96, ES96, HKC^{+}95, JRLS95, KKAN^{+}96, KI96, MBRN95, NKI^{+}95, NSC96, OWA95, PXR^{+}95, SSN96, ST96, SZB^{+}95, ZPS96, vdBLCvM96, ALM97b, BBL^{+}99b, BVKP98, BF97b, BGN^{+}97, DSM^{+}98a, HH98, HJJJ97, JR98, LRG^{+}98, MBW^{+}99, MWF^{+}99, OWM^{+}99, RDS^{+}99, SWS97a, SDHM99, TTG95, TCV^{+}98, UvMJ^{+}99, WNB97, vMNS97]. Two-dimensional [MML96, CFK95, DFC^{+}96]. two-hybrid [SZB^{+}95]. Type [FCTS99, FHY99, GJW^{+}97, KSY97, LGdC95, NCW^{+}99, RST^{+}99, TOSF^{+}98, YF99, ZOKS99, BPS99, BLL^{+}96, BH98, GSJ98, GLF^{+}96, GGR^{+}96, KDvFP96, KKFN^{+}95, KSK^{+}96, LOD96, LWB^{+}95, MHLB96, MLO^{+}95, MM96, PHO^{+}95, REdC95, SPBCM97, STS96, SM95c, YtDH^{+}95, ZL96, dHvPL^{+}95, ABE^{+}98, BNJ99, BGL^{+}99, CMD^{+}98, CJH^{+}98, CIRG99, EvdESvdH^{+}97, ETP^{+}99, FCTS99, GWLH98, GGK97, KGF^{+}97, LEB^{+}97, OTL^{+}97, OZF^{+}99, PDS^{+}97b, RKR^{+}97, SJF^{+}97, SSSS97, SEAB^{+}99, TSH^{+}98, WFS97, WBR^{+}98]. Type-1 [OTL^{+}97]. type-specific [MM96]. Types [DFL^{+}99, FHWV97, MKS^{+}97]. typhimurium [MHK^{+}95, dPF95]. Typical [CGSY^{+}99]. tyrosinase [BBOB96]. Tyrosine [ALM^{+}97a, CCS^{+}98, CH95, FSDA97, GRTB97, KCDB95, LKSW98, MN98, SWVR98, AS96, BGB^{+}95a, BKRT95, CDRB^{+}96, FDH96, FWG^{+}96, GJ95a, GHP^{+}96, GLS96, GZK^{+}95, KSR96, LS95a, LDF95b, MWOB96, MBS95, MPW95, MTGY96, MA95, RSRK96, RDM^{+}96, SWB^{+}96, SHT^{+}95, SHPP96, SR96, Wal95, XGC96, YYH^{+}96, ZL95a, ZMG96b, ALCC^{+}99, BALL98, BATP99, BD99, BSM^{+}99, CWMD97, DP98, HIWL97, KJR^{+}98, LHB^{+}99, LWSL^{+}98, OTS98, OWW^{+}99, PRV98, SZFM98, SLN^{+}97, SC97, SBGR98, SHB99, SDC^{+}97, VC97, cXP97, ZDK^{+}99]. tyrosine- [MWOB96]. Tyrosine-based [LKSW98, GLS96, BD99]. tyrosine-phosphorylated [CDRB^{+}96].

U2 [KGGK99]. U2AF [GCKC^{+}97]. U937 [vdBCH^{+}95]. UBF [JMTCF96]. Ubiquitin [KSF^{+}99, BIR96, MCB96, HBPJ98, JB99, SSR^{+}97, SSRP99]. Ubiquitin-Conjugating [KSF^{+}99, HBPJ98]. ubiquitin-like [BIR96, MCB96]. Ubiquitination [RD96b, FY99, HZR98, RSD98]. ubiquitous [CK96, SS95a, GDL98]. ubiquitously [HKL96]. UDP [FFDP98, LCP95]. Ultrastructural [AGB^{+}96, PAN97, WMO^{+}95, DSvDH^{+}98]. ultrastructure [FGA^{+}96, KSM96, ETP^{+}99]. Ultraviolet [AKM^{+}98]. umbilical [GHP^{+}96]. unassisted [KIT^{+}97]. unattached [RCKS95, CSMM98]. unc [BTTB96, MCA96, OFY^{+}95, SLP^{+}96, KNT^{+}99, OBB99, BBOE98]. unc-44 [OFY^{+}95]. Unc-45 [BBOE98]. UNC-60B [OBB99]. unc-68 [MCA96]. unc-7 [SLP^{+}96]. UNC-76 [KNT^{+}99]. unc-89 [BTTB96]. Unc104 [PdHTV99]. Unc104/Kif1a [PdHTV99]. uncapped [MMP^{+}95]. Uncapping [SB99b, SJC96]. Unconventional [HGG^{+}97, MM95a, LM98, LG99, SD98]. Uncovers [MM97a]. underglycosylation [KKG95]. undergo [SSL^{+}95, KCV^{+}98]. undergoing [DH96]. underlie [VBZ96]. Underlies [BHC99]. Undifferentiated [PHM^{+}98]. Unexpected [hYGYF97]. Unfolded [TGL^{+}98]. Unfolding [TSB99]. unified [DEG^{+}96]. Unique [SWD^{+}98, PMCT96, BINRM97, NMT97, PWGG98, SSM98, VC97]. unit [LBCB96]. units [LCP95, JP98, LGM98, PFAA98]. unmasks [WHM^{+}96, LZK^{+}97]. Unmyelinated [HSL^{+}99]. Unstable [NNSN97, KTC95, RGS99]. until [FdEP99, MPT^{+}98]. untranslated [RZZ96, BM98]. Unusual [KQB^{+}98, VSRP97]. up-regulation [SRR96a]. upon [AW96, BM98, CLF^{+}97a, GBBSO96, KGMF95, LHS97b, MMR99a, NCZ^{+}95, PL96, SLPE^{+}98, TYT^{+}97, ZN95a]. Upregulates [vHGB^{+}97]. Upregulation [IRW^{+}99, RPS^{+}97, SSS^{+}99b]. upstream [BGGR^{+}96]. Uptake [WSG^{+}98, FKS95, HP95, SSF^{+}95, BJR^{+}99, GJB^{+}98, MGB^{+}97]. urchin [CCP96, MMS^{+}96b, STRM95, WTCV95, WWP^{+}96, HCRS98, LSM^{+}98, MS97b]. Urokinase [FIC^{+}97, NCW^{+}99, NCV^{+}98, CNP^{+}95, DCW^{+}96, KvEdV^{+}96, SM95c, CMD^{+}98, GKO99, WYL^{+}99, YKO97]. Urokinase-Type [NCW^{+}99, SM95c]. Uropod [SALdP^{+}97]. Uropods [dPCM^{+}97]. Us11 [SSRP99]. Us11-Dependent [SSRP99]. Use [CJR^{+}97]. Used [RCE^{+}99]. Uses [LHS97b]. Using [GXE^{+}99, HFP98a, ABN96, CRLM96, HP95, KE98, KLW^{+}96, MHLB96, MAM95, NKI^{+}95, OMO95, RSL^{+}96, ASP^{+}97, EKS97, KQB^{+}98, RDP^{+}99, RV98, SSI^{+}97, WLK97]. Uso1p [SLSW96]. Utrophin [DPT^{+}97, GMS97]. Utrophin-deficient [DPT^{+}97, GMS97]. UV [HWFA96, MRM^{+}98].

V [TNY^{+}95, BLLB95, EKSC98, GLD^{+}96, GZC^{+}99, JH96, LCS96, MBS^{+}96b, PHU^{+}95, SAC^{+}98, WMR^{+}98, ZMM^{+}95, DNC^{+}95, GBN95, GSP^{+}98, MHLB96, PT97, RKR^{+}99, SHPB99, WBR^{+}98, CRG^{+}95, NOR^{+}96, POR^{+}97, TKVC95, UNPW98, UvMJ^{+}99, vMNS97]. v-fos [JH96]. V-like [GSP^{+}98]. V-src [TNY^{+}95]. V159N [BD98]. v3 [JBD^{+}95, BJS^{+}95]. VA [Bri99]. Vac8p [WCW98]. VacA [CHH97]. Vaccinia [LG99]. Vacuolar [DBE98, GHS98, CS96b, MBRN95, NCS95, PCYS95, VE95, vWDG^{+}95, BSH99, GWB^{+}98, KQB^{+}98, RSD98, SMCE97, UNPW98, WCW98]. vacuolating [LHM^{+}96, CHH97]. Vacuolation [AFG^{+}98]. Vacuole [BOS^{+}97, BPWS98, PBS97, VG99, XMMW97, CF95, HMSK95, HCW96, HDK96, NWPW95, NCS95, OSHG^{+}96, SK95b, SDTE95, XW96, CC98, DRE97, HC97, KSOK97, SBOK97, SHL97, WCW98]. Vacuole/Lysosome [BOS^{+}97, VG99]. Vacuoles [JR98, LGGS97, MW97, SWS97b, UNPW98, UvMJ^{+}99]. Valency [SLS^{+}99].

Vam3p [DRE97, DBE98]. VAMP [MTL^{+}96, GK96, RGS^{+}96, AYP^{+}99]. VAMP-2 [MTL^{+}96]. Vamp-7 [AYP^{+}99]. VAMP/ [RGS^{+}96]. Vap [SBD^{+}99]. Vap-33 [SBD^{+}99]. variable [CFK95]. variably [BMG^{+}95]. variant [CLR^{+}96, SRH^{+}96, AKCC99, GBvdN^{+}98, LSM^{+}98, WSR^{+}97]. variants [JBD^{+}95, KBM^{+}95, BGH^{+}98, GF97, vHGB^{+}97]. Variations [MWL97, HW96b]. variety [BFS^{+}95]. Varying [NTAN98]. Vascular [AEWG^{+}97, EWW^{+}98, HTK^{+}99, JCR97, SPK^{+}97, HMB95a, LCC^{+}95, LMMO96, YWW^{+}96, BP98, CMD^{+}98, FHSM^{+}97, JLID98, KZS^{+}98, LVN^{+}98, SPR^{+}98, TCYS99, qXLCH97]. vasodilator [STP96]. vasodilator-stimulated [STP96]. VASP [LLH^{+}99]. Vault [ARR^{+}98, HGI^{+}99, KSK^{+}99a]. VCAM [AKC^{+}95, AFFS96]. VCAM-1 [AKC^{+}95, AFFS96]. VE [LCC^{+}95, NRD98]. VE- [NRD98]. VE-cadherin [LCC^{+}95]. vegetative [SCK^{+}95]. VEGF [SPR^{+}98, KvEdV^{+}96, KKT^{+}98]. vein [GHP^{+}96, SV98]. Velocity [ZLL^{+}98]. Ventral [GS99, MRL^{+}99, ZhTA^{+}97, LTR^{+}97]. ventricle [CFC^{+}95]. Ventricular [HWPC96, CMS98]. venules [PFGS95, CFS98]. Vero [MBS96a]. verotoxin [KLW^{+}96]. Verprolin [ABE^{+}98, VMH97]. versus [CQB^{+}96, SK95a, SJ99]. Vertebrate [HC99, ARM^{+}95, MH95b, PMMF95, XHUC96, BKM^{+}98, CBS^{+}97, GCA^{+}98, MBTW97, PFDL97, RC98]. Vesicle [BML^{+}97, MM95b, PT97, SW99, SDTE95, ABRB95a, EGH^{+}95, GEK95, GHF^{+}95, GK96, MTL^{+}96, PR95, RLC^{+}96, SJSJ95, TMDD96, WSBE^{+}95, CBTZ98, DFL^{+}99, FHK97, NTH98, NBJ^{+}99, POR^{+}97, PB98, SHPB99, TMM97, VCL^{+}98, VCS^{+}99, WNB97, YHO^{+}98, ZCOP99, vMNS97]. Vesicle-associated [PT97, MTL^{+}96]. Vesicle-mediated [SDTE95]. Vesicles [EGKC^{+}99, RWOA97, WSCN97, AGPG96, AS96, CGB^{+}95, DCOK95, GHF^{+}95, HKL96, KI96, KBW^{+}96, KSL96, LHS96, MGRS96, NDBR95, dPF95, RAM^{+}96, SOR^{+}95, SIAS96, SOR^{+}96, SWC^{+}96, SOG96, WTCV95, CSA^{+}97, CC98, DSvDH^{+}98, DSG98, DSM^{+}97, EGS^{+}98, FGA^{+}98, HC97, JMM^{+}98, JH97, KNSP97, KKG^{+}98, LH97, LLSO98, MCRB97, OMS98, PSB98, SFR^{+}98, SLL^{+}98, SBD^{+}99, YHMS^{+}98, YPDM98]. Vesicular [ZWT^{+}97, Dav95, FLPS95, GBN95, MXSRB96, WFBO^{+}96, AYP^{+}99, LE97, RLVM98, SBOK97, SMS^{+}98]. vesiculated [AMJM95]. Vesiculation [HTK^{+}98]. VI [BFP^{+}96, BKJL^{+}98, KsYA98, LM98]. VI-mediated [KsYA98]. Via [ZLPL99, ANK97, AKM^{+}98, BM99, BPWS98, DH99, DBG^{+}97, FFLP99, HOS^{+}97, HWFA96, IMW^{+}99, JLM99, KKG^{+}98, KSJW95, MM99b, MLO^{+}95, NCS95, NCV^{+}98, OSPJ^{+}96, PWG^{+}95, PKCS98, SKA^{+}96, SHR^{+}95, SLS^{+}99, TSD^{+}97, TCM96, VNTM^{+}99, WMC^{+}95a, WLF99, vHGB^{+}97]. Viability [FFDP98, HY95, HCM96a, MVM^{+}99, MLC99]. Viable [ULMT99]. Vid24p [CC98]. videomicroscopy [CGB^{+}95]. View [FPRL97]. VII [BTNZ^{+}95, KCN^{+}96, RKR^{+}97]. Vik1p [MBW^{+}99]. Villin [FCTPA^{+}99, PPWL97]. Villin/Gelsolin [PPWL97]. Villus [CCMG97]. Vimentin [PJE^{+}99, GCGE^{+}96, GG95b, NNMW96, OIT^{+}95, TOT^{+}96, CCC^{+}99, PYM^{+}98, YMPG98]. Vinculin [LZK^{+}97, MNS^{+}99, WUUI^{+}98, WKR^{+}98, BPT96, OSPJ^{+}96, SSW^{+}96, KSK^{+}99b, WKR^{+}98]. Vinculin-binding [MNS^{+}99, KSK^{+}99b]. Vinexin [KSK^{+}99b]. Viral [MFM^{+}95, MB99, ASB^{+}99]. Virus [MB99, PPW^{+}95, BWG95, DPHW96, IM96, KKGD96, KGMF95, LNR^{+}96, LGdC95, MXSRB96, SLL96, SPMB^{+}96, SHG96, CTB^{+}97, KKM^{+}97, KSV^{+}99, KS98, LNRR98, LHWH97, LG99, MM97a, MBDS^{+}98, PDO^{+}99, PMBM^{+}99, RLVM98, SEH97b, VPM^{+}98]. Virus-mediated [PPW^{+}95]. Visceral [HTK^{+}99, LMMO96]. Visual [RST^{+}99]. Visualization [ASP99, DFC^{+}96, NTH98, SWM^{+}98, VB98, WTM^{+}99, WWD^{+}97, WBR^{+}98, RSL^{+}96, GBMS95, SSI^{+}97, SCPPW98, WSSB99]. visualized [SHS96, SPS^{+}97a]. visualizing [VE95]. vital [VE95, MBDBB98]. Vitelline [TGTL97]. vitro [All95, AHM^{+}95, BEG^{+}95, CGJ^{+}95, DCOK95, FBD^{+}95, KSL96, LNY^{+}95, MB96, PPW^{+}95, SKA^{+}96, SJC96, SK95b, SSA^{+}96, VKP^{+}95, DPZ^{+}97, DKN^{+}97, EWW^{+}98, FTB97, FSA^{+}99, GCML98, IFSV^{+}98, KDG98, LL97, LTMR97, LLG^{+}99, MSH^{+}98, MDV98, NBS^{+}98, SFR^{+}98, SS98, TGW99, WLK97, WFS97, YLD97]. vitronectin [LGRB96, HAB^{+}99]. Vivo [BSAJ97, CT97a, YTO^{+}99, BKRT95, CGJ^{+}95, FBD^{+}95, HSWC96, HG95b, JMTCF96, MGS^{+}95, NA96, OP96a, PFF^{+}95, RSL^{+}96, SKA^{+}96, SRG^{+}96, VKP^{+}95, WPZ^{+}96, WSSM95, WB96, BD98, BKM^{+}98, BF97a, BKMN^{+}98, BW97, CJR^{+}97, CHW^{+}99, FCTPA^{+}99, GBvdN^{+}98, GDM^{+}99, HC99, HMB^{+}98, HVL^{+}99, HHS^{+}99, KFO^{+}99, MLN^{+}98, MLB98, MCC^{+}98, OSR^{+}98, PWGG98, RBRB99, qSmLY97, TCV^{+}98, UNPW98, WPW^{+}97, WHG^{+}97, WBR^{+}98, YPB^{+}98, dBCKS98]. VLA [AKC^{+}95, RAH^{+}95]. VLA-4 [AKC^{+}95]. VLA-5 [RAH^{+}95]. Vma12p [GHS98]. Vma12p/Vma22p [GHS98]. Vma22p [GHS98]. Voltage [SSS99a, WS98, ZLM^{+}98]. Voltage-dependent [SSS99a]. Voltage-gated [WS98, ZLM^{+}98]. Volume [YGG97b]. volumes [EDB^{+}96]. VP16 [TSB99]. Vparp [KSK^{+}99a]. Vpr [JMWG98]. VPS [PBS97]. VPS-dependent [PBS97]. vps1 [NCS95]. Vps10p [CS96b, SMCE97]. Vps15 [SDTE95]. VPS27 [PCYS95]. VPS29 [SMCE97]. VPS30 [SMCE97]. Vps33p [VG99]. Vps33p-Dependent [VG99]. Vps34 [SDTE95]. VPS35 [SMCE97]. vs [HHB^{+}97]. Vti1p [vMNS97]. vulgaris [RS95, KMI^{+}97]. vzg [HWPC96]. vzg-1 [HWPC96].

WAF1 [FLW^{+}95]. walk [YvdEH^{+}95]. Wall [OBB99, LMB^{+}95, MCA96, DAF^{+}97, DH99]. warthog [PAT99]. Wasp [ELL^{+}99, BRS98]. WASP-related [BRS98]. water [RW95]. water-soluble [RW95]. Wave [GCE^{+}97]. waves [LZC^{+}95, MKI^{+}96, BRD99]. WD [CBB^{+}96, WKK^{+}95, ZL95b]. WD-repeat [CBB^{+}96, WKK^{+}95]. weak [TNY^{+}95]. Weaver [LW96]. Web [BHC99]. weight [BKP^{+}95]. Weights [HFE^{+}98]. well [MFMW95, RRRA95]. Werner [SSS^{+}99b]. Wheat [ABAAS98]. Where [JLID98]. Whereas [BGH^{+}97]. whey [LDCB95]. which [CLR^{+}96, GEW^{+}95, KGMF95, OCS^{+}95, TN95, LPM^{+}99, VZNR98]. while [VLZR96, LPM^{+}98]. White [BPS99]. Whole [PvDA^{+}98]. widely [GCM^{+}95, PJL^{+}95, RGS^{+}96, WKRB96]. widespread [FJT^{+}95]. WIF [IMS^{+}98, PCD^{+}98]. Wild [YF99, GLF^{+}96, PHO^{+}95, dHvPL^{+}95, BNJ99, LEB^{+}97, WBR^{+}98]. Wild-Type [YF99, GLF^{+}96, PHO^{+}95, dHvPL^{+}95]. wing [BMLM96, EAL^{+}95, EWS96, MZL^{+}99]. wingless [OP96a, RSC96, MDK^{+}99]. Wiscott [Li97]. withdrawal [AW96, SLW^{+}99]. within [BWWT97, DMG96, FAO^{+}96, FST99, GBS99, HBJ95, NRP98, NNMW96, SMB^{+}99, WBR^{+}98, ZSB99, GB98]. without [GH95, KKAN^{+}96, Lan97, Ols99, PPF96, RS95, dS98a]. Wnt [FhJZ^{+}99, LTR^{+}97, SGR97, TYSP^{+}96]. Wnt-1 [TYSP^{+}96]. Wnt-5A [TYSP^{+}96]. Wnt-signaling [SGR97]. Work [Ste97]. works [YNOH95]. wortmannin [MBRN95, SJHC96, Dav95]. wound [PTBC96, ZMM^{+}95, CMD^{+}98, OTL^{+}97]. wounds [BMLM96].

X [NNK^{+}97, CCMC97, CMWL96, CCBL98, EDB^{+}96, FHY99, HST^{+}98, KPZ^{+}97, LGdC95, MCPB99]. XAC [RAA^{+}97]. XCTK2 [WVM97]. xenografts [SKA^{+}96]. Xenopus [KSG^{+}96, LAW^{+}98, AOMB96, AK97b, BK96, hCMPG97, CLL^{+}98, DMMS98, DFL^{+}99, EJ99, FFGG96, FB96, FRHS98, FKS95, GHJJ95, GHW99, HTH^{+}97, KGE^{+}98, LTR^{+}97, LKWG96, LGM98, MCJK98, MM97b, MRL^{+}99, MBF^{+}97, MKI^{+}96, NG99, RD96a, RAA^{+}97, SRH^{+}95, SGR97, TGN97, TGTL97, TYSP^{+}96, TZL^{+}98, VLAHV98, VWG^{+}97, WVM97, WSGPS95]. Xgrip109 [MGIZ98]. XI [TKM^{+}96]. XII [KBM^{+}95]. XIIIb [FLPS95, LLVS98]. XIST [CMWL96, CCBL98]. Xklp2 [WBA^{+}98]. Xklp3 [LAW^{+}98]. Xlrbpa [EJ97]. Xmad1 [CSMM98]. Xmad2 [CSMM98]. XMAP310 [AK97b]. XTC [WZF97].

Yarrowia [ESR97, STVR95, TSSR98]. Yeast [BIR96, BPWS98, EGH^{+}95, FK95, FWL^{+}98, FKNM97, GHS98, HER95, wJTSSL98, KBI^{+}99, LRAB99, MFL^{+}99, MMATe96, SBS^{+}99, SME98, TOB^{+}99, VG99, WH97, XMMW97, ZC99, ZWT^{+}97, ABR95, ARM^{+}95, BGN^{+}96, BM96, CF95, CP95, CMM^{+}95, CPL^{+}95, CS96b, DBBB95, EM96, GKVR95, GEK95, GAW^{+}96, GDC95, GBN95, HY95, HMSK95, HM95, HW96a, HB95a, HDK96, HB95b, IWW95, JRLS95, KTC95, KM95, LR95, LZD95, LK95, LW95, LHS96, MJMS96, MBM95, NRM^{+}96, NCS95, RD96b, STN^{+}96, SSY^{+}96, SH96, SHR^{+}95, SSSKR95, SK95b, SOR^{+}95, SCK^{+}95, SCKG95, SSP95, SFNRH95, SMC^{+}96, SCG^{+}95, SDTE95, SZB^{+}95, dCBR95, SVT96, STVR95, VPH95, VH96, VMN95, VE95, WKWC96, WM96, WMXE96, YGK96a, YGK96b, YSC^{+}95, ABE^{+}98, AL97, ASP^{+}97, BSW^{+}98, BSD97, BSH99, BBAN^{+}99, BHHF97, BD97, BD98, BLL^{+}96]. Yeast [BSY97, BHJ98, BG97, BVKP98, BS97, BW97, CS97a, CDN97, CYH98, CIRG99, CYSD99, DSG98, DS98b, DDRC99, EM98, ELMS98, FHAP98, FN98, FFLP99, GE97, GDL98, GWB^{+}99, GCD97, HBS^{+}97, HKS97, HTM^{+}98, HTW97, JMS97, JB99, KMMC98, KHA^{+}99, KWS^{+}99, LL97, Li97, LL98b, LEB^{+}97, LFP98, jLC97, MCM^{+}99, MFF97, MW97, MB97, MSL98b, MR98, MUS98, Nei98, NA99, OAB^{+}98, OC98, OKB^{+}98, PNEH98, PSB98, RHLG98, RSD98, SS97, SHC^{+}97, SN98, SHPB99, SGSW97, SYM^{+}97, SY99, SSSG98, SR97, SB99b, Ste97, SHL97, SSM98, SBR98, TNFK^{+}98, THS^{+}98, TCMB98, TSSR98, VMH97, VCS^{+}99, WE98, YWMH99, YYSB97, vMNS97]. Yes [MKB^{+}99]. Yes-Associated [MKB^{+}99]. yields [FFZ^{+}95]. yolk [GHJJ95]. Ypt [JMS97]. Ypt1 [JRLS95, JLRS95]. Ypt7p [MW97].

Z [BSMH95, FPF^{+}99, GTC^{+}98, SBSG96]. Z-band [FPF^{+}99]. ZAP [HIWL97, SLPE^{+}98, SWVR98]. ZAP-70 [HIWL97, SLPE^{+}98, SWVR98]. ZASP [FPF^{+}99]. zebrafish [CM95a]. Zellweger [FH97]. zero [DSC^{+}95]. zeta [XZSC96]. Zinc [DSVL^{+}98, HDV^{+}99, BLG95, RHA^{+}97, SYT97]. zinc-binding [BLG95]. Zinc-finger [DSVL^{+}98]. Zip1 [SR95]. Zip1-induced [SR95]. Zipper [WBA^{+}98]. ZO [HGW^{+}98, INMT97, YHK^{+}97, CDH^{+}99, IFM^{+}99, MSFA99]. ZO-1 [HGW^{+}98, INMT97, YHK^{+}97, CDH^{+}99, IFM^{+}99]. Zo-2 [CDH^{+}99, IFM^{+}99]. ZO-3 [HGW^{+}98, CDH^{+}99, IFM^{+}99]. Zona [GRB^{+}97]. Zone [DSVL^{+}98, HWPC96, FSH^{+}99]. zones [UHJ^{+}96, WKL^{+}99]. zonula [MW96, RHHRB96]. Zonular [WGP98]. ZP3 [AZF96]. ZP3-dependent [AZF96]. ZPR1 [GMGGD98]. zVAD [LNM^{+}98]. zVAD-fmk [LNM^{+}98]. ZW10 [SWL^{+}97, SWHG98, WGG96]. ZYG [MCTMK98]. ZYG-9 [MCTMK98]. zygote [ETF95, HCRS98]. zygotes [STRM95, OPB98]. Zygotic [MCM^{+}99]. Zyxin [DAB99, NB97].

References

Addya:1997:TNT

[AAB^{+}97] Sankar Addya, Hindupur K. Anandatheerthavarada, Gopa Biswas, Shripad V. Bhagwat, Jayati Mullick, and Narayan G. Avadhani. Targeting of NH_{2}-terminal-processed microsomal protein to mitochondria: A novel pathway for the biogenesis of hepatic mitochondrial P450MT2. Journal of Cell Biology, 139(3):589–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

              rupress.org/content/139/3/589.

Ainger:1997:TLE

[AAD^{+}97] Kevin Ainger, Daniela Avossa, Amy S. Diana, Christopher Barry, Elisa Barbarese, and John H. Carson. Transport and Localization Elements in Myelin Basic Protein mRNA. Journal of Cell Biology, 138(5):1077–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

              URL http://jcb.rupress.org/content/138/5/1077.

Adams:1998:RPP

[AAM^{+}98] Christopher M. Adams, Michael G. Anderson, David G. Motto, Margaret P. Price, Wayne A. Johnson, and Michael J. Welsh. Ripped pocket and pickpocket, novel Drosophila DEG/ENaC subunits expressed in early development and in mechanosensory neurons. Journal of Cell Biology, 140(1): 143–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/143.

Adam:1995:CRF

[AAP^{+}95] T. Adam, M. Arpin, M. C. Prévost, P. Gounon, and P. J. Sansonetti. Cytoskeletal rearrangements and the functional role of T-plastin during entry of Shigella flexneri into HeLa cells. Journal of Cell Biology, 129(2):367–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/2/367.

Abu-Amer:1997:SRO

[AARS^{+}97] Yousef Abu-Amer, F. Patrick Ross, Paul Schlesinger, M. Mehrdad Tondravi, and Steven L. Teitelbaum. Substrate recognition by osteoclast precursors induces C-src/ microtubule association. Journal of Cell Biology, 137(1): 247–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/1/247.

Ando-Akatsuka:1996:IDO

[AASH^{+}96] Y. Ando-Akatsuka, M. Saitou, T. Hirase, M. Kishi, A. Sakakibara, M. Itoh, S. Yonemura, M. Furuse, and S. Tsukita. Interspecies diversity of the occludin sequence: cDNA cloning of human, mouse, dog, and rat-kangaroo homologues. Journal of Cell Biology, 133(1):43–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/1/43.

Abranches:1998:TSC

[ABAAS98] Rita Abranches, Alison F. Beven, Luis Aragón-Alcaide, and Peter J. Shaw. Transcription Sites Are Not Correlated with Chromosome Territories in Wheat Nuclei. Journal of Cell Biology, 143(1):5–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/1/5.

Anderson:1998:SHD

[ABE^{+}98] Blake L. Anderson, Istvan Boldogh, Marie Evangelista, Charles Boone, Lloyd A. Greene, and Liza A. Pon. The Src Homology Domain 3 (SH3) of a Yeast Type I Myosin, Myo5p, Binds to Verprolin and Is Required for Targeting to Sites of Actin Polarization. Journal of Cell Biology, 141(6):1357–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/6/1357.

Annaert:1997:ECE

[ABK^{+}97] Wim G. Annaert, Bernd Becker, Ute Kistner, Michael Reth, and Reinhard Jahn. Export of Cellubrevin from the Endoplasmic Reticulum Is Controlled by BAP31. Journal of Cell Biology, 139(6):1397–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/6/1397.

Arregui:1998:IIM

[ABL98] Carlos O. Arregui, Janne Balsamo, and Jack Lilien. Impaired Integrin-mediated Adhesion and Signaling in Fibroblasts Expressing a Dominant-negative Mutant PTP1B. Journal of Cell Biology, 143(3):861–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/3/861.

Alonso:1999:CIC

[ABM^{+}99] Maria Teresa Alonso, Maria José Barrero, Pedro Michelena, Estela Carnicero, Inmaculada Cuchillo, Antonio G. García, Javier García-Sancho, Mayte Montero, and Javier Alvarez. Ca^{2+}-induced Ca^{2+} release in chromaffin cells seen from inside the ER with targeted aequorin. Journal of Cell Biology, 144 (2):241–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/2/241.

Angres:1996:MTI

[ABN96] B. Angres, A. Barth, and W. J. Nelson. Mechanism for transition from initial to stable cell–cell adhesion: kinetic analysis of E-cadherin-mediated adhesion using a quantitative adhesion assay. Journal of Cell Biology, 134(2):549–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/549.

Aloyz:1998:PED

[ABP^{+}98] Raquel S. Aloyz, Shernaz X. Bamji, Christine D. Pozniak, Jean G. Toma, Jasvinder Atwal, David R. Kaplan, and Freda D. Miller. P53 is essential for developmental neuron death as regulated by the TrkA and p75 neurotrophin receptors. Journal of Cell Biology, 143(6):1691–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1691.

Aitchison:1995:NYN

[ABR95] J. D. Aitchison, G. Blobel, and M. P. Rout. Nup120p: a yeast nucleoporin required for NPC distribution and mRNA transport. Journal of Cell Biology, 131(6):1659–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1659.

Aridor:1995:SCB

[ABRB95a] M. Aridor, S. I. Bannykh, T. Rowe, and W. E. Balch. Sequential coupling between COPII and COPI vesicle coats in endoplasmic reticulum to Golgi transport. Journal of Cell Biology, 131(4):875–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/875.

Armeanu:1995:CFI

[ABRB^{+}95b] S. Armeanu, H. J. Bühring, M. Reuss-Borst, C. A. Müller, and G. Klein. E-cadherin is functionally involved in the maturation of the erythroid lineage. Journal of Cell Biology, 131 (1):243–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/243.

Altschuler:1998:RDF

[ABT^{+}98] Yoram Altschuler, Shana M. Barbas, Laura J. Terlecky, Kitty Tang, Stephen Hardy, Keith E. Mostov, and Sandra L. Schmid. Redundant and Distinct Functions for Dynamin-1 and Dynamin-2 Isoforms. Journal of Cell Biology, 143(7): 1871–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/7/1871.

Aigner:1995:APS

[AC95a] L. Aigner and P. Caroni. Absence of persistent spreading, branching, and adhesion in GAP-43-depleted growth cones. Journal of Cell Biology, 128(4):647–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/647.

Arber:1995:TEM

[AC95b] S. Arber and P. Caroni. Thrombospondin-4, an extracellular matrix protein expressed in the developing and adult nervous system promotes neurite outgrowth. Journal of Cell Biology, 131(4):1083–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/1083.

Abney:1997:CDI

[ACF^{+}97] James R. Abney, Bryan Cutler, Misty L. Fillbach, Daniel Axelrod, and Bethe A. Scalettar. Chromatin Dynamics in Interphase Nuclei and Its Implications for Nuclear Structure. Journal of Cell Biology, 137(7):1459–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/7/1459.

Aszodi:1998:CIE

[ACH^{+}98] Attila Aszódi, Danny Chan, Ernst Hunziker, John F. Bateman, and Reinhard Fässler. Collagen II Is Essential for the Removal of the Notochord and the Formation of Intervertebral Discs. Journal of Cell Biology, 143(5):1399–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1399.

Adams:1995:ANA

[ACHG95] L. Adams, B. M. Carlson, L. Henderson, and D. Goldman. Adaptation of nicotinic acetylcholine receptor, myogenin, and MRF4 gene expression to long-term muscle denervation. Journal of Cell Biology, 131(5):1341–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/5/1341.

Alon:1997:KST

[ACP^{+}97] Ronen Alon, Shuqi Chen, Kamal D. Puri, Erik B. Finger, and Timothy A. Springer. The Kinetics of L-selectin Tethers and the Mechanics of Selectin-mediated Rolling. Journal of Cell Biology, 138(5):1169–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/5/1169.

Adams:1998:MEC

[ACSN98] Cynthia L. Adams, Yih-Tai Chen, Stephen J. Smith, and W. James Nelson. Mechanisms of epithelial cell–cell adhesion and cell compaction revealed by high-resolution tracking of E-Cadherin/green fluorescent protein. Journal of Cell Biology, 142(4):1105–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/4/1105.

Albers:1995:EEK

[ADP^{+}95] K. M. Albers, F. E. Davis, T. N. Perrone, E. Y. Lee, Y. Liu, and M. Vore. Expression of an epidermal keratin protein in liver of transgenic mice causes structural and functional abnormalities. Journal of Cell Biology, 128(1):157–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
1/157.

Ahmad:1998:CDD

[AEVB98] Fridoon J. Ahmad, Christophe J. Echeverri, Richard B. Vallee, and Peter W. Baas. Cytoplasmic Dynein and Dynactin Are Required for the Transport of Microtubules into the Axon. Journal of Cell Biology, 140(2):391–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/2/391.

Aird:1997:VBS

[AEWG^{+}97] William C. Aird, Jay M. Edelberg, Hartmut Weiler-Guettler, William W. Simmons, Thomas W. Smith, and Robert D. Rosenberg. Vascular Bed–specific Expression of an Endothelial Cell Gene Is Programmed by the Tissue Microenvironment. Journal of Cell Biology, 138(5):1117–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
5/1117.

Alon:1996:ITS

[AFFS96] R. Alon, R. C. Fuhlbrigge, E. B. Finger, and T. A. Springer. Interactions through L-selectin between leukocytes and adherent leukocytes nucleate rolling adhesions on selectins and VCAM-1 in shear flow. Journal of Cell Biology, 135(3):849– ??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/849.

Abrami:1998:PFT

[AFG^{+}98] Laurence Abrami, Marc Fivaz, Pierre-Etienne Glauser, Robert G. Parton, and F. van der Goot. A Pore-forming Toxin Interacts with a GPI-anchored Protein and Causes Vacuolation of the Endoplasmic Reticulum. Journal of Cell Biology, 140(3):525–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/525.

Alvarez:1999:GT

[AFHS99] Cecilia Alvarez, Hideaki Fujita, Ann Hubbard, and Elizabeth Sztul. ER to Golgi Transport. Journal of Cell Biology, 147(6): 1205–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/6/1205.

Aubert:1996:UBC

[AGB^{+}96] S. Aubert, E. Gout, R. Bligny, D. Marty-Mazars, F. Barrieu, J. Alabouvette, F. Marty, and R. Douce. Ultrastructural and biochemical characterization of autophagy in higher plant cells subjected to carbon deprivation: control by the supply of mitochondria with respiratory substrates. Journal of Cell Biology, 133(6):1251–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/6/1251.

Albertinazzi:1998:ONS

[AGP^{+}98] Chiara Albertinazzi, Daniela Gilardelli, Simona Paris, Renato Longhi, and Ivan de Curtis. Overexpression of a Neural-specific Rho Family GTPase, cRac1B, Selectively Induces Enhanced Neuritogenesis and Neurite Branching in Primary Neurons. Journal of Cell Biology, 142(3):815–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/3/815.

Aniento:1996:EBC

[AGPG96] F. Aniento, F. Gu, R. G. Parton, and J. Gruenberg. An endosomal beta COP is involved in the pH-dependent formation of transport vesicles destined for late endosomes. Journal of Cell Biology, 133(1):29–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/1/29.

Anastasi:1997:NHG

[AGS^{+}97] Sergio Anastasi, Silvia Giordano, Olga Sthandier, Giovanna Gambarotta, Rossella Maione, Paolo Comoglio, and Paolo Amati. A natural hepatocyte growth factor/scatter factor autocrine loop in myoblast cells and the effect of the constitutive met kinase activation on myogenic differentiation. Journal of Cell Biology, 137(5):1057–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/5/1057.

Alexander:1996:RME

[AHBW96] C. M. Alexander, E. W. Howard, M. J. Bissell, and Z. Werb. Rescue of mammary epithelial cell apoptosis and entactin degradation by a tissue inhibitor of metalloproteinases-1 transgene. Journal of Cell Biology, 135(6):1669–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
6/1669.

Antony:1995:DRM

[AHCC95] C. Antony, M. Huchet, J. P. Changeux, and J. Cartaud. Developmental regulation of membrane traffic organization during synaptogenesis in mouse diaphragm muscle. Journal of Cell Biology, 130(4):959–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/4/959.

Allinquant:1995:DAP

[AHM^{+}95] B. Allinquant, P. Hantraye, P. Mailleux, K. Moya, C. Bouillot, and A. Prochiantz. Downregulation of amyloid precursor protein inhibits neurite outgrowth in vitro. Journal of Cell Biology, 128(5):919–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/5/919.

Allet:1996:DPA

[AHM^{+}96] B. Allet, A. Hochmann, I. Martinou, A. Berger, M. Missotten, B. Antonsson, R. Sadoul, J. C. Martinou, and L. Bernasconi. Dissecting processing and apoptotic activity of a cysteine protease by mutant analysis. Journal of Cell Biology, 135(2): 479–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/2/479.

Araki:1996:RPK

[AJS96] N. Araki, M. T. Johnson, and J. A. Swanson. A role for phosphoinositide 3-kinase in the completion of macropinocytosis and phagocytosis by macrophages. Journal of Cell Biology, 135(5):1249–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/5/1249.

Ahn:1995:SBA

[AK95] A. H. Ahn and L. M. Kunkel. Syntrophin binds to an alternatively spliced exon of dystrophin. Journal of Cell Biology, 128 (3):363–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/3/363.

Altman:1997:CME

[AK97a] Roger Altman and Douglas Kellogg. Control of Mitotic Events by Nap1 and the Gin4 Kinase. Journal of Cell Biology, 138 (1):119–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/1/119.

Andersen:1997:XXR

[AK97b] Søren S. L. Andersen and Eric Karsenti. XMAP310: a Xenopus Rescue-promoting Factor Localized to the Mitotic Spindle. Journal of Cell Biology, 139(4):975–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
4/975.

Adams:1999:LCS

[AK99] Ian R. Adams and John V. Kilmartin. Localization of Core Spindle Pole Body (SPB) Components during SPB Duplication in Saccharomyces cerevisiae. Journal of Cell Biology, 145 (4):809–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/809.

Alon:1995:IVS

[AKC^{+}95] R. Alon, P. D. Kassner, M. W. Carr, E. B. Finger, M. E. Hemler, and T. A. Springer. The integrin VLA-4 supports tethering and rolling in flow on VCAM-1. Journal of Cell Biology, 128(6):1243–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/6/1243.

Adair-Kirk:1999:ITV

[AKCC99] Tracy L. Adair-Kirk, Kathleen H. Cox, and John V. Cox. Intracellular Trafficking of Variant Chicken Kidney Ae1 Anion Exchangers. Journal of Cell Biology, 147(6):1237–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/6/1237.

Ainsztein:1998:ICS

[AKLME98] Alexandra M. Ainsztein, Stefanie E. Kandels-Lewis, Alastair M. Mackay, and William C. Earnshaw. INCENP Centromere and Spindle Targeting: Identification of Essential Conserved Motifs and Involvement of Heterochromatin Protein HP1. Journal of Cell Biology, 143(7):1763–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1763.

Aragane:1998:ULI

[AKM^{+}98] Yoshinori Aragane, Dagmar Kulms, Dieter Metze, Gabriele Wilkes, Birgit Pöppelmann, Thomas A. Luger, and Thomas Schwarz. Ultraviolet light induces apoptosis via direct activation of CD95 (Fas/APO-1) independently of its ligand CD95L. Journal of Cell Biology, 140(1):171–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/1/171.

Aeschlimann:1995:TCM

[AKP95] D. Aeschlimann, O. Kaupp, and M. Paulsson. Transglutaminasecatalyzed matrix cross-linking in differentiating cartilage: identification of osteonectin as a major glutaminyl substrate. Journal of Cell Biology, 129(3):881–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/3/881.

Arkowitz:1997:SCD

[AL97] Robert A. Arkowitz and Nick Lowe. A Small Conserved Domain in the Yeast Spa2p Is Necessary and Sufficient for Its Polarized Localization. Journal of Cell Biology, 138(1):17–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/1/17.

Annaert:1999:PCS

[ALC^{+}99] Wim G. Annaert, Lyne Levesque, Kathleen Craessaerts, Inge Dierinck, Greet Snellings, David Westaway, Peter St. George-Hyslop, Barbara Cordell, Paul Fraser, and Bart De Strooper. Presenilin 1 controls γ-secretase processing of amyloid precursor protein in pre-Golgi compartments of hippocampal neurons. Journal of Cell Biology, 147(2):277–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
2/277.

Angers-Loustau:1999:PTP

[ALCC^{+}99] Alexandre Angers-Loustau, Jean-François Côté, Alain Charest, Donald Dowbenko, Susan Spencer, Laurence A. Lasky, and Michel L. Tremblay. Protein Tyrosine Phosphatase–PEST Regulates Focal Adhesion Disassembly, Migration, and Cytokinesis in Fibroblasts. Journal of Cell Biology, 144(5): 1019–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/5/1019.

Altschuler:1999:ARF

[ALK^{+}99] Y. Altschuler, S.-H. Liu, L. Katz, K. Tang, S. Hardy, F. Brodsky, G. Apodaca, and K. Mostov. Adp-ribosylation Factor 6 and Endocytosis at the Apical Surface of Madin–Darby Canine Kidney Cells. Journal of Cell Biology, 147(1):7–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
1/7.

Allan:1995:PPR

[All95] V. Allan. Protein phosphatase 1 regulates the cytoplasmic dynein-driven formation of endoplasmic reticulum networks in vitro. Journal of Cell Biology, 128(5):879–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/879.

Aicher:1997:CRP

[ALM^{+}97a] Babette Aicher, Markus M. Lerch, Thomas Müller, James Schilling, and Axel Ullrich. Cellular redistribution of protein tyrosine phosphatases LAR and PTPσ by inducible proteolytic processing. Journal of Cell Biology, 138(3):681–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/681.

Andreassen:1997:CTI

[ALM97b] Paul R. Andreassen, Françoise B. Lacroix, and Robert L. Margolis. Chromosomes with Two Intact Axial Cores Are Induced by G_{2} Checkpoint Override: Evidence That DNA Decatenation Is not Required to Template the Chromosome Structure. Journal of Cell Biology, 136(1):29–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/1/29.

Andreassen:1998:DSL

[ALVMM98] Paul R. Andreassen, Françoise B. Lacroix, Emma Villa-Moruzzi, and Robert L. Margolis. Differential Subcellular Localization of Protein Phosphatase-1 α, γ1, and δ Isoforms during Both Interphase and Mitosis in Mammalian Cells. Journal of Cell Biology, 141(5):1207–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/5/1207.

Arneson:1995:EEL

[AM95a] L. S. Arneson and J. Miller. Efficient endosomal localization of major histocompatibility complex class II-invariant chain complexes requires multimerization of the invariant chain targeting sequence. Journal of Cell Biology, 129(5):1217–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1217.

Arribas:1995:TGF

[AM95b] J. Arribas and J. Massagué. Transforming growth factoralpha and beta-amyloid precursor protein share a secretory mechanism. Journal of Cell Biology, 128(3):433–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
3/433.

Acharya:1995:RVG

[AMJM95] U. Acharya, J. M. McCaffery, R. Jacobs, and V. Malhotra. Reconstitution of vesiculated Golgi membranes into stacks of cisternae: requirement of NSF in stack formation. Journal of Cell Biology, 129(3):577–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/3/577.

Adebanjo:1998:MAI

[AMY^{+}98] Olugbenga A. Adebanjo, Baljit S. Moonga, Tomoo Yamate, Li Sun, Cedric Minkin, Etsuko Abe, and Mone Zaidi. Mode of Action of Interleukin-6 on Mature Osteoclasts. Novel Interactions with Extracellular Ca^{2+} Sensing in the Regulation of Osteoclastic Bone Resorption. Journal of Cell Biology, 142(5): 1347–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/5/1347.

Allison:1999:ERR

[AN99] David C. Allison and Andrea L. Nestor. Evidence for a Relatively Random Array of Human Chromosomes on the Mitotic Ring. Journal of Cell Biology, 145(1):1–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/145/1/1.

Akita:1997:IPT

[ANK97] Mitsuru Akita, Erik Nielsen, and Kenneth Keegstra. Identification of Protein Transport Complexes in the Chloroplastic Envelope Membranes via Chemical Cross–Linking. Journal of Cell Biology, 136(5):983–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/5/983.

Aono:1999:PAI

[ANRT99] Shinya Aono, Shinichi Nakagawa, Albert B. Reynolds, and Masatoshi Takeichi. p120^{ctn} Acts as an Inhibitory Regulator of Cadherin Function in Colon Carcinoma Cells. Journal of Cell Biology, 145(3):551–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/3/551.

Adams:1996:QAC

[ANS96] C. L. Adams, W. J. Nelson, and S. J. Smith. Quantitative analysis of cadherin-catenin-actin reorganization during development of cell–cell adhesion. Journal of Cell Biology, 135(6): 1899–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1899.

Airas:1997:DRF

[ANS^{+}97] Laura Airas, Jussi Niemelä, Marko Salmi, Tarja Puurunen, David J. Smith, and Sirpa Jalkanen. Differential Regulation and Function of CD73, a Glycosyl–Phosphatidylinositol– linked 70-kD Adhesion Molecule, on Lymphocytes and Endothelial Cells. Journal of Cell Biology, 136(2):421–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/2/421.

Amling:1997:BLD

[ANT^{+}97] Michael Amling, Lynn Neff, Sakae Tanaka, Daisuke Inoue, Keisuke Kuida, Eleanor Weir, William M. Philbrick, Arthur E. Broadus, and Roland Baron. Bcl-2 Lies Downstream of Parathyroid Hormone–related Peptide in a Signaling Pathway That Regulates Chondrocyte Maturation during Skeletal Development. Journal of Cell Biology, 136(1):205–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
1/205.

Abe:1996:XLA

[AOMB96] H. Abe, T. Obinata, L. S. Minamide, and J. R. Bamburg. Xenopus laevis actin-depolymerizing factor/cofilin: a phosphorylation-regulated protein essential for development. Journal of Cell Biology, 132(5):871–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/5/871.

Abitorabi:1997:PIL

[APF^{+}97] M. Abi Abitorabi, Russell K. Pachynski, Ronald E. Ferrando, Mark Tidswell, and David J. Erle. Presentation of Integrins on Leukocyte Microvilli: a Role for the Extracellular Domain in Determining Membrane Localization. Journal of Cell Biology, 139(2):563–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/2/563.

Albers:1996:CON

[APG^{+}96] K. M. Albers, T. N. Perrone, T. P. Goodness, M. E. Jones, M. A. Green, and B. M. Davis. Cutaneous overexpression of NT-3 increases sensory and sympathetic neuron number and enhances touch dome and hair follicle innervation. Journal of Cell Biology, 134(2):487–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/487.

Alphey:1997:KMK

[APH^{+}97] Luke Alphey, Louise Parker, Gillian Hawcroft, Yiquan Guo, Kim Kaiser, and Gareth Morgan. KLP38B: a Mitotic Kinesinrelated Protein That Binds PP1. Journal of Cell Biology, 138 (2):395–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/2/395.

Albertini:1998:NNI

[APRB98] Markus Albertini, Lucy F. Pemberton, Jonathan S. Rosenblum, and Günter Blobel. A Novel Nuclear Import Pathway for the Transcription Factor TFIIS. Journal of Cell Biology, 143(6):1447–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1447.

Aitchison:1995:TNR

[ARM^{+}95] J. D. Aitchison, M. P. Rout, M. Marelli, G. Blobel, and R. W. Wozniak. Two novel related yeast nucleoporins Nup170p and Nup157p: complementation with the vertebrate homologue Nup155p and functional interactions with the yeast nuclear pore-membrane protein Pom152p. Journal of Cell Biology, 131(5):1133–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/5/1133.

Abbondanza:1998:IVP

[ARR^{+}98] Ciro Abbondanza, Valentina Rossi, Annarita Roscigno, Luigi Gallo, Angela Belsito, Giulio Piluso, Nicola Medici, Vincenzo Nigro, Anna Maria Molinari, Bruno Moncharmont, and Giovanni A. Puca. Interaction of Vault Particles with Estrogen Receptor in the MCF-7 Breast Cancer Cell. Journal of Cell Biology, 141(6):1301–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/6/1301.

Austin:1996:FNS

[AS96] C. D. Austin and D. Shields. Formation of nascent secretory vesicles from the trans-Golgi network of endocrine cells is inhibited by tyrosine kinase and phosphatase inhibitors. Journal of Cell Biology, 135(6):1471–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/6/1471.

Almeida:1998:MAC

[ASACF98] Fátima Almeida, Rainer Saffrich, Wilhelm Ansorge, and Maria Carmo-Fonseca. Microinjection of Anti-coilin Antibodies Affects the Structure of Coiled Bodies. Journal of Cell Biology, 142(4):899–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/4/899.

Anton:1999:ILP

[ASB^{+}99] Luis C. Antón, Ulrich Schubert, Igor Bacík, Michael F. Princiotta, Pamela A. Wearsch, James Gibbs, Patricia M. Day, Claudio Realini, Martin C. Rechsteiner, Jack R. Bennink, and Jonathan W. Yewdell. Intracellular Localization of Proteasomal Degradation of a Viral Antigen. Journal of Cell Biology, 146(1):113–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/113.

Acakpo-Satchivi:1997:GMD

[ASES^{+}97] Leslie J. R. Acakpo-Satchivi, Winfried Edelmann, Carol Sartorius, Brian D. Lu, Philip A. Wahr, Simon C. Watkins, Joseph M. Metzger, Leslie Leinwand, and Raju Kucherlapati. Growth and Muscle Defects in Mice Lacking Adult Myosin Heavy Chain Genes. Journal of Cell Biology, 139(5): 1219–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1219.

Afshar:1995:ICL

[ASH95] K. Afshar, J. Scholey, and R. S. Hawley. Identification of the chromosome localization domain of the Drosophila nod kinesin-like protein. Journal of Cell Biology, 131(4):833– ??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/4/833.

Ayscough:1997:HRA

[ASP^{+}97] Kathryn R. Ayscough, Joel Stryker, Navin Pokala, Miranda Sanders, Phil Crews, and David G. Drubin. High Rates of Actin Filament Turnover in Budding Yeast and Roles for Actin in Establishment and Maintenance of Cell Polarity Revealed Using the Actin Inhibitor Latrunculin–A. Journal of Cell Biology, 137(2):399–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/2/399.

Andjelkovic:1999:VCB

[ASP99] Anuska V. Andjelkovic, Dennis D. Spencer, and Joel S. Pachter. Visualization of Chemokine Binding Sites on Human Brain Microvessels. Journal of Cell Biology, 145(2):403–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/403.

Assoian:1997:ADC

[Ass97] Richard K. Assoian. Anchorage-dependent Cell Cycle Progression. Journal of Cell Biology, 136(1):1–??, January 1997.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/1/1.

Ayme-Southgate:1995:BSA

[ASSS^{+}95] A. Ayme-Southgate, R. Southgate, J. Saide, G. M. Benian, and M. L. Pardue. Both synchronous and asynchronous muscle isoforms of projectin (the Drosophila bent locus product) contain functional kinase domains. Journal of Cell Biology, 128 (3):393–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/3/393.

Aizawa:1996:OCS

[ASY96] H. Aizawa, K. Sutoh, and I. Yahara. Overexpression of cofilin stimulates bundling of actin filaments, membrane ruffling, and cell movement in Dictyostelium. Journal of Cell Biology, 132 (3):335–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/335.

Agarraberes:1997:IHR

[ATD97] Fernando A. Agarraberes, Stanley R. Terlecky, and J. Fred Dice. An Intralysosomal hsp70 Is Required for a Selective Pathway of Lysosomal Protein Degradation. Journal of Cell Biology, 137(4):825–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/4/825.

Adachi:1997:DIR

[ATH^{+}97] Hiroyuki Adachi, Yasuhiro Takahashi, Takeshi Hasebe, Mikako Shirouzu, Shigeyuki Yokoyama, and Kazuo Sutoh. Dictyostelium IQGAP-related Protein Specifically Involved in the Completion of Cytokinesis. Journal of Cell Biology, 137 (4):891–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/4/891.

Ahting:1999:TCC

[ATH^{+}99] Uwe Ahting, Clemens Thun, Reiner Hegerl, Dieter Typke, Frank E. Nargang, Walter Neupert, and Stephan Nussberger. The Tom Core Complex. Journal of Cell Biology, 147(5): 959–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/5/959.

Abrami:1999:PMM

[AvdG99] Laurence Abrami and F. Gisou van der Goot. Plasma Membrane Microdomains Act as Concentration Platforms to Facilitate Intoxication by Aerolysin. Journal of Cell Biology, 147 (1):175–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/1/175.

Andres:1996:MEC

[AW96] V. Andrés and K. Walsh. Myogenin expression, cell cycle withdrawal, and phenotypic differentiation are temporally separable events that precede cell fusion upon myogenesis. Journal of Cell Biology, 132(4):657–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/4/657.

Aridor:1998:CSC

[AWB^{+}98] Meir Aridor, Jacques Weissman, Sergei Bannykh, Claude Nuoffer, and William E. Balch. Cargo Selection by the COPII Budding Machinery during Export from the ER. Journal of Cell Biology, 141(1):61–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/141/1/61.

Astoul:1999:DPK

[AWC99] Emmanuelle Astoul, Sandra Watton, and Doreen Cantrell. The dynamics of protein kinase B regulation during B cell antigen receptor engagement. Journal of Cell Biology, 145 (7):1511–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/7/1511.

Anderson:1996:CPT

[AWS96] K. I. Anderson, Y. L. Wang, and J. V. Small. Coordination of protrusion and translocation of the keratocyte involves rolling of the cell body. Journal of Cell Biology, 134(5):1209– ??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/5/1209.

Allen:1996:MEM

[AYF96] E. Allen, Q. C. Yu, and E. Fuchs. Mice expressing a mutant desmosomal cadherin exhibit abnormalities in desmosomes, proliferation, and epidermal differentiation. Journal of Cell Biology, 133(6):1367–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/6/1367.

Ahmad:1999:ERK

[AYMB99] Fridoon J. Ahmad, Wenqian Yu, Francis J. McNally, and Peter W. Baas. An Essential Role for Katanin in Severing Microtubules in the Neuron. Journal of Cell Biology, 145(2):305–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/305.

Advani:1999:VMV

[AYP^{+}99] Raj J. Advani, Bin Yang, Rytis Prekeris, Kelly C. Lee, Judith Klumperman, and Richard H. Scheller. Vamp-7 Mediates Vesicular Transport from Endosomes to Lysosomes. Journal of Cell Biology, 146(4):765–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/4/765.

Arnoult:1996:ZDA

[AZF96] C. Arnoult, Y. Zeng, and H. M. Florman. ZP3-dependent activation of sperm cation channels regulates acrosomal secretion during mammalian fertilization. Journal of Cell Biology, 134 (3):637–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/3/637.

Allen:1998:RCM

[AZRJ98] William E. Allen, Daniel Zicha, Anne J. Ridley, and Gareth E. Jones. A Role for Cdc42 in Macrophage Chemotaxis. Journal of Cell Biology, 141(5):1147–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/5/1147.

Borgese:1996:RMP

[BAC^{+}96] N. Borgese, D. Aggujaro, P. Carrera, G. Pietrini, and M. Bassetti. A role for N-myristoylation in protein targeting: NADHcytochrome b5 reductase requires myristic acid for association with outer mitochondrial but not ER membranes. Journal of Cell Biology, 135(6):1501–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/6/1501.

Baumeister:1997:AMA

[BAC97] Anja Baumeister, Silvia Arber, and Pico Caroni. Accumulation of Muscle Ankyrin Repeat Protein Transcript Reveals Local Activation of Primary Myotube Endcompartments during Muscle Morphogenesis. Journal of Cell Biology, 139(5): 1231–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1231.

Bertolotto:1998:MGP

[BAH^{+}98] Corine Bertolotto, Patricia Abbe, Timothy J. Hemesath, Karine Bille, David E. Fisher, Jean-Paul Ortonne, and Robert Ballotti. Microphthalmia Gene Product as a Signal Transducer in cAMP–Induced Differentiation of Melanocytes. Journal of Cell Biology, 142(3):827–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/3/827.

Balsamo:1998:NPT

[BALL98] Janne Balsamo, Carlos Arregui, TinChung Leung, and Jack Lilien. The Nonreceptor Protein Tyrosine Phosphatase PTP1B Binds to the Cytoplasmic Domain of N–Cadherin and Regulates the Cadherin–Actin Linkage. Journal of Cell Biology, 143(2):523–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/523.

Barlowe:1997:CGT

[Bar97] Charles Barlowe. Coupled ER to Golgi Transport Reconstituted with Purified Cytosolic Proteins. Journal of Cell Biology, 139(5):1097–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/5/1097.

Bi:1995:CRE

[BAS95] G. Q. Bi, J. M. Alderton, and R. A. Steinhardt. Calciumregulated exocytosis is required for cell membrane resealing. Journal of Cell Biology, 131(6):1747–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1747.

Boudreau:1997:IAP

[BAS^{+}97] Nancy Boudreau, Catherine Andrews, Anabella Srebrow, Ali Ravanpay, and David A. Cheresh. Induction of the Angiogenic Phenotype by Hox D3. Journal of Cell Biology, 139(1): 257–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/257.

Blanco-Aparicio:1999:NRM

[BATP99] Carmen Blanco-Aparicio, Josema Torres, and Rafael Pulido. A Novel Regulatory Mechanism of Map Kinases Activation and Nuclear Translocation Mediated by Pka and the Ptp–Sl Tyrosine Phosphatase. Journal of Cell Biology, 147(6):1129–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/6/1129.

Bannykh:1997:MDE

[BB97] Sergei I. Bannykh and William E. Balch. Membrane dynamics at the endoplasmic reticulum-Golgi interface. Journal of Cell Biology, 138(1):1–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/1/1.

Becker:1999:EKR

[BBAN^{+}99] Frank Becker, Laura Block-Alper, Gerald Nakamura, Josephine Harada, K. Dane Wittrup, and David I. Meyer. Expression of the 180-kD Ribosome Receptor Induces Membrane Proliferation and Increased Secretory Activity in Yeast. Journal of Cell Biology, 146(2):273–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/2/273.

Brugarolas:1998:PCC

[BBJ98] James Brugarolas, Roderick T. Bronson, and Tyler Jacks. p21 Is a Critical CDK2 Regulator Essential for Proliferation Control in Rb -deficient Cells. Journal of Cell Biology, 141 (2):503–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/2/503.

Basu:1999:MES

[BBL^{+}99a] Joydeep Basu, Hassan Bousbaa, Elsa Logarinho, ZeXiao Li, Byron C. Williams, Carla Lopes, Claudio E. Sunkel, and Michael L. Goldberg. Mutations in the essential spindle checkpoint gene bub1 cause chromosome missegregation and fail to block apoptosis in Drosophila. Journal of Cell Biology, 146 (1):13–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/13.

Bistrup:1999:STS

[BBL^{+}99b] Annette Bistrup, Sunil Bhakta, Jin Kyu Lee, Yevgeniy Y. Belov, Michael Dee Gunn, Feng-Rong Zuo, Chiao-Chain Huang, Reiji Kannagi, Steven D. Rosen, and Stefan Hemmerich. Sulfotransferases of Two Specificities Function in the Reconstitution of High Endothelial Cell Ligands for L-selectin. Journal of Cell Biology, 145(4):899–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/4/899.

Bertolotto:1996:RTG

[BBOB96] C. Bertolotto, K. Bille, J. P. Ortonne, and R. Ballotti. Regulation of tyrosinase gene expression by cAMP in B16 melanoma cells involves two CATGTG motifs surrounding the TATA box: implication of the microphthalmia gene product. Journal of Cell Biology, 134(3):747–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/3/747.

Barral:1998:UMC

[BBOE98] José M. Barral, Christopher C. Bauer, Irving Ortiz, and Henry F. Epstein. Unc-45 mutations in Caenorhabditis elegans implicate a CRO1/She4p-like domain in myosin assembly. Journal of Cell Biology, 143(5):1215–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1215.

Bangs:1998:FAT

[BBP^{+}98] Peter Bangs, Brian Burke, Christine Powers, Roger Craig, Aruna Purohit, and Stephen Doxsey. Functional Analysis of Tpr: Identification of Nuclear Pore Complex Association and Nuclear Localization Domains and a Role in mRNA Export. Journal of Cell Biology, 143(7):1801–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1801.

Beh:1997:KEN

[BBR97] Christopher T. Beh, Valeria Brizzio, and Mark D. Rose. KAR5 Encodes a Novel Pheromone-inducible Protein Required for Homotypic Nuclear Fusion. Journal of Cell Biology, 139(5): 1063–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1063.

Baldini:1998:ERN

[BBW^{+}98] Giulia Baldini, Giovanna Baldini, Guangyi Wang, Mattew Weber, Marina Zweyer, Renato Bareggi, Joan W. Witkin, and Alberto M. Martelli. Expression of Rab3D N135I Inhibits Regulated Secretion of ACTH in AtT-20 Cells. Journal of Cell Biology, 140(2):305–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/2/305.

Balczon:1995:DCR

[BBZ^{+}95] R. Balczon, L. Bao, W. E. Zimmer, K. Brown, R. P. Zinkowski, and B. R. Brinkley. Dissociation of centrosome replication events from cycles of DNA synthesis and mitotic division in hydroxyurea-arrested Chinese hamster ovary cells. Journal of Cell Biology, 130(1):105–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/1/105.

Bass:1998:FIR

[BCAS98] Joseph Bass, Gavin Chiu, Yair Argon, and Donald F. Steiner. Folding of Insulin Receptor Monomers Is Facilitated by the Molecular Chaperones Calnexin and Calreticulin and Impaired by Rapid Dimerization. Journal of Cell Biology, 141 (3):637–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/3/637.

Bastianutto:1995:OCI

[BCC^{+}95] C. Bastianutto, E. Clementi, F. Codazzi, P. Podini, F. De Giorgi, R. Rizzuto, J. Meldolesi, and T. Pozzan. Overexpression of calreticulin increases the Ca^{2+} capacity of rapidly exchanging Ca^{2+} stores and reveals aspects of their lumenal microenvironment and function. Journal of Cell Biology, 130 (4):847–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/847.

Baines:1995:QLP

[BCMK95] I. C. Baines, A. Corigliano-Murphy, and E. D. Korn. Quantification and localization of phosphorylated myosin I isoforms in Acanthamoeba castellanii. Journal of Cell Biology, 130(3): 591–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/3/591.

Bornslaeger:1996:BCD

[BCSG96] E. A. Bornslaeger, C. M. Corcoran, T. S. Stappenbeck, and K. J. Green. Breaking the connection: displacement of the desmosomal plaque protein desmoplakin from cell–cell interfaces disrupts anchorage of intermediate filament bundles and alters intercellular junction assembly. Journal of Cell Biology, 134(4):985–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/4/985.

Burleigh:1997:CSE

[BCWA97] Barbara A. Burleigh, Elisabet V. Caler, Paul Webster, and Norma W. Andrews. A Cytosolic Serine Endopeptidase from Trypanosoma cruzi Is Required for the Generation of Ca^{2+} Signaling in Mammalian Cells. Journal of Cell Biology, 136 (3):609–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/609.

Belgareh:1997:DNP

[BD97] Naïma Belgareh and Valérie Doye. Dynamics of Nuclear Pore Distribution in Nucleoporin Mutant Yeast Cells. Journal of Cell Biology, 136(4):747–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/4/747.

Belmont:1998:YVA

[BD98] Lisa D. Belmont and David G. Drubin. The Yeast V159N Actin Mutant Reveals Roles for Actin Dynamics In Vivo. Journal of Cell Biology, 142(5):1289–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/5/1289.

Bonifacino:1999:MBR

[BD99] Juan S. Bonifacino and Esteban C. Dell’Angelica. Molecular Bases for the Recognition of Tyrosine-based Sorting Signals. Journal of Cell Biology, 145(5):923–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/5/923.

Brown:1995:RPK

[BDE^{+}95] W. J. Brown, D. B. DeWald, S. D. Emr, H. Plutner, and W. E. Balch. Role for phosphatidylinositol 3-kinase in the sorting and transport of newly synthesized lysosomal enzymes in mammalian cells. Journal of Cell Biology, 130(4):781–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
4/781.

Butler:1997:AIS

[BDO^{+}97] Margaret Husta Butler, Carol David, Gian-Carlo Ochoa, Zachary Freyberg, Laurie Daniell, Detlev Grabs, Ottavio Cremona, and Pietro De Camilli. Amphiphysin II (SH3P9; BIN1), a member of the amphiphysin/Rvs family, is concentrated in the cortical cytomatrix of axon initial segments and nodes of ranvier in brain and around T tubules in skeletal muscle. Journal of Cell Biology, 137(6):1355–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/6/1355.

Bastos:1997:NNN

[BdPE^{+}97] Ricardo Bastos, Lluis Ribas de Pouplana, Mark Enarson, Khaldon Bodoor, and Brian Burke. Nup84, a novel nucleoporin that is associated with CAN/Nup214 on the cytoplasmic face of the nuclear pore complex. Journal of Cell Biology, 137(5):989–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/5/989.

Burgess:1997:SPS

[BDS97] Robert W. Burgess, David L. Deitcher, and Thomas L. Schwarz. The synaptic protein syntaxin1 is required for cellularization of Drosophila embryos. Journal of Cell Biology, 138 (4):861–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/4/861.

Bregman:1995:TDR

[BDvdZW95] D. B. Bregman, L. Du, S. van der Zee, and S. L. Warren. Transcription-dependent redistribution of the large subunit of RNA polymerase II to discrete nuclear domains. Journal of Cell Biology, 129(2):287–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/2/287.

Beauvais:1995:CFS

[BEG^{+}95] A. Beauvais, C. A. Erickson, T. Goins, S. E. Craig, M. J. Humphries, J. P. Thiery, and S. Dufour. Changes in the fibronectin-specific integrin expression pattern modify the migratory behavior of sarcoma S180 cells in vitro and in the embryonic environment. Journal of Cell Biology, 128(4): 699–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/699.

Burkhardt:1997:ODP

[BENV97] Janis K. Burkhardt, Christophe J. Echeverri, Tommy Nilsson, and Richard B. Vallee. Overexpression of the Dynamitin (p50) Subunit of the Dynactin Complex Disrupts Dynein-dependent Maintenance of Membrane Organelle Distribution. Journal of Cell Biology, 139(2):469–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/2/469.

Balsamo:1995:IRC

[BEZ^{+}95] J. Balsamo, H. Ernst, M. K. Zanin, S. Hoffman, and J. Lilien. The interaction of the retina cell surface Nacetylgalactosaminylphosphotransferase with an endogenous proteoglycan ligand results in inhibition of cadherin-mediated adhesion. Journal of Cell Biology, 129(5):1391–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1391.

Boche:1997:NRN

[BF97a] Irene Boche and Ellen Fanning. Nucleocytoplasmic Recycling of the Nuclear Localization Signal Receptor α Subunit In Vivo Is Dependent on a Nuclear Export Signal, Energy, and RCC1. Journal of Cell Biology, 139(2):313–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
2/313.

Brickner:1997:SEN

[BF97b] Jason H. Brickner and Robert S. Fuller. SOI1 Encodes a Novel, Conserved Protein That Promotes TGN–Endosomal Cycling of Kex2p and Other Membrane Proteins by Modulating the Function of Two TGN Localization Signals. Journal of Cell Biology, 139(1):23–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/1/23.

Bashour:1997:IRC

[BFHB97] Anne-Marie Bashour, Aaron T. Fullerton, Matthew J. Hart, and George S. Bloom. IQGAP1, a Rac- and Cdc42-binding Protein, Directly Binds and Cross-links Microfilaments. Journal of Cell Biology, 137(7):1555–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/7/1555.

Bohmann:1995:MAP

[BFL95] K. Bohmann, J. A. Ferreira, and A. I. Lamond. Mutational analysis of p80 coilin indicates a functional interaction between coiled bodies and the nucleolus. Journal of Cell Biology, 131(4):817–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/817.

Bloch:1997:IET

[BFL^{+}97] Wilhelm Bloch, Erik Forsberg, Sylvia Lentini, Cord Brakebusch, Karl Martin, Hans W. Krell, Ulrich H. Weidle, Klaus Addicks, and Reinhard Fässler. β1 Integrin Is Essential for Teratoma Growth and Angiogenesis. Journal of Cell Biology, 139(1):265–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/1/265.

Braghetta:1996:DRC

[BFP^{+}96] P. Braghetta, C. Fabbro, S. Piccolo, D. Marvulli, P. Bonaldo, D. Volpin, and G. M. Bressan. Distinct regions control transcriptional activation of the alpha1(VI) collagen promoter in different tissues of transgenic mice. Journal of Cell Biology, 135(4):1163–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/4/1163.

Bonilha:1999:EPM

[BFRB99] Vera Lúcia Bonilha, Silvia C. Finnemann, and Enrique Rodriguez-Boulan. Ezrin Promotes Morphogenesis of Apical Microvilli and Basal Infoldings in Retinal Pigment Epithelium. Journal of Cell Biology, 147(7):1533–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1533.

Brinkmann:1995:HGF

[BFS^{+}95] V. Brinkmann, H. Foroutan, M. Sachs, K. M. Weidner, and W. Birchmeier. Hepatocyte growth factor/scatter factor induces a variety of tissue-specific morphogenic programs in epithelial cells. Journal of Cell Biology, 131(6):1573–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1573.

Berry:1997:FYD

[BG97] Lynne D. Berry and Kathleen L. Gould. Fission Yeast dim1^{+} Encodes a Functionally Conserved Polypeptide Essential for Mitosis. Journal of Cell Biology, 137(6):1337–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
6/1337.

Bloom:1998:CAM

[BG98] George S. Bloom and Lawrence S. B. Goldstein. Cruising along Microtubule Highways: How Membranes Move through the Secretory Pathway. Journal of Cell Biology, 140(6):1277–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/6/1277.

Benmerah:1995:TKS

[BGB^{+}95a] A. Benmerah, J. Gagnon, B. Bègue, B. Mégarbané, A. Dautry-Varsat, and N. Cerf-Bensussan. The tyrosine kinase substrate eps15 is constitutively associated with the plasma membrane adaptor AP-2. Journal of Cell Biology, 131(6):1831–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1831.

Berryman:1995:EOM

[BGB95b] M. Berryman, R. Gary, and A. Bretscher. Ezrin oligomers are major cytoskeletal components of placental microvilli: a proposal for their involvement in cortical morphogenesis. Journal of Cell Biology, 131(5):1231–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/5/1231.

Burden-Gulley:1999:PRC

[BGBK99] Susan M. Burden-Gulley and Susann M. Brady-Kalnay. PTPμ regulates N–Cadherin-dependent neurite outgrowth. Journal of Cell Biology, 144(6):1323–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/6/1323.

Balice-Gordon:1998:FGJ

[BGBS98] Rita J. Balice-Gordon, Linda J. Bone, and Steven S. Scherer. Functional Gap Junctions in the Schwann Cell Myelin Sheath. Journal of Cell Biology, 142(4):1095–??, August 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/142/4/
1095.

Bonaccorsi:1998:SSO

[BGG98] Silvia Bonaccorsi, Maria Grazia Giansanti, and Maurizio Gatti. Spindle Self-organization and Cytokinesis During Male Meiosis in asterless Mutants of Drosophila melanogaster. Journal of Cell Biology, 142(3):751–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/3/751.

Bou-Gharios:1996:PFU

[BGGR^{+}96] G. Bou-Gharios, L. A. Garrett, J. Rossert, K. Niederreither, H. Eberspaecher, C. Smith, C. Black, and B. Crombrugghe. A potent far-upstream enhancer in the mouse pro alpha 2(I) collagen gene regulates expression of reporter genes in transgenic mice. Journal of Cell Biology, 134(5):1333–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/5/1333.

Becker:1997:PMH

[BGH^{+}97] K. David Becker, Kim R. Gottshall, Reed Hickey, Jean-Claude Perriard, and Kenneth R. Chien. Point mutations in human β cardiac myosin heavy chain have differential effects on sarcomeric structure and assembly: an ATP binding site change disrupts both thick and thin filaments, whereas hypertrophic cardiomyopathy mutations display normal assembly. Journal of Cell Biology, 137(1):131–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/1/131.

Burkin:1998:FRS

[BGH^{+}98] Dean J. Burkin, Maojian Gu, Bradley L. Hodges, James T. Campanelli, and Stephen J. Kaufman. A Functional Role for Specific Spliced Variants of the α7β1 Integrin in Acetylcholine Receptor Clustering. Journal of Cell Biology, 143(4): 1067–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/4/1067.

Brakenhoff:1995:HEA

[BGK^{+}95] R. H. Brakenhoff, M. Gerretsen, E. M. Knippels, M. van Dijk, H. van Essen, D. O. Weghuis, R. J. Sinke, G. B. Snow, and G. A. van Dongen. The human E48 antigen, highly homologous to the murine Ly-6 antigen ThB, is a GPI-anchored molecule apparently involved in keratinocyte cell–cell adhesion. Journal of Cell Biology, 129(6):1677–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/6/1677.

Blocker:1999:TTI

[BGL^{+}99] Ariel Blocker, Pierre Gounon, Eric Larquet, Kirsten Niebuhr, Véronique Cabiaux, Claude Parsot, and Philippe Sansonetti.

The Tripartite Type III Secreton of Shigella flexneri Inserts Ipab and Ipac into Host Membranes. Journal of Cell Biology, 147(3):683–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/3/683.

Brizzio:1996:CFD

[BGN^{+}96] V. Brizzio, A. E. Gammie, G. Nijbroek, S. Michaelis, and M. D. Rose. Cell fusion during yeast mating requires high levels of a-factor mating pheromone. Journal of Cell Biology, 135(6):1727–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1727.

Buczynski:1997:ITD

[BGN^{+}97] Greg Buczynski, Bryon Grove, Anson Nomura, Maurice Kleve, John Bush, Richard A. Firtel, and James Cardelli. Inactivation of Two Dictyostelium discoideum Genes, DdPIK1 and DdPIK2, Encoding Proteins Related to Mammalian Phosphatidylinositide 3-kinases, Results in Defects in Endocytosis, Lysosome to Postlysosome Transport, and Actin Cytoskeleton Organization. Journal of Cell Biology, 136(6):1271–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
6/1271.

Bornfeldt:1995:SPI

[BGR^{+}95] K. E. Bornfeldt, L. M. Graves, E. W. Raines, Y. Igarashi, G. Wayman, S. Yamamura, Y. Yatomi, J. S. Sidhu, E. G. Krebs, and S. Hakomori. Sphingosine-1-phosphate inhibits PDGF-induced chemotaxis of human arterial smooth muscle cells: spatial and temporal modulation of PDGF chemotactic signal transduction. Journal of Cell Biology, 130(1):193–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/1/193.

Brizzio:1998:RIF

[BGR98] Valeria Brizzio, Alison E. Gammie, and Mark D. Rose. Rvs161p Interacts with Fus2p to Promote Cell Fusion in Saccharomyces cerevisiae. Journal of Cell Biology, 141(3):567–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/3/567.

Borowsky:1998:LNT

[BH98] Mark L. Borowsky and Richard O. Hynes. Layilin, A Novel Talin-binding Transmembrane Protein Homologous with Ctype Lectins, is Localized in Membrane Ruffles. Journal of Cell Biology, 143(2):429–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/2/429.

Blake:1999:DDL

[BHBB99] Derek J. Blake, Richard Hawkes, Matthew A. Benson, and Phillip W. Beesley. Different Dystrophin-like Complexes Are Expressed in Neurons and Glia. Journal of Cell Biology, 147(3):645–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/3/645.

Blagoveshchenskaya:1999:CWS

[BHC99] Anastasiya D. Blagoveshchenskaya, Eric W. Hewitt, and Daniel F. Cutler. A Complex Web of Signal-dependent Trafficking Underlies the Triorganellar Distribution of P–Selectin in Neuroendocrine PC12 Cells. Journal of Cell Biology, 145 (7):1419–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/7/1419.

Boeuf:1997:LIF

[BHD^{+}97] Hélène Boeuf, Charlotte Hauss, Fabienne De Graeve, Nathalie Baran, and Claude Kedinger. Leukemia inhibitory factordependent transcriptional activation in embryonic stem cells. Journal of Cell Biology, 138(6):1207–??, September 1997. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/138/6/
1207.

Bednar:1995:CCS

[BHDW95] J. Bednar, R. A. Horowitz, J. Dubochet, and C. L. Woodcock. Chromatin conformation and salt-induced compaction: three-dimensional structural information from cryoelectron microscopy. Journal of Cell Biology, 131(6):1365–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1365.

Babcock:1997:MPI

[BHG^{+}97] Donner F. Babcock, James Herrington, Paul C. Goodwin, Young Bae Park, and Bertil Hille. Mitochondrial Participation in the Intracellular Ca^{2+} Network. Journal of Cell Biology, 136(4):833–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/4/833.

Beinhauer:1997:MFY

[BHHF97] Jens D. Beinhauer, Iain M. Hagan, Johannes H. Hegemann, and Ursula Fleig. Mal3, the Fission Yeast Homologue of the Human APC-interacting Protein EB-1 Is Required for Microtubule Integrity and the Maintenance of Cell Form. Journal of Cell Biology, 139(3):717–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/3/717.

Bernard:1998:FYB

[BHJ98] Pascal Bernard, Kevin Hardwick, and Jean-Paul Javerzat. Fission Yeast Bub1 Is a Mitotic Centromere Protein Essential for the Spindle Checkpoint and the Preservation of Correct Ploidy through Mitosis. Journal of Cell Biology, 143(7):1775–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/7/1775.

Burgoon:1995:FAP

[BHP^{+}95] M. P. Burgoon, R. B. Hazan, G. R. Phillips, K. L. Crossin, G. M. Edelman, and B. A. Cunningham. Functional analysis of posttranslational cleavage products of the neuron-glia cell adhesion molecule, Ng–CAM. Journal of Cell Biology, 130 (3):733–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/3/733.

Buchenau:1998:DPG

[BHSAJ98] Peter Buchenau, Jacob Hodgson, Helen Strutt, and Donna J. Arndt-Jovin. The Distribution of Polycomb–Group Proteins During Cell Division and Development in Drosophila Embryos: Impact on Models for Silencing. Journal of Cell Biology, 141 (2):469–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/2/469.

Bijlmakers:1997:ISU

[BINRM97] Marie-José J. E. Bijlmakers, Misako Isobe-Nakamura, Lindsay J. Ruddock, and Mark Marsh. Intrinsic Signals in the Unique Domain Target p56 ^{lck} to the Plasma Membrane Independently of CD4. Journal of Cell Biology, 137(5):1029–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/5/1029.

Biggins:1996:YUL

[BIR96] S. Biggins, I. Ivanovska, and M. D. Rose. Yeast ubiquitinlike genes are involved in duplication of the microtubule organizing center. Journal of Cell Biology, 133(6):1331–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
6/1331.

Bauren:1996:DDT

[BJB^{+}96] G. Baurén, W. Q. Jiang, K. Bernholm, F. Gu, and L. Wieslander. Demonstration of a dynamic, transcription-dependent organization of pre-mRNA splicing factors in polytene nuclei. Journal of Cell Biology, 133(5):929–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/5/929.

Bray:1999:CUC

[BJR^{+}99] Patrick G. Bray, Omar Janneh, Kaylene J. Raynes, Mathirut Mungthin, Hagai Ginsburg, and Stephen A. Ward. Cellular Uptake of Chloroquine Is Dependent on Binding to Ferriprotoporphyrin IX and Is Independent of NHE Activity in Plasmodium falciparum. Journal of Cell Biology, 145(2):363–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/363.

Bennett:1995:CIC

[BJS^{+}95] K. L. Bennett, D. G. Jackson, J. C. Simon, E. Tanczos, R. Peach, B. Modrell, I. Stamenkovic, G. Plowman, and A. Aruffo. CD44 isoforms containing exon V3 are responsible for the presentation of heparin-binding growth factor. Journal of Cell Biology, 128(4):687–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/4/687.

Bachant:1996:NSM

[BK96] J. B. Bachant and M. W. Klymkowsky. A nontetrameric species is the major soluble form of keratin in Xenopus oocytes and rabbit reticulocyte lysates. Journal of Cell Biology, 132 (1):153–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/1/153.

Buchstaller:1996:CAM

[BKB^{+}96] A. Buchstaller, S. Kunz, P. Berger, B. Kunz, U. Ziegler, C. Rader, and P. Sonderegger. Cell adhesion molecules Ng-CAM and axonin-1 form heterodimers in the neuronal membrane and cooperate in neurite outgrowth promotion. Journal of Cell Biology, 135(6):1593–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/6/1593.

Brandenberger:1996:NCL

[BKEC96] R. Brandenberger, R. A. Kammerer, J. Engel, and M. Chiquet. Native chick laminin-4 containing the beta 2 chain (s-laminin) promotes motor axon growth. Journal of Cell Biology, 135(6): 1583–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1583.

Buss:1998:LMV

[BKJL^{+}98] Folma Buss, John Kendrick-Jones, Corinne Lionne, Alex E. Knight, Graham P. Côté, and J. Paul Luzio. The Localization of Myosin VI at the Golgi Complex and Leading Edge of Fibroblasts and Its Phosphorylation and Recruitment into Membrane Ruffles of A431 Cells after Growth Factor Stimulation. Journal of Cell Biology, 143(6):1535–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1535.

Belliveau:1997:NNB

[BKK^{+}97] Daniel J. Belliveau, Irena Krivko, Judi Kohn, Christian Lachance, Christine Pozniak, Dmitri Rusakov, David Kaplan, and Freda D. Miller. NGF and Neurotrophin-3 Both Activate TrkA on Sympathetic Neurons but Differentially Regulate Survival and Neuritogenesis. Journal of Cell Biology, 136 (2):375–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/2/375.

Bobinnec:1998:CDV

[BKM^{+}98] Y. Bobinnec, A. Khodjakov, L. M. Mir, C. L. Rieder, B. Eddé, and M. Bornens. Centriole Disassembly In Vivo and Its Effect on Centrosome Structure and Function in Vertebrate Cells. Journal of Cell Biology, 143(6):1575–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1575.

Brady-Kalnay:1998:DIP

[BKMN^{+}98] Susann M. Brady-Kalnay, Tracy Mourton, Joseph P. Nixon, Gregory E. Pietz, Michael Kinch, Haiyan Chen, Robert Brackenbury, David L. Rimm, Robert L. Del Vecchio, and Nicholas K. Tonks. Dynamic Interaction of PTPμ with Multiple Cadherins In Vivo. Journal of Cell Biology, 141(1):287–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/1/287.

Borradori:1997:LBP

[BKN^{+}97] Luca Borradori, Peter J. Koch, Carien M. Niessen, Stefan Erkeland, Manuel R. van Leusden, and Arnoud Sonnenberg. The Localization of Bullous Pemphigoid Antigen 180 (BP180) in Hemidesmosomes Is Mediated by Its Cytoplasmic Domain and Seems to be Regulated by the β4 Integrin Subunit. Journal of Cell Biology, 136(6):1333–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/6/1333.

Bikfalvi:1995:DMC

[BKP^{+}95] A. Bikfalvi, S. Klein, G. Pintucci, N. Quarto, P. Mignatti, and D. B. Rifkin. Differential modulation of cell phenotype by different molecular weight forms of basic fibroblast growth factor: possible intracellular signaling by the high molecular weight forms. Journal of Cell Biology, 129(1):233–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
1/233.

Brady-Kalnay:1995:RPT

[BKRT95] S. M. Brady-Kalnay, D. L. Rimm, and N. K. Tonks. Receptor protein tyrosine phosphatase PTPmu associates with cadherins and catenins in vivo. Journal of Cell Biology, 130 (4):977–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/977.

Bagai:1996:TCT

[BL96] S. Bagai and R. A. Lamb. Truncation of the COOH-terminal region of the paramyxovirus SV5 fusion protein leads to hemifusion but not complete fusion. Journal of Cell Biology, 135 (1):73–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/73.

Bogan:1999:TCI

[BL99] Jonathan S. Bogan and Harvey F. Lodish. Two Compartments for Insulin–Stimulated Exocytosis in 3t3-L1 Adipocytes Defined by Endogenous Acrp30 and Glut4. Journal of Cell Biology, 146(3):609–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/3/609.

Block:1996:PEC

[BLB^{+}96] G. D. Block, J. Locker, W. C. Bowen, B. E. Petersen, S. Katyal, S. C. Strom, T. Riley, T. A. Howard, and G. K. Michalopoulos. Population expansion, clonal growth, and specific differentiation patterns in primary cultures of hepatocytes induced by HGF/SF, EGF and TGF alpha in a chemically defined (HGM) medium. Journal of Cell Biology, 132 (6):1133–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/6/1133.

Benmerah:1998:AEI

[BLB^{+}98] Alexandre Benmerah, Christophe Lamaze, Bernadette Bègue, Sandra L. Schmid, Alice Dautry-Varsat, and Nadine Cerf-Bensussan. AP-2/ eps15 Interaction Is Required for Receptormediated Endocytosis. Journal of Cell Biology, 140(5): 1055–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/5/1055.

Benghezal:1995:ISC

[BLC95] M. Benghezal, P. N. Lipke, and A. Conzelmann. Identification of six complementation classes involved in the biosynthesis of glycosylphosphatidylinositol anchors in Saccharomyces cerevisiae. Journal of Cell Biology, 130(6):1333–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1333.

Balsamo:1996:RBP

[BLE^{+}96] J. Balsamo, T. Leung, H. Ernst, M. K. Zanin, S. Hoffman, and J. Lilien. Regulated binding of PTP1B-like phosphatase to N-cadherin: control of cadherin-mediated adhesion by dephosphorylation of beta-catenin. Journal of Cell Biology, 134 (3):801–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/3/801.

Bastos:1996:TFM

[BLEB96] R. Bastos, A. Lin, M. Enarson, and B. Burke. Targeting and function in mRNA export of nuclear pore complex protein Nup153. Journal of Cell Biology, 134(5):1141–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
5/1141.

Bellini:1995:ZBD

[BLG95] M. Bellini, J. C. Lacroix, and J. G. Gall. A zinc-binding domain is required for targeting the maternal nuclear protein PwA33 to lampbrush chromosome loops. Journal of Cell Biology, 131(3):563–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/3/563.

Brandt:1995:ITN

[BLL95] R. Brandt, J. Léger, and G. Lee. Interaction of tau with the neural plasma membrane mediated by tau’s amino-terminal projection domain. Journal of Cell Biology, 131(5):1327– ??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/5/1327.

Bender:1996:AAP

[BLL^{+}96] L. Bender, H. S. Lo, H. Lee, V. Kokojan, V. Peterson, and A. Bender. Associations among PH and SH3 domaincontaining proteins and Rho-type GTPases in Yeast. Journal of Cell Biology, 133(4):879–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/4/879.

Blystone:1995:IBC

[BLLB95] S. D. Blystone, F. P. Lindberg, S. E. LaFlamme, and E. J. Brown. Integrin beta 3 cytoplasmic tail is necessary and sufficient for regulation of alpha 5 beta 1 phagocytosis by alpha v beta 3 and integrin-associated protein. Journal of Cell Biology, 130(3):745–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/3/745.

Block:1998:LPN

[Blo98] Steven M. Block. Leading the Procession: New Insights into Kinesin Motors. Journal of Cell Biology, 140(6):1281–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/6/1281.

Brancolini:1997:DCC

[BLRS97] Claudio Brancolini, Dean Lazarevic, Joe Rodriguez, and Claudio Schneider. Dismantling Cell–Cell Contacts during Apoptosis Is Coupled to a Caspase-dependent Proteolytic Cleavage of β-Catenin. Journal of Cell Biology, 139(3):759–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/759.

Bush:1996:RAY

[BM96] G. L. Bush and D. I. Meyer. The refolding activity of the yeast heat shock proteins Ssa1 and Ssa2 defines their role in protein translocation. Journal of Cell Biology, 135(5):1229–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/5/1229.

Bassnett:1997:CDD

[BM97] Steven Bassnett and Danijela Mataic. Chromatin Degradation in Differentiating Fiber Cells of the Eye Lens. Journal of Cell Biology, 137(1):37–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/1/37.

Bruick:1998:PPU

[BM98] Richard K. Bruick and Stephen P. Mayfield. Processing of the psbA 5’ Untranslated Region in Chlamydomonas reinhardtii Depends upon Factors Mediating Ribosome Association. Journal of Cell Biology, 143(5):1145–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1145.

Bijlmakers:1999:TAC

[BM99] Marie-José J. E. Bijlmakers and Mark Marsh. Trafficking of an Acylated Cytosolic Protein: Newly Synthesized p56^{lck} Travels to the Plasma Membrane via the Exocytic Pathway. Journal of Cell Biology, 145(3):457–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/3/457.

Barnard:1997:SNA

[BMB97] Richard J. O. Barnard, Alan Morgan, and Robert D. Burgoyne. Stimulation of NSF ATPase Activity by α-SNAP Is Required for SNARE Complex Disassembly and Exocytosis. Journal of Cell Biology, 139(4):875–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
4/875.

Bailly:1999:RBA

[BMC^{+}99] Maryse Bailly, Frank Macaluso, Michael Cammer, Amanda Chan, Jeffrey E. Segall, and John S. Condeelis. Relationship between Arp2/3 complex and the barbed ends of actin filaments at the leading edge of carcinoma cells after epidermal growth factor stimulation. Journal of Cell Biology, 145 (2):331–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/2/331.

Bennett:1995:RCB

[BMG^{+}95] K. L. Bennett, B. Modrell, B. Greenfield, A. Bartolazzi, I. Stamenkovic, R. Peach, D. G. Jackson, F. Spring, and A. Aruffo. Regulation of CD44 binding to hyaluronan by glycosylation of variably spliced exons. Journal of Cell Biology, 131(6): 1623–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1623.

Beatty:1999:TSL

[BMG^{+}99a] Wandy L. Beatty, Stéphane Méresse, Pierre Gounon, Jean Davoust, Joëlle Mounier, Philippe J. Sansonetti, and Jean-Pierre Gorvel. Trafficking of Shigella Lipopolysaccharide in Polarized Intestinal Epithelial Cells. Journal of Cell Biology, 145(4):689–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/689.

Brunet:1999:KFI

[BMG^{+}99b] Stéphane Brunet, Angélica Santa Maria, Philippe Guillaud, Denis Dujardin, Jacek Z. Kubiak, and Bernard Maro. Kinetochore fibers are not involved in the formation of the first meiotic spindle in mouse oocytes, but control the exit from the first meiotic M phase. Journal of Cell Biology, 146(1):1–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
146/1/1.

Braga:1997:SGR

[BMHH97] Vania M. M. Braga, Laura M. Machesky, Alan Hall, and Neil A. Hotchin. The Small GTPases Rho and Rac Are Required for the Establishment of Cadherin-dependent Cell– Cell Contacts. Journal of Cell Biology, 137(6):1421–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
6/1421.

Bi:1997:KMD

[BML^{+}97] Guo-Qiang Bi, Robert L. Morris, Guochun Liao, Janet M. Alderton, Jonathan M. Scholey, and Richard A. Steinhardt. Kinesin- and myosin-driven steps of vesicle recruitment for Ca^{2+}-regulated exocytosis. Journal of Cell Biology, 138(5): 999–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/5/999.

Bi:1998:IAC

[BML^{+}98] Erfei Bi, Paul Maddox, Daniel J. Lew, E. D. Salmon, John N. McMillan, Elaine Yeh, and John R. Pringle. Involvement of an Actomyosin Contractile Ring in Saccharomyces cerevisiae Cytokinesis. Journal of Cell Biology, 142(5):1301–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
5/1301.

Brock:1996:HIW

[BMLM96] J. Brock, K. Midwinter, J. Lewis, and P. Martin. Healing of incisional wounds in the embryonic chick wing bud: characterization of the actin purse-string and demonstration of a requirement for Rho activation. Journal of Cell Biology, 135(4): 1097–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/1097.

Bamji:1998:PNR

[BMP^{+}98] Shernaz X. Bamji, Marta Majdan, Christine D. Pozniak, Daniel J. Belliveau, Raquel Aloyz, Judi Kohn, Carrie G. Causing, and Freda D. Miller. The p75 Neurotrophin Receptor Mediates Neuronal Apoptosis and Is Essential for Naturally Occurring Sympathetic Neuron Death. Journal of Cell Biology, 140(4):911–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/4/911.

Beauchamp:1999:DMT

[BMPP99] Jonathan R. Beauchamp, Jennifer E. Morgan, Charles N. Pagel, and Terence A. Partridge. Dynamics of myoblast transplantation reveal a discrete minority of precursors with stem cell-like properties as the myogenic source. Journal of Cell Biology, 144(6):1113–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/6/1113.

Berthiaume:1995:MSF

[BMS95] E. P. Berthiaume, C. Medina, and J. A. Swanson. Molecular size-fractionation during endocytosis in macrophages. Journal of Cell Biology, 129(4):989–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/4/989.

Bass:1997:TCN

[BMS^{+}97] Hank W. Bass, Wallace F. Marshall, John W. Sedat, David A. Agard, and W. Zacheus Cande. Telomeres Cluster De Novo before the Initiation of Synapsis: a Three-dimensional Spatial Analysis of Telomere Positions before and during Meiotic Prophase. Journal of Cell Biology, 137(1):5–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/137/1/5.

Bartolazzi:1996:GCI

[BNA^{+}96] A. Bartolazzi, A. Nocks, A. Aruffo, F. Spring, and I. Stamenkovic. Glycosylation of CD44 is implicated in CD44mediated cell adhesion to hyaluronan. Journal of Cell Biology, 132(6):1199–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/6/1199.

Beaulieu:1999:LOD

[BNJ99] Jean-Martin Beaulieu, Minh Dang Nguyen, and Jean-Pierre Julien. Late Onset Death of Motor Neurons in Mice Overexpressing Wild–Type Peripherin. Journal of Cell Biology, 147(3):531–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/3/531.

Banerji:1999:LNH

[BNW^{+}99] Suneale Banerji, Jian Ni, Shu-Xia Wang, Steven Clasper, Jeffrey Su, Raija Tammi, Margaret Jones, and David G. Jackson. LYVE-1, a New Homologue of the CD44 Glycoprotein, Is a Lymph-specific Receptor for Hyaluronan. Journal of Cell Biology, 144(4):789–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/4/789.

Berditchevski:1999:CIT

[BO99] Fedor Berditchevski and Elena Odintsova. Characterization of Integrin–Tetraspanin Adhesion Complexes. Journal of Cell Biology, 146(2):477–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/2/477.

Bastmeyer:1995:FEG

[BOLS95] M. Bastmeyer, H. Ott, C. A. Leppert, and C. A. Stuermer. Fish E587 glycoprotein, a member of the L1 family of cell adhesion molecules, participates in axonal fasciculation and the age-related order of ganglion cell axons in the goldfish retina. Journal of Cell Biology, 130(4):969–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/4/969.

Bourdoulous:1998:FMR

[BOM^{+}98] Sandrine Bourdoulous, Gertraud Orend, Deidre A. MacKenna, Renata Pasqualini, and Erkki Ruoslahti. Fibronectin Matrix Regulates Activation of RHO and CDC42 GTPases and Cell Cycle Progression. Journal of Cell Biology, 143(1):267–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/1/267.

Bornstein:1995:DFI

[Bor95] P. Bornstein. Diversity of function is inherent in matricellular proteins: an appraisal of thrombospondin 1. Journal of Cell Biology, 130(3):503–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/3/503.

Baba:1997:TDP

[BOS^{+}97] Misuzu Baba, Masako Osumi, Sidney V. Scott, Daniel J. Klionsky, and Yoshinori Ohsumi. Two distinct pathways for targeting proteins from the cytoplasm to the vacuole/ lysosome. Journal of Cell Biology, 139(7):1687–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
7/1687.

Byzova:1998:AVC

[BP98] Tatiana V. Byzova and Edward F. Plow. Activation of α_{V} β_{3} on Vascular Cells Controls Recognition of Prothrombin. Journal of Cell Biology, 143(7):2081–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/7/2081.

Barth:1997:NTD

[BPA^{+}97] Angela I. M. Barth, Anne L. Pollack, Yoram Altschuler, Keith E. Mostov, and W. James Nelson. NH_{2}-terminal deletion of β-catenin results in stable colocalization of mutant β-catenin with adenomatous polyposis coli protein and altered MDCK cell adhesion. Journal of Cell Biology, 136(3): 693–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/693.

Bacher:1999:RRG

[BPD99] Gerald Bacher, Martin Pool, and Bernhard Dobberstein. The Ribosome Regulates the Gtpase of the β-Subunit of the Signal Recognition Particle Receptor. Journal of Cell Biology, 146 (4):723–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/723.

Bowman:1999:DRC

[BPKB^{+}99] Aaron B. Bowman, Ramila S. Patel-King, Sharon E. Benashski, J. Michael McCaffery, Lawrence S. B. Goldstein, and Stephen M. King. Drosophila roadblock and Chlamydomonas Lc7. Journal of Cell Biology, 146(1):165–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
1/165.

Belien:1999:MTM

[BPS99] Ann T. J. Beliën, Paolo A. Paganetti, and Martin E. Schwab. Membrane-type 1 Matrix Metalloprotease (MT1-MMP) Enables Invasive Migration of Glioma Cells in Central Nervous System White Matter. Journal of Cell Biology, 144(2): 373–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/2/373.

Brown:1996:ILP

[BPT96] M. C. Brown, J. A. Perrotta, and C. E. Turner. Identification of LIM3 as the principal determinant of paxillin focal adhesion localization and characterization of a novel motif on paxillin directing vinculin and focal adhesion kinase binding. Journal of Cell Biology, 135(4):1109–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/1109.

Bryant:1998:RTY

[BPWS98] Nia J. Bryant, Robert C. Piper, Lois S. Weisman, and Tom H. Stevens. Retrograde traffic out of the yeast vacuole to the TGN occurs via the prevacuolar/endosomal compartment. Journal of Cell Biology, 142(3):651–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/3/651.

Bo:1995:EDM

[BQF^{+}95] L. Bö, R. H. Quarles, N. Fujita, Z. Bartoszewicz, S. Sato, and B. D. Trapp. Endocytic depletion of L–MAG from CNS myelin in quaking mice. Journal of Cell Biology, 131(6):1811–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1811.

Bruses:1998:RNC

[BR98] Juan L. Brusés and Urs Rutishauser. Regulation of Neural Cell Adhesion Molecule Polysialylation: Evidence for Nontranscriptional Control and Sensitivity to an Intracellular Pool of Calcium. Journal of Cell Biology, 140(5):1177–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
5/1177.

Barres:1999:ACO

[BR99] Ben A. Barres and Martin C. Raff. Axonal Control of Oligodendrocyte Development. Journal of Cell Biology, 147(6): 1123–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/6/1123.

Brachet:1997:ICC

[BRAM97] Valérie Brachet, Graça Raposo, Sebastian Amigorena, and Ira Mellman. Ii Chain Controls the Transport of Major Histocompatibility Complex Class II Molecules to and from Lysosomes. Journal of Cell Biology, 137(1):51–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/137/1/51.

Bannykh:1996:OER

[BRB96] S. I. Bannykh, T. Rowe, and W. E. Balch. The organization of endoplasmic reticulum export complexes. Journal of Cell Biology, 135(1):19–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/1/19.

Boitier:1999:MEN

[BRD99] Eric Boitier, Ruth Rea, and Michael R. Duchen. Mitochondria Exert a Negative Feedback on the Propagation of Intracellular Ca^{2+} Waves in Rat Cortical Astrocytes. Journal of Cell Biology, 145(4):795–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/4/795.

Burns:1995:ELM

[BRHD95] C. G. Burns, M. Reedy, J. Heuser, and A. De Lozanne. Expression of light meromyosin in Dictyostelium blocks normal myosin II function. Journal of Cell Biology, 130(3):605–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/3/605.

Bridgman:1999:MVM

[Bri99] P. C. Bridgman. Myosin VA Movements in Normal and Dilute– Lethal Axons Provide Support for a Dual Filament Motor Complex. Journal of Cell Biology, 146(5):1045–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
5/1045.

Bridge:1996:SEA

[BRJP96] E. Bridge, K. U. Riedel, B. M. Johansson, and U. Pettersson. Spliced exons of adenovirus late RNAs colocalize with snRNP in a specific nuclear domain. Journal of Cell Biology, 135(2): 303–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/2/303.

Beckwith:1998:KCD

[BRLM98] Susan M. Beckwith, Christian H. Roghi, Bo Liu, and N. Ronald Morris. The “8-kD” Cytoplasmic Dynein Light Chain Is Required for Nuclear Migration and for Dynein Heavy Chain Localization in Aspergillus nidulans. Journal of Cell Biology, 143(5):1239–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/5/1239.

Bachelder:1999:PII

[BRM^{+}99] Robin E. Bachelder, Mark J. Ribick, Alessandra Marchetti, Rita Falcioni, Silvia Soddu, Kathryn R. Davis, and Arthur M. Mercurio. P53 inhibits α6β4 integrin survival signaling by promoting the caspase 3-dependent cleavage of Akt/PKB. Journal of Cell Biology, 147(5):1063–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/5/1063.

Belkin:1997:MIR

[BRP^{+}97] Alexey M. Belkin, S. Francesco Retta, Olga Y. Pletjushkina, Fiorella Balzac, Lorenzo Silengo, Reinhard Fassler, Victor E. Koteliansky, Keith Burridge, and Guido Tarone. Muscle β1D Integrin Reinforces the Cytoskeleton–Matrix Link: Modulation of Integrin Adhesive Function by Alternative Splicing. Journal of Cell Biology, 139(6):1583–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
6/1583.

Bear:1998:SWR

[BRS98] James E. Bear, John F. Rawls, and Charles L. Saxe. SCAR, a WASP-related Protein, Isolated as a Suppressor of Receptor Defects in Late Dictyostelium Development. Journal of Cell Biology, 142(5):1325–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/5/1325.

Benting:1999:GMA

[BRS99] Jürgen H. Benting, Anton G. Rietveld, and Kai Simons. Nglycans Mediate the Apical Sorting of a Gpi–Anchored, Raft– Associated Protein in Madin–Darby Canine Kidney Cells. Journal of Cell Biology, 146(2):313–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/2/313.

Bottger:1995:PNP

[BS95a] A. Böttger and B. A. Spruce. Proenkephalin is a nuclear protein responsive to growth arrest and differentiation signals. Journal of Cell Biology, 130(6):1251–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1251.

Buss:1995:MIN

[BS95b] F. Buss and M. Stewart. Macromolecular interactions in the nucleoporin p62 complex of rat nuclear pores: binding of nucleoporin p54 to the rod domain of p62. Journal of Cell Biology, 128(3):251–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/3/251.

Bryant:1997:TSS

[BS97] Nia J. Bryant and Tom H. Stevens. Two Separate Signals Act Independently to Localize a Yeast Late Golgi Membrane Protein through a Combination of Retrieval and Retention. Journal of Cell Biology, 136(2):287–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/2/287.

Buchenau:1997:DNR

[BSAJ97] Peter Buchenau, Harald Saumweber, and Donna J. Arndt-Jovin. The dynamic nuclear redistribution of an hnRNP Khomologous protein during Drosophila embryo development and heat shock. Flexibility of transcription sites in vivo. Journal of Cell Biology, 137(2):291–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/291.

Blocker:1997:MRB

[BSB^{+}97] Ariel Blocker, Fedor F. Severin, Janis K. Burkhardt, James B. Bingham, Hanry Yu, Jean-Christophe Olivo, Trina A. Schroer, Anthony A. Hyman, and Gareth Griffiths. Molecular Requirements for Bi-directional Movement of Phagosomes Along Microtubules. Journal of Cell Biology, 137(1):113–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
1/113.

Blumenthal:1996:DIH

[BSD^{+}96] R. Blumenthal, D. P. Sarkar, S. Durell, D. E. Howard, and S. J. Morris. Dilation of the influenza hemagglutinin fusion pore revealed by the kinetics of individual cell–cell fusion events. Journal of Cell Biology, 135(1):63–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/135/1/63.

Bascom-Slack:1997:YMP

[BSD97] Carol A. Bascom-Slack and Dean S. Dawson. The Yeast Motor Protein, Kar3p, Is Essential for Meiosis I. Journal of Cell Biology, 139(2):459–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/2/459.

Beck:1999:SIV

[BSH99] Thomas Beck, Anja Schmidt, and Michael N. Hall. Starvation Induces Vacuolar Targeting and Degradation of the Tryptophan Permease in Yeast. Journal of Cell Biology, 146(6): 1227–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1227.

Briggs:1995:CNC

[BSJJ95] K. K. Briggs, A. J. Silvers, K. M. Johansen, and J. Johansen. Calsensin: a novel calcium-binding protein expressed in a subset of peripheral leech neurons fasciculating in a single axon tract. Journal of Cell Biology, 129(5):1355–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1355.

Blessing:1996:OBM

[BSK96] M. Blessing, P. Schirmacher, and S. Kaiser. Overexpression of bone morphogenetic protein-6 (BMP-6) in the epidermis of transgenic mice: inhibition or stimulation of proliferation depending on the pattern of transgene expression and formation of psoriatic lesions. Journal of Cell Biology, 135(1):227–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/1/227.

Buchert:1999:JAP

[BSM^{+}99] Michael Buchert, Stefan Schneider, Virginia Meskenaite, Mark T. Adams, Eli Canaani, Thomas Baechi, Karin Moelling, and Christopher M. Hovens. The Junction-associated Protein AF-6 Interacts and Clusters with Specific Eph Receptor Tyrosine Kinases at Specialized Sites of Cell–Cell Contact in the Brain. Journal of Cell Biology, 144(2):361–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
2/361.

Byeon:1995:NPH

[BSMH95] M. K. Byeon, Y. Sugi, R. R. Markwald, and S. Hoffman. NCAM polypeptides in heart development: association with Z discs of forms that contain the muscle-specific domain. Journal of Cell Biology, 128(1):209–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/1/209.

Benlimame:1995:AMF

[BSN95] N. Benlimame, D. Simard, and I. R. Nabi. Autocrine motility factor receptor is a marker for a distinct membranous tubular organelle. Journal of Cell Biology, 129(2):459–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
2/459.

Bauer:1999:ANT

[BSRG99] Manuel K. A. Bauer, Alexis Schubert, Oliver Rocks, and Stefan Grimm. Adenine Nucleotide Translocase-1, a Component of the Permeability Transition Pore, Can Dominantly Induce Apoptosis. Journal of Cell Biology, 147(7):1493–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1493.

Bashkirov:1997:MCE

[BSS^{+}97] Vladimir I. Bashkirov, Harry Scherthan, Jachen A. Solinger, Jean-Marie Buerstedde, and Wolf-Dietrich Heyer. A Mouse Cytoplasmic Exoribonuclease (mXRN1p) with Preference for G4 Tetraplex Substrates. Journal of Cell Biology, 136(4): 761–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/4/761.

Bahler:1998:RPK

[BSW^{+}98] Jürg Bähler, Alexander B. Steever, Sally Wheatley, Yu li Wang, John R. Pringle, Kathleen L. Gould, and Dannel McCollum. Role of Polo Kinase and Mid1p in Determining the Site of Cell Division in Fission Yeast. Journal of Cell Biology, 143(6): 1603–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1603.

Blystone:1999:MMI

[BSW^{+}99] Scott D. Blystone, Suzanne E. Slater, Matthew P. Williams, Michael T. Crow, and Eric J. Brown. A molecular mechanism of integrin crosstalk: α_{v}β_{3} suppression of calcium/calmodulindependent protein kinase II regulates α_{5}β_{1} function. Journal of Cell Biology, 145(4):889–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/4/889.

Berger:1997:MCR

[BSY97] Karen H. Berger, L. Farah Sogo, and Michael P. Yaffe. Mdm12p, a Component Required for Mitochondrial Inheritance That Is Conserved between Budding and Fission Yeast. Journal of Cell Biology, 136(3):545–??, February 1997. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/3/
545.

Bruckner-Tuderman:1995:IMA

[BTNZ^{+}95] L. Bruckner-Tuderman, O. Nilssen, D. R. Zimmermann, M. T. Dours-Zimmermann, D. U. Kalinke, T. Gedde-Dahl, and J. O. Winberg. Immunohistochemical and mutation analyses demonstrate that procollagen VII is processed to collagen VII through removal of the NC-2 domain. Journal of Cell Biology, 131(2):551–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/2/551.

Benian:1996:CEG

[BTTB96] G. M. Benian, T. L. Tinley, X. Tang, and M. Borodovsky. The Caenorhabditis elegans gene unc-89, required fpr muscle Mline assembly, encodes a giant modular protein composed of Ig and signal transduction domains. Journal of Cell Biology, 132 (5):835–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/835.

Burger:1996:TSG

[BvdBvM96] K. N. Burger, P. van der Bijl, and G. van Meer. Topology of sphingolipid galactosyltransferases in ER and Golgi: transbilayer movement of monohexosyl sphingolipids is required for higher glycosphingolipid biosynthesis. Journal of Cell Biology, 133(1):15–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/1/15.

Boldogh:1998:IBM

[BVKP98] Istvan Boldogh, Nikola Vojtov, Sharon Karmon, and Liza A. Pon. Interaction between Mitochondria and the Actin Cytoskeleton in Budding Yeast Requires Two Integral Mitochondrial Outer Membrane Proteins, Mmm1p and Mdm10p. Journal of Cell Biology, 141(6):1371–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/6/1371.

Bucci:1997:VDN

[BW97] Mirella Bucci and Susan R. Wente. In Vivo Dynamics of Nuclear Pore Complexes in Yeast. Journal of Cell Biology, 136 (6):1185–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/6/1185.

Balda:1996:FDP

[BWF^{+}96] M. S. Balda, J. A. Whitney, C. Flores, S. González, M. Cereijido, and K. Matter. Functional dissociation of paracellular permeability and transepithelial electrical resistance and disruption of the apical-basolateral intramembrane diffusion barrier by expression of a mutant tight junction membrane protein. Journal of Cell Biology, 134(4):1031–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
4/1031.

Barth:1995:ORS

[BWG95] B. U. Barth, J. M. Wahlberg, and H. Garoff. The oligomerization reaction of the Semliki Forest virus membrane protein subunits. Journal of Cell Biology, 128(3):283–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
3/283.

Barkalow:1996:CRP

[BWKH96] K. Barkalow, W. Witke, D. J. Kwiatkowski, and J. H. Hartwig. Coordinated regulation of platelet actin filament barbed ends by gelsolin and capping protein. Journal of Cell Biology, 134 (2):389–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/389.

Burgess:1997:FWM

[BWWT97] S. A. Burgess, M. L. Walker, H. D. White, and J. Trinick. Flexibility within Myosin Heads Revealed by Negative Stain and Single–Particle Analysis. Journal of Cell Biology, 139(3): 675–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/3/675.

Burgess:1999:LOO

[BxQK^{+}99] Teresa L. Burgess, Yi xin Qian, Stephen Kaufman, Brian D. Ring, Gwyneth Van, Charles Capparelli, Michael Kelley, Hailing Hsu, William J. Boyle, Colin R. Dunstan, Sylvia Hu, and David L. Lacey. The Ligand for Osteoprotegerin (OPGL) Directly Activates Mature Osteoclasts. Journal of Cell Biology, 145(3):527–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/3/527.

Brieher:1996:LDR

[BYG96] W. M. Brieher, A. S. Yap, and B. M. Gumbiner. Lateral dimerization is required for the homophilic binding activity of C-cadherin. Journal of Cell Biology, 135(2):487–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
2/487.

Belkin:1996:BID

[BZB^{+}96] A. M. Belkin, N. I. Zhidkova, F. Balzac, F. Altruda, D. Tomatis, A. Maier, G. Tarone, V. E. Koteliansky, and K. Burridge. Beta 1D integrin displaces the beta 1A isoform in striated muscles: localization at junctional structures and signaling potential in nonmuscle cells. Journal of Cell Biology, 132(1): 211–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/1/211.

Bottazzi:1999:RPE

[BZBA99] Maria Elena Bottazzi, Xiaoyun Zhu, Ralph M. Böhmer, and Richard K. Assoian. Regulation of P21^{cip1} Expression by Growth Factors and the Extracellular Matrix Reveals a Role for Transient ERK Activity in G1 Phase. Journal of Cell Biology, 146(6):1255–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/6/1255.

Bartles:1998:SET

[BZL^{+}98] James R. Bartles, Lili Zheng, Anli Li, Allison Wierda, and Bin Chen. Small espin: a third actin-bundling protein and potential forked protein ortholog in brush border microvilli. Journal of Cell Biology, 143(1):107–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/1/107.

Chi:1995:SCC

[CAA95] N. C. Chi, E. J. Adam, and S. A. Adam. Sequence and characterization of cytoplasmic nuclear protein import factor p97. Journal of Cell Biology, 130(2):265–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/2/265.

Carvalho:1999:SMA

[CAC^{+}99] Teresa Carvalho, Fátima Almeida, Alexandre Calapez, Miguel Lafarga, Maria T. Berciano, and Maria Carmo-Fonseca. The Spinal Muscular Atrophy Disease Gene Product, Smn. Journal of Cell Biology, 147(4):715–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/4/715.

Chen:1999:EDS

[CAH^{+}99] M. S. Chen, E. A. C. Almeida, A.-P. J. Huovila, Y. Takahashi, L. M. Shaw, A. M. Mercurio, and J. M. White. Evidence that Distinct States of the Integrin α6β1 Interact with Laminin and an ADAM. Journal of Cell Biology, 144(3):549–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
3/549.

Carlevaro:1997:TPE

[CAR^{+}97] Mariella F. Carlevaro, Adriana Albini, Domenico Ribatti, Chiara Gentili, Roberto Benelli, Silvia Cermelli, Ranieri Cancedda, and Fiorella Descalzi Cancedda. Transferrin Promotes Endothelial Cell Migration and Invasion: Implication in Cartilage Neovascularization. Journal of Cell Biology, 136(6): 1375–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/6/1375.

Carpenter:1999:EEG

[Car99] Graham Carpenter. Employment of the Epidermal Growth Factor Receptor in Growth Factor–Independent Signaling Pathways. Journal of Cell Biology, 146(4):697–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
4/697.

Clark:1996:CHD

[CAS96] R. A. Clark, R. Alon, and T. A. Springer. CD44 and hyaluronan-dependent rolling interactions of lymphocytes on tonsillar stroma. Journal of Cell Biology, 134(4):1075–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/4/1075.

Caroni:1997:IND

[CAS97] Pico Caroni, Ludwig Aigner, and Corinna Schneider. Intrinsic Neuronal Determinants Locally Regulate Extrasynaptic and Synaptic Growth at the Adult Neuromuscular Junction. Journal of Cell Biology, 136(3):679–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/3/679.

Carroll:1998:SRM

[CAS^{+}98] Christopher W. Carroll, Roger Altman, David Schieltz, John R. Yates, and Douglas Kellogg. The Septins Are Required for the Mitosis-specific Activation of the Gin4 Kinase. Journal of Cell Biology, 143(3):709–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/3/709.

Chi:1996:RSI

[CAVA96] N. C. Chi, E. J. Adam, G. D. Visser, and S. A. Adam. RanBP1 stabilizes the interaction of Ran with p97 nuclear protein import. Journal of Cell Biology, 135(3):559–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
3/559.

Castets:1996:NCB

[CBB^{+}96] F. Castets, M. Bartoli, J. V. Barnier, G. Baillat, P. Salin, A. Moqrich, J. P. Bourgeois, F. Denizot, G. Rougon, G. Calothy, and A. Monneron. A novel calmodulin-binding protein, belonging to the WD-repeat family, is localized in dendrites of a subset of CNS neurons. Journal of Cell Biology, 134(4):1051–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/4/1051.

Croft:1999:DLM

[CBB^{+}99] Jenny A. Croft, Joanna M. Bridger, Shelagh Boyle, Paul Perry, Peter Teague, and Wendy A. Bickmore. Differences in the Localization and Morphology of Chromosomes in the Human Nucleus. Journal of Cell Biology, 145(6):1119–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
6/1119.

Campbell:1998:CKS

[CBM^{+}98] James J. Campbell, Edward P. Bowman, Kristine Murphy, Kenneth R. Youngman, Michael A. Siani, Darren A. Thompson, Lijun Wu, Albert Zlotnik, and Eugene C. Butcher. 6-c-kine (SLC), a lymphocyte adhesion-triggering chemokine expressed by high endothelium, is an agonist for the MIP-3β receptor CCR7. Journal of Cell Biology, 141(4):1053–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
4/1053.

Cash:1997:RAR

[CBS^{+}97] David E. Cash, Cheryl B. Bock, Klaus Schughart, Elwood Linney, and T. Michael Underhill. Retinoic Acid Receptor α Function in Vertebrate Limb Skeletogenesis: a Modulator of Chondrogenesis. Journal of Cell Biology, 136(2):445–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
2/445.

Coorssen:1998:BFS

[CBTZ98] Jens R. Coorssen, Paul S. Blank, Masahiro Tahara, and Joshua Zimmerberg. Biochemical and Functional Studies of Cortical Vesicle Fusion: The SNARE Complex and Ca^{2+} Sensitivity. Journal of Cell Biology, 143(7):1845–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1845.

Chiang:1998:VNP

[CC98] Meng-Chieh Chiang and Hui-Ling Chiang. Vid24p, a Novel Protein Localized to the Fructose-1,6-bisphosphatase– containing Vesicles, Regulates Targeting of Fructose-1,6-bisphosphatase from the Vesicles to the Vacuole for Degradation. Journal of Cell Biology, 140(6):1347–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1347.

Chalupowicz:1995:FII

[CCB^{+}95] D. G. Chalupowicz, Z. A. Chowdhury, T. L. Bach, C. Barsigian, and J. Martinez. Fibrin II induces endothelial cell capillary tube formation. Journal of Cell Biology, 130(1):207–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/1/207.

Cirulli:1998:KAE

[CCB^{+}98] V. Cirulli, L. Crisa, G. M. Beattie, M. I. Mally, A. D. Lopez, A. Fannon, A. Ptasznik, L. Inverardi, C. Ricordi, T. Deerinck, M. Ellisman, R. A. Reisfeld, and A. Hayek. KSA Antigen Ep–CAM Mediates Cell–Cell Adhesion of Pancreatic Epithelial Cells: Morphoregulatory Roles in Pancreatic Islet Development. Journal of Cell Biology, 140(6):1519–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1519.

Cleves:1996:NPP

[CCBK96] A. E. Cleves, D. N. Cooper, S. H. Barondes, and R. B. Kelly. A new pathway for protein export in Saccharomyces cerevisiae. Journal of Cell Biology, 133(5):1017–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/5/1017.

Clemson:1998:SLX

[CCBL98] Christine Moulton Clemson, Jennifer C. Chow, Carolyn J. Brown, and Jeanne Bentley Lawrence. Stabilization and Localization of Xist RNA are Controlled by Separate Mechanisms and are Not Sufficient for X Inactivation. Journal of Cell Biology, 142(1):13–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/1/13.

Correia:1999:IAV

[CCC^{+}99] Ivan Correia, Donald Chu, Ying-Hao Chou, Robert D. Goldman, and Paul Matsudaira. Integrating the Actin and Vimentin Cytoskeletons. Journal of Cell Biology, 146(4): 831–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/831.

Chiba:1997:SRF

[CCMC97] Hideki Chiba, John Clifford, Daniel Metzger, and Pierre Chambon. Specific and redundant functions of retinoid X receptor/retinoic acid receptor heterodimers in differentiation, proliferation, and apoptosis of F9 embryonal carcinoma cells. Journal of Cell Biology, 139(3):735–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/735.

Coopersmith:1997:BTM

[CCMG97] Craig M. Coopersmith, Chitra Chandrasekaran, M. Shane Mc-Nevin, and Jeffrey I. Gordon. Bi-transgenic mice reveal that K-ras^{Val12} augments a p53-independent apoptosis when small intestinal villus enterocytes reenter the cell cycle. Journal of Cell Biology, 138(1):167–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/1/167.

Collas:1996:TMS

[CCP96] P. Collas, J. C. Courvalin, and D. Poccia. Targeting of membranes to sea urchin sperm chromatin is mediated by a lamin B receptor-like integral membrane protein. Journal of Cell Biology, 135(6):1715–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1715.

Constam:1996:SSN

[CCR96] D. B. Constam, M. Calfon, and E. J. Robertson. SPC4, SPC6, and the novel protease SPC7 are coexpressed with bone morphogenetic proteins at distinct sites during embryogenesis. Journal of Cell Biology, 134(1):181–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/1/181.

Calautti:1998:TPS

[CCS^{+}98] Enzo Calautti, Sara Cabodi, Paul L. Stein, Mechthild Hatzfeld, Nancy Kedersha, and G. Paolo Dotto. Tyrosine Phosphorylation and Src Family Kinases Control Keratinocyte Cell–Cell Adhesion. Journal of Cell Biology, 141(6):1449–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/6/1449.

Cullen:1999:MS

[CDGO99] C. Fiona Cullen, Peter Deák, David M. Glover, and Hiroyuki Ohkura. mini spindles. Journal of Cell Biology, 146(5):1005– ??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/1005.

Cole:1998:CKI

[CDH^{+}98] Douglas G. Cole, Dennis R. Diener, Amy L. Himelblau, Peter L. Beech, Jason C. Fuster, and Joel L. Rosenbaum. Chlamydomonas kinesin-II-dependent intraflagellar transport (IFT): IFT particles contain proteins required for ciliary assembly in Caenorhabditis elegans sensory neurons. Journal of Cell Biology, 141(4):993–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/993.

Cordenonsi:1999:CCG

[CDH^{+}99] Michelangelo Cordenonsi, Fabio D’Atri, Eva Hammar, David A. D. Parry, John Kendrick-Jones, David Shore, and Sandra Citi. Cingulin Contains Globular and Coiled–Coil Domains and Interacts with Zo-1, Zo-2, Zo-3, and Myosin. Journal of Cell Biology, 147(7):1569–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/7/1569.

Chang:1997:CPR

[CDN97] Fred Chang, David Drubin, and Paul Nurse. cdc12p, a Protein Required for Cytokinesis in Fission Yeast, Is a Component of the Cell Division Ring and Interacts with Profilin. Journal of Cell Biology, 137(1):169–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/1/169.

Cabrera:1996:TRE

[CDRB^{+}96] N. Cabrera, E. Díaz-Rodríguez, E. Becker, D. Martín-Zanca, and A. Pandiella. TrkA receptor ectodomain cleavage generates a tyrosine-phosphorylated cell-associated fragment. Journal of Cell Biology, 132(3):427–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/3/427.

Clark:1997:TSL

[CES97] Rachael A. Clark, Harold P. Erickson, and Timothy A. Springer. Tenascin Supports Lymphocyte Rolling. Journal of Cell Biology, 137(3):755–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/3/755.

Cole:1998:RTG

[CES^{+}98] Nelson B. Cole, Jan Ellenberg, Jia Song, Diane DiEuliis, and Jennifer Lippincott-Schwartz. Retrograde Transport of Golgilocalized Proteins to the ER. Journal of Cell Biology, 140 (1):1–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/1.

Cimato:1997:NGF

[CEZA97] Thomas R. Cimato, Murray J. Ettinger, Xianbo Zhou, and John M. Aletta. Nerve Growth Factor–specific Regulation of Protein Methylation during Neuronal Differentiation of PC12 Cells. Journal of Cell Biology, 138(5):1089–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
5/1089.

Chang:1995:TYP

[CF95] A. Chang and G. R. Fink. Targeting of the yeast plasma membrane [H+]ATPase: a novel gene AST1 prevents mislocalization of mutant ATPase to the vacuole. Journal of Cell Biology, 128(1):39–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/1/39.

Chung:1999:PPP

[CF99] Chang Y. Chung and Richard A. Firtel. Paka, a Putative Pak Family Member, Is Required for Cytokinesis and the Regulation of the Cytoskeleton in Dictyostelium discoideum Cells during Chemotaxis. Journal of Cell Biology, 147(3): 559–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/3/559.

Chanson:1998:ESA

[CFB^{+}98] Marc Chanson, Marjorie Fanjul, Domenico Bosco, Eric Nelles, Susanne Suter, Klaus Willecke, and Paolo Meda. Enhanced Secretion of Amylase from Exocrine Pancreas of Connexin32deficient Mice. Journal of Cell Biology, 141(5):1267–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
5/1267.

Carl:1995:ISD

[CFC^{+}95] S. L. Carl, K. Felix, A. H. Caswell, N. R. Brandt, W. J. Ball, P. L. Vaghy, G. Meissner, and D. G. Ferguson. Immunolocalization of sarcolemmal dihydropyridine receptor and sarcoplasmic reticular triadin and ryanodine receptor in rabbit ventricle and atrium. Journal of Cell Biology, 129(3):673–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/673.

Chretien:1995:SGM

[CFK95] D. Chrétien, S. D. Fuller, and E. Karsenti. Structure of growing microtubule ends: two-dimensional sheets close into tubes at variable rates. Journal of Cell Biology, 129(5):1311–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1311.

Chernomordik:1998:PMF

[CFL^{+}98] Leonid V. Chernomordik, Vadim A. Frolov, Eugenia Leikina, Peter Bronk, and Joshua Zimmerberg. The Pathway of Membrane Fusion Catalyzed by Influenza Hemagglutinin: Restriction of Lipids, Hemifusion, and Lipidic Fusion Pore Formation. Journal of Cell Biology, 140(6):1369–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1369.

Clark:1998:SLG

[CFS98] Rachael A. Clark, Robert C. Fuhlbrigge, and Timothy A. Springer. L-selectin Ligands That Are O -glycoprotease Resistant and Distinct from MECA-79 Antigen are Sufficient for Tethering and Rolling of Lymphocytes on Human High Endothelial Venules. Journal of Cell Biology, 140(3):721– ??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/3/721.

Campbell:1995:MMC

[CG95] M. S. Campbell and G. J. Gorbsky. Microinjection of mitotic cells with the 3F3/2 anti-phosphoepitope antibody delays the onset of anaphase. Journal of Cell Biology, 129(5):1195–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/5/1195.

Costell:1999:PMI

[CGA^{+}99] Mercedes Costell, Erika Gustafsson, Attila Aszódi, Matthias Mörgelin, Wilhelm Bloch, Ernst Hunziker, Klaus Addicks, Rupert Timpl, and Reinhard Fässler. Perlecan Maintains the Integrity of Cartilage and Some Basement Membranes. Journal of Cell Biology, 147(5):1109–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/5/1109.

Ceccaldi:1995:DSA

[CGB^{+}95] P. E. Ceccaldi, F. Grohovaz, F. Benfenati, E. Chieregatti, P. Greengard, and F. Valtorta. Dephosphorylated synapsin I anchors synaptic vesicles to actin cytoskeleton: an analysis by videomicroscopy. Journal of Cell Biology, 128(5):905–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/5/905.

Chyung:1997:NSC

[CGC^{+}97a] Abraham S. C. Chyung, Barry D. Greenberg, David G. Cook, Robert W. Doms, and Virginia M.-Y. Lee. Novel β-secretase cleavage of β-amyloid precursor protein in the endoplasmic reticulum/intermediate compartment of NT2N cells. Journal of Cell Biology, 138(3):671–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/3/671.

Crepaldi:1997:EEH

[CGC^{+}97b] Tiziana Crepaldi, Alexis Gautreau, Paolo M. Comoglio, Daniel Louvard, and Monique Arpin. Ezrin Is an Effector of Hepatocyte Growth Factor–mediated Migration and Morphogenesis in Epithelial Cells. Journal of Cell Biology, 138(2):423–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
2/423.

Chaponnier:1995:SNT

[CGJ^{+}95] C. Chaponnier, M. Goethals, P. A. Janmey, F. Gabbiani, G. Gabbiani, and J. Vandekerckhove. The specific NH2terminal sequence Ac–EEED of alpha-smooth muscle actin plays a role in polymerization in vitro and in vivo. Journal of Cell Biology, 130(4):887–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/4/887.

Chilcote:1995:CSS

[CGM^{+}95] T. J. Chilcote, T. Galli, O. Mundigl, L. Edelmann, P. S. McPherson, K. Takei, and P. De Camilli. Cellubrevin and synaptobrevins: similar subcellular localization and biochemical properties in PC12 cells. Journal of Cell Biology, 129 (1):219–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/1/219.

Carr:1999:SBS

[CGM^{+}99] Chavela M. Carr, Eric Grote, Mary Munson, Frederick M. Hughson, and Peter J. Novick. Sec1p Binds to SNARE Complexes and Concentrates at Sites of Secretion. Journal of Cell Biology, 146(2):333–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/2/333.

Cottingham:1999:NRS

[CGMH99] Frank R. Cottingham, Larisa Gheber, Dana L. Miller, and M. Andrew Hoyt. Novel Roles for Saccharomyces cerevisiae Mitotic Spindle Motors. Journal of Cell Biology, 147(2): 335–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/2/335.

Chang:1995:CSR

[CGSP95] J. H. Chang, S. Gill, J. Settleman, and S. J. Parsons. c-Src regulates the simultaneous rearrangement of actin cytoskeleton, p190RhoGAP, and p120RasGAP following epidermal growth factor stimulation. Journal of Cell Biology, 130(2):355–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/2/355.

Cano-Gauci:1999:GDM

[CGSY^{+}99] Danielle F. Cano-Gauci, Howard H. Song, Huiling Yang, Colin McKerlie, Barbara Choo, Wen Shi, Rose Pullano, Tino D. Piscione, Silviu Grisaru, Shawn Soon, Larisa Sedlackova, A. Keith Tanswell, Tak W. Mak, Herman Yeger, Gina A. Lockwood, Norman D. Rosenblum, and Jorge Filmus. Glypican-3–deficient Mice Exhibit Developmental Overgrowth and Some of the Abnormalities Typical of Simpson–Golabi– Behmel Syndrome. Journal of Cell Biology, 146(1):255–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
146/1/255.

Chasserot-Golaz:1996:AIE

[CGVS^{+}96] S. Chasserot-Golaz, N. Vitale, I. Sagot, B. Delouche, S. Dirrig, L. A. Pradel, J. P. Henry, D. Aunis, and M. F. Bader. Annexin II in exocytosis: catecholamine secretion requires the translocation of p36 to the subplasmalemmal region in chromaffin cells. Journal of Cell Biology, 133(6):1217–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
6/1217.

Crowley:1995:TPC

[CH95] E. Crowley and A. F. Horwitz. Tyrosine phosphorylation and cytoskeletal tension regulate the release of fibroblast adhesions. Journal of Cell Biology, 131(2):525–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/2/525.

Cottingham:1997:MSP

[CH97] Frank R. Cottingham and M. Andrew Hoyt. Mitotic Spindle Positioning in Saccharomyces cerevisiae Is Accomplished by Antagonistically Acting Microtubule Motor Proteins. Journal of Cell Biology, 138(5):1041–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/5/1041.

Csink:1998:LSC

[CH98] Amy K. Csink and Steven Henikoff. Large-scale Chromosomal Movements During Interphase Progression in Drosophila. Journal of Cell Biology, 143(1):13–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/1/13.

Castle:1997:PSM

[CHC97] Anna M. Castle, Amy Y. Huang, and J. David Castle. Passive Sorting in Maturing Granules of AtT-20 Cells: The Entry and Exit of Salivary Amylase and Proline-rich Protein. Journal of Cell Biology, 138(1):45–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/1/45.

Cover:1997:AID

[CHH97] Timothy L. Cover, Phyllis I. Hanson, and John E. Heuser. Acid-induced Dissociation of VacA, the Helicobacter pylori Vacuolating Cytotoxin, Reveals Its Pattern of Assembly. Journal of Cell Biology, 138(4):759–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/4/759.

Camper:1997:IRC

[CHLÅ97] Lisbet Camper, Dick Heinegård, and Evy Lundgren-Åkerlund. Integrin α2β1 is a receptor for the cartilage matrix protein chondroadherin. Journal of Cell Biology, 138(5):1159– ??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/5/1159.

Cary:1998:IPM

[CHP^{+}98] Leslie A. Cary, Dong Cho Han, Thomas R. Polte, Steven K. Hanks, and Jun-Lin Guan. Identification of p130^{Cas} as a mediator of focal adhesion kinase-promoted cell migration. Journal of Cell Biology, 140(1):211–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/1/211.

Chao:1999:SMT

[CHW^{+}99] Daniel S. Chao, Jesse C. Hay, Shawn Winnick, Rytis Prekeris, Judith Klumperman, and Richard H. Scheller. SNARE Membrane Trafficking Dynamics In Vivo. Journal of Cell Biology, 144(5):869–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/5/869.

Cohen:1999:DKP

[CIK^{+}99] Ofer Cohen, Boaz Inbal, Joseph L. Kissil, Tal Raveh, Hanna Berissi, Taly Spivak-Kroizaman, Elena Feinstein, and Adi Kimchi. Dap-kinase Participates in TNF-α-And FAS–Induced Apoptosis and Its Function Requires the Death Domain. Journal of Cell Biology, 146(1):141–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/1/141.

Chen:1999:RDP

[CIRG99] Chih-Ying Chen, Michael F. Ingram, Peter H. Rosal, and Todd R. Graham. Role for Drs2p, a P–Type Atpase and Potential Aminophospholipid Translocase, in Yeast Late Golgi Function. Journal of Cell Biology, 147(6):1223–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1223.

Cohen:1995:DAD

[CJG^{+}95] M. W. Cohen, C. Jacobson, E. W. Godfrey, K. P. Campbell, and S. Carbonetto. Distribution of alpha-dystroglycan during embryonic nerve-muscle synaptogenesis. Journal of Cell Biology, 129(4):1093–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/4/1093.

Chen:1998:DRB

[CJH^{+}98] D. Chen, X. Ji, M. A. Harris, J. Q. Feng, G. Karsenty, A. J. Celeste, V. Rosen, G. R. Mundy, and S. E. Harris. Differential Roles for Bone Morphogenetic Protein (BMP) Receptor Type IB and IA in Differentiation and Specification of Mesenchymal Precursor Cells to Osteoblast and Adipocyte Lineages. Journal of Cell Biology, 142(1):295–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/1/295.

Cardoso:1997:MUS

[CJR^{+}97] M. Cristina Cardoso, Cuthbert Joseph, Hans-Peter Rahn, Regina Reusch, Bernardo Nadal-Ginard, and Heinrich Leonhardt. Mapping and Use of a Sequence that Targets DNA Ligase I to Sites of DNA Replication In Vivo. Journal of Cell Biology, 139(3):579–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/3/579.

Chan:1999:HBM

[CJS^{+}99] G. K. T. Chan, S. A. Jablonski, V. Sudakin, J. C. Hittle, and T. J. Yen. Human Bubr1 is a mitotic checkpoint kinase that monitors Cenp-E functions at kinetochores and binds the cyclosome/APC. Journal of Cell Biology, 146(5): 941–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/941.

Cohen:1997:LIC

[CJY^{+}97] Monroe W. Cohen, Christian Jacobson, Peter D. Yurchenco, Glenn E. Morris, and Salvatore Carbonetto. Laminin-induced Clustering of Dystroglycan on Embryonic Muscle Cells: Comparison with Agrin-induced Clustering. Journal of Cell Biology, 136(5):1047–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/5/1047.

Cooper:1996:SMF

[CK96] J. A. Cooper and D. P. Kiehart. Septins may form a ubiquitous family of cytoskeletal filaments. Journal of Cell Biology, 134(6):1345–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/6/1345.

Clark:1998:IMS

[CKB^{+}98] Edwin A. Clark, Warren G. King, Joan S. Brugge, Marc Symons, and Richard O. Hynes. Integrin-mediated Signals Regulated by Members of the Rho Family of GTPases. Journal of Cell Biology, 142(2):573–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/573.

Cheng:1995:CGR

[CKM95] L. Cheng, M. Khan, and A. W. Mudge. Calcitonin gene-related peptide promotes Schwann cell proliferation. Journal of Cell Biology, 129(3):789–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/3/789.

Cox:1996:ARA

[CKP96] R. T. Cox, C. Kirkpatrick, and M. Peifer. Armadillo is required for adherens junction assembly, cell polarity, and morphogenesis during Drosophila embryogenesis. Journal of Cell Biology, 134(1):133–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/1/133.

Corbett:1995:RRT

[CKS^{+}95] A. H. Corbett, D. M. Koepp, G. Schlenstedt, M. S. Lee, A. K. Hopper, and P. A. Silver. RNA1p, a Ran/TC4 GTPase activating protein, is required for nuclear import. Journal of Cell Biology, 130(5):1017–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/5/1017.

Cardoso:1999:DMA

[CL99] M. Cristina Cardoso and Heinrich Leonhardt. DNA Methyltransferase Is Actively Retained in the Cytoplasm during Early Development. Journal of Cell Biology, 147(1):25–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
1/25.

Chernomordik:1997:ESM

[CLF^{+}97a] Leonid V. Chernomordik, Eugenia Leikina, Vadim Frolov, Peter Bronk, and Joshua Zimmerberg. An Early Stage of Membrane Fusion Mediated by the Low pH Conformation of Influenza Hemagglutinin Depends upon Membrane Lipids. Journal of Cell Biology, 136(1):81–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/1/81.

Cunningham:1997:MAP

[CLF^{+}97b] C. Casey Cunningham, Nicole Leclerc, Lisa A. Flanagan, Mei Lu, Paul A. Janmey, and Kenneth S. Kosik. Microtubuleassociated Protein 2c Reorganizes Both Microtubules and Microfilaments into Distinct Cytological Structures in an Actinbinding Protein-280–deficient Melanoma Cell Line. Journal of Cell Biology, 136(4):845–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/4/845.

Crosbie:1999:MTS

[CLH^{+}99] Rachelle H. Crosbie, Connie S. Lebakken, Kathleen H. Holt, David P. Venzke, Volker Straub, Jane C. Lee, R. Mark Grady, Jeffery S. Chamberlain, Joshua R. Sanes, and Kevin P. Campbell. Membrane Targeting and Stabilization of Sarcospan Is Mediated by the Sarcoglycan Subcomplex. Journal of Cell Biology, 145(1):153–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/1/153.

Cheresh:1999:RCC

[CLK99] David A. Cheresh, Jie Leng, and Richard L. Klemke. Regulation of cell contraction and membrane ruffling by distinct signals in migratory cells. Journal of Cell Biology, 146(5): 1107–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/1107.

Chuang:1995:MCE

[CLL95] J. Z. Chuang, D. C. Lin, and S. Lin. Molecular cloning, expression, and mapping of the high affinity actin-capping domain of chicken cardiac tensin. Journal of Cell Biology, 128 (6):1095–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/6/1095.

Cubizolles:1998:PNX

[CLL^{+}98] Fabien Cubizolles, Vincent Legagneux, René Le Guellec, Isabelle Chartrain, Rustem Uzbekov, Chris Ford, and Katherine Le Guellec. pEg7, a New Xenopus Protein Required for Mitotic Chromosome Condensation in Egg Extracts. Journal of Cell Biology, 143(6):1437–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/6/1437.

Champliaud:1996:HAC

[CLR^{+}96] M. F. Champliaud, G. P. Lunstrum, P. Rousselle, T. Nishiyama, D. R. Keene, and R. E. Burgeson. Human amnion contains a novel laminin variant, laminin 7, which like laminin 6, covalently associates with laminin 5 to promote stable epithelialstromal attachment. Journal of Cell Biology, 132(6):1189–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
132/6/1189.

Carragher:1999:DCF

[CLRR99] Neil O. Carragher, Bodo Levkau, Russell Ross, and Elaine W. Raines. Degraded Collagen Fragments Promote Rapid Disassembly of Smooth Muscle Focal Adhesions That Correlates with Cleavage of Pp125^{FAK}, Paxillin, and Talin. Journal of Cell Biology, 147(3):619–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/3/619.

Campenot:1996:DNS

[CLS96] B. Campenot, K. Lund, and D. L. Senger. Delivery of newly synthesized tubulin to rapidly growing distal axons of sympathetic neurons in compartmented cultures. Journal of Cell Biology, 135(3):701–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/3/701.

Carlier:1997:ADF

[CLS^{+}97] Marie-France Carlier, Valérie Laurent, Jérôme Santolini, Ronald Melki, Dominique Didry, Gui-Xian Xia, Yan Hong, Nam-Hai Chua, and Dominique Pantaloni. Actin depolymerizing factor (ADF/cofilin) enhances the rate of filament turnover: Implication in actin-based motility. Journal of Cell Biology, 136(6):1307–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/6/1307.

Cheah:1995:HCD

[CLT^{+}95] K. S. Cheah, A. Levy, P. A. Trainor, A. W. Wai, T. Kuffner, C. L. So, K. K. Leung, R. H. Lovell-Badge, and P. P. Tam. Human COL2A1-directed SV40 T antigen expression in transgenic and chimeric mice results in abnormal skeletal development. Journal of Cell Biology, 128(1):223–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
1/223.

Chitaev:1996:BPD

[CLT^{+}96] N. A. Chitaev, R. E. Leube, R. B. Troyanovsky, L. G. Eshkind, W. W. Franke, and S. M. Troyanovsky. The binding of plakoglobin to desmosomal cadherins: patterns of binding sites and topogenic potential. Journal of Cell Biology, 133 (2):359–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/2/359.

Collas:1999:KAP

[CLT99] Philippe Collas, Katherine Le Guellec, and Kjetil Taskén. The a–Kinase–Anchoring Protein Akap95 Is a Multivalent Protein with a Key Role in Chromatin Condensation at Mitosis. Journal of Cell Biology, 147(6):1167–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/6/1167.

Chang:1995:LEC

[CM95a] D. C. Chang and C. Meng. A localized elevation of cytosolic free calcium is associated with cytokinesis in the zebrafish embryo. Journal of Cell Biology, 131(6):1539–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1539.

Cramer:1995:MIP

[CM95b] L. P. Cramer and T. J. Mitchison. Myosin is involved in postmitotic cell spreading. Journal of Cell Biology, 131(1): 179–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/179.

Clark:1998:ILP

[CMC^{+}98] Sharon F. Clark, Sally Martin, Amanda J. Carozzi, Michelle M. Hill, and David E. James. Intracellular Localization of Phosphatidylinositide 3-kinase and Insulin Receptor Substrate-1 in Adipocytes: Potential Involvement of a Membrane Skeleton. Journal of Cell Biology, 140(5):1211–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/5/1211.

Carmeliet:1998:RIR

[CMD^{+}98] Peter Carmeliet, Lieve Moons, Mieke Dewerchin, Steven Rosenberg, Jean-Marc Herbert, Florea Lupu, and Désiré Collen. Receptor-independent Role of Urokinase–Type Plasminogen Activator in Pericellular Plasmin and Matrix Metalloproteinase Proteolysis during Vascular Wound Healing in Mice. Journal of Cell Biology, 140(1):233–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
1/233.

Chilcoat:1996:GLP

[CMHT96] N. D. Chilcoat, S. M. Melia, A. Haddad, and A. P. Turkewitz. Granule lattice protein 1 (Grl1p), an acidic, calcium-binding protein in Tetrahymena thermophila dense-core secretory granules, influences granule size, shape, content organization, and release but not protein sorting or condensation. Journal of Cell Biology, 135(6):1775–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1775.

Chakravarti:1998:LRC

[CML^{+}98] Shukti Chakravarti, Terry Magnuson, Jonathan H. Lass, Karl J. Jepsen, Christian LaMantia, and Heidi Carroll. Lumican Regulates Collagen Fibril Assembly: Skin Fragility and Corneal Opacity in the Absence of Lumican. Journal of Cell Biology, 141(5):1277–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/5/1277.

Chant:1995:RBP

[CMM^{+}95] J. Chant, M. Mischke, E. Mitchell, I. Herskowitz, and J. R. Pringle. Role of Bud3p in producing the axial budding pattern of yeast. Journal of Cell Biology, 129(3):767–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/767.

Chavez:1996:BRS

[CMM96] R. A. Chavez, S. G. Miller, and H. P. Moore. A biosynthetic regulated secretory pathway in constitutive secretory cells. Journal of Cell Biology, 133(6):1177–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/6/1177.

Charlton:1997:FCM

[CMR^{+}97] Carol A. Charlton, William A. Mohler, Glenn L. Radice, Richard O. Hynes, and Helen M. Blau. Fusion Competence of Myoblasts Rendered Genetically Null for N–Cadherin in Culture. Journal of Cell Biology, 138(2):331–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
2/331.

Caceres:1997:RMD

[CMS^{+}97] Javier F. Cáceres, Tom Misteli, Gavin R. Screaton, David L. Spector, and Adrian R. Krainer. Role of the Modular Domains of SR Proteins in Subnuclear Localization and Alternative Splicing Specificity. Journal of Cell Biology, 138(2):225–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/2/225.

Clerk:1998:SPM

[CMS98] Angela Clerk, Ashour Michael, and Peter H. Sugden. Stimulation of the p38 Mitogen-activated Protein Kinase Pathway in Neonatal Rat Ventricular Myocytes by the G Protein–coupled Receptor Agonists, Endothelin-1 and Phenylephrine: a Role in Cardiac Myocyte Hypertrophy? Journal of Cell Biology, 142(2):523–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/523.

Chen:1996:RGA

[CMUW96] P. Chen, J. E. Murphy-Ullrich, and A. Wells. A role for gelsolin in actuating epidermal growth factor receptor-mediated cell motility. Journal of Cell Biology, 134(3):689–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
3/689.

Clemson:1996:XRP

[CMWL96] C. M. Clemson, J. A. McNeil, H. F. Willard, and J. B. Lawrence. XIST RNA paints the inactive X chromosome at interphase: evidence for a novel RNA involved in nuclear/chromosome structure. Journal of Cell Biology, 132(3): 259–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/259.

Cook:1998:RGT

[CNG98] Tiffani A. Cook, Takayuki Nagasaki, and Gregg G. Gundersen. Rho Guanosine Triphosphatase Mediates the Selective Stabilization of Microtubules Induced by Lysophosphatidic Acid. Journal of Cell Biology, 141(1):175–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/1/175.

Conese:1995:AMR

[CNP^{+}95] M. Conese, A. Nykjaer, C. M. Petersen, O. Cremona, R. Pardi, P. A. Andreasen, J. Gliemann, E. I. Christensen, and F. Blasi. alpha-2 Macroglobulin receptor/Ldl receptorrelated protein(Lrp)-dependent internalization of the urokinase receptor. Journal of Cell Biology, 131(6):1609–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1609.

Cole:1999:KIC

[Col99] Douglas G. Cole. Kinesin-ii, Coming and Going. Journal of Cell Biology, 147(3):463–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/3/463.

Castellino:1995:ERC

[COML95] F. Castellino, S. Ono, F. Matsumura, and A. Luini. Essential role of caldesmon in the actin filament reorganization induced by glucocorticoids. Journal of Cell Biology, 131(5): 1223–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/5/1223.

Chant:1995:PBS

[CP95] J. Chant and J. R. Pringle. Patterns of bud-site selection in the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 129(3):751–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/3/751.

Cockell:1995:CTS

[CPL^{+}95] M. Cockell, F. Palladino, T. Laroche, G. Kyrion, C. Liu, A. J. Lustig, and S. M. Gasser. The carboxy termini of Sir4 and Rap1 affect Sir3 localization: evidence for a multicomponent complex required for yeast telomeric silencing. Journal of Cell Biology, 129(4):909–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/4/909.

Charroux:1999:G

[CPP^{+}99] Bernard Charroux, Livio Pellizzoni, Robert A. Perkinson, Andrej Shevchenko, Matthias Mann, and Gideon Dreyfuss. Gemin3. Journal of Cell Biology, 147(6):1181–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1181.

Campbell:1996:BCC

[CQB^{+}96] J. J. Campbell, S. Qin, K. B. Bacon, C. R. Mackay, and E. C. Butcher. Biology of chemokine and classical chemoattractant receptors: differential requirements for adhesion-triggering versus chemotactic responses in lymphoid cells. Journal of Cell Biology, 134(1):255–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/1/255.

Constam:1999:RBM

[CR99] Daniel B. Constam and Elizabeth J. Robertson. Regulation of Bone Morphogenetic Protein Activity by Pro Domains and Proprotein Convertases. Journal of Cell Biology, 144(1): 139–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/1/139.

Costa:1998:PCC

[CRA^{+}98] Michael Costa, William Raich, Cristina Agbunag, Ben Leung, Jeff Hardin, and James R. Priess. A Putative Catenin– Cadherin System Mediates Morphogenesis of the Caenorhabditis elegans Embryo. Journal of Cell Biology, 141(1):297–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/1/297.

Cox:1995:GDA

[CRCH95] D. Cox, J. A. Ridsdale, J. Condeelis, and J. Hartwig. Genetic deletion of ABP-120 alters the three-dimensional organization of actin filaments in Dictyostelium pseudopods. Journal of Cell Biology, 128(5):819–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/5/819.

Castellani:1995:MDS

[CRG^{+}95] L. Castellani, M. C. Reedy, M. C. Gauzzi, C. Provenzano, S. Alemà, and G. Falcone. Maintenance of the differentiated state in skeletal muscle: activation of v–Src disrupts sarcomeres in quail myotubes. Journal of Cell Biology, 130 (4):871–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/871.

Chial:1998:SCN

[CRGW98] Heidi J. Chial, Michael P. Rout, Thomas H. Giddings, and Mark Winey. Saccharomyces cerevisiae Ndc1p Is a Shared Component of Nuclear Pore Complexes and Spindle Pole Bodies. Journal of Cell Biology, 143(7):1789–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1789.

Carey:1996:EUG

[CRLM96] K. L. Carey, S. A. Richards, K. M. Lounsbury, and I. G. Macara. Evidence using a green fluorescent proteinglucocorticoid receptor chimera that the Ran/TC4 GTPase mediates an essential function independent of nuclear protein import. Journal of Cell Biology, 133(5):985–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
5/985.

Corrado:1996:TMM

[CRM^{+}96] K. Corrado, J. A. Rafael, P. L. Mills, N. M. Cole, J. A. Faulkner, K. Wang, and J. S. Chamberlain. Transgenic mdx mice expressing dystrophin with a deletion in the actin-binding domain display a “mild Becker” phenotype. Journal of Cell Biology, 134(4):873–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/4/873.

Carroll:1997:CRF

[CRM^{+}97] David J. Carroll, Chodavarapu S. Ramarao, Lisa M. Mehlmann, Serge Roche, Mark Terasaki, and Laurinda A. Jaffe. Calcium Release at Fertilization in Starfish Eggs Is Mediated by Phospholipase Cγ. Journal of Cell Biology, 138(6): 1303–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1303.

Carmena:1998:DPK

[CRM^{+}98] Mar Carmena, Maria Giovanna Riparbelli, Gianluca Minestrini, Álvaro M. Tavares, Richard Adams, Giuliano Callaini, and David M. Glover. Drosophila Polo Kinase Is Required for Cytokinesis. Journal of Cell Biology, 143(3):659–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/659.

Chamberlain:1995:DEA

[CRMB95] L. H. Chamberlain, D. Roth, A. Morgan, and R. D. Burgoyne. Distinct effects of alpha–SNAP, 14-3-3 proteins, and calmodulin on priming and triggering of regulated exocytosis. Journal of Cell Biology, 130(5):1063–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/5/1063.

Cordes:1997:IPP

[CRRF97] Volker C. Cordes, Sonja Reidenbach, Hans-Richard Rackwitz, and Werner W. Franke. Identification of Protein p270/Tpr as a Constitutive Component of the Nuclear Pore Complex– attached Intranuclear Filaments. Journal of Cell Biology, 136 (3):515–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/515.

Collins:1999:BCM

[CRYC99] Lila R. Collins, William A. Ricketts, Linda Yeh, and David Cheresh. Bifurcation of Cell Migratory and Proliferative Signaling by the Adaptor Protein Shc. Journal of Cell Biology, 147(7):1561–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/7/1561.

Chuang:1996:DTT

[CS96a] J. S. Chuang and R. W. Schekman. Differential trafficking and timed localization of two chitin synthase proteins, Chs2p and Chs3p. Journal of Cell Biology, 135(3):597–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
3/597.

Cooper:1996:VCB

[CS96b] A. A. Cooper and T. H. Stevens. Vps10p cycles between the late-Golgi and prevacuolar compartments in its function as the sorting receptor for multiple yeast vacuolar hydrolases. Journal of Cell Biology, 133(3):529–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/3/529.

Carminati:1997:MOM

[CS97a] Janet L. Carminati and Tim Stearns. Microtubules Orient the Mitotic Spindle in Yeast through Dynein-dependent Interactions with the Cell Cortex. Journal of Cell Biology, 138 (3):629–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/3/629.

Corsi:1997:LDS

[CS97b] Ann K. Corsi and Randy Schekman. The Lumenal Domain of Sec63p Stimulates the ATPase Activity of BiP and Mediates BiP Recruitment to the Translocon in Saccharomyces cerevisiae. Journal of Cell Biology, 137(7):1483–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
7/1483.

Chuang:1998:CTR

[CS98] Jen-Zen Chuang and Ching-Hwa Sung. The Cytoplasmic Tail of Rhodopsin Acts as a Novel Apical Sorting Signal in Polarized MDCK Cells. Journal of Cell Biology, 142(5):1245– ??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/5/1245.

Chen:1999:ABS

[CS99] Shuqi Chen and Timothy A. Springer. An Automatic Braking System That Stabilizes Leukocyte Rolling by an Increase in Selectin Bond Number with Shear. Journal of Cell Biology, 144 (1):185–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/1/185.

Chen:1997:PDS

[CSA^{+}97] Ye-Guang Chen, Anirban Siddhanta, Cary D. Austin, Scott M. Hammond, Tsung-Chang Sung, Michael A. Frohman, Andrew J. Morris, and Dennis Shields. Phospholipase D Stimulates Release of Nascent Secretory Vesicles from the trans - Golgi Network. Journal of Cell Biology, 138(3):495–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/495.

Claas:1998:ABR

[CSC^{+}98] Christoph Claas, Simone Seiter, Andreas Claas, Larissa Savelyeva, Manfred Schwab, and Margot Zöller. Association Between the Rat Homologue of CO-029, a Metastasis-associated Tetraspanin Molecule and Consumption Coagulopathy. Journal of Cell Biology, 141(1):267–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/1/267.

Chen:1997:BSC

[CSCM97] Peng Chen, Stephanie K. Sapperstein, Jonathan D. Choi, and Susan Michaelis. Biogenesis of the Saccharomyces cerevisiae Mating Pheromone a–Factor. Journal of Cell Biology, 136(2): 251–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/2/251.

Creutz:1996:ILF

[CSDR96] C. E. Creutz, S. L. Snyder, S. N. Daigle, and J. Redick. Identification, localization, and functional implications of an abundant nematode annexin. Journal of Cell Biology, 132(6): 1079–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/6/1079.

Ciccotosto:1999:IIM

[CSEM99] Giuseppe D. Ciccotosto, Martin R. Schiller, Betty A. Eipper, and Richard E. Mains. Induction of Integral Membrane PAM Expression in AtT-20 Cells Alters the Storage and Trafficking of POMC and PC1. Journal of Cell Biology, 144(3): 459–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/459.

Chiozzi:1997:SCF

[CSF^{+}97] Paola Chiozzi, Juana M. Sanz, Davide Ferrari, Simonetta Falzoni, Arrigo Aleotti, Gary N. Buell, Ginetta Collo, and Francesco Di Virgilio. Spontaneous cell fusion in macrophage cultures expressing high levels of the P2Z/P2X_{7} receptor. Journal of Cell Biology, 138(3):697–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/3/697.

Cardone:1996:STP

[CSM^{+}96] M. H. Cardone, B. L. Smith, P. A. Mennitt, D. Mochly-Rosen, R. B. Silver, and K. E. Mostov. Signal transduction by the polymeric immunoglobulin receptor suggests a role in regulation of receptor transcytosis. Journal of Cell Biology, 133 (5):997–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/5/997.

Cramer:1997:ING

[CSM97] Louise P. Cramer, Margaret Siebert, and Timothy J. Mitchison. Identification of Novel Graded Polarity Actin Filament Bundles in Locomoting Heart Fibroblasts: Implications for the Generation of Motile Force. Journal of Cell Biology, 136 (6):1287–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/6/1287.

Chen:1998:SCP

[CSMM98] Rey-Huei Chen, Andrej Shevchenko, Matthias Mann, and Andrew W. Murray. Spindle Checkpoint Protein Xmad1 Recruits Xmad2 to Unattached Kinetochores. Journal of Cell Biology, 143(2):283–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/283.

Chen:1999:CAC

[CSN99] Yih-Tai Chen, Daniel B. Stewart, and W. James Nelson. Coupling assembly of the E-cadherin/β–catenin complex to efficient endoplasmic reticulum exit and basal-lateral membrane targeting of E-cadherin in polarized MDCK cells. Journal of Cell Biology, 144(4):687–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/4/687.

Carvalho:1995:TAE

[CSO^{+}95] T. Carvalho, J. S. Seeler, K. Ohman, P. Jordan, U. Pettersson, G. Akusjärvi, M. Carmo-Fonseca, and A. Dejean. Targeting of adenovirus E1A and E4-ORF3 proteins to nuclear matrix-associated PML bodies. Journal of Cell Biology, 131 (1):45–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/45.

Caulin:1997:CCK

[CSO97] Carlos Caulín, Guy S. Salvesen, and Robert G. Oshima. Caspase Cleavage of Keratin 18 and Reorganization of Intermediate Filaments during Epithelial Cell Apoptosis. Journal of Cell Biology, 138(6):1379–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/6/1379.

Conrad:1995:CCB

[CSY^{+}95] P. A. Conrad, E. J. Smart, Y. S. Ying, R. G. Anderson, and G. S. Bloom. Caveolin cycles between plasma membrane caveolae and the Golgi complex by microtubule-dependent and microtubule-independent steps. Journal of Cell Biology, 131(6):1421–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/6/1421.

Chan:1998:CKB

[CSY98] G. K. T. Chan, B. T. Schaar, and T. J. Yen. Characterization of the Kinetochore Binding Domain of CENP–E Reveals Interactions with the Kinetochore Proteins CENP–F and hBUBR1. Journal of Cell Biology, 143(1):49–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/143/1/49.

Chilcoat:1997:VAM

[CT97a] N. Doane Chilcoat and Aaron P. Turkewitz. In vivo analysis of the major exocytosis-sensitive phosphoprotein in Tetrahymena. Journal of Cell Biology, 139(5):1197–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
5/1197.

Chitaev:1997:DCD

[CT97b] Nikolai A. Chitaev and Sergey M. Troyanovsky. Direct Ca^{2+}-dependent heterophilic interaction between desmosomal cadherins, desmoglein and desmocollin, contributes to cell– cell adhesion. Journal of Cell Biology, 138(1):193–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
1/193.

Chitaev:1998:ALC

[CT98] Nikolai A. Chitaev and Sergey M. Troyanovsky. Adhesive But Not Lateral E-cadherin Complexes Require Calcium and Catenins for Their Formation. Journal of Cell Biology, 142 (3):837–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/3/837.

Caston:1997:SVS

[CTB^{+}97] José R. Castón, Benes L. Trus, Frank P. Booy, Reed B. Wickner, Joseph S. Wall, and Alasdair C. Steven. Structure of L–A Virus: a Specialized Compartment for the Transcription and Replication of Double-stranded RNA. Journal of Cell Biology, 138(5):975–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/5/975.

Cunningham:1995:API

[Cun95] C. C. Cunningham. Actin polymerization and intracellular solvent flow in cell surface blebbing. Journal of Cell Biology, 129(6):1589–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1589.

Chrzanowska-Wodnicka:1996:RSC

[CWB96] M. Chrzanowska-Wodnicka and K. Burridge. Rho-stimulated contractility drives the formation of stress fibers and focal adhesions. Journal of Cell Biology, 133(6):1403–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
6/1403.

Cohen:1998:HCL

[CWM^{+}98] Alexandra R. Cohen, Daniel F. Wood, Shirin M. Marfatia, Zenta Walther, Athar H. Chishti, and James Melvin Anderson. Human CASK/LIN-2 binds Syndecan-2 and protein 4.1 and localizes to the basolateral membrane of epithelial cells. Journal of Cell Biology, 142(1):129–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/1/129.

Chen:1997:TDS

[CWMD97] Lucinda I. Chen, Melanie K. Webster, April N. Meyer, and Daniel J. Donoghue. Transmembrane Domain Sequence Requirements for Activation of the p185^{c−neu} Receptor Tyrosine Kinase. Journal of Cell Biology, 137(3):619–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
3/619.

Chang:1999:PIP

[CWSG99] Chia-Che Chang, Daniel S. Warren, Katherine A. Sacksteder, and Stephen J. Gould. Pex12 Interacts with Pex5 and Pex10 and Acts Downstream of Receptor Docking in Peroxisomal Matrix Protein Import. Journal of Cell Biology, 147(4): 761–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/4/761.

Chang:1997:INI

[CWSL97] David D. Chang, Carol Wong, Healy Smith, and Jenny Liu. ICAP-1, a novel β_{1} integrin cytoplasmic domain-associated protein, binds to a conserved and functionally important NPXY sequence motif of β_{1} integrin. Journal of Cell Biology, 138(5):1149–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/5/1149.

Cherry:1998:DDD

[CWT^{+}98] Richard J. Cherry, Keith M. Wilson, Kathy Triantafilou, Peter O’Toole, Ian E. G. Morrison, Patricia R. Smith, and Nelson Fernández. Detection of Dimers of Dimers of Human Leukocyte Antigen (HLA)–DR on the Surface of Living Cells by Single–Particle Fluorescence Imaging. Journal of Cell Biology, 140(1):71–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/1/71.

Colognato:1999:LPI

[CWY99] Holly Colognato, Donald A. Winkelmann, and Peter D. Yurchenco. Laminin Polymerization Induces a Receptor– Cytoskeleton Network. Journal of Cell Biology, 145(3):619–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/3/619.

Xiong:1997:IAA

[cXP97] Wen cheng Xiong and J. Thomas Parsons. Induction of Apoptosis after Expression of PYK2, a Tyrosine Kinase Structurally Related to Focal Adhesion Kinase. Journal of Cell Biology, 139 (2):529–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/529.

Chen:1998:YSP

[CYH98] Xiaoyue Peter Chen, Hongwei Yin, and Tim C. Huffaker. The Yeast Spindle Pole Body Component Spc72p Interacts with Stu2p and Is Required for Proper Microtubule Assembly. Journal of Cell Biology, 141(5):1169–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/5/1169.

Cope:1999:NPK

[CYSD99] M. Jamie T. V. Cope, Shirley Yang, Ching Shang, and David G. Drubin. Novel Protein Kinases Ark1p and Prk1p Associate with and Regulate the Cortical Actin Cytoskeleton in Budding Yeast. Journal of Cell Biology, 144(6):1203–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
144/6/1203.

Claude:1999:G

[CZK^{+}99] Alejandro Claude, Bao-Ping Zhao, Craig E. Kuziemsky, Sophie Dahan, Scott J. Berger, Jian-Ping Yan, Adrian D. Armold, Eric M. Sullivan, and Paul Melançon. Gbf1. Journal of Cell Biology, 146(1):71–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/1/71.

Drees:1999:MDZ

[DAB99] Beth E. Drees, Katy M. Andrews, and Mary C. Beckerle. Molecular Dissection of Zyxin Function Reveals Its Involvement in Cell Motility. Journal of Cell Biology, 147(7):1549– ??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/7/1549.

DeMarini:1997:SBH

[DAF^{+}97] Douglas J. DeMarini, Alison E. M. Adams, Hanna Fares, Claudio De Virgilio, Giorgio Valle, John S. Chuang, and John R. Pringle. A Septin-based Hierarchy of Proteins Required for Localized Deposition of Chitin in the Saccharomyces cerevisiae Cell Wall. Journal of Cell Biology, 139(1):75–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/139/1/75.

Diamond:1995:HAL

[DAP^{+}95] M. S. Diamond, R. Alon, C. A. Parkos, M. T. Quinn, and T. A. Springer. Heparin is an adhesive ligand for the leukocyte integrin Mac-1 (CD11b/CD1). Journal of Cell Biology, 130(6): 1473–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1473.

Davidson:1995:WCM

[Dav95] H. W. Davidson. Wortmannin causes mistargeting of procathepsin D. Evidence for the involvement of a phosphatidylinositol 3-kinase in vesicular transport to lysosomes. Journal of Cell Biology, 130(4):797–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/4/797.

Drees:1995:TEY

[DBBB95] B. Drees, C. Brown, B. G. Barrell, and A. Bretscher. Tropomyosin is essential in yeast, yet the TPM1 and TPM2 products perform distinct functions. Journal of Cell Biology, 128(3):383–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/3/383.

DeAngelis:1999:SMP

[DBC^{+}99] Luciana De Angelis, Libera Berghella, Marcello Coletta, Laura Lattanzi, Malvina Zanchi, M. Gabriella, Carola Ponzetto, and Giulio Cossu. Skeletal Myogenic Progenitors Originating from Embryonic Dorsal Aorta Coexpress Endothelial and Myogenic Markers and Contribute to Postnatal Muscle Growth and Regeneration. Journal of Cell Biology, 147(4):869–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
4/869.

deBelle:1998:GSB

[dBCKS98] Ian de Belle, Shutao Cai, and Terumi Kohwi-Shigematsu. The Genomic Sequences Bound to Special AT-rich Sequencebinding Protein 1 (SATB1) In Vivo in Jurkat T Cells Are Tightly Associated with the Nuclear Matrix at the Bases of the Chromatin Loops. Journal of Cell Biology, 141(2):335–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/2/335.

Darsow:1998:ALM

[DBE98] Tamara Darsow, Christopher G. Burd, and Scott D. Emr. Acidic di-leucine motif essential for AP-3-dependent sorting and restriction of the functional specificity of the Vam3p vacuolar t-SNARE. Journal of Cell Biology, 142(4):913–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
142/4/913.

Denzer:1997:ABN

[DBG^{+}97] Alain J. Denzer, Ralph Brandenberger, Matthias Gesemann, Matthias Chiquet, and Markus A. Ruegg. Agrin Binds to the Nerve–Muscle Basal Lamina via Laminin. Journal of Cell Biology, 137(3):671–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/3/671.

Danielson:1997:TDD

[DBH^{+}97] Keith G. Danielson, Helene Baribault, David F. Holmes, Helen Graham, Karl E. Kadler, and Renato V. Iozzo. Targeted Disruption of Decorin Leads to Abnormal Collagen Fibril Morphology and Skin Fragility. Journal of Cell Biology, 136(3): 729–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/729.

Dufour:1999:DFC

[DBJDT99] Sylvie Dufour, Alice Beauvais-Jouneau, Annie Delouvée, and Jean Paul Thiery. Differential Function of N–Cadherin and Cadherin-7 in the Control of Embryonic Cell Motility. Journal of Cell Biology, 146(2):501–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/2/501.

Driessen:1999:CCT

[DBLD^{+}99] Christoph Driessen, Rebecca A. R. Bryant, Ana-Maria Lennon-Duménil, José A. Villadangos, Paula Wolf Bryant, Guo-Ping Shi, Harold A. Chapman, and Hidde L. Ploegh. Cathepsin S Controls the Trafficking and Maturation of Mhc Class II Molecules in Dendritic Cells. Journal of Cell Biology, 147(4):775–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/4/775.

Davis:1998:ITL

[DBOM98] Elaine C. Davis, Thomas J. Broekelmann, Yuji Ozawa, and Robert P. Mecham. Identification of Tropoelastin as a Ligand for the 65-kD FK506-binding Protein, FKBP65, in the Secretory Pathway. Journal of Cell Biology, 140(2):295–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
2/295.

Damke:1995:CIP

[DBvdBS95] H. Damke, T. Baba, A. M. van der Bliek, and S. L. Schmid. Clathrin-independent pinocytosis is induced in cells overexpressing a temperature-sensitive mutant of dynamin. Journal of Cell Biology, 131(1):69–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/1/69.

Davis:1996:SLM

[DC96a] A. F. Davis and D. A. Clayton. In situ localization of mitochondrial DNA replication in intact mammalian cells. Journal of Cell Biology, 135(4):883–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/4/883.

DeBernardo:1996:HID

[DC96b] A. P. DeBernardo and S. Chang. Heterophilic interactions of DM–GRASP: GRASP–NgCAM interactions involved in neurite extension. Journal of Cell Biology, 133(3):657–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/657.

Strambio-de-Castillia:1995:ICN

[dCBR95] C. Strambio de Castillia, G. Blobel, and M. P. Rout. Isolation and characterization of nuclear envelopes from the yeast Saccharomyces. Journal of Cell Biology, 131(1):19–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
1/19.

Strambio-de-Castillia:1999:PCN

[dCBR99] Caterina Strambio de Castillia, Günter Blobel, and Michael P. Rout. Proteins Connecting the Nuclear Pore Complex with the Nuclear Interior. Journal of Cell Biology, 144(5):839–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
144/5/839.

Dejucq:1997:TAD

[DCJ97] Nathalie Dejucq, Suzanne Chousterman, and Bernard Jégou. The Testicular Antiviral Defense System: Localization, Expression, and Regulation of 2’5’ Oligoadenylate Synthetase, Double–Stranded RNA-activated Protein Kinase, and Mx Proteins in the Rat Seminiferous Tubule. Journal of Cell Biology, 139(4):865–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/4/865.

Desnos:1995:BSV

[DCOK95] C. Desnos, L. Clift-O’Grady, and R. B. Kelly. Biogenesis of synaptic vesicles in vitro. Journal of Cell Biology, 130(5): 1041–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/5/1041.

Debiec:1998:CAM

[DCR98] Hanna Debiec, Erik Ilsø Christensen, and Pierre Marie Ronco. The Cell Adhesion Molecule L1 Is Developmentally Regulated in the Renal Epithelium and Is Involved in Kidney Branching Morphogenesis. Journal of Cell Biology, 143(7):2067–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/7/2067.

deChaves:1995:ASP

[dCVCV95] E. Posse de Chaves, D. E. Vance, R. B. Campenot, and J. E. Vance. Axonal synthesis of phosphatidylcholine is required for normal axonal growth in rat sympathetic neurons. Journal of Cell Biology, 128(5):913–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/5/913.

Decoster:1998:ATN

[DCVF98] Els Decoster, Sigrid Cornelis, Bart Vanhaesebroeck, and Walter Fiers. Autocrine tumor necrosis factor (TNF) and lymphotoxin (LT) α differentially modulate cellular sensitivity to TNF/LT-α cytotoxicity in L929 cells. Journal of Cell Biology, 143(7):2057–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/7/2057.

Deng:1996:PAI

[DCW^{+}96] G. Deng, S. A. Curriden, S. Wang, S. Rosenberg, and D. J. Loskutoff. Is plasminogen activator inhibitor-1 the molecular switch that governs urokinase receptor-mediated cell adhesion and release? Journal of Cell Biology, 134(6):1563–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/6/1563.

Dominguez:1998:GEP

[DDF^{+}98] Michel Dominguez, Kurt Dejgaard, Joachim Füllekrug, Sophie Dahan, Ali Fazel, Jean-Pierre Paccaud, David Y. Thomas, John J. M. Bergeron, and Tommy Nilsson. gp25L/ emp24/ p24 Protein Family Members of the cis- Golgi Network Bind Both COP I and II Coatomer. Journal of Cell Biology, 140(4): 751–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/4/751.

Dominov:1998:BEI

[DDM98] Janice A. Dominov, Jonathan J. Dunn, and Jeffrey Boone Miller. Bcl-2 Expression Identifies an Early Stage of Myogenesis and Promotes Clonal Expansion of Muscle Cells. Journal of Cell Biology, 142(2):537–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/537.

Drgonova:1999:GBP

[DDRC99] Jana Drgonová, Tomás Drgon, Dong-Hyun Roh, and Enrico Cabib. The GTP-binding Protein Rho1p Is Required for Cell Cycle Progression and Polarization of the Yeast Cell. Journal of Cell Biology, 146(2):373–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/2/373.

Demaurex:1996:IPR

[DDWG96] N. Demaurex, G. P. Downey, T. K. Waddell, and S. Grinstein. Intracellular pH regulation during spreading of human neutrophils. Journal of Cell Biology, 133(6):1391–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
6/1391.

Du:1997:PNR

[DDWM97] Shao Jun Du, Stephen H. Devoto, Monte Westerfield, and Randall T. Moon. Positive and Negative Regulation of Muscle Cell Identity by Members of the hedgehog and TGF-β Gene Families. Journal of Cell Biology, 139(1):145–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/1/145.

Distel:1996:UNP

[DEG^{+}96] B. Distel, R. Erdmann, S. J. Gould, G. Blobel, D. I. Crane, J. M. Cregg, G. Dodt, Y. Fujiki, J. M. Goodman, W. W. Just, J. A. Kiel, W. H. Kunau, P. B. Lazarow, G. P. Mannaerts, H. W. Moser, T. Osumi, R. A. Rachubinski, A. Roscher, S. Subramani, H. F. Tabak, T. Tsukamoto, D. Valle, I. van der Klei, P. P. van Veldhoven, and M. Veenhuis. A unified nomenclature for peroxisome biogenesis factors. Journal of Cell Biology, 135(1):1–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/1/1.

DeZwaan:1997:KRK

[DEPR97] Todd M. DeZwaan, Eric Ellingson, David Pellman, and David M. Roof. Kinesin-related KIP3 of Saccharomyces cerevisiae Is Required for a Distinct Step in Nuclear Migration. Journal of Cell Biology, 138(5):1023–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
5/1023.

Dustin:1996:VCI

[DFC^{+}96] M. L. Dustin, L. M. Ferguson, P. Y. Chan, T. A. Springer, and D. E. Golan. Visualization of CD2 interaction with LFA-3 and determination of the two-dimensional dissociation constant for adhesion receptors in a contact area. Journal of Cell Biology, 132(3):465–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/3/465.

Dehouck:1997:NFL

[DFD^{+}97] Bénédicte Dehouck, Laurence Fenart, Marie-Pierre Dehouck, Annick Pierce, Gérard Torpier, and Roméo Cecchelli. A New Function for the LDL Receptor: Transcytosis of LDL across the Blood–Brain Barrier. Journal of Cell Biology, 138(4): 877–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/4/877.

Dominguez:1999:FDG

[DFD^{+}99] Michel Dominguez, Ali Fazel, Sophie Dahan, Jacque Lovell, Louis Hermo, Alejandro Claude, Paul Melançon, and J. J. M. Bergeron. Fusogenic domains of Golgi membranes are sequestered into specialized regions of the stack that can be released by mechanical fragmentation. Journal of Cell Biology, 145(4):673–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/673.

Dogterom:1996:IMP

[DFGL96] M. Dogterom, M. A. Félix, C. C. Guet, and S. Leibler. Influence of M-phase chromatin on the anisotropy of microtubule asters. Journal of Cell Biology, 133(1):125–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
1/125.

Drummond:1999:TDA

[DFL^{+}99] Sheona Drummond, Paul Ferrigno, Carol Lyon, Jackie Murphy, Martin Goldberg, Terry Allen, Carl Smythe, and Christopher J. Hutchison. Temporal Differences in the Appearance of NEP–B78 and an LBR-like Protein during Xenopus Nuclear Envelope Reassembly Reflect the Ordered Recruitment of Functionally Discrete Vesicle Types. Journal of Cell Biology, 144(2):225–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/2/225.

Doberstein:1997:GAM

[DFMG97] Stephen K. Doberstein, Richard D. Fetter, Anand Y. Mehta, and Corey S. Goodman. Genetic Analysis of Myoblast Fusion: blown fuse Is Required for Progression Beyond the Prefusion Complex. Journal of Cell Biology, 136(6):1249–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
6/1249.

Dodt:1996:MPG

[DG96] G. Dodt and S. J. Gould. Multiple PEX genes are required for proper subcellular distribution and stability of Pex5p, the PTS1 receptor: evidence that PTS1 protein import is mediated by a cycling receptor. Journal of Cell Biology, 135(6): 1763–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1763.

Dupree:1999:AGI

[DGP99] Jeffrey L. Dupree, Jean-Antoine Girault, and Brian Popko. Axo-glial Interactions Regulate the Localization of Axonal Paranodal Proteins. Journal of Cell Biology, 147(6):1145– ??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/6/1145.

Denzer:1995:ATE

[DGSR95] A. J. Denzer, M. Gesemann, B. Schumacher, and M. A. Ruegg. An amino-terminal extension is required for the secretion of chick agrin and its binding to extracellular matrix. Journal of Cell Biology, 131(6):1547–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1547.

Didenko:1996:PDS

[DH96] V. V. Didenko and P. J. Hornsby. Presence of double-strand breaks with single-base 3’ overhangs in cells undergoing apoptosis but not necrosis. Journal of Cell Biology, 135(5):1369– ??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/5/1369.

Delley:1999:CWS

[DH99] Pierre-Alain Delley and Michael N. Hall. Cell Wall Stress Depolarizes Cell Growth via Hyperactivation of Rho1. Journal of Cell Biology, 147(1):163–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/147/1/163.

DiPersio:1997:IRN

[DHDJ^{+}97] C. Michael DiPersio, Kairbaan M. Hodivala-Dilke, Rudolf Jaenisch, Jordan A. Kreidberg, and Richard O. Hynes. α3β1 Integrin Is Required for Normal Development of the Epidermal Basement Membrane. Journal of Cell Biology, 137(3):729–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/3/729.

Dittie:1996:AAC

[DHT96] A. S. Dittie, N. Hajibagheri, and S. A. Tooze. The AP-1 adaptor complex binds to immature secretory granules from PC12 cells, and is regulated by ADP-ribosylation factor. Journal of Cell Biology, 132(4):523–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/4/523.

deHoop:1995:IRW

[dHvPL^{+}95] M. de Hoop, C. von Poser, C. Lange, E. Ikonen, W. Hunziker, and C. G. Dotti. Intracellular routing of wild-type and mutated polymeric immunoglobulin receptor in hippocampal neurons in culture. Journal of Cell Biology, 130(6):1447– ??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1447.

Donaldson:1996:SPC

[DK96] A. D. Donaldson and J. V. Kilmartin. Spc42p: a phosphorylated component of the S. cerevisiae spindle pole body (SPD) with an essential function during SPB duplication. Journal of Cell Biology, 132(5):887–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/5/887.

Debec:1996:LAF

[DKD^{+}96] A. Debec, R. F. Kalpin, D. R. Daily, P. D. McCallum, W. F. Rothwell, and W. Sullivan. Live analysis of free centrosomes in normal and aphidicolin-treated Drosophila embryos. Journal of Cell Biology, 134(1):103–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/1/103.

Dechampesme:1999:ARR

[DKLS^{+}99] Anne-Marie Dechampesme, Olga Koroleva, Isabelle Leger-Silvestre, Nicole Gas, and Sylvie Camier. Assembly of 5S Ribosomal RNA Is Required at a Specific Step of the Pre-rRNA Processing Pathway. Journal of Cell Biology, 145(7):1369–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/7/1369.

Dietrich:1997:RFC

[DKN^{+}97] Jes Dietrich, Jesper Kastrup, Bodil L. Nielsen, Niels Ødum, and Carsten Geisler. Regulation and Function of the CD3γ DxxxLL Motif: a Binding Site for Adaptor Protein-1 and Adaptor Protein-2 in Vitro. Journal of Cell Biology, 138 (2):271–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/2/271.

Davis:1996:MCN

[DLB96] J. Q. Davis, S. Lambert, and V. Bennett. Molecular composition of the node of Ranvier: identification of ankyrinbinding cell adhesion molecules neurofascin (mucin+/third FNIII domain-) and NrCAM at nodal axon segments. Journal of Cell Biology, 135(5):1355–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/5/1355.

Duchen:1998:TMD

[DLC98] Michael R. Duchen, Anne Leyssens, and Martin Crompton. Transient Mitochondrial Depolarizations Reflect Focal Sarcoplasmic Reticular Calcium Release in Single Rat Cardiomyocytes. Journal of Cell Biology, 142(4):975–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
4/975.

Deng:1995:DDR

[DLGB95] H. Deng, J. K. Lee, L. S. Goldstein, and D. Branton. Drosophila development requires spectrin network formation. Journal of Cell Biology, 128(1):71–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/1/71.

Durbeej:1995:NMA

[DLIB^{+}95] M. Durbeej, E. Larsson, O. Ibraghimov-Beskrovnaya, S. L. Roberds, K. P. Campbell, and P. Ekblom. Non-muscle alphadystroglycan is involved in epithelial development. Journal of Cell Biology, 130(1):79–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/1/79.

Desai:1995:NRM

[DM95] A. Desai and T. J. Mitchison. A new role for motor proteins as couplers to depolymerizing microtubules. Journal of Cell Biology, 128(1):1–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/1/1.

Dubreuil:1996:NMC

[DMD^{+}96] R. R. Dubreuil, G. MacVicar, S. Dissanayake, C. Liu, D. Homer, and M. Hortsch. Neuroglian-mediated cell adhesion induces assembly of the membrane skeleton at cell contact sites. Journal of Cell Biology, 133(3):647–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/647.

Dyer:1996:SSP

[DMG96] J. M. Dyer, J. A. McNew, and J. M. Goodman. The sorting sequence of the peroxisomal integral membrane protein PMP47 is contained within a short hydrophilic loop. Journal of Cell Biology, 133(2):269–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/2/269.

DeBie:1996:IPC

[DMM^{+}96] I. De Bie, M. Marcinkiewicz, D. Malide, C. Lazure, K. Nakayama, M. Bendayan, and N. G. Seidah. The isoforms of proprotein convertase PC5 are sorted to different subcellular compartments. Journal of Cell Biology, 135(5): 1261–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/5/1261.

Desai:1998:ACM

[DMMS98] Arshad Desai, Paul S. Maddox, Timothy J. Mitchison, and E. D. Salmon. Anaphase A Chromosome Movement and Poleward Spindle Microtubule Flux Occur At Similar Rates in Xenopus Extract Spindles. Journal of Cell Biology, 141 (3):703–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/3/703.

Denker:1996:DDA

[DMP^{+}96] S. P. Denker, J. M. McCaffery, G. E. Palade, P. A. Insel, and M. G. Farquhar. Differential distribution of alpha subunits and beta gamma subunits of heterotrimeric G proteins on Golgi membranes of the exocrine pancreas. Journal of Cell Biology, 133(5):1027–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/5/1027.

Dallas:1995:DRL

[DMS^{+}95] S. L. Dallas, K. Miyazono, T. M. Skerry, G. R. Mundy, and L. F. Bonewald. Dual role for the latent transforming growth factor-beta binding protein in storage of latent TGF-beta in the extracellular matrix and as a structural matrix protein. Journal of Cell Biology, 131(2):539–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/2/539.

Dempsey:1997:AEH

[DMY^{+}97] Peter J. Dempsey, Katherine S. Meise, Yoshino Yoshitake, Katsuzo Nishikawa, and Robert J. Coffey. Apical Enrichment of Human EGF Precursor in Madin–Darby Canine Kidney Cells Involves Preferential Basolateral Ectodomain Cleavage Sensitive to a Metalloprotease Inhibitor. Journal of Cell Biology, 138(4):747–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/4/747.

DeArcangelis:1996:ILA

[DNB^{+}96] A. De Arcangelis, P. Neuville, R. Boukamel, O. Lefebvre, M. Kedinger, and P. Simon-Assmann. Inhibition of laminin alpha 1-chain expression leads to alteration of basement membrane assembly and cell differentiation. Journal of Cell Biology, 133(2):417–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/2/417.

Demetriou:1995:RCI

[DNC^{+}95] M. Demetriou, I. R. Nabi, M. Coppolino, S. Dedhar, and J. W. Dennis. Reduced contact-inhibition and substratum adhesion in epithelial cells expressing GlcNAc-transferase V. Journal of Cell Biology, 130(2):383–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/2/383.

Dietrich:1996:RCG

[DNH^{+}96] J. Dietrich, A. Neisig, X. Hou, A. M. Wegener, M. Gajhede, and C. Geisler. Role of CD3 gamma in T cell receptor assembly. Journal of Cell Biology, 132(3):299–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
3/299.

Dumenil:1998:IIS

[DOP^{+}98] G. Duménil, J. C. Olivo, S. Pellegrini, M. Fellous, P. J. Sansonetti, and G. Tran Van Nhieu. Interferon α Inhibits a Srcmediated Pathway Necessary for Shigella -induced Cytoskeletal Rearrangements in Epithelial Cells. Journal of Cell Biology, 143(4):1003–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/4/1003.

Desagher:1999:BIC

[DOSN^{+}99] Solange Desagher, Astrid Osen-Sand, Anthony Nichols, Robert Eskes, Sylvie Montessuit, Sandra Lauper, Kinsey Maundrell, Bruno Antonsson, and Jean-Claude Martinou. Bid-induced Conformational Change of Bax Is Responsible for Mitochondrial Cytochrome c Release during Apoptosis. Journal of Cell Biology, 144(5):891–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/5/891.

Dove:1999:B

[Dov99] Alan W. Dove. In brief. Journal of Cell Biology, 146(1):0–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
146/1/0.

Dai:1998:RTP

[DP98] Zhengshan Dai and H. Benjamin Peng. A Role of Tyrosine Phosphatase in Acetylcholine Receptor Cluster Dispersal and Formation. Journal of Cell Biology, 141(7):1613–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
7/1613.

DiPaolo:1996:PSE

[DPC^{+}96] G. Di Paolo, V. Pellier, M. Catsicas, B. Antonsson, S. Catsicas, and G. Grenningloh. The phosphoprotein stathmin is essential for nerve growth factor-stimulated differentiation. Journal of Cell Biology, 133(6):1383–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/6/1383.

delPozo:1997:IRC

[dPCM^{+}97] Miguel Angel del Pozo, Carlos Cabañas, María C. Montoya, Ann Ager, Paloma Sánchez-Mateos, and Francisco Sánchez-Madrid. ICAMs Redistributed by Chemokines to Cellular Uropods as a Mechanism for Recruitment of T Lymphocytes. Journal of Cell Biology, 137(2):493–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/493.

Portillo:1995:TST

[dPF95] F. Garcia del Portillo and B. B. Finlay. Targeting of Salmonella typhimurium to vesicles containing lysosomal membrane glycoproteins bypasses compartments with mannose 6-phosphate receptors. Journal of Cell Biology, 129(1):81–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/1/81.

Diamantopoulos:1999:DLC

[DPG^{+}99] Georgios S. Diamantopoulos, Franck Perez, Holly V. Goodson, Gérard Batelier, Ronald Melki, Thomas E. Kreis, and Janet E. Rickard. Dynamic Localization of CLIP-170 to Microtubule Plus Ends Is Coupled to Microtubule Assembly. Journal of Cell Biology, 144(1):99–??, January 1999. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/1/99.

Dorer:1995:SCC

[DPH95] R. Dorer, P. M. Pryciak, and L. H. Hartwell. Saccharomyces cerevisiae cells execute a default pathway to select a mate in the absence of pheromone gradients. Journal of Cell Biology, 131(4):845–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/845.

Danieli:1996:MFM

[DPHW96] T. Danieli, S. L. Pelletier, Y. I. Henis, and J. M. White. Membrane fusion mediated by the influenza virus hemagglutinin requires the concerted action of at least three hemagglutinin trimers. Journal of Cell Biology, 133(3):559–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/559.

Dicou:1997:TPD

[DPM^{+}97] Eleni Dicou, Beth Pflug, Marilyn Magazin, Thérèse Lehy, Daniel Djakiew, Pascual Ferrara, Véronique Nerrière, and Douglas Harvie. Two Peptides Derived from the Nerve Growth Factor Precursor Are Biologically Active. Journal of Cell Biology, 136(2):389–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/2/389.

delPozo:1995:CRC

[dPSMNSM95] M. A. del Pozo, P. Sánchez-Mateos, M. Nieto, and F. Sánchez-Madrid. Chemokines regulate cellular polarization and adhesion receptor redistribution during lymphocyte interaction with endothelium and extracellular matrix. Involvement of cAMP signaling pathway. Journal of Cell Biology, 131(2): 495–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/2/495.

Deconinck:1997:PAN

[DPT^{+}97] Anne E. Deconinck, Allyson C. Potter, Jonathon M. Tinsley, Sarah J. Wood, Ruth Vater, Carol Young, Laurent Metzinger, Angela Vincent, Clarke R. Slater, and Kay E. Davies.

Postsynaptic Abnormalities at the Neuromuscular Junctions of Utrophin-deficient Mice. Journal of Cell Biology, 136(4): 883–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/4/883.

Dawson:1997:CMV

[DPZ^{+}97] David W. Dawson, S. Frieda A. Pearce, Ruiqin Zhong, Roy L. Silverstein, William A. Frazier, and Noël P. Bouck. CD36 Mediates the In Vitro Inhibitory Effects of Thrombospondin-1 on Endothelial Cells. Journal of Cell Biology, 138(3):707–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/3/707.

Dawson:1995:DCC

[DRAT95] I. A. Dawson, S. Roth, and S. Artavanis-Tsakonas. The Drosophila cell cycle gene fizzy is required for normal degradation of cyclins A and B during mitosis and has homology to the CDC20 gene of Saccharomyces cerevisiae. Journal of Cell Biology, 129(3):725–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/3/725.

Darsow:1997:MSH

[DRE97] Tamara Darsow, Stephanie E. Rieder, and Scott D. Emr. A Multispecificity Syntaxin Homologue, Vam3p, Essential for Autophagic and Biosynthetic Protein Transport to the Vacuole. Journal of Cell Biology, 138(3):517–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/517.

Dembinsky:1996:CMI

[DRR96] A. Dembinsky, H. Rubin, and S. Ravid. Chemoattractantmediated increases in cGMP induce changes in Dictyostelium myosin II heavy chain-specific protein kinase C activities. Journal of Cell Biology, 134(4):911–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/4/911.

deSaintPhalle:1998:SAM

[dS98a] Brigitte de Saint Phalle and William Sullivan. Spindle Assembly and Mitosis without Centrosomes in Parthenogenetic Sciara Embryos. Journal of Cell Biology, 141(6):1383–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
6/1383.

Demeter:1998:INC

[DS98b] Janos Demeter and Shelley Sazer. imp2, a New Component of the Actin Ring in the Fission Yeast Schizosaccharomyces pombe. Journal of Cell Biology, 143(2):415–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/2/415.

Doyle:1995:PZN

[DSC^{+}95] J. P. Doyle, J. G. Stempak, P. Cowin, D. R. Colman, and D. D’Urso. Protein zero, a nervous system adhesion molecule, triggers epithelial reversion in host carcinoma cells. Journal of Cell Biology, 131(2):465–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/2/465.

Doria:1999:EHE

[DSC^{+}99] Margherita Doria, Anna Elisabetta Salcini, Emanuela Colombo, Tristram G. Parslow, Pier Giuseppe Pelicci, and Pier Paolo Di Fiore. The Eps15 Homology (Eh) Domain–Based Interaction between Eps15 and Hrb Connects the Molecular Machinery of Endocytosis to That of Nucleocytosolic Transport. Journal of Cell Biology, 147(7):1379–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/7/1379.

Driessens:1995:TDC

[DSE^{+}95] M. H. Driessens, P. J. Stroeken, N. F. Rodriguez Erena, M. A. van der Valk, E. A. van Rijthoven, and E. Roos. Targeted disruption of CD44 in MDAY–D2 lymphosarcoma cells has no effect on subcutaneous growth or metastatic capacity. Journal of Cell Biology, 131(6):1849–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1849.

Daro:1997:IEF

[DSG^{+}97] Elizabeth Daro, David Sheff, Marie Gomez, Thomas Kreis, and Ira Mellman. Inhibition of Endosome Function in CHO Cells Bearing a Temperature-sensitive Defect in the Coatomer (COPI) Component ϵ-COP. Journal of Cell Biology, 139(7): 1747–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1747.

David:1998:ILC

[DSG98] Doris David, Sumathy Sundarababu, and Jeffrey E. Gerst. Involvement of Long Chain Fatty Acid Elongation in the Trafficking of Secretory Vesicles in Yeast. Journal of Cell Biology, 143(5):1167–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/5/1167.

Demlehner:1995:CAH

[DSGF95] M. P. Demlehner, S. Schäfer, C. Grund, and W. W. Franke. Continual assembly of half-desmosomal structures in the absence of cell contacts and their frustrated endocytosis: a coordinated Sisyphus cycle. Journal of Cell Biology, 131(3): 745–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/3/745.

Devarajan:1996:ISC

[DSM^{+}96] P. Devarajan, P. R. Stabach, A. S. Mann, T. Ardito, M. Kashgarian, and J. S. Morrow. Identification of a small cytoplasmic ankyrin (AnkG119) in the kidney and muscle that binds beta I sigma spectrin and associates with the Golgi apparatus. Journal of Cell Biology, 133(4):819–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/4/819.

Dharmawardhane:1997:LPA

[DSM^{+}97] Suranganie Dharmawardhane, Luraynne C. Sanders, Stuart S. Martin, R. Hugh Daniels, and Gary M. Bokoch. Localization of p21-Activated Kinase 1 (PAK1) to Pinocytic Vesicles and Cortical Actin Structures in Stimulated Cells. Journal of Cell Biology, 138(6):1265–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/6/1265.

Du:1998:ITR

[DSM^{+}98a] Xiaobo Du, J. Daniel Stoops, James R. Mertz, C. Michael Stanley, and Joseph L. Dixon. Identification of Two Regions in Apolipoprotein B100 that are Exposed on the Cytosolic Side of the Endoplasmic Reticulum Membrane. Journal of Cell Biology, 141(3):585–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/3/585.

Duclos:1998:PMD

[DSM^{+}98b] Franck Duclos, Volker Straub, Steven A. Moore, David P. Venzke, Ron F. Hrstka, Rachelle H. Crosbie, Madeleine Durbeej, Connie S. Lebakken, Audrey J. Ettinger, Jack van der Meulen, Kathleen H. Holt, Leland E. Lim, Joshua R. Sanes, Beverly L. Davidson, John A. Faulkner, Roger Williamson, and Kevin P. Campbell. Progressive Muscular Dystrophy in α-Sarcoglycan–deficient Mice. Journal of Cell Biology, 142(6): 1461–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/6/1461.

Diaz:1997:NRE

[DSP97] Elva Díaz, Frauke Schimmöller, and Suzanne R. Pfeffer. A Novel Rab9 Effector Required for Endosome–to–TGN Transport. Journal of Cell Biology, 138(2):283–??, July 1997. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/138/2/
283.

DSouza-Schorey:1998:ATR

[DSvDH^{+}98] Crislyn D’Souza-Schorey, Elly van Donselaar, Victor W. Hsu, Chunzhi Yang, Philip D. Stahl, and Peter J. Peters. ARF6 Targets Recycling Vesicles to the Plasma Membrane: Insights from an Ultrastructural Investigation. Journal of Cell Biology, 140(3):603–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/603.

Dunia:1996:HMR

[DSvdV^{+}96] I. Dunia, J. J. Smit, M. A. van der Valk, H. Bloemendal, P. Borst, and E. L. Benedetti. Human multidrug resistance 3-P-glycoprotein expression in transgenic mice induces lens membrane alterations leading to cataract. Journal of Cell Biology, 132(4):701–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/701.

Dieck:1998:BNZ

[DSVL^{+}98] Susannetom Dieck, Lydia Sanmartí-Vila, Kristina Langnaese, Karin Richter, Stefan Kindler, Antje Soyke, Heike Wex, Karl-Heinz Smalla, Udo Kämpf, Jürgen-Theodor Fränzer, Markus Stumm, Craig C. Garner, and Eckart D. Gundelfinger. Bassoon, a novel zinc-finger CAG/glutamine-repeat protein selectively localized at the active zone of presynaptic nerve terminals. Journal of Cell Biology, 142(2):499–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/142/2/
499.

Deckert:1996:EGA

[DTB96] M. Deckert, M. Ticchioni, and A. Bernard. Endocytosis of GPI-anchored proteins in human lymphocytes: role of glycolipid-based domains, actin cytoskeleton, and protein kinases. Journal of Cell Biology, 133(4):791–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
4/791.

Dimitrova:1999:MRC

[DTMG99] Daniela S. Dimitrova, Ivan T. Todorov, Thomas Melendy, and David M. Gilbert. Mcm2, but Not Rpa, Is a Component of the Mammalian Early G1-Phase Prereplication Complex. Journal of Cell Biology, 146(4):709–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/4/709.

Danielsen:1995:TLG

[DvD95] E. M. Danielsen and B. van Deurs. A transferrin-like GPIlinked iron-binding protein in detergent-insoluble noncaveolar microdomains at the apical surface of fetal intestinal epithelial cells. Journal of Cell Biology, 131(4):939–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/939.

Deshmukh:1996:GMS

[DVD^{+}96] M. Deshmukh, J. Vasilakos, T. L. Deckwerth, P. A. Lampe, B. D. Shivers, and E. M. Johnson. Genetic and metabolic status of NGF-deprived sympathetic neurons saved by an inhibitor of ICE family proteases. Journal of Cell Biology, 135(5):1341–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/5/1341.

Depoortere:1998:RCD

[DVL^{+}98] Fabienne Depoortere, Alexandra Van Keymeulen, Jiri Lukas, Sabine Costagliola, Jirina Bartkova, Jacques E. Dumont, Jiri Bartek, Pierre P. Roger, and Sarah Dremier. A requirement for cyclin D3–cyclin-dependent kinase (cdk)-4 assembly in the cyclic adenosine monophosphate-dependent proliferation of thyrocytes. Journal of Cell Biology, 140(6):1427–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1427.

Du:1997:FIB

[DW97] Lei Du and Stephen L. Warren. A Functional Interaction between the Carboxy–Terminal Domain of RNA Polymerase II and Pre-mRNA Splicing. Journal of Cell Biology, 136(1): 5–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/5.

Dujardin:1998:ERC

[DWM^{+}98] Denis Dujardin, U. Irene Wacker, Anne Moreau, Trina A. Schroer, Janet E. Rickard, and Jan R. De Mey. Evidence for a Role of CLIP-170 in the Establishment of Metaphase Chromosome Alignment. Journal of Cell Biology, 141(4):849–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
4/849.

Discher:1995:MPS

[DWS^{+}95] D. E. Discher, R. Winardi, P. O. Schischmanoff, M. Parra, J. G. Conboy, and N. Mohandas. Mechanochemistry of protein 4.1’s spectrin-actin-binding domain: ternary complex interactions, membrane binding, network integration, structural strengthening. Journal of Cell Biology, 130(4):897–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
4/897.

Dowling:1996:BIR

[DYF96] J. Dowling, Q. C. Yu, and E. Fuchs. Beta4 integrin is required for hemidesmosome formation, cell adhesion and cell survival. Journal of Cell Biology, 134(2):559–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/2/559.

DelMaschio:1996:PLA

[DZC^{+}96] A. Del Maschio, A. Zanetti, M. Corada, Y. Rival, L. Ruco, M. G. Lampugnani, and E. Dejana. Polymorphonuclear leukocyte adhesion triggers the disorganization of endothelial cellto-cell adherens junctions. Journal of Cell Biology, 135(2): 497–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/2/497.

Dictenberg:1998:PTF

[DZS^{+}98] Jason B. Dictenberg, Wendy Zimmerman, Cynthia A. Sparks, Aaron Young, Charles Vidair, Yixian Zheng, Walter Carrington, Fredric S. Fay, and Stephen J. Doxsey. Pericentrin and γ-Tubulin Form a Protein Complex and Are Organized into a Novel Lattice at the Centrosome. Journal of Cell Biology, 141 (1):163–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/163.

Eaton:1995:CRC

[EAL^{+}95] S. Eaton, P. Auvinen, L. Luo, Y. N. Jan, and K. Simons. CDC42 and Rac1 control different actin-dependent processes in the Drosophila wing disc epithelium. Journal of Cell Biology, 131(1):151–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/1/151.

Eckley:1997:CPC

[EAM^{+}97] D. Mark Eckley, Alexandra M. Ainsztein, Alastair M. Mackay, Ilya G. Goldberg, and William C. Earnshaw. Chromosomal Proteins and Cytokinesis: Patterns of Cleavage Furrow Formation and Inner Centromere Protein Positioning in Mitotic Heterokaryons and Mid-anaphase Cells. Journal of Cell Biology, 136(6):1169–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/6/1169.

Eskes:1998:BIC

[EAOS^{+}98] Robert Eskes, Bruno Antonsson, Astrid Osen-Sand, Sylvie Montessuit, Christoph Richter, Rémy Sadoul, Gonzalo Mazzei, Anthony Nichols, and Jean-Claude Martinou. Bax-induced Cytochrome C Release from Mitochondria Is Independent of the Permeability Transition Pore but Highly Dependent on Mg^{2+} Ions. Journal of Cell Biology, 143(1):217–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
1/217.

Erdmann:1995:GPO

[EB95] R. Erdmann and G. Blobel. Giant peroxisomes in oleic acid-induced Saccharomyces cerevisiae lacking the peroxisomal membrane protein Pmp27p. Journal of Cell Biology, 128 (4):509–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/509.

Erdmann:1996:IPP

[EB96] R. Erdmann and G. Blobel. Identification of Pex13p a peroxisomal membrane receptor for the PTS1 recognition factor. Journal of Cell Biology, 135(1):111–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/1/111.

Engert:1996:PPD

[EBR96] J. C. Engert, E. B. Berglund, and N. Rosenthal. Proliferation precedes differentiation in IGF–I–stimulated myogenesis. Journal of Cell Biology, 135(2):431–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/2/431.

Erlebacher:1996:IET

[ED96] A. Erlebacher and R. Derynck. Increased expression of TGFbeta 2 in osteoblasts results in an osteoporosis-like phenotype.

Journal of Cell Biology, 132(1):195–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/1/195.

Eils:1996:TDR

[EDB^{+}96] R. Eils, S. Dietzel, E. Bertin, E. Schröck, M. R. Speicher, T. Ried, M. Robert-Nicoud, C. Cremer, and T. Cremer. Threedimensional reconstruction of painted human interphase chromosomes: active and inactive X chromosome territories have similar volumes but differ in shape and surface structure. Journal of Cell Biology, 135(6):1427–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/6/1427.

Elgersma:1998:MPR

[EEHW^{+}98] Ype Elgersma, Minetta Elgersma-Hooisma, Thibaut Wenzel, J. Michael McCaffery, Marilyn G. Farquhar, and Suresh Subramani. A Mobile PTS2 Receptor for Peroxisomal Protein Import in Pichia pastoris. Journal of Cell Biology, 140(4): 807–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/4/807.

Elfgang:1995:SPS

[EELF^{+}95] C. Elfgang, R. Eckert, H. Lichtenberg-Fraté, A. Butterweck, O. Traub, R. A. Klein, D. F. Hülser, and K. Willecke. Specific permeability and selective formation of gap junction channels in connexin-transfected HeLa cells. Journal of Cell Biology, 129(3):805–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/3/805.

Elder:1998:AMS

[EFB^{+}98] Gregory A. Elder, Victor L. Friedrich, Paolo Bosco, Chulho Kang, Andrei Gourov, Pang-Hsien Tu, Virginia M.-Y. Lee, and Robert A. Lazzarini. Absence of the Mid-sized Neurofilament Subunit Decreases Axonal Calibers, Levels of Light Neurofilament (NF–L), and Neurofilament Content. Journal of Cell Biology, 141(3):727–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/3/727.

Elder:1998:RHN

[EFK^{+}98] Gregory A. Elder, Victor L. Friedrich, Chulho Kang, Paolo Bosco, Andrei Gourov, Pang-Hsien Tu, Bin Zhang, Virginia M.-Y. Lee, and Robert A. Lazzarini. Requirement of Heavy Neurofilament Subunit in the Development of Axons with Large Calibers. Journal of Cell Biology, 143(1):195–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/1/195.

Elder:1999:ARA

[EFML99] Gregory A. Elder, Victor L. Friedrich, Alla Margita, and Robert A. Lazzarini. Age-related Atrophy of Motor Axons in Mice Deficient in the Mid–Sized Neurofilament Subunit. Journal of Cell Biology, 146(1):181–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/1/181.

Erickson:1997:DMC

[EGA97] Mary Ruth S. Erickson, Brian J. Galletta, and Susan M. Abmayr. Drosophila myoblast city Encodes a Conserved Protein That Is Essential for Myoblast Fusion, Dorsal Closure, and Cytoskeletal Organization. Journal of Cell Biology, 138 (3):589–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/3/589.

Espenshade:1995:YSG

[EGH^{+}95] P. Espenshade, R. E. Gimeno, E. Holzmacher, P. Teung, and C. A. Kaiser. Yeast SEC16 gene encodes a multidomain vesicle coat protein that interacts with Sec23p. Journal of Cell Biology, 131(2):311–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/2/311.

Engqvist-Goldstein:1999:ABP

[EGKC^{+}99] Åsa E. Y. Engqvist-Goldstein, Michael M. Kessels, Vikramjit S. Chopra, Michael R. Hayden, and David G. Drubin. An actinbinding protein of the Sla2/Huntingtin interacting protein 1 family is a novel component of clathrin-coated pits and vesicles. Journal of Cell Biology, 147(7):1503–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1503.

Eckley:1999:ADS

[EGM^{+}99] D. Mark Eckley, Steven R. Gill, Karin A. Melkonian, James B. Bingham, Holly V. Goodson, John E. Heuser, and Trina A. Schroer. Analysis of Dynactin Subcomplexes Reveals a Novel Actin–Related Protein Associated with the Arp1 Minifilament Pointed End. Journal of Cell Biology, 147(2):307–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
2/307.

Ebneth:1998:OTP

[EGS^{+}98] A. Ebneth, R. Godemann, K. Stamer, S. Illenberger, B. Trinczek, E.-M. Mandelkow, and E. Mandelkow. Overexpression of Tau Protein Inhibits Kinesin-dependent Trafficking of Vesicles, Mitochondria, and Endoplasmic Reticulum: Implications for Alzheimer’s Disease. Journal of Cell Biology, 143(3):777–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/3/777.

Eddy:1997:CPT

[EHC97] R. J. Eddy, J. Han, and J. S. Condeelis. Capping Protein Terminates but Does Not Initiate Chemoattractant-induced Actin Assembly in Dictyostelium. Journal of Cell Biology, 139(5): 1243–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1243.

Einheber:1995:TGF

[EHM^{+}95] S. Einheber, M. J. Hannocks, C. N. Metz, D. B. Rifkin, and J. L. Salzer. Transforming growth factor-beta 1 regulates axon/Schwann cell interactions. Journal of Cell Biology, 129 (2):443–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/443.

Enomoto-Iwamoto:1998:BMP

[EIIM^{+}98] Motomi Enomoto-Iwamoto, Masahiro Iwamoto, Yoshiki Mukudai, Yasuhiko Kawakami, Tsutomu Nohno, Yoshinobu Higuchi, Seiji Takemoto, Hideyo Ohuchi, Sumihare Noji, and Kojiro Kurisu. Bone Morphogenetic Protein Signaling Is Required for Maintenance of Differentiated Phenotype, Control of Proliferation, and Hypertrophy in Chondrocytes. Journal of Cell Biology, 140(2):409–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/2/409.

Eckmann:1997:XDS

[EJ97] Christian R. Eckmann and Michael F. Jantsch. Xlrbpa, a Double-stranded RNA-binding Protein Associated with Ribosomes and Heterogeneous Nuclear RNPs. Journal of Cell Biology, 138(2):239–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/2/239.

Eckmann:1999:REE

[EJ99] Christian R. Eckmann and Michael F. Jantsch. The RNAediting Enzyme ADAR1 Is Localized to the Nascent Ribonucleoprotein Matrix on Xenopus Lampbrush Chromosomes but Specifically Associates with an Atypical Loop. Journal of Cell Biology, 144(4):603–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/144/4/603.

Endow:1997:SDD

[EK97] Sharyn A. Endow and Donald J. Komma. Spindle Dynamics during Meiosis in Drosophila Oocytes. Journal of Cell Biology, 137(6):1321–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/6/1321.

Erxleben:1997:MCA

[EKF^{+}97] Christian Erxleben, Norbert Klauke, Matthias Flötenmeyer, Marie-Pierre Blanchard, Claudia Braun, and Helmut Plattner. Microdomain Ca^{2+} Activation during Exocytosis in Paramecium Cells. Superposition of Local Subplasmalemmal Calcium Store Activation by Local Ca^{2+} Influx. Journal of Cell Biology, 136(3):597–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/3/597.

Elgersma:1996:SDS

[EKK^{+}96] Y. Elgersma, L. Kwast, A. Klein, T. Voorn-Brouwer, M. van den Berg, B. Metzig, T. America, H. F. Tabak, and B. Distel. The SH3 domain of the Saccharomyces cerevisiae peroxisomal membrane protein Pex13p functions as a docking site for Pex5p, a mobile receptor for the import PTS1containing proteins. Journal of Cell Biology, 135(1):97–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/1/97.

Ehlers:1995:NSR

[EKPK95] M. D. Ehlers, D. R. Kaplan, D. L. Price, and V. E. Koliatsos. NGF-stimulated retrograde transport of trkA in the mammalian nervous system. Journal of Cell Biology, 130(1):149–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/1/149.

Espelin:1997:PAS

[EKS97] Christopher W. Espelin, Kenneth B. Kaplan, and Peter K. Sorger. Probing the Architecture of a Simple Kinetochore Using DNA–Protein Crosslinking. Journal of Cell Biology, 139(6):1383–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/6/1383.

Eliceiri:1998:IRS

[EKSC98] Brian P. Eliceiri, Richard Klemke, Staffan Strömblad, and David A. Cheresh. Integrin α vβ3 Requirement for Sustained Mitogen-activated Protein Kinase Activity during Angiogenesis. Journal of Cell Biology, 140(5):1255–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
5/1255.

Egile:1999:ACE

[ELL^{+}99] Coumaran Egile, Thomas P. Loisel, Valérie Laurent, Rong Li, Dominique Pantaloni, Philippe J. Sansonetti, and Marie-France Carlier. Activation of the Cdc42 effector N-wasp by the Shigella flexneri Icsa protein promotes actin nucleation by Arp2/3 complex and bacterial actin-based motility. Journal of Cell Biology, 146(6):1319–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/6/1319.

Erdman:1998:PRG

[ELMS98] Scott Erdman, Li Lin, Michael Malczynski, and Michael Snyder. Pheromone-regulated Genes Required for Yeast Mating Differentiation. Journal of Cell Biology, 140(3):461–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/3/461.

Elia:1996:RAT

[EM96] L. Elia and L. Marsh. Role of the ABC transporter Ste6 in cell fusion during yeast conjugation. Journal of Cell Biology, 135(3):741–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/3/741.

Elia:1998:RPM

[EM98] Lisa Elia and Lorraine Marsh. A Role for a Protease in Morphogenic Responses during Yeast Cell Fusion. Journal of Cell Biology, 142(6):1473–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/6/1473.

Erdmann:1996:DEP

[EMB^{+}96] S. Erdmann, W. Müller, S. Bahrami, S. I. Vornehm, H. Mayer, P. Bruckner, K. von der Mark, and H. Burkhardt. Differential effects of parathyroid hormone fragments on collagen gene expression in chondrocytes. Journal of Cell Biology, 135(4): 1179–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/1179.

Espada:1999:HRA

[EPMB^{+}99] Jesús Espada, Mirna Pérez-Moreno, Vania M. M. Braga, Pablo Rodriguez-Viciana, and Amparo Cano. H-Ras Activation Promotes Cytoplasmic Accumulation and Phosphoinositide 3-Oh Kinase Association of β-Catenin in Epidermal Keratinocytes. Journal of Cell Biology, 146(5):967–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
5/967.

Echeverri:1996:MCK

[EPVV96] C. J. Echeverri, B. M. Paschal, K. T. Vaughan, and R. B. Vallee. Molecular characterization of the 50-kD subunit of dynactin reveals function for the complex in chromosome alignment and spindle organization during mitosis. Journal of Cell Biology, 132(4):617–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/132/4/617.

Erickson:1996:PRT

[ES96] H. P. Erickson and D. Stoffler. Protofilaments and rings, two conformations of the tubulin family conserved from bacterial FtsZ to alpha/beta and gamma tubulin. Journal of Cell Biology, 135(1):5–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/1/5.

Eferl:1999:FCJ

[ESH^{+}99] Robert Eferl, Maria Sibilia, Frank Hilberg, Andrea Fuchsbichler, Iris Kufferath, Barbara Guertl, Rainer Zenz, Erwin F. Wagner, and Kurt Zatloukal. Functions of c–Jun in Liver and Heart Development. Journal of Cell Biology, 145(5):1049–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/5/1049.

Ellenberg:1997:NMD

[ESM^{+}97] Jan Ellenberg, Eric D. Siggia, Jorge E. Moreira, Carolyn L. Smith, John F. Presley, Howard J. Worman, and Jennifer Lippincott-Schwartz. Nuclear Membrane Dynamics and Reassembly in Living Cells: Targeting of an Inner Nuclear Membrane Protein in Interphase and Mitosis. Journal of Cell Biology, 138(6):1193–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/6/1193.

Eitzen:1997:EPP

[ESR97] Gary A. Eitzen, Rachel K. Szilard, and Richard A. Rachubinski. Enlarged Peroxisomes Are Present in Oleic Acid–grown Yarrowia lipolytica Overexpressing the PEX16 Gene Encoding an Intraperoxisomal Peripheral Membrane Peroxin. Journal of Cell Biology, 137(6):1265–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/6/1265.

Elion:1995:FLN

[ETF95] E. A. Elion, J. Trueheart, and G. R. Fink. Fus2 localizes near the site of cell fusion and is required for both cell fusion and nuclear alignment during zygote formation. Journal of Cell Biology, 130(6):1283–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/6/1283.

Esni:1999:NCA

[ETP^{+}99] Farzad Esni, Inge-Bert Täljedal, Anne-Karina Perl, Harold Cremer, Gerhard Christofori, and Henrik Semb. Neural Cell Adhesion Molecule (N-CAM) Is Required for Cell Type Segregation and Normal Ultrastructure in Pancreatic Islets. Journal of Cell Biology, 144(2):325–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/2/325.

Emeis:1997:ESG

[EvdESvdH^{+}97] J. J. Emeis, Y. van den Eijnden-Schrauwen, C. M. van den Hoogen, W. de Priester, A. Westmuckett, and F. Lupu. An Endothelial Storage Granule for Tissue–Type Plasminogen Activator. Journal of Cell Biology, 139(1):245–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
1/245.

Eaton:1996:RRC

[EWS96] S. Eaton, R. Wepf, and K. Simons. Roles for Rac1 and Cdc42 in planar polarization and hair outgrowth in the wing of Drosophila. Journal of Cell Biology, 135(5):1277–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/5/1277.

Esser:1998:VEG

[EWW^{+}98] Sybille Esser, Karen Wolburg, Hartwig Wolburg, Georg Breier, Teymuras Kurzchalia, and Werner Risau. Vascular Endothelial Growth Factor Induces Endothelial Fenestrations In Vitro. Journal of Cell Biology, 140(4):947–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
4/947.

Ethell:1999:CSH

[EY99] Iryna M. Ethell and Yu Yamaguchi. Cell Surface Heparan Sulfate Proteoglycan Syndecan-2 Induces the Maturation of Dendritic Spines in Rat Hippocampal Neurons. Journal of Cell Biology, 144(3):575–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/3/575.

Einheber:1997:AMP

[EZC^{+}97] Steven Einheber, George Zanazzi, William Ching, Steven Scherer, Teresa A. Milner, Elior Peles, and James L. Salzer. The Axonal Membrane Protein Caspr, a Homologue of Neurexin IV, Is a Component of the Septate-like Paranodal Junctions That Assemble during Myelination. Journal of Cell Biology, 139(6):1495–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/6/1495.

Field:1995:ACR

[FA95] C. M. Field and B. M. Alberts. Anillin, a contractile ring protein that cycles from the nucleus to the cell cortex. Journal of Cell Biology, 131(1):165–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/1/165.

Field:1996:PDS

[FaAR^{+}96] C. M. Field, O. al Awar, J. Rosenblatt, M. L. Wong, B. Alberts, and T. J. Mitchison. A purified Drosophila septin complex forms filaments and exhibits GTPase activity. Journal of Cell Biology, 133(3):605–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/3/605.

Faulstich:1996:ACM

[FAO^{+}96] D. Faulstich, S. Auerbach, L. Orci, M. Ravazzola, S. Wegchingel, F. Lottspeich, G. Stenbeck, C. Harter, F. T. Wieland, and H. Tschochner. Architecture of coatomer: molecular characterization of delta–COP and protein interactions within the complex. Journal of Cell Biology, 135(1):53–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/135/1/53.

Fuhlbrigge:1996:SFL

[FAP^{+}96] R. C. Fuhlbrigge, R. Alon, K. D. Puri, J. B. Lowe, and T. A. Springer. Sialylated, fucosylated ligands for L-selectin expressed on leukocytes mediate tethering and rolling adhesions in physiologic flow conditions. Journal of Cell Biology, 135(3): 837–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/837.

Fremion:1999:HPG

[FAZ^{+}99] F. Frémion, M. Astier, S. Zaffran, A. Guillèn, V. Homburger, and M. Sémériva. The Heterotrimeric Protein G_{o} Is Required for the Formation of Heart Epithelium in Drosophila. Journal of Cell Biology, 145(5):1063–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/5/1063.

Fang:1996:NPQ

[FB96] J. Fang and R. M. Benbow. Nuclear proteins of quiescent Xenopus laevis cells inhibit DNA replication in intact and permeabilized nuclei. Journal of Cell Biology, 133(5):955–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
5/955.

Filardo:1995:RNM

[FBD^{+}95] E. J. Filardo, P. C. Brooks, S. L. Deming, C. Damsky, and D. A. Cheresh. Requirement of the NPXY motif in the integrin beta 3 subunit cytoplasmic tail for melanoma cell migration in vitro and in vivo. Journal of Cell Biology, 130(2):441–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/441.

Fawcett:1995:MHB

[FBH^{+}95] J. Fawcett, C. Buckley, C. L. Holness, I. N. Bird, J. H. Spragg, J. Saunders, A. Harris, and D. L. Simmons. Mapping the homotypic binding sites in CD31 and the role of CD31 adhesion in the formation of interendothelial cell contacts. Journal of Cell Biology, 128(6):1229–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/6/1229.

Fontoura:1999:CBP

[FBM99] Beatriz M. A. Fontoura, Günter Blobel, and Michael J. Matunis. A Conserved Biogenesis Pathway for Nucleoporins: Proteolytic Processing of a 186-Kilodalton Precursor Generates Nup98 and the Novel Nucleoporin, Nup96. Journal of Cell Biology, 144(6):1097–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/6/1097.

Fang:1995:TRL

[FC95] G. Fang and T. R. Cech. Telomerase RNA localized in the replication band and spherical subnuclear organelles in hypotrichous ciliates. Journal of Cell Biology, 130(2):243–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/2/243.

Foxman:1997:MNC

[FCB97] Ellen F. Foxman, James J. Campbell, and Eugene C. Butcher. Multistep Navigation and the Combinatorial Control of Leukocyte Chemotaxis. Journal of Cell Biology, 139(5):1349–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/5/1349.

Farschon:1997:TPA

[FCMN97] David M. Farschon, Clément Couture, Tomas Mustelin, and Donald D. Newmeyer. Temporal Phases in Apoptosis Defined by the Actions of Src Homology 2 Domains, Ceramide, Bcl-2, Interleukin-1β Converting Enzyme Family Proteases, and a Dense Membrane Fraction. Journal of Cell Biology, 137 (5):1117–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/5/1117.

Ferrary:1999:VVR

[FCTPA^{+}99] Evelyne Ferrary, Michel Cohen-Tannoudji, Gérard Pehau-Arnaudet, Alexandre Lapillonne, Rafika Athman, Tereza Ruiz, Lilia Boulouha, Fatima El Marjou, Anne Doye, Jean-Jacques Fontaine, Claude Antony, Charles Babinet, Daniel Louvard, Frédéric Jaisser, and Sylvie Robine. In Vivo, Villin Is Required for Ca^{2+} -Dependent F–Actin Disruption in Intestinal Brush Borders. Journal of Cell Biology, 146(4):819–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
4/819.

Flucher:1999:TTR

[FCTS99] Bernhard E. Flucher, Antonio Conti, Hiroshi Takeshima, and Vincenzo Sorrentino. Type 3 and Type 1 Ryanodine Receptors Are Localized in Triads of the Same Mammalian Skeletal Muscle Fibers. Journal of Cell Biology, 146(3):621–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
3/621.

Furuno:1999:HCR

[FdEP99] Nobuaki Furuno, Nicole den Elzen, and Jonathon Pines. Human Cyclin a Is Required for Mitosis until Mid Prophase. Journal of Cell Biology, 147(2):295–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/2/295.

Ferns:1996:AIA

[FDH96] M. Ferns, M. Deiner, and Z. Hall. Agrin-induced acetylcholine receptor clustering in mammalian muscle requires tyrosine phosphorylation. Journal of Cell Biology, 132(5):937–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
132/5/937.

Furukawa:1997:NAC

[FEF^{+}97] Katsutoshi Furukawa, Steven Estus, Weiming Fu, Robert J. Mark, and Mark P. Mattson. Neuroprotective Action of Cycloheximide Involves Induction of Bcl-2 and Antioxidant Pathways. Journal of Cell Biology, 136(5):1137–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
5/1137.

Felgner:1997:DNM

[FFB^{+}97] Harald Felgner, Rainer Frank, Jacek Biernat, Eva-Maria Mandelkow, Eckhard Mandelkow, Beat Ludin, Andrew Matus, and Manfred Schliwa. Domains of Neuronal Microtubuleassociated Proteins and Flexural Rigidity of Microtubules. Journal of Cell Biology, 138(5):1067–??, September 1997. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/138/5/
1067.

Fanchiotti:1998:UGG

[FFDP98] Sandra Fanchiotti, Fabiana Fernández, Cecilia D’Alessio, and Armando J. Parodi. The UDP–Glc: Glycoprotein glucosyltransferase is essential for Schizosaccharomyces pombe viability under conditions of extreme endoplasmic reticulum stress. Journal of Cell Biology, 143(3):625–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/625.

Fagotto:1996:BCA

[FFGG96] F. Fagotto, N. Funayama, U. Gluck, and B. M. Gumbiner. Binding to cadherins antagonizes the signaling activity of betacatenin during axis formation in Xenopus. Journal of Cell Biology, 132(6):1105–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/6/1105.

Furuse:1998:CNI

[FFH^{+}98] Mikio Furuse, Kohji Fujita, Takashi Hiiragi, Kazushi Fujimoto, and Shoichiro Tsukita. Claudin-1 and -2: Novel Integral Membrane Proteins Localizing at Tight Junctions with No Sequence Similarity to Occludin. Journal of Cell Biology, 141 (7):1539–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/7/1539.

Fraschini:1999:BYB

[FFLP99] Roberta Fraschini, Elisa Formenti, Giovanna Lucchini, and Simonetta Piatti. Budding Yeast Bub2 Is Localized at Spindle Pole Bodies and Activates the Mitotic Checkpoint via a Different Pathway from Mad2. Journal of Cell Biology, 145 (5):979–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/5/979.

Funayama:1995:EAI

[FFMG95] N. Funayama, F. Fagotto, P. McCrea, and B. M. Gumbiner. Embryonic axis induction by the armadillo repeat domain of beta-catenin: evidence for intracellular signaling. Journal of Cell Biology, 128(5):959–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/5/959.

FitzGerald:1995:PEM

[FFZ^{+}95] D. J. FitzGerald, C. M. Fryling, A. Zdanovsky, C. B. Saelinger, M. Kounnas, J. A. Winkles, D. Strickland, and S. Leppla. Pseudomonas exotoxin-mediated selection yields cells with altered expression of low-density lipoprotein receptorrelated protein. Journal of Cell Biology, 129(6):1533–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1533.

Friend:1996:MCR

[FGA^{+}96] D. S. Friend, N. Ghildyal, K. F. Austen, M. F. Gurish, R. Matsumoto, and R. L. Stevens. Mast cells that reside at different locations in the jejunum of mice infected with Trichinella spiralis exhibit sequential changes in their granule ultrastructure and chymase phenotype. Journal of Cell Biology, 135(1): 279–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/279.

Futter:1998:PMC

[FGA^{+}98] C. E. Futter, A. Gibson, E. H. Allchin, S. Maxwell, L. J. Ruddock, G. Odorizzi, D. Domingo, I. S. Trowbridge, and C. R. Hopkins. In Polarized MDCK Cells Basolateral Vesicles Arise from Clathrin-γ-adaptin-coated Domains on Endosomal Tubules. Journal of Cell Biology, 141(3):611–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
3/611.

Fares:1996:IDR

[FGP96] H. Fares, L. Goetsch, and J. R. Pringle. Identification of a developmentally regulated septin and involvement of the septins in spore formation in Saccharomyces cerevisiae. Journal of Cell Biology, 132(3):399–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/3/399.

Faust:1997:TDP

[FH97] Phyllis L. Faust and Mary E. Hatten. Targeted Deletion of the PEX2 Peroxisome Assembly Gene in Mice Provides a Model for Zellweger Syndrome, a Human Neuronal Migration Disorder. Journal of Cell Biology, 139(5):1293–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
5/1293.

Fahrenkrog:1998:MAY

[FHAP98] Birthe Fahrenkrog, Eduard C. Hurt, Ueli Aebi, and Nelly Panté. Molecular Architecture of the Yeast Nuclear Pore Complex: Localization of Nsp1p Subcomplexes. Journal of Cell Biology, 143(3):577–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/3/577.

Fanger:1995:CCM

[FHCL95] C. M. Fanger, M. Hoth, G. R. Crabtree, and R. S. Lewis. Characterization of T cell mutants with defects in capacitative calcium entry: genetic evidence for the physiological roles of CRAC channels. Journal of Cell Biology, 131(3): 655–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/3/655.

Fagotto:1999:DAI

[FhJZ^{+}99] François Fagotto, Eek hoon Jho, Li Zeng, Thomas Kurth, Thomas Joos, Christine Kaufmann, and Frank Costantini. Domains of Axin Involved in Protein–Protein Interactions, Wnt Pathway Inhibition, and Intracellular Localization. Journal of Cell Biology, 145(4):741–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/4/741.

Faundez:1997:ARF

[FHK97] Victor Faúndez, Jim-Tong Horng, and Regis B. Kelly. ADP Ribosylation Factor 1 Is Required for Synaptic Vesicle Budding in PC12 Cells. Journal of Cell Biology, 138(3):505–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/505.

Felding-Habermann:1997:SIL

[FHSM^{+}97] Brunhilde Felding-Habermann, Steve Silletti, Fang Mei, Chi-Hung Siu, Paul M. Yip, Peter C. Brooks, David A. Cheresh, Timothy E. O’Toole, Mark H. Ginsberg, and Anthony M. P. Montgomery. A Single Immunoglobulin-like Domain of the Human Neural Cell Adhesion Molecule L1 Supports Adhesion by Multiple Vascular and Platelet Integrins. Journal of Cell Biology, 139(6):1567–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/6/1567.

French:1996:FFL

[FHV^{+}96] L. E. French, M. Hahne, I. Viard, G. Radlgruber, R. Zanone, K. Becker, C. Müller, and J. Tschopp. Fas and Fas ligand in embryos and adult mice: ligand expression in several immuneprivileged tissues and coexpression in adult tissues characterized by apoptotic cell turnover. Journal of Cell Biology, 133 (2):335–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/2/335.

Fricker:1997:INM

[FHWV97] Mark Fricker, Michael Hollinshead, Nick White, and David Vaux. Interphase Nuclei of Many Mammalian Cell Types Contain Deep, Dynamic, Tubular Membrane-bound Invaginations of the Nuclear Envelope. Journal of Cell Biology, 136(3): 531–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/531.

Fukuda:1999:CTL

[FHY99] Minoru Fukuda, Nobuyoshi Hiraoka, and Jiunn-Chern Yeh. C-type Lectins and Sialyl Lewis X Oligosaccharides. Journal of Cell Biology, 147(3):467–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/3/467.

Franco:1997:PHP

[FIC^{+}97] Paola Franco, Ciro Iaccarino, Ferdinando Chiaradonna, Anna Brandazza, Carlo Iavarone, M. Rosaria Mastronicola, M. Luisa Nolli, and M. Patrizia Stoppelli. Phosphorylation of human pro–urokinase on ser^{138/303} impairs its receptor-dependent ability to promote myelomonocytic adherence and motility. Journal of Cell Biology, 137(3):779–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/3/779.

Franklin:1998:CNS

[FJ98] James L. Franklin and Eugene M. Johnson. Control of Neuronal Size Homeostasis by Trophic Factor–mediated Coupling of Protein Degradation to Protein Synthesis. Journal of Cell Biology, 142(5):1313–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/5/1313.

Frisen:1995:RWL

[FJT^{+}95] J. Frisén, C. B. Johansson, C. Török, M. Risling, and U. Lendahl. Rapid, widespread, and longlasting induction of nestin contributes to the generation of glial scar tissue after CNS injury. Journal of Cell Biology, 131(2):453–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
2/453.

Farkasovsky:1995:YNA

[FK95] M. Farkasovsky and H. Küntzel. Yeast Num1p associates with the mother cell cortex during S/G2 phase and affects microtubular functions. Journal of Cell Biology, 131(4):1003– ??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/4/1003.

Fan:1999:SHO

[FK99] Hongran Fan and Paul A. Khavari. Sonic Hedgehog Opposes Epithelial Cell Cycle Arrest. Journal of Cell Biology, 147(1): 71–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/1/71.

Foxman:1999:ICC

[FKB99] Ellen F. Foxman, Eric J. Kunkel, and Eugene C. Butcher. Integrating Conflicting Chemotactic Signals. Journal of Cell Biology, 147(3):577–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/147/3/577.

Feng:1999:GAC

[FKM^{+}99] Guoping Feng, Eric Krejci, Jordi Molgo, Jeanette M. Cunningham, Jean Massoulié, and Joshua R. Sanes. Genetic Analysis of Collagen Q: Roles in Acetylcholinesterase and Butyrylcholinesterase Assembly and in Synaptic Structure and Function. Journal of Cell Biology, 144(6):1349–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
6/1349.

Fujimura-Kamada:1997:NMA

[FKNM97] Konomi Fujimura-Kamada, Franklin J. Nouvet, and Susan Michaelis. A novel membrane-associated metalloprotease, Ste24p, is required for the first step of NH_{2}-terminal processing of the yeast a-factor precursor. Journal of Cell Biology, 136 (2):271–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/2/271.

Fukata:1998:AMB

[FKO^{+}98] Yuko Fukata, Kazushi Kimura, Noriko Oshiro, Hideyuki Saya, Yoshiharu Matsuura, and Kozo Kaibuchi. Association of the Myosin-binding Subunit of Myosin Phosphatase and Moesin: Dual Regulation of Moesin Phosphorylation by Rho-associated Kinase and Myosin Phosphatase. Journal of Cell Biology, 141 (2):409–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/2/409.

Firmbach-Kraft:1995:ANL

[FKS95] I. Firmbach-Kraft and R. Stick. Analysis of nuclear lamin isoprenylation in Xenopus oocytes: isoprenylation of lamin B3 precedes its uptake into the nucleus. Journal of Cell Biology, 129(1):17–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/1/17.

Ferrari:1997:DIC

[FKWP97] Giorgio Ferrari, Andrew M. Knight, Colin Watts, and Jean Pieters. Distinct Intracellular Compartments Involved in Invariant Chain Degradation and Antigenic Peptide Loading of Major Histocompatibility Complex (MHC) Class II Molecules. Journal of Cell Biology, 139(6):1433–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
6/1433.

Fang:1999:CAP

[FLA99] Li Fang, Sam W. Lee, and Stuart A. Aaronson. Comparative Analysis of P73 and P53 Regulation and Effector Functions. Journal of Cell Biology, 147(4):823–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
4/823.

Farrelly:1999:EMR

[FLO^{+}99] Nadia Farrelly, Yi-Ju Lee, Janine Oliver, Caroline Dive, and Charles H. Streuli. Extracellular Matrix Regulates Apoptosis in Mammary Epithelium through a Control on Insulin Signaling. Journal of Cell Biology, 144(6):1337–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
6/1337.

Fiedler:1995:AXN

[FLPS95] K. Fiedler, F. Lafont, R. G. Parton, and K. Simons. Annexin XIIIb: a novel epithelial specific annexin is implicated in vesicular traffic to the apical plasma membrane. Journal of Cell Biology, 128(6):1043–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/6/1043.

Fan:1995:PIP

[FLW^{+}95] Z. Fan, Y. Lu, X. Wu, A. DeBlasio, A. Koff, and J. Mendelsohn. Prolonged induction of p21Cip1/WAF1/CDK2/PCNA complex by epidermal growth factor receptor activation mediates ligand-induced A431 cell growth inhibition. Journal of Cell Biology, 131(1):235–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/1/235.

Fung:1998:HCP

[FMD^{+}98] Jennifer C. Fung, Wallace F. Marshall, Abby Dernburg, David A. Agard, and John W. Sedat. Homologous chromosome pairing in Drosophila melanogaster proceeds through multiple independent initiations. Journal of Cell Biology, 141(1):5–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/1/5.

Feng:1995:GPI

[FMJ^{+}95] Z. Feng, A. Marti, B. Jehn, H. J. Altermatt, G. Chicaiza, and R. Jaggi. Glucocorticoid and progesterone inhibit involution and programmed cell death in the mouse mammary gland. Journal of Cell Biology, 131(4):1095–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/1095.

Fry:1998:CNN

[FMM^{+}98] Andrew M. Fry, Thibault Mayor, Patrick Meraldi, York-Dieter Stierhof, Kayoko Tanaka, and Erich A. Nigg. C-Nap1, a Novel Centrosomal Coiled–Coil Protein and Candidate Substrate of the Cell Cycle–regulated Protein Kinase Nek2. Journal of Cell Biology, 141(7):1563–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/7/1563.

Foti:1997:NMC

[FMP^{+}97] Michelangelo Foti, Aram Mangasarian, Vincent Piguet, Daniel P. Lew, Karl-Heinz Krause, Didier Trono, and Jean-Louis Carpentier. Nef-mediated Clathrin-coated Pit Formation. Journal of Cell Biology, 139(1):37–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/139/1/37.

Finger:1998:SRE

[FN98] Fern P. Finger and Peter Novick. Spatial Regulation of Exocytosis: Lessons from Yeast. Journal of Cell Biology, 142 (3):609–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/3/609.

Feierbach:1999:ACD

[FNDS99] Becket Feierbach, Eva Nogales, Kenneth H. Downing, and Tim Stearns. Alf1p, a CLIP-170 Domain-containing Protein, Is Functionally and Physically Associated with α-Tubulin. Journal of Cell Biology, 144(1):113–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/1/113.

Fukasawa:1999:GEA

[FNH99] Masayoshi Fukasawa, Masahiro Nishijima, and Kentaro Hanada. Genetic evidence for ATP-dependent endoplasmic reticulum-to-Golgi apparatus trafficking of ceramide for sphingomyelin synthesis in Chinese hamster ovary cells. Journal of Cell Biology, 144(4):673–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/4/673.

Fukata:1999:PAR

[FOK^{+}99] Yuko Fukata, Noriko Oshiro, Nagatoki Kinoshita, Yoji Kawano, Yoichiro Matsuoka, Vann Bennett, Yoshiharu Matsuura, and Kozo Kaibuchi. Phosphorylation of Adducin by Rho–Kinase Plays a Crucial Role in Cell Motility. Journal of Cell Biology, 145(2):347–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/2/347.

Faulkner:1999:ZNZ

[FPF^{+}99] Georgine Faulkner, Alberto Pallavicini, Elide Formentin, Anna Comelli, Chiara Ievolella, Silvia Trevisan, Gladis Bortoletto, Paolo Scannapieco, Michela Salamon, Vincent Mouly, Giorgio Valle, and Gerolamo Lanfranchi. ZASP: a New Z-band Alternatively Spliced PDZ-motif Protein. Journal of Cell Biology, 146(2):465–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/2/465.

Futter:1996:MEC

[FPHH96] C. E. Futter, A. Pearse, L. J. Hewlett, and C. R. Hopkins. Multivesicular endosomes containing internalized EGF– EGF receptor complexes mature and then fuse directly with lysosomes. Journal of Cell Biology, 132(6):1011–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
6/1011.

Fassler:1995:LBI

[FPM^{+}95] R. Fässler, M. Pfaff, J. Murphy, A. A. Noegel, S. Johansson, R. Timpl, and R. Albrecht. Lack of beta 1 integrin gene in embryonic stem cells affects morphology, adhesion, and migration but not integration into the inner cell mass of blastocysts. Journal of Cell Biology, 128(5):979–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/979.

Ferreira:1997:SOL

[FPRL97] João Ferreira, Giovanni Paolella, Carlos Ramos, and Angus I. Lamond. Spatial organization of large-scale chromatin domains in the nucleus: a magnified view of single chromosome territories. Journal of Cell Biology, 139(7):1597–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
7/1597.

Feng:1995:RIR

[FPWN95] Y. Feng, B. Press, and A. Wandinger-Ness. Rab 7: an important regulator of late endocytic membrane traffic. Journal of Cell Biology, 131(6):1435–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1435.

Frank:1998:NRR

[FR98] Deborah J. Frank and Mark B. Roth. ncl-1 Is Required for the Regulation of Cell Size and Ribosomal RNA Synthesis in Caenorhabditis elegans. Journal of Cell Biology, 140(6): 1321–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/6/1321.

Ferrari:1998:CSC

[FRHS98] Michael B. Ferrari, Katharina Ribbeck, Donald J. Hagler, and Nicholas C. Spitzer. A Calcium Signaling Cascade Essential for Myosin Thick Filament Assembly in Xenopus Myocytes. Journal of Cell Biology, 141(6):1349–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/6/1349.

Fava:1999:H

[FRMLT^{+}99] Fabienne Fava, Brigitte Raynaud-Messina, Jeanne Leung-Tack, Laurent Mazzolini, Min Li, Jean Claude Guillemot, Didier Cachot, Yvette Tollon, Pascual Ferrara, and Michel Wright. Human 76p. Journal of Cell Biology, 147(4):857– ??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/4/857.

Faigle:1998:DPL

[FRT^{+}98] Wolfgang Faigle, Graça Raposo, Daniele Tenza, Valérie Pinet, Anne B. Vogt, Harald Kropshofer, Alain Fischer, Geneviève de Saint-Basile, and Sebastian Amigorena. Deficient Peptide Loading and MHC Class II Endosomal Sorting in a Human Genetic Immunodeficiency Disease: the Chediak–Higashi Syndrome. Journal of Cell Biology, 141(5):1121–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
5/1121.

French:1999:EHS

[FSA^{+}99] M. M. French, S. E. Smith, K. Akanbi, T. Sanford, J. Hecht, M. C. Farach-Carson, and D. D. Carson. Expression of the Heparan Sulfate Proteoglycan, Perlecan, during Mouse Embryogenesis and Perlecan Chondrogenic Activity In Vitro. Journal of Cell Biology, 145(5):1103–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/5/1103.

Franch:1995:IPF

[FSAP95] H. A. Franch, J. W. Shay, R. J. Alpern, and P. A. Preisig. Involvement of pRB family in TGF beta-dependent epithelial cell hypertrophy. Journal of Cell Biology, 129(1):245–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
1/245.

Falciola:1997:HMG

[FSC^{+}97] Luca Falciola, Fabio Spada, Sabina Calogero, Gernot Längst, Renate Voit, Ingrid Grummt, and Marco E. Bianchi. High Mobility Group 1 Protein Is Not Stably Associated with the Chromosomes of Somatic Cells. Journal of Cell Biology, 137 (1):19–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/1/19.

Famiglietti:1997:TRE

[FSDA97] Julie Famiglietti, Jing Sun, Horace M. DeLisser, and Steven M. Albelda. Tyrosine residue in exon 14 of the cytoplasmic domain of platelet endothelial cell adhesion molecule-1 (PECAM-1/CD31) regulates ligand binding specificity. Journal of Cell Biology, 138(6):1425–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/6/1425.

Furuse:1998:SGP

[FSFT98] Mikio Furuse, Hiroyuki Sasaki, Kazushi Fujimoto, and Shoichiro Tsukita. A Single Gene Product, Claudin-1 or -2, Reconstitutes Tight Junction Strands and Recruits Occludin in Fibroblasts. Journal of Cell Biology, 143(2):391–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
2/391.

Feijo:1999:GPT

[FSH^{+}99] J. A. Feijó, J. Sainhas, G. R. Hackett, J. G. Kunkel, and P. K. Hepler. Growing Pollen Tubes Possess a Constitutive Alkaline Band in the Clear Zone and a Growth-dependent Acidic Tip. Journal of Cell Biology, 144(3):483–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
3/483.

Friedman:1996:KSP

[FSHD96] D. B. Friedman, H. A. Sundberg, E. Y. Huang, and T. N. Davis. The 110-kD spindle pole body component of Saccharomyces cerevisiae is a phosphoprotein that is modified in a cell cycle-dependent manner. Journal of Cell Biology, 132 (5):903–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/903.

Fannon:1995:NCM

[FSIG^{+}95] A. M. Fannon, D. L. Sherman, G. Ilyina-Gragerova, P. J. Brophy, V. L. Friedrich, and D. R. Colman. Novel E-cadherin-mediated adhesion in peripheral nerve: Schwann cell architecture is stabilized by autotypic adherens junctions. Journal of Cell Biology, 129(1):189–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/1/189.

Fink:1999:MCI

[FSM^{+}99] Charles C. Fink, Boris Slepchenko, Ion I. Moraru, James Schaff, James Watras, and Leslie M. Loew. Morphological Control of Inositol–1,4,5–Trisphosphate–Dependent Signals. Journal of Cell Biology, 147(5):929–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
5/929.

Fialka:1996:EDC

[FSR^{+}96] I. Fialka, H. Schwarz, E. Reichmann, M. Oft, M. Busslinger, and H. Beug. The estrogen-dependent c–JunER protein causes a reversible loss of mammary epithelial cell polarity involving a destabilization of adherens junctions. Journal of Cell Biology, 132(6):1115–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/6/1115.

Furuse:1999:MIH

[FST99] Mikio Furuse, Hiroyuki Sasaki, and Shoichiro Tsukita. Manner of Interaction of Heterogeneous Claudin Species within and between Tight Junction Strands. Journal of Cell Biology, 147(4): 891–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/4/891.

Fischer:1995:ANA

[FT95] R. Fischer and W. E. Timberlake. Aspergillus nidulans apsA (anucleate primary sterigmata) encodes a coiled-coil protein required for nuclear positioning and completion of asexual development. Journal of Cell Biology, 128(4):485–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/485.

Fath:1997:MMS

[FTB97] Karl R. Fath, Gina M. Trimbur, and David R. Burgess. Molecular Motors and a Spectrin Matrix Associate with Golgi Membranes In Vitro. Journal of Cell Biology, 139(5):1169–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/5/1169.

Funakoshi:1996:ATP

[FTH96] T. Funakoshi, S. Takeda, and N. Hirokawa. Active transport of photoactivated tubulin molecules in growing axons revealed by a new electron microscopic analysis. Journal of Cell Biology, 133(6):1347–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/6/1347.

Fincham:1996:TSK

[FUB^{+}96] V. J. Fincham, M. Unlu, V. G. Brunton, J. D. Pitts, J. A. Wyke, and M. C. Frame. Translocation of Src kinase to the cell periphery is mediated by the actin cytoskeleton under the control of the Rho family of small G proteins. Journal of Cell Biology, 135(6):1551–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1551.

Frisch:1996:RJT

[FVKS96] S. M. Frisch, K. Vuori, D. Kelaita, and S. Sicks. A role for Jun–N–terminal kinase in anoikis; suppression by bcl-2 and crmA. Journal of Cell Biology, 135(5):1377–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
5/1377.

Frisch:1996:CAD

[FVRCH96] S. M. Frisch, K. Vuori, E. Ruoslahti, and P. Y. Chan-Hui. Control of adhesion-dependent cell survival by focal adhesion kinase. Journal of Cell Biology, 134(3):793–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
3/793.

Ferrari:1997:EAA

[FWBSO97] Davide Ferrari, Sebastian Wesselborg, Manuel K. A. Bauer, and Klaus Schulze-Osthoff. Extracellular ATP Activates Transcription Factor NF-κB through the P2Z Purinoreceptor by Selectively Targeting NF-κB p65 (RelA). Journal of Cell Biology, 139(7):1635–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/7/1635.

Fukushima:1996:HEA

[FWG^{+}96] T. Fukushima, T. K. Waddell, S. Grinstein, G. G. Goss, J. Orlowski, and G. P. Downey. Na^{+} /H^{+} exchange activity during phagocytosis in human neutrophils: role of Fcgamma receptors and tyrosine kinases. Journal of Cell Biology, 132(6): 1037–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/6/1037.

Fazeli:1996:ASM

[FWHW96] S. Fazeli, D. J. Wells, C. Hobbs, and F. S. Walsh. Altered secondary myogenesis in transgenic animals expressing the neural cell adhesion molecule under the control of a skeletal muscle alpha-actin promoter. Journal of Cell Biology, 135(1): 241–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/241.

Frazier:1998:PPY

[FWL^{+}98] Jennifer A. Frazier, Mei Lie Wong, Mark S. Longtine, John R. Pringle, Matthias Mann, Timothy J. Mitchison, and Christine Field. Polymerization of purified yeast septins: Evidence that organized filament arrays may not be required for septin function. Journal of Cell Biology, 143(3):737–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/737.

Fisk:1997:MAM

[FY97] Harold A. Fisk and Michael P. Yaffe. Mutational Analysis of Mdm1p Function in Nuclear and Mitochondrial Inheritance. Journal of Cell Biology, 138(3):485–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/3/485.

Fisk:1999:RUM

[FY99] Harold A. Fisk and Michael P. Yaffe. A Role for Ubiquitination in Mitochondrial Inheritance in Saccharomyces cerevisiae. Journal of Cell Biology, 145(6):1199–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/6/1199.

Foote:1998:PDL

[FZZN98] Cynthia I. Foote, Lan Zhou, Xing Zhu, and Bruce J. Nicholson. The Pattern of Disulfide Linkages in the Extracellular Loop Regions of Connexin 32 Suggests a Model for the Docking Interface of Gap Junctions. Journal of Cell Biology, 140(5): 1187–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/5/1187.

Gimenez-Abian:1995:PTI

[GACM^{+}95] J. F. Giménez-Abián, D. J. Clarke, A. M. Mullinger, C. S. Downes, and R. T. Johnson. A postprophase topoisomerase II-dependent chromatid core separation step in the formation of metaphase chromosomes. Journal of Cell Biology, 131 (1):7–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/7.

Gu:1997:FDC

[GAPG97] Feng Gu, Fernando Aniento, Robert G. Parton, and Jean Gruenberg. Functional Dissection of COP–I Subunits in the Biogenesis of Multivesicular Endosomes. Journal of Cell Biology, 139(5):1183–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/5/1183.

Goodson:1996:SLS

[GAW^{+}96] H. V. Goodson, B. L. Anderson, H. M. Warrick, L. A. Pon, and J. A. Spudich. Synthetic lethality screen identifies a novel yeast myosin I gene (MYO5): myosin I proteins are required for polarization of the actin cytoskeleton. Journal of Cell Biology, 133(6):1277–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/6/1277.

Gabathuler:1998:SAP

[GAZ^{+}98] Reinhard Gabathuler, Judie Alimonti, Qian-Jin Zhang, Gerassimos Kolaitis, Gregor Reid, and Wilfred A. Jefferies. Surrogate Antigen Processing Mediated by TAP-dependent Antigenic Peptide Secretion. Journal of Cell Biology, 140(1): 17–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/17.

Gregory:1998:KGR

[GB98] Stephen L. Gregory and Nicholas H. Brown. kakapo, a Gene Required for Adhesion Between and Within Cell Layers in Drosophila, Encodes a Large Cytoskeletal Linker Protein Related to Plectin and Dystrophin. Journal of Cell Biology, 143(5):1271–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/5/1271.

Grimm:1996:BRA

[GBBSO96] S. Grimm, M. K. Bauer, P. A. Baeuerle, and K. Schulze-Osthoff. Bcl-2 down-regulates the activity of transcription factor NF-kappaB induced upon apoptosis. Journal of Cell Biology, 134(1):13–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/1/13.

Guillaud:1998:SPR

[GBFL^{+}98] Laurent Guillaud, Christophe Bosc, Anne Fourest-Lieuvin, Eric Denarier, Fabienne Pirollet, Laurence Lafanechère, and Didier Job. STOP Proteins are Responsible for the High Degree of Microtubule Stabilization Observed in Neuronal Cells. Journal of Cell Biology, 142(1):167–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/1/167.

Geering:1996:OMK

[GBG^{+}96] K. Geering, A. Beggah, P. Good, S. Girardet, S. Roy, D. Schaer, and P. Jaunin. Oligomerization and maturation of Na,K–ATPase: functional interaction of the cytoplasmic NH2 terminus of the beta subunit with the alpha subunit. Journal of Cell Biology, 133(6):1193–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/6/1193.

Garcia-Blanco:1995:NSV

[GBMS95] M. A. Garcia-Blanco, D. D. Miller, and M. P. Sheetz. Nuclear spreads: I. Visualization of bipartite ribosomal RNA domains. Journal of Cell Biology, 128(1):15–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/128/1/15.

Govindan:1995:RMY

[GBN95] B. Govindan, R. Bowser, and P. Novick. The role of Myo2, a yeast class V myosin, in vesicular transport. Journal of Cell Biology, 128(6):1055–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/6/1055.

Goder:1999:GCI

[GBS99] Veit Goder, Christoph Bieri, and Martin Spiess. Glycosylation Can Influence Topogenesis of Membrane Proteins and Reveals Dynamic Reorientation of Nascent Polypeptides within the Translocon. Journal of Cell Biology, 147(2):257–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
2/257.

Gimond:1998:CLM

[GBvdN^{+}98] Clotilde Gimond, Christian Baudoin, Ronald van der Neut, Duco Kramer, Jero Calafat, and Arnoud Sonnenberg. CreloxP –mediated Inactivation of the α6A Integrin Splice Variant In Vivo: Evidence for a Specific Functional Role of α6A in Lymphocyte Migration but Not in Heart Development. Journal of Cell Biology, 143(1):253–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/1/253.

Gunsalus:1995:MTD

[GBW^{+}95] K. C. Gunsalus, S. Bonaccorsi, E. Williams, F. Verni, M. Gatti, and M. L. Goldberg. Mutations in twinstar, a Drosophila gene encoding a cofilin/ADF homologue, result in defects in centrosome migration and cytokinesis. Journal of Cell Biology, 131(5):1243–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/5/1243.

Gondre:1998:ANR

[GBW98] Marjorie Gondré, Patrick Burrola, and David E. Weinstein. Accelerated Nerve Regeneration Mediated by Schwann Cells Expressing a Mutant Form of the POU Protein SCIP. Journal of Cell Biology, 141(2):493–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/2/493.

Gigliotti:1998:NND

[GCA^{+}98] Silvia Gigliotti, Giuliano Callaini, Silvia Andone, Maria Giovanna Riparbelli, Roberto Pernas-Alonso, Gyula Hoffmann, Franco Graziani, and Carla Malva. Nup154, a New Drosophila Gene Essential for Male and Female Gametogenesis Is Related to the Nup155 Vertebrate Nucleoporin Gene. Journal of Cell Biology, 142(5):1195–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/5/1195.

Garcia-Castro:1997:IBG

[GCAHW97] Martín I. García-Castro, Robert Anderson, Janet Heasman, and Christopher Wylie. Interactions between Germ Cells and Extracellular Matrix Glycoproteins during Migration and Gonad Assembly in the Mouse Embryo. Journal of Cell Biology, 138(2):471–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/2/471.

Gaertig:1995:ALA

[GCB^{+}95] J. Gaertig, M. A. Cruz, J. Bowen, L. Gu, D. G. Pennock, and M. A. Gorovsky. Acetylation of lysine 40 in alpha-tubulin is not essential in Tetrahymena thermophila. Journal of Cell Biology, 129(5):1301–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/5/1301.

Gopalan:1997:NMM

[GCD97] Ganesan Gopalan, Clarence S. M. Chan, and Peter J. Donovan. A Novel Mammalian, Mitotic Spindle–associated Kinase Is Related to Yeast and Fly Chromosome Segregation Regulators. Journal of Cell Biology, 138(3):643–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/643.

Guan:1997:SIL

[GCE^{+}97] Xiaojun Guan, Benjamin F. Cravatt, George R. Ehring, James E. Hall, Dale L. Boger, Richard A. Lerner, and Norton B. Gilula. The Sleep-inducing Lipid Oleamide Deconvolutes Gap Junction Communication and Calcium Wave Transmission in Glial Cells. Journal of Cell Biology, 139(7):1785– ??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1785.

Galou:1996:DGF

[GCGE^{+}96] M. Galou, E. Colucci-Guyon, D. Ensergueix, J. L. Ridet, M. Gimenez y Ribotta, A. Privat, C. Babinet, and P. Dupouey. Disrupted glial fibrillary acidic protein network in astrocytes from vimentin knockout mice. Journal of Cell Biology, 133 (4):853–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/4/853.

Gama-Carvalho:1997:TUS

[GCKC^{+}97] Margarida Gama-Carvalho, Randy D. Krauss, Lijian Chiang, Juan Valcárcel, Michael R. Green, and Maria Carmo-Fonseca. Targeting of U2AF^{65} to Sites of Active Splicing in the Nucleus. Journal of Cell Biology, 137(5):975–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/5/975.

Giannini:1995:RRC

[GCM^{+}95] G. Giannini, A. Conti, S. Mammarella, M. Scrobogna, and V. Sorrentino. The ryanodine receptor/calcium channel genes are widely and differentially expressed in murine brain and peripheral tissues. Journal of Cell Biology, 128(5):893–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/893.

Giuffre:1997:MAA

[GCM^{+}97] Laura Giuffrè, Anne-Sophie Cordey, Natacha Monai, Yanik Tardy, Marc Schapira, and Olivier Spertini. Monocyte Adhesion to Activated Aortic Endothelium: Role of L–Selectin and Heparan Sulfate Proteoglycans. Journal of Cell Biology, 136 (4):945–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/4/945.

Gorbsky:1998:MAM

[GCM98] Gary J. Gorbsky, Rey-Huei Chen, and Andrew W. Murray. Microinjection of Antibody to Mad2 Protein into Mammalian Cells in Mitosis Induces Premature Anaphase. Journal of Cell Biology, 141(5):1193–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/5/1193.

Geimer:1998:NKP

[GCML98] Stefan Geimer, Judith Clees, Michael Melkonian, and Karl-Ferdinand Lechtreck. A Novel 95-kD Protein Is Located in a Linker between Cytoplasmic Microtubules and Basal Bodies in a Green Flagellate and Forms Striated Filaments In Vitro. Journal of Cell Biology, 140(5):1149–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/5/1149.

Guild:1997:AFC

[GCST97] Gregory M. Guild, Patricia S. Connelly, Michael K. Shaw, and Lewis G. Tilney. Actin filament cables in Drosophila nurse cells are composed of modules that slide passively past one another during dumping. Journal of Cell Biology, 138(4):783–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/4/783.

Gorlich:1997:NCR

[GDB^{+}97] Dirk Görlich, Marylena Dabrowski, F. Ralf Bischoff, Ulrike Kutay, Peer Bork, Enno Hartmann, Siegfried Prehn, and Elisa Izaurralde. A Novel Class of RanGTP Binding Proteins. Journal of Cell Biology, 138(1):65–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/1/65.

Gindhart:1998:KLC

[GDB^{+}98] Joseph G. Gindhart, Chand J. Desai, Sven Beushausen, Kai Zinn, and Lawrence S. B. Goldstein. Kinesin Light Chains Are Essential for Axonal Transport in Drosophila. Journal of Cell Biology, 141(2):443–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/2/443.

Gorsch:1995:CAN

[GDC95] L. C. Gorsch, T. C. Dockendorff, and C. N. Cole. A conditional allele of the novel repeat-containing yeast nucleoporin RAT7/ NUP159 causes both rapid cessation of mRNA export and reversible clustering of nuclear pore complexes. Journal of Cell Biology, 129(4):939–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/4/939.

Gaglio:1997:MSP

[GDC97] Tirso Gaglio, Mary A. Dionne, and Duane A. Compton. Mitotic Spindle Poles are Organized by Structural and Motor Proteins in Addition to Centrosomes. Journal of Cell Biology, 138(5):1055–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/5/1055.

Guinet:1996:DDP

[GDF^{+}96] F. Guinet, J. A. Dvorak, H. Fujioka, D. B. Keister, O. Muratova, D. C. Kaslow, M. Aikawa, A. B. Vaidya, and T. E. Wellems. A developmental defect in Plasmodium falciparum male gametogenesis. Journal of Cell Biology, 135(1):269– ??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/269.

Goode:1998:RCA

[GDL98] Bruce L. Goode, David G. Drubin, and Pekka Lappalainen. Regulation of the Cortical Actin Cytoskeleton in Budding Yeast by Twinfilin, a Ubiquitous Actin Monomer-sequestering Protein. Journal of Cell Biology, 142(3):723–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
3/723.

Griffiths:1999:CDI

[GDM^{+}99] Gareth J. Griffiths, Laurence Dubrez, Clive P. Morgan, Neil A. Jones, Jenna Whitehouse, Bernard M. Corfe, Caroline Dive, and John A. Hickman. Cell Damage-induced Conformational Changes of the Pro–Apoptotic Protein Bak In Vivo Precede the Onset of Apoptosis. Journal of Cell Biology, 144(5):903–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
144/5/903.

Gesemann:1995:ARA

[GDR95] M. Gesemann, A. J. Denzer, and M. A. Ruegg. Acetylcholine receptor-aggregating activity of agrin isoforms and mapping of the active site. Journal of Cell Biology, 128(4):625–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/4/625.

Gerald:1998:RDR

[GDTBD98] Noel Gerald, Jianwu Dai, H. Ping Ting-Beall, and Arturo De Lozanne. A Role for Dictyostelium RacE in Cortical Tension and Cleavage Furrow Progression. Journal of Cell Biology, 141 (2):483–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/2/483.

Gaynor:1997:CIA

[GE97] Erin C. Gaynor and Scott D. Emr. COPI-independent Anterograde Transport: Cargo-selective ER to Golgi Protein Transport in Yeast COPI Mutants. Journal of Cell Biology, 136(4): 789–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/4/789.

Gnessi:1995:TDI

[GEJ^{+}95] L. Gnessi, A. Emidi, E. A. Jannini, E. Carosa, M. Maroder, M. Arizzi, S. Ulisse, and G. Spera. Testicular development involves the spatiotemporal control of PDGFs and PDGF receptors gene expression and action. Journal of Cell Biology, 131(4):1105–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/1105.

Gimeno:1995:SEY

[GEK95] R. E. Gimeno, P. Espenshade, and C. A. Kaiser. SED4 encodes a yeast endoplasmic reticulum protein that binds Sec16p and participates in vesicle formation. Journal of Cell Biology, 131 (2):325–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/2/325.

Grandi:1995:NNP

[GEW^{+}95] P. Grandi, S. Emig, C. Weise, F. Hucho, T. Pohl, and E. C. Hurt. A novel nuclear pore protein Nup82p which specifically binds to a fraction of Nsp1p. Journal of Cell Biology, 130(6): 1263–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1263.

Gregorio:1995:MTF

[GF95] C. C. Gregorio and V. M. Fowler. Mechanisms of thin filament assembly in embryonic chick cardiac myocytes: tropomodulin requires tropomyosin for assembly. Journal of Cell Biology, 129(3):683–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/3/683.

Gough:1997:DSS

[GF97] Nancy R. Gough and Douglas M. Fambrough. Different Steady State Subcellular Distributions of the Three Splice Variants of Lysosome-associated Membrane Protein LAMP-2 Are Determined Largely by the COOH-terminal Amino Acid Residue. Journal of Cell Biology, 137(5):1161–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/5/1161.

Gibson:1998:SMR

[GFM^{+}98] A. Gibson, C. E. Futter, S. Maxwell, E. H. Allchin, M. Shipman, J.-P. Kraehenbuhl, D. Domingo, G. Odorizzi, I. S. Trowbridge, and C. R. Hopkins. Sorting Mechanisms Regulating Membrane Protein Traffic in the Apical Transcytotic Pathway of Polarized MDCK Cells. Journal of Cell Biology, 143 (1):81–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/1/81.

Geerts:1999:BIP

[GFN^{+}99] Dirk Geerts, Lionel Fontao, Mirjam G. Nievers, Roel Q. J. Schaapveld, Patricia E. Purkis, Grant N. Wheeler, E. Birgitte Lane, Irene M. Leigh, and Arnoud Sonnenberg. Binding of Integrin α6β4 to Plectin Prevents Plectin Association with F–Actin but Does Not Interfere with Intermediate Filament Binding. Journal of Cell Biology, 147(2):417–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
2/417.

Greber:1995:DCL

[GG95a] U. F. Greber and L. Gerace. Depletion of calcium from the lumen of endoplasmic reticulum reversibly inhibits passive diffusion and signal-mediated transport into the nucleus. Journal of Cell Biology, 128(1):5–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/1/5.

Gurland:1995:SDM

[GG95b] G. Gurland and G. G. Gundersen. Stable, detyrosinated microtubules function to localize vimentin intermediate filaments in fibroblasts. Journal of Cell Biology, 131(5):1275–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
5/1275.

Gonzalez-Garay:1996:ATL

[GGC96] M. L. Gonzalez-Garay and F. Cabral. alpha-tubulin limits its own synthesis: evidence for a mechanism involving translational repression. Journal of Cell Biology, 135(6):1525–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/6/1525.

Goetz:1997:IPS

[GGD^{+}97] Douglas J. Goetz, Daniel M. Greif, Han Ding, Raymond T. Camphausen, Steven Howes, Kenneth M. Comess, Karen R. Snapp, Geoffrey S. Kansas, and Francis W. Luscinskas. Isolated P-selectin Glycoprotein Ligand-1 Dynamic Adhesion to P- and E-selectin. Journal of Cell Biology, 137(2):509–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/2/509.

Gupta:1997:CIC

[GGK97] Malini C. Gupta, Patricia L. Graham, and James M. Kramer. Characterization of α1(IV) Collagen Mutations in Caenorhabditis elegans and the Effects of α1 and α2(IV) Mutations on Type IV Collagen Distribution. Journal of Cell Biology, 137 (5):1185–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/5/1185.

Goping:1998:RTB

[GGL^{+}98] Ing Swie Goping, Atan Gross, Josée N. Lavoie, Mai Nguyen, Ronald Jemmerson, Kevin Roth, Stanley J. Korsmeyer, and Gordon C. Shore. Regulated Targeting of BAX to Mitochondria. Journal of Cell Biology, 143(1):207–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
1/207.

Gillham:1999:IMG

[GGPG99] Helen Gillham, Matthew C. H. M. Golding, Rainer Pepperkok, and William J. Gullick. Intracellular Movement of Green Fluorescent Protein–Tagged Phosphatidylinositol 3-Kinase in Response to Growth Factor Receptor Signaling. Journal of Cell Biology, 146(4):869–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/4/869.

Goulielmos:1996:FPF

[GGR^{+}96] G. Goulielmos, F. Gounari, S. Remington, S. Müller, M. Häner, U. Aebi, and S. D. Georgatos. Filensin and phakinin form a novel type of beaded intermediate filaments and coassemble de novo in cultured cells. Journal of Cell Biology, 132(4):643–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/643.

Gorodinsky:1995:GAP

[GH95] A. Gorodinsky and D. A. Harris. Glycolipid-anchored proteins in neuroblastoma cells form detergent-resistant complexes without caveolin. Journal of Cell Biology, 129(3): 619–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/3/619.

Gross:1995:PBS

[GHF^{+}95] S. D. Gross, D. P. Hoffman, P. L. Fisette, P. Baas, and R. A. Anderson. A phosphatidylinositol 4,5-bisphosphate-sensitive casein kinase I alpha associates with synaptic vesicles and phosphorylates a subset of vesicle proteins. Journal of Cell Biology, 130(3):711–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/3/711.

Gallo:1995:SXO

[GHJJ95] C. J. Gallo, A. R. Hand, T. L. Jones, and L. A. Jaffe. Stimulation of Xenopus oocyte maturation by inhibition of the Gprotein alpha S subunit, a component of the plasma membrane and yolk platelet membranes. Journal of Cell Biology, 130(2):275–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/2/275.

Garfinkel:1996:FDP

[GHP^{+}96] S. Garfinkel, X. Hu, I. A. Prudovsky, G. A. McMahon, E. M. Kapnik, S. D. McDowell, and T. Maciag. FGF-1-dependent proliferative and migratory responses are impaired in senescent human umbilical vein endothelial cells and correlate with the inability to signal tyrosine phosphorylation of fibroblast growth factor receptor-1 substrates. Journal of Cell Biology, 134(3):783–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/3/783.

Graham:1998:AYV

[GHS98] Laurie A. Graham, Kathryn J. Hill, and Tom H. Stevens. Assembly of the yeast vacuolar H^{+}-ATPase occurs in the endoplasmic reticulum and requires a Vma12p/Vma22p assembly complex. Journal of Cell Biology, 142(1):39–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/142/1/39.

Gant:1999:RLP

[GHW99] Tracey Michele Gant, Crafford A. Harris, and Katherine L. Wilson. Roles of LAP2 Proteins in Nuclear Assembly and DNA Replication: Truncated LAP2β Proteins Alter Lamina Assembly, Envelope Formation, Nuclear Size, and DNA Replication Efficiency in Xenopus laevis Extracts. Journal of Cell Biology, 144(6):1083–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/6/1083.

Gadella:1995:OEG

[GJ95a] T. W. Gadella and T. M. Jovin. Oligomerization of epidermal growth factor receptors on A431 cells studied by time-resolved fluorescence imaging microscopy. A stereochemical model for tyrosine kinase receptor activation. Journal of Cell Biology, 129(6):1543–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1543.

Goto:1995:CDC

[GJ95b] K. Goto and C. H. Johnson. Is the cell division cycle gated by a circadian clock? The case of Chlamydomonas reinhardtii. Journal of Cell Biology, 129(4):1061–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/4/1061.

Gabriel:1998:TUG

[GJB^{+}98] Heinz-Dieter Gabriel, Dirk Jung, Christoph Bützler, Achim Temme, Otto Traub, Elke Winterhager, and Klaus Willecke. Transplacental Uptake of Glucose Is Decreased in Embryonic Lethal Connexin26-deficient Mice. Journal of Cell Biology, 140 (6):1453–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/6/1453.

Graham:1997:TIC

[GJW^{+}97] Patricia L. Graham, Jeffrey J. Johnson, Shaoru Wang, Marion H. Sibley, Malini C. Gupta, and James M. Kramer. Type IV Collagen Is Detectable in Most, but Not All, Basement Membranes of Caenorhabditis elegans and Assembles on Tissues That Do Not Express It. Journal of Cell Biology, 137 (5):1171–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/5/1171.

Grote:1996:EVF

[GK96] E. Grote and R. B. Kelly. Endocytosis of VAMP is facilitated by a synaptic vesicle targeting signal. Journal of Cell Biology, 132(4):537–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/537.

Gaidarov:1999:PAI

[GK99] Ibragim Gaidarov and James H. Keen. Phosphoinositide– Ap-2 interactions required for targeting to plasma membrane clathrin-coated pits. Journal of Cell Biology, 146(4): 755–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/755.

Gallicano:1998:DRE

[GKB^{+}98] G. Ian Gallicano, Panos Kouklis, Christoph Bauer, Mei Yin, Valeri Vasioukhin, Linda Degenstein, and Elaine Fuchs. Desmoplakin Is Required Early in Development for Assembly of Desmosomes and Cytoskeletal Linkage. Journal of Cell Biology, 143(7):2009–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/7/2009.

Goldman:1996:FIF

[GKC^{+}96] R. D. Goldman, S. Khuon, Y. H. Chou, P. Opal, and P. M. Steinert. The function of intermediate filaments in cell shape and cytoskeletal integrity. Journal of Cell Biology, 134(4): 971–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/971.

Gould:1996:PSP

[GKM^{+}96] S. J. Gould, J. E. Kalish, J. C. Morrell, J. Bjorkman, A. J. Urquhart, and D. I. Crane. Pex13p is an SH3 protein of the peroxisome membrane and a docking factor for the predominantly cytoplasmic PTs1 receptor. Journal of Cell Biology, 135 (1):85–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/85.

Ghiso:1999:TDI

[GKO99] Julio A. Aguirre Ghiso, Katherine Kovalski, and Liliana Ossowski. Tumor Dormancy Induced by Downregulation of Urokinase Receptor in Human Carcinoma Involves Integrin and MAPK Signaling. Journal of Cell Biology, 147(1):89– ??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/1/89.

Gebbink:1997:INP

[GKP^{+}97] Martijn F. B. G. Gebbink, Onno Kranenburg, Mieke Poland, Francis P. G. van Horck, Brahim Houssa, and Wouter H. Moolenaar. Identification of a Novel, Putative Rho-specific GDP/GTP Exchange Factor and a RhoA-binding Protein: Control of Neuronal Morphology. Journal of Cell Biology, 137 (7):1603–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/7/1603.

Gettner:1995:CBP

[GKR95] S. N. Gettner, C. Kenyon, and L. F. Reichardt. Characterization of beta pat-3 heterodimers, a family of essential integrin receptors in C. elegans. Journal of Cell Biology, 129 (4):1127–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/4/1127.

Gonzalez:1998:OGG

[GKS^{+}98] Alfonso Dueñas Gonzalez, Mitsunori Kaya, Wen Shi, Howard Song, Joseph R. Testa, Linda Z. Penn, and Jorge Filmus. OCI-5/ GPC3, a Glypican Encoded by a Gene That Is Mutated in the Simpson–Golabi–Behmel Overgrowth Syndrome, Induces Apoptosis in a Cell Line–specific Manner. Journal of Cell Biology, 141(6):1407–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/6/1407.

Gammie:1995:DDR

[GKVR95] A. E. Gammie, L. J. Kurihara, R. B. Vallee, and M. D. Rose. DNM1, a dynamin-related gene, participates in endosomal trafficking in yeast. Journal of Cell Biology, 130(3): 553–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/3/553.

Geng:1995:IIC

[GL95] Y. Geng and M. Lotz. Increased intracellular Ca^{2+} selectively suppresses IL-1-induced NO production by reducing iNOS mRNA stability. Journal of Cell Biology, 129(6):1651–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1651.

Gratzer:1995:MNR

[GLB^{+}95] S. Gratzer, T. Lithgow, R. E. Bauer, E. Lamping, F. Paltauf, S. D. Kohlwein, V. Haucke, T. Junne, G. Schatz, and M. Horst. Mas37p, a novel receptor subunit for protein import into mitochondria. Journal of Cell Biology, 129(1):25–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/1/25.

Gao:1996:TMA

[GLD^{+}96] A. G. Gao, F. P. Lindberg, J. M. Dimitry, E. J. Brown, and W. A. Frazier. Thrombospondin modulates alpha v beta 3 function through integrin-associated protein. Journal of Cell Biology, 135(2):533–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/135/2/533.

Gotta:1996:CTC

[GLF^{+}96] M. Gotta, T. Laroche, A. Formenton, L. Maillet, H. Scherthan, and S. M. Gasser. The clustering of telomeres and colocalization with Rap1, Sir3, and Sir4 proteins in wild-type Saccharomyces cerevisiae. Journal of Cell Biology, 134(6):1349– ??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1349.

Gillece:1999:ECF

[GLL^{+}99] Pauline Gillece, José Manuel Luz, William J. Lennarz, Francisco Javier de la Cruz, and Karin Römisch. Export of a Cysteine–Free Misfolded Secretory Protein from the Endoplasmic Reticulum for Degradation Requires Interaction with Protein Disulfide Isomerase. Journal of Cell Biology, 147(7): 1443–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/7/1443.

Gradin:1998:RMD

[GLMG98] Helena Melander Gradin, Niklas Larsson, Ulrica Marklund, and Martin Gullberg. Regulation of Microtubule Dynamics by Extracellular Signals: cAMP-dependent Protein Kinase Switches Off the Activity of Oncoprotein 18 in Intact Cells. Journal of Cell Biology, 140(1):131–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/1/131.

Godyna:1995:ILD

[GLP^{+}95] S. Godyna, G. Liau, I. Popa, S. Stefansson, and W. S. Argraves. Identification of the low density lipoprotein receptor-related protein (LRP) as an endocytic receptor for thrombospondin-1. Journal of Cell Biology, 129(5):1403–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/5/1403.

Gauen:1996:MFP

[GLS96] L. K. Timson Gauen, M. E. Linder, and A. S. Shaw. Multiple features of the p59fyn src homology 4 domain define a motif for immune-receptor tyrosine-based activation motif (ITAM) binding and for plasma membrane localization. Journal of Cell Biology, 133(5):1007–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/5/1007.

Li:1999:FRR

[gLSE^{+}99] Min gang Li, Madeline Serr, Kevin Edwards, Susan Ludmann, Daisuke Yamamoto, Lewis G. Tilney, Christine M. Field, and Thomas S. Hays. Filamin is required for ring canal assembly and actin organization during Drosophila oogenesis. Journal of Cell Biology, 146(5):1061–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/5/1061.

Ghosh:1995:ENR

[GM95] R. N. Ghosh and F. R. Maxfield. Evidence for nonvectorial, retrograde transferrin trafficking in the early endosomes of HEp2 cells. Journal of Cell Biology, 128(4):549–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/549.

Gotoh:1999:AID

[GM99] Tomomi Gotoh and Masataka Mori. Arginase II Downregulates Nitric Oxide (NO) Production and Prevents NOmediated Apoptosis in Murine Macrophage-derived RAW 264.7 Cells. Journal of Cell Biology, 144(3):427–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
3/427.

Garcia-Mata:1999:CDA

[GMBSS99] Rafael García-Mata, Zsuzsa Bebök, Eric J. Sorscher, and Elizabeth S. Sztul. Characterization and dynamics of aggresome formation by a cytosolic Gfp–chimera. Journal of Cell Biology, 146(6):1239–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/6/1239.

Garcia:1995:RBC

[GMC^{+}95] E. P. Garcia, P. S. McPherson, T. J. Chilcote, K. Takei, and P. De Camilli. rbSec1A and B colocalize with syntaxin 1 and SNAP-25 throughout the axon, but are not in a stable complex with syntaxin. Journal of Cell Biology, 129(1):105–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
1/105.

Golsteyn:1995:CCR

[GMFN95] R. M. Golsteyn, K. E. Mundt, A. M. Fry, and E. A. Nigg. Cell cycle regulation of the activity and subcellular localization of Plk1, a human protein kinase implicated in mitotic spindle function. Journal of Cell Biology, 129(6):1617–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1617.

Gangwani:1998:IZT

[GMGGD98] Laxman Gangwani, Monique Mikrut, Zoya Galcheva-Gargova, and Roger J. Davis. Interaction of ZPR1 with Translation Elongation Factor-1α in Proliferating Cells. Journal of Cell Biology, 143(6):1471–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/6/1471.

Glickman:1996:BMC

[GMS^{+}96] J. N. Glickman, P. A. Morton, J. W. Slot, S. Kornfeld, and H. J. Geuze. The biogenesis of the MHC class II compartment in human I-cell disease B lymphoblasts. Journal of Cell Biology, 132(5):769–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/5/769.

Grady:1997:SND

[GMS97] R. Mark Grady, John P. Merlie, and Joshua R. Sanes. Subtle Neuromuscular Defects in Utrophin-deficient Mice. Journal of Cell Biology, 136(4):871–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/4/871.

Ghosh:1998:ETC

[GMS^{+}98] Richik N. Ghosh, William G. Mallet, Thwe T. Soe, Timothy E. McGraw, and Frederick R. Maxfield. An Endocytosed TGN38 Chimeric Protein Is Delivered to the TGN after Trafficking through the Endocytic Recycling Compartment in CHO Cells. Journal of Cell Biology, 142(4):923–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
4/923.

Goldstein:1995:CCO

[Gol95] B. Goldstein. Cell contacts orient some cell division axes in the Caenorhabditis elegans embryo. Journal of Cell Biology, 129 (4):1071–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/4/1071.

Glover:1996:PKC

[GOT96] D. M. Glover, H. Ohkura, and A. Tavares. Polo kinase: the choreographer of the mitotic stage? Journal of Cell Biology, 135(6):1681–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1681.

Gonczy:1999:CDR

[GPKH99] Pierre Gönczy, Silke Pichler, Matthew Kirkham, and Anthony A. Hyman. Cytoplasmic Dynein Is Required for Distinct Aspects of Mtoc Positioning, Including Centrosome Separation, in the One Cell Stage Caenorhabditis elegans Embryo. Journal of Cell Biology, 147(1):135–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/1/135.

Guinebault:1995:IDT

[GPRS^{+}95] C. Guinebault, B. Payrastre, C. Racaud-Sultan, H. Mazarguil, M. Breton, G. Mauco, M. Plantavid, and H. Chap. Integrindependent translocation of phosphoinositide 3-kinase to the cytoskeleton of thrombin-activated platelets involves specific interactions of p85 alpha with actin filaments and focal adhesion kinase. Journal of Cell Biology, 129(3):831–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/831.

Giordano:1999:MDG

[GPSF99] Ennio Giordano, Ivana Peluso, Stefania Senger, and Maria Furia. minifly, A Drosophila Gene Required for Ribosome Biogenesis. Journal of Cell Biology, 144(6):1123–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
6/1123.

Gao:1997:CSC

[GR97] Fen-Biao Gao and Martin Raff. Cell Size Control and a Cellintrinsic Maturation Program in Proliferating Oligodendrocyte Precursor Cells. Journal of Cell Biology, 138(6):1367– ??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1367.

Geng:1997:EPS

[GRB^{+}97] Jian-Guo Geng, Thomas J. Raub, Carolyn A. Baker, Geri A. Sawada, Li Ma, and Åke P. Elhammer. Expression of a P-selectin ligand in Zona Pellucida of porcine oocytes and P-selectin on acrosomal membrane of porcine sperm cells. Potential implications for their involvement in sperm–egg interactions. Journal of Cell Biology, 137(3):743–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
3/743.

Girzalsky:1999:IPP

[GRS^{+}99] Wolfgang Girzalsky, Peter Rehling, Katharina Stein, Julia Kipper, Lars Blank, Wolf-Hubert Kunau, and Ralf Erdmann. Involvement of Pex13p in Pex14p Localization and Peroxisomal Targeting Signal 2-dependent Protein Import into Peroxisomes. Journal of Cell Biology, 144(6):1151–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
6/1151.

Garver:1997:TPS

[GRTB97] Timothy D. Garver, Qun Ren, Shmuel Tuvia, and Vann Bennett. Tyrosine Phosphorylation at a Site Highly Conserved in the L1 Family of Cell Adhesion Molecules Abolishes Ankyrin Binding and Increases Lateral Mobility of Neurofascin. Journal of Cell Biology, 137(3):703–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/3/703.

Galbraith:1999:KPS

[GS99] Catherine G. Galbraith and Michael P. Sheetz. Keratocytes Pull with Similar Forces on Their Dorsal and Ventral Surfaces. Journal of Cell Biology, 147(6):1313–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1313.

Gaglio:1996:OMA

[GSB^{+}96] T. Gaglio, A. Saredi, J. B. Bingham, M. J. Hasbani, S. R. Gill, T. A. Schroer, and D. A. Compton. Opposing motor activities are required for the organization of the mammalian mitotic spindle pole. Journal of Cell Biology, 135(2):399–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
2/399.

Green:1997:RGP

[GSB97] Jennifer M. Green, Alan D. Schreiber, and Eric J. Brown. Role for a Glycan Phosphoinositol Anchor in Fcγ Receptor Synergy. Journal of Cell Biology, 139(5):1209–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
5/1209.

Gaglio:1995:NRO

[GSC95] T. Gaglio, A. Saredi, and D. A. Compton. NuMA is required for the organization of microtubules into aster-like mitotic arrays. Journal of Cell Biology, 131(3):693–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/693.

Gotz:1996:TCC

[GSC^{+}96] B. Götz, A. Scholze, A. Clement, A. Joester, K. Schütte, F. Wigger, R. Frank, E. Spiess, P. Ekblom, and A. Faissner. Tenascin-c contains distinct adhesive, anti-adhesive, and neurite outgrowth promoting sites for neurons. Journal of Cell Biology, 132(4):681–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/681.

Goldfinger:1998:PLF

[GSJ98] Lawrence E. Goldfinger, M. Sharon Stack, and Jonathan C. R. Jones. Processing of Laminin-5 and Its Functional Consequences: Role of Plasmin and Tissue-type Plasminogen Activator. Journal of Cell Biology, 141(1):255–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
1/255.

Gonczy:1999:DCD

[GSK^{+}99] Pierre Gönczy, Heinke Schnabel, Titus Kaletta, Ana Duran Amores, Tony Hyman, and Ralf Schnabel. Dissection of Cell Division Processes in the One Cell Stage Caenorhabditis elegans Embryo by Mutational Analysis. Journal of Cell Biology, 144(5):927–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/5/927.

Gijon:1999:RPS

[GSKL99] Miguel A. Gijón, Diane M. Spencer, Alan L. Kaiser, and Christina C. Leslie. Role of Phosphorylation Sites and the C2 Domain in Regulation of Cytosolic Phospholipase A_{2}. Journal of Cell Biology, 145(6):1219–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/6/1219.

Gow:1998:DPP

[GSL98] Alexander Gow, Cherie M. Southwood, and Robert A. Lazzarini. Disrupted Proteolipid Protein Trafficking Results in Oligodendrocyte Apoptosis in an Animal Model of Pelizaeus– Merzbacher Disease. Journal of Cell Biology, 140(4):925– ??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/4/925.

Guttinger:1998:EVL

[GSP^{+}98] Maria Guttinger, Francesca Sutti, Maddalena Panigada, Simona Porcellini, Barbara Merati, Margherita Mariani, Tambet Teesalu, G. Giacomo Consalez, and Fabio Grassi. Epithelial V-like Antigen (EVA), a Novel Member of the Immunoglobulin Superfamily, Expressed in Embryonic Epithelia with a Potential Role as Homotypic Adhesion Molecule in Thymus Histogenesis. Journal of Cell Biology, 141(4):1061–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
4/1061.

Gunthert:1996:EDC

[GSvR^{+}96] A. R. Günthert, J. Sträter, U. von Reyher, C. Henne, S. Joos, K. Koretz, G. Moldenhauer, P. H. Krammer, and P. Möller. Early detachment of colon carcinoma cells during CD95(APO-1/Fas)-mediated apoptosis. I. De-adhesion from hyaluronate by shedding of CD44. Journal of Cell Biology, 134(4): 1089–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/1089.

Gregorio:1998:NTT

[GTC^{+}98] Carol C. Gregorio, Karoly Trombitás, Thomas Centner, Bernhard Kolmerer, Gunter Stier, Kathleen Kunke, Koichi Suzuki, Franz Obermayr, Bernhard Herrmann, Henk Granzier, Hiroyuki Sorimachi, and Siegfried Labeit. The NH_{2} Terminus of Titin Spans the Z–Disc: Its Interaction with a Novel 19-kD Ligand (T-cap) Is Required for Sarcomeric Integrity. Journal of Cell Biology, 143(4):1013–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/4/1013.

Gogos:1996:GTD

[GTL^{+}96] J. A. Gogos, R. Thompson, W. Lowry, B. F. Sloane, H. Weintraub, and M. Horwitz. Gene trapping in differentiating cell lines: regulation of the lysosomal protease cathepsin B in skeletal myoblast growth and fusion. Journal of Cell Biology, 134 (4):837–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/837.

Gu:1999:SFD

[GTP^{+}99] Jianguo Gu, Masahito Tamura, Roumen Pankov, Erik H. J. Danen, Takahisa Takino, Kazue Matsumoto, and Kenneth M. Yamada. Shc and Fak Differentially Regulate Cell Motility and Directionality Modulated by Pten. Journal of Cell Biology, 146(2):389–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/2/389.

Gu:1998:TSP

[GTY98] Jianguo Gu, Masahito Tamura, and Kenneth M. Yamada. Tumor Suppressor PTEN Inhibits Integrin- and Growth Factor– mediated Mitogen-activated Protein (MAP) Kinase Signaling Pathways. Journal of Cell Biology, 143(5):1375–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1375.

Gimond:1999:ICS

[GvdFvD^{+}99] Clotilde Gimond, Arjan van der Flier, Sanne van Delft, Cord Brakebusch, Ingrid Kuikman, John G. Collard, Reinhard Fässler, and Arnoud Sonnenberg. Induction of Cell Scattering by Expression of β1 Integrins in β1-Deficient Epithelial Cells Requires Activation of Members of the Rho Family of Gtpases and Downregulation of Cadherin and Catenin Function. Journal of Cell Biology, 147(6):1325–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1325.

Goeckeler:1995:MLC

[GW95] Z. M. Goeckeler and R. B. Wysolmerski. Myosin light chain kinase-regulated endothelial cell contraction: the relationship between isometric tension, actin polymerization, and myosin phosphorylation. Journal of Cell Biology, 130(3):613–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/3/613.

Gary:1998:FEP

[GWB^{+}98] Jonathan D. Gary, Andrew E. Wurmser, Cecilia J. Bonangelino, Lois S. Weisman, and Scott D. Emr. Fab1p Is Essential for PtdIns(3)P 5-Kinase Activity and the Maintenance of Vacuolar Size and Membrane Homeostasis. Journal of Cell Biology, 143(1):65–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/1/65.

Goode:1999:CPR

[GWB^{+}99] Bruce L. Goode, Jonathan J. Wong, Anne-Christine Butty, Matthias Peter, Ashley L. McCormack, John R. Yates, David G. Drubin, and Georjana Barnes. Coronin Promotes the Rapid Assembly and Cross-linking of Actin Filaments and May Link the Actin and Microtubule Cytoskeletons in Yeast. Journal of Cell Biology, 144(1):83–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/144/1/83.

Gilboa:1998:OST

[GWLH98] Lilach Gilboa, Rebecca G. Wells, Harvey F. Lodish, and Yoav I. Henis. Oligomeric Structure of Type I and Type II Transforming Growth Factor β Receptors: Homodimers Form in the ER and Persist at the Plasma Membrane. Journal of Cell Biology, 140(4):767–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/4/767.

Gu:1999:ARP

[GXE^{+}99] Minyi Gu, Xiaodong Xi, Graham D. Englund, Michael C. Berndt, and Xiaoping Du. Analysis of the roles of 14-3-3 in the platelet glycoprotein Ib–IX–Mediated activation of integrin α_{IIb} β_{3} using a reconstituted mammalian cell expression model. Journal of Cell Biology, 147(5):1085–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
5/1085.

Gutierrez:1997:FEC

[GYRWR97] Jesus A. Gutierrez, Jianming Yu, Susan Rivera, and Marianne Wessling-Resnick. Functional Expression Cloning and Characterization of SFT, a Stimulator of Fe Transport. Journal of Cell Biology, 139(4):895–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/4/895.

Goodearl:1995:ACS

[GYS^{+}95] A. D. Goodearl, A. G. Yee, A. W. Sandrock, G. Corfas, and G. D. Fischbach. ARIA is concentrated in the synaptic basal lamina of the developing chick neuromuscular junction. Journal of Cell Biology, 130(6):1423–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/6/1423.

Green:1999:RCF

[GZC^{+}99] Jennifer M. Green, Alexander Zheleznyak, Jun Chung, Frederik P. Lindberg, Marika Sarfati, William A. Frazier, and Eric J. Brown. Role of Cholesterol in Formation and Function of a Signaling Complex Involving α vβ3, Integrin–Associated Protein (Cd47), and Heterotrimeric G Proteins. Journal of Cell Biology, 146(3):673–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/3/673.

Gebbink:1995:CSE

[GZK^{+}95] M. F. Gebbink, G. C. Zondag, G. M. Koningstein, E. Feiken, R. W. Wubbolts, and W. H. Moolenaar. Cell surface expression of receptor protein tyrosine phosphatase RPTP mu is regulated by cell–cell contact. Journal of Cell Biology, 131(1): 251–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/251.

Hermann:1999:VRA

[HAB^{+}99] P. Hermann, M. Armant, E. Brown, M. Rubio, H. Ishihara, D. Ulrich, R. G. Caspary, F. P. Lindberg, R. Armitage, C. Maliszewski, G. Delespesse, and M. Sarfati. The Vitronectin Receptor and its Associated CD47 Molecule Mediates Proinflammatory Cytokine Synthesis in Human Monocytes by Interaction with Soluble CD23. Journal of Cell Biology, 144(4):

767–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/4/767.

Henry:1995:MDR

[HAS95] M. D. Henry, C. Gonzalez Agosti, and F. Solomon. Molecular dissection of radixin: distinct and interdependent functions of the amino- and carboxy-terminal domains. Journal of Cell Biology, 129(4):1007–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/4/1007.

Henderson:1996:SAC

[HAS^{+}96] S. Henderson, R. Allsopp, D. Spector, S. S. Wang, and C. Harley. In situ analysis of changes in telomere size during replicative aging and cell transformation. Journal of Cell Biology, 134(1):1–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/1/1.

Harsay:1995:PSP

[HB95a] E. Harsay and A. Bretscher. Parallel secretory pathways to the cell surface in yeast. Journal of Cell Biology, 131(2): 297–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/2/297.

Hurwitz:1995:NEY

[HB95b] M. E. Hurwitz and G. Blobel. NUP82 is an essential yeast nucleoporin required for poly(A)+ RNA export. Journal of Cell Biology, 130(6):1275–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/6/1275.

Heidemann:1998:CCF

[HB98] Steven R. Heidemann and Robert E. Buxbaum. Cell crawling: First the motor, now the transmission. Journal of Cell Biology, 141(1):1–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/1.

Holly:1999:PFK

[HB99] Stephen P. Holly and Kendall J. Blumer. Pak-family Kinases Regulate Cell and Actin Polarization Throughout the Cell Cycle of Saccharomyces cerevisiae. Journal of Cell Biology, 147(4):845–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/4/845.

Hagg:1998:CFM

[HBH98] Rupert Hagg, Peter Bruckner, and Erik Hedbom. Cartilage Fibrils of Mammals are Biochemically Heterogeneous: Differential Distribution of Decorin and Collagen IX. Journal of Cell Biology, 142(1):285–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/1/285.

Hopkinson:1995:MGS

[HBJ95] S. B. Hopkinson, S. E. Baker, and J. C. Jones. Molecular genetic studies of a human epidermal autoantigen (the 180-kd bullous pemphigoid antigen/bp180): identification of functionally important sequences within the bp180 molecule and evidence for an interaction between bp180 and alpha 6 integrin. Journal of Cell Biology, 130(1):117–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/1/117.

Hauser:1998:GUC

[HBPJ98] Hans-Peter Hauser, Michael Bardroff, George Pyrowolakis, and Stefan Jentsch. A Giant Ubiquitin-conjugating Enzyme Related to IAP Apoptosis Inhibitors. Journal of Cell Biology, 141(6):1415–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/6/1415.

Halverson:1997:CDB

[HBS^{+}97] Dana Halverson, Mary Baum, Janet Stryker, John Carbon, and Louise Clarke. A Centromere DNA-binding Protein from Fission Yeast Affects Chromosome Segregation and Has Homology to Human CENP–B. Journal of Cell Biology, 136(3): 487–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/487.

Huang:1997:INV

[HC97] Pei-Hsin Huang and Hui-Ling Chiang. Identification of Novel Vesicles in the Cytosol to Vacuole Protein Degradation Pathway. Journal of Cell Biology, 136(4):803–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/4/
803.

Hart:1999:VIA

[HC99] Marilyn C. Hart and John A. Cooper. Vertebrate Isoforms of Actin Capping Protein β Have Distinct Functions in Vivo. Journal of Cell Biology, 147(6):1287–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1287.

Heath:1995:NPC

[HCA^{+}95] C. V. Heath, C. S. Copeland, D. C. Amberg, V. Del Priore, M. Snyder, and C. N. Cole. Nuclear pore complex clustering and nuclear accumulation of poly(A)+ RNA associated with mutation of the Saccharomyces cerevisiae RAT2/ NUP120 gene. Journal of Cell Biology, 131(6):1677–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1677.

Heil-Chapdelaine:1999:FCB

[HCAC99] Richard A. Heil-Chapdelaine, Neil R. Adames, and John A. Cooper. Formin’ the Connection between Microtubules and the Cell Cortex. Journal of Cell Biology, 144(5):809–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
5/809.

Hyman:1995:SCA

[HCAW95] A. A. Hyman, D. Chrétien, I. Arnal, and R. H. Wade. Structural changes accompanying GTP hydrolysis in microtubules: information from a slowly hydrolyzable analogue guanylyl- (alpha,beta)-methylene-diphosphonate. Journal of Cell Biology, 128(1):117–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/1/117.

Hofmann:1998:SCD

[HCG^{+}98] Christian Hofmann, Iain M. Cheeseman, Bruce L. Goode, Kent L. McDonald, Georjana Barnes, and David G. Drubin. Saccharomyces cerevisiae Duo1p and Dam1p, Novel Proteins Involved in Mitotic Spindle Function. Journal of Cell Biology, 143(4):1029–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/4/1029.

Hughes:1999:MIS

[HCLG99] Simon M. Hughes, Maggie M.-Y. Chi, Oliver H. Lowry, and Kristian Gundersen. Myogenin Induces a Shift of Enzyme Activity from Glycolytic to Oxidative Metabolism in Muscles of Transgenic Mice. Journal of Cell Biology, 145(3):633–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
3/633.

Hopmann:1996:AOB

[HCM96a] R. Hopmann, J. A. Cooper, and K. G. Miller. Actin organization, bristle morphology, and viability are affected by actin capping protein mutations in Drosophila. Journal of Cell Biology, 133(6):1293–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/6/1293.

Hungerford:1996:IPC

[HCM^{+}96b] J. E. Hungerford, M. T. Compton, M. L. Matter, B. G. Hoffstrom, and C. A. Otey. Inhibition of pp125FAK in cultured fibroblasts results in apoptosis. Journal of Cell Biology, 135(5): 1383–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/5/1383.

Henry:1997:TDP

[HCM^{+}97] Ralph Henry, Matthew Carrigan, Michael McCaffery, Xianyue Ma, and Kenneth Cline. Targeting Determinants and Proposed Evolutionary Basis for the Sec and the Delta pH Protein Transport Systems in Chloroplast Thylakoid Membranes. Journal of Cell Biology, 136(4):823–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
4/823.

Chen:1997:CTO

[hCMPG97] Yan hua Chen, Christa Merzdorf, David L. Paul, and Daniel A. Goodenough. COOH Terminus of Occludin Is Required for Tight Junction Barrier Function in Early Xenopus Embryos. Journal of Cell Biology, 138(4):891–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/4/891.

Haug-Collet:1999:CCP

[HCPW^{+}99] K. Haug-Collet, B. Pearson, R. Webel, R. T. Szerencsei, R. J. Winkfein, P. P. M. Schnetkamp, and N. J. Colley. Cloning and characterization of a potassium-dependent sodium/calcium exchanger in Drosophila. Journal of Cell Biology, 147(3):659– ??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/3/659.

Hinchcliffe:1998:CCR

[HCRS98] Edward H. Hinchcliffe, Grizzel O. Cassels, Conly L. Rieder, and Greenfield Sluder. The Coordination of Centrosome Reproduction with Nuclear Events of the Cell Cycle in the Sea Urchin Zygote. Journal of Cell Biology, 140(6):1417–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1417.

Hill:1996:AMF

[HCW96] K. L. Hill, N. L. Catlett, and L. S. Weisman. Actin and myosin function in directed vacuole movement during cell division in Saccharomyces cerevisiae. Journal of Cell Biology, 135(6): 1535–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1535.

Huang:1998:GJM

[HCW^{+}98] G. Y. Huang, E. S. Cooper, K. Waldo, M. L. Kirby, N. B. Gilula, and C. W. Lo. Gap Junction–mediated Cell–Cell Communication Modulates Mouse Neural Crest Migration. Journal of Cell Biology, 143(6):1725–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/6/1725.

Hansen:1999:PNC

[HCW99] William J. Hansen, Nicholas J. Cowan, and William J. Welch. Prefoldin–nascent Chain Complexes in the Folding of Cytoskeletal Proteins. Journal of Cell Biology, 145(2):265–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/265.

Hayes:1996:RMC

[HD96] S. A. Hayes and J. F. Dice. Roles of molecular chaperones in protein degradation. Journal of Cell Biology, 132(3): 255–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/255.

Houchmandzadeh:1999:EMS

[HD99] Bahram Houchmandzadeh and Stefan Dimitrov. Elasticity Measurements Show the Existence of Thin Rigid Cores Inside Mitotic Chromosomes. Journal of Cell Biology, 145(2):215–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/215.

Hodivala-Dilke:1998:NRI

[HDDKH98] Kairbaan M. Hodivala-Dilke, C. Michael DiPersio, Jordan A. Kreidberg, and Richard O. Hynes. Novel Roles for α3β1 Integrin as a Regulator of Cytoskeletal Assembly and as a Trans-dominant Inhibitor of Integrin Receptor Function in Mouse Keratinocytes. Journal of Cell Biology, 142(5):1357– ??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/5/1357.

Huang:1997:DOP

[HDES97] Sui Huang, Thomas J. Deerinck, Mark H. Ellisman, and David L. Spector. The Dynamic Organization of the Perinucleolar Compartment in the Cell Nucleus. Journal of Cell Biology, 137(5):965–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/5/965.

Huang:1998:PCT

[HDES98] Sui Huang, Thomas J. Deerinck, Mark H. Ellisman, and David L. Spector. The Perinucleolar Compartment and Transcription. Journal of Cell Biology, 143(1):35–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/143/1/35.

Hong:1996:PTS

[HDK96] E. Hong, A. R. Davidson, and C. A. Kaiser. A pathway for targeting soluble misfolded proteins to the yeast vacuole. Journal of Cell Biology, 135(3):623–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/3/623.

Heyninck:1999:ZFP

[HDV^{+}99] Karen Heyninck, Dirk De Valck, Wim Vanden Berghe, Wim Van Criekinge, Roland Contreras, Walter Fiers, Guy Haegeman, and Rudi Beyaert. The zinc finger protein A20 inhibits TNF-induced NF-κB-dependent gene expression by interfering with an RIP- or TRAF2-mediated transactivation signal and directly binds to a novel NF-κb-inhibiting protein ABIN. Journal of Cell Biology, 145(7):1471–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/7/1471.

Hannan:1996:TPD

[HE96] L. A. Hannan and M. Edidin. Traffic, polarity, and detergent solubility of a glycosylphosphatidylinositol-anchored protein after LDL-deprivation of MDCK cells. Journal of Cell Biology, 133(6):1265–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/6/1265.

Hamburger:1995:YGR

[HER95] D. Hamburger, M. Egerton, and H. Riezman. Yeast Gaa1p is required for attachment of a completed GPI anchor onto proteins. Journal of Cell Biology, 129(3):629–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/3/

Hudson:1998:CPB

[HFE^{+}98] Damien F. Hudson, Kerry J. Fowler, Elizabeth Earle, Richard Saffery, Paul Kalitsis, Helen Trowell, Joanne Hill, Nigel G. Wreford, David M. de Kretser, Michael R. Cancilla, Emily Howman, Linda Hii, Suzanne M. Cutts, Danielle V. Irvine, and K. H. A. Choo. Centromere Protein B Null Mice are Mitotically and Meiotically Normal but Have Lower Body and Testis Weights. Journal of Cell Biology, 141(2):309–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
2/309.

Hoth:1997:MRS

[HFL97] Markus Hoth, Christopher M. Fanger, and Richard S. Lewis. Mitochondrial Regulation of Store-operated Calcium Signaling in T Lymphocytes. Journal of Cell Biology, 137(3):633–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/3/633.

Hofer:1998:CCE

[HFP98a] Aldebaran M. Hofer, Cristina Fasolato, and Tullio Pozzan. Capacitative Ca^{2+} entry is closely linked to the filling state of internal Ca^{2+} stores: a study using simultaneous measurements of I_{CRAC} and intraluminal [Ca^{2+}]. Journal of Cell Biology, 140 (2):325–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/2/325.

Huang:1998:DCN

[HFP^{+}98b] Da Wei Huang, Laura Fanti, Daniel T. S. Pak, Michael R. Botchan, Sergio Pimpinelli, and Rebecca Kellum. Distinct Cytoplasmic and Nuclear Fractions of Drosophila Heterochromatin Protein 1: Their Phosphorylation Levels and Associations with Origin Recognition Complex Proteins. Journal of Cell Biology, 142(2):307–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/307.

Hirokawa:1996:SST

[HFSHK96] N. Hirokawa, T. Funakoshi, R. Sato-Harada, and Y. Kanai. Selective stabilization of tau in axons and microtubule-associated protein 2C in cell bodies and dendrites contributes to polarized localization of cytoskeletal proteins in mature neurons. Journal of Cell Biology, 132(4):667–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/4/667.

He:1995:RPD

[HG95a] Y. He and F. Grinnell. Role of phospholipase D in the cAMP signal transduction pathway activated during fibroblast contraction of collagen matrices. Journal of Cell Biology, 130(5): 1197–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/5/1197.

Hermiston:1995:VAC

[HG95b] M. L. Hermiston and J. I. Gordon. In vivo analysis of cadherin function in the mouse intestinal epithelium: essential roles in adhesion, maintenance of differentiation, and regulation of programmed cell death. Journal of Cell Biology, 129 (2):489–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/489.

Hu:1996:MFC

[HGG96] T. Hu, T. Guan, and L. Gerace. Molecular and functional characterization of the p62 complex, an assembly of nuclear pore complex glycoproteins. Journal of Cell Biology, 134(3): 589–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/3/589.

Hasson:1997:UMI

[HGG^{+}97] Tama Hasson, Peter G. Gillespie, Jesus A. Garcia, Richard B. MacDonald, Yi dong Zhao, Ann G. Yee, Mark S. Mooseker, and David P. Corey. Unconventional Myosins in Inner–Ear Sensory Epithelia. Journal of Cell Biology, 137(6):1287–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/6/1287.

Huttenlocher:1996:MCM

[HGH96] A. Huttenlocher, M. H. Ginsberg, and A. F. Horwitz. Modulation of cell migration by integrin-mediated cytoskeletal linkages and ligand-binding affinity. Journal of Cell Biology, 134(6):1551–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/6/1551.

Herrmann:1999:RMV

[HGI^{+}99] Christine Herrmann, Elaheh Golkaramnay, Elisabeth Inman, Leonard Rome, and Walter Volknandt. Recombinant Major Vault Protein Is Targeted to Neuritic Tips of PC12 Cells. Journal of Cell Biology, 144(6):1163–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/6/1163.

Huttenlauch:1995:MEP

[HGP^{+}95] I. Huttenlauch, N. Geisler, U. Plessmann, R. K. Peck, K. Weber, and R. Stick. Major epiplasmic proteins of ciliates are articulins: cloning, recombinant expression, and structural characterization. Journal of Cell Biology, 130(6):1401–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1401.

Hotchin:1995:RCS

[HGW95] N. A. Hotchin, A. Gandarillas, and F. M. Watt. Regulation of cell surface beta 1 integrin levels during keratinocyte terminal differentiation. Journal of Cell Biology, 128(6):1209–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
6/1209.

Haskins:1998:ZNM

[HGW^{+}98] Julie Haskins, Lijie Gu, Erika S. Wittchen, Jennifer Hibbard, and Bruce R. Stevenson. ZO-3, a Novel Member of the MAGUK Protein Family Found at the Tight Junction, Interacts with ZO-1 and Occludin. Journal of Cell Biology, 141 (1):199–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/199.

Honing:1995:CDI

[HH95a] S. Höning and W. Hunziker. Cytoplasmic determinants involved in direct lysosomal sorting, endocytosis, and basolateral targeting of rat lgp120 (lamp–I) in MDCK cells. Journal of Cell Biology, 128(3):321–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/3/321.

Hotchin:1995:AIA

[HH95b] N. A. Hotchin and A. Hall. The assembly of integrin adhesion complexes requires both extracellular matrix and intracellular rho/rac GTPases. Journal of Cell Biology, 131(6): 1857–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1857.

Hancock:1998:PMP

[HH98] William O. Hancock and Jonathon Howard. Processivity of the Motor Protein Kinesin Requires Two Heads. Journal of Cell Biology, 140(6):1395–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/6/1395.

Hungerford:1997:INM

[HHB^{+}97] Jill E. Hungerford, James P. Hoeffler, Chauncey W. Bowers, Lisa M. Dahm, Rocco Falchetto, Jeffrey Shabanowitz, Donald F. Hunt, and Charles D. Little. Identification of a Novel Marker for Primordial Smooth Muscle and Its Differential Expression Pattern in Contractile vs Noncontractile Cells. Journal of Cell Biology, 137(4):925–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/4/925.

Huhtala:1995:CSA

[HHM^{+}95] P. Huhtala, M. J. Humphries, J. B. McCarthy, P. M. Tremble, Z. Werb, and C. H. Damsky. Cooperative signaling by alpha 5 beta 1 and alpha 4 beta 1 integrins regulates metalloproteinase gene expression in fibroblasts adhering to fibronectin. Journal of Cell Biology, 129(3):867–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/3/867.

Hortsch:1998:SRO

[HHM^{+}98] Michael Hortsch, Diahann Homer, Jyoti Dhar Malhotra, Sherry Chang, Jason Frankel, Gregory Jefford, and Ronald R. Dubreuil. Structural Requirements for Outside–In and Inside– Out Signaling by Drosophila Neuroglian, a Member of the L1 Family of Cell Adhesion Molecules. Journal of Cell Biology, 142(1):251–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/1/251.

Huot:1998:SPD

[HHR^{+}98] Jacques Huot, François Houle, Simon Rousseau, Réna G. Deschesnes, Girish M. Shah, and Jacques Landry. SAPK2/ p38-dependent F–Actin Reorganization Regulates Early Membrane Blebbing during Stress-induced Apoptosis. Journal of Cell Biology, 143(5):1361–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/5/1361.

Huet:1998:GBI

[HHRdB^{+}98] Guillemette Huet, Sylviane Hennebicq-Reig, Carmen de Bolos, Fausto Ulloa, Thécla Lesuffleur, Alain Barbat, Véronique Carrière, Isabelle Kim, Francisco X. Real, Philippe Delannoy, and Alain Zweibaum. GalNAc-α-O-benzyl inhibits NeuAcα2-3 glycosylation and blocks the intracellular transport of apical glycoproteins and mucus in differentiated HT-29 cells. Journal of Cell Biology, 141(6):1311–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/6/1311.

Hurt:1999:NVA

[HHS^{+}99] Ed Hurt, Stefan Hannus, Birgit Schmelzl, Denise Lau, David Tollervey, and George Simos. A Novel In Vivo Assay Reveals Inhibition of Ribosomal Nuclear Export in Ran–Cycle and Nucleoporin Mutants. Journal of Cell Biology, 144(3): 389–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/389.

Houmeida:1995:SIB

[HHTT95] A. Houmeida, J. Holt, L. Tskhovrebova, and J. Trinick. Studies of the interaction between titin and myosin. Journal of Cell Biology, 131(6):1471–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1471.

Higashiyama:1995:MPC

[HIG^{+}95] S. Higashiyama, R. Iwamoto, K. Goishi, G. Raab, N. Taniguchi, M. Klagsbrun, and E. Mekada. The membrane protein CD9/ DRAP 27 potentiates the juxtacrine growth factor activity of the membrane-anchored heparin-binding EGF-like growth factor. Journal of Cell Biology, 128(5):929–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/128/5/
929.

Hirose:1998:MDR

[HIW^{+}98] Masaya Hirose, Toshimasa Ishizaki, Naoki Watanabe, Masayoshi Uehata, Onno Kranenburg, Wouter H. Moolenaar, Fumio Matsumura, Midori Maekawa, Haruhiko Bito, and Shuh Narumiya. Molecular Dissection of the Rho-associated Protein Kinase (p160ROCK)-regulated Neurite Remodeling in Neuroblastoma N1E-115 Cells. Journal of Cell Biology, 141(7):1625–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
7/1625.

Huby:1997:ZPT

[HIWL97] Russell D. J. Huby, Makio Iwashima, Arthur Weiss, and Steven C. Ley. ZAP-70 Protein Tyrosine Kinase Is Constitutively Targeted to the T Cell Cortex Independently of its SH2 Domains. Journal of Cell Biology, 137(7):1639–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
7/1639.

Huang:1997:DGT

[HJJJ97] Yueqiao Huang, John Jellies, Kristen M. Johansen, and Jørgen Johansen. Differential Glycosylation of Tractin and Leech-CAM, Two Novel Ig Superfamily Members, Regulates Neurite Extension and Fascicle Formation. Journal of Cell Biology, 138(1):143–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/1/143.

Heidemann:1999:DOM

[HKBM99] Steven R. Heidemann, Stefanie Kaech, Robert E. Buxbaum, and Andrew Matus. Direct Observations of the Mechanical Behaviors of the Cytoskeleton in Living Fibroblasts. Journal of Cell Biology, 145(1):109–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/1/109.

Haass:1995:PSB

[HKC^{+}95] C. Haass, E. H. Koo, A. Capell, D. B. Teplow, and D. J. Selkoe. Polarized sorting of beta-amyloid precursor protein and its proteolytic products in MDCK cells is regulated by two independent signals. Journal of Cell Biology, 128(4): 537–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/537.

Harada:1999:LPS

[HKJ^{+}99] Hidemitsu Harada, Päivi Kettunen, Han-Sung Jung, Tuija Mustonen, Y. Alan Wang, and Irma Thesleff. Localization of Putative Stem Cells in Dental Epithelium and Their Association with Notch and Fgf Signaling. Journal of Cell Biology, 147(1):105–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/1/105.

Hyland:1999:CNK

[HKKH99] Katherine M. Hyland, Jeffrey Kingsbury, Doug Koshland, and Philip Hieter. Ctf19p: a Novel Kinetochore Protein in Saccharomyces cerevisiae and a Potential Link between the Kinetochore and Mitotic Spindle. Journal of Cell Biology, 145 (1):15–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/1/15.

Haass:1996:PUE

[HKL96] N. K. Haass, M. A. Kartenbeck, and R. E. Leube. Pantophysin is a ubiquitously expressed synaptophysin homologue and defines constitutive transport vesicles. Journal of Cell Biology, 134(3):731–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/3/731.

Hannon:1996:DEF

[HKM^{+}96] K. Hannon, A. J. Kudla, M. J. McAvoy, K. L. Clase, and B. B. Olwin. Differentially expressed fibroblast growth factors regulate skeletal muscle development through autocrine and paracrine mechanisms. Journal of Cell Biology, 132(6): 1151–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/6/1151.

Hay:1998:LDP

[HKO^{+}98] Jesse C. Hay, Judith Klumperman, Viola Oorschot, Martin Steegmaier, Christin S. Kuo, and Richard H. Scheller. Localization, Dynamics, and Protein Interactions Reveal Distinct Roles for ER and Golgi SNAREs. Journal of Cell Biology, 141 (7):1489–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/7/1489.

Henley:1998:DMI

[HKOM98] John R. Henley, Eugene W. A. Krueger, Barbara J. Oswald, and Mark A. McNiven. Dynamin-mediated Internalization of Caveolae. Journal of Cell Biology, 141(1):85–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/1/85.

Hagios:1996:TPN

[HKS^{+}96] C. Hagios, M. Koch, J. Spring, M. Chiquet, and R. Chiquet-Ehrismann. Tenascin-y: a protein of novel domain structure is secreted by differentiated fibroblasts of muscle connective tissue. Journal of Cell Biology, 134(6):1499–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
6/1499.

Hermann:1997:YGM

[HKS97] Greg J. Hermann, Edward J. King, and Janet M. Shaw. The Yeast Gene, MDM20, Is Necessary for Mitochondrial Inheritance and Organization of the Actin Cytoskeleton. Journal of Cell Biology, 137(1):141–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/1/141.

Kim:1999:SAI

[hKsKC99] Jae hyun Kim, Jung seog Kang, and Clarence S. M. Chan. Sli15 Associates with the Ipl1 Protein Kinase to Promote Proper Chromosome Segregation in Saccharomyces cerevisiae. Journal of Cell Biology, 145(7):1381–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/7/1381.

Hirai:1998:EFK

[HLG^{+}98] Yohei Hirai, André Lochter, Sybille Galosy, Shogo Koshida, Shinichiro Niwa, and Mina J. Bissell. Epimorphin Functions as a Key Morphoregulator for Mammary Epithelial Cells. Journal of Cell Biology, 140(1):159–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/1/159.

Hack:1998:SDL

[HLJ^{+}98] Andrew A. Hack, Chantal T. Ly, Fang Jiang, Cynthia J. Clendenin, Kirsten S. Sigrist, Robert L. Wollmann, and Elizabeth M. McNally. γ-sarcoglycan Deficiency Leads to Muscle Membrane Defects and Apoptosis Independent of Dystrophin. Journal of Cell Biology, 142(5):1279–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
5/1279.

Huttenlocher:1998:ICS

[HLK^{+}98] Anna Huttenlocher, Margot Lakonishok, Melissa Kinder, Stanley Wu, Tho Truong, Karen A. Knudsen, and Alan F. Horwitz. Integrin and Cadherin Synergy Regulates Contact Inhibition of Migration and Motile Activity. Journal of Cell Biology, 141(2):515–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/2/515.

He:1997:MCH

[HLMD97] Huiling He, Minh Lam, Thomas S. McCormick, and Clark W. Distelhorst. Maintenance of Calcium Homeostasis in the Endoplasmic Reticulum by Bcl-2. Journal of Cell Biology, 138(6): 1219–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1219.

Haney:1995:IST

[HLSM95] P. M. Haney, M. A. Levy, M. S. Strube, and M. Mueckler. Insulin-sensitive targeting of the GLUT4 glucose transporter in L6 myoblasts is conferred by its COOH-terminal cytoplasmic tail. Journal of Cell Biology, 129(3):641–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/641.

Hardwick:1995:MPC

[HM95] K. G. Hardwick and A. W. Murray. Mad1p, a phosphoprotein component of the spindle assembly checkpoint in budding yeast. Journal of Cell Biology, 131(3):709–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/709.

Henley:1996:ADL

[HM96] J. R. Henley and M. A. McNiven. Association of a dynaminlike protein with the Golgi apparatus in mammalian cells. Journal of Cell Biology, 133(4):761–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/4/761.

Hanneken:1995:HAI

[HMB95a] A. Hanneken, P. A. Maher, and A. Baird. High affinity immunoreactive FGF receptors in the extracellular matrix of vascular endothelial cells–implications for the modulation of FGF-2. Journal of Cell Biology, 128(6):1221–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
6/1221.

Hu:1995:EFD

[HMB95b] R. J. Hu, S. Moorthy, and V. Bennett. Expression of functional domains of beta G-spectrin disrupts epithelial morphology in cultured cells. Journal of Cell Biology, 128(6):1069–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
6/1069.

Hartwell:1998:RPS

[HMB^{+}98] Daqing W. Hartwell, Tanya N. Mayadas, Gaëtan Berger, Paul S. Frenette, Helen Rayburn, Richard O. Hynes, and Denisa D. Wagner. Role of P–Selectin Cytoplasmic Domain in Granular Targeting In Vivo and in Early Inflammatory Responses. Journal of Cell Biology, 143(4):1129–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/1129.

Hobert:1999:CLP

[HMC^{+}99] Oliver Hobert, Donald G. Moerman, Kathleen A. Clark, Mary C. Beckerle, and Gary Ruvkun. A Conserved LIM Protein That Affects Muscular Adherens Junction Integrity and Mechanosensory Function in Caenorhabditis elegans. Journal of Cell Biology, 144(1):45–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/1/45.

Houchmandzadeh:1997:ESE

[HMCL97] Bahram Houchmandzadeh, John F. Marko, Didier Chatenay, and Albert Libchaber. Elasticity and Structure of Eukaryote Chromosomes Studied by Micromanipulation and Micropipette Aspiration. Journal of Cell Biology, 139(1): 1–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/1.

Hanein:1997:ECC

[HMD97] Dorit Hanein, Paul Matsudaira, and David J. DeRosier. Evidence for a Conformational Change in Actin Induced by Fimbrin (N375) Binding. Journal of Cell Biology, 139(2): 387–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/387.

Hirschberg:1998:KAS

[HME^{+}98] Koret Hirschberg, Chad M. Miller, Jan Ellenberg, John F. Presley, Eric D. Siggia, Robert D. Phair, and Jennifer Lippincott-Schwartz. Kinetic Analysis of Secretory Protein Traffic and Characterization of Golgi to Plasma Membrane Transport Intermediates in Living Cells. Journal of Cell Biology, 143(6):1485–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1485.

Higgins:1998:DRI

[HMS^{+}98] Jonathan M. G. Higgins, Didier A. Mandlebrot, Sunil K. Shaw, Gary J. Russell, Elizabeth A. Murphy, Yih-Tai Chen, W. James Nelson, Christina M. Parker, and Michael B. Brenner. Direct and regulated interaction of integrin α_{E}β_{7} with E-cadherin. Journal of Cell Biology, 140(1):197–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
1/197.

Harding:1995:ICY

[HMSK95] T. M. Harding, K. A. Morano, S. V. Scott, and D. J. Klionsky. Isolation and characterization of yeast mutants in the cytoplasm to vacuole protein targeting pathway. Journal of Cell Biology, 131(3):591–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/3/591.

Hua:1998:IPS

[HN98] Xuequn Helen Hua and John Newport. Identification of a Preinitiation Step in DNA Replication That Is Independent of Origin Recognition Complex and cdc6, but Dependent on cdk2. Journal of Cell Biology, 140(2):271–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
2/271.

Harrison:1998:ICE

[HOCK98] Alistair Harrison, Patricia Olds-Clarke, and Stephen M. King. Identification of the t Complex–encoded Cytoplasmic Dynein Light Chain Tctex1 in Inner Arm I1 Supports the Involvement of Flagellar Dyneins in Meiotic Drive. Journal of Cell Biology, 140(5):1137–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/5/1137.

Heijnen:1997:TSG

[HOS^{+}97] Harry F. G. Heijnen, Viola Oorschot, Jan J. Sixma, Jan W. Slot, and David E. James. Thrombin Stimulates Glucose Transport in Human Platelets via the Translocation of the Glucose Transporter GLUT-3 from α-Granules to the Cell Surface. Journal of Cell Biology, 138(2):323–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/138/2/

Hanada:1995:CHO

[HP95] K. Hanada and R. E. Pagano. A Chinese hamster ovary cell mutant defective in the non-endocytic uptake of fluorescent analogs of phosphatidylserine: isolation using a cytosol acidification protocol. Journal of Cell Biology, 128(5):793–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/793.

Hernandez:1997:ARM

[HPD^{+}97] Lorraine D. Hernandez, Reuben J. Peters, Sue E. Delos, John A. T. Young, David A. Agard, and Judith M. White. Activation of a Retroviral Membrane Fusion Protein: Soluble Receptor-induced Liposome Binding of the ALSV Envelope Glycoprotein. Journal of Cell Biology, 139(6):1455–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/6/1455.

Han:1997:MMP

[HPKL97] Ji Woong Han, Jong Ho Park, Misook Kim, and JooHun Lee. mRNAs for Microtubule Proteins Are Specifically Colocalized during the Sequential Formation of Basal Body, Flagella, and Cytoskeletal Microtubules in the Differentiation of Naegleria gruberi. Journal of Cell Biology, 137(4):871–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
4/871.

Hoock:1997:IAL

[HPL97] Thomas C. Hoock, Luanne L. Peters, and Samuel E. Lux. Isoforms of ankyrin-3 that lack the NH_{2}-terminal repeats associate with mouse macrophage lysosomes. Journal of Cell Biology, 136(5):1059–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/5/1059.

Hall:1999:HRS

[HPL99] Randy A. Hall, Richard T. Premont, and Robert J. Lefkowitz. Heptahelical Receptor Signaling: Beyond the G Protein Paradigm. Journal of Cell Biology, 145(5):927–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
5/927.

Hato:1998:CRR

[HPS98] Takaaki Hato, Nisar Pampori, and Sanford J. Shattil. Complementary Roles for Receptor Clustering and Conformational Change in the Adhesive and Signaling Functions of Integrin α_{IIb} β_{3}. Journal of Cell Biology, 141(7):1685–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
7/1685.

Hutton:1998:FDB

[HPY^{+}98] Elizabeth Hutton, Rudolph D. Paladini, Qian-Chun Yu, Mei Yen, Pierre A. Coulombe, and Elaine Fuchs. Functional Differences between Keratins of Stratified and Simple Epithelia. Journal of Cell Biology, 143(2):487–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/2/487.

Hof:1999:DFA

[HR99] Woutervan’t Hof and Marilyn D. Resh. Dual Fatty Acylation of p59 ^{Fyn} Is Required for Association with the T Cell Receptor ζ Chain through Phosphotyrosine–Src Homology Domain-2 Interactions. Journal of Cell Biology, 145(2):377–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
2/377.

Huhse:1998:PSC

[HRA^{+}98] Bettina Huhse, Peter Rehling, Markus Albertini, Lars Blank, Karl Meller, and Wolf-H. Kunau. Pex17p of Saccharomyces cerevisiae Is a Novel Peroxin and Component of the Peroxisomal Protein Translocation Machinery. Journal of Cell Biology, 140(1):49–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/1/49.

Hirschi:1998:PTH

[HRD98] Karen K. Hirschi, Stephanie A. Rohovsky, and Patricia A. D’Amore. PDGF, TGF-β, and heterotypic cell–cell interactions mediate endothelial cell-induced recruitment of 10T1/2 cells and their differentiation to a smooth muscle fate. Journal of Cell Biology, 141(3):805–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/3/805.

Hettema:1998:CDL

[HRK^{+}98] Ewald H. Hettema, Caroline C. M. Ruigrok, Marian Groot Koerkamp, Marlene van den Berg, Henk F. Tabak, Ben Distel, and Ineke Braakman. The Cytosolic DnaJ-like Protein Djp1p Is Involved Specifically in Peroxisomal Protein Import. Journal of Cell Biology, 142(2):421–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/421.

Haaf:1999:SMR

[HRR^{+}99] Thomas Haaf, Elke Raderschall, Gurucharan Reddy, David C. Ward, Charles M. Radding, and Efim I. Golub. Sequestration of Mammalian Rad51-Recombination Protein into Micronuclei. Journal of Cell Biology, 144(1):11–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/144/1/11.

Huang:1996:IDR

[HS96] S. Huang and D. L. Spector. Intron-dependent recruitment of pre-mRNA splicing factors to sites of transcription. Journal of Cell Biology, 133(4):719–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/4/719.

Habermacher:1997:RFD

[HS97] Geoffrey Habermacher and Winfield S. Sale. Regulation of Flagellar Dynein by Phosphorylation of a 138-kD Inner Arm Dynein Intermediate Chain. Journal of Cell Biology, 136(1): 167–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/167.

Hock:1998:CPH

[HSB98] Robert Hock, Ulrich Scheer, and Michael Bustin. Chromosomal Proteins HMG-14 and HMG-17 Are Released from Mitotic Chromosomes and Imported into the Nucleus by Active Transport. Journal of Cell Biology, 143(6):1427–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1427.

Hirao:1996:RME

[HSK^{+}96] M. Hirao, N. Sato, T. Kondo, S. Yonemura, M. Monden, T. Sasaki, Y. Takai, S. Tsukita, and S. Tsukita. Regulation mechanism of ERM (ezrin/radixin/moesin) protein/ plasma membrane association: possible involvement of phosphatidylinositol turnover and Rho-dependent signaling pathway. Journal of Cell Biology, 135(1):37–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/135/1/37.

Hemar:1995:EIR

[HSL^{+}95] A. Hémar, A. Subtil, M. Lieb, E. Morelon, R. Hellio, and A. Dautry-Varsat. Endocytosis of interleukin 2 receptors in human T lymphocytes: distinct intracellular localization and fate of the receptor alpha, beta, and gamma chains. Journal of Cell Biology, 129(1):55–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/1/55.

Haney:1999:HBL

[HSL^{+}99] C. A. Haney, Z. Sahenk, C. Li, V. P. Lemmon, J. Roder, and B. D. Trapp. Heterophilic Binding of L1 on Unmyelinated Sensory Axons Mediates Schwann Cell Adhesion and Is Required for Axonal Survival. Journal of Cell Biology, 146(5): 1173–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/1173.

Hande:1999:TLD

[HSLB99] M. Prakash Hande, Enrique Samper, Peter Lansdorp, and María A. Blasco. Telomere Length Dynamics and Chromosomal Instability in Cells Derived from Telomerase Null Mice. Journal of Cell Biology, 144(4):589–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
4/589.

Hocking:1996:NRI

[HSML96] D. C. Hocking, R. K. Smith, and P. J. McKeown-Longo. A novel role for the integrin-binding III-10 module in fibronectin matrix assembly. Journal of Cell Biology, 133(2):431–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
2/431.

Hocking:1998:ADM

[HSML98] Denise C. Hocking, Jane Sottile, and Paula J. McKeown-Longo. Activation of distinct α_{5}β_{1}-mediated signaling pathways by fibronectin’s cell adhesion and matrix assembly domains. Journal of Cell Biology, 141(1):241–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/1/
241.

Hresko:1999:MNH

[HSSW99] Michelle Coutu Hresko, Lawrence A. Schriefer, Paresh Shrimankar, and Robert H. Waterston. Myotactin, a Novel Hypodermal Protein Involved in Muscle–Cell Adhesion in Caenorhabditis elegans. Journal of Cell Biology, 146(3): 659–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/3/659.

Hoenger:1998:IRM

[HST^{+}98] A. Hoenger, S. Sack, M. Thormählen, A. Marx, J. Müller, H. Gross, and E. Mandelkow. Image Reconstructions of Microtubules Decorated with Monomeric and Dimeric Kinesins: Comparison with X–Ray Structure and Implications for Motility. Journal of Cell Biology, 141(2):419–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
2/419.

Harder:1998:LDS

[HSVS98] Thomas Harder, Peter Scheiffele, Paul Verkade, and Kai Simons. Lipid Domain Structure of the Plasma Membrane Revealed by Patching of Membrane Components. Journal of Cell Biology, 141(4):929–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/4/929.

Hackam:1998:IHP

[HSW^{+}98] Abigail S. Hackam, Roshni Singaraja, Cheryl L. Wellington, Martina Metzler, Krista McCutcheon, Taiqi Zhang, Michael Kalchman, and Michael R. Hayden. The Influence of Huntingtin Protein Size on Nuclear Localization and Cellular Toxicity. Journal of Cell Biology, 141(5):1097–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/5/
1097.

Hausler:1996:IDH

[HSWC96] T. Häusler, Y. D. Stierhof, E. Wirtz, and C. Clayton. Import of a DHFR hybrid protein into glycosomes in vivo is not inhibited by the folate-analogue aminopterin. Journal of Cell Biology, 132(3):311–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/3/311.

Hirokawa:1998:GTS

[HT98] Nobutaka Hirokawa and Sen Takeda. Gene Targeting Studies Begin to Reveal the Function of Neurofilament Proteins. Journal of Cell Biology, 143(1):1–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/1/1.

Heald:1997:SAX

[HTH^{+}97] Rebecca Heald, Régis Tournebize, Anja Habermann, Eric Karsenti, and Anthony Hyman. Spindle Assembly in Xenopus Egg Extracts: Respective Roles of Centrosomes and Microtubule Self–Organization. Journal of Cell Biology, 138(3): 615–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/3/615.

Harada:1998:GVL

[HTK^{+}98] A. Harada, Y. Takei, Y. Kanai, Y. Tanaka, S. Nonaka, and N. Hirokawa. Golgi Vesiculation and Lysosome Dispersion in Cells Lacking Cytoplasmic Dynein. Journal of Cell Biology, 141(1):51–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/51.

Hayashi:1999:CBP

[HTK^{+}99] Ken’ichiro Hayashi, Masanori Takahashi, Kazuhiro Kimura, Wataru Nishida, Hiroshi Saga, and Kenji Sobue. Changes in the balance of phosphoinositide 3-kinase/protein kinase B (Akt) and the mitogen-activated protein kinases (ERK/ p38MAPK) determine a phenotype of visceral and vascular smooth muscle cells. Journal of Cell Biology, 145(4):727–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/4/727.

Holleran:1996:CAA

[HTKH96] E. A. Holleran, M. K. Tokito, S. Karki, and E. L. Holzbaur. Centractin (ARP1) associates with spectrin revealing a potential mechanism to link dynactin to intracellular organelles. Journal of Cell Biology, 135(6):1815–??, December 1996. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/135/6/
1815.

Hermann:1998:MFY

[HTM^{+}98] Greg J. Hermann, John W. Thatcher, John P. Mills, Karen G. Hales, Margaret T. Fuller, Jodi Nunnari, and Janet M. Shaw. Mitochondrial Fusion in Yeast Requires the Transmembrane GTPase Fzo1p. Journal of Cell Biology, 143(2):359–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/2/359.

Haas:1995:KDB

[HTS^{+}95] S. Haas, P. Thatikunta, A. Steplewski, E. M. Johnson, K. Khalili, and S. Amini. A 39-kD DNA-binding protein from mouse brain stimulates transcription of myelin basic protein gene in oligodendrocytic cells. Journal of Cell Biology, 130(5): 1171–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/5/1171.

Holst:1997:ASM

[HTW97] Bjørn Holst, Christine Tachibana, and Jakob R. Winther. Active Site Mutations in Yeast Protein Disulfide Isomerase Cause Dithiothreitol Sensitivity and a Reduced Rate of Protein Folding in the Endoplasmic Reticulum. Journal of Cell Biology, 138(6):1229–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/6/1229.

Herold:1998:DFD

[HTWC98] Andrea Herold, Ray Truant, Heather Wiegand, and Bryan R. Cullen. Determination of the Functional Domain Organization of the Importin α Nuclear Import Factor. Journal of Cell Biology, 143(2):309–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/309.

Hieda:1999:MAC

[HTY^{+}99] Miki Hieda, Taro Tachibana, Fumihiko Yokoya, Shingo Kose, Naoko Imamoto, and Yoshihiro Yoneda. A Monoclonal Antibody to the COOH-terminal Acidic Portion of Ran Inhibits Both the Recycling of Ran and Nuclear Protein Import in Living Cells. Journal of Cell Biology, 144(4):645–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
4/645.

Holstein:1996:MCB

[HUU96] S. E. Holstein, H. Ungewickell, and E. Ungewickell. Mechanism of clathrin basket dissociation: separate functions of protein domains of the DnaJ homologue auxilin. Journal of Cell Biology, 135(4):925–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/4/925.

Hellman:1999:VAB

[HVL^{+}99] Rachel Hellman, Marc Vanhove, Annabelle Lejeune, Fred J. Stevens, and Linda M. Hendershot. The In Vivo Association of BiP with Newly Synthesized Proteins Is Dependent on the Rate and Stability of Folding and Not Simply on the Presence of Sequences That Can Bind to BiP. Journal of Cell Biology, 144(1):21–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/1/21.

Harris:1996:LYE

[HW96a] S. L. Harris and M. G. Waters. Localization of a yeast early Golgi mannosyltransferase, Och1p, involves retrograde transport. Journal of Cell Biology, 132(6):985–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
6/985.

Hoppe:1996:HVA

[HW96b] P. E. Hoppe and R. H. Waterston. Hydrophobicity variations along the surface of the coiled-coil rod may mediate striated muscle myosin assembly in Caenorhabditis elegans. Journal of Cell Biology, 135(2):371–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/2/371.

Happe:1998:CFT

[HW98] Scott Happe and Peggy Weidman. Cell-free Transport to Distinct Golgi Cisternae Is Compartment Specific and ARF Independent. Journal of Cell Biology, 140(3):511–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
3/511.

Huang:1996:IIB

[HWC^{+}96] X. Z. Huang, J. F. Wu, D. Cass, D. J. Erle, D. Corry, S. G. Young, R. V. Farese, and D. Sheppard. Inactivation of the integrin beta 6 subunit gene reveals a role of epithelial integrins in regulating inflammation in the lung and skin. Journal of Cell Biology, 133(4):921–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/4/921.

Huang:1996:UAG

[HWFA96] R. P. Huang, J. X. Wu, Y. Fan, and E. D. Adamson. UV activates growth factor receptors via reactive oxygen intermediates. Journal of Cell Biology, 133(1):211–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
1/211.

Hecht:1996:VZG

[HWPC96] J. H. Hecht, J. A. Weiner, S. R. Post, and J. Chun. Ventricular zone gene-1 (vzg-1) encodes a lysophosphatidic acid receptor expressed in neurogenic regions of the developing cerebral cortex. Journal of Cell Biology, 135(4):1071–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/1071.

Hagan:1995:PSF

[HY95] I. Hagan and M. Yanagida. The product of the spindle formation gene sad1+ associates with the fission yeast spindle pole body and is essential for viability. Journal of Cell Biology, 129 (4):1033–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/4/1033.

Honda:1998:ANA

[HYE^{+}98] Kazufumi Honda, Tesshi Yamada, Ritsuko Endo, Yoshinori Ino, Masahiro Gotoh, Hitoshi Tsuda, Yozo Yamada, Hiroshige Chiba, and Setsuo Hirohashi. Actinin-4, a Novel Actinbundling Protein Associated with Cell Motility and Cancer Invasion. Journal of Cell Biology, 140(6):1383–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1383.

Yang:1997:ULB

[hYGYF97] Zhao hui Yang, G. Ian Gallicano, Qian-Chun Yu, and Elaine Fuchs. An Unexpected Localization of Basonuclin in the Centrosome, Mitochondria, and Acrosome of Developing Spermatids. Journal of Cell Biology, 137(3):657–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/137/3/
657.

Hsueh:1998:DIC

[HYK^{+}98] Yi-Ping Hsueh, Fu-Chia Yang, Viktor Kharazia, Scott Naisbitt, Alexandra R. Cohen, Richard J. Weinberg, and Morgan Sheng. Direct interaction of CASK/LIN-2 and syndecan heparan sulfate proteoglycan and their overlapping distribution in neuronal synapses. Journal of Cell Biology, 142(1):139–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
142/1/139.

Hua:1997:RCK

[HYN97] Xuequn Helen Hua, Hong Yan, and John Newport. A Role for Cdk2 Kinase in Negatively Regulating DNA Replication during S Phase of the Cell Cycle. Journal of Cell Biology, 137 (1):183–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/1/183.

Hebert:1997:NLG

[HZC^{+}97] Daniel N. Hebert, Jian-Xin Zhang, Wei Chen, Brigitte Foellmer, and Ari Helenius. The Number and Location of Glycans on Influenza Hemagglutinin Determine Folding and Association with Calnexin and Calreticulin. Journal of Cell Biology, 139(3):613–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/3/613.

Hicke:1998:CTP

[HZR98] Linda Hicke, Bettina Zanolari, and Howard Riezman. Cytoplasmic Tail Phosphorylation of the α-Factor Receptor Is Required for Its Ubiquitination and Internalization. Journal of Cell Biology, 141(2):349–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/2/349.

Ilic:1998:EMS

[IAS^{+}98] Duško Ilić, Eduardo A. C. Almeida, David D. Schlaepfer, Paul Dazin, Shinichi Aizawa, and Caroline H. Damsky. Extracellular Matrix Survival Signals Transduced by Focal Adhesion Kinase Suppress p53-mediated Apoptosis. Journal of Cell Biology, 143(2):547–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/547.

Islam:1996:ECH

[ICW^{+}96] S. Islam, T. E. Carey, G. T. Wolf, M. J. Wheelock, and K. R. Johnson. Expression of N-cadherin by human squamous carcinoma cells induces a scattered fibroblastic phenotype with disrupted cell–cell adhesion. Journal of Cell Biology, 135(6): 1643–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1643.

Irminger-Finger:1996:MEG

[IFHR^{+}96] I. Irminger-Finger, E. Hurt, A. Roebuck, M. A. Collart, and S. J. Edelstein. MHP1, an essential gene in Saccharomyces cerevisiae required for microtubule function. Journal of Cell Biology, 135(5):1323–??, December 1996. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/5/1323.

Itoh:1999:DBT

[IFM^{+}99] Masahiko Itoh, Mikio Furuse, Kazumasa Morita, Koji Kubota, Mitinori Saitou, and Shoichiro Tsukita. Direct Binding of Three Tight Junction–Associated Maguks, Zo-1, Zo-2, and Zo-3, with the Cooh Termini of Claudins. Journal of Cell Biology, 147(6):1351–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/6/1351.

Irminger-Finger:1998:VRB

[IFSV^{+}98] Irmgard Irminger-Finger, Jesus V. Soriano, Geneviève Vaudan, Roberto Montesano, and André-Pascal Sappino. In Vitro Repression of Brca1-associated RING Domain Gene, Bard1, Induces Phenotypic Changes in Mammary Epithelial Cells. Journal of Cell Biology, 143(5):1329–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1329.

Izumi:1998:APD

[IHT^{+}98] Yasushi Izumi, Tomonori Hirose, Yoko Tamai, Syu ichi Hirai, Yoji Nagashima, Toyoshi Fujimoto, Yo Tabuse, Kenneth J. Kemphues, and Shigeo Ohno. An Atypical PKC Directly Associates and Colocalizes at the Epithelial Tight Junction with ASIP, a Mammalian Homologue of Caenorhabditis elegans Polarity Protein PAR-3. Journal of Cell Biology, 143 (1):95–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/1/95.

Imamura:1999:FDC

[IIM^{+}99] Yuzo Imamura, Masahiko Itoh, Yoshito Maeno, Shoichiro Tsukita, and Akira Nagafuchi. Functional Domains of α-Catenin Required for the Strong State of Cadherin-based Cell Adhesion. Journal of Cell Biology, 144(6):1311–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
6/1311.

Izaurralde:1997:RMT

[IJB^{+}97] Elisa Izaurralde, Artur Jarmolowski, Christina Beisel, Iain W. Mattaj, Gideon Dreyfuss, and Utz Fischer. A Role for the M9 Transport Signal of hnRNP A1 in mRNA Nuclear Export. Journal of Cell Biology, 137(1):27–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/1/27.

Ishizaki:1998:RCL

[IJR98] Yasuki Ishizaki, Michael D. Jacobson, and Martin C. Raff. A Role for Caspases in Lens Fiber Differentiation. Journal of Cell Biology, 140(1):153–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/1/153.

Imai:1998:ORB

[IKR^{+}98] Shinji Imai, Marko Kaksonen, Erkki Raulo, Tarja Kinnunen, Carole Fages, Xiaojuan Meng, Merja Lakso, and Heikki Rauvala. Osteoblast Recruitment and Bone Formation Enhanced by Cell Matrix–associated Heparin-binding Growth-associated Molecule (HB–GAM). Journal of Cell Biology, 143(4):1113– ??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/4/1113.

Igarashi:1996:GCC

[IKT^{+}96] M. Igarashi, S. Kozaki, S. Terakawa, S. Kawano, C. Ide, and Y. Komiya. Growth cone collapse and inhibition of neurite growth by Botulinum neurotoxin C1: a t–SNARE is involved in axonal growth. Journal of Cell Biology, 134(1):205–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
1/205.

Italiano:1999:BPA

[ILSH99] Joseph E. Italiano, Patrick Lecine, Ramesh A. Shivdasani, and John H. Hartwig. Blood Platelets Are Assembled Principally at the Ends of Proplatelet Processes Produced by Differentiated Megakaryocytes. Journal of Cell Biology, 147(6): 1299–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/6/1299.

Ishov:1996:PND

[IM96] A. M. Ishov and G. G. Maul. The periphery of nuclear domain 10 (ND10) as site of DNA virus deposition. Journal of Cell Biology, 134(4):815–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/4/815.

Iwasaki:1996:ECE

[IMFS96] K. Iwasaki, J. McCarter, R. Francis, and T. Schedl. emo-1, a Caenorhabditis elegans Sec61p gamma homologue, is required for oocyte development and ovulation. Journal of Cell Biology, 134(3):699–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/3/699.

Ito:1999:KCM

[IMK99] Koichi Ito, Yasushi Miyashita, and Haruo Kasai. Kinetic Control of Multiple Forms of Ca^{2+} Spikes by Inositol Trisphosphate in Pancreatic Acinar Cells. Journal of Cell Biology, 146 (2):405–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/2/405.

Manabe:1997:MCA

[iMOeM^{+}97] Ri ichiroh Manabe, Naoko Oh-e, Toshinaga Maeda, Tomohiko Fukuda, and Kiyotoshi Sekiguchi. Modulation of Celladhesive Activity of Fibronectin by the Alternatively Spliced EDA Segment. Journal of Cell Biology, 139(1):295–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/1/295.

Ihrke:1998:APM

[IMS^{+}98] Gudrun Ihrke, Greg V. Martin, Michael R. Shanks, Michael Schrader, Trina A. Schroer, and Ann L. Hubbard. Apical Plasma Membrane Proteins and Endolyn-78 Travel through a Subapical Compartment in Polarized WIF–B Hepatocytes. Journal of Cell Biology, 141(1):115–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/1/115.

Imafuku:1999:PSF

[IMW^{+}99] I. Imafuku, T. Masaki, M. Waragai, S. Takeuchi, M. Kawabata, S. i. Hirai, S. Ohno, L. E. Nee, C. F. Lippa, I. Kanazawa, M. Imagawa, and H. Okazawa. Presenilin 1 suppresses the function of C-Jun homodimers via interaction with Qm/Jif-1. Journal of Cell Biology, 147(1):121–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/1/121.

Ikeda:1999:A

[INM^{+}99] Wataru Ikeda, Hiroyuki Nakanishi, Jun Miyoshi, Kenji Mandai, Hiroyoshi Ishizaki, Miki Tanaka, Atushi Togawa, Kenichi Takahashi, Hideo Nishioka, Hisahiro Yoshida, Akira Mizoguchi, Shin ichi Nishikawa, and Yoshimi Takai. Afadin. Journal of Cell Biology, 146(5):1117–??, September 1999. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/146/5/
1117.

Itoh:1997:IZC

[INMT97] Masahiko Itoh, Akira Nagafuchi, Seiji Moroi, and Shoichiro Tsukita. Involvement of ZO-1 in Cadherin-based Cell Adhesion through Its Direct Binding to α Catenin and Actin Filaments. Journal of Cell Biology, 138(1):181–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
1/181.

Infante:1999:GCG

[IRMF^{+}99] Carlos Infante, Francisco Ramos-Morales, Concepción Fedriani, Michel Bornens, and Rosa M. Rios. GMAP-210, a cis-Golgi network-associated protein, is a minus end microtubulebinding protein. Journal of Cell Biology, 145(1):83–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
1/83.

Ivaska:1999:IMI

[IRW^{+}99] Johanna Ivaska, Hilkka Reunanen, Jukka Westermarck, Leeni Koivisto, Veli-Matti Kähäri, and Jyrki Heino. Integrin α2β1 Mediates Isoform–Specific Activation of p38 and Upregulation of Collagen Gene Transcription by a Mechanism Involving the α2 Cytoplasmic Tail. Journal of Cell Biology, 147(2):

401–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/2/401.

Imaizumi-Scherrer:1996:AFM

[ISFB^{+}96] T. Imaizumi-Scherrer, D. M. Faust, J. C. Bénichou, R. Hellio, and M. C. Weiss. Accumulation in fetal muscle and localization to the neuromuscular junction of cAMP-dependent protein kinase A regulatory and catalytic subunits RI alpha and C alpha. Journal of Cell Biology, 134(5):1241–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
5/1241.

Itin:1995:TPE

[ISH95] C. Itin, R. Schindler, and H. P. Hauri. Targeting of protein ERGIC-53 to the ER/ERGIC/cis-Golgi recycling pathway. Journal of Cell Biology, 131(1):57–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/131/1/57.

Ishov:1997:HCI

[ISM97] Alexander M. Ishov, Richard M. Stenberg, and Gerd G. Maul. Human Cytomegalovirus Immediate Early Interaction with Host Nuclear Structures: Definition of an Immediate Transcript Environment. Journal of Cell Biology, 138(1):5–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
1/5.

Ishov:1999:PCN

[ISN^{+}99] Alexander M. Ishov, Alexey G. Sotnikov, Dmitri Negorev, Olga V. Vladimirova, Norma Neff, Tetsu Kamitani, Edward T. H. Yeh, Jerome F. Strauss, and Gerd G. Maul. Pml Is Critical for Nd10 Formation and Recruits the Pml–Interacting Protein Daxx to This Nuclear Structure When Modified by Sumo-1. Journal of Cell Biology, 147(2):221–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
2/221.

Italiano:1999:LDM

[ISR99] Joseph E. Italiano, Murray Stewart, and Thomas M. Roberts. Localized depolymerization of the major sperm protein cytoskeleton correlates with the forward movement of the cell body in the amoeboid movement of nematode sperm. Journal of Cell Biology, 146(5):1087–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/5/1087.

Iovine:1997:NES

[IW97] M. Kathryn Iovine and Susan R. Wente. A Nuclear Export Signal in Kap95p Is Required for Both Recycling the Import Factor and Interaction with the Nucleoporin GLFG Repeat Regions of Nup116p and Nup100p. Journal of Cell Biology, 137(4):797–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/4/797.

Iovine:1995:GRR

[IWW95] M. K. Iovine, J. L. Watkins, and S. R. Wente. The GLFG repetitive region of the nucleoporin Nup116p interacts with Kap95p, an essential yeast nuclear import factor. Journal of Cell Biology, 131(6):1699–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1699.

Isaac:1998:NFM

[IYM98] Cynthia Isaac, Yunfeng Yang, and U. Thomas Meier. Nopp140 Functions as a Molecular Link Between the Nucleolus and the Coiled Bodies. Journal of Cell Biology, 142(2):319–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
2/319.

Johnson:1999:CCR

[JB99] Erica S. Johnson and Günter Blobel. Cell Cycle–Regulated Attachment of the Ubiquitin–Related Protein Sumo to the Yeast Septins. Journal of Cell Biology, 147(5):981–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
5/981.

Jackson:1995:PFL

[JBD^{+}95] D. G. Jackson, J. I. Bell, R. Dickinson, J. Timans, J. Shields, and N. Whittle. Proteoglycan forms of the lymphocyte homing receptor CD44 are alternatively spliced variants containing the v3 exon. Journal of Cell Biology, 128(4):673–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/673.

Jakob:1998:DME

[JBRA98] Claude A. Jakob, Patricie Burda, Jürgen Roth, and Markus Aebi. Degradation of Misfolded Endoplasmic Reticulum Glycoproteins in Saccharomyces cerevisiae Is Determined by a Specific Oligosaccharide Structure. Journal of Cell Biology, 142(5):1223–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/5/1223.

Jimenez:1998:MBC

[JCC^{+}98] Javier Jiménez, Víctor J. Cid, Rosa Cenamor, María Yuste, Gloria Molero, César Nombela, and Miguel Sánchez. Morphogenesis beyond Cytokinetic Arrest in Saccharomyces cerevisiae. Journal of Cell Biology, 143(6):1617–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1617.

Jackson:1995:EED

[JCPK95] P. K. Jackson, S. Chevalier, M. Philippe, and M. W. Kirschner. Early events in DNA replication require cyclin E and are blocked by p21CIP1. Journal of Cell Biology, 130(4):755– ??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/755.

Jones:1997:RTC

[JCR97] Peter Lloyd Jones, Julie Crack, and Marlene Rabinovitch. Regulation of Tenascin-C, a vascular smooth muscle cell survival factor that interacts with the α_{v}β_{3} integrin to promote epidermal growth factor receptor phosphorylation and growth. Journal of Cell Biology, 139(1):279–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/1/279.

Jacobson:1995:DLR

[JCWP95] M. R. Jacobson, L. G. Cao, Y. L. Wang, and T. Pederson. Dynamic localization of RNase MRP RNA in the nucleolus observed by fluorescent RNA cytochemistry in living cells. Journal of Cell Biology, 131(6):1649–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1649.

Jones:1997:EEP

[JDK97] D. Holstead Jones, Tyler C. Davies, and Gerald M. Kidder. Embryonic Expression of the Putative γ Subunit of the Sodium Pump Is Required for Acquisition of Fluid Transport Capacity during Mouse Blastocyst Development. Journal of Cell Biology, 139(6):1545–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/6/1545.

Johnson:1998:COC

[JE98] Catharine E. Johnson and Paul T. Englund. Changes in Organization of Crithidia fasciculata Kinetoplast DNA Replication Proteins during the Cell Cycle. Journal of Cell Biology, 143(4): 911–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/4/911.

Jorgensen:1997:AGJ

[JGCS97] Niklas R. Jørgensen, Steven T. Geist, Roberto Civitelli, and Thomas H. Steinberg. ATP- and gap junction-dependent intercellular calcium signaling in osteoblastic cells. Journal of Cell Biology, 139(2):497–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/2/497.

Johnson:1996:TRC

[JGD^{+}96] A. O. Johnson, R. N. Ghosh, K. W. Dunn, R. Garippa, J. Park, S. Mayor, F. R. Maxfield, and T. E. McGraw. Transferrin receptor containing the SDYQRL motif of TGN38 causes a reorganization of the recycling compartment but is not targeted to the TGN. Journal of Cell Biology, 135(6):1749–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
6/1749.

Jin:1996:RIC

[JGM96] P. Jin, Y. Gu, and D. O. Morgan. Role of inhibitory CDC2 phosphorylation in radiation-induced G2 arrest in human cells.

Journal of Cell Biology, 134(4):963–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/4/963.

Jotte:1996:MFV

[JH96] R. M. Jotte and J. T. Holt. Myristylation of FBR v-fos dictates the differentiation pathways in malignant osteosarcoma. Journal of Cell Biology, 135(2):457–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/2/457.

Jones:1997:PKR

[JH97] Steven M. Jones and Kathryn E. Howell. Phosphatidylinositol 3–Kinase Is Required for the Formation of Constitutive Transport Vesicles from the TGN. Journal of Cell Biology, 139(2): 339–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/339.

Jin:1998:NLC

[JHM98] Pei Jin, Stephen Hardy, and David O. Morgan. Nuclear Localization of Cyclin B1 Controls Mitotic Entry After DNA Damage. Journal of Cell Biology, 141(4):875–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/4/
875.

Jordan:1997:MLC

[JK97] Pascale Jordan and Roger Karess. Myosin Light Chain– activating Phosphorylation Sites Are Required for Oogenesis in Drosophila. Journal of Cell Biology, 139(7):1805–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/7/1805.

Job:1998:CPK

[JL98] Christy Job and Leon Lagnado. Calcium and Protein Kinase C Regulate the Actin Cytoskeleton in the Synaptic Terminal of Retinal Bipolar Cells. Journal of Cell Biology, 143(6): 1661–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1661.

Luo:1997:NGI

[jLC97] Wen jie Luo and Amy Chang. Novel Genes Involved in Endosomal Traffic in Yeast Revealed by Suppression of a Targetingdefective Plasma Membrane ATPase Mutant. Journal of Cell Biology, 138(4):731–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/4/731.

John:1998:DMS

[JLC98] Linu M. John, James D. Lechleiter, and Patricia Camacho. Differential Modulation of SERCA2 Isoforms by Calreticulin. Journal of Cell Biology, 142(4):963–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/4/963.

Jesse:1998:IRF

[JLID98] Traci L. Jesse, Rhonda LaChance, Michael F. Iademarco, and Douglas C. Dean. Interferon Regulatory Factor-2 Is a Transcriptional Activator in Muscle Where It Regulates Expression of Vascular Cell Adhesion Molecule-1. Journal of Cell Biology, 140(5):1265–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/5/1265.

Janes:1999:ALR

[JLM99] Peter W. Janes, Steven C. Ley, and Anthony I. Magee. Aggregation of Lipid Rafts Accompanies Signaling via the T Cell Antigen Receptor. Journal of Cell Biology, 147(2):447–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/2/447.

Jones:1995:RNE

[JLRS95] S. Jones, R. J. Litt, C. J. Richardson, and N. Segev. Requirement of nucleotide exchange factor for Ypt1 GTPase mediated protein transport. Journal of Cell Biology, 130(5): 1051–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/5/1051.

Janicki:1997:IAA

[JM97a] Susan Janicki and Mervyn J. Monteiro. Increased Apoptosis Arising from Increased Expression of the Alzheimer’s Disease– associated Presenilin-2 Mutation (N141I). Journal of Cell Biology, 139(2):485–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/2/485.

Jontes:1997:BBM

[JM97b] James D. Jontes and Ronald A. Milligan. Brush Border Myosin–I Structure and ADP-dependent Conformational Changes Revealed by Cryoelectron Microscopy and Image Analysis. Journal of Cell Biology, 139(3):683–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/683.

Jiang:1998:CFT

[JMD^{+}98] Qinshi Jiang, Daniel Mak, Sreenivas Devidas, Erik M. Schwiebert, Alvina Bragin, Yulong Zhang, William R. Skach, William B. Guggino, J. Kevin Foskett, and John F. Engelhardt. Cystic Fibrosis Transmembrane Conductance Regulator–associated ATP Release Is Controlled by a Chloride Sensor. Journal of Cell Biology, 143(3):645–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/645.

Jeng:1998:TCC

[JMM^{+}98] Chung-Jiuan Jeng, Steven A. McCarroll, Thomas F. J. Martin, Erik Floor, James Adams, David Krantz, Stefan Butz, Robert Edwards, and Erik S. Schweitzer. Thy-1 Is a Component Common to Multiple Populations of Synaptic Vesicles. Journal of Cell Biology, 140(3):685–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/3/685.

Jedd:1997:TNY

[JMS97] Gregory Jedd, Jon Mulholland, and Nava Segev. Two New Ypt GTPases Are Required for Exit From the Yeast trans -Golgi Compartment. Journal of Cell Biology, 137(3):563–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
3/563.

Jordan:1996:VET

[JMTCF96] P. Jordan, M. Mannervik, L. Tora, and M. Carmo-Fonseca. In vivo evidence that TATA-binding protein/SL1 colocalizes with UBF and RNA polymerase I when rRNA synthesis is either active or inactive. Journal of Cell Biology, 133(2):225–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
2/225.

Jenkins:1998:CHV

[JMWG98] Yonchu Jenkins, Michele McEntee, Karsten Weis, and Warner C. Greene. Characterization of HIV-1 Vpr Nuclear Import: Analysis of Signals and Pathways. Journal of Cell Biology, 143(4):875–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/4/875.

Jou:1998:ERE

[JN98] Tzuu-Shuh Jou and W. James Nelson. Effects of Regulated Expression of Mutant RhoA and Rac1 Small GTPases on the Development of Epithelial (MDCK) Cell Polarity. Journal of Cell Biology, 142(1):85–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/1/85.

Jontes:1998:TDS

[JOPM98] James D. Jontes, E. Michael Ostap, Thomas D. Pollard, and Ronald A. Milligan. Three-dimensional Structure of Acanthamoeba castellanii Myosin–IB (MIB) Determined by Cryoelectron Microscopy of Decorated Actin Filaments. Journal of Cell Biology, 141(1):155–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/1/155.

Jackson:1998:RCS

[JP98] Dean A. Jackson and Ana Pombo. Replicon Clusters Are Stable Units of Chromosome Structure: Evidence That Nuclear Organization Contributes to the Efficient Activation and Propagation of S Phase in Human Cells. Journal of Cell Biology, 140(6):1285–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/6/1285.

Jiang:1998:IMP

[JR98] Liwen Jiang and John C. Rogers. Integral Membrane Protein Sorting to Vacuoles in Plant Cells: Evidence for Two Pathways. Journal of Cell Biology, 143(5):1183–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1183.

Jedd:1995:YGE

[JRLS95] G. Jedd, C. Richardson, R. Litt, and N. Segev. The Ypt1 GTPase is essential for the first two steps of the yeast secretory pathway. Journal of Cell Biology, 131(3):583–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/583.

Jou:1998:SFR

[JSN98] Tzuu-Shuh Jou, Eveline E. Schneeberger, and W. James Nelson. Structural and Functional Regulation of Tight Junctions by RhoA and Rac1 Small GTPases. Journal of Cell Biology, 142(1):101–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/1/101.

James:1996:ILG

[JSR96] P. L. James, C. E. Stewart, and P. Rotwein. Insulinlike growth factor binding protein-5 modulates muscle differentiation through an insulin-like growth factor-dependent mechanism. Journal of Cell Biology, 133(3):683–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/683.

Jolly:1999:IIA

[JVRNM99] Caroline Jolly, Claire Vourc’h, Michel Robert-Nicoud, and Richard I. Morimoto. Intron-independent Association of Splicing Factors with Active Genes. Journal of Cell Biology, 145 (6):1133–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/6/1133.

Jones:1999:ESM

[JWFS99] Frank E. Jones, Thomas Welte, Xin-Yuan Fu, and David F. Stern. Erbb4 Signaling in the Mammary Gland Is Required for Lobuloalveolar Development and Stat5 Activation during Lactation. Journal of Cell Biology, 147(1):77–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/147/1/77.

Jung:1996:DML

[JWH96] G. Jung, X. Wu, and J. A. Hammer. Dictyostelium mutants lacking multiple classic myosin I isoforms reveal combinations of shared and distinct functions. Journal of Cell Biology, 133 (2):305–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/2/305.

Johnston:1998:ACR

[JWK98] Jennifer A. Johnston, Cristina L. Ward, and Ron R. Kopito. Aggresomes: a Cellular Response to Misfolded Proteins. Journal of Cell Biology, 143(7):1883–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/7/1883.

Jacobsen:1996:RCI

[JWR96] M. D. Jacobsen, M. Weil, and M. C. Raff. Role of Ced-3/ ICE-family proteases in staurosporine-induced programmed cell death. Journal of Cell Biology, 133(5):1041–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
5/1041.

Jarrous:1999:LNC

[JWW^{+}99] Nayef Jarrous, Joseph S. Wolenski, Donna Wesolowski, Christopher Lee, and Sidney Altman. Localization in the Nucleolus and Coiled Bodies of Protein Subunits of the Ribonucleoprotein Ribonuclease P. Journal of Cell Biology, 146 (3):559–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/3/559.

Kim:1995:CBA

[KA95] P. S. Kim and P. Arvan. Calnexin and BiP act as sequential molecular chaperones during thyroglobulin folding in the endoplasmic reticulum. Journal of Cell Biology, 128(1):29–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/1/29.

Korff:1998:IEC

[KA98] Thomas Korff and Hellmut G. Augustin. Integration of Endothelial Cells in Multicellular Spheroids Prevents Apoptosis and Induces Differentiation. Journal of Cell Biology, 143(5): 1341–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/5/1341.

Kundra:1995:CRP

[KAAFZ95] V. Kundra, B. Anand-Apte, L. A. Feig, and B. R. Zetter. The chemotactic response to PDGF–BB: evidence of a role for Ras. Journal of Cell Biology, 130(3):725–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/3/725.

Kronert:1999:SMH

[KAFB99] William A. Kronert, Angel Acebes, Alberto Ferrús, and Sanford I. Bernstein. Specific Myosin Heavy Chain Mutations Suppress Troponin I Defects in Drosophila Muscles. Journal of Cell Biology, 144(5):989–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/5/989.

Kelleher:1995:SSM

[KAP95] J. F. Kelleher, S. J. Atkinson, and T. D. Pollard. Sequences, structural models, and cellular localization of the actin-related proteins Arp2 and Arp3 from Acanthamoeba. Journal of Cell Biology, 131(2):385–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/2/385.

Krebs:1999:AMA

[KAS^{+}99] Joseph F. Krebs, Robert C. Armstrong, Anu Srinivasan, Teresa Aja, Angela M. Wong, Aileen Aboy, Rob Sayers, Bryan Pham, Tam Vu, Kim Hoang, Donald S. Karanewsky, Christian Leist, Albert Schmitz, Joe C. Wu, Kevin J. Tomaselli, and Lawrence C. Fritz. Activation of Membrane-associated Procaspase-3 Is Regulated by Bcl-2. Journal of Cell Biology, 144(5):915–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/5/915.

Kelly:1995:MLC

[KAT^{+}95] R. Kelly, S. Alonso, S. Tajbakhsh, G. Cossu, and M. Buckingham. Myosin light chain 3F regulatory sequences confer regionalized cardiac and skeletal muscle expression in transgenic mice. Journal of Cell Biology, 129(2):383–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
2/383.

Keresztes:1999:IAG

[KB99a] Margit Keresztes and Johannes Boonstra. Import(ance) of Growth Factors in(to) the Nucleus. Journal of Cell Biology, 145(3):421–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/3/421.

Koehnle:1999:SAT

[KB99b] Thomas J. Koehnle and Anthony Brown. Slow Axonal Transport of Neurofilament Protein in Cultured Neurons. Journal of Cell Biology, 144(3):447–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/3/447.

Kowalczyk:1997:ATD

[KBB^{+}97a] Andrew P. Kowalczyk, Elayne A. Bornslaeger, Jeffrey E. Borgwardt, Helena L. Palka, Avninder S. Dhaliwal, Connie M. Corcoran, Mitchell F. Denning, and Kathleen J. Green. The Amino-terminal Domain of Desmoplakin Binds to Plakoglobin and Clusters Desmosomal Cadherin–Plakoglobin Complexes. Journal of Cell Biology, 139(3):773–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/773.

Kreft:1997:LCM

[KBB^{+}97b] Bertolt Kreft, Dietmar Berndorff, Anja Böttinger, Silvia Finnemann, Doris Wedlich, Michael Hortsch, Rudolf Tauber, and Reinhard Geßner. LI-Cadherin-mediated cell–cell adhesion does not require cytoplasmic interactions. Journal of Cell Biology, 136(5):1109–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/5/1109.

Kwon:1999:CAR

[KBC99] Sunjong Kwon, Elisa Barbarese, and John H. Carson. The Cis–Acting RNA Trafficking Signal from Myelin Basic Protein mRNA and Its Cognate Trans–Acting Ligand Hnrnp A2 Enhance CaP–Dependent Translation. Journal of Cell Biology, 147(2):247–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/2/247.

Kataoka:1999:TSN

[KBD99] Naoyuki Kataoka, Jennifer L. Bachorik, and Gideon Dreyfuss. Transportin-SR, a Nuclear Import Receptor for SR Proteins. Journal of Cell Biology, 145(6):1145–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/6/1145.

Kok:1998:PBB

[KBF^{+}98] Jan Willem Kok, Teresa Babia, Catalin M. Filipeanu, Adriaan Nelemans, Gustavo Egea, and Dick Hoekstra. PDMP Blocks Brefeldin A–induced Retrograde Membrane Transport from Golgi to ER: Evidence for Involvement of Calcium Homeostasis and Dissociation from Sphingolipid Metabolism. Journal of Cell Biology, 142(1):25–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/1/25.

Kappe:1999:CGM

[KBG^{+}99] Stefan Kappe, Thomas Bruderer, Soren Gantt, Hisashi Fujioka, Victor Nussenzweig, and Robert Ménard. Conservation of a Gliding Motility and Cell Invasion Machinery in Apicomplexan Parasites. Journal of Cell Biology, 147(5): 937–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/5/937.

Kirisako:1999:FPA

[KBI^{+}99] Takayoshi Kirisako, Misuzu Baba, Naotada Ishihara, Kouichi Miyazawa, Mariko Ohsumi, Tamotsu Yoshimori, Takeshi Noda, and Yoshinori Ohsumi. Formation process of autophagosome is traced with Apg8/Aut7p in yeast. Journal of Cell Biology, 147(2):435–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/147/2/435.

Koch:1995:LSS

[KBM^{+}95] M. Koch, B. Bohrmann, M. Matthison, C. Hagios, B. Trueb, and M. Chiquet. Large and small splice variants of collagen XII: differential expression and ligand binding. Journal of Cell Biology, 130(4):1005–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/4/1005.

Kozminski:1995:CKL

[KBR95] K. G. Kozminski, P. L. Beech, and J. L. Rosenbaum. The Chlamydomonas kinesin-like protein FLA10 is involved in motility associated with the flagellar membrane. Journal of Cell Biology, 131(6):1517–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1517.

Klages:1999:AGG

[KBS^{+}99] Birgit Klages, Ursula Brandt, Melvin I. Simon, Günter Schultz, and Stefan Offermanns. Activation of G_{12} /G_{13} results in shape change and rho/rho-kinase-mediated myosin light chain phosphorylation in mouse platelets. Journal of Cell Biology, 144(4):745–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/4/745.

Kues:1995:LNR

[KBSW95] W. A. Kues, H. R. Brenner, B. Sakmann, and V. Witzemann. Local neurotrophic repression of gene transcripts encoding fetal AChRs at rat neuromuscular synapses. Journal of Cell Biology, 130(4):949–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/4/949.

Ktistakis:1996:EPD

[KBW^{+}96] N. T. Ktistakis, H. A. Brown, M. G. Waters, P. C. Sternweis, and M. G. Roth. Evidence that phospholipase D mediates ADP ribosylation factor-dependent formation of Golgi coated vesicles. Journal of Cell Biology, 134(2):295–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/295.

Kahana:1999:BNT

[KC99] Jason A. Kahana and Don W. Cleveland. Beyond Nuclear Transport. Journal of Cell Biology, 146(6):1205–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
6/1205.

Khodjakov:1996:FPC

[KCBR96] A. Khodjakov, R. W. Cole, A. S. Bajer, and C. L. Rieder. The force for poleward chromosome motion in Haemanthus cells acts along the length of the chromosome during metaphase but only at the kinetochore during anaphase. Journal of Cell Biology, 132(6):1093–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/6/1093.

Kinch:1995:TPR

[KCDB95] M. S. Kinch, G. J. Clark, C. J. Der, and K. Burridge. Tyrosine phosphorylation regulates the adhesions of ras-transformed breast epithelia. Journal of Cell Biology, 130(2):461–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/461.

Kouranov:1998:TTN

[KCFS98] Andrei Kouranov, Xuejun Chen, Bruno Fuks, and Danny J. Schnell. Tic20 and Tic22 Are New Components of the Protein Import Apparatus at the Chloroplast Inner Envelope Membrane. Journal of Cell Biology, 143(4):991–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/991.

Klemke:1997:RCM

[KCG^{+}97] Richard L. Klemke, Shuang Cai, Ana L. Giannini, Patricia J. Gallagher, Primal de Lanerolle, and David A. Cheresh. Regulation of Cell Motility by Mitogen-activated Protein Kinase. Journal of Cell Biology, 137(2):481–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/481.

Khodjakov:1997:CFP

[KCM^{+}97] Alexey Khodjakov, Richard W. Cole, Bruce F. McEwen, Karolyn F. Buttle, and Conly L. Rieder. Chromosome Fragments Possessing Only One Kinetochore Can Congress to the Spindle Equator. Journal of Cell Biology, 136(2):229–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/2/229.

Knibbs:1996:FFV

[KCN^{+}96] R. N. Knibbs, R. A. Craig, S. Natsuka, A. Chang, M. Cameron, J. B. Lowe, and L. M. Stoolman. The fucosyltransferase FucT– VII regulates E-selectin ligand synthesis in human T cells. Journal of Cell Biology, 133(4):911–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/4/911.

Kitzmann:1998:MRF

[KCV^{+}98] Magali Kitzmann, Gilles Carnac, Marie Vandromme, Michael Primig, Ned J. C. Lamb, and Anne Fernandez. The Muscle Regulatory Factors MyoD and Myf-5 Undergo Distinct Cell Cycle–specific Expression in Muscle Cells. Journal of Cell Biology, 142(6):1447–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/142/6/1447.

Kislauskis:1997:AMR

[KcZS97] Edward H. Kislauskis, Xiao chun Zhu, and Robert H. Singer. β-actin Messenger RNA Localization and Protein Synthesis Augment Cell Motility. Journal of Cell Biology, 136(6): 1263–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/6/1263.

Khwaja:1997:LCB

[KD97a] Asim Khwaja and Julian Downward. Lack of correlation between activation of Jun–NH_{2}-terminal kinase and induction of apoptosis after detachment of epithelial cells. Journal of Cell Biology, 139(4):1017–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/4/1017.

King:1997:PID

[KD97b] Stephen J. King and Susan K. Dutcher. Phosphoregulation of an Inner Dynein Arm Complex in Chlamydomonas reinhardtii Is Altered in Phototactic Mutant Strains. Journal of Cell Biology, 136(1):177–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/1/177.

Kim:1997:SFA

[KDBW97] Euikyung Kim, Lei Du, David B. Bregman, and Stephen L. Warren. Splicing Factors Associate with Hyperphosphorylated RNA Polymerase II in the Absence of Pre-mRNA. Journal of Cell Biology, 136(1):19–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/1/19.

Kehlenbach:1998:NSF

[KDG98] Ralph H. Kehlenbach, Achim Dickmanns, and Larry Gerace. Nucleocytoplasmic Shuttling Factors Including Ran and CRM1 Mediate Nuclear Export of NFAT In Vitro. Journal of Cell Biology, 141(4):863–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/863.

Kehlenbach:1999:RRR

[KDK^{+}99] Ralph H. Kehlenbach, Achim Dickmanns, Angelika Kehlenbach, Tinglu Guan, and Larry Gerace. A Role for RanBP1 in the Release of CRM1 from the Nuclear Pore Complex in a Terminal Step of Nuclear Export. Journal of Cell Biology, 145(4):645–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/645.

Kiosses:1999:RPA

[KDO^{+}99] William B. Kiosses, R. Hugh Daniels, Carol Otey, Gary M. Bokoch, and Martin Alexander Schwartz. A Role for P21-Activated Kinase in Endothelial Cell Migration. Journal of Cell Biology, 147(4):831–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/4/831.

Kasper:1996:NTM

[KDvFP96] D. Kasper, F. Dittmer, K. von Figura, and R. Pohlmann. Neither type of mannose 6-phosphate receptor is sufficient for targeting of lysosomal enzymes along intracellular routes. Journal of Cell Biology, 134(3):615–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/3/615.

Kenworthy:1998:DGA

[KE98] A. K. Kenworthy and M. Edidin. Distribution of a glycosylphosphatidylinositol-anchored protein at the apical surface of MDCK cells examined at a resolution of < 100å using imaging fluorescence resonance energy transfer. Journal of Cell Biology, 142(1):69–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/1/69.

Kallunki:1997:TSE

[KEJ97] Pekka Kallunki, Gerald M. Edelman, and Frederick S. Jones. Tissue-specific Expression of the L1 Cell Adhesion Molecule Is Modulated by the Neural Restrictive Silencer Element. Journal of Cell Biology, 138(6):1343–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/6/1343.

Kluck:1999:PAP

[KEP^{+}99] Ruth M. Kluck, Mauro Degli Esposti, Guy Perkins, Christian Renken, Tomomi Kuwana, Ella Bossy-Wetzel, Martin Goldberg, Terry Allen, Michael J. Barber, Douglas R. Green, and Donald D. Newmeyer. The Pro–Apoptotic Proteins, Bid and Bax, Cause a Limited Permeabilization of the Mitochondrial Outer Membrane That Is Enhanced by Cytosol. Journal of Cell Biology, 147(4):809–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/4/809.

Kussel:1995:PDP

[KF95] P. Küssel and M. Frasch. Pendulin, a Drosophila protein with cell cycle-dependent nuclear localization, is required for normal cell proliferation. Journal of Cell Biology, 129(6):1491–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/6/1491.

Kolman:1996:DMH

[KFE96] M. F. Kolman, L. M. Futey, and T. T. Egelhoff. Dictyostelium myosin heavy chain kinase A regulates myosin localization during growth and development. Journal of Cell Biology, 132(1): 101–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/1/101.

Kolossov:1998:FCC

[KFL^{+}98] E. Kolossov, B. K. Fleischmann, Q. Liu, W. Bloch, S. Viatchenko-Karpinski, O. Manzke, G. J. Ji, H. Bohlen, K. Addicks, and J. Hescheler. Functional Characteristics of ES Cell–derived Cardiac Precursor Cells Identified by Tissuespecific Expression of the Green Fluorescent Protein. Journal of Cell Biology, 143(7):2045–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/7/2045.

Kachinsky:1999:PCS

[KFM99] Amy M. Kachinsky, Stanley C. Froehner, and Sharon L. Milgram. A PDZ-containing Scaffold Related to the Dystrophin Complex at the Basolateral Membrane of Epithelial Cells. Journal of Cell Biology, 145(2):391–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/2/391.

Kawano:1999:PMB

[KFO^{+}99] Yoji Kawano, Yuko Fukata, Noriko Oshiro, Mutsuki Amano, Toshikazu Nakamura, Masaaki Ito, Fumio Matsumura, Masaki Inagaki, and Kozo Kaibuchi. Phosphorylation of Myosin– Binding Subunit (Mbs) of Myosin Phosphatase by Rho–Kinase in Vivo. Journal of Cell Biology, 147(5):1023–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
5/1023.

Kean:1997:PMT

[KGA^{+}97] Leslie S. Kean, Althea M. Grant, Cesar Angeletti, Yannick Mahé, Karl Kuchler, Robert S. Fuller, and J. Wylie Nichols. Plasma Membrane Translocation of Fluorescent-labeled Phosphatidylethanolamine Is Controlled by Transcription Regulators, PDR1 and PDR3. Journal of Cell Biology, 138(2):255–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/2/255.

Kumagai:1998:XCP

[KGE^{+}98] Akiko Kumagai, Zijian Guo, Katayoon H. Emami, Sophie X. Wang, and William G. Dunphy. The Xenopus Chk1 Protein Kinase Mediates a Caffeine-sensitive Pathway of Checkpoint Control in Cell-free Extracts. Journal of Cell Biology, 142(6): 1559–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/6/1559.

Kang:1997:COS

[KGF^{+}97] Jong-Sun Kang, Min Gao, Jessica L. Feinleib, Philip D. Cotter, Sarah N. Guadagno, and Robert S. Krauss. CDO: an oncogene-, serum-, and anchorage-regulated member of the Ig/fibronectin Type III repeat family. Journal of Cell Biology, 138(1):203–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/1/203.

Kramer:1999:CBE

[KGGK99] Angela Krämer, Patric Grüter, Karsten Gröning, and Berthold Kastner. Combined Biochemical and Electron Microscopic Analyses Reveal the Architecture of the Mammalian U2 snRNP. Journal of Cell Biology, 145(7):1355–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/145/7/
1355.

Kromer:1998:ERD

[KGHG98] Andreas Krömer, Michael M. Glombik, Wieland B. Huttner, and Hans-Hermann Gerdes. Essential Role of the Disulfidebonded Loop of Chromogranin B for Sorting to Secretory Granules Is Revealed by Expression of a Deletion Mutant in the Absence of Endogenous Granin Synthesis. Journal of Cell Biology, 140(6):1331–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/6/1331.

Kreitmeier:1995:THD

[KGHMT95] M. Kreitmeier, G. Gerisch, C. Heizer, and A. Müller-Taubenberger. A talin homologue of Dictyostelium rapidly assembles at the leading edge of cells in response to chemoattractant. Journal of Cell Biology, 129(1):179–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/1/
179.

Kedersha:1999:RBP

[KGL^{+}99] Nancy L. Kedersha, Mita Gupta, Wei Li, Ira Miller, and Paul Anderson. RNA-binding Proteins Tia-1 and Tiar Link the Phosphorylation of Eif-2α to the Assembly of Mammalian Stress Granules. Journal of Cell Biology, 147(7):1431–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/7/1431.

Korioth:1995:MCN

[KGMF95] F. Korioth, C. Gieffers, G. G. Maul, and J. Frey. Molecular characterization of NDP52, a novel protein of the nuclear domain 10, which is redistributed upon virus infection and interferon treatment. Journal of Cell Biology, 130(1):1–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
1/1.

Kurvari:1998:GSS

[KGS98] Venkatesh Kurvari, Nick V. Grishin, and William J. Snell. A Gamete-specific, Sex-limited Homeodomain Protein in Chlamydomonas. Journal of Cell Biology, 143(7):1971–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/7/1971.

Koval:1995:TCA

[KGW^{+}95] M. Koval, S. T. Geist, E. M. Westphale, A. E. Kemendy, R. Civitelli, E. C. Beyer, and T. H. Steinberg. Transfected connexin45 alters gap junction permeability in cells expressing endogenous connexin43. Journal of Cell Biology, 130(4): 987–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/987.

Katayama:1999:FYG

[KHA^{+}99] Satoshi Katayama, Dai Hirata, Manuel Arellano, Pilar Pérez, and Takashi Toda. Fission Yeast α-Glucan Synthase Mok1 Requires the Actin Cytoskeleton to Localize the Sites of Growth and Plays an Essential Role in Cell Morphogenesis Downstream of Protein Kinase C Function. Journal of Cell Biology, 144(6):1173–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/6/1173.

Koval:1997:CRM

[KHHS97] Michael Koval, James E. Harley, Elizabeth Hick, and Thomas H. Steinberg. Connexin46 Is Retained as Monomers in a trans -Golgi Compartment of Osteoblastic Cells. Journal of Cell Biology, 137(4):847–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/4/847.

Kerscher:1997:TTC

[KHS^{+}97] Oliver Kerscher, Jason Holder, Maithreyan Srinivasan, Roxanne S. Leung, and Robert E. Jensen. The Tim54p–Tim22p Complex Mediates Insertion of Proteins into the Mitochondrial Inner Membrane. Journal of Cell Biology, 139(7):1663–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/7/1663.

Koenig:1996:SVT

[KI96] J. H. Koenig and K. Ikeda. Synaptic vesicles have two distinct recycling pathways. Journal of Cell Biology, 135(3): 797–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/797.

Katahira:1997:MCF

[KIH^{+}97] Jun Katahira, Norimitsu Inoue, Yasuhiko Horiguchi, Morihiro Matsuda, and Nakaba Sugimoto. Molecular Cloning and Functional Characterization of the Receptor for Clostridium perfringens Enterotoxin. Journal of Cell Biology, 136(6):1239–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/6/1239.

Kose:1997:RUN

[KIT^{+}97] Shingo Kose, Naoko Imamoto, Taro Tachibana, Takuya Shimamoto, and Yoshihiro Yoneda. Ran-unassisted Nuclear Migration of a 97-kD Component of Nuclear Pore–targeting Complex. Journal of Cell Biology, 139(4):841–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
4/841.

Kono:1996:CCA

[KJBM^{+}96] T. Kono, K. T. Jones, A. Bos-Mikich, D. G. Whittingham, and J. Carroll. A cell cycle-associated change in Ca^{2+} releasing activity leads to the generation of Ca^{2+} transients in mouse embryos during the first mitotic division. Journal of Cell Biology, 132(5):915–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/5/915.

Karakesisoglou:1999:ISD

[KJE^{+}99] Iakowos Karakesisoglou, Klaus-Peter Janssen, Ludwig Eichinger, Angelika A. Noegel, and Michael Schleicher. Identification of a suppressor of the dictyostelium profilin-minus phenotype as a CD36/LIMP–II homologue. Journal of Cell Biology, 145 (1):167–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/1/167.

Kudla:1998:FRT

[KJR^{+}98] Arthur J. Kudla, Nathan C. Jones, R. Scott Rosenthal, Kirstin Arthur, Kari L. Clase, and Bradley B. Olwin. The FGF Receptor–1 Tyrosine Kinase Domain Regulates Myogenesis but Is Not Sufficient to Stimulate Proliferation. Journal of Cell Biology, 142(1):241–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/1/241.

Kim:1996:ERS

[KKA96] P. S. Kim, O. Y. Kwon, and P. Arvan. An endoplasmic reticulum storage disease causing congenital goiter with hypothyroidism. Journal of Cell Biology, 133(3):517–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/517.

Kanai:1996:ITS

[KKAN^{+}96] Y. Kanai, M. Kanai-Azuma, T. Noce, T. C. Saido, T. Shiroishi, Y. Hayashi, and K. Yazaki. Identification of two Sox17 messenger RNA isoforms, with and without the high mobility group box region, and their differential expression in mouse spermatogenesis. Journal of Cell Biology, 133(3):667–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/667.

Kadoya:1995:AAD

[KKD^{+}95] Y. Kadoya, K. Kadoya, M. Durbeej, K. Holmvall, L. Sorokin, and P. Ekblom. Antibodies against domain E3 of laminin-1 and integrin alpha 6 subunit perturb branching epithelial morphogenesis of submandibular gland, but by different modes. Journal of Cell Biology, 129(2):521–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/2/521.

Kuriyama:1995:CME

[KKE^{+}95] R. Kuriyama, M. Kofron, R. Essner, T. Kato, S. Dragas-Granoic, C. K. Omoto, and A. Khodjakov. Characterization of a minus end-directed kinesin-like motor protein from cultured mammalian cells. Journal of Cell Biology, 129(4):1049–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/4/1049.

Kellogg:1995:MNS

[KKFN^{+}95] D. R. Kellogg, A. Kikuchi, T. Fujii-Nakata, C. W. Turck, and A. W. Murray. Members of the NAP/SET family of proteins interact specifically with B-type cyclins. Journal of Cell Biology, 130(3):661–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/3/661.

Karaoglu:1995:FCO

[KKG95] D. Karaoglu, D. J. Kelleher, and R. Gilmore. Functional characterization of Ost3p. Loss of the 34-kD subunit of the Saccharomyces cerevisiae oligosaccharyltransferase results in biased underglycosylation of acceptor substrates. Journal of Cell Biology, 130(3):567–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/3/567.

Klumperman:1998:MPR

[KKG^{+}98] Judith Klumperman, Regina Kuliawat, Janice M. Griffith, Hans J. Geuze, and Peter Arvan. Mannose 6–Phosphate Receptors Are Sorted from Immature Secretory Granules via Adaptor Protein AP-1, Clathrin, and Syntaxin 6–positive Vesicles. Journal of Cell Biology, 141(2):359–??, April 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/2/
359.

Kielian:1996:MMI

[KKGD96] M. Kielian, M. R. Klimjack, S. Ghosh, and W. A. Duffus. Mechanisms of mutations inhibiting fusion and infection by Semliki Forest virus. Journal of Cell Biology, 134 (4):863–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/863.

Karlsson:1999:CCD

[KKH^{+}99] Christina Karlsson, Stephanie Katich, Anja Hagting, Ingrid Hoffmann, and Jonathon Pines. Cdc25b and Cdc25c Differ Markedly in Their Properties as Initiators of Mitosis. Journal of Cell Biology, 146(3):573–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/3/573.

Kuliawat:1997:DSL

[KKLA97] Regina Kuliawat, Judith Klumperman, Thomas Ludwig, and Peter Arvan. Differential Sorting of Lysosomal Enzymes Out of the Regulated Secretory Pathway in Pancreatic β-Cells. Journal of Cell Biology, 137(3):595–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/3/595.

Kanaseki:1997:SFM

[KKM^{+}97] Toku Kanaseki, Kazunori Kawasaki, Masayuki Murata, Yoko Ikeuchi, and Shun ichi Ohnishi. Structural Features of Membrane Fusion between Influenza Virus and Liposome as Revealed by Quick–Freezing Electron Microscopy. Journal of Cell Biology, 137(5):1041–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/5/1041.

Kaverina:1999:MTS

[KKS99] Irina Kaverina, Olga Krylyshkina, and J. Victor Small. Microtubule Targeting of Substrate Contacts Promotes Their Relaxation and Dissociation. Journal of Cell Biology, 146(5): 1033–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/1033.

Korpelainen:1998:OVT

[KKT^{+}98] Eija I. Korpelainen, Marika J. Karkkainen, Auri Tenhunen, Merja Lakso, Heikki Rauvala, Matti Vierula, Martti Parvinen, and Kari Alitalo. Overexpression of VEGF in Testis and Epididymis Causes Infertility in Transgenic Mice: Evidence for Nonendothelial Targets for VEGF. Journal of Cell Biology, 143(6):1705–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1705.

Karki:1998:CPS

[KLH98] Sher Karki, Bernadette LaMonte, and Erika L. F. Holzbaur. Characterization of the p22 Subunit of Dynactin Reveals the Localization of Cytoplasmic Dynein and Dynactin to the Midbody of Dividing Cells. Journal of Cell Biology, 142(4): 1023–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/4/1023.

Krauss:1997:SPN

[KLL^{+}97] Sharon Wald Krauss, Carolyn A. Larabell, Stephen Lockett, Philippe Gascard, Sheldon Penman, Narla Mohandas, and Joel Anne Chasis. Structural Protein 4.1 in the Nucleus of Human Cells: Dynamic Rearrangements during Cell Division. Journal of Cell Biology, 137(2):275–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/275.

Klemke:1998:CCC

[KLM^{+}98] Richard L. Klemke, Jie Leng, Rachel Molander, Peter C. Brooks, Kristiina Vuori, and David A. Cheresh. CAS/ crk Coupling Serves as a “Molecular Switch” for Induction of Cell Migration. Journal of Cell Biology, 140(4):961–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
4/961.

Killick:1995:MCC

[KLMR95] R. Killick, P. K. Legan, C. Malenczak, and G. P. Richardson. Molecular cloning of chick beta-tectorin, an extracellular matrix molecule of the inner ear. Journal of Cell Biology, 129 (2):535–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/535.

Kurz:1996:AIG

[KLN^{+}96] A. Kurz, S. Lampel, J. E. Nickolenko, J. Bradl, A. Benner, R. M. Zirbel, T. Cremer, and P. Lichter. Active and inactive genes localize preferentially in the periphery of chromosome territories. Journal of Cell Biology, 135(5):1195–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
5/1195.

Kim:1998:NBN

[KLO^{+}98] Tae-Aug Kim, Jinkyu Lim, Setsuo Ota, Sandhya Raja, Rick Rogers, Benjamin Rivnay, Hava Avraham, and Shalom Avraham. NRP/ b, a Novel Nuclear Matrix Protein, Associates With p110^{RB} and Is Involved in Neuronal Differentiation. Journal of Cell Biology, 141(3):553–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/3/553.

Kost:1999:RHC

[KLS^{+}99] Benedikt Kost, Emmanuel Lemichez, Pius Spielhofer, Yan Hong, Kimberly Tolias, Christopher Carpenter, and Nam-Hai Chua. Rac Homologues and Compartmentalized Phosphatidylinositol 4, 5-Bisphosphate Act in a Common Pathway to Regulate Polar Pollen Tube Growth. Journal of Cell Biology, 145(2):317–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/2/317.

Kim:1996:DMP

[KLW^{+}96] J. H. Kim, C. A. Lingwood, D. B. Williams, W. Furuya, M. F. Manolson, and S. Grinstein. Dynamic measurement of the pH of the Golgi complex in living cells using retrograde transport of the verotoxin receptor. Journal of Cell Biology, 134(6): 1387–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1387.

Kellogg:1995:NAC

[KM95] D. R. Kellogg and A. W. Murray. NAP1 acts with Clb1 to perform mitotic functions and to suppress polar bud growth in budding yeast. Journal of Cell Biology, 130(3):675–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
3/675.

Kleijmeer:1997:MHC

[KMG^{+}97] Monique J. Kleijmeer, Stanislaw Morkowski, Janice M. Griffith, Alexander Y. Rudensky, and Hans J. Geuze. Major Histocompatibility Complex Class II Compartments in Human and Mouse B Lymphoblasts Represent Conventional Endocytic Compartments. Journal of Cell Biology, 139(3):639– ??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/3/639.

Kato-Minoura:1997:CIA

[KMHK97] Takako Kato-Minoura, Masafumi Hirono, and Ritsu Kamiya. Chlamydomonas Inner–Arm Dynein Mutant, ida5, Has a Mutation in an Actin-encoding Gene. Journal of Cell Biology, 137(3):649–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/3/649.

Koch:1997:TDP

[KMI^{+}97] Peter J. Koch, M. G. Mahoney, Hiroyasu Ishikawa, Leena Pulkkinen, Jouni Uitto, Leonard Shultz, George F. Murphy, Diana Whitaker-Menezes, and John R. Stanley. Targeted disruption of the Pemphigus Vulgaris antigen (desmoglein 3) gene in mice causes loss of keratinocyte cell adhesion with a phenotype similar to Pemphigus Vulgaris. Journal of Cell Biology, 137(5):1091–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/5/1091.

Kuznetsov:1998:FIM

[KMK^{+}98] Andrey V. Kuznetsov, Oleg Mayboroda, Dagmar Kunz, Kirstin Winkler, Walter Schubert, and Wolfram S. Kunz. Functional Imaging of Mitochondria in Saponin-permeabilized Mice Muscle Fibers. Journal of Cell Biology, 140(5):1091–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/5/1091.

Kang:1998:CRR

[KMM^{+}98] Jong-Sun Kang, Philip J. Mulieri, Cary Miller, David A. Sassoon, and Robert S. Krauss. CDO, A Robo-related Cell Surface Protein that Mediates Myogenic Differentiation. Journal of Cell Biology, 143(2):403–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/2/403.

Karpova:1998:AFA

[KMMC98] Tatiana S. Karpova, James G. McNally, Samuel L. Moltz, and John A. Cooper. Assembly and Function of the Actin Cytoskeleton of Yeast: Relationships between Cables and Patches. Journal of Cell Biology, 142(6):1501–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1501.

Ku:1995:CHH

[KMOO95] N. O. Ku, S. Michie, R. G. Oshima, and M. B. Omary. Chronic hepatitis, hepatocyte fragility, and increased soluble phosphoglycokeratins in transgenic mice expressing a keratin 18 conserved arginine mutant. Journal of Cell Biology, 131(5): 1303–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/5/1303.

Ku:1998:MMK

[KMS^{+}98] Nam-On Ku, Sara A. Michie, Roy M. Soetikno, Evelyn Z. Resurreccion, Rosemary L. Broome, and M. Bishr Omary. Mutation of a Major Keratin Phosphorylation Site Predisposes to Hepatotoxic Injury in Transgenic Mice. Journal of Cell Biology, 143(7):2023–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/7/2023.

Kueng:1997:NFS

[KND^{+}97] Peter Kueng, Zariana Nikolova, Valentin Djonov, Andrew Hemphill, Valeria Rohrbach, Dominik Boehlen, Gisela Zuercher, Anne-Catherine Andres, and Andrew Ziemiecki. A novel family of serine/threonine kinases participating in spermiogenesis. Journal of Cell Biology, 139(7):1851–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/7/1851.

Kim:1997:GNS

[KNH^{+}97] Eunjoon Kim, Scott Naisbitt, Yi-Ping Hsueh, Anuradha Rao, Adam Rothschild, Ann Marie Craig, and Morgan Sheng. GKAP, a novel synaptic protein that interacts with the guanylate kinase-like domain of the PSD-95/SAP90 family of channel clustering molecules. Journal of Cell Biology, 136(3): 669–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/669.

Kirsch:1997:RPM

[KNSP97] Thorsten Kirsch, Hyun-Duck Nah, Irving M. Shapiro, and Maurizio Pacifici. Regulated Production of Mineralizationcompetent Matrix Vesicles in Hypertrophic Chondrocytes. Journal of Cell Biology, 137(5):1149–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/5/1149.

Kuroda:1999:MHC

[KNT^{+}99] Shun’ichi Kuroda, Noritaka Nakagawa, Chiharu Tokunaga, Kenji Tatematsu, and Katsuyuki Tanizawa. Mammalian Homologue of the Caenorhabditis elegans UNC-76 Protein Involved in Axonal Outgrowth Is a Protein Kinase C ζ– interacting Protein. Journal of Cell Biology, 144(3):403– ??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/403.

Koch:1999:CEL

[KOA^{+}99] Manuel Koch, Pamela F. Olson, Anne Albus, William Jin, Dale D. Hunter, William J. Brunken, Robert E. Burgeson, and Marie-France Champliaud. Characterization and expression of the laminin γ3 chain: A novel, non-basement membraneassociated, laminin chain. Journal of Cell Biology, 145(3): 605–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/3/605.

Kawahara:1998:CIC

[KOM^{+}98] Atsuo Kawahara, Yoshiyuki Ohsawa, Hirotaka Matsumura, Yasuo Uchiyama, and Shigekazu Nagata. Caspase-independent Cell Killing by Fas-associated Protein with Death Domain. Journal of Cell Biology, 143(5):1353–??, November 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/143/5/
1353.

Kramer:1996:MDH

[KP96] H. Krämer and M. Phistry. Mutations in the Drosophila hook gene inhibit endocytosis of the boss transmembrane ligand into multivesicular bodies. Journal of Cell Biology, 133(6):1205–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/6/1205.

King:1995:MRI

[KPKWW95] S. M. King, R. S. Patel-King, C. G. Wilkerson, and G. B. Witman. The 78,000-M(r) intermediate chain of Chlamydomonas outer arm dynein is a microtubule-binding protein. Journal of Cell Biology, 131(2):399–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/2/399.

Koutoulis:1997:CRO

[KPW^{+}97] Anthony Koutoulis, Gregory J. Pazour, Curtis G. Wilkerson, Kazuo Inaba, Hong Sheng, Saeko Takada, and George B. Witman. The Chlamydomonas reinhardtii ODA3 Gene Encodes a Protein of the Outer Dynein Arm Docking Complex. Journal of Cell Biology, 137(5):1069–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/5/1069.

Kwan:1997:ACC

[KPZ^{+}97] Kin Ming Kwan, Michael K. M. Pang, Sheila Zhou, Soot Keng Cowan, Richard Y. C. Kong, Tim Pfordte, Bjorn R. Olsen, David O. Sillence, Patrick P. L. Tam, and Kathryn S. E. Cheah. Abnormal Compartmentalization of Cartilage Matrix Components in Mice Lacking Collagen X: Implications for Function. Journal of Cell Biology, 136(2):459–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
2/459.

Karsten:1998:PPT

[KQB^{+}98] Verena Karsten, Huilin Qi, Con J. M. Beckers, Anita Reddy, Jean-Francois Dubremetz, Paul Webster, and Keith A. Joiner. The Protozoan Parasite Toxoplasma gondii Targets Proteins to Dense Granules and the Vacuolar Space Using Both Conserved and Unusual Mechanisms. Journal of Cell Biology, 141 (6):1323–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/6/1323.

Khodjakov:1996:KMT

[KR96] A. Khodjakov and C. L. Rieder. Kinetochores moving away from their associated pole do not exert a significant pushing force on the chromosome. Journal of Cell Biology, 135(2): 315–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/2/315.

Khodjakov:1999:SRT

[KR99] Alexey Khodjakov and Conly L. Rieder. The Sudden Recruitment of γ-Tubulin to the Centrosome at the Onset of Mitosis and Its Dynamic Exchange Throughout the Cell Cycle, Do Not Require Microtubules. Journal of Cell Biology, 146(3): 585–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/3/585.

Keely:1999:RRS

[KRCP99] Patricia J. Keely, Elena V. Rusyn, Adrienne D. Cox, and Leslie V. Parise. R-Ras Signals through Specific Integrin α Cytoplasmic Domains to Promote Migration and Invasion of Breast Epithelial Cells. Journal of Cell Biology, 145(5): 1077–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/5/1077.

Kibbey:1998:LRC

[KRGA98] Richard G. Kibbey, Josep Rizo, Lila M. Gierasch, and Richard G. W. Anderson. The LDL Receptor Clustering Motif Interacts with the Clathrin Terminal Domain in a Reverse Turn Conformation. Journal of Cell Biology, 142(1):59–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
1/59.

Kalies:1998:SSC

[KRH98] Kai-Uwe Kalies, Tom A. Rapoport, and Enno Hartmann. The β Subunit of the Sec61 Complex Facilitates Cotranslational Protein Transport and Interacts with the Signal Peptidase during Translocation. Journal of Cell Biology, 141(4):887–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
4/887.

Kobayashi:1998:NMP

[KRK^{+}98] Naoto Kobayashi, Jochen Reiser, Wilhelm Kriz, Ryoko Kuriyama, and Peter Mundel. Nonuniform microtubular polarity established by CHO1/MKLP1 motor protein is necessary for process formation of podocytes. Journal of Cell Biology, 143(7):1961–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/7/1961.

Kretz-Remy:1996:IKB

[KRMMA96] C. Kretz-Remy, P. Mehlen, M. E. Mirault, and A. P. Arrigo. Inhibition of I kappa B-alpha phosphorylation and degradation and subsequent NF-kappa B activation by glutathione peroxidase overexpression. Journal of Cell Biology, 133(5):1083–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/5/1083.

Kaverina:1998:TCS

[KRS98a] Irina Kaverina, Klemens Rottner, and J. Victor Small. Targeting, Capture, and Stabilization of Microtubules at Early Focal Adhesions. Journal of Cell Biology, 142(1):181–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
1/181.

Kielty:1998:TSM

[KRS^{+}98b] Cay M. Kielty, Michael Raghunath, Linda D. Siracusa, Michael J. Sherratt, Reiner Peters, C. Adrian Shuttleworth, and Sergio A. Jimenez. The Tight Skin Mouse: Demonstration of Mutant Fibrillin-1 Production and Assembly into Abnormal Microfibrils. Journal of Cell Biology, 140(5):1159–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
5/1159.

Kouranov:1997:AIP

[KS97] Andrei Kouranov and Danny J. Schnell. Analysis of the Interactions of Preproteins with the Import Machinery over the Course of Protein Import into Chloroplasts. Journal of Cell Biology, 139(7):1677–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/7/1677.

Keller:1998:CRS

[KS98] Patrick Keller and Kai Simons. Cholesterol Is Required for Surface Transport of Influenza Virus Hemagglutinin. Journal of Cell Biology, 140(6):1357–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/6/1357.

Kashiwagi:1997:AMP

[KSE^{+}97] Hirokazu Kashiwagi, Martin A. Schwartz, Martin Eigenthaler, K. A. Davis, Mark H. Ginsberg, and Sanford J. Shattil. Affinity modulation of platelet integrin α_{IIb}β_{3} by β_{3}-endonexin, a selective binding partner of the β_{3} integrin cytoplasmic tail. Journal of Cell Biology, 137(6):1433–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/6/1433.

Koller:1999:PPP

[KSF^{+}99] Antonius Koller, William B. Snyder, Klaas Nico Faber, Thibaut J. Wenzel, Linda Rangell, Gilbert A. Keller, and Suresh Subramani. Pex22p of Pichia pastoris, essential for peroxisomal matrix protein import, anchors the ubiquitinconjugating enzyme, Pex4p, on the peroxisomal membrane. Journal of Cell Biology, 146(1):99–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/1/99.

Kelley:1996:XNM

[KSG^{+}96] C. A. Kelley, J. R. Sellers, D. L. Gard, D. Bui, R. S. Adelstein, and I. C. Baines. Xenopus nonmuscle myosin heavy chain isoforms have different subcellular localizations and enzymatic activities. Journal of Cell Biology, 134(3):675–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/3/675.

Kunz:1998:NFM

[KSG^{+}98] Stefan Kunz, Marianne Spirig, Claudia Ginsburg, Andrea Buchstaller, Philipp Berger, Rainer Lanz, Christoph Rader, Lorenz Vogt, Beat Kunz, and Peter Sonderegger. Neurite Fasciculation Mediated by Complexes of Axonin-1 and Ng Cell Adhesion Molecule. Journal of Cell Biology, 143(6): 1673–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1673.

Knudsen:1995:IAA

[KSJW95] K. A. Knudsen, A. P. Soler, K. R. Johnson, and M. J. Wheelock. Interaction of alpha-actinin with the cadherin/catenin cell–cell adhesion complex via alpha-catenin. Journal of Cell Biology, 130(1):67–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/1/67.

Keon:1996:SNT

[KSK^{+}96] B. H. Keon, S. Schäfer, C. Kuhn, C. Grund, and W. W. Franke. Symplekin, a novel type of tight junction plaque protein. Journal of Cell Biology, 134(4):1003–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/4/1003.

Kickhoefer:1999:KVP

[KSK^{+}99a] Valerie A. Kickhoefer, Amara C. Siva, Nancy L. Kedersha, Elisabeth M. Inman, Cristina Ruland, Michel Streuli, and Leonard H. Rome. The 193-Kd Vault Protein, Vparp, Is a Novel Poly(Adp–Ribose) Polymerase. Journal of Cell Biology, 146(5):917–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/5/917.

Kioka:1999:VNV

[KSK^{+}99b] Noriyuki Kioka, Shohei Sakata, Takeshi Kawauchi, Teruo Amachi, Steven K. Akiyama, Kenji Okazaki, Christopher Yaen, Kenneth M. Yamada, and Shin ichi Aota. Vinexin: a Novel Vinculin-binding Protein with Multiple SH3 Domains Enhances Actin Cytoskeletal Organization. Journal of Cell Biology, 144(1):59–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/1/59.

Kruth:1995:SAL

[KSL^{+}95] H. S. Kruth, S. I. Skarlatos, K. Lilly, J. Chang, and I. Ifrim. Sequestration of acetylated LDL and cholesterol crystals by human monocyte-derived macrophages. Journal of Cell Biology, 129(1):133–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/1/133.

Kuehn:1996:AAP

[KSL96] M. J. Kuehn, R. Schekman, and P. O. Ljungdahl. Amino acid permeases require COPII components and the ER resident membrane protein Shr3p for packaging into transport vesicles in vitro. Journal of Cell Biology, 135(3):585–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
3/585.

Kreuz:1996:AFM

[KSM96] A. J. Kreuz, A. Simcox, and D. Maughan. Alterations in flight muscle ultrastructure and function in Drosophila tropomyosin mutants. Journal of Cell Biology, 135(3):673–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
3/673.

Kim:1997:TLO

[KSOK97] John Kim, Sidney V. Scott, Michael N. Oda, and Daniel J. Klionsky. Transport of a Large Oligomeric Protein by the Cytoplasm to Vacuole Protein Targeting Pathway. Journal of Cell Biology, 137(3):609–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/3/609.

Kandror:1995:ECC

[KSP95] K. V. Kandror, J. M. Stephens, and P. F. Pilch. Expression and compartmentalization of caveolin in adipose cells: coordinate regulation with and structural segregation from GLUT4. Journal of Cell Biology, 129(4):999–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/4/999.

Kypta:1996:ABT

[KSR96] R. M. Kypta, H. Su, and L. F. Reichardt. Association between a transmembrane protein tyrosine phosphatase and the cadherin-catenin complex. Journal of Cell Biology, 134(6): 1519–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1519.

Kang:1995:ITM

[KSS^{+}95] S. J. Kang, K. S. Shin, W. K. Song, D. B. Ha, C. H. Chung, and M. S. Kang. Involvement of transglutaminase in myofibril assembly of chick embryonic myoblasts in culture. Journal of Cell Biology, 130(5):1127–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/5/1127.

Kann:1999:PDB

[KSV^{+}99] Michael Kann, Beate Sodeik, Angelika Vlachou, Wolfram H. Gerlich, and Ari Helenius. Phosphorylation-dependent Binding of Hepatitis B Virus Core Particles to the Nuclear Pore Complex. Journal of Cell Biology, 145(1):45–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/145/1/45.

Kranenburg:1995:ICD

[KSVZ95] O. Kranenburg, V. Scharnhorst, A. J. Van der Eb, and A. Zantema. Inhibition of cyclin-dependent kinase activity triggers neuronal differentiation of mouse neuroblastoma cells. Journal of Cell Biology, 131(1):227–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/1/227.

Kitayama:1997:TIM

[KSY97] Chikako Kitayama, Asako Sugimoto, and Masayuki Yamamoto. Type II Myosin Heavy Chain Encoded by the myo2 Gene Composes the Contractile Ring during Cytokinesis in Schizosaccharomyces pombe. Journal of Cell Biology, 137(6): 1309–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/6/1309.

Kutzleb:1998:PPP

[KSY^{+}98] Christian Kutzleb, Gabriele Sanders, Raina Yamamoto, Xiaolu Wang, Beate Lichte, Elisabeth Petrasch-Parwez, and Manfred W. Kilimann. Paralemmin, a Prenyl–Palmitoyl– anchored Phosphoprotein Abundant in Neurons and Implicated in Plasma Membrane Dynamics and Cell Process Formation. Journal of Cell Biology, 143(3):795–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/795.

Kamal:1998:AVM

[KsYA98] Adeela Kamal, Yun shu Ying, and Richard G. W. Anderson. Annexin VI-mediated Loss of Spectrin during Coated Pit Budding Is Coupled to Delivery of LDL to Lysosomes. Journal of Cell Biology, 142(4):937–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/4/937.

Kubo:1999:CS

[KSYK^{+}99] Akiharu Kubo, Hiroyuki Sasaki, Akiko Yuba-Kubo, Shoichiro Tsukita, and Nobuyuki Shiina. Centriolar Satellites. Journal of Cell Biology, 147(5):969–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/5/969.

Kong:1998:ARN

[KTAX98] Jiming Kong, Vivian W.-Y. Tung, John Aghajanian, and Zuoshang Xu. Antagonistic Roles of Neurofilament Subunits NF– H and NF–M Against NF–L in Shaping Dendritic Arborization in Spinal Motor Neurons. Journal of Cell Biology, 140 (5):1167–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/5/1167.

Karpova:1995:AFY

[KTC95] T. S. Karpova, K. Tatchell, and J. A. Cooper. Actin filaments in yeast are unstable in the absence of capping protein or fimbrin. Journal of Cell Biology, 131(6):1483–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1483.

Kondo:1997:EER

[KTD^{+}97] Takahisa Kondo, Kosei Takeuchi, Yoshinori Doi, Shigenobu Yonemura, Shigekazu Nagata, Shoichiro Tsukita, and Sachiko Tsukita. ERM (Ezrin/Radixin/Moesin)-based molecular mechanism of microvillar breakdown at an early stage of apoptosis. Journal of Cell Biology, 139(3):749–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/749.

Kotani:1999:RAA

[KTYT99] Shuji Kotani, Hirofumi Tanaka, Hideyo Yasuda, and Kazuo Todokoro. Regulation of APC Activity by Phosphorylation and Regulatory Factors. Journal of Cell Biology, 146(4): 791–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/791.

Kunimoto:1995:NSI

[Kun95] M. Kunimoto. A neuron-specific isoform of brain ankyrin, 440-kD ankyrinB, is targeted to the axons of rat cerebellar neurons. Journal of Cell Biology, 131(6):1821–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1821.

Koolwijk:1996:CET

[KvEdV^{+}96] P. Koolwijk, M. G. van Erck, W. J. de Vree, M. A. Vermeer, H. A. Weich, R. Hanemaaijer, and V. W. van Hinsbergh. Cooperative effect of TNFalpha, bFGF, and VEGF on the formation of tubular structures of human microvascular endothelial cells in a fibrin matrix. Role of urokinase activity. Journal of Cell Biology, 132(6):1177–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/6/1177.

Koepp:1996:DLN

[KWCS96] D. M. Koepp, D. H. Wong, A. H. Corbett, and P. A. Silver. Dynamic localization of the nuclear import receptor and its interactions with transport factors. Journal of Cell Biology, 133(6):1163–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/6/1163.

Kallio:1998:MPM

[KWD^{+}98] Marko Kallio, Jasminder Weinstein, John R. Daum, Daniel J. Burke, and Gary J. Gorbsky. Mammalian p55CDC mediates association of the spindle checkpoint protein Mad2 with the cyclosome/anaphase-promoting complex, and is involved in regulating anaphase onset and late mitotic events. Journal of Cell Biology, 141(6):1393–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/6/1393.

Kinose:1996:GPM

[KWK^{+}96] F. Kinose, S. X. Wang, U. S. Kidambi, C. L. Moncman, and D. A. Winkelmann. Glycine 699 is pivotal for the motor activity of skeletal muscle myosin. Journal of Cell Biology, 134 (4):895–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/895.

Kircher:1999:NIP

[KWN^{+}99] Stefan Kircher, Frank Wellmer, Peter Nick, Alexander Rügner, Eberhard Schäfer, and Klaus Harter. Nuclear Import of the Parsley bZIP Transcription Factor CPRF2 Is Regulated by Phytochrome Photoreceptors. Journal of Cell Biology, 144(2): 201–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/2/201.

Konzok:1999:DDH

[KWS^{+}99] Angelika Konzok, Igor Weber, Evelyn Simmeth, Ulrike Hacker, Markus Maniak, and Annette Müller-Taubenberger. Daip1, a Dictyostelium Homologue of the Yeast Actin–Interacting Protein 1, Is Involved in Endocytosis, Cytokinesis, and Motility. Journal of Cell Biology, 146(2):453–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/2/453.

Kunz:1996:ISC

[KZKS96] S. Kunz, U. Ziegler, B. Kunz, and P. Sonderegger. Intracellular signaling is changed after clustering of the neural cell adhesion molecules axonin-1 and NgCAM during neurite fasciculation. Journal of Cell Biology, 135(1):253–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
1/253.

Kyriakides:1998:MLT

[KZS^{+}98] Themis R. Kyriakides, Yu-Hong Zhu, Lynne T. Smith, Steven D. Bain, Zhantao Yang, Ming T. Lin, Keith G. Danielson, Renato V. Iozzo, Mary LaMarca, Cindy E. McKinney, Edward I. Ginns, and Paul Bornstein. Mice That Lack Thrombospondin 2 Display Connective Tissue Abnormalities That Are Associated with Disordered Collagen Fibrillogenesis, an Increased Vascular Density, and a Bleeding Diathesis. Journal of Cell Biology, 140(2):419–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/2/419.

Lansdorp:1997:LMT

[Lan97] Peter M. Lansdorp. Lessons from Mice without Telomerase. Journal of Cell Biology, 139(2):309–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/2/309.

LeMaire-Adkins:1997:LCC

[LARH97] Renée LeMaire-Adkins, Kristi Radke, and Patricia A. Hunt. Lack of checkpoint control at the metaphase/anaphase transition: a mechanism of meiotic nondisjunction in mammalian females. Journal of Cell Biology, 139(7):1611–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
7/1611.

Lin:1997:IMA

[LAS^{+}97] Tsung H. Lin, Andrew E. Aplin, Yu Shen, Qiming Chen, Michael Schaller, Lewis Romer, Ikramuddin Aukhil, and R. L. Juliano. Integrin-mediated Activation of MAP Kinase Is Independent of FAK: Evidence for Dual Integrin Signaling Pathways in Fibroblasts. Journal of Cell Biology, 136(6):1385–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/6/1385.

LeBot:1998:RXS

[LAW^{+}98] Nathalie Le Bot, Claude Antony, Jamie White, Eric Karsenti, and Isabelle Vernos. Role of Xklp3, a subunit of the xenopus kinesin II heterotrimeric complex, in membrane transport between the endoplasmic reticulum and the Golgi apparatus. Journal of Cell Biology, 143(6):1559–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1559.

Lillie:1998:SKR

[LB98] S. H. Lillie and S. S. Brown. Smy1p, a Kinesin-related Protein That Does Not Require Microtubules. Journal of Cell Biology, 140(4):873–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/4/873.

Littleton:1997:DFN

[LBB97] J. Troy Littleton, Manzoor A. Bhat, and Hugo J. Bellen. Deciphering the Function of Neurexins at Cellular Junctions. Journal of Cell Biology, 137(4):793–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/4/793.

Lenormand:1998:GFI

[LBBP98] Philippe Lenormand, Jean-Marc Brondello, Anne Brunet, and Jacques Pouysségur. Growth Factor–induced p42/p44 MAPK Nuclear Translocation and Retention Requires Both MAPK Activation and Neosynthesis of Nuclear Anchoring Proteins. Journal of Cell Biology, 142(3):625–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/3/625.

Lue:1996:TID

[LBCB96] R. A. Lue, E. Brandin, E. P. Chan, and D. Branton. Two independent domains of hDlg are sufficient for subcellular targeting: the PDZ1-2 conformational unit and an alternatively spliced domain. Journal of Cell Biology, 135(4):1125– ??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/1125.

Lollike:1995:EFP

[LBL95] K. Lollike, N. Borregaard, and M. Lindau. The exocytotic fusion pore of small granules has a conductance similar to an ion channel. Journal of Cell Biology, 129(1):99–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/1/99.

Liu:1998:FNI

[LBO^{+}98] Feizhou Liu, Christopher C. Bauer, Irving Ortiz, Richard G. Cook, Michael F. Schmid, and Henry F. Epstein. β-filagenin, a Newly Identified Protein Coassembling with Myosin and Paramyosin in Caenorhabditis elegans. Journal of Cell Biology, 140(2):347–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/2/347.

Litvinov:1997:ECA

[LBW^{+}97] Sergey V. Litvinov, Maarten Balzar, Manon J. Winter, Hellen A. M. Bakker, Inge H. Briaire de Bruijn, Frans Prins, Gert Jan Fleuren, and Sven O. Warnaar. Epithelial Cell Adhesion Molecule (Ep–CAM) Modulates Cell–Cell Interactions Mediated by Classic Cadherins. Journal of Cell Biology, 139(5): 1337–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1337.

Lenz-Bohme:1997:IMD

[LBWF^{+}97] Bettina Lenz-Böhme, Jasmine Wismar, Silke Fuchs, Rita Reifegerste, Erich Buchner, Heinrich Betz, and Bertram Schmitt. Insertional Mutation of the Drosophila Nuclear Lamin Dm_{0} Gene Results in Defective Nuclear Envelopes, Clustering of Nuclear Pore Complexes, and Accumulation of Annulate Lamellae. Journal of Cell Biology, 137(5):1001–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/5/1001.

Loike:1999:DRI

[LCB^{+}99] John D. Loike, Long Cao, Sadna Budhu, Eugene E. Marcantonio, Joseph El Khoury, Stanley Hoffman, Ted A. Yednock, and Samuel C. Silverstein. Differential Regulation of β_{1} Integrins by Chemoattractants Regulates Neutrophil Migration through Fibrin. Journal of Cell Biology, 144(5):1047–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
144/5/1047.

Lampugnani:1995:MOE

[LCC^{+}95] M. G. Lampugnani, M. Corada, L. Caveda, F. Breviario, O. Ayalon, B. Geiger, and E. Dejana. The molecular organization of endothelial cell to cell junctions: differential association of plakoglobin, beta-catenin, and alpha-catenin with vascular endothelial cadherin (VE-cadherin). Journal of Cell Biology, 129(1):203–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/1/203.

Landolfi:1998:CHA

[LCD^{+}98] Barbara Landolfi, Silvana Curci, Lucantonio Debellis, Tullio Pozzan, and Aldebaran M. Hofer. Ca^{2+} homeostasis in the agonist-sensitive internal store: Functional interactions between mitochondria and the ER measured in situ in intact cells. Journal of Cell Biology, 142(5):1235–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
5/1235.

Li:1996:PMH

[LCH96] F. Q. Li, A. Coonrod, and M. Horwitz. Preferential MyoD homodimer formation demonstrated by a general method of dominant negative mutation employing fusion with a lysosomal protease. Journal of Cell Biology, 135(4):1043–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/4/1043.

Laird:1995:GJT

[LCK95] D. W. Laird, M. Castillo, and L. Kasprzak. Gap junction turnover, intracellular trafficking, and phosphorylation of connexin43 in brefeldin A-treated rat mammary tumor cells. Journal of Cell Biology, 131(5):1193–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/5/1193.

Lencer:1995:TCT

[LCM^{+}95] W. I. Lencer, C. Constable, S. Moe, M. G. Jobling, H. M. Webb, S. Ruston, J. L. Madara, T. R. Hirst, and R. K. Holmes. Targeting of cholera toxin and Escherichia coli heat labile toxin in polarized epithelia: role of COOH-terminal KDEL. Journal of Cell Biology, 131(4):951–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/951.

Litwin:1997:NCI

[LCN^{+}97] Marek Litwin, Katherine Clark, Leanne Noack, Jill Furze, Michael Berndt, Steven Albelda, Mathew Vadas, and Jennifer Gamble. Novel cytokine-independent induction of endothelial adhesion molecules regulated by platelet/endothelial cell adhesion molecule (CD31). Journal of Cell Biology, 139(1): 219–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/219.

Labriola:1995:RGU

[LCP95] C. Labriola, J. J. Cazzulo, and A. J. Parodi. Retention of glucose units added by the UDP–GLC: glycoprotein glucosyltransferase delays exit of glycoproteins from the endoplasmic reticulum. Journal of Cell Biology, 130(4):771–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
4/771.

Lewis:1996:PKC

[LCS96] J. M. Lewis, D. A. Cheresh, and M. A. Schwartz. Protein kinase C regulates alpha v beta 5-dependent cytoskeletal associations and focal adhesion kinase phosphorylation. Journal of Cell Biology, 134(5):1323–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/5/1323.

Low:1998:SMI

[LCW^{+}98] Seng Hui Low, Steven J. Chapin, Christian Wimmer, Sidney W. Whiteheart, László G. Kömüves, Keith E. Mostov, and Thomas Weimbs. The SNARE Machinery Is Involved in Apical Plasma Membrane Trafficking in MDCK Cells. Journal of Cell Biology, 141(7):1503–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/7/1503.

Lilly:1995:CGG

[LD95] P. J. Lilly and P. N. Devreotes. Chemoattractant and GTP gamma S-mediated stimulation of adenylyl cyclase in Dictyostelium requires translocation of CRAC to membranes. Journal of Cell Biology, 129(6):1659–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/6/1659.

Lin:1995:EMR

[LDCB95] C. Q. Lin, P. J. Dempsey, R. J. Coffey, and M. J. Bissell. Extracellular matrix regulates whey acidic protein gene expression by suppression of TGF-alpha in mouse mammary epithelial cells: studies in culture and in transgenic mice. Journal of Cell Biology, 129(4):1115–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/4/1115.

Lorentz:1997:KRC

[LDD^{+}97] Olivier Lorentz, Isabelle Duluc, Adèle De Arcangelis, Patricia Simon-Assmann, Michèle Kedinger, and Jean-Noël Freund. Key Role of the Cdx2 Homeobox Gene in Extracellular Matrix–mediated Intestinal Cell Differentiation. Journal of Cell Biology, 139(6):1553–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/6/1553.

LaRochelle:1995:SRD

[LDF^{+}95a] W. J. LaRochelle, O. R. Dirsch, P. W. Finch, H. G. Cheon, M. May, C. Marchese, J. H. Pierce, and S. A. Aaronson. Specific receptor detection by a functional keratinocyte growth factor-immunoglobulin chimera. Journal of Cell Biology, 129 (2):357–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/357.

Lorenzen:1995:CTS

[LDF95b] J. A. Lorenzen, C. Y. Dadabay, and E. H. Fischer. COOHterminal sequence motifs target the T cell protein tyrosine phosphatase to the ER and nucleus. Journal of Cell Biology, 131(3):631–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/3/631.

Lin:1998:SSS

[LDH^{+}98] Kuo-I Lin, Joseph A. DiDonato, Alexander Hoffmann, J. Marie Hardwick, and Rajiv R. Ratan. Suppression of Steady-state, but not Stimulus-induced NF-κB Activity Inhibits Alphavirus-induced Apoptosis. Journal of Cell Biology, 141(7):1479–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/7/1479.

Leppert:1999:NID

[LDP^{+}99] Christian A. Leppert, Heike Diekmann, Claudia Paul, Ute Laessing, Monika Marx, Martin Bastmeyer, and Claudia A. O. Stuermer. Neurolin Ig Domain 2 Participates in Retinal Axon Guidance and Ig Domains 1 and 3 in Fasciculation. Journal of Cell Biology, 144(2):339–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/2/339.

Liu:1997:DLV

[LE97] Yongjian Liu and Robert H. Edwards. Differential Localization of Vesicular Acetylcholine and Monoamine Transporters in PC12 Cells but Not CHO Cells. Journal of Cell Biology, 139(4):907–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/4/907.

Longo:1997:HBR

[LEB^{+}97] Valter D. Longo, Lisa M. Ellerby, Dale E. Bredesen, Joan S. Valentine, and Edith B. Gralla. Human Bcl-2 Reverses Survival Defects in Yeast Lacking Superoxide Dismutase and Delays Death of Wild–Type Yeast. Journal of Cell Biology, 137 (7):1581–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/7/1581.

Loeb:1995:ACR

[LF95] J. A. Loeb and G. D. Fischbach. ARIA can be released from extracellular matrix through cleavage of a heparinbinding domain. Journal of Cell Biology, 130(1):127–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
1/127.

Lowell:1996:DSF

[LFB96] C. A. Lowell, L. Fumagalli, and G. Berton. Deficiency of Src family kinases p59/61hck and p58c-fgr results in defective adhesion-dependent neutrophil functions. Journal of Cell Biology, 133(4):895–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/4/895.

Longtine:1998:RYG

[LFP98] Mark S. Longtine, Hanna Fares, and John R. Pringle. Role of the Yeast Gin4p Protein Kinase in Septin Assembly and the Relationship between Septin Assembly and Septin Function. Journal of Cell Biology, 143(3):719–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/719.

Lange:1995:MMC

[LG95] B. M. Lange and K. Gull. A molecular marker for centriole maturation in the mammalian cell cycle. Journal of Cell Biology, 130(4):919–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/4/919.

Locker:1999:URC

[LG99] Jacomine Krijnse Locker and Gareth Griffiths. An Unconventional Role for Cytoplasmic Disulfide Bonds in Vaccinia Virus Proteins. Journal of Cell Biology, 144(2):267–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
144/2/267.

Lefebvre:1995:TXC

[LGdC95] V. Lefebvre, S. Garofalo, and B. de Crombrugghe. Type X collagen gene expression in mouse chondrocytes immortalized by a temperature-sensitive simian virus 40 large tumor antigen. Journal of Cell Biology, 128(1):239–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
1/239.

Liou:1997:AEP

[LGGS97] Willisa Liou, Hans J. Geuze, Math. J. H. Geelen, and Jan W. Slot. The Autophagic and Endocytic Pathways Converge at the Nascent Autophagic Vacuoles. Journal of Cell Biology, 136 (1):61–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/61.

Lochter:1997:MMS

[LGM^{+}97] André Lochter, Sybille Galosy, John Muschler, Neal Freedman, Zena Werb, and Mina J. Bissell. Matrix Metalloproteinase Stromelysin-1 Triggers a Cascade of Molecular Alterations That Leads to Stable Epithelial–to–Mesenchymal Conversion and a Premalignant Phenotype in Mammary Epithelial Cells. Journal of Cell Biology, 139(7):1861–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
7/1861.

Lemaitre:1998:DGD

[LGM98] Jean-Marc Lemaitre, Gérard Géraud, and Marcel Méchali. Dynamics of the Genome during Early Xenopus laevis Development: Karyomeres As Independent Units of Replication. Journal of Cell Biology, 142(5):1159–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/5/1159.

Lindberg:1996:IAP

[LGRB96] F. P. Lindberg, H. D. Gresham, M. I. Reinhold, and E. J. Brown. Integrin-associated protein immunoglobulin domain is necessary for efficient vitronectin bead binding. Journal of Cell Biology, 134(5):1313–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/5/1313.

Legrand:1999:AEC

[LGZ^{+}99] Claire Legrand, Christine Gilles, Jean-Marie Zahm, Myriam Polette, Anne-Cécile Buisson, Hervé Kaplan, Philippe Birembaut, and Jean-Marie Tournier. Airway Epithelial Cell Migration Dynamics: MMP-9 Role in Cell–Extracellular Matrix Remodeling. Journal of Cell Biology, 146(2):517–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
2/517.

LeBorgne:1997:MPR

[LH97] Roland Le Borgne and Bernard Hoflack. Mannose 6-Phosphate Receptors Regulate the Formation of Clathrin-coated Vesicles in the TGN. Journal of Cell Biology, 137(2):335–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
2/335.

Ledig:1999:RTP

[LHB^{+}99] Matthias M. Ledig, Fawaz Haj, John L. Bixby, Andrew W. Stoker, and Bernhard K. Mueller. The Receptor Tyrosine Phosphatase Crypα Promotes Intraretinal Axon Growth. Journal of Cell Biology, 147(2):375–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/2/375.

Letourneur:1995:SMD

[LHDC95] F. Letourneur, S. Hennecke, C. Démollière, and P. Cosson. Steric masking of a dilysine endoplasmic reticulum retention motif during assembly of the human high affinity receptor for immunoglobulin E. Journal of Cell Biology, 129(4):971–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/4/971.

Lai:1996:STT

[LHK^{+}96] C. C. Lai, K. Hong, M. Kinnell, M. Chalfie, and M. Driscoll. Sequence and transmembrane topology of MEC-4, an ion channel subunit required for mechanotransduction in Caenorhabditis elegans. Journal of Cell Biology, 133(5):1071–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
5/1071.

Lupetti:1996:OSS

[LHM^{+}96] P. Lupetti, J. E. Heuser, R. Manetti, P. Massari, S. Lanzavecchia, P. L. Bellon, R. Dallai, R. Rappuoli, and J. L. Telford. Oligomeric and subunit structure of the Helicobacter pylori vacuolating cytotoxin. Journal of Cell Biology, 133(4):801–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/4/801.

Liang:1997:GBN

[LHR^{+}97] Yu Liang, Monika Häring, Peter J. Roughley, Renée K. Margolis, and Richard U. Margolis. Glypican and Biglycan in the Nuclei of Neurons and Glioma Cells: Presence of Functional Nuclear Localization Signals and Dynamic Changes in Glypican During the Cell Cycle. Journal of Cell Biology, 139(4): 851–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/851.

Lupashin:1996:BRT

[LHS96] V. V. Lupashin, S. Hamamoto, and R. W. Schekman. Biochemical requirements for the targeting and fusion of ERderived transport vesicles with purified yeast Golgi membranes. Journal of Cell Biology, 132(3):277–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
3/277.

Liu:1997:FAA

[LHS97a] Jianwei Liu, Thomas E. Hughes, and William C. Sessa. The First 35 Amino Acids and Fatty Acylation Sites Determine the Molecular Targeting of Endothelial Nitric Oxide Synthase into the Golgi Region of Cells: a Green Fluorescent Protein Study. Journal of Cell Biology, 137(7):1525–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
7/1525.

Lubeck:1997:NCC

[LHS97b] Jens Lübeck, Lisa Heins, and Jürgen Soll. A Nuclear-coded Chloroplastic Inner Envelope Membrane Protein Uses a Soluble Sorting Intermediate upon Import into the Organelle. Journal of Cell Biology, 137(6):1279–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/6/1279.

Lingappa:1997:MAD

[LHWH97] Jaisri R. Lingappa, Rebecca L. Hill, Mei Lie Wong, and Ramanujan S. Hegde. A Multistep, ATP-dependent Pathway for Assembly of Human Immunodeficiency Virus Capsids in a Cell-free System. Journal of Cell Biology, 136(3): 567–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/567.

Li:1997:BYP

[Li97] Rong Li. Bee1, a Yeast Protein with Homology to Wiscott– Aldrich Syndrome Protein, Is Critical for the Assembly of Cortical Actin Cytoskeleton. Journal of Cell Biology, 136(3): 649–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/649.

Li:1995:GTM

[LJ95] Q. Li and H. C. Joshi. gamma-tubulin is a minus end-specific microtubule binding protein. Journal of Cell Biology, 131(1): 207–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/207.

Lohret:1997:TPI

[LJK97] Timothy A. Lohret, Robert E. Jensen, and Kathleen W. Kinnally. Tim23, a Protein Import Component of the Mitochondrial Inner Membrane, Is Required for Normal Activity of the Multiple Conductance Channel, MCC. Journal of Cell Biology, 137(2):377–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/2/377.

Loukin:1995:MES

[LK95] S. Loukin and C. Kung. Manganese effectively supports yeast cell-cycle progression in place of calcium. Journal of Cell Biology, 131(4):1025–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/1025.

Liu:1996:IDS

[LK96] Y. Liu and R. N. Kitsis. Induction of DNA synthesis and apoptosis in cardiac myocytes by E1A oncoprotein. Journal of Cell Biology, 133(2):325–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/2/325.

Li:1995:LEC

[LKD^{+}95] W. C. Li, J. R. Kuszak, K. Dunn, R. R. Wang, W. Ma, G. M. Wang, A. Spector, M. Leib, A. M. Cotliar, and M. Weiss. Lens epithelial cell apoptosis appears to be a common cellular basis for non-congenital cataract development in humans and animals. Journal of Cell Biology, 130(1):169–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
1/169.

Lee:1999:CMS

[LKE^{+}99] Laifong Lee, Saskia K. Klee, Marie Evangelista, Charles Boone, and David Pellman. Control of Mitotic Spindle Position by the Saccharomyces cerevisiae Formin Bni1p. Journal of Cell Biology, 144(5):947–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/5/947.

Lawrence:1997:TLF

[LKKL97] Michael B. Lawrence, Geoffrey S. Kansas, Eric J. Kunkel, and Klaus Ley. Threshold Levels of Fluid Shear Promote Leukocyte Adhesion through Selectins (CD62L,P,E). Journal of Cell Biology, 136(3):717–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/136/3/717.

Lowin-Kropf:1998:CPC

[LKSW98] Bente Lowin-Kropf, Virginia Smith Shapiro, and Arthur Weiss. Cytoskeletal polarization of T cells is regulated by an immunoreceptor tyrosine-based activation motif-dependent mechanism. Journal of Cell Biology, 140(4):861–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
4/861.

Lawlis:1996:CAC

[LKWG96] S. J. Lawlis, S. M. Keezer, J. R. Wu, and D. M. Gilbert. Chromosome architecture can dictate site-specific initiation of DNA replication in Xenopus egg extracts. Journal of Cell Biology, 135(5):1207–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/5/1207.

Lechler:1997:VRC

[LL97] Terry Lechler and Rong Li. In Vitro Reconstitution of Cortical Actin Assembly Sites in Budding Yeast. Journal of Cell Biology, 138(1):95–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/1/95.

Lippincott:1998:DFC

[LL98a] John Lippincott and Rong Li. Dual Function of Cyk2, a cdc15/ PSTPIP Family Protein, in Regulating Actomyosin Ring Dynamics and Septin Distribution. Journal of Cell Biology, 143(7):1947–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/7/1947.

Lippincott:1998:SAM

[LL98b] John Lippincott and Rong Li. Sequential Assembly of Myosin II, an IQGAP-like Protein, and Filamentous Actin to a Ring Structure Involved in Budding Yeast Cytokinesis. Journal of Cell Biology, 140(2):355–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/2/355.

Lin:1999:GIA

[LLG^{+}99] Chung-Chih Lin, Harold D. Love, Jennifer N. Gushue, John J. M. Bergeron, and Joachim Ostermann. ER/ Golgi Intermediates Acquire Golgi Enzymes by Brefeldin a–Sensitive Retrograde Transport in Vitro. Journal of Cell Biology, 147(7): 1457–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/7/1457.

Laurent:1999:RPE

[LLH^{+}99] Valérie Laurent, Thomas P. Loisel, Birgit Harbeck, Ann Wehman, Lothar Gröbe, Brigitte M. Jockusch, Jürgen Wehland, Frank B. Gertler, and Marie-France Carlier. Role of proteins of the Ena/VASP family in actin-based motility of Listeria monocytogenes. Journal of Cell Biology, 144(6): 1245–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/6/1245.

Liao:1995:DHK

[LLK^{+}95] J. Liao, L. A. Lowthert, N. O. Ku, R. Fernandez, and M. B. Omary. Dynamics of human keratin 18 phosphorylation: polarized distribution of phosphorylated keratins in simple epithelial tissues. Journal of Cell Biology, 131(5):1291–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/5/1291.

Li:1996:PRG

[LLL^{+}96] H. Li, T. F. Liu, A. Lazrak, C. Peracchia, G. S. Goldberg, P. D. Lampe, and R. G. Johnson. Properties and regulation of gap junctional hemichannels in the plasma membranes of cultured cells. Journal of Cell Biology, 134(4):1019–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
4/1019.

Lin:1995:TAB

[LLN^{+}95] K. I. Lin, S. H. Lee, R. Narayanan, J. M. Baraban, J. M. Hardwick, and R. R. Ratan. Thiol agents and Bcl-2 identify an alphavirus-induced apoptotic pathway that requires activation of the transcription factor NF-kappa B. Journal of Cell Biology, 131(5):1149–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/5/1149.

Lin:1998:MCB

[LLNH^{+}98] Ping Lin, Helen Le-Niculescu, Robert Hofmeister, J. Michael McCaffery, Mingjie Jin, Hanjo Hennemann, Tammie McQuistan, Luc De Vries, and Marilyn Gist Farquhar. The Mammalian Calcium-binding Protein, Nucleobindin (CALNUC), Is a Golgi Resident Protein. Journal of Cell Biology, 141 (7):1515–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/7/1515.

Lu:1995:BOI

[LLRTP95] P. J. Lu, Q. L. Lu, A. Rughetti, and J. Taylor-Papadimitriou. bcl-2 overexpression inhibits cell death and promotes the morphogenesis, but not tumorigenesis of human mammary epithelial cells. Journal of Cell Biology, 129(5):1363–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1363.

Love:1998:IFG

[LLSO98] Harold D. Love, Chung-Chih Lin, Craig S. Short, and Joachim Ostermann. Isolation of Functional Golgi-derived Vesicles with a Possible Role in Retrograde Transport. Journal of Cell Biology, 140(3):541–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/541.

Lafont:1998:AXA

[LLVS98] Frank Lafont, Sandra Lecat, Paul Verkade, and Kai Simons. Annexin XIIIb Associates with Lipid Microdomains to Function in Apical Delivery. Journal of Cell Biology, 142(6):1413– ??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/6/1413.

Lastres:1996:EMC

[LLZ^{+}96] P. Lastres, A. Letamendía, H. Zhang, C. Rius, N. Almendro, U. Raab, L. A. López, C. Langa, A. Fabra, M. Letarte, and C. Bernabéu. Endoglin modulates cellular responses to TGF-beta 1. Journal of Cell Biology, 133(5):1109–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
5/1109.

Lantz:1998:CVU

[LM98] Valerie A. Lantz and Kathryn G. Miller. A Class VI Unconventional Myosin Is Associated with a Homologue of a Microtubule-binding Protein, Cytoplasmic Linker Protein– 170, in Neurons and at the Posterior Pole of Drosophila Embryos. Journal of Cell Biology, 140(4):897–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
4/897.

Li:1997:DET

[LMA^{+}97] Zhenlin Li, Mathias Mericskay, Onnik Agbulut, Gillian Butler-Browne, Lena Carlsson, Lars-Eric Thornell, Charles Babinet, and Denise Paulin. Desmin Is Essential for the Tensile Strength and Integrity of Myofibrils but Not for Myogenic Commitment, Differentiation, and Fusion of Skeletal Muscle. Journal of Cell Biology, 139(1):129–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/1/129.

Lu:1995:GPD

[LMB^{+}95] C. F. Lu, R. C. Montijn, J. L. Brown, F. Klis, J. Kurjan, H. Bussey, and P. N. Lipke. Glycosyl phosphatidylinositoldependent cross-linking of alpha-agglutinin and beta 1,6glucan in the Saccharomyces cerevisiae cell wall. Journal of Cell Biology, 128(3):333–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/3/333.

Liu:1998:DNC

[LMB98] Shu-Hui Liu, Michael S. Marks, and Frances M. Brodsky. A Dominant-negative Clathrin Mutant Differentially Affects Trafficking of Molecules with Distinct Sorting Motifs in the Class II Major Histocompatibility Complex (MHC) Pathway. Journal of Cell Biology, 140(5):1023–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/5/1023.

Lipardi:1998:CTR

[LMC^{+}98] Concetta Lipardi, Rosalia Mora, Veronica Colomer, Simona Paladino, Lucio Nitsch, Enrique Rodriguez-Boulan, and Chiara Zurzolo. Caveolin Transfection Results in Caveolae Formation but Not Apical Sorting of Glycosylphosphatidylinositol (GPI)-anchored Proteins in Epithelial Cells. Journal of Cell Biology, 140(3):617–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/3/617.

LaJeunesse:1998:SAD

[LMF98] Dennis R. LaJeunesse, Brooke M. McCartney, and Richard G. Fehon. Structural Analysis of Drosophila Merlin Reveals Functional Domains Important for Growth Control and Subcellular Localization. Journal of Cell Biology, 141(7):1589–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
7/1589.

Ladinsky:1999:GST

[LMM^{+}99] Mark S. Ladinsky, David N. Mastronarde, J. Richard McIntosh, Kathryn E. Howell, and L. Andrew Staehelin. Golgi Structure in Three Dimensions: Functional Insights from the Normal Rat Kidney Cell. Journal of Cell Biology, 144(6): 1135–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/6/1135.

Li:1996:ESP

[LMMO96] L. Li, J. M. Miano, B. Mercer, and E. N. Olson. Expression of the SM22alpha promoter in transgenic mice provides evidence for distinct transcriptional regulatory programs in vascular and visceral smooth muscle cells. Journal of Cell Biology, 132(5):849–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/5/849.

Lindon:1998:CCR

[LMP98] Catherine Lindon, Didier Montarras, and Christian Pinset. Cell Cycle–regulated Expression of the Muscle Determination Factor Myf5 in Proliferating Myoblasts. Journal of Cell Biology, 140(1):111–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/1/111.

Li:1997:RCN

[LMS97] Yonghong Li, Pamela Maher, and David Schubert. Requirement for cGMP in Nerve Cell Death Caused by Glutathione Depletion. Journal of Cell Biology, 139(5):1317–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
5/1317.

Li:1999:NMS

[LMW99] Ji Li, Richard Mayne, and Chuanyue Wu. A Novel Muscle– Specific β1 Integrin Binding Protein (Mibp) That Modulates Myogenic Differentiation. Journal of Cell Biology, 147(7): 1391–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/7/1391.

Lane:1996:AMR

[LN96] H. A. Lane and E. A. Nigg. Antibody microinjection reveals an essential role for human polo-like kinase 1 (Plk1) in the functional maturation of mitotic centrosomes. Journal of Cell Biology, 135(6):1701–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1701.

Lampe:1998:CII

[LNG^{+}98] Paul D. Lampe, Beth P. Nguyen, Susana Gil, Marcia Usui, John Olerud, Yoshikazu Takada, and William G. Carter. Cellular Interaction of Integrin α3β1 with Laminin 5 Promotes Gap Junctional Communication. Journal of Cell Biology, 143(6):1735–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1735.

Leung:1998:DCR

[LNH^{+}98] Keith K. H. Leung, Ling Jim Ng, Ken K. Y. Ho, Patrick P. L. Tam, and Kathryn S. E. Cheah. Different cis -Regulatory DNA Elements Mediate Developmental Stage- and Tissuespecific Expression of the Human COL2A1 Gene in Transgenic Mice. Journal of Cell Biology, 141(6):1291–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
6/1291.

Lavoie:1998:ENA

[LNM^{+}98] Josée N. Lavoie, M. Nguyen, R. C. Marcellus, P. E. Branton, and G. C. Shore. E4orf4, a Novel Adenovirus Death Factor That Induces p53-independent Apoptosis by a Pathway That Is Not Inhibited by zVAD-fmk. Journal of Cell Biology, 140 (3):637–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/3/637.

Lazarovits:1996:ECI

[LNR^{+}96] J. Lazarovits, H. Y. Naim, A. C. Rodriguez, R. H. Wang, E. Fire, C. Bird, Y. I. Henis, and M. G. Roth. Endocytosis of chimeric influenza virus hemagglutinin proteins that lack a cytoplasmic recognition feature for coated pits. Journal of Cell Biology, 134(2):339–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/2/339.

Lin:1998:MMT

[LNRR98] Sasa Lin, Hussein Y. Naim, A. Chapin Rodriguez, and Michael G. Roth. Mutations in the Middle of the Transmembrane Domain Reverse the Polarity of Transport of the Influenza Virus Hemagglutinin in MDCK Epithelial Cells. Journal of Cell Biology, 142(1):51–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/1/51.

Lombillo:1995:AKM

[LNY^{+}95] V. A. Lombillo, C. Nislow, T. J. Yen, V. I. Gelfand, and J. R. McIntosh. Antibodies to the kinesin motor domain and CENP–E inhibit microtubule depolymerizationdependent motion of chromosomes in vitro. Journal of Cell Biology, 128(1):107–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/1/107.

Liao:1996:PAP

[LO96] J. Liao and M. B. Omary. 14-3-3 proteins associate with phosphorylated simple epithelial keratins during cell cycle progression and act as a solubility cofactor. Journal of Cell Biology, 133(2):345–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/2/345.

Leiper:1996:IAG

[LOD96] J. M. Leiper, P. B. Oatey, and C. J. Danpure. Inhibition of alanine: glyoxylate aminotransferase 1 dimerization is a prerequisite for its peroxisome-to-mitochondrion mistargeting in primary hyperoxaluria type 1. Journal of Cell Biology, 135(4): 939–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/939.

Lavoie:1999:RPC

[LPD^{+}99] C. Lavoie, J. Paiement, M. Dominguez, L. Roy, S. Dahan, J. N. Gushue, and J. J. M. Bergeron. Roles for α_{2} p24 and COPI in Endoplasmic Reticulum Cargo Exit Site Formation. Journal of Cell Biology, 146(2):285–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/2/285.

LaVallee:1998:AMK

[LPM^{+}98] Theresa M. LaVallee, Igor A. Prudovsky, Grainne A. McMahon, Xiaoguo Hu, and Thomas Maciag. Activation of the MAP Kinase Pathway by FGF-1 Correlates with Cell Proliferation Induction While Activation of the Src Pathway Correlates with Migration. Journal of Cell Biology, 141(7):1647–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/7/1647.

Lu:1999:CAJ

[LPM^{+}99] Qun Lu, Mercedes Paredes, Miguel Medina, Jianhua Zhou, Robert Cavallo, Mark Peifer, Lisa Orecchio, and Kenneth S. Kosik. δ-catenin, an Adhesive Junction–associated Protein Which Promotes Cell Scattering. Journal of Cell Biology, 144 (3):519–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/519.

Lew:1995:CCC

[LR95] D. J. Lew and S. I. Reed. A cell cycle checkpoint monitors cell morphogenesis in budding yeast. Journal of Cell Biology, 129(3):739–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/3/739.

Lebakken:1996:SMC

[LR96] C. S. Lebakken and A. C. Rapraeger. Syndecan-1 mediates cell spreading in transfected human lymphoblastoid (Raji) cells. Journal of Cell Biology, 132(6):1209–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/6/1209.

Lord:1998:TER

[LR98] J. Michael Lord and Lynne M. Roberts. Toxin entry: Retrograde transport through the secretory pathway. Journal of Cell Biology, 140(4):733–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/4/733.

Lehman:1999:YHT

[LRAB99] Kevin Lehman, Guendalina Rossi, Joan E. Adamo, and Patrick Brennwald. Yeast Homologues of Tomosyn and lethal giant larvae Function in Exocytosis and Are Associated with the Plasma Membrane Snare, Sec9. Journal of Cell Biology, 146(1):125–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/125.

Linke:1999:BTC

[LRC^{+}99] Wolfgang A. Linke, Diane E. Rudy, Thomas Centner, Mathias Gautel, Christian Witt, Siegfried Labeit, and Carol C. Gregorio. I-band Titin in Cardiac Muscle Is a Three–Element Molecular Spring and Is Critical for Maintaining Thin Filament Structure. Journal of Cell Biology, 146(3):631–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
146/3/631.

Li:1998:TFS

[LRG^{+}98] Rui Li, Philippe Rieu, Diana L. Griffith, David Scott, and M. Amin Arnaout. Two Functional States of the CD11b A–Domain: Correlations with Key Features of Two Mn^{2+}- complexed Crystal Structures. Journal of Cell Biology, 143(6): 1523–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1523.

Langholz:1995:CCG

[LRM^{+}95] O. Langholz, D. Röckel, C. Mauch, E. Kozlowska, I. Bank, T. Krieg, and B. Eckes. Collagen and collagenase gene expression in three-dimensional collagen lattices are differentially regulated by alpha 1 beta 1 and alpha 2 beta 1 integrins. Journal of Cell Biology, 131(6):1903–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1903.

Llorente:1998:EMD

[LRS^{+}98] Alicia Llorente, Andrzej Rapak, Sandra L. Schmid, Bo van Deurs, and Kirsten Sandvig. Expression of Mutant Dynamin Inhibits Toxicity and Transport of Endocytosed Ricin to the Golgi Apparatus. Journal of Cell Biology, 140(3): 553–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/3/553.

Lamaze:1995:REG

[LS95a] C. Lamaze and S. L. Schmid. Recruitment of epidermal growth factor receptors into coated pits requires their activated tyrosine kinase. Journal of Cell Biology, 129(1):47–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/1/47.

Lyman:1995:IBB

[LS95b] S. K. Lyman and R. Schekman. Interaction between BiP and Sec63p is required for the completion of protein translocation into the ER of Saccharomyces cerevisiae. Journal of Cell Biology, 131(5):1163–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/5/1163.

Li:1998:ICC

[LSB98] Gang Li, Gail Sudlow, and Andrew S. Belmont. Interphase cell cycle dynamics of a late-replicating, heterochromatic homogeneously staining region: Precise choreography of condensation/decondensation and nuclear positioning. Journal of Cell Biology, 140(5):975–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/5/975.

Lippincott-Schwartz:1995:KMM

[LSCM^{+}95] J. Lippincott-Schwartz, N. B. Cole, A. Marotta, P. A. Conrad, and G. S. Bloom. Kinesin is the motor for microtubulemediated Golgi–to–ER membrane traffic. Journal of Cell Biology, 128(3):293–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/3/293.

Larsson:1999:OSM

[LSH^{+}99] Niklas Larsson, Bo Segerman, Bonnie Howell, Kajsa Fridell, Lynne Cassimeris, and Martin Gullberg. Op18/ stathmin Mediates Multiple Region–Specific Tubulin and Microtubule– Regulating Activities. Journal of Cell Biology, 146(6):1289– ??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1289.

Lussier:1995:LTS

[LSKB95] M. Lussier, A. M. Sdicu, T. Ketela, and H. Bussey. Localization and targeting of the Saccharomyces cerevisiae Kre2p/ Mnt1p alpha 1,2-mannosyltransferase to a media-Golgi compartment. Journal of Cell Biology, 131(4):913–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/913.

Leung:1999:IFP

[LSL99] Conrad L. Leung, Dongming Sun, and Ronald K. H. Liem. The Intermediate Filament Protein Peripherin Is the Specific Interaction Partner of Mouse BPAG1-n (Dystonin) in Neurons. Journal of Cell Biology, 144(3):435–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
3/435.

Lozano:1998:PDR

[LSM^{+}98] Jean-Claude Lozano, Philippe Schatt, François Marquès, Gérard Peaucellier, Philippe Fort, Jean-Pierre Féral, Anne-Marie Genevière, and André Picard. A Presumptive Developmental Role for a Sea Urchin Cyclin B Splice Variant. Journal of Cell Biology, 140(2):283–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/2/283.

Leung:1999:MAC

[LSZ^{+}99a] Conrad L. Leung, Dongming Sun, Min Zheng, David R. Knowles, and Ronald K. H. Liem. Microtubule Actin Cross– Linking Factor (Macf). Journal of Cell Biology, 147(6):1275– ??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/6/1275.

Luetterforst:1999:MCC

[LSZ^{+}99b] Robert Luetterforst, Espen Stang, Natasha Zorzi, Amanda Carozzi, Michael Way, and Robert G. Parton. Molecular characterization of caveolin association with the Golgi complex: Identification of a cis-Golgi targeting domain in the caveolin molecule. Journal of Cell Biology, 145(7):1443–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
7/1443.

Liu:1996:FAS

[LTE^{+}96] G. Liu, J. Tang, B. T. Edmonds, J. Murray, S. Levin, and J. Condeelis. F-actin sequesters elongation factor 1alpha from interaction with aminoacyl-tRNA in a pH-dependent reaction. Journal of Cell Biology, 135(4):953–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/953.

Lelongt:1997:MMM

[LTMR97] Brigitte Lelongt, Germain Trugnan, Gillian Murphy, and Pierre M. Ronco. Matrix Metalloproteinases MMP2 and MMP9 Are Produced in Early Stages of Kidney Morphogenesis but Only MMP9 Is Required for Renal Organogenesis In Vitro. Journal of Cell Biology, 136(6):1363–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
6/1363.

Larabell:1997:EDV

[LTR^{+}97] Carolyn A. Larabell, Monica Torres, Brian A. Rowning, Cynthia Yost, Jeffrey R. Miller, Mike Wu, David Kimelman, and Randall T. Moon. Establishment of the Dorso-ventral Axis in Xenopus Embryos Is Presaged by Early Asymmetries in β-Catenin That Are Modulated by the Wnt Signaling Pathway.

Journal of Cell Biology, 136(5):1123–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/5/1123.

Lewis:1996:CMF

[LTVC96] S. A. Lewis, G. Tian, I. E. Vainberg, and N. J. Cowan. Chaperonin-mediated folding of actin and tubulin. Journal of Cell Biology, 132(1):1–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/1/1.

Liu:1997:CPA

[LTW^{+}97] Gseping Liu, Laurel Thomas, Robin A. Warren, Caroline A. Enns, C. Casey Cunningham, John H. Hartwig, and Gary Thomas. Cytoskeletal Protein ABP-280 Directs the Intracellular Trafficking of Furin and Modulates Proprotein Processing in the Endocytic Pathway. Journal of Cell Biology, 139(7): 1719–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1719.

Larochelle:1996:NMR

[LVD96] D. A. Larochelle, K. K. Vithalani, and A. De Lozanne. A novel member of the rho family of small GTP-binding proteins is specifically required for cytokinesis. Journal of Cell Biology, 133(6):1321–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/6/1321.

Lorenzon:1998:ECP

[LVN^{+}98] P. Lorenzon, E. Vecile, E. Nardon, E. Ferrero, J. M. Harlan, F. Tedesco, and A. Dobrina. Endothelial Cell E- and P–Selectin and Vascular Cell Adhesion Molecule-1 Function as Signaling Receptors. Journal of Cell Biology, 142(5):1381– ??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/5/1381.

Luderus:1996:SSD

[LvSC^{+}96] M. E. Ludérus, B. van Steensel, L. Chong, O. C. Sibon, F. F. Cremers, and T. de Lange. Structure, subnuclear distribution, and nuclear matrix association of the mammalian telomeric complex. Journal of Cell Biology, 135(4):867–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/867.

Lum:1995:DHC

[LW95] P. Y. Lum and R. Wright. Degradation of HMG–CoA reductase-induced membranes in the fission yeast, Schizosaccharomyces pombe. Journal of Cell Biology, 131(1):81–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/1/81.

Liesi:1996:WGN

[LW96] P. Liesi and J. M. Wright. Weaver granule neurons are rescued by calcium channel antagonists and antibodies against a neurite outgrowth domain of the B2 chain of laminin. Journal of Cell Biology, 134(2):477–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/477.

Liu:1995:TMK

[LWB^{+}95] X. Liu, H. Wu, M. Byrne, J. Jeffrey, S. Krane, and R. Jaenisch. A targeted mutation at the known collagenase cleavage site in mouse type I collagen impairs tissue remodeling. Journal of Cell Biology, 130(1):227–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/1/227.

Lu:1997:DRF

[LWKK97] Mei Lu, Walter Witke, David J. Kwiatkowski, and Kenneth S. Kosik. Delayed Retraction of Filopodia in Gelsolin Null Mice. Journal of Cell Biology, 138(6):1279–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
6/1279.

Liao:1995:CFP

[LWM^{+}95] H. Liao, R. J. Winkfein, G. Mack, J. B. Rattner, and T. J. Yen. CENP-f is a protein of the nuclear matrix that assembles onto kinetochores at late G2 and is rapidly degraded after mitosis. Journal of Cell Biology, 130(3):507–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
3/507.

Lauber:1997:AKP

[LWS^{+}97a] Martina H. Lauber, Irene Waizenegger, Thomas Steinmann, Heinz Schwarz, Ulrike Mayer, Inwhan Hwang, Wolfgang Lukowitz, and Gerd Jürgens. The Arabidopsis KNOLLE Protein Is a Cytokinesis-specific Syntaxin. Journal of Cell Biology, 139(6):1485–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/6/1485.

Lewis:1997:CTB

[LWS^{+}97b] Jani E. Lewis, James K. Wahl, Kristin M. Sass, Pamela J. Jensen, Keith R. Johnson, and Margaret J. Wheelock. Crosstalk between Adherens Junctions and Desmosomes Depends on Plakoglobin. Journal of Cell Biology, 136(4):919–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/4/919.

Liang:1999:SMP

[LWS99] Wenchuan Liang, Hans M. Warrick, and James A. Spudich. A Structural Model for Phosphorylation Control of Dictyostelium Myosin II Thick Filament Assembly. Journal of Cell Biology, 147(5):1039–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/5/1039.

Lepple-Wienhues:1998:TKP

[LWSL^{+}98] Albrecht Lepple-Wienhues, Ildikò Szabò, Tilmann Laun, Nubia Kristen Kaba, Erich Gulbins, and Florian Lang. The Tyrosine Kinase p56 ^{lck} Mediates Activation of Swelling-induced Chloride Channels in Lymphocytes. Journal of Cell Biology, 141(1):281–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/281.

Li:1998:DCP

[LXC^{+}98] Kaijun Li, Eugene Yujun Xu, Jeffrey K. Cecil, F. Rudolf Turner, Timothy L. Megraw, and Thomas C. Kaufman. Drosophila Centrosomin Protein is Required for Male Meiosis and Assembly of the Flagellar Axoneme. Journal of Cell Biology, 141(2):455–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/2/455.

Li:1995:PDI

[LXK^{+}95] R. Li, J. Xie, C. Kantor, V. Koistinen, D. C. Altieri, P. Nortamo, and C. G. Gahmberg. A peptide derived from the intercellular adhesion molecule-2 regulates the avidity of the leukocyte integrins CD11b/CD18 and CD11c/CD18. Journal of Cell Biology, 129(4):1143–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/4/1143.

Lloyd:1995:BKN

[LYC^{+}95] C. Lloyd, Q. C. Yu, J. Cheng, K. Turksen, L. Degenstein, E. Hutton, and E. Fuchs. The basal keratin network of stratified squamous epithelia: defining K15 function in the absence of K14. Journal of Cell Biology, 129(5):1329–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1329.

Lo:1997:PKD

[LYD^{+}97] Su Hao Lo, Qian-Chun Yu, Linda Degenstein, Lan Bo Chen, and Elaine Fuchs. Progressive Kidney Degeneration in Mice Lacking Tensin. Journal of Cell Biology, 136(6):1349–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/6/1349.

Lin:1999:OCN

[LYH^{+}99] Ping Lin, Yong Yao, Robert Hofmeister, Roger Y. Tsien, and Marilyn Gist Farquhar. Overexpression of CALNUC (Nucleobindin) Increases Agonist and Thapsigargin Releasable Ca^{2+} Storage in the Golgi. Journal of Cell Biology, 145(2):279–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/279.

Lu:1998:RAI

[LYLR98] Di Lu, Hong Yang, Robert H. Lenox, and Mohan K. Raizada. Regulation of Angiotensin II–induced Neuromodulation by MARCKS in Brain Neurons. Journal of Cell Biology, 142 (1):217–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/1/217.

Lu:1996:AIR

[LYR96] D. Lu, H. Yang, and M. K. Raizada. Angiotensin II regulation of neuromodulation: downstream signaling mechanism from activation of mitogen-activated protein kinase. Journal of Cell Biology, 135(6):1609–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1609.

Le:1999:RC

[LYS99] Tam Luan Le, Alpha S. Yap, and Jennifer L. Stow. Recycling of E–Cadherin. Journal of Cell Biology, 146(1):219–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
1/219.

Lorenzon:1995:CWP

[LZC^{+}95] P. Lorenzon, D. Zacchetti, F. Codazzi, G. Fumagalli, J. Meldolesi, and F. Grohovaz. Ca^{2+} waves in PC12 neurites: a bidirectional, receptor-oriented form of Ca^{2+} signaling. Journal of Cell Biology, 129(3):797–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/3/797.

Li:1995:RCA

[LZD95] R. Li, Y. Zheng, and D. G. Drubin. Regulation of cortical actin cytoskeleton assembly during polarized cell growth in budding yeast. Journal of Cell Biology, 128(4):599–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/4/599.

Laine:1997:VPU

[LZK^{+}97] Roney O. Laine, William Zeile, Fan Kang, Daniel L. Purich, and Frederick S. Southwick. Vinculin Proteolysis Unmasks an ActA Homolog for Actin-based Shigella Motility. Journal of Cell Biology, 138(6):1255–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/6/1255.

Mustonen:1995:ERT

[MA95] T. Mustonen and K. Alitalo. Endothelial receptor tyrosine kinases involved in angiogenesis. Journal of Cell Biology, 129 (4):895–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/4/895.

McArdle:1998:CDM

[MAB98] Kristen McArdle, Taylor StC. Allen, and Elizabeth A. Bucher. Ca^{2+}-dependent muscle dysfunction caused by mutation of the Caenorhabditis elegans troponin T-1 gene. Journal of Cell Biology, 143(5):1201–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/5/1201.

Mancini:1997:MPP

[MAC^{+}97] Mariangela Mancini, Benjamin O. Anderson, Elizabeth Caldwell, Monireh Sedghinasab, Philip B. Paty, and David M. Hockenbery. Mitochondrial Proliferation and Paradoxical Membrane Depolarization during Terminal Differentiation and Apoptosis in a Human Colon Carcinoma Cell Line. Journal of Cell Biology, 138(2):449–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/2/449.

Milligan:1997:PTP

[MAE^{+}97] Scott C. Milligan, James G. Alb, Raya B. Elagina, Vytas A. Bankaitis, and David R. Hyde. The Phosphatidylinositol Transfer Protein Domain of Drosophila Retinal Degeneration B Protein Is Essential for Photoreceptor Cell Survival and Recovery from Light Stimulation. Journal of Cell Biology, 139 (2):351–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/351.

Mackay:1998:DMI

[MAEE98] Alastair M. Mackay, Alexandra M. Ainsztein, D. Mark Eckley, and William C. Earnshaw. A Dominant Mutant of Inner Centromere Protein (INCENP), a Chromosomal Protein, Disrupts Prometaphase Congression and Cytokinesis. Journal of Cell Biology, 140(5):991–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/5/991.

Marrs:1995:PEC

[MAFJ^{+}95] J. A. Marrs, C. Andersson-Fisone, M. C. Jeong, L. Cohen-Gould, C. Zurzolo, I. R. Nabi, E. Rodriguez-Boulan, and W. J. Nelson. Plasticity in epithelial cell phenotype: modulation by expression of different cadherin cell adhesion molecules. Journal of Cell Biology, 129(2):507–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/2/507.

Maroto:1996:DPM

[MAG^{+}96] M. Maroto, J. Arredondo, D. Goulding, R. Marco, B. Bullard, and M. Cervera. Drosophila paramyosin/miniparamyosin gene products show a large diversity in quantity, localization, and isoform pattern: a possible role in muscle maturation and function. Journal of Cell Biology, 134(1):81–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/134/1/81.

Maher:1996:NTF

[Mah96] P. A. Maher. Nuclear Translocation of fibroblast growth factor (FGF) receptors in response to FGF-2. Journal of Cell Biology, 134(2):529–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/529.

Martineau:1995:DHC

[MAM95] S. N. Martineau, P. R. Andreassen, and R. L. Margolis. Delay of HeLa cell cleavage into interphase using dihydrocytochalasin B: retention of a postmitotic spindle and telophase disc correlates with synchronous cleavage recovery. Journal of Cell Biology, 131(1):191–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/1/191.

Michele:1999:TFP

[MAM99] Daniel E. Michele, Faris P. Albayya, and Joseph M. Metzger. Thin filament protein dynamics in fully differentiated adult cardiac myocytes: Toward a model of sarcomere maintenance. Journal of Cell Biology, 145(7):1483–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/7/1483.

Marsh:1995:MRG

[MAMJ95] B. J. Marsh, R. A. Alm, S. R. McIntosh, and D. E. James. Molecular regulation of GLUT-4 targeting in 3T3-L1 adipocytes. Journal of Cell Biology, 130(5):1081–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/5/1081.

Martin:1995:ENT

[MAS^{+}95] M. Martin, C. Andréoli, A. Sahuquet, P. Montcourrier, M. Algrain, and P. Mangeat. Ezrin NH2-terminal domain inhibits the cell extension activity of the COOH-terminal domain. Journal of Cell Biology, 128(6):1081–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/6/1081.

Montero:1997:CHE

[MAS^{+}97] Mayte Montero, Javier Alvarez, Wilhelm J. J. Scheenen, Rosario Rizzuto, Jacopo Meldolesi, and Tullio Pozzan. Ca^{2+} homeostasis in the endoplasmic reticulum: Coexistence of high and low [Ca^{2+}] subcompartments in intact HeLa cells. Journal of Cell Biology, 139(3):601–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/3/601.

Mallard:1998:DPE

[MAT^{+}98] Frédéric Mallard, Claude Antony, Danièle Tenza, Jean Salamero, Bruno Goud, and Ludger Johannes. Direct pathway from early/recycling endosomes to the Golgi apparatus revealed through the study of Shiga toxin B-fragment transport. Journal of Cell Biology, 143(4):973–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/973.

Marelli:1998:SBK

[MAW98] Marcello Marelli, John D. Aitchison, and Richard W. Wozniak. Specific Binding of the Karyopherin Kap121p to a Subunit of the Nuclear Pore Complex Containing Nup53p, Nup59p, and Nup170p. Journal of Cell Biology, 143(7):1813–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1813.

McCracken:1996:AAP

[MB96] A. A. McCracken and J. L. Brodsky. Assembly of ERassociated protein degradation in vitro: dependence on cytosol, calnexin, and ATP. Journal of Cell Biology, 132(3): 291–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/291.

McDonald:1997:PCS

[MB97] Heather B. McDonald and Breck Byers. A Proteasome Cap Subunit Required for Spindle Pole Body Duplication in Yeast. Journal of Cell Biology, 137(3):539–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/3/539.

Mas:1999:RTM

[MB99] Paloma Más and Roger N. Beachy. Replication of tobacco mosaic virus on endoplasmic reticulum and role of the cytoskeleton and virus movement protein in intracellular distribution of viral RNA. Journal of Cell Biology, 147(5):945–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
5/945.

Ma:1997:PAP

[MBA97] Alice D. Ma, Lawrence F. Brass, and Charles S. Abrams. Pleckstrin associates with plasma membranes and induces the formation of membrane projections: Requirements for phosphorylation and the NH_{2}-terminal PH domain. Journal of Cell Biology, 136(5):1071–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/5/1071.

Martin-Bermudo:1996:ISD

[MBB96] M. D. Martin-Bermudo and N. H. Brown. Intracellular signals direct integrin localization to sites of function in embryonic muscles. Journal of Cell Biology, 134(1):217–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
1/217.

Molina:1997:CAK

[MBB^{+}97] Isabel Molina, Sigrid Baars, Julie A. Brill, Karen G. Hales, Margaret T. Fuller, and Pedro Ripoll. A Chromatin-associated Kinesin-related Protein Required for Normal Mitotic Chromosome Segregation in Drosophila. Journal of Cell Biology, 139(6):1361–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/6/1361.

Muslimov:1998:ADR

[MBBT98] Ilham A. Muslimov, Gary Banker, Jürgen Brosius, and Henri Tiedge. Activity-dependent Regulation of Dendritic BC1 RNA in Hippocampal Neurons in Culture. Journal of Cell Biology, 141(7):1601–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/7/1601.

Miralles:1998:TPK

[MBCS98] Francisco Miralles, Tadej Battelino, Paul Czernichow, and Raphael Scharfmann. TGF-β Plays a Key Role in Morphogenesis of the Pancreatic Islets of Langerhans by Controlling the Activity of the Matrix Metalloproteinase MMP-2. Journal of Cell Biology, 143(3):827–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/3/827.

Menegazzi:1996:CIE

[MBD^{+}96] R. Menegazzi, S. Busetto, P. Dri, R. Cramer, and P. Patriarca. Chloride ion efflux regulates adherence, spreading, and respiratory burst of neutrophils stimulated by tumor necrosis factor-alpha (TNF) on biologic surfaces. Journal of Cell Biology, 135(2):511–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/2/511.

Martin-Bermudo:1998:MIA

[MBDBB98] Maria D. Martin-Bermudo, Olga M. Dunin-Borkowski, and Nicholas H. Brown. Modulation of Integrin Activity is Vital for Morphogenesis. Journal of Cell Biology, 141(4):1073–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/4/1073.

Munoz-Barroso:1998:DHI

[MBDS^{+}98] Isabel Muñoz-Barroso, Stewart Durell, Kazuyasu Sakaguchi, Ettore Appella, and Robert Blumenthal. Dilation of the Human Immunodeficiency Virus–1 Envelope Glycoprotein Fusion Pore Revealed by the Inhibitory Action of a Synthetic Peptide from gp41. Journal of Cell Biology, 140(2):315–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
2/315.

Merlo:1995:PDP

[MBF^{+}95a] G. R. Merlo, F. Basolo, L. Fiore, L. Duboc, and N. E. Hynes. p53-dependent and p53-independent activation of apoptosis in mammary epithelial cells reveals a survival function of EGF and insulin. Journal of Cell Biology, 128(6):1185–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
6/1185.

Moritz:1995:TDS

[MBF^{+}95b] M. Moritz, M. B. Braunfeld, J. C. Fung, J. W. Sedat, B. M. Alberts, and D. A. Agard. Three-dimensional structural characterization of centrosomes from early Drosophila embryos. Journal of Cell Biology, 130(5):1149–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
5/1149.

Mukhopadhyay:1997:SAR

[MBF^{+}97] Amitabha Mukhopadhyay, Alejandro M. Barbieri, Kouichi Funato, Richard Roberts, and Philip D. Stahl. Sequential Actions of Rab5 and Rab7 Regulate Endocytosis in the Xenopus Oocyte. Journal of Cell Biology, 136(6):1227–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
6/1227.

Mullock:1998:FLL

[MBF^{+}98] Barbara M. Mullock, Nicholas A. Bright, Clare W. Fearon, Sally R. Gray, and J. Luzio. Fusion of Lysosomes with Late Endosomes Produces a Hybrid Organelle of Intermediate Density and Is NSF Dependent. Journal of Cell Biology, 140(3):

591–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/3/591.

Mayinger:1995:SMA

[MBM95] P. Mayinger, V. A. Bankaitis, and D. I. Meyer. Sac1p mediates the adenosine triphosphate transport into yeast endoplasmic reticulum that is required for protein translocation. Journal of Cell Biology, 131(6):1377–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1377.

Melikyan:1997:IOM

[MBOC97] Grigory B. Melikyan, Sofya A. Brener, Dong C. Ok, and Fredric S. Cohen. Inner but Not Outer Membrane Leaflets Control the Transition from Glycosylphosphatidylinositolanchored Influenza Hemagglutinin-induced Hemifusion to Full Fusion. Journal of Cell Biology, 136(5):995–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/5/
995.

McCollum:1995:SPC

[MBP^{+}95] D. McCollum, M. K. Balasubramanian, L. E. Pelcher, S. M. Hemmingsen, and K. L. Gould. Schizosaccharomyces pombe cdc4+ gene encodes a novel EF-hand protein essential for cytokinesis. Journal of Cell Biology, 130(3):651–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
3/651.

Matsuoka:1995:DSW

[MBRN95] K. Matsuoka, D. C. Bassham, N. V. Raikhel, and K. Nakamura. Different sensitivity to wortmannin of two vacuolar sorting signals indicates the presence of distinct sorting machineries in tobacco cells. Journal of Cell Biology, 130(6): 1307–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1307.

Mastick:1995:IST

[MBS95] C. C. Mastick, M. J. Brady, and A. R. Saltiel. Insulin stimulates the tyrosine phosphorylation of caveolin. Journal of Cell Biology, 129(6):1523–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/6/1523.

Majoul:1996:TEL

[MBS96a] I. V. Majoul, P. I. Bastiaens, and H. D. Söling. Transport of an external Lys–Asp–Glu–Leu (KDEL) protein from the plasma membrane to the endoplasmic reticulum: studies with cholera toxin in Vero cells. Journal of Cell Biology, 133(4):777–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
4/777.

Montgomery:1996:HNC

[MBS^{+}96b] A. M. Montgomery, J. C. Becker, C. H. Siu, V. P. Lemmon, D. A. Cheresh, J. D. Pancook, X. Zhao, and R. A. Reisfeld. Human neural cell adhesion molecule L1 and rat homologue NILE are ligands for integrin alpha v beta 3. Journal of Cell Biology, 132(3):475–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/3/475.

Makarenkova:1997:FGF

[MBTW97] H. Makarenkova, D. L. Becker, C. Tickle, and A. E. Warner. Fibroblast Growth Factor 4 Directs Gap Junction Expression in the Mesenchyme of the Vertebrate Limb Bud. Journal of Cell Biology, 138(5):1125–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/5/1125.

Manning:1999:DRK

[MBW^{+}99] Brendan D. Manning, Jennifer G. Barrett, Julie A. Wallace, Howard Granok, and Michael Snyder. Differential Regulation of the Kar3p Kinesin-related Protein by Two Associated Proteins, Cik1p and Vik1p. Journal of Cell Biology, 144(6): 1219–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/6/1219.

Marschall:1995:NCA

[MC95] L. G. Marschall and L. Clarke. A novel cis-acting centromeric DNA element affects S. pombe centromeric chromatin structure at a distance. Journal of Cell Biology, 128(4):

445–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/445.

Merdes:1997:PSP

[MC97] Andreas Merdes and Don W. Cleveland. Pathways of Spindle Pole Formation: Different Mechanisms; Conserved Components. Journal of Cell Biology, 138(5):953–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
5/953.

McGowan:1998:OKE

[MC98] Kevin M. McGowan and Pierre A. Coulombe. Onset of Keratin 17 Expression Coincides with the Definition of Major Epithelial Lineages during Skin Development. Journal of Cell Biology, 143(2):469–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/469.

Maryon:1996:UER

[MCA96] E. B. Maryon, R. Coronado, and P. Anderson. unc-68 encodes a ryanodine receptor involved in regulating C. elegans body-wall muscle contraction. Journal of Cell Biology, 134 (4):885–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/885.

Matunis:1996:NUL

[MCB96] M. J. Matunis, E. Coutavas, and G. Blobel. A novel ubiquitinlike modification modulates the partitioning of the Ran– GTPase–activating protein RanGAP1 between the cytosol and the nuclear pore complex. Journal of Cell Biology, 135(6): 1457–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1457.

Melby:1998:SSS

[MCBE98] Thomas E. Melby, Charles N. Ciampaglio, Gina Briscoe, and Harold P. Erickson. The Symmetrical Structure of Structural Maintenance of Chromosomes (SMC) and MukB Proteins: Long, Antiparallel Coiled Coils, Folded at a Flexible Hinge. Journal of Cell Biology, 142(6):1595–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1595.

Chan:1998:MOS

[mCBLK98] Yiu mo Chan, Carsten G. Bönnemann, Hart G. W. Lidov, and Louis M. Kunkel. Molecular Organization of Sarcoglycan Complex in Mouse Myotubes in Culture. Journal of Cell Biology, 143(7):2033–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/7/2033.

Mahbubani:1997:CCR

[MCC^{+}97] Hiro M. Mahbubani, James P. J. Chong, Stephane Chevalier, Pia Thömmes, and J. Julian Blow. Cell Cycle Regulation of the Replication Licensing System: Involvement of a Cdk-dependent Inhibitor. Journal of Cell Biology, 136(1): 125–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/125.

Misteli:1998:SPS

[MCC^{+}98] Tom Misteli, Javier F. Cáceres, Jade Q. Clement, Adrian R. Krainer, Miles F. Wilkinson, and David L. Spector. Serine Phosphorylation of SR Proteins Is Required for Their Recruitment to Sites of Transcription In Vivo. Journal of Cell Biology, 143(2):297–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/297.

Ma:1998:CSP

[MCJK98] Le Ma, Lewis C. Cantley, Paul A. Janmey, and Marc W. Kirschner. Corequirement of Specific Phosphoinositides and Small GTP-binding Protein Cdc42 in Inducing Actin Assembly in Xenopus Egg Extracts. Journal of Cell Biology, 140 (5):1125–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/5/1125.

Marfatia:1996:MOP

[MCL^{+}96] S. M. Marfatia, J. H. Morais Cabral, L. Lin, C. Hough, P. J. Bryant, L. Stolz, and A. H. Chishti. Modular organization of the PDZ domains in the human discs-large protein suggests a mechanism for coupling PDZ domain-binding proteins to ATP and the membrane cytoskeleton. Journal of Cell Biology, 135(3):753–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/3/753.

Maddox:1999:MDM

[MCM^{+}99] Paul Maddox, E. Chin, A. Mallavarapu, E. Yeh, E. D. Salmon, and K. Bloom. Microtubule Dynamics from Mating through the First Zygotic Division in the Budding Yeast Saccharomyces cerevisiae. Journal of Cell Biology, 144(5):977–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
5/977.

Mermoud:1999:HMR

[MCPB99] Jacqueline E. Mermoud, Carl Costanzi, John R. Pehrson, and Neil Brockdorff. Histone Macroh2a1.2 Relocates to the Inactive X Chromosome after Initiation and Propagation of X– Inactivation. Journal of Cell Biology, 147(7):1399–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/7/1399.

Musch:1997:MII

[MCRB97] Anne Müsch, David Cohen, and Enrique Rodriguez-Boulan. Myosin II Is Involved in the Production of Constitutive Transport Vesicles from the TGN. Journal of Cell Biology, 138 (2):291–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/2/291.

MacFarlane:1997:PAL

[MCS^{+}97a] Marion MacFarlane, Kelvin Cain, Xiao-Ming Sun, Emad S. Alnemri, and Gerald M. Cohen. Processing/ activation of At Least Four Interleukin-1β Converting Enzyme-like Proteases Occurs during the Execution Phase of Apoptosis in Human Monocytic Tumor Cells. Journal of Cell Biology, 137 (2):469–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/2/469.

Mironov:1997:RNA

[MCS^{+}97b] Alexander Mironov, Antonino Colanzi, Maria Giuseppina Silletta, Giusy Fiucci, Silvio Flati, Aurora Fusella, Roman Polishchuk, Alexander Mironov, Giuseppe Di Tullio, Roberto Weigert, Vivek Malhotra, Daniela Corda, Maria Antonietta De Matteis, and Alberto Luini. Role of NAD^{+} and ADP– Ribosylation in the Maintenance of the Golgi Structure. Journal of Cell Biology, 139(5):1109–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/5/1109.

Miner:1998:RLE

[MCS98] Jeffrey H. Miner, Jeanette Cunningham, and Joshua R. Sanes. Roles for Laminin in Embryogenesis: Exencephaly, Syndactyly, and Placentopathy in Mice Lacking the Laminin α5 Chain. Journal of Cell Biology, 143(6):1713–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1713.

Matthews:1998:ZCE

[MCTMK98] Lisa R. Matthews, Philip Carter, Danielle Thierry-Mieg, and Ken Kemphues. ZYG-9, A Caenorhabditis elegans Protein Required for Microtubule Organization and Function, Is a Component of Meiotic and Mitotic Spindle Poles. Journal of Cell Biology, 141(5):1159–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/5/1159.

Murray:1999:OPK

[MDC^{+}99] Nicole R. Murray, Laurie A. Davidson, Robert S. Chapkin, W. Clay Gustafson, Diane G. Schattenberg, and Alan P. Fields. Overexpression of Protein Kinase C β_{II} Induces Colonic Hyperproliferation and Increased Sensitivity to Colon Carcinogenesis. Journal of Cell Biology, 145(4):699–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
4/699.

Martinou:1999:RCC

[MDE^{+}99] Isabelle Martinou, Solange Desagher, Robert Eskes, Bruno Antonsson, Elisabeth André, Stanislav Fakan, and Jean-Claude Martinou. The Release of Cytochrome c from Mitochondria during Apoptosis of NGF-deprived Sympathetic Neurons Is a Reversible Event. Journal of Cell Biology, 144 (5):883–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/5/883.

McCartney:1999:DAC

[MDK^{+}99] Brooke M. McCartney, Herman A. Dierick, Catherine Kirkpatrick, Melissa M. Moline, Annette Baas, Mark Peifer, and Amy Bejsovec. Drosophila Apc2 Is a Cytoskeletally– Associated Protein That Regulates Wingless Signaling in the Embryonic Epidermis. Journal of Cell Biology, 146(6):1303– ??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1303.

Marshall:1996:RSH

[MDQG96] P. A. Marshall, J. M. Dyer, M. E. Quick, and J. M. Goodman. Redox-sensitive homodimerization of Pex11p: a proposed mechanism to regulate peroxisomal division. Journal of Cell Biology, 135(1):123–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/1/123.

Mehlin:1995:SIB

[MDS95] H. Mehlin, B. Daneholt, and U. Skoglund. Structural interaction between the nuclear pore complex and a specific translocating RNP particle. Journal of Cell Biology, 129(5):1205–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/5/1205.

Moll:1998:VDE

[MDV98] Thomas Moll, Elisabetta Dejana, and Dietmar Vestweber. In Vitro Degradation of Endothelial Catenins by a Neutrophil Protease. Journal of Cell Biology, 140(2):403–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
2/403.

Mackay:1997:RRD

[MEFH97] Deborah J. G. Mackay, Fred Esch, Heinz Furthmayr, and Alan Hall. Rho- and Rac-dependent assembly of focal adhesion complexes and actin filaments in permeabilized fibroblasts: an essential role for ezrin/radixin/moesin proteins. Journal of Cell Biology, 138(4):927–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/4/927.

Murray:1996:BAF

[MELC96] J. W. Murray, B. T. Edmonds, G. Liu, and J. Condeelis. Bundling of actin filaments by elongation factor 1 alpha inhibits polymerization at filament ends. Journal of Cell Biology, 135(5):1309–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/5/1309.

Murphy-Erdosh:1995:CBS

[MEYPR95] C. Murphy-Erdosh, C. K. Yoshida, N. Paradies, and L. F. Reichardt. The cadherin-binding specificities of B-cadherin and LCAM. Journal of Cell Biology, 129(5):1379–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1379.

Macaulay:1996:ANP

[MF96a] C. Macaulay and D. J. Forbes. Assembly of the nuclear pore: biochemically distinct steps revealed with NEM, GTP gamma S, and BAPTA. Journal of Cell Biology, 132(1):5–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
1/5.

McCartney:1996:DCS

[MF96b] B. M. McCartney and R. G. Fehon. Distinct cellular and subcellular patterns of expression imply distinct functions for the Drosophila homologues of moesin and the neurofibromatosis 2 tumor suppressor, merlin. Journal of Cell Biology, 133 (4):843–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/4/843.

Martin:1999:ILG

[MFBK99] David C. Martin, John L. Fowlkes, Bojana Babic, and Rama Khokha. Insulin-like Growth Factor II Signaling in Neoplastic Proliferation Is Blocked by Transgenic Expression of the Metalloproteinase Inhibitor Timp-1. Journal of Cell Biology, 146 (4):881–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/881.

Madeo:1997:YMS

[MFF97] Frank Madeo, Eleonore Fröhlich, and Kai-Uwe Fröhlich. A Yeast Mutant Showing Diagnostic Markers of Early and Late Apoptosis. Journal of Cell Biology, 139(3):729–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/729.

Madeo:1999:OSR

[MFL^{+}99] Frank Madeo, Eleonore Fröhlich, Martin Ligr, Martin Grey, Stephan J. Sigrist, Dieter H. Wolf, and Kai-Uwe Fröhlich. Oxygen stress: a regulator of apoptosis in yeast. Journal of Cell Biology, 145(4):757–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/4/757.

Martinou:1995:VPE

[MFM^{+}95] I. Martinou, P. A. Fernandez, M. Missotten, E. White, B. Allet, R. Sadoul, and J. C. Martinou. Viral proteins E1B19K and p35 protect sympathetic neurons from cell death induced by NGF deprivation. Journal of Cell Biology, 128(1):201–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/1/201.

Matera:1995:PCC

[MFMW95] A. G. Matera, M. R. Frey, K. Margelot, and S. L. Wolin. A perinucleolar compartment contains several RNA polymerase III transcripts as well as the polypyrimidine tract-binding protein, hnRNP I. Journal of Cell Biology, 129(5):1181–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1181.

Morris:1996:INA

[MG96] E. J. Morris and H. M. Geller. Induction of neuronal apoptosis by camptothecin, an inhibitor of DNA topoisomerase–I: evidence for cell cycle-independent toxicity. Journal of Cell Biology, 134(3):757–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/3/757.

Myers:1996:DGA

[MGA^{+}96] C. D. Myers, P. Y. Goh, T. S. Allen, E. A. Bucher, and T. Bogaert. Developmental genetic analysis of troponin T mutations in striated and nonstriated muscle cells of Caenorhabditis elegans. Journal of Cell Biology, 132(6):1061–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
6/1061.

Magez:1997:SUT

[MGB^{+}97] Stefan Magez, Maurice Geuskens, Alain Beschin, Herwig del Favero, Hendrik Verschueren, Ralf Lucas, Etienne Pays, and Patrick de Baetselier. Specific Uptake of Tumor Necrosis Factor-α Is Involved in Growth Control of Trypanosoma brucei. Journal of Cell Biology, 137(3):715–??, May 1997. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/137/3/
715.

Marmorstein:1998:APC

[MGB^{+}98] Alan D. Marmorstein, Yunbo C. Gan, Vera L. Bonilha, Silvia C. Finnemann, Karl G. Csaky, and Enrique Rodriguez-Boulan. Apical Polarity of N-CAM and EMMPRIN in Retinal Pigment Epithelium Resulting from Suppression of Basolateral Signal Recognition. Journal of Cell Biology, 142(3): 697–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/3/697.

Monier-Gavelle:1997:CTB

[MGD97] Frédérique Monier-Gavelle and Jean-Loup Duband. Cross Talk between Adhesion Molecules: Control of N-cadherin Activity by Intracellular Signals Elicited by β1 and β3 Integrins in Migrating Neural Crest Cells. Journal of Cell Biology, 137 (7):1663–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/7/1663.

Martin:1998:XTA

[MGIZ98] Ona C. Martin, Ruwanthi N. Gunawardane, Akihiro Iwamatsu, and Yixian Zheng. Xgrip109: a γ Tubulin–Associated Protein with an Essential Role in γ Tubulin Ring Complex (γTuRC) Assembly and Centrosome Function. Journal of Cell Biology, 141(3):675–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/3/675.

McGoldrick:1995:MAN

[MGM95] C. A. McGoldrick, C. Gruver, and G. S. May. myoA of Aspergillus nidulans encodes an essential myosin I required for secretion and polarized growth. Journal of Cell Biology, 128 (4):577–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/577.

Mahajan:1998:MCS

[MGM98] Rohit Mahajan, Larry Gerace, and Frauke Melchior. Molecular Characterization of the SUMO-1 Modification of RanGAP1 and Its Role in Nuclear Envelope Association. Journal of Cell Biology, 140(2):259–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/2/259.

Marquis:1997:PPA

[MGP97] Hélène Marquis, Howard Goldfine, and Daniel A. Portnoy. Proteolytic Pathways of Activation and Degradation of a Bacterial Phospholipase C during Intracellular Infection by Listeria monocytogenes. Journal of Cell Biology, 137(6):1381–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/6/1381.

Muresan:1996:PEM

[MGRS96] V. Muresan, C. P. Godek, T. S. Reese, and B. J. Schnapp. Plus-end motors override minus-end motors during transport of squid axon vesicles on microtubules. Journal of Cell Biology, 135(2):383–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/2/383.

McGrail:1995:RCD

[MGS^{+}95] M. McGrail, J. Gepner, A. Silvanovich, S. Ludmann, M. Serr, and T. S. Hays. Regulation of cytoplasmic dynein function in vivo by the Drosophila Glued complex. Journal of Cell Biology, 131(2):411–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/2/411.

McBride:1996:HMI

[MGS96] H. M. McBride, I. S. Goping, and G. C. Shore. The human mitochondrial import receptor, hTom20p, prevents a cryptic matrix targeting sequence from gaining access to the protein translocation machinery. Journal of Cell Biology, 134 (2):307–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/307.

Melchior:1995:GHR

[MGY^{+}95] F. Melchior, T. Guan, N. Yokoyama, T. Nishimoto, and L. Gerace. GTP hydrolysis by Ran occurs at the nuclear pore complex in an early step of protein import. Journal of Cell Biology, 131(3):571–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/3/571.

Mickey:1995:RMI

[MH95a] B. Mickey and J. Howard. Rigidity of microtubules is increased by stabilizing agents. Journal of Cell Biology, 130 (4):909–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/909.

Morris:1995:ATM

[MH95b] R. L. Morris and P. J. Hollenbeck. Axonal transport of mitochondria along microtubules and F-actin in living vertebrate neurons. Journal of Cell Biology, 131(5):1315–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
5/1315.

Machesky:1997:RAP

[MH97] Laura M. Machesky and Alan Hall. Role of Actin Polymerization and Adhesion to Extracellular Matrix in Rac- and Rhoinduced Cytoskeletal Reorganization. Journal of Cell Biology, 138(4):913–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/4/913.

McEwen:1997:KFM

[MHC^{+}97] Bruce F. McEwen, Amy B. Heagle, Grisel O. Cassels, Karolyn F. Buttle, and Conly L. Rieder. Kinetochore Fiber Maturation in PtK_{1} Cells and Its Implications for the Mechanisms of Chromosome Congression and Anaphase Onset. Journal of Cell Biology, 137(7):1567–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/7/1567.

Mattagajasingh:1999:NIP

[MHH^{+}99] Subhendra N. Mattagajasingh, Shu-Ching Huang, Julia S. Hartenstein, Michael Snyder, Vincent T. Marchesi, and Edward J. Benz. A Nonerythroid Isoform of Protein 4.1R Interacts with the Nuclear Mitotic Apparatus (NuMA) Protein. Journal of Cell Biology, 145(1):29–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/1/29.

McCormick:1995:SAS

[MHK^{+}95] B. A. McCormick, P. M. Hofman, J. Kim, D. K. Carnes, S. I. Miller, and J. L. Madara. Surface attachment of Salmonella typhimurium to intestinal epithelia imprints the subepithelial matrix with gradients chemotactic for neutrophils. Journal of Cell Biology, 131(6):1599–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1599.

Marchant:1996:RTV

[MHLB96] J. K. Marchant, R. A. Hahn, T. F. Linsenmayer, and D. E. Birk. Reduction of type V collagen using a dominant-negative strategy alters the regulation of fibrillogenesis and results in the loss of corneal-specific fibril morphology. Journal of Cell Biology, 135(5):1415–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/5/1415.

Mundel:1997:SAA

[MHM^{+}97] Peter Mundel, Hans W. Heid, Thomas M. Mundel, Meike Krüger, Jochen Reiser, and Wilhelm Kriz. Synaptopodin: an Actin-associated Protein in Telencephalic Dendrites and Renal Podocytes. Journal of Cell Biology, 139(1):193–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
1/193.

Miguelez:1999:HDD

[MHM99] Elisa M. Miguélez, Carlos Hardisson, and Manuel B. Manzanal. Hyphal death during colony development in Streptomyces antibioticus: Morphological evidence for the existence of a process of cell deletion in a multicellular prokaryote. Journal of Cell Biology, 145(3):515–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/3/515.

Maney:1998:MCA

[MHWW98] Todd Maney, Andrew W. Hunter, Mike Wagenbach, and Linda Wordeman. Mitotic Centromere–associated Kinesin Is Important for Anaphase Chromosome Segregation. Journal of Cell Biology, 142(3):787–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/3/787.

Machesky:1999:SAD

[MI99] Laura M. Machesky and Robert H. Insall. Signaling to Actin Dynamics. Journal of Cell Biology, 146(2):267–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
2/267.

Miller:1996:TTN

[MJ96] K. E. Miller and H. C. Joshi. Tubulin transport in neurons. Journal of Cell Biology, 133(6):1355–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/6/1355.

Mothes:1998:SSR

[MJBR98] Walther Mothes, Berit Jungnickel, Josef Brunner, and Tom A. Rapoport. Signal Sequence Recognition in Cotranslational Translocation by Protein Components of the Endoplasmic Reticulum Membrane. Journal of Cell Biology, 142(2):355–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
142/2/355.

Marschall:1996:ATY

[MJMS96] L. G. Marschall, R. L. Jeng, J. Mulholland, and T. Stearns. Analysis of Tub4p, a yeast gamma-tubulin-like protein: implications for microtubule-organizing center function. Journal of Cell Biology, 134(2):443–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/443.

Masson:1995:BMM

[MK95a] D. Masson and T. E. Kreis. Binding of E–MAP–115 to microtubules is regulated by cell cycle-dependent phosphorylation. Journal of Cell Biology, 131(4):1015–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/1015.

Miller:1995:FD

[MK95b] K. G. Miller and D. P. Kiehart. Fly division. Journal of Cell Biology, 131(1):1–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/1/1.

Mayer:1995:EMG

[MKB^{+}95] R. Mayer, J. Kartenbeck, M. Büchler, G. Jedlitschky, I. Leier, and D. Keppler. Expression of the MRP gene-encoded conjugate export pump in liver and its selective absence from the canalicular membrane in transport-deficient mutant hepatocytes. Journal of Cell Biology, 131(1):137–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
1/137.

Mohler:1999:YAP

[MKB^{+}99] Peter J. Mohler, Silvia M. Kreda, Richard C. Boucher, Marius Sudol, M. Jackson Stutts, and Sharon L. Milgram. Yes-associated Protein 65 Localizes P62^{c−−Y es} to the Apical Compartment of Airway Epithelia by Association with Ebp50. Journal of Cell Biology, 147(4):879–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
4/879.

Manders:1999:DID

[MKC99] Erik M. M. Manders, Hiroshi Kimura, and Peter R. Cook. Direct Imaging of DNA in Living Cells Reveals the Dynamics of Chromosome Formation. Journal of Cell Biology, 144 (5):813–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/5/813.

Mertens:1996:PCP

[MKF96] C. Mertens, C. Kuhn, and W. W. Franke. Plakophilins 2a and 2b: constitutive proteins of dual location in the karyoplasm and the desmosomal plaque. Journal of Cell Biology, 135(4): 1009–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/1009.

Muto:1996:CWA

[MKI^{+}96] A. Muto, S. Kume, T. Inoue, H. Okano, and K. Mikoshiba. Calcium waves along the cleavage furrows in cleavage-stage Xenopus embryos and its inhibition by heparin. Journal of Cell Biology, 135(1):181–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/1/181.

Marshall:1995:PPP

[MKL^{+}95] P. A. Marshall, Y. I. Krimkevich, R. H. Lark, J. M. Dyer, M. Veenhuis, and J. M. Goodman. Pmp27 promotes peroxisomal proliferation. Journal of Cell Biology, 129(2):345–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
2/345.

Ma:1996:TCC

[MKLS96] Y. Ma, A. Kouranov, S. E. LaSala, and D. J. Schnell. Two components of the chloroplast protein import apparatus, IAP86 and IAP75, interact with the transit sequence during the recognition and translocation of precursor proteins at the outer envelope. Journal of Cell Biology, 134(2):315–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/315.

Meyer:1997:CCC

[MKS^{+}97] Kay N. Meyer, Eigil Kjeldsen, Tobias Straub, Birgitta R. Knudsen, Ian D. Hickson, Akihiko Kikuchi, Hans Kreipe, and Fritz Boege. Cell Cycle–coupled Relocation of Types I and II Topoisomerases and Modulation of Catalytic Enzyme Activities. Journal of Cell Biology, 136(4):775–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
4/775.

Myster:1999:DDH

[MKW^{+}99] Steven H. Myster, Julie A. Knott, Katrina M. Wysocki, Eileen O’Toole, and Mary E. Porter. Domains in the 1α Dynein Heavy Chain Required for Inner Arm Assembly and Flagellar Motility in Chlamydomonas. Journal of Cell Biology, 146 (4):801–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/801.

Muallem:1995:AFD

[MKXY95] S. Muallem, K. Kwiatkowska, X. Xu, and H. L. Yin. Actin filament disassembly is a sufficient final trigger for exocytosis in nonexcitable cells. Journal of Cell Biology, 128(4): 589–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/589.

Mayer:1997:SGE

[ML97] D. C. Ghislaine Mayer and Leslie A. Leinwand. Sarcomeric Gene Expression and Contractility in Myofibroblasts. Journal of Cell Biology, 139(6):1477–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/6/1477.

Matsuoka:1998:AVS

[MLB98] Yoichiro Matsuoka, Xiaolin Li, and Vann Bennett. Adducin Is an In Vivo Substrate for Protein Kinase C: Phosphorylation in the MARCKS-related Domain Inhibits Activity in Promoting Spectrin–Actin Complexes and Occurs in Many Cells, Including Dendritic Spines of Neurons. Journal of Cell Biology, 142 (2):485–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/485.

Montanaro:1999:DLR

[MLC99] Federica Montanaro, Michael Lindenbaum, and Salvatore Carbonetto. α-dystroglycan Is a Laminin Receptor Involved in Extracellular Matrix Assembly on Myotubes and Muscle Cell Viability. Journal of Cell Biology, 145(6):1325–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
6/1325.

Marsh:1995:OTR

[MLJM95] E. W. Marsh, P. L. Leopold, N. L. Jones, and F. R. Maxfield. Oligomerized transferrin receptors are selectively retained by a lumenal sorting signal in a long-lived endocytic recycling compartment. Journal of Cell Biology, 129(6):1509–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1509.

Mitsiadis:1995:ENR

[MLLT95] T. A. Mitsiadis, M. Lardelli, U. Lendahl, and I. Thesleff. Expression of Notch 1, 2 and 3 is regulated by epithelialmesenchymal interactions and retinoic acid in the developing mouse tooth and associated with determination of ameloblast cell fate. Journal of Cell Biology, 130(2):407–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/130/2/
407.

Masson:1998:VES

[MLN^{+}98] Régis Masson, Olivier Lefebvre, Agnès Noël, Mostapha El Fahime, Marie-Pierre Chenard, Corinne Wendling, Florence Kebers, Marianne LeMeur, Andrée Dierich, Jean-Michel Foidart, Paul Basset, and Marie-Christine Rio. In Vivo Evidence That the Stromelysin-3 Metalloproteinase Contributes in a Paracrine Manner to Epithelial Cell Malignancy. Journal of Cell Biology, 140(6):1535–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/6/1535.

Motley:1995:MAG

[MLO^{+}95] A. Motley, M. J. Lumb, P. B. Oatey, P. R. Jennings, P. A. De Zoysa, R. J. Wanders, H. F. Tabak, and C. J. Danpure. Mammalian alanine/glyoxylate aminotransferase 1 is imported into peroxisomes via the PTS1 translocation pathway. Increased degeneracy and context specificity of the mammalian PTS1 motif and implications for the peroxisome-to-mitochondrion mistargeting of AGT in primary hyperoxaluria type 1. Journal of Cell Biology, 131(1):95–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/1/95.

Mark:1995:SNO

[MLW^{+}95] M. D. Mark, Y. Liu, S. T. Wong, T. R. Hinds, and D. R. Storm. Stimulation of neurite outgrowth in PC12 cells by EGF and KCl depolarization: a Ca(2+)-independent phenomenon. Journal of Cell Biology, 130(3):701–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/3/701.

Lau:1999:IFD

[mLZB99] Pak ming Lau, Robert S. Zucker, and David Bentley. Induction of Filopodia by Direct Local Elevation of Intracellular Calcium Ion Concentration. Journal of Cell Biology, 145(6):1265–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/6/1265.

Mermall:1995:UMR

[MM95a] V. Mermall and K. G. Miller. The 95F unconventional myosin is required for proper organization of the Drosophila syncytial blastoderm. Journal of Cell Biology, 129(6):1575–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1575.

Miyake:1995:VAE

[MM95b] K. Miyake and P. L. McNeil. Vesicle accumulation and exocytosis at sites of plasma membrane disruption. Journal of Cell Biology, 131(6):1737–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1737.

Murphy:1996:CTS

[MM96] A. M. Murphy and D. J. Montell. Cell type-specific roles for Cdc42, Rac, and RhoL in Drosophila oogenesis. Journal of Cell Biology, 133(3):617–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/3/617.

Maceyka:1997:CAU

[MM97a] Michael Maceyka and Carolyn E. Machamer. Ceramide Accumulation Uncovers a Cycling Pathway for the cis -Golgi Network Marker, Infectious Bronchitis Virus M Protein. Journal of Cell Biology, 139(6):1411–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/6/1411.

Miller:1997:ASA

[MM97b] Jeffrey R. Miller and Randall T. Moon. Analysis of the Signaling Activities of Localization Mutants of β-Catenin during Axis Specification in Xenopus. Journal of Cell Biology, 139 (1):229–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/229.

Mallavarapu:1999:RAC

[MM99a] Aneil Mallavarapu and Tim Mitchison. Regulated actin cytoskeleton assembly at filopodium tips controls their extension and retraction. Journal of Cell Biology, 146(5):1097– ??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/1097.

Mallet:1999:CFF

[MM99b] William G. Mallet and Frederick R. Maxfield. Chimeric forms of furin and Tgn38 are transported from the plasma membrane to the trans-Golgi network via distinct endosomal pathways. Journal of Cell Biology, 146(2):345–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/2/345.

Mariani:1997:ICE

[MMAK97] Sara M. Mariani, Bernd Matiba, Elena A. Armandola, and Peter H. Krammer. Interleukin 1β-converting enzyme related proteases/caspases are involved in TRAIL-induced apoptosis of myeloma and leukemia cells. Journal of Cell Biology, 137 (1):221–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/1/221.

Matsui:1996:YSH

[MMATe96] Y. Matsui, R. Matsui, R. Akada, and A. Toh-e. Yeast src homology region 3 domain-binding proteins involved in bud formation. Journal of Cell Biology, 133(4):865–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
4/865.

Matsui:1998:RKP

[MMD^{+}98] Takeshi Matsui, Masato Maeda, Yoshinori Doi, Shigenobu Yonemura, Mutsuki Amano, Kozo Kaibuchi, Sachiko Tsukita, and Shoichiro Tsukita. Rho-kinase phosphorylates COOHterminal threonines of ezrin/radixin/moesin (ERM) proteins and regulates their head-to-tail association. Journal of Cell Biology, 140(3):647–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/647.

Meier:1995:NPC

[MMF95] E. Meier, B. R. Miller, and D. J. Forbes. Nuclear pore complex assembly studied with a biochemical assay for annulate lamellae formation. Journal of Cell Biology, 129(6):1459–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1459.

Matthies:1996:AMS

[MMGT96] H. J. Matthies, H. B. McDonald, L. S. Goldstein, and W. E. Theurkauf. Anastral meiotic spindle morphogenesis: role of the non-claret disjunctional kinesin-like protein. Journal of Cell Biology, 134(2):455–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/455.

Miller:1997:BDF

[MMK^{+}97] Timothy M. Miller, Krista L. Moulder, C. Michael Knudson, Douglas J. Creedon, Mohanish Deshmukh, Stanley J. Korsmeyer, and Eugene M. Johnson. Bax Deletion Further Orders the Cell Death Pathway in Cerebellar Granule Cells and Suggests a Caspase-independent Pathway to Cell Death. Journal of Cell Biology, 139(1):205–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/1/205.

Marr:1996:TDC

[MML96] K. M. Marr, D. N. Mastronarde, and M. K. Lyon. Twodimensional crystals of photosystem II: biochemical characterization, cryoelectron microscopy and localization of the D1 and cytochrome b559 polypeptides. Journal of Cell Biology, 132 (5):823–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/823.

Meier:1998:ACM

[MMM^{+}98] Thomas Meier, Fabrizio Masciulli, Chris Moore, Fabrice Schoumacher, Urs Eppenberger, Alain J. Denzer, Graham Jones, and Hans Rudolf Brenner. Agrin Can Mediate Acetylcholine Receptor Gene Expression in Muscle by Aggregation of Muscle-derived Neuregulins. Journal of Cell Biology, 141 (3):715–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/3/715.

Mitsiadis:1995:MMH

[MMMT95] T. A. Mitsiadis, T. Muramatsu, H. Muramatsu, and I. Thesleff. Midkine (MK), a heparin-binding growth/ differentiation factor, is regulated by retinoic acid and epithelial-mesenchymal interactions in the developing mouse tooth, and affects cell proliferation and morphogenesis. Journal of Cell Biology, 129(1):267–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/1/267.

Moreau:1996:SCA

[MMMW96] V. Moreau, A. Madania, R. P. Martin, and B. Winson. The Saccharomyces cerevisiae actin-related protein Arp2 is involved in the actin cytoskeleton. Journal of Cell Biology, 134 (1):117–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/1/117.

Matthies:1999:MTG

[MMN^{+}99] Heinrich J. G. Matthies, Lisa G. Messina, Ruria Namba, Kimberly J. Greer, M. Y. Walker, and R. Scott Hawley. Mutations in the α-tubulin 67C gene specifically impair achiasmate segregation in Drosophila melanogaster. Journal of Cell Biology, 147(6):1137–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/6/1137.

Marchand:1995:ABM

[MMP^{+}95] J. B. Marchand, P. Moreau, A. Paoletti, P. Cossart, M. F. Carlier, and D. Pantaloni. Actin-based movement of Listeria monocytogenes: actin assembly results from the local maintenance of uncapped filament barbed ends at the bacterium surface. Journal of Cell Biology, 130(2):331–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/331.

Miller:1999:CLM

[MMR99a] Rita K. Miller, Dina Matheos, and Mark D. Rose. The Cortical Localization of the Microtubule Orientation Protein, Kar9p, Is Dependent upon Actin and Proteins Required for Polarization. Journal of Cell Biology, 144(5):963–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/5/963.

Moore:1999:HPI

[MMR99b] Landon L. Moore, Mike Morrison, and Mark B. Roth. Hcp-1, a Protein Involved in Chromosome Segregation, Is Localized to the Centromere of Mitotic Chromosomes in Caenorhabditis elegans. Journal of Cell Biology, 147(3):471–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
3/471.

Morasso:1996:RED

[MMS96a] M. I. Morasso, N. G. Markova, and T. D. Sargent. Regulation of epidermal differentiation by a Distal-less homeodomain gene. Journal of Cell Biology, 135(6):1879–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
6/1879.

Moy:1996:SUS

[MMS^{+}96b] G. W. Moy, L. M. Mendoza, J. R. Schulz, W. J. Swanson, C. G. Glabe, and V. D. Vacquier. The sea urchin sperm receptor for egg jelly is a modular protein with extensive homology to the human polycystic kidney disease protein, PKD1. Journal of Cell Biology, 133(4):809–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/4/809.

Maeda:1998:IRL

[MN98] Nobuaki Maeda and Masaharu Noda. Involvement of receptorlike protein tyrosine phosphatase ζ /RPTPβ and its ligand pleiotrophin/heparin-binding growth-associated molecule (HB-GAM) in neuronal migration. Journal of Cell Biology, 142(1):203–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/1/203.

Moore:1998:TMC

[MNG^{+}98] R. A. Moore, H. Nguyen, J. Galceran, I. N. Pessah, and P. D. Allen. A Transgenic Myogenic Cell Line Lacking Ryanodine Receptor Protein for Homologous Expression Studies: Reconstitution of Ry_{1} R Protein and Function. Journal of Cell Biology, 140(4):843–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/4/843.

Mbalaviele:1998:CMH

[MNM^{+}98] Gabriel Mbalaviele, Riko Nishimura, Akira Myoi, Maria Niewolna, Sakamuri V. Reddy, Di Chen, Jian Feng, David Roodman, Gregory R. Mundy, and Toshiyuki Yoneda. Cadherin-6 Mediates the Heterotypic Interactions between the Hemopoietic Osteoclast Cell Lineage and Stromal Cells in a Murine Model of Osteoclast Differentiation. Journal of Cell Biology, 141(6):1467–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/6/1467.

Mancini:1998:CPC

[MNR^{+}98] Marie Mancini, Donald W. Nicholson, Sophie Roy, Nancy A. Thornberry, Erin P. Peterson, Livia A. Casciola-Rosen, and Antony Rosen. The Caspase-3 Precursor Has a Cytosolic and Mitochondrial Distribution: Implications for Apoptotic Signaling. Journal of Cell Biology, 140(6):1485–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1485.

Mandai:1997:ANA

[MNS^{+}97] Kenji Mandai, Hiroyuki Nakanishi, Ayako Satoh, Hiroshi Obaishi, Manabu Wada, Hideo Nishioka, Masahiko Itoh, Akira Mizoguchi, Takeo Aoki, Toyoshi Fujimoto, Yoichi Matsuda, Shoichiro Tsukita, and Yoshimi Takai. Afadin: a Novel Actin Filament–binding Protein with One PDZ Domain Localized at Cadherin-based Cell–to–Cell Adherens Junction. Journal of Cell Biology, 139(2):517–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/2/517.

Mandai:1999:PSA

[MNS^{+}99] Kenji Mandai, Hiroyuki Nakanishi, Ayako Satoh, Kenichi Takahashi, Keiko Satoh, Hideo Nishioka, Akira Mizoguchi, and Yoshimi Takai. Ponsin/SH3P12: An l-afadin- and vinculinbinding protein localized at cell–cell and cell–matrix adherens junctions. Journal of Cell Biology, 144(5):1001–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
5/1001.

Mesaeli:1999:CEC

[MNZ^{+}99] Nasrin Mesaeli, Kimitoshi Nakamura, Elena Zvaritch, Peter Dickie, Ewa Dziak, Karl-Heinz Krause, Michal Opas, David H. MacLennan, and Marek Michalak. Calreticulin Is Essential for Cardiac Development. Journal of Cell Biology, 144(5):857–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
144/5/857.

Matsumura:1998:SLS

[MOY^{+}98] Fumio Matsumura, Shoichiro Ono, Yoshihiko Yamakita, Go Totsukawa, and Shigeko Yamashiro. Specific Localization of Serine 19 Phosphorylated Myosin II during Cell Locomotion and Mitosis of Cultured Cells. Journal of Cell Biology, 140(1): 119–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/119.

Meldolesi:1998:HCS

[MP98] Jacopo Meldolesi and Tullio Pozzan. The Heterogeneity of ER Ca^{2+} Stores Has a Key Role in Nonmuscle Cell Signaling and Function. Journal of Cell Biology, 142(6):1395–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1395.

Moore:1995:PSG

[MPB^{+}95] K. L. Moore, K. D. Patel, R. E. Bruehl, F. Li, D. A. Johnson, H. S. Lichenstein, R. D. Cummings, D. F. Bainton, and R. P. McEver. P-selectin glycoprotein ligand-1 mediates rolling of human neutrophils on P-selectin. Journal of Cell Biology, 128 (4):661–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/661.

McGough:1997:CCT

[MPCW97] Amy McGough, Brian Pope, Wah Chiu, and Alan Weeds. Cofilin Changes the Twist of F–Actin: Implications for Actin Filament Dynamics and Cellular Function. Journal of Cell Biology, 138(4):771–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/4/771.

Miner:1997:LCE

[MPL^{+}97] Jeffrey H. Miner, Bruce L. Patton, Stephen I. Lentz, Debra J. Gilbert, William D. Snider, Nancy A. Jenkins, Neal G. Copeland, and Joshua R. Sanes. The Laminin α Chains: Expression, Developmental Transitions, and Chromosomal Locations of α1-5, Identification of Heterotrimeric Laminins 8–11, and Cloning of a Novel α3 Isoform. Journal of Cell Biology, 137(3):685–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/3/685.

Martin-Padura:1998:JAM

[MPLS^{+}98] Inés Martìn-Padura, Susan Lostaglio, Markus Schneemann, Lisa Williams, Maria Romano, Paolo Fruscella, Carla Panzeri, Antonella Stoppacciaro, Luigi Ruco, Antonello Villa, David Simmons, and Elisabetta Dejana. Junctional Adhesion Molecule, a Novel Member of the Immunoglobulin Superfamily That Distributes at Intercellular Junctions and Modulates Monocyte Transmigration. Journal of Cell Biology, 142 (1):117–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/1/117.

McNagny:1997:TNC

[MPR^{+}97] Kelly M. McNagny, Inger Pettersson, Fabio Rossi, Ingo Flamme, Andrej Shevchenko, Matthias Mann, and Thomas Graf. Thrombomucin, a Novel Cell Surface Protein that Defines Thrombocytes and Multipotent Hematopoietic Progenitors. Journal of Cell Biology, 138(6):1395–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
6/1395.

Markovic:1998:MFM

[MPSC98] Ingrid Markovic, Helena Pulyaeva, Alexander Sokoloff, and Leonid V. Chernomordik. Membrane Fusion Mediated by Baculovirus gp64 Involves Assembly of Stable gp64 Trimers into Multiprotein Aggregates. Journal of Cell Biology, 143(5): 1155–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/5/1155.

Moore:1998:CPM

[MPT^{+}98] Daniel P. Moore, Andrea W. Page, Tracy Tzu-Ling Tang, Anne W. Kerrebrock, and Terry L. Orr-Weaver. The Cohesion Protein MEI–S332 Localizes to Condensed Meiotic and Mitotic Centromeres until Sister Chromatids Separate. Journal of Cell Biology, 140(5):1003–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/5/1003.

Meier:1995:INA

[MPW95] T. Meier, G. M. Perez, and B. G. Wallace. Immobilization of nicotinic acetylcholine receptors in mouse C2 myotubes by agrin-induced protein tyrosine phosphorylation. Journal of Cell Biology, 131(2):441–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/2/441.

Mainiero:1996:IFA

[MPY^{+}96] F. Mainiero, A. Pepe, M. Yeon, Y. Ren, and F. G. Giancotti. The intracellular functions of alpha6beta4 integrin are regulated by EGF. Journal of Cell Biology, 134(1):241–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
1/241.

Morfini:1997:SKP

[MQR^{+}97] Gerardo Morfini, Santiago Quiroga, Alberto Rosa, Kenneth Kosik, and Alfredo Cáceres. Suppression of KIF2 in PC12 Cells Alters the Distribution of a Growth Cone Nonsynaptic Membrane Receptor and Inhibits Neurite Extension. Journal of Cell Biology, 138(3):657–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/3/657.

Miesenbock:1995:CRE

[MR95] G. Miesenböck and J. E. Rothman. The capacity to retrieve escaped ER proteins extends to the trans-most cisterna of the Golgi stack. Journal of Cell Biology, 129(2):309–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
2/309.

Mondesert:1996:BGE

[MR96] G. Mondésert and S. I. Reed. BED1, a gene encoding a galactosyltransferase homologue, is required for polarized growth and efficient bud emergence in Saccharomyces cerevisiae. Journal of Cell Biology, 132(1):137–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/1/137.

Miller:1998:KNC

[MR98] Rita K. Miller and Mark D. Rose. Kar9p Is a Novel Cortical Protein Required for Cytoplasmic Microtubule Orientation in Yeast. Journal of Cell Biology, 140(2):377–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
2/377.

Miller:1999:EDV

[MRL^{+}99] Jeffrey R. Miller, Brian A. Rowning, Carolyn A. Larabell, Julia A. Yang-Snyder, Rebecca L. Bates, and Randall T. Moon. Establishment of the Dorsal–Ventral Axis in Xenopus Embryos Coincides with the Dorsal Enrichment of Dishevelled That Is Dependent on Cortical Rotation. Journal of Cell Biology, 146 (2):427–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/2/427.

Martini:1998:RPC

[MRM^{+}98] Emmanuelle Martini, Danièle M. J. Roche, Kathrin Marheineke, Alain Verreault, and Geneviève Almouzni. Recruitment of Phosphorylated Chromatin Assembly Factor 1 to Chromatin after UV Irradiation of Human Cells. Journal of Cell Biology, 143(3):563–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/3/563.

Marks:1995:LTS

[MRvD^{+}95] M. S. Marks, P. A. Roche, E. van Donselaar, L. Woodruff, P. J. Peters, and J. S. Bonifacino. A lysosomal targeting signal in the cytoplasmic tail of the beta chain directs HLA–DM to MHC class II compartments. Journal of Cell Biology, 131(2): 351–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/2/351.

Miner:1996:MFD

[MS96] J. H. Miner and J. R. Sanes. Molecular and functional defects in kidneys of mice lacking collagen alpha 3(IV): implications for Alport syndrome. Journal of Cell Biology, 135(5): 1403–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/5/1403.

McNeil:1997:LRM

[MS97a] Paul L. McNeil and Richard A. Steinhardt. Loss, Restoration, and Maintenance of Plasma Membrane Integrity. Journal of Cell Biology, 137(1):1–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/1/1.

Morris:1997:HKI

[MS97b] Robert L. Morris and Jonathan M. Scholey. Heterotrimeric Kinesin–II Is Required for the Assembly of Motile 9+2 Ciliary Axonemes on Sea Urchin Embryos. Journal of Cell Biology, 138(5):1009–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/5/1009.

May:1998:PCD

[MS98] Timo May and Jürgen Soll. Positive Charges Determine the Topology and Functionality of the Transmembrane Domain in the Chloroplastic Outer Envelope Protein Toc34. Journal of Cell Biology, 141(4):895–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/895.

Mitchell:1999:CCI

[MS99] David R. Mitchell and Winfield S. Sale. Characterization of a Chlamydomonas Insertional Mutant that Disrupts Flagellar Central Pair Microtubule-associated Structures. Journal of Cell Biology, 144(2):293–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/2/293.

Machado:1998:HAR

[MSA98] Cristina Machado, Claudio E. Sunkel, and Deborah J. Andrew. Human Autoantibodies Reveal Titin as a Chromosomal Protein. Journal of Cell Biology, 141(2):321–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
2/321.

McGill:1997:LMA

[MSB^{+}97] Gaël McGill, Akiko Shimamura, Richard C. Bates, Robert E. Savage, and David E. Fisher. Loss of Matrix Adhesion Triggers Rapid Transformation–Selective Apoptosis in Fibroblasts. Journal of Cell Biology, 138(4):901–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/4/901.

Ma:1999:ACT

[MSB99] Hong Ma, Alan J. Siegel, and Ronald Berezney. Association of Chromosome Territories with the Nuclear Matrix. Journal of Cell Biology, 146(3):531–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/3/531.

Ma:1998:STD

[MSD^{+}98] Hong Ma, Jagath Samarabandu, Rekandu S. Devdhar, Raj Acharya, Ping chin Cheng, Chunling Meng, and Ronald Berezney. Spatial and Temporal Dynamics of DNA Replication Sites in Mammalian Cells. Journal of Cell Biology, 143(6): 1415–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1415.

Mills:1998:AMB

[MSEP98] Jason C. Mills, Nicole L. Stone, Joseph Erhardt, and Randall N. Pittman. Apoptotic Membrane Blebbing Is Regulated by Myosin Light Chain Phosphorylation. Journal of Cell Biology, 140(3):627–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/627.

Mays:1995:HMI

[MSF^{+}95] R. W. Mays, K. A. Siemers, B. A. Fritz, A. W. Lowe, G. van Meer, and W. J. Nelson. Hierarchy of mechanisms involved in generating Na/K–ATPase polarity in MDCK epithelial cells. Journal of Cell Biology, 130(5):1105–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
5/1105.

Miao:1999:NMC

[MSF^{+}99a] Hua-Quan Miao, Shay Soker, Leonard Feiner, José Luis Alonso, Jonathan A. Raper, and Michael Klagsbrun. Neuropilin-1 mediates collapsin-1/semaphorin III inhibition of endothelial cell motility. Journal of Cell Biology, 146(1):233–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
1/233.

Morita:1999:COB

[MSF^{+}99b] Kazumasa Morita, Hiroyuki Sasaki, Kazushi Fujimoto, Mikio Furuse, and Shoichiro Tsukita. Claudin-11/ OSP-based Tight Junctions of Myelin Sheaths in Brain and Sertoli Cells in Testis. Journal of Cell Biology, 145(3):579–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/145/3/
579.

Mitic:1999:COC

[MSFA99] Laura L. Mitic, Eveline E. Schneeberger, Alan S. Fanning, and James Melvin Anderson. Connexin-occludin Chimeras Containing the Zo–Binding Domain of Occludin Localize at Mdck Tight Junctions and Nrk Cell Contacts. Journal of Cell Biology, 146(3):683–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/3/683.

Morita:1999:EC

[MSFT99] Kazumasa Morita, Hiroyuki Sasaki, Mikio Furuse, and Shoichiro Tsukita. Endothelial Claudin. Journal of Cell Biology, 147(1):185–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/1/185.

Michiels:1997:RML

[MSH^{+}97] Frits Michiels, Jord C. Stam, Peter L. Hordijk, Rob A. van der Kammen, Lisette Ruuls-Van Stalle, Constance A. Feltkamp, and John G. Collard. Regulated membrane localization of Tiam1, mediated by the NH_{2}-terminal pleckstrin homology domain, is required for Rac-dependent membrane ruffling and C–Jun NH_{2}-terminal kinase activation. Journal of Cell Biology, 137(2):387–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/2/387.

Malliri:1998:TFA

[MSH^{+}98] Angeliki Malliri, Marc Symons, Robert F. Hennigan, Adam F. L. Hurlstone, Richard F. Lamb, Tricia Wheeler, and Bradford W. Ozanne. The Transcription Factor AP-1 Is Required for EGF-induced Activation of Rho-like GTPases, Cytoskeletal Rearrangements, Motility, and In Vitro Invasion of A431 Cells. Journal of Cell Biology, 143(4):1087–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/1087.

Mountain:1999:KRP

[MSH^{+}99] Vicki Mountain, Calvin Simerly, Louisa Howard, Asako Ando, Gerald Schatten, and Duane A. Compton. The Kinesin– Related Protein, Hset, Opposes the Activity of Eg5 and Cross– Links Microtubules in the Mammalian Mitotic Spindle. Journal of Cell Biology, 147(2):351–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/2/351.

Magin:1998:LKK

[MSL^{+}98a] Thomas M. Magin, Rolf Schröder, Sabine Leitgeb, Frederique Wanninger, Kurt Zatloukal, Christine Grund, and David W. Melton. Lessons from Keratin 18 Knockout Mice: Formation of Novel Keratin Filaments, Secondary Loss of Keratin 7 and Accumulation of Liver-specific Keratin 8-Positive Aggregates. Journal of Cell Biology, 140(6):1441–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/6/1441.

McMillan:1998:MCM

[MSL98b] John N. McMillan, Rey A. L. Sia, and Daniel J. Lew. A Morphogenesis Checkpoint Monitors the Actin Cytoskeleton in Yeast. Journal of Cell Biology, 142(6):1487–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1487.

Mukherjee:1999:ESL

[MSM99] Sushmita Mukherjee, Thwe Thwe Soe, and Frederick R. Maxfield. Endocytic Sorting of Lipid Analogues Differing Solely in the Chemistry of Their Hydrophobic Tails. Journal of Cell Biology, 144(6):1271–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/6/1271.

Mori:1999:CST

[MSMC99] Hiroki Mori, Elizabeth J. Summer, Xianyue Ma, and Kenneth Cline. Component specificity for the thylakoidal sec and delta Ph-dependent protein transport pathways. Journal of Cell Biology, 146(1):45–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/1/45.

Mullins:1997:SST

[MSP97] R. Dyche Mullins, Walter F. Stafford, and Thomas D. Pollard. Structure, subunit topology, and actin-binding activity of the Arp2/3 complex from Acanthamoeba. Journal of Cell Biology, 136(2):331–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/2/331.

Mills:1999:EA

[MSP99] Jason C. Mills, Nicole L. Stone, and Randall N. Pittman. Extranuclear Apoptosis. Journal of Cell Biology, 146(4): 703–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/703.

Marg:1999:NNN

[MSS^{+}99] Andreas Marg, Pinar Sirim, Frank Spaltmann, Antonius Plagge, Gunther Kauselmann, Friedrich Buck, Fritz G. Rathjen, and Thomas Brümmendorf. Neurotractin, A Novel Neurite Outgrowth-promoting Ig-like Protein that Interacts with CEPU-1 and LAMP. Journal of Cell Biology, 145(4):865–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/4/865.

Masumoto:1996:IRC

[MST^{+}96] N. Masumoto, T. Sasaki, M. Tahara, A. Mammoto, Y. Ikebuchi, K. Tasaka, M. Tokunaga, Y. Takai, and A. Miyake. Involvement of Rabphilin-3A in cortical granule exocytosis in mouse eggs. Journal of Cell Biology, 135(6):1741–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/6/1741.

Majoul:1998:KRE

[MSW^{+}98] Irina Majoul, Kai Sohn, Felix Theodor Wieland, Rainer Pepperkok, Mariagrazia Pizza, Jörg Hillemann, and Hans-Dieter Söling. KDEL Receptor (Erd2p)-mediated Retrograde Transport of the Cholera Toxin A Subunit from the Golgi Involves COPI, p23, and the COOH Terminus of Erd2p. Journal of Cell Biology, 143(3):601–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/3/601.

Millward-Sadler:1999:IRS

[MSWL^{+}99] S. J. Millward-Sadler, M. O. Wright, H.-S. Lee, K. Nishida, H. Caldwell, G. Nuki, and D. M. Salter. Integrin-regulated Secretion of Interleukin 4: a Novel Pathway of Mechanotransduction in Human Articular Chondrocytes. Journal of Cell Biology, 145(1):183–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/1/183.

Miyamoto:1995:IFM

[MTC^{+}95] S. Miyamoto, H. Teramoto, O. A. Coso, J. S. Gutkind, P. D. Burbelo, S. K. Akiyama, and K. M. Yamada. Integrin function: molecular hierarchies of cytoskeletal and signaling molecules. Journal of Cell Biology, 131(3):791–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/791.

Miyamoto:1996:ICC

[MTGY96] S. Miyamoto, H. Teramoto, J. S. Gutkind, and K. M. Yamada. Integrins can collaborate with growth factors for phosphorylation of receptor tyrosine kinases and MAP kinase activation: roles of integrin aggregation and occupancy of receptors. Journal of Cell Biology, 135(6):1633–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
6/1633.

Molloy:1998:RES

[MTK^{+}98] Sean S. Molloy, Laurel Thomas, Craig Kamibayashi, Marc C. Mumby, and Gary Thomas. Regulation of Endosome Sorting by a Specific PP2A Isoform. Journal of Cell Biology, 142(6): 1399–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/6/1399.

Martin:1996:GTG

[MTL^{+}96] S. Martin, J. Tellam, C. Livingstone, J. W. Slot, G. W. Gould, and D. E. James. The glucose transporter (GLUT-4) and vesicle-associated membrane protein-2 (VAMP-2) are segregated from recycling endosomes in insulin-sensitive cells. Journal of Cell Biology, 134(3):625–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/3/625.

Ma:1999:DIC

[MTLGC99] Shuo Ma, Leda Triviños-Lagos, Ralph Gräf, and Rex L. Chisholm. Dynein Intermediate Chain Mediated Dynein– Dynactin Interaction Is Required for Interphase Microtubule Organization and Centrosome Replication and Separation in Dictyostelium. Journal of Cell Biology, 147(6):1261–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/6/1261.

Miller:1995:BSS

[MTLW95] J. D. Miller, S. Tajima, L. Lauffer, and P. Walter. The beta subunit of the signal recognition particle receptor is a transmembrane GTPase that anchors the alpha subunit, a peripheral membrane GTPase, to the endoplasmic reticulum membrane. Journal of Cell Biology, 128(3):273–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
3/273.

Meeusen:1999:MNC

[MTW^{+}99] Shelly Meeusen, Quinton Tieu, Edith Wong, Eric Weiss, David Schieltz, John R. Yates, and Jodi Nunnari. Mgm101p Is a Novel Component of the Mitochondrial Nucleoid That Binds DNA and Is Required for the Repair of Oxidatively Damaged Mitochondrial DNA. Journal of Cell Biology, 145(2):291–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/291.

Murphy:1998:MTC

[MUS98] Steven M. Murphy, Lenore Urbani, and Tim Stearns. The Mammalian γ-Tubulin Complex Contains Homologues of the Yeast Spindle Pole Body Components Spc97p and Spc98p. Journal of Cell Biology, 141(3):663–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/3/663.

Merlin:1999:JRS

[MVM^{+}99] Alessio Merlin, Wolfgang Voos, Ammy C. Maarse, Michiel Meijer, Nikolaus Pfanner, and Joachim Rassow. The J-related Segment of Tim44 Is Essential for Cell Viability: a Mutant Tim44 Remains in the Mitochondrial Import Site, but Inefficiently Recruits mtHsp70 and Impairs Protein Translocation. Journal of Cell Biology, 145(5):961–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/5/961.

Mertens:1996:HSE

[MVvdBD96] G. Mertens, B. Van der Schueren, H. van den Berghe, and G. David. Heparan sulfate expression in polarized epithelial cells: the apical sorting of glypican (GPI-anchored proteoglycan) is inversely related to its heparan sulfate content. Journal of Cell Biology, 132(3):487–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/3/487.

McIntosh:1995:CBP

[MW95a] J. R. McIntosh and R. R. West. A cell biological perspective on genome research. Journal of Cell Biology, 131(6):1361–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/6/1361.

Misteli:1995:RTN

[MW95b] T. Misteli and G. Warren. A role for tubular networks and a COP I-independent pathway in the mitotic fragmentation of Golgi stacks in a cell-free system. Journal of Cell Biology, 130(5):1027–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/5/1027.

Muller:1996:ABS

[MW96] H. A. Müller and E. Wieschaus. armadillo, bazooka, and stardust are critical for early stages in formation of the zonula adherens and maintenance of the polarized blastoderm epithelium in Drosophila. Journal of Cell Biology, 134(1):149–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/1/149.

Mayer:1997:DYV

[MW97] Andreas Mayer and William Wickner. Docking of Yeast Vacuoles Is Catalyzed by the Ras-like GTPase Ypt7p after Symmetric Priming by Sec18p (NSF). Journal of Cell Biology, 136 (2):307–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/2/307.

Matunis:1998:SMR

[MWB98] Michael J. Matunis, Jian Wu, and Günter Blobel. SUMO-1 Modification and Its Role in Targeting the Ran GTPaseactivating Protein, RanGAP1, to the Nuclear Pore Complex. Journal of Cell Biology, 140(3):499–??, February 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/140/3/
499.

Melikyan:1995:GAI

[MWC95] G. B. Melikyan, J. M. White, and F. S. Cohen. GPI-anchored influenza hemagglutinin induces hemifusion to both red blood cell and planar bilayer membranes. Journal of Cell Biology, 131(3):679–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/3/679.

Marszalek:1999:NDK

[MWF^{+}99] Joseph R. Marszalek, Joshua A. Weiner, Samuel J. Farlow, Jerold Chun, and Lawrence S. B. Goldstein. Novel Dendritic Kinesin Sorting Identified by Different Process Targeting of Two Related Kinesins: KIF21A and KIF21B. Journal of Cell Biology, 145(3):469–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/3/469.

McCarthy:1997:ICI

[MWGE97] Nicola J. McCarthy, Moira K. B. Whyte, Christopher S. Gilbert, and Gerard I. Evan. Inhibition of Ced-3/ICE-related proteases does not prevent cell death induced by oncogenes, DNA damage, or the Bcl-2 homologue bak. Journal of Cell Biology, 136(1):215–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/1/215.

Marszalek:1996:NSN

[MWL^{+}96] J. R. Marszalek, T. L. Williamson, M. K. Lee, Z. Xu, P. N. Hoffman, M. W. Becher, T. O. Crawford, and D. W. Cleveland. Neurofilament subunit NF–H modulates axonal diameter by selectively slowing neurofilament transport. Journal of Cell Biology, 135(3):711–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/3/711.

Mironov:1997:VIT

[MWL97] Alexander A. Mironov, Peggy Weidman, and Alberto Luini. Variations on the Intracellular Transport Theme: Maturing Cisternae and Trafficking Tubules. Journal of Cell Biology, 138 (3):481–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/3/481.

Marks:1996:PTT

[MWOB96] M. S. Marks, L. Woodruff, H. Ohno, and J. S. Bonifacino. Protein targeting by tyrosine- and di-leucine-based signals: evidence for distinct saturable components. Journal of Cell Biology, 135(2):341–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/2/341.

Milner:1996:DMA

[MWT^{+}96] D. J. Milner, G. Weitzer, D. Tran, A. Bradley, and Y. Capetanaki. Disruption of muscle architecture and myocardial degeneration in mice lacking desmin. Journal of Cell Biology, 134(5):1255–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/5/1255.

Musch:1996:TVS

[MXSRB96] A. Müsch, H. Xu, D. Shields, and E. Rodriguez-Boulan. Transport of vesicular stomatitis virus G protein to the cell surface is signal mediated in polarized and nonpolarized cells. Journal of Cell Biology, 133(3):543–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/3/543.

Mineo:1998:TPK

[MYC^{+}98] Chieko Mineo, Yun-Shu Ying, Christine Chapline, Susan Jaken, and Richard G. W. Anderson. Targeting of Protein Kinase Cα to Caveolae. Journal of Cell Biology, 141(3):601–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/3/601.

Moore:1999:NIC

[MYTK99] Jonathan D. Moore, Jing Yang, Ray Truant, and Sally Kornbluth. Nuclear import of Cdk/cyclin complexes: Identification of distinct mechanisms for import of Cdk2/Cyclin E and Cdc2/ Cyclin B1. Journal of Cell Biology, 144(2):213–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
2/213.

Moritz:1998:RTR

[MZAO98] Michelle Moritz, Yixian Zheng, Bruce M. Alberts, and Karen Oegema. Recruitment of the γ-Tubulin Ring Complex to Drosophila Salt-stripped Centrosome Scaffolds. Journal of Cell Biology, 142(3):775–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/3/775.

Muller:1997:MFP

[MZL97] Laurent Muller, Xiaorong Zhu, and Iris Lindberg. Mechanism of the Facilitation of PC2 Maturation by 7B2: Involvement in ProPC2 Transport and Activation but Not Folding. Journal of Cell Biology, 139(3):625–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/3/625.

Martin:1999:WBN

[MZL^{+}99] Doris Martin, Susan Zusman, Xitong Li, Erin L. Williams, Narmada Khare, Sol DaRocha, Ruth Chiquet-Ehrismann, and Stefan Baumgartner. wing blister, A New Drosophila Laminin α Chain Required for Cell Adhesion and Migration during Embryonic and Imaginal Development. Journal of Cell Biology, 145(1):191–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/1/191.

Murphy:1997:INS

[MZN97] Edwin C. Murphy, Tianli Zheng, and Christopher V. Nicchitta. Identification of a novel stage of ribosome/nascent chain association with the endoplasmic reticulum membrane. Journal of Cell Biology, 136(6):1213–??, March 1997. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/6/1213.

Matter:1998:IME

[MZNR98] Michelle L. Matter, Zhuohua Zhang, Christer Nordstedt, and Erkki Ruoslahti. The α5β1 Integrin Mediates Elimination of Amyloid-β Peptide and Protects Against Apoptosis. Journal of Cell Biology, 141(4):1019–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/1019.

Neveu:1996:NPS

[NA96] I. Neveu and E. Arenas. Neurotrophins promote the survival and development of neurons in the cerebellum of hypothyroid rats in vivo. Journal of Cell Biology, 133(3):631–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
3/631.

Nern:1999:CFG

[NA99] Aljoscha Nern and Robert A. Arkowitz. A Cdc24p–Far1p– Gβγ Protein Complex Required for Yeast Orientation during Mating. Journal of Cell Biology, 144(6):1187–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
6/1187.

Nathke:1996:APC

[NAP^{+}96] I. S. Näthke, C. L. Adams, P. Polakis, J. H. Sellin, and W. J. Nelson. The adenomatous polyposis coli tumor suppressor protein localizes to plasma membrane sites involved in active cell migration. Journal of Cell Biology, 134(1):165–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
1/165.

Nelson:1998:MTF

[NAsG^{+}98] David S. Nelson, Cecilia Alvarez, Ya sheng Gao, Rafael García-Mata, Elizabeth Fialkowski, and Elizabeth Sztul. The membrane transport factor TAP/p115 cycles between the Golgi and earlier secretory compartments and contains distinct domains required for its localization and function. Journal of Cell Biology, 143(2):319–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/319.

Niclas:1996:CCR

[NAV96] J. Niclas, V. J. Allan, and R. D. Vale. Cell cycle regulation of dynein association with membranes modulates microtubulebased organelle transport. Journal of Cell Biology, 133(3): 585–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/3/585.

Nix:1997:NCS

[NB97] David A. Nix and Mary C. Beckerle. Nuclear–cytoplasmic Shuttling of the Focal Contact Protein, Zyxin: a Potential Mechanism for Communication between Sites of Cell Adhesion and the Nucleus. Journal of Cell Biology, 138(5):1139– ??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/5/1139.

Neale:1999:BNB

[NBJ^{+}99] Elaine A. Neale, Linda M. Bowers, Min Jia, Karen E. Bateman, and Lura C. Williamson. Botulinum Neurotoxin a Blocks Synaptic Vesicle Exocytosis but Not Endocytosis at the Nerve Terminal. Journal of Cell Biology, 147(6):1249–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1249.

Niemann:1998:RMG

[NBS^{+}98] Catherin Niemann, Volker Brinkmann, Eva Spitzer, Guido Hartmann, Martin Sachs, Helga Naundorf, and Walter Birchmeier. Reconstitution of Mammary Gland Development In Vitro: Requirement of c-met and c-erbB2 Signaling for Branching and Alveolar Morphogenesis. Journal of Cell Biology, 143(2):533–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/533.

Ng:1996:SSS

[NBW96] D. T. Ng, J. D. Brown, and P. Walter. Signal sequences specify the targeting route to the endoplasmic reticulum membrane. Journal of Cell Biology, 134(2):269–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/2/269.

Nothwehr:1995:GVM

[NCS95] S. F. Nothwehr, E. Conibear, and T. H. Stevens. Golgi and vacuolar membrane proteins reach the vacuole in vps1 mutant yeast cells via the plasma membrane. Journal of Cell Biology, 129(1):35–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/1/35.

Nykjaer:1998:MPI

[NCV^{+}98] Anders Nykjær, Erik I. Christensen, Henrik Vorum, Henrik Hager, Claus M. Petersen, Hans Røigaard, Hye Y. Min, Frederik Vilhardt, Lisbeth B. Møller, Stuart Kornfeld, and Jørgen Gliemann. Mannose 6-phosphate/insulin-like growth factor-II receptor targets the urokinase receptor to lysosomes via a novel binding interaction. Journal of Cell Biology, 141(3):815–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/3/815.

Nguyen:1999:MLC

[NCW^{+}99] Diem H. D. Nguyen, Andrew D. Catling, Donna J. Webb, Mauricio Sankovic, Lori A. Walker, Avril V. Somlyo, Michael J. Weber, and Steven L. Gonias. Myosin light chain kinase functions downstream of Ras/ERK to promote migration of urokinase-type plasminogen activator-stimulated cells in an integrin–selective manner. Journal of Cell Biology, 146 (1):149–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/149.

Nakagawa:1995:TDF

[NCZ^{+}95] T. Nakagawa, J. Chen, Z. Zhang, Y. Kanai, and N. Hirokawa. Two distinct functions of the carboxyl-terminal tail domain of NF–M upon neurofilament assembly: cross-bridge formation and longitudinal elongation of filaments. Journal of Cell Biology, 129(2):411–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/2/411.

Nakielny:1996:HCP

[ND96] S. Nakielny and G. Dreyfuss. The hnRNP C proteins contain a nuclear retention sequence that can override nuclear export signals. Journal of Cell Biology, 134(6):1365–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
6/1365.

Naim:1995:ABC

[NDBR95] H. Y. Naim, D. T. Dodds, C. B. Brewer, and M. G. Roth. Apical and basolateral coated pits of MDCK cells differ in their rates of maturation into coated vesicles, but not in the ability to distinguish between mutant hemagglutinin proteins with different internalization signals. Journal of Cell Biology, 129 (5):1241–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/5/1241.

Neiman:1998:PMF

[Nei98] Aaron M. Neiman. Prospore Membrane Formation Defines a Developmentally Regulated Branch of the Secretory Pathway in Yeast. Journal of Cell Biology, 140(1):29–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/140/1/29.

Nelson:1999:CFA

[NG99] Richard W. Nelson and Barry M. Gumbiner. A Cell–Free Assay System for β-Catenin Signaling That Recapitulates Direct Inductive Events in the Early Xenopus laevis Embryo. Journal of Cell Biology, 147(2):367–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/2/367.

Nunes:1997:LTG

[NGMR97] Irene Nunes, Pierre-Emmanuel Gleizes, Christine N. Metz, and Daniel B. Rifkin. Latent Transforming Growth Factorβ Binding Protein Domains Involved in Activation and Transglutaminase-dependent Cross–Linking of Latent Transforming Growth Factor-β. Journal of Cell Biology, 136(5): 1151–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/5/1151.

Niewiadomska:1999:DCR

[NGT99] Paulina Niewiadomska, Dorothea Godt, and Ulrich Tepass. D E–Cadherin Is Required for Intercellular Motility during Drosophila Oogenesis. Journal of Cell Biology, 144(3):533– ??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/533.

Nakata:1995:PMA

[NH95] T. Nakata and N. Hirokawa. Point mutation of adenosine triphosphate-binding motif generated rigor kinesin that selectively blocks anterograde lysosome membrane transport. Journal of Cell Biology, 131(4):1039–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/4/1039.

Nobes:1999:RGC

[NH99] Catherine D. Nobes and Alan Hall. Rho GTPases Control Polarity, Protrusion, and Adhesion during Cell Movement. Journal of Cell Biology, 144(6):1235–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/6/1235.

Neujahr:1997:TDP

[NHA^{+}97] Ralph Neujahr, Christina Heizer, Richard Albrecht, Maria Ecke, Jean-Marc Schwartz, Igor Weber, and Günther Gerisch. Three-dimensional Patterns and Redistribution of Myosin II and Actin in Mitotic Dictyostelium Cells. Journal of Cell Biology, 139(7):1793–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/7/1793.

Norenberg:1995:CFD

[NHB^{+}95] U. Nörenberg, M. Hubert, T. Brümmendorf, A. Tárnok, and F. G. Rathjen. Characterization of functional domains of the tenascin–R (restrictin) polypeptide: cell attachment site, binding with F11, and enhancement of F11-mediated neurite outgrowth by tenascin–R. Journal of Cell Biology, 130(2):473–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/2/473.

Nishiyama:1997:GRR

[NIK^{+}97] Hiroyuki Nishiyama, Katsuhiko Itoh, Yoshiyuki Kaneko, Masamichi Kishishita, Osamu Yoshida, and Jun Fujita. A Glycine-rich RNA-binding Protein Mediating Cold-inducible Suppression of Mammalian Cell Growth. Journal of Cell Biology, 137(4):899–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/4/899.

Nakamura:1995:MAH

[NIM95] K. Nakamura, R. Iwamoto, and E. Mekada. Membraneanchored heparin-binding EGF-like growth factor (HB–EGF) and diphtheria toxin receptor-associated protein (DRAP27)/ CD9 form a complex with integrin alpha 3 beta 1 at cell–cell contact sites. Journal of Cell Biology, 129(6):1691–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
6/1691.

Norman:1998:AMP

[NJB^{+}98] J. C. Norman, D. Jones, S. T. Barry, M. R. Holt, S. Cockcroft, and D. R. Critchley. ARF1 Mediates Paxillin Recruitment to Focal Adhesions and Potentiates Rho-stimulated Stress Fiber Formation in Intact and Permeabilized Swiss 3T3 Fibroblasts. Journal of Cell Biology, 143(7):1981–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1981.

Ninomiya:1995:DET

[NKI^{+}95] Y. Ninomiya, M. Kagawa, K. Iyama, I. Naito, Y. Kishiro, J. M. Seyer, M. Sugimoto, T. Oohashi, and Y. Sado. Differential expression of two basement membrane collagen genes, COL4A6 and COL4A5, demonstrated by immunofluorescence staining using peptide-specific monoclonal antibodies. Journal of Cell Biology, 130(5):1219–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/5/1219.

Nishi:1999:AFS

[NKK^{+}99] Miyuki Nishi, Shinji Komazaki, Nagomi Kurebayashi, Yasuo Ogawa, Tetsuo Noda, Masamitsu Iino, and Hiroshi Takeshima.

Abnormal Features in Skeletal Muscle from Mice Lacking Mitsugumin29. Journal of Cell Biology, 147(7):1473–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1473.

Nobes:1998:NMR

[NLM^{+}98] Catherine D. Nobes, Inger Lauritzen, Marie-Geneviève Mattei, Sonia Paris, Alan Hall, and Pierre Chardin. A New Member of the Rho Family, Rnd1, Promotes Disassembly of Actin Filament Structures and Loss of Cell Adhesion. Journal of Cell Biology, 141(1):187–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/1/187.

Nakamura:1998:WON

[NMH^{+}98] Masafumi Nakamura, Hirohisa Masuda, Johji Horii, Kei ichi Kuma, Nobuhiko Yokoyama, Tomoyuki Ohba, Hideo Nishitani, Takashi Miyata, Masao Tanaka, and Takeharu Nishimoto. When Overexpressed, a Novel Centrosomal Protein, RanBPM, Causes Ectopic Microtubule Nucleation Similar to γ-Tubulin. Journal of Cell Biology, 143(4):1041–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/1041.

Nicchitta:1995:SRS

[NMHS95] C. V. Nicchitta, E. C. Murphy, R. Haynes, and G. S. Shelness. Stage- and ribosome-specific alterations in nascent chain–Sec61p interactions accompany translocation across the ER membrane. Journal of Cell Biology, 129(4):957–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
4/957.

Novitch:1996:SMC

[NMJL96] B. G. Novitch, G. J. Mulligan, T. Jacks, and A. B. Lassar. Skeletal muscle cells lacking the retinoblastoma protein display defects in muscle gene expression and accumulate in S and G2 phases of the cell cycle. Journal of Cell Biology, 135(2): 441–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/2/441.

Nakajima:1997:EFS

[NMK^{+}97] Yuji Nakajima, Kohei Miyazono, Mitsuyasu Kato, Masao Takase, Toshiyuki Yamagishi, and Hiroaki Nakamura. Extracellular Fibrillar Structure of Latent TGFβ Binding Protein-1: Role in TGFβ-dependent Endothelial–Mesenchymal Transformation during Endocardial Cushion Tissue Formation in Mouse Embryonic Heart. Journal of Cell Biology, 136(1): 193–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/193.

Nurminskaya:1998:PTH

[NMNL98] Maria Nurminskaya, Cordula Magee, Dmitry Nurminsky, and Thomas F. Linsenmayer. Plasma Transglutaminase in Hypertrophic Chondrocytes: Expression and Cell-specific Intracellular Activation Produce Cell Death and Externalization. Journal of Cell Biology, 142(4):1135–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/4/1135.

Nagata:1997:TIP

[NMT97] Yuka Nagata, Yoshinao Muro, and Kazuo Todokoro. Thrombopoietininduced Polyploidization of Bone Marrow Megakaryocytes Is Due to a Unique Regulatory Mechanism in Late Mitosis. Journal of Cell Biology, 139(2):449–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/2/449.

Ng:1997:PBB

[NNK^{+}97] Florence W. H. Ng, Mai Nguyen, Tony Kwan, Philip E. Branton, Donald W. Nicholson, James A. Cromlish, and Gordon C. Shore. p28 Bap31, a Bcl-2/Bcl–X_{L}- and procaspase-8–associated protein in the endoplasmic reticulum. Journal of Cell Biology, 139(2):327–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/2/327.

Nikolic:1996:BAA

[NNMW96] B. Nikolic, E. Mac Nulty, B. Mir, and G. Wiche. Basic amino acid residue cluster within nuclear targeting sequence motif is essential for cytoplasmic plectin-vimentin network junctions. Journal of Cell Biology, 134(6):1455–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
6/1455.

Nishijima:1997:DSP

[NNSN97] Hitoshi Nishijima, Hideo Nishitani, Takashi Seki, and Takeharu Nishimoto. A dual-specificity phosphatase Cdc25B is an unstable protein and triggers p34^{cdc2} /cyclin B activation in hamster BHK21 cells arrested with hydroxyurea. Journal of Cell Biology, 138(5):1105–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/5/1105.

Nagahama:1996:VSI

[NOR^{+}96] M. Nagahama, L. Orci, M. Ravazzola, M. Amherdt, L. Lacomis, P. Tempst, J. E. Rothman, and T. H. Söllner. A v– SNARE implicated in intra-Golgi transport. Journal of Cell Biology, 133(3):507–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/3/507.

Nakanishi:1997:NNN

[NOS^{+}97] Hiroyuki Nakanishi, Hiroshi Obaishi, Ayako Satoh, Manabu Wada, Kenji Mandai, Keiko Satoh, Hideo Nishioka, Yoshiharu Matsuura, Akira Mizoguchi, and Yoshimi Takai. Neurabin: a Novel Neural Tissue–specific Actin Filament–binding Protein Involved in Neurite Formation. Journal of Cell Biology, 139(4): 951–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/951.

Nieman:1999:CPM

[NPJW99] Marvin T. Nieman, Ryan S. Prudoff, Keith R. Johnson, and Margaret J. Wheelock. N-cadherin Promotes Motility in Human Breast Cancer Cells Regardless of Their E–Cadherin Expression. Journal of Cell Biology, 147(3):631–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
3/631.

Novak:1995:DMD

[NPRT95] K. D. Novak, M. D. Peterson, M. C. Reedy, and M. A. Titus. Dictyostelium myosin I double mutants exhibit conditional defects in pinocytosis. Journal of Cell Biology, 131(5): 1205–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/5/1205.

Navarro:1998:DLV

[NRD98] Pilar Navarro, Luigi Ruco, and Elisabetta Dejana. Differential Localization of VE- and N–Cadherins in Human Endothelial Cells: VE–Cadherin Competes with N–Cadherin for Junctional Localization. Journal of Cell Biology, 140(6):1475–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/6/1475.

Nehrbass:1996:YNN

[NRM^{+}96] U. Nehrbass, M. P. Rout, S. Maguire, G. Blobel, and R. W. Wozniak. The yeast nucleoporin Nup188p interacts genetically and physically with the core structures of the nuclear pore complex. Journal of Cell Biology, 133(6):1153–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
6/1153.

Neame:1998:BCC

[NRP98] Stephen J. Neame, Lee L. Rubin, and Karen L. Philpott. Blocking Cytochrome c Activity within Intact Neurons Inhibits Apoptosis. Journal of Cell Biology, 142(6):1583– ??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/6/1583.

Nakamura:1995:CCG

[NRW^{+}95] N. Nakamura, C. Rabouille, R. Watson, T. Nilsson, N. Hui, P. Slusarewicz, T. E. Kreis, and G. Warren. Characterization of a cis-Golgi matrix protein, GM130. Journal of Cell Biology, 131(6):1715–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/6/1715.

Norcott:1996:TPS

[NSC96] J. P. Norcott, R. Solari, and D. F. Cutler. Targeting of P-selectin to two regulated secretory organelles in PC12 cells. Journal of Cell Biology, 134(5):1229–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
5/1229.

Neugebauer:1995:CES

[NSR95] K. M. Neugebauer, J. A. Stolk, and M. B. Roth. A conserved epitope on a subset of SR proteins defines a larger family of Pre-mRNA splicing factors. Journal of Cell Biology, 129 (4):899–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/4/899.

Noda:1995:KNM

[NSYK^{+}95] Y. Noda, R. Sato-Yoshitake, S. Kondo, M. Nangaku, and N. Hirokawa. KIF2 is a new microtubule-based anterograde motor that transports membranous organelles distinct from those carried by kinesin heavy chain or KIF3A/B. Journal of Cell Biology, 129(1):157–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/1/157.

Novak:1997:MOI

[NT97] Kristine D. Novak and Margaret A. Titus. Myosin I Overexpression Impairs Cell Migration. Journal of Cell Biology, 136 (3):633–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/633.

Naora:1998:ACR

[NTAN98] Honami Naora, Izumi Takai, Masakazu Adachi, and Hiroto Naora. Altered Cellular Responses by Varying Expression of a Ribosomal Protein Gene: Sequential Coordination of Enhancement and Suppression of Ribosomal Protein S3a Gene Expression Induces Apoptosis. Journal of Cell Biology, 141 (3):741–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/3/741.

Nakata:1998:VDS

[NTH98] Takao Nakata, Sumio Terada, and Nobutaka Hirokawa. Visualization of the Dynamics of Synaptic Vesicle and Plasma Membrane Proteins in Living Axons. Journal of Cell Biology, 140(3):659–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/659.

Ng:1996:MPC

[NW96] D. T. Ng and P. Walter. ER membrane protein complex required for nuclear fusion. Journal of Cell Biology, 132(4): 499–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/4/499.

Nicklas:1995:KCS

[NWG95] R. B. Nicklas, S. C. Ward, and G. J. Gorbsky. Kinetochore chemistry is sensitive to tension and may link mitotic forces to a cell cycle checkpoint. Journal of Cell Biology, 130(4): 929–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/929.

Niewohner:1997:TNC

[NWM^{+}97] Jens Niewöhner, Igor Weber, Markus Maniak, Annette Müller-Taubenberger, and Günther Gerisch. Talin-null Cells of Dictyostelium Are Strongly Defective in Adhesion to Particle and Substrate Surfaces and Slightly Impaired in Cytokinesis. Journal of Cell Biology, 138(2):349–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/2/349.

Nicolson:1995:TFP

[NWPW95] T. A. Nicolson, L. S. Weisman, G. S. Payne, and W. T. Wickner. A truncated form of the Pho80 cyclin redirects the Pho85 kinase to disrupt vacuole inheritance in S. cerevisiae. Journal of Cell Biology, 130(4):835–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/4/835.

Nakamura:1995:DCC

[NYN^{+}95] T. Y. Nakamura, I. Yamamoto, H. Nishitani, T. Matozaki, T. Suzuki, S. Wakabayashi, M. Shigekawa, and K. Goshima. Detachment of cultured cells from the substratum induced by the neutrophil-derived oxidant NH2Cl: synergistic role of phosphotyrosine and intracellular Ca^{2+} concentration. Journal of Cell Biology, 131(2):509–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/2/509.

Nicchitta:1997:RRM

[NZ97] Christopher V. Nicchitta and Tianli Zheng. Regulation of the Ribosome–Membrane Junction at Early Stages of Presecretory Protein Translocation in the Mammalian Endoplasmic Reticulum. Journal of Cell Biology, 139(7):1697–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
7/1697.

Oakes:1998:MAS

[OAB^{+}98] M. Oakes, J. P. Aris, J. S. Brockenbrough, H. Wai, L. Vu, and M. Nomura. Mutational Analysis of the Structure and Localization of the Nucleolus in the Yeast Saccharomyces cerevisiae. Journal of Cell Biology, 143(1):23–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/143/1/23.

Oulad-Abdelghani:1996:CPG

[OABD^{+}96] M. Oulad-Abdelghani, P. Bouillet, D. Décimo, A. Gansmuller, S. Heyberger, P. Dollé, S. Bronner, Y. Lutz, and P. Chambon. Characterization of a premeiotic germ cell-specific cytoplasmic protein encoded by Stra8, a novel retinoic acidresponsive gene. Journal of Cell Biology, 135(2):469–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/2/469.

Ono:1999:UAC

[OBB99] Shoichiro Ono, David L. Baillie, and Guy M. Benian. UNC-60B, an ADF/cofilin family protein, is required for proper assembly of actin into myofibrils in Caenorhabditis elegans body wall muscle. Journal of Cell Biology, 145(3):491–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
3/491.

Ogg:1998:FGD

[OBW98] Stephen C. Ogg, Wolfgang P. Barz, and Peter Walter. A Functional GTPase Domain, but not its Transmembrane Domain, is Required for Function of the SRP Receptor β-subunit. Journal of Cell Biology, 142(2):341–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/341.

Osman:1998:IYH

[OC98] Mahasin A. Osman and Richard A. Cerione. Iqg1p, a Yeast Homologue of the Mammalian IQGAPs, Mediates Cdc42p Effects on the Actin Cytoskeleton. Journal of Cell Biology, 142 (2):443–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/443.

Owen:1995:CSB

[OCS^{+}95] C. A. Owen, M. A. Campbell, P. L. Sannes, S. S. Boukedes, and E. J. Campbell. Cell surface-bound elastase and cathepsin G on human neutrophils: a novel, non-oxidative mechanism by which neutrophils focus and preserve catalytic activity of serine proteinases. Journal of Cell Biology, 131(3): 775–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/3/775.

Ooi:1998:ARF

[ODB98] Chean Eng Ooi, Esteban C. Dell’Angelica, and Juan S. Bonifacino. ADP-ribosylation Factor 1 (ARF1) Regulates Recruitment of the AP-3 Adaptor Complex to Membranes. Journal of Cell Biology, 142(2):391–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/391.

Oliver:1999:SPA

[ODJ99] Tim Oliver, Micah Dembo, and Ken Jacobson. Separation of Propulsive and Adhesive Traction Stresses in Locomoting Keratocytes. Journal of Cell Biology, 145(3):589–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
3/589.

Orlandi:1998:FDI

[OF98] Palmer A. Orlandi and Peter H. Fishman. Filipin-dependent Inhibition of Cholera Toxin: Evidence for Toxin Internalization and Activation through Caveolae-like Domains. Journal of Cell Biology, 141(4):905–??, May 1998. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/905.

Ott:1998:RTF

[OFM^{+}98] Ilka Ott, Edgar G. Fischer, Yohei Miyagi, Barbara M. Mueller, and Wolfram Ruf. A Role for Tissue Factor in Cell Adhesion and Migration Mediated by Interaction with Actin-binding Protein 280. Journal of Cell Biology, 140(5):1241–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
5/1241.

Otsuka:1995:ARG

[OFY^{+}95] A. J. Otsuka, R. Franco, B. Yang, K. H. Shim, L. Z. Tang, Y. Y. Zhang, P. Boontrakulpoontawee, A. Jeyaprakash, E. Hedgecock, and V. I. Wheaton. An ankyrin-related gene (unc-44) is necessary for proper axonal guidance in Caenorhabditis elegans. Journal of Cell Biology, 129(4):1081–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
4/1081.

Obermann:1996:SSM

[OGS^{+}96] W. M. Obermann, M. Gautel, F. Steiner, P. F. van der Ven, K. Weber, and D. O. Fürst. The structure of the sarcomeric M band: localization of defined domains of myomesin, M-protein, and the 250-kD carboxy-terminal region of titin by immunoelectron microscopy. Journal of Cell Biology, 134(6): 1441–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1441.

Ookata:1995:CBI

[OHB^{+}95] K. Ookata, S. Hisanaga, J. C. Bulinski, H. Murofushi, H. Aizawa, T. J. Itoh, H. Hotani, E. Okumura, K. Tachibana, and T. Kishimoto. Cyclin B interaction with microtubuleassociated protein 4 (MAP4) targets p34cdc2 kinase to microtubules and is a potential regulator of M-phase microtubule dynamics. Journal of Cell Biology, 128(5):849–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/849.

Ogawara:1995:DTP

[OIT^{+}95] M. Ogawara, N. Inagaki, K. Tsujimura, Y. Takai, M. Sekimata, M. H. Ha, S. Imajoh-Ohmi, S. Hirai, S. Ohno, and H. Sugiura. Differential targeting of protein kinase C and CaM kinase II signalings to vimentin. Journal of Cell Biology, 131(4):1055–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/1055.

Oritani:1996:ISC

[OK96] K. Oritani and P. W. Kincade. Identification of stromal cell products that interact with pre–B cells. Journal of Cell Biology, 134(3):771–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/3/771.

Ozawa:1998:MPR

[OK98] Masayuki Ozawa and Rolf Kemler. The Membrane-proximal Region of the E–Cadherin Cytoplasmic Domain Prevents Dimerization and Negatively Regulates Adhesion Activity. Journal of Cell Biology, 142(6):1605–??, September 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/142/6/
1605.

Otsuga:1998:DRG

[OKB^{+}98] Denichiro Otsuga, Brian R. Keegan, Ellen Brisch, John W. Thatcher, Greg J. Hermann, William Bleazard, and Janet M. Shaw. The Dynamin-related GTPase, Dnm1p, Controls Mitochondrial Morphology in Yeast. Journal of Cell Biology, 143 (2):333–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/2/333.

Olivera:1999:SKE

[OKE^{+}99] Ana Olivera, Takafumi Kohama, Lisa Edsall, Victor Nava, Olivier Cuvillier, Samantha Poulton, and Sarah Spiegel. Sphingosine Kinase Expression Increases Intracellular Sphingosine-1-Phosphate and Promotes Cell Growth and Survival. Journal of Cell Biology, 147(3):545–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/3/545.

Olsen:1999:LPP

[Ols99] Bjorn R. Olsen. Life without Perlecan Has Its Problems. Journal of Cell Biology, 147(5):909–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/5/909.

Olson:1995:AMF

[OMO95] K. R. Olson, J. R. McIntosh, and J. B. Olmsted. Analysis of MAP 4 function in living cells using green fluorescent protein (GFP) chimeras. Journal of Cell Biology, 130(3):639–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/3/639.

OMalley:1997:SAR

[OMS97] James P. O’Malley, Charlotte T. Moore, and Miriam M. Salpeter. Stabilization of Acetylcholine Receptors by Exogenous ATP and Its Reversal by cAMP and Calcium. Journal of Cell Biology, 138(1):159–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/1/159.

Oh:1998:DNC

[OMS98] Phil Oh, Deirdre P. McIntosh, and Jan E. Schnitzer. Dynamin at the Neck of Caveolae Mediates Their Budding to Form Transport Vesicles by GTP-driven Fission from the Plasma Membrane of Endothelium. Journal of Cell Biology, 141 (1):101–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/101.

Osheim:1996:MRE

[OMW^{+}96] Y. N. Osheim, E. B. Mougey, J. Windle, M. Anderson, M. O’Reilly, O. L. Miller, A. Beyer, and B. Sollner-Webb. Metazoan rDNA enhancer acts by making more genes transcriptionally active. Journal of Cell Biology, 133(5):943–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/5/943.

Ohtsuka:1998:NNA

[ONI^{+}98] Toshihisa Ohtsuka, Hiroyuki Nakanishi, Wataru Ikeda, Ayako Satoh, Yumiko Momose, Hideo Nishioka, and Yoshimi Takai.

Nexilin: a Novel Actin Filament-binding Protein Localized at Cell–Matrix Adherens Junction. Journal of Cell Biology, 143(5):1227–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/5/1227.

Orford:1999:EEC

[OOB99] Keith Orford, Caroline C. Orford, and Stephen W. Byers. Exogenous Expression of β-Catenin Regulates Contact Inhibition, Anchorage–Independent Growth, Anoikis, and Radiation–Induced Cell Cycle Arrest. Journal of Cell Biology, 146(4):855–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/4/855.

Orsulic:1996:VSF

[OP96a] S. Orsulic and M. Peifer. An in vivo structure-function study of armadillo, the beta-catenin homologue, reveals both separate and overlapping regions of the protein required for cell adhesion and for wingless signaling. Journal of Cell Biology, 134(5):1283–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/5/1283.

Ostap:1996:BKC

[OP96b] E. M. Ostap and T. D. Pollard. Biochemical kinetic characterization of the Acanthamoeba myosin–I ATPase. Journal of Cell Biology, 132(6):1053–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/6/1053.

Ostap:1996:OFM

[OP96c] E. M. Ostap and T. D. Pollard. Overlapping functions of myosin–I isoforms? Journal of Cell Biology, 133(2):221–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/2/221.

Okamoto:1998:SMD

[OPB98] Koji Okamoto, Philip S. Perlman, and Ronald A. Butow. The Sorting of Mitochondrial DNA and Mitochondrial Proteins in Zygotes: Preferential Transmission of Mitochondrial DNA to the Medial Bud. Journal of Cell Biology, 142(3):613–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
142/3/613.

Odorizzi:1996:ABE

[OPD^{+}96] G. Odorizzi, A. Pearse, D. Domingo, I. S. Trowbridge, and C. R. Hopkins. Apical and basolateral endosomes of MDCK cells are interconnected and contain a polarized sorting mechanism. Journal of Cell Biology, 135(1):139–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
1/139.

Oliferenko:1999:ACC

[OPH^{+}99] Snezhana Oliferenko, Karin Paiha, Thomas Harder, Volker Gerke, Christoph Schwärzler, Heinz Schwarz, Hartmut Beug, Ursula Günthert, and Lukas A. Huber. Analysis of Cd44-Containing Lipid Rafts. Journal of Cell Biology, 146(4): 843–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/4/843.

Ordahl:1999:MSS

[Ord99] Charles P. Ordahl. Myogenic Shape–Shifters. Journal of Cell Biology, 147(4):695–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/147/4/695.

Opstelten:1995:EGI

[ORW^{+}95] D. J. Opstelten, M. J. Raamsman, K. Wolfs, M. C. Horzinek, and P. J. Rottier. Envelope glycoprotein interactions in coronavirus assembly. Journal of Cell Biology, 131(2):339–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/2/339.

Oh:1996:DFP

[OS96a] Y. K. Oh and J. A. Swanson. Different fates of phagocytosed particles after delivery into macrophage lysosomes. Journal of Cell Biology, 132(4):585–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/4/585.

Ouyang:1996:CPN

[OS96b] P. Ouyang and S. P. Sugrue. Characterization of pinin, a novel protein associated with the desmosome-intermediate filament complex. Journal of Cell Biology, 135(4):1027–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/1027.

Okumura:1996:ITM

[OSH^{+}96] E. Okumura, T. Sekiai, S. Hisanaga, K. Tachibana, and T. Kishimoto. Initial triggering of M-phase in starfish oocytes: a possible novel component of maturation-promoting factor besides cdc2 kinase. Journal of Cell Biology, 132(1):125–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
132/1/125.

Oda:1996:ICV

[OSHG^{+}96] M. N. Oda, S. V. Scott, A. Hefner-Gravink, A. D. Caffarelli, and D. J. Klionsky. Identification of a cytoplasm to vacuole targeting determinant in aminopeptidase I. Journal of Cell Biology, 132(6):999–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/6/999.

Okazawa:1996:BIR

[OSK^{+}96] H. Okazawa, J. Shimizu, M. Kamei, I. Imafuku, H. Hamada, and I. Kanazawa. Bcl-2 inhibits retinoic acid-induced apoptosis during the neural differentiation of embryonal stem cells. Journal of Cell Biology, 132(5):955–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/5/955.

OConnor:1998:RCG

[OSM98] Kathleen L. O’Connor, Leslie M. Shaw, and Arthur M. Mercurio. Release of cAMP Gating by the α6β4 Integrin Stimulates Lamellae Formation and the Chemotactic Migration of Invasive Carcinoma Cells. Journal of Cell Biology, 143(6): 1749–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1749.

Opas:1996:CMC

[OSPJ^{+}96] M. Opas, M. Szewczenko-Pawlikowski, G. K. Jass, N. Mesaeli, and M. Michalak. Calreticulin modulates cell adhesiveness via regulation of vinculin expression. Journal of Cell Biology, 135(6):1913–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1913.

Obermann:1998:VFH

[OSR^{+}98] Wolfgang M. J. Obermann, Holger Sondermann, Alicia A. Russo, Nikola P. Pavletich, and F. Ulrich Hartl. In Vivo Function of Hsp90 Is Dependent on ATP Binding and ATP Hydrolysis. Journal of Cell Biology, 143(4):901–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/901.

Ochiya:1995:HFA

[OST^{+}95] T. Ochiya, H. Sakamoto, M. Tsukamoto, T. Sugimura, and M. Terada. Hst-1 (FGF-4) antisense oligonucleotides block murine limb development. Journal of Cell Biology, 130(4): 997–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/997.

Odorizzi:1997:SRB

[OT97] Greg Odorizzi and Ian S. Trowbridge. Structural Requirements for Basolateral Sorting of the Human Transferrin Receptor in the Biosynthetic and Endocytic Pathways of Madin–Darby Canine Kidney Cells. Journal of Cell Biology, 137(6):1255–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/6/1255.

Okada:1997:EMM

[OTL^{+}97] Akiko Okada, Catherine Tomasetto, Yves Lutz, Jean-Pierre Bellocq, Marie-Christine Rio, and Paul Basset. Expression of Matrix Metalloproteinases during Rat Skin Wound Healing: Evidence that Membrane Type-1 Matrix Metalloproteinase Is a Stromal Activator of Pro–Gelatinase A. Journal of Cell Biology, 137(1):67–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/1/67.

Oancea:1998:GFP

[OTQM98] Elena Oancea, Mary N. Teruel, Andrew F. G. Quest, and Tobias Meyer. Green Fluorescent Protein (GFP)-tagged Cysteine-rich Domains from Protein Kinase C as Fluorescent Indicators for Diacylglycerol Signaling in Living Cells. Journal of Cell Biology, 140(3):485–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/3/485.

OGrady:1998:LNC

[OTS98] Pauline O’Grady, Tran Cam Thai, and Haruo Saito. The Laminin–Nidogen Complex is a Ligand for a Specific Splice Isoform of the Transmembrane Protein Tyrosine Phosphatase LAR. Journal of Cell Biology, 141(7):1675–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/7/
1675.

Oegema:1995:CCD

[OWA95] K. Oegema, W. G. Whitfield, and B. Alberts. The cell cycledependent localization of the CP190 centrosomal protein is determined by the coordinate action of two separable domains. Journal of Cell Biology, 131(5):1261–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
5/1261.

OConnell:1999:SGB

[OWAlW99] Christopher B. O’Connell, Sally P. Wheatley, Sohail Ahmed, and Yu li Wang. The Small GTP-binding Protein Rho Regulates Cortical Activities in Cultured Cells during Division. Journal of Cell Biology, 144(2):305–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/2/305.

Oktay:1999:IMA

[OWD^{+}99] Maja Oktay, Kishore K. Wary, Michael Dans, Raymond B. Birge, and Filippo G. Giancotti. Integrin-mediated Activation of Focal Adhesion Kinase Is Required for Signaling to Jun NH_{2}-terminal Kinase and Progression through the G1 Phase of the Cell Cycle. Journal of Cell Biology, 145(7):1461–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/7/1461.

Oegema:1999:CTR

[OWM^{+}99] Karen Oegema, Christiane Wiese, Ona C. Martin, Ronald A. Milligan, Akihiro Iwamatsu, Timothy J. Mitchison, and Yixian Zheng. Characterization of Two Related Drosophila γtubulin Complexes that Differ in Their Ability to Nucleate Microtubules. Journal of Cell Biology, 144(4):721–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
4/721.

Osterhout:1999:MDO

[OWW^{+}99] Donna J. Osterhout, Amy Wolven, Rebecca M. Wolf, Marilyn D. Resh, and Moses V. Chao. Morphological Differentiation of Oligodendrocytes Requires Activation of Fyn Tyrosine Kinase. Journal of Cell Biology, 145(6):1209–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
6/1209.

Oohashi:1999:MTM

[OZF^{+}99] Toshitaka Oohashi, Xiao-Hong Zhou, Kang Feng, Brigitta Richter, Matthias Mörgelin, Maria Thereza Perez, Wei-Dong Su, Ruth Chiquet-Ehrismann, Uwe Rauch, and Reinhard Fässler. Mouse ten-m/Odz is a new family of dimeric Type II transmembrane proteins expressed in many tissues. Journal of Cell Biology, 145(3):563–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/3/563.

Peters:1997:DAS

[PAF97] Matthew F. Peters, Marvin E. Adams, and Stanley C. Froehner. Differential Association of Syntrophin Pairs with the Dystrophin Complex. Journal of Cell Biology, 138(1):81–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/1/81.

Passreiter:1998:PBI

[PAL^{+}98] Michael Passreiter, Markus Anton, Dorothee Lay, Rainer Frank, Cordula Harter, Felix T. Wieland, Karin Gorgas, and Wilhelm W. Just. Peroxisome biogenesis: Involvement of ARF and coatomer. Journal of Cell Biology, 141(2):373–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
2/373.

Plattner:1997:UOB

[PAN97] Helmut Plattner, Antonio R. Artalejo, and Erwin Neher. Ultrastructural Organization of Bovine Chromaffin Cell Cortex — Analysis by Cryofixation and Morphometry of Aspects Pertinent to Exocytosis. Journal of Cell Biology, 139(7):1709– ??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1709.

Purcell:1999:DRW

[PAT99] Karen Purcell and Spyros Artavanis-Tsakonas. The Developmental Role of warthog, the Notch Modifier Encoding Drab6. Journal of Cell Biology, 146(4):731–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/4/731.

Peranen:1996:RPP

[PAV^{+}96] J. Peränen, P. Auvinen, H. Virta, R. Wepf, and K. Simons. Rab8 promotes polarized membrane transport through reorganization of actin and microtubules in fibroblasts. Journal of Cell Biology, 135(1):153–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/1/153.

Powers:1998:TAD

[PB98] Jacqueline Powers and Charles Barlowe. Transport of Axl2p Depends on Erv14p, an ER–Vesicle Protein Related to the Drosophila cornichon Gene Product. Journal of Cell Biology, 142(5):1209–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/5/1209.

Pollack:1997:DCI

[PBA^{+}97] Anne L. Pollack, Angela I. M. Barth, Yoram Altschuler, W. James Nelson, and Keith E. Mostov. Dynamics of β-Catenin Interactions with APC Protein Regulate Epithelial Tubulogenesis. Journal of Cell Biology, 137(7):1651–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
7/1651.

Paillart:1996:AII

[PBB^{+}96] C. Paillart, J. L. Boudier, J. A. Boudier, H. Rochat, F. Couraud, and B. Dargent. Activity-induced internalization and rapid degradation of sodium channels in cultured fetal neurons. Journal of Cell Biology, 134(2):499–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/499.

Pardo:1998:CCR

[PBCS98] Luis A. Pardo, Andrea Brüggemann, Javier Camacho, and Walter Stühmer. Cell Cycle–related Changes in the Conducting Properties of r-eag K^{+} Channels. Journal of Cell Biology, 143(3):767–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/3/767.

Peracino:1998:GPS

[PBJ^{+}98] Barbara Peracino, Jane Borleis, Tian Jin, Monika Westphal, Jean-Marc Schwartz, Lijun Wu, Enrico Bracco, Günther Gerisch, Peter Devreotes, and Salvatore Bozzaro. G Protein β Subunit–null Mutants Are Impaired in Phagocytosis and Chemotaxis Due to Inappropriate Regulation of the Actin Cytoskeleton. Journal of Cell Biology, 141(7):1529–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
7/1529.

Ptasznik:1997:PKN

[PBM^{+}97] Andrzej Ptasznik, Gillian M. Beattie, Martin I. Mally, Vincenzo Cirulli, Ana Lopez, and Alberto Hayek. Phosphatidylinositol 3-Kinase Is a Negative Regulator of Cellular Differentiation. Journal of Cell Biology, 137(5):1127–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
5/1127.

Piper:1997:MPA

[PBS97] Robert C. Piper, Nia J. Bryant, and Tom H. Stevens. The membrane protein alkaline phosphatase is delivered to the vacuole by a route that is distinct from the VPS-dependent pathway. Journal of Cell Biology, 138(3):531–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/531.

Pellman:1995:TMA

[PBT^{+}95] D. Pellman, M. Bagget, Y. H. Tu, G. R. Fink, and H. Tu. Two microtubule-associated proteins required for anaphase spindle movement in Saccharomyces cerevisiae. Journal of Cell Biology, 130(6):1373–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/6/1373.

Paladini:1998:DEK

[PC98] Rudolph D. Paladini and Pierre A. Coulombe. Directed Expression of Keratin 16 to the Progenitor Basal Cells of Transgenic Mouse Skin Delays Skin Maturation. Journal of Cell Biology, 142(4):1035–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/4/1035.

Paladini:1999:FDE

[PC99] Rudolph D. Paladini and Pierre A. Coulombe. The Functional Diversity of Epidermal Keratins Revealed by the Partial Rescue of the Keratin 14 Null Phenotype by Keratin 16. Journal of Cell Biology, 146(5):1185–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/5/1185.

Pous:1998:FSS

[PCD^{+}98] C. Poüs, K. Chabin, A. Drechou, L. Barbot, T. Phung-Koskas, C. Settegrana, M. L. Bourguet-Kondracki, M. Maurice, D. Cassio, M. Guyot, and G. Durand. Functional Specialization of Stable and Dynamic Microtubules in Protein Traffic in WIF–B Cells. Journal of Cell Biology, 142(1):153–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
1/153.

Parkos:1996:CMP

[PCL^{+}96] C. A. Parkos, S. P. Colgan, T. W. Liang, A. Nusrat, A. E. Bacarra, D. K. Carnes, and J. L. Madara. CD47 mediates post-adhesive events required for neutrophil migration across polarized intestinal epithelia. Journal of Cell Biology, 132(3): 437–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/437.

Platero:1998:CCL

[PCQH98] J. Suso Platero, Amy K. Csink, Adrian Quintanilla, and Steven Henikoff. Changes in Chromosomal Localization of Heterochromatin-binding Proteins during the Cell Cycle in Drosophila. Journal of Cell Biology, 140(6):1297–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1297.

Piper:1995:VCV

[PCYS95] R. C. Piper, A. A. Cooper, H. Yang, and T. H. Stevens. VPS27 controls vacuolar and endocytic traffic through a prevacuolar compartment in Saccharomyces cerevisiae. Journal of Cell Biology, 131(3):603–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/3/603.

Pollock:1999:RMT

[PdHTV99] Nira Pollock, Eugenio L. de Hostos, Christoph W. Turck, and Ronald D. Vale. Reconstitution of membrane transport powered by a novel dimeric kinesin motor of the Unc104/Kif1a family purified from Dictyostelium. Journal of Cell Biology, 147(3):493–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/3/493.

Poranen:1999:NVH

[PDO^{+}99] Minna M. Poranen, Rimantas Daugelavičius, Päivi M. Ojala, Michael W. Hess, and Dennis H. Bamford. A Novel Virus–Host Cell Membrane Interaction. Journal of Cell Biology, 147(3): 671–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/3/671.

Pereira:1997:MAP

[PDS^{+}97a] Andrea J. Pereira, Brian Dalby, Russell J. Stewart, Stephen J. Doxsey, and Lawrence S. B. Goldstein. Mitochondrial Association of a Plus End–Directed Microtubule Motor Expressed during Mitosis in Drosophila. Journal of Cell Biology, 136 (5):1081–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/5/1081.

Pilcher:1997:ACR

[PDS^{+}97b] Brian K. Pilcher, Jo Ann Dumin, Barry D. Sudbeck, Stephen M. Krane, Howard G. Welgus, and William C. Parks. The Activity of Collagenase-1 Is Required for Keratinocyte Migration on a Type I Collagen Matrix. Journal of Cell Biology, 137(6):1445–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/6/1445.

Pazour:1999:DDI

[PDW99] Gregory J. Pazour, Bethany L. Dickert, and George B. Witman. The DHC1b (DHC2) Isoform of Cytoplasmic Dynein Is Required for Flagellar Assembly. Journal of Cell Biology, 144 (3):473–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/473.

Pederson:1998:GFN

[Ped98] Thoru Pederson. Growth Factors in the Nucleolus? Journal of Cell Biology, 143(2):279–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/2/279.

Protasi:1998:RRR

[PFAA98] Feliciano Protasi, Clara Franzini-Armstrong, and Paul D. Allen. Role of Ryanodine Receptors in the Assembly of Calcium Release Units in Skeletal Muscle. Journal of Cell Biology, 140(4):831–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/4/831.

Protasi:1997:CIS

[PFAF97] Feliciano Protasi, Clara Franzini-Armstrong, and Bernhard E. Flucher. Coordinated Incorporation of Skeletal Muscle Dihydropyridine Receptors and Ryanodine Receptors in Peripheral Couplings of BC_{3} H1 Cells. Journal of Cell Biology, 137 (4):859–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/4/859.

Powers:1997:VGN

[PFDL97] Maureen A. Powers, Douglass J. Forbes, James E. Dahlberg, and Elsebet Lund. The Vertebrate GLFG Nucleoporin, Nup98, Is an Essential Component of Multiple RNA Export Pathways. Journal of Cell Biology, 136(2):241–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/2/241.

Pittler:1995:VRP

[PFF^{+}95] S. J. Pittler, S. J. Fliesler, P. L. Fisher, P. K. Keller, and L. M. Rapp. In vivo requirement of protein prenylation for maintenance of retinal cytoarchitecture and photoreceptor structure. Journal of Cell Biology, 130(2):431–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/2/431.

Puri:1995:SCM

[PFGS95] K. D. Puri, E. B. Finger, G. Gaudernack, and T. A. Springer. Sialomucin CD34 is the major L-selectin ligand in human tonsil high endothelial venules. Journal of Cell Biology, 131(1): 261–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/261.

Press:1998:MRC

[PFHWN98] Barry Press, Yan Feng, Bernard Hoflack, and Angela Wandinger-Ness. Mutant Rab7 Causes the Accumulation of Cathepsin D and Cation-independent Mannose 6–Phosphate Receptor in an Early Endocytic Compartment. Journal of Cell Biology, 140(5):1075–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/5/1075.

Pritchard:1999:RSM

[PFKvD99] Colin E. J. Pritchard, Maarten Fornerod, Lawryn H. Kasper, and Jan M. A. van Deursen. RAE1 Is a Shuttling mRNA Export Factor That Binds to a GLEBS-like NUP98 Motif at the Nuclear Pore Complex through Multiple Domains. Journal of Cell Biology, 145(2):237–??, April 1999. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/2/237.

Pizzo:1997:DPI

[PFP97] Paola Pizzo, Cristina Fasolato, and Tullio Pozzan. Dynamic Properties of an Inositol 1,4,5-Trisphosphate- and Thapsigargin-insensitive Calcium Pool in Mammalian Cell Lines. Journal of Cell Biology, 136(2):355–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/2/
355.

Paschal:1995:INC

[PG95] B. M. Paschal and L. Gerace. Identification of NTF2, a cytosolic factor for nuclear import that interacts with nuclear pore complex protein p62. Journal of Cell Biology, 129(4):925–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/4/925.

Poulat:1995:NLT

[PGC^{+}95] F. Poulat, F. Girard, M. P. Chevron, C. Gozé, X. Rebillard, B. Calas, N. Lamb, and P. Berta. Nuclear localization of the testis determining gene product SRY. Journal of Cell Biology, 128(5):737–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/5/737.

Paululat:1997:MER

[PGD^{+}97] Achim Paululat, Anette Goubeaud, Christine Damm, Stefan Knirr, Susanne Burchard, and Renate Renkawitz-Pohl. The Mesodermal Expression of rolling stone (rost) Is Essential for Myoblast Fusion in Drosophila and Encodes a Potential Transmembrane Protein. Journal of Cell Biology, 138(2):337–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/2/337.

Pellegrini:1999:LCA

[PGP^{+}99] Graziella Pellegrini, Osvaldo Golisano, Patrizia Paterna, Alessandro Lambiase, Stefano Bonini, Paolo Rama, and Michele De Luca. Location and Clonal Analysis of Stem Cells and Their Differentiated Progeny in the Human Ocular Surface. Journal of Cell Biology, 145(4):769–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
4/769.

Parpal:1995:IPP

[PGS95] S. Parpal, J. Gustavsson, and P. Strålfors. Isolation of phosphooligosaccharide/phosphoinositol glycan from caveolae and cytosol of insulin-stimulated cells. Journal of Cell Biology, 131 (1):125–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/1/125.

Philips:1997:OBR

[PH97a] Jennifer Philips and Ira Herskowitz. Osmotic Balance Regulates Cell Fusion during Mating in Saccharomyces cerevisiae. Journal of Cell Biology, 138(5):961–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
5/961.

Porter:1997:ICT

[PH97b] Joanna C. Porter and Nancy Hogg. Integrin Cross Talk: Activation of Lymphocyte Function-associated Antigen-1 on Human T Cells Alters α4β1- and α5β1-mediated Function. Journal of Cell Biology, 138(6):1437–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/6/1437.

Philips:1998:IKK

[PH98] Jennifer Philips and Ira Herskowitz. Identification of Kel1p, a Kelch Domain-containing Protein Involved in Cell Fusion and Morphology in Saccharomyces cerevisiae. Journal of Cell Biology, 143(2):375–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/375.

Postma:1998:ALC

[PHA^{+}98] Friso R. Postma, Trudi Hengeveld, Jacqueline Alblas, Ben N. G. Giepmans, Gerben C. M. Zondag, Kees Jalink, and Wouter H. Moolenaar. Acute loss of Cell–Cell Communication Caused by G Protein–coupled Receptors: a Critical Role for c–Src. Journal of Cell Biology, 140(5):1199–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
5/1199.

Pierini:1996:FER

[PHB96] L. Pierini, D. Holowka, and B. Baird. Fc epsilon RI-mediated association of 6-micron beads with RBL-2H3 mast cells results in exclusion of signaling proteins from the forming phagosome and abrogation of normal downstream signaling. Journal of Cell Biology, 134(6):1427–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/6/1427.

Priddle:1998:DTG

[PHM^{+}98] Helen Priddle, Lance Hemmings, Susan Monkley, Alison Woods, Bipin Patel, Deborah Sutton, Graham A. Dunn, Daniel Zicha, and David R. Critchley. Disruption of the Talin Gene Compromises Focal Adhesion Assembly in Undifferentiated but Not Differentiated Embryonic Stem Cells. Journal of Cell Biology, 142(4):1121–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/4/1121.

Pierce:1998:APK

[PHN^{+}98] Andrew Pierce, Clare M. Heyworth, Sian E. Nicholls, Elaine Spooncer, T. Michael Dexter, Janet M. Lord, P. Jane Owen-Lynch, Gwen Wark, and Anthony D. Whetton. An Activated Protein Kinase C α Gives a Differentiation Signal for Hematopoietic Progenitor Cells and Mimicks Macrophage Colony-stimulating Factor–stimulated Signaling Events. Journal of Cell Biology, 140(6):1511–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/6/1511.

Peters:1995:OWT

[PHO^{+}95] P. J. Peters, V. W. Hsu, C. E. Ooi, D. Finazzi, S. B. Teal, V. Oorschot, J. G. Donaldson, and R. D. Klausner. Overexpression of wild-type and mutant ARF1 and ARF6: distinct perturbations of nonoverlapping membrane compartments. Journal of Cell Biology, 128(6):1003–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/6/1003.

Piali:1995:CPL

[PHU^{+}95] L. Piali, P. Hammel, C. Uherek, F. Bachmann, R. H. Gisler, D. Dunon, and B. A. Imhof. CD31/ PECAM-1 is a ligand for alpha v beta 3 integrin involved in adhesion of leukocytes to endothelium. Journal of Cell Biology, 130(2):451–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/451.

Paraskeva:1999:CMR

[PIB^{+}99] Efrosyni Paraskeva, Elisa Izaurralde, F. Ralf Bischoff, Jochen Huber, Ulrike Kutay, Enno Hartmann, Reinhard Lührmann, and Dirk Görlich. CRM1-mediated Recycling of Snurportin 1 to the Cytoplasm. Journal of Cell Biology, 145(2):255–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/255.

Pekny:1999:ARC

[PJE^{+}99] Milos Pekny, Clas B. Johansson, Camilla Eliasson, Josefina Stakeberg, Åsa Wallén, Thomas Perlmann, Urban Lendahl, Christer Betsholtz, Claes-Henric Berthold, and Jonas Frisén. Abnormal reaction to central nervous system injury in mice lacking glial fibrillary acidic protein and vimentin. Journal of Cell Biology, 145(3):503–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/3/503.

Peters:1995:AEA

[PJL^{+}95] L. L. Peters, K. M. John, F. M. Lu, E. M. Eicher, A. Higgins, M. Yialamas, L. C. Turtzo, A. J. Otsuka, and S. E. Lux. Ank3 (epithelial ankyrin), a widely distributed new member of the ankyrin gene family and the major ankyrin in kidney, is expressed in alternatively spliced forms, including forms that lack the repeat domain. Journal of Cell Biology, 130(2):313–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/2/313.

Patel-King:1997:CHP

[PKBHK97] Ramila S. Patel-King, Sharon E. Benashski, Alistair Harrison, and Stephen M. King. A Chlamydomonas Homologue of the Putative Murine t Complex Distorter Tctex-2 Is an Outer Arm Dynein Light Chain. Journal of Cell Biology, 137(5):1081–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/5/1081.

Prekeris:1998:SMC

[PKCS98] Rytis Prekeris, Judith Klumperman, Yu A. Chen, and Richard H. Scheller. Syntaxin 13 Mediates Cycling of Plasma Membrane Proteins via Tubulovesicular Recycling Endosomes. Journal of Cell Biology, 143(4):957–??, November 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/143/4/
957.

Paglini:1998:SRM

[PKQ^{+}98] Gabriela Paglini, Patricia Kunda, Santiago Quiroga, Kenneth Kosik, and Alfredo Cáceres. Suppression of Radixin and Moesin Alters Growth Cone Morphology, Motility, and Process Formation In Primary Cultured Neurons. Journal of Cell Biology, 143(2):443–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/443.

Pasqualini:1995:PIP

[PKR95] R. Pasqualini, E. Koivunen, and E. Ruoslahti. A peptide isolated from phage display libraries is a structural and functional mimic of an RGD-binding site on integrins. Journal of Cell Biology, 130(5):1189–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/5/1189.

Purdue:1996:THC

[PL96] P. E. Purdue and P. B. Lazarow. Targeting of human catalase to peroxisomes is dependent upon a novel COOH-terminal peroxisomal targeting sequence. Journal of Cell Biology, 134 (4):849–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/849.

Pomies:1997:CLD

[PLB97] Pascal Pomiès, Heather A. Louis, and Mary C. Beckerle. CRP1, a LIM Domain Protein Implicated in Muscle Differentiation, Interacts with α-Actinin. Journal of Cell Biology, 139 (1):157–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/157.

Porter:1996:GTM

[PLM^{+}96] R. M. Porter, S. Leitgeb, D. W. Melton, O. Swensson, R. A. Eady, and T. M. Magin. Gene targeting at the mouse cytokeratin 10 locus: severe skin fragility and changes of cytokeratin expression in the epidermis. Journal of Cell Biology, 132 (5):925–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/925.

Pfau:1995:LAD

[PLR^{+}95] S. Pfau, D. Leitenberg, H. Rinder, B. R. Smith, R. Pardi, and J. R. Bender. Lymphocyte adhesion-dependent calcium signaling in human endothelial cells. Journal of Cell Biology, 128 (5):969–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/5/969.

Petit:1995:AMS

[PLZ^{+}95] P. X. Petit, H. Lecoeur, E. Zorn, C. Dauguet, B. Mignotte, and M. L. Gougeon. Alterations in mitochondrial structure and function are early events of dexamethasone-induced thymocyte apoptosis. Journal of Cell Biology, 130(1):157–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
1/157.

Puertollano:1999:MPN

[PMBM^{+}99] Rosa Puertollano, Fernando Martín-Belmonte, Jaime Millán, María del Carmen de Marco, Juan P. Albar, Leonor Kremer, and Miguel A. Alonso. The MAL Proteolipid Is Necessary for Normal Apical Transport and Accurate Sorting of the Influenza Virus Hemagglutinin in Madin–Darby Canine Kidney Cells. Journal of Cell Biology, 145(1):141–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
1/141.

Patton:1997:DFL

[PMCS97] Bruce L. Patton, Jeffrey H. Miner, Arlene Y. Chiu, and Joshua R. Sanes. Distribution and Function of Laminins in the Neuromuscular System of Developing, Adult, and Mutant Mice. Journal of Cell Biology, 139(6):1507–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
6/1507.

Prekeris:1996:ILA

[PMCT96] R. Prekeris, M. W. Mayhew, J. B. Cooper, and D. M. Terrian. Identification and localization of an actin-binding motif that is unique to the epsilon isoform of protein kinase C and participates in the regulation of synaptic function. Journal of Cell Biology, 132(1):77–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/1/77.

Piperno:1996:IDA

[PMH96] G. Piperno, K. Mead, and S. Henderson. Inner dynein arms but not outer dynein arms require the activity of kinesin homologue protein KHP1(FLA10) to reach the distal part of flagella in Chlamydomonas. Journal of Cell Biology, 133(2):371–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/2/371.

Philpott:1997:APK

[PMKR97] Karen L. Philpott, Mary Jane McCarthy, Anke Klippel, and Lee L. Rubin. Activated Phosphatidylinositol 3-Kinase and Akt Kinase Promote Survival of Superior Cervical Neurons. Journal of Cell Biology, 139(3):809–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/809.

Powers:1995:RND

[PMMF95] M. A. Powers, C. Macaulay, F. R. Masiarz, and D. J. Forbes. Reconstituted nuclei depleted of a vertebrate GLFG nuclear pore protein, p97, import but are defective in nuclear growth and replication. Journal of Cell Biology, 128(5):721–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/721.

Park:1998:CDK

[PMP^{+}98] David S. Park, Erick J. Morris, Jaya Padmanabhan, Michael L. Shelanski, Herbert M. Geller, and Lloyd A. Greene. Cyclindependent Kinases Participate in Death of Neurons Evoked by DNA-damaging Agents. Journal of Cell Biology, 143(2): 457–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/2/457.

Petersen:1998:FDF

[PNEH98] Janni Petersen, Olaf Nielsen, Richard Egel, and Iain M. Hagan. FH3, A Domain Found in Formins, Targets the Fission Yeast Formin Fus1 to the Projection Tip During Conjugation. Journal of Cell Biology, 141(5):1217–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/5/1217.

Poliard:1995:CCI

[PNL^{+}95] A. Poliard, A. Nifuji, D. Lamblin, E. Plee, C. Forest, and O. Kellermann. Controlled conversion of an immortalized mesodermal progenitor cell towards osteogenic, chondrogenic, or adipogenic pathways. Journal of Cell Biology, 130(6):1461– ??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1461.

Patel:1995:PSM

[PNM95] K. D. Patel, M. U. Nollert, and R. P. McEver. P-selectin must extend a sufficient length from the plasma membrane to mediate rolling of neutrophils. Journal of Cell Biology, 131(6): 1893–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1893.

Paek:1997:EMV

[POR^{+}97] Inbok Paek, Lelio Orci, Mariella Ravazzola, Hediye Erdjument-Bromage, Mylene Amherdt, Paul Tempst, Thomas H. Söllner, and James E. Rothman. ERS-24, a Mammalian v–SNARE Implicated in Vesicle Traffic between the ER and the Golgi. Journal of Cell Biology, 137(5):1017–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/5/1017.

Partikian:1998:RDG

[PÖS^{+}98] Arthur Partikian, Bence Ölveczky, R. Swaminathan, Yuxin Li, and A. S. Verkman. Rapid Diffusion of Green Fluorescent Protein in the Mitochondrial Matrix. Journal of Cell Biology, 140(4):821–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/4/821.

Prestle:1996:PKC

[PPBJ96] J. Prestle, K. Pfizenmaier, J. Brenner, and F. J. Johannes. Protein kinase C mu is located at the Golgi compartment. Journal of Cell Biology, 134(6):1401–??, September 1996. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/134/6/
1401.

Powell:1996:FTD

[PPF96] J. A. Powell, L. Petherbridge, and B. E. Flucher. Formation of triads without the dihydropyridine receptor alpha subunits in cell lines from dysgenic skeletal muscle. Journal of Cell Biology, 134(2):375–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/2/375.

Perez:1998:RRN

[PPP^{+}98] Matthew K. Perez, Henry L. Paulson, Sagun J. Pendse, Sarah J. Saionz, Nancy M. Bonini, and Randall N. Pittman. Recruitment and the Role of Nuclear Localization in Polyglutamine-mediated Aggregation. Journal of Cell Biology, 143(6):1457–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1457.

Prchla:1995:VMR

[PPW^{+}95] E. Prchla, C. Plank, E. Wagner, D. Blaas, and R. Fuchs. Virusmediated release of endosomal content in vitro: different behavior of adenovirus and rhinovirus serotype 2. Journal of Cell Biology, 131(1):111–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/1/111.

Pestonjamasp:1997:SPN

[PPWL97] Kersi N. Pestonjamasp, Robert K. Pope, Julia D. Wulfkuhle, and Elizabeth J. Luna. Supervillin (p205): a novel membraneassociated, F-actin-binding protein in the Villin/Gelsolin superfamily. Journal of Cell Biology, 139(5):1255–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
5/1255.

Page:1995:TSS

[PR95] L. J. Page and M. S. Robinson. Targeting signals and subunit interactions in coated vesicle adaptor complexes. Journal of Cell Biology, 131(3):619–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/3/619.

Pemberton:1997:DPP

[PRB97] Lucy F. Pemberton, Jonathan S. Rosenblum, and Günter Blobel. A Distinct and Parallel Pathway for the Nuclear Import of an mRNA-binding Protein. Journal of Cell Biology, 139(7): 1645–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1645.

Pemberton:1999:NIT

[PRB99] Lucy F. Pemberton, Jonathan S. Rosenblum, and Günter Blobel. Nuclear Import of the TATA-binding Protein: Mediation by the Karyopherin Kap114p and a Possible Mechanism for Intranuclear Targeting. Journal of Cell Biology, 145(7):1407–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/7/1407.

Puceat:1998:SFT

[PRV98] Michel Pucéat, Serge Roche, and Guy Vassort. Src Family Tyrosine Kinase Regulates Intracellular pH in Cardiomyocytes. Journal of Cell Biology, 141(7):1637–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/7/1637.

Pruyne:1998:TCA

[PSB98] David W. Pruyne, Daniel H. Schott, and Anthony Bretscher. Tropomyosin-containing Actin Cables Direct the Myo2pdependent Polarized Delivery of Secretory Vesicles in Budding Yeast. Journal of Cell Biology, 143(7):1931–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1931.

Pahl:1996:ATF

[PSBB96] H. L. Pahl, M. Sester, H. G. Burgert, and P. A. Baeuerle. Activation of transcription factor NF-kappaB by the adenovirus E3/19K protein requires its ER retention. Journal of Cell Biology, 132(4):511–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/511.

Piperno:1998:DMR

[PSH^{+}98] Gianni Piperno, Edward Siuda, Scott Henderson, Margarethe Segil, Heikki Vaananen, and Massimo Sassaroli. Distinct Mutants of Retrograde Intraflagellar Transport (IFT) Share Similar Morphological and Molecular Defects. Journal of Cell Biology, 143(6):1591–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1591.

Page:1999:S

[PSLR99] Lesley J. Page, Penelope J. Sowerby, Winnie W. Y. Lui, and Margaret S. Robinson. γ-synergin. Journal of Cell Biology, 146(5):993–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/5/993.

Peters:1998:DML

[PSPG^{+}98] Matthew F. Peters, Hélène M. Sadoulet-Puccio, R. Mark Grady, Neal R. Kramarcy, Louis M. Kunkel, Joshua R. Sanes, Robert Sealock, and Stanley C. Froehner. Differential Membrane Localization and Intermolecular Associations of α-Dystrobrevin Isoforms in Skeletal Muscle. Journal of Cell Biology, 142(5):1269–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/5/1269.

Pazour:1995:MAP

[PSW95] G. J. Pazour, O. A. Sineshchekov, and G. B. Witman. Mutational analysis of the phototransduction pathway of Chlamydomonas reinhardtii. Journal of Cell Biology, 131(2):427– ??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/2/427.

Prekeris:1997:BMV

[PT97] Rytis Prekeris and David M. Terrian. Brain myosin V is a synaptic vesicle-associated motor protein: Evidence for a Ca^{2}+-dependent interaction with the synaptobrevin– synaptophysin complex. Journal of Cell Biology, 137(7): 1589–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/7/1589.

Paladini:1996:ORE

[PTBC96] R. D. Paladini, K. Takahashi, N. S. Bravo, and P. A. Coulombe. Onset of re-epithelialization after skin injury correlates with a reorganization of keratin filaments in wound edge keratinocytes: defining a potential role for keratin 16. Journal of Cell Biology, 132(3):381–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/3/381.

Potter:1998:CRA

[PTJ^{+}98] David A. Potter, Jennifer S. Tirnauer, Richard Janssen, Dorothy E. Croall, Christina N. Hughes, Kerry A. Fiacco, James W. Mier, Masatoshi Maki, and Ira M. Herman. Calpain Regulates Actin Remodeling during Cell Spreading. Journal of Cell Biology, 141(3):647–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/3/647.

Purohit:1999:DIP

[PTVD99] Aruna Purohit, Sharon H. Tynan, Richard Vallee, and Stephen J. Doxsey. Direct Interaction of Pericentrin with Cytoplasmic Dynein Light Intermediate Chain Contributes to Mitotic Spindle Organization. Journal of Cell Biology, 147(3): 481–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/3/481.

Prokop:1998:KMA

[PURB98] Andreas Prokop, Jay Uhler, John Roote, and Michael Bate. The kakapo Mutation Affects Terminal Arborization and Central Dendritic Sprouting of Drosophila Motorneurons. Journal of Cell Biology, 143(5):1283–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/5/1283.

Ploug:1998:AGD

[PvDA^{+}98] Thorkil Ploug, Bo van Deurs, Hua Ai, Samuel W. Cushman, and Evelyn Ralston. Analysis of GLUT4 Distribution in Whole Skeletal Muscle Fibers: Identification of Distinct Storage Compartments That Are Recruited by Insulin and Muscle Contractions. Journal of Cell Biology, 142(6):1429–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1429.

Perez:1998:EIF

[PW98] Denise Perez and Eileen White. E1B 19K inhibits Fasmediated apoptosis through FADD-dependent sequestration of FLICE. Journal of Cell Biology, 141(5):1255–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
5/1255.

Pabla:1999:IDC

[PWD^{+}99] Ravinder Pabla, Andrew S. Weyrich, Dan A. Dixon, Paul F. Bray, Thomas M. McIntyre, Stephen M. Prescott, and Guy A. Zimmerman. Integrin-dependent Control of Translation: Engagement of Integrin α_{IIb} β_{3} Regulates Synthesis of Proteins in Activated Human Platelets. Journal of Cell Biology, 144 (1):175–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/1/175.

Pasternak:1995:MFD

[PWE95] C. Pasternak, S. Wong, and E. L. Elson. Mechanical function of dystrophin in muscle cells. Journal of Cell Biology, 128(3): 355–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/3/355.

Pavalko:1995:CDS

[PWG^{+}95] F. M. Pavalko, D. M. Walker, L. Graham, M. Goheen, C. M. Doerschuk, and G. S. Kansas. The cytoplasmic domain of Lselectin interacts with cytoskeletal proteins via alpha-actinin: receptor positioning in microvilli does not require interaction with alpha-actinin. Journal of Cell Biology, 129(4):1155–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/4/1155.

Pozzi:1998:IMU

[PWGG98] Ambra Pozzi, Kishore K. Wary, Filippo G. Giancotti, and Humphrey A. Gardner. Integrin α1β1 Mediates a Unique Collagen-dependent Proliferation Pathway In Vivo. Journal of Cell Biology, 142(2):587–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/587.

Palacios:1996:RTM

[PWK^{+}96] I. Palacios, K. Weis, C. Klebe, I. W. Mattaj, and C. Dingwall. RAN/ TC4 mutants identify a common requirement for snRNP and protein import into the nucleus. Journal of Cell Biology, 133(3):485–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/3/485.

Pazour:1998:DLC

[PWW98] Gregory J. Pazour, Curtis G. Wilkerson, and George B. Witman. A Dynein Light Chain Is Essential for the Retrograde Particle Movement of Intraflagellar Transport (IFT). Journal of Cell Biology, 141(4):979–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/979.

Parton:1997:CAD

[PWZS97] Robert G. Parton, Michael Way, Natasha Zorzi, and Espen Stang. Caveolin-3 Associates with Developing T-tubules during Muscle Differentiation. Journal of Cell Biology, 136(1): 137–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/137.

Piper:1995:DTT

[PXR^{+}95] R. C. Piper, X. Xu, D. G. Russell, B. M. Little, and S. M. Landfear. Differential targeting of two glucose transporters from Leishmania enriettii is mediated by an NH2-terminal domain. Journal of Cell Biology, 128(4):499–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/499.

Peng:1999:ACN

[PXRR99] H. Benjamin Peng, Hongbo Xie, Susanna G. Rossi, and Richard L. Rotundo. Acetylcholinesterase Clustering at the Neuromuscular Junction Involves Perlecan and Dystroglycan. Journal of Cell Biology, 145(4):911–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/4/911.

Purdue:1998:PPN

[PYL98] P. Edward Purdue, Xudong Yang, and Paul B. Lazarow. Pex18p and Pex21p, a Novel Pair of Related Peroxins Essential for Peroxisomal Targeting by the PTS2 Pathway. Journal of Cell Biology, 143(7):1859–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/7/1859.

Prahlad:1998:RMV

[PYM^{+}98] Veena Prahlad, Miri Yoon, Robert D. Moir, Ronald D. Vale, and Robert D. Goldman. Rapid Movements of Vimentin on Microtubule Tracks: Kinesin-dependent Assembly of Intermediate Filament Networks. Journal of Cell Biology, 143(1): 159–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/1/159.

Plonsky:1996:IFP

[PZ96] I. Plonsky and J. Zimmerberg. The initial fusion pore induced by baculovirus GP64 is large and forms quickly. Journal of Cell Biology, 135(6):1831–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1831.

Qu:1998:DAI

[QBvD^{+}98] Zhuqing Qu, Levent Balkir, Judith C. T. van Deutekom, Paul D. Robbins, Ryan Pruchnic, and Johnny Huard. Development of Approaches to Improve Cell Survival in Myoblast Transfer Therapy. Journal of Cell Biology, 142(5): 1257–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/5/1257.

Quintyne:1999:DRM

[QGE^{+}99] N. J. Quintyne, S. R. Gill, D. M. Eckley, C. L. Crego, D. A. Compton, and T. A. Schroer. Dynactin Is Required for Microtubule Anchoring at Centrosomes. Journal of Cell Biology, 147 (2):321–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/2/321.

Qiao:1998:SSD

[QPH^{+}98] Hui Qiao, Sandra L. Pelletier, Lucas Hoffman, Jill Hacker, R. Todd Armstrong, and Judith M. White. Specific Single or Double Proline Substitutions in the “Spring-loaded” Coiled– Coil Region of the Influenza Hemagglutinin Impair or Abolish Membrane Fusion Activity. Journal of Cell Biology, 141 (6):1335–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/6/1335.

Sun:1997:GMP

[qSmLY97] Hui qiao Sun, Keng mean Lin, and Helen L. Yin. Gelsolin Modulates Phospholipase C Activity In Vivo through Phospholipid Binding. Journal of Cell Biology, 138(4):811–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/4/811.

Qu:1998:TCN

[QWS^{+}98] Zhican Qu, Lawrence A. Wolfraim, John Svaren, Markus U. Ehrengruber, Norman Davidson, and Jeffrey Milbrandt. The Transcriptional Corepressor NAB2 Inhibits NGF-induced Differentiation of PC12 Cells. Journal of Cell Biology, 142(4): 1075–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/4/1075.

Xie:1997:MSP

[qXLCH97] Han qing Xie, Dale W. Laird, Tsg-Hui Chang, and Valerie W. Hu. A Mitosis-specific Phosphorylation of the Gap Junction Protein Connexin43 in Human Vascular Cells: Biochemical Characterization and Localization. Journal of Cell Biology, 137(1):203–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/1/203.

Rosenblatt:1997:XAD

[RAA^{+}97] Jody Rosenblatt, Brian J. Agnew, Hiroshi Abe, James R. Bamburg, and Timothy J. Mitchison. Xenopus actin depolymerizing factor/cofilin (XAC) is responsible for the turnover of actin filaments in Listeria monocytogenes tails. Journal of Cell Biology, 136(6):1323–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/6/1323.

Roussel:1996:RTM

[RACHV96] P. Roussel, C. André, L. Comai, and D. Hernandez-Verdun. The rDNA transcription machinery is assembled during mitosis in active NORs and absent in inactive NORs. Journal of Cell Biology, 133(2):235–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/2/235.

Rybakova:1996:NMI

[RAE96] I. N. Rybakova, K. J. Amann, and J. M. Ervasti. A new model for the interaction of dystrophin with F-actin. Journal of Cell Biology, 135(3):661–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/3/661.

Ruppert:1995:LAM

[RAH^{+}95] M. Ruppert, S. Aigner, M. Hubbe, H. Yagita, and P. Altevogt. The L1 adhesion molecule is a cellular ligand for VLA-5. Journal of Cell Biology, 131(6):1881–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1881.

Rowe:1996:CVD

[RAM^{+}96] T. Rowe, M. Aridor, J. M. McCaffery, H. Plutner, C. Nuoffer, and W. E. Balch. COPII vesicles derived from mammalian endoplasmic reticulum microsomes recruit COPI. Journal of Cell Biology, 135(4):895–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/4/895.

Roy:1998:SHR

[RB98] Laura M. Roy and Alice Barkan. A SecY Homologue Is Required for the Elaboration of the Chloroplast Thylakoid Membrane and for Normal Chloroplast Gene Expression. Journal of Cell Biology, 141(2):385–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/2/385.

Reczek:1997:IEP

[RBB97] David Reczek, Mark Berryman, and Anthony Bretscher. Identification of EBP50: a PDZ-containing Phosphoprotein that Associates with Members of the Ezrin–Radixin–Moesin Family. Journal of Cell Biology, 139(1):169–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/139/1/
169.

Ruiz:1996:TMP

[RBL^{+}96] P. Ruiz, V. Brinkmann, B. Ledermann, M. Behrend, C. Grund, C. Thalhammer, F. Vogel, C. Birchmeier, U. Günthert, W. W. Franke, and W. Birchmeier. Targeted mutation of plakoglobin in mice reveals essential functions of desmosomes in the embryonic heart. Journal of Cell Biology, 135(1):215–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
1/215.

Robbins:1999:LME

[RBM^{+}99] Jennifer R. Robbins, Angela I. Barth, Hélène Marquis, Eugenio L. de Hostos, W. James Nelson, and Julie A. Theriot. Listeria monocytogenes exploits normal host cell processes to spread from cell to cell. Journal of Cell Biology, 146(6): 1333–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1333.

Rogakou:1999:MCD

[RBRB99] Emmy P. Rogakou, Chye Boon, Christophe Redon, and William M. Bonner. Megabase Chromatin Domains Involved in DNA Double–Strand Breaks in Vivo. Journal of Cell Biology, 146(5):905–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/5/905.

Robinson:1997:DKO

[RC97] Douglas N. Robinson and Lynn Cooley. Drosophila Kelch Is an Oligomeric Ring Canal Actin Organizer. Journal of Cell Biology, 138(4):799–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/4/799.

Rieder:1998:EMV

[RC98] Conly L. Rieder and Richard W. Cole. Entry into Mitosis in Vertebrate Somatic Cells Is Guarded by a Chromosome Damage Checkpoint That Reverses the Cell Cycle When Triggered during Early but Not Late Prophase. Journal of Cell Biology, 142(4):1013–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/4/1013.

Rafael:1996:FED

[RCC^{+}96] J. A. Rafael, G. A. Cox, K. Corrado, D. Jung, K. P. Campbell, and J. S. Chamberlain. Forced expression of dystrophin deletion constructs reveals structure-function correlations. Journal of Cell Biology, 134(1):93–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/1/93.

Rosenbaum:1999:ITE

[RCD99] Joel L. Rosenbaum, Douglas G. Cole, and Dennis R. Diener. Intraflagellar transport: The eyes have it. Journal of Cell Biology, 144(3):385–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/3/385.

Roberg:1999:LSH

[RCE^{+}99] Kevin J. Roberg, Michelle Crotwell, Peter Espenshade, Ruth Gimeno, and Chris A. Kaiser. LST1 Is a SEC24 Homologue Used for Selective Export of the Plasma Membrane ATPase from the Endoplasmic Reticulum. Journal of Cell Biology, 145(4):659–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/659.

Rieder:1995:CDA

[RCKS95] C. L. Rieder, R. W. Cole, A. Khodjakov, and G. Sluder. The checkpoint delaying anaphase in response to chromosome monoorientation is mediated by an inhibitory signal produced by unattached kinetochores. Journal of Cell Biology, 130(4): 941–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/941.

Ramos:1996:XEC

[RD96a] J. W. Ramos and D. W. DeSimone. Xenopus embryonic cell adhesion to fibronectin: position-specific activation of RGD/ synergy site-dependent migratory behavior at gastrulation. Journal of Cell Biology, 134(1):227–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/1/227.

Roth:1996:UYF

[RD96b] A. F. Roth and N. G. Davis. Ubiquitination of the yeast afactor receptor. Journal of Cell Biology, 134(3):661–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
3/661.

Radhakrishna:1997:ARF

[RD97] Harish Radhakrishna and Julie G. Donaldson. ADPribosylation Factor 6 Regulates a Novel Plasma Membrane Recycling Pathway. Journal of Cell Biology, 139(1):49–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/1/49.

Ross:1996:NRG

[RDM^{+}96] A. H. Ross, M. C. Daou, C. A. McKinnon, P. J. Condon, M. B. Lachyankar, R. M. Stephens, D. R. Kaplan, and D. E. Wolf. The neurotrophin receptor, gp75, forms a complex with the receptor tyrosine kinase TrkA. Journal of Cell Biology, 132 (5):945–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/945.

Rosenberg:1999:EAS

[RDP^{+}99] Arielle R. Rosenberg, Lélia Delamarre, Claudine Pique, Isabelle Le Blanc, Graziella Griffith, and Marie-Christine Dokhélar. Early Assembly Step of a Retroviral Envelope Glycoprotein: Analysis Using a Dominant Negative Assay. Journal of Cell Biology, 145(1):57–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/1/57.

Rezniczek:1998:LIB

[RdPRW98] Günther A. Rezniczek, José M. de Pereda, Siegfried Reipert, and Gerhard Wiche. Linking Integrin α_{6}β_{4}-based cell adhesion to the intermediate filament cytoskeleton: Direct interaction between the β_{4} subunit and plectin at multiple molecular sites. Journal of Cell Biology, 141(1):209–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/1/209.

Rakhilin:1999:BRC

[RDS^{+}99] Sergey Rakhilin, Renaldo C. Drisdel, Daphna Sagher, Daniel S. McGehee, Yolanda Vallejo, and William N. Green. α-bungarotoxin Receptors Contain α7 Subunits in Two Different Disulfide–Bonded Conformations. Journal of Cell Biology, 146(1):203–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/203.

Roach:1995:ODH

[REA95] H. I. Roach, J. Erenpreisa, and T. Aigner. Osteogenic differentiation of hypertrophic chondrocytes involves asymmetric cell divisions and apoptosis. Journal of Cell Biology, 131(2): 483–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/2/483.

Ross:1997:DAI

[REC97] Paul E. Ross, George R. Ehring, and Michael D. Cahalan. Dynamics of ATP-induced Calcium Signaling in Single Mouse Thymocytes. Journal of Cell Biology, 138(5):987–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
138/5/987.

Rossert:1995:SCA

[REdC95] J. Rossert, H. Eberspaecher, and B. de Crombrugghe. Separate cis-acting DNA elements of the mouse pro-alpha 1(I) collagen promoter direct expression of reporter genes to different type I collagen-producing cells in transgenic mice. Journal of Cell Biology, 129(5):1421–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/5/1421.

Radice:1997:PMG

[RFCR^{+}97] Glenn L. Radice, M. Celeste Ferreira-Cornwell, Stephen D. Robinson, Helen Rayburn, Lewis A. Chodosh, Masatoshi Takeichi, and Richard O. Hynes. Precocious Mammary Gland Development in P–Cadherin–deficient Mice. Journal of Cell Biology, 139(4):1025–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/4/1025.

Ruepp:1997:STB

[RFK^{+}97] Stefan Ruepp, André Furger, Ursula Kurath, Christina Kunz Renggli, Andrew Hemphill, Reto Brun, and Isabel Roditi. Survival of Trypanosoma brucei in the Tsetse Fly Is Enhanced by the Expression of Specific Forms of Procyclin. Journal of Cell Biology, 137(6):1369–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/6/1369.

Rivero:1996:DDC

[RFNF96] F. Rivero, R. Furukawa, A. A. Noegel, and M. Fechheimer. Dictyostelium discoideum cells lacking the 34,000-dalton actinbinding protein can grow, locomote, and develop, but exhibit defects in regulation of cell structure and movement: a case of partial redundancy. Journal of Cell Biology, 135(4): 965–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/965.

Rodriguez-Frade:1999:SDR

[RFVCM^{+}99] José M. Rodríguez-Frade, Antonio J. Vila-Coro, Ana Martín, Marta Nieto, Francisco Sánchez-Madrid, Amanda E. I. Proudfoot, Timothy N. C. Wells, Carlos Martínez-A, and Mario Mellado. Similarities and Differences in RANTES- and (AOP)- RANTES–triggered Signals: Implications for Chemotaxis. Journal of Cell Biology, 144(4):755–??, February 1999. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/144/4/
755.

Rossetto:1996:VSI

[RGS^{+}96] O. Rossetto, L. Gorza, G. Schiavo, N. Schiavo, R. H. Scheller, and C. Montecucco. VAMP/ synaptobrevin isoforms 1 and 2 are widely and differentially expressed in nonneuronal tissues. Journal of Cell Biology, 132(1):167–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/1/167.

Russell:1999:UFB

[RGS99] Iain D. Russell, Adam S. Grancell, and Peter K. Sorger. The Unstable F-box Protein p58-Ctf13 Forms the Structural Core of the CBF3 Kinetochore Complex. Journal of Cell Biology, 145(5):933–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/5/933.

Roof:1997:MCA

[RHA^{+}97] Dorothy J. Roof, Annmarie Hayes, Michael Adamian, Athar H. Chishti, and Tiansen Li. Molecular Characterization of abLIM, a Novel Actin-binding and Double Zinc Finger Protein. Journal of Cell Biology, 138(3):575–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/575.

Rattner:1996:TIA

[RHF^{+}96] J. B. Rattner, M. J. Hendzel, C. S. Furbee, M. T. Muller, and D. P. Bazett-Jones. Topoisomerase II alpha is associated with the mammalian centromere in a cell cycle- and species-specific manner and is required for proper centromere/kinetochore structure. Journal of Cell Biology, 134(5):1097–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
5/1097.

Rajasekaran:1996:CZO

[RHHRB96] A. K. Rajasekaran, M. Hojo, T. Huima, and E. Rodriguez-Boulan. Catenins and zonula occludens-1 form a complex during early stages in the assembly of tight junctions. Journal of Cell Biology, 132(3):451–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/3/451.

Reynolds:1998:HOG

[RHLG98] Todd B. Reynolds, B. Diane Hopkins, Matthew R. Lyons, and Todd R. Graham. The High Osmolarity Glycerol Response (HOG) MAP Kinase Pathway Controls Localization of a Yeast Golgi Glycosyltransferase. Journal of Cell Biology, 143(4): 935–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/4/935.

Raoul:1999:PCD

[RHP99] Cédric Raoul, Christopher E. Henderson, and Brigitte Pettmann. Programmed Cell Death of Embryonic Motoneurons Triggered through the FAS Death Receptor. Journal of Cell Biology, 147(5):1049–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/5/1049.

Ruhrberg:1997:PNC

[RHPW97] Christiana Ruhrberg, M. A. Nasser Hajibagheri, David A. D. Parry, and Fiona M. Watt. Periplakin, a Novel Component of Cornified Envelopes and Desmosomes That Belongs to the Plakin Family and Forms Complexes with Envoplakin. Journal of Cell Biology, 139(7):1835–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/7/1835.

Ruhrberg:1996:ENP

[RHS^{+}96] C. Ruhrberg, M. A. Hajibagheri, M. Simon, T. P. Dooley, and F. M. Watt. Envoplakin, a novel precursor of the cornified envelope that has homology to desmoplakin. Journal of Cell Biology, 134(3):715–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/3/715.

Rao:1998:NDR

[RHW^{+}98] Mala V. Rao, Megan K. Houseweart, Toni L. Williamson, Thomas O. Crawford, Janet Folmer, and Don W. Cleveland. Neurofilament-dependent Radial Growth of Motor Axons and Axonal Organization of Neurofilaments Does Not Require the Neurofilament Heavy Subunit (NF–H) or Its Phosphorylation. Journal of Cell Biology, 143(1):171–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/1/171.

Rijnboutt:1996:EGL

[RJP^{+}96a] S. Rijnboutt, G. Jansen, G. Posthuma, J. B. Hynes, J. H. Schornagel, and G. J. Strous. Endocytosis of GPI-linked membrane folate receptor-alpha. Journal of Cell Biology, 132(1): 35–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/1/35.

Ronnov-Jessen:1996:FFA

[RJP96b] L. Rønnov-Jessen and O. W. Petersen. A function for filamentous alpha-smooth muscle actin: retardation of motility in fibroblasts. Journal of Cell Biology, 134(1):67–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/134/1/67.

Rosette:1995:CCG

[RK95] C. Rosette and M. Karin. Cytoskeletal control of gene expression: depolymerization of microtubules activates NF-kappa B. Journal of Cell Biology, 128(6):1111–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/6/1111.

Ruf:1997:TIT

[RKB97] Stephanie Ruf, Hans Kössel, and Ralph Bock. Targeted Inactivation of a Tobacco Intron–containing Open Reading Frame Reveals a Novel Chloroplast-encoded Photosystem I–related Gene. Journal of Cell Biology, 139(1):95–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/139/1/95.

Rivero:1998:IAB

[RKB^{+}98] Francisco Rivero, Adam Kuspa, Regine Brokamp, Monika Matzner, and Angelika A. Noegel. Interaptin, an Actinbinding Protein of the α-Actinin Superfamily in Dictyostelium discoideum, Is Developmentally and cAMP-regulated and Associates with Intracellular Membrane Compartments. Journal of Cell Biology, 142(3):735–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/3/735.

Radhakrishna:1996:AFS

[RKD96] H. Radhakrishna, R. D. Klausner, and J. G. Donaldson. Aluminum fluoride stimulates surface protrusions in cells overexpressing the ARF6 GTPase. Journal of Cell Biology, 134 (4):935–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/935.

Rousselle:1997:LBN

[RKR^{+}97] Patricia Rousselle, Douglas R. Keene, Florence Ruggiero, Marie-France Champliaud, Michel van der Rest, and Robert E. Burgeson. Laminin 5 Binds the NC-1 Domain of Type VII Collagen. Journal of Cell Biology, 138(3):719–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/719.

Rogers:1999:RMM

[RKR^{+}99] Stephen L. Rogers, Ryan L. Karcher, Joseph T. Roland, Alexander A. Minin, Walter Steffen, and Vladimir I. Gelfand. Regulation of Melanosome Movement in the Cell Cycle by Reversible Association with Myosin V. Journal of Cell Biology, 146(6):1265–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/6/1265.

Rahman:1999:DKH

[RKRG99] Amena Rahman, Adeela Kamal, Elizabeth A. Roberts, and Lawrence S. B. Goldstein. Defective Kinesin Heavy Chain Behavior in Mouse Kinesin Light Chain Mutants. Journal of Cell Biology, 146(6):1277–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/6/1277.

Rafuse:1996:CAR

[RL96] V. F. Rafuse and L. Landmesser. Contractile activity regulates isoform expression and polysialylation of NCAM in cultured myotubes: involvement of Ca^{2+} and protein kinase C. Journal of Cell Biology, 132(5):969–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/5/969.

Richter:1999:SPP

[RL99] Stefan Richter and Gayle K. Lamppa. Stromal processing peptidase binds transit peptides and initiates their ATPdependent turnover in chloroplasts. Journal of Cell Biology, 147(1):33–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/1/33.

Ryan:1996:SVR

[RLC^{+}96] T. A. Ryan, L. Li, L. S. Chin, P. Greengard, and S. J. Smith. Synaptic vesicle recycling in synapsin I knock-out mice. Journal of Cell Biology, 134(5):1219–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/134/5/1219.

Richards:1996:NES

[RLCM96] S. A. Richards, K. M. Lounsbury, K. L. Carey, and I. G. Macara. A nuclear export signal is essential for the cytosolic localization of the Ran binding protein, RanBP1. Journal of Cell Biology, 134(5):1157–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/5/1157.

Ryan:1999:TDL

[RLMC99] Maureen C. Ryan, Keesook Lee, Yuko Miyashita, and William G. Carter. Targeted Disruption of the LAMA3 Gene in Mice Reveals Abnormalities in Survival and Late Stage Differentiation of Epithelial Cells. Journal of Cell Biology, 145 (6):1309–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/6/1309.

Rahkila:1998:DTV

[RLVM98] P. Rahkila, V. Luukela, K. Väänänen, and K. Metsikkö. Differential Targeting of Vesicular Stomatitis Virus G Protein and Influenza Virus Hemagglutinin Appears During Myogenesis of L6 Muscle Cells. Journal of Cell Biology, 140(5):1101–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/5/1101.

Rabinovitz:1997:IFC

[RM97] Isaac Rabinovitz and Arthur M. Mercurio. The Integrin α6β4 Functions in Carcinoma Cell Migration on Laminin-1 by Mediating the Formation and Stabilization of Actin-containing Motility Structures. Journal of Cell Biology, 139(7):1873– ??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1873.

Rak:1995:MPC

[RME^{+}95] J. Rak, Y. Mitsuhashi, V. Erdos, S. N. Huang, J. Filmus, and R. S. Kerbel. Massive programmed cell death in intestinal epithelial cells induced by three-dimensional growth conditions: suppression by mutant c–H–ras oncogene expression. Journal of Cell Biology, 131(6):1587–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/6/1587.

Rudge:1998:RPD

[RME98] Simon A. Rudge, Andrew J. Morris, and JoAnne Engebrecht. Relocalization of Phospholipase D Activity Mediates Membrane Formation During Meiosis. Journal of Cell Biology, 140 (1):81–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/81.

Rabouille:1995:RGS

[RMWW95] C. Rabouille, T. Misteli, R. Watson, and G. Warren. Reassembly of Golgi stacks from mitotic Golgi fragments in a cell-free system. Journal of Cell Biology, 129(3):605–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/605.

Reyes:1997:CHS

[RMY97] Jose C. Reyes, Christian Muchardt, and Moshe Yaniv. Components of the human SWI/SNF complex are enriched in active chromatin and are associated with the nuclear matrix. Journal of Cell Biology, 137(2):263–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/2/263.

Rapaport:1999:BTC

[RN99] Doron Rapaport and Walter Neupert. Biogenesis of Tom40, Core Component of the Tom Complex of Mitochondria. Journal of Cell Biology, 146(2):321–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/2/321.

Redfield:1997:CPS

[RNK97] Ann Redfield, Marvin T. Nieman, and Karen A. Knudsen. Cadherins Promote Skeletal Muscle Differentiation in Threedimensional Cultures. Journal of Cell Biology, 138(6):1323– ??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1323.

Rupp:1996:SPM

[ROG^{+}96] G. Rupp, E. O’Toole, L. C. Gardner, B. F. Mitchell, and M. E. Porter. The sup-pf-2 mutations of Chlamydomonas alter the activity of the outer dynein arms by modification of the gamma-dynein heavy chain. Journal of Cell Biology, 135(6): 1853–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1853.

Rehm:1997:AIT

[RP97] Armin Rehm and Hidde L. Ploegh. Assembly and intracellular targeting of the βγ subunits of heterotrimeric G proteins. Journal of Cell Biology, 137(2):305–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/305.

Rosenblum:1997:NIP

[RPB97] Jonathan S. Rosenblum, Lucy F. Pemberton, and Günter Blobel. A Nuclear Import Pathway for a Protein Involved in tRNA Maturation. Journal of Cell Biology, 139(7):1655–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/7/1655.

Rosenblum:1998:NIE

[RPBB98] Jonathan S. Rosenblum, Lucy F. Pemberton, Neris Bonifaci, and Günter Blobel. Nuclear Import and the Evolution of a Multifunctional RNA-binding Protein. Journal of Cell Biology, 143(4):887–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/4/887.

Raichman:1995:DLF

[RPC^{+}95] M. Raichman, M. C. Panzeri, E. Clementi, P. Papazafiri, M. Eckley, D. O. Clegg, A. Villa, and J. Meldolesi. Differential localization and functional role of calsequestrin in growing and differentiated myoblasts. Journal of Cell Biology, 128(3): 341–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/3/341.

Rojo:1997:ITP

[RPE^{+}97] Manuel Rojo, Rainer Pepperkok, Gregory Emery, Roland Kellner, Espen Stang, Robert G. Parton, and Jean Gruenberg. Involvement of the Transmembrane Protein p23 in Biosynthetic Protein Transport. Journal of Cell Biology, 139(5): 1119–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1119.

Rauch:1997:AAL

[RPS^{+}97] Frank Rauch, Bernhard Polzar, Harald Stephan, Silvia Zanotti, Renate Paddenberg, and Hans Georg Mannherz. Androgen Ablation Leads to an Upregulation and Intranuclear Accumulation of Deoxyribonuclease I in Rat Prostate Epithelial Cells Paralleling Their Apoptotic Elimination. Journal of Cell Biology, 137(4):909–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/4/909.

Renshaw:1999:FAK

[RPS99] Mark W. Renshaw, Leo S. Price, and Martin Alexander Schwartz. Focal Adhesion Kinase Mediates the Integrin Signaling Requirement for Growth Factor Activation of Map Kinase. Journal of Cell Biology, 147(3):611–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
3/611.

Romberg:1998:RKN

[RPV98] Laura Romberg, Daniel W. Pierce, and Ronald D. Vale. Role of the Kinesin Neck Region in Processive Microtubule-based Motility. Journal of Cell Biology, 140(6):1407–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1407.

Rao:1996:LPF

[RPW96] L. Rao, D. Perez, and E. White. Lamin proteolysis facilitates nuclear events during apoptosis. Journal of Cell Biology, 135(6):1441–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1441.

Ray:1999:KIR

[RPY^{+}99] Krishanu Ray, Sharon E. Perez, Zhaohuai Yang, Jenny Xu, Bruce W. Ritchings, Hermann Steller, and Lawrence S. B. Goldstein. Kinesin-II Is Required for Axonal Transport of Choline Acetyltransferase in Drosophila. Journal of Cell Biology, 147(3):507–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/3/507.

Rodgers:1996:ECI

[RR96] W. Rodgers and J. K. Rose. Exclusion of CD45 inhibits activity of p56lck associated with glycolipid-enriched membrane domains. Journal of Cell Biology, 135(6):1515–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
6/1515.

Rotundo:1997:TQC

[RRA97] Richard L. Rotundo, Susana G. Rossi, and Lili Anglister. Transplantation of Quail Collagen-tailed Acetylcholinesterase Molecules Onto the Frog Neuromuscular Synapse. Journal of Cell Biology, 136(2):367–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/2/367.

Roettger:1995:IGP

[RRH^{+}95] B. F. Roettger, R. U. Rentsch, E. M. Hadac, E. H. Hellen, T. P. Burghardt, and L. J. Miller. Insulation of a G proteincoupled receptor on the plasmalemmal surface of the pancreatic acinar cell. Journal of Cell Biology, 130(3):579–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
3/579.

Roberg:1997:PRM

[RRK97] Kevin J. Roberg, Neil Rowley, and Chris A. Kaiser. Physiological Regulation of Membrane Protein Sorting Late in the Secretory Pathway of Saccharomyces cerevisiae. Journal of Cell Biology, 137(7):1469–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/7/1469.

Rodriguez:1995:TSF

[RROA95] A. Rodríguez, M. G. Rioult, A. Ora, and N. W. Andrews. A trypanosome-soluble factor induces IP3 formation, intracellular Ca^{2+} mobilization and microfilament rearrangement in host cells. Journal of Cell Biology, 129(5):1263–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
5/1263.

Roettger:1995:DPI

[RRP^{+}95] B. F. Roettger, R. U. Rentsch, D. Pinon, E. Holicky, E. Hadac, J. M. Larkin, and L. J. Miller. Dual pathways of internalization of the cholecystokinin receptor. Journal of Cell Biology, 128 (6):1029–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/6/1029.

Reinbothe:1995:EPF

[RRRA95] S. Reinbothe, C. Reinbothe, S. Runge, and K. Apel. Enzymatic product formation impairs both the chloroplast receptor-binding function as well as translocation competence of the NADPH: protochlorophyllide oxidoreductase, a nuclearencoded plastid precursor protein. Journal of Cell Biology, 129 (2):299–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/299.

Roh:1995:IDD

[RS95] J. Y. Roh and J. R. Stanley. Intracellular domain of desmoglein 3 (pemphigus vulgaris antigen) confers adhesive function on the extracellular domain of E-cadherin without binding catenins. Journal of Cell Biology, 128(5):939–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/939.

Raucher:1999:MEI

[RS99] Drazen Raucher and Michael P. Sheetz. Membrane Expansion Increases Endocytosis Rate during Mitosis. Journal of Cell Biology, 144(3):497–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/3/497.

Rossanese:1999:GSC

[RSB^{+}99] Olivia W. Rossanese, Jon Soderholm, Brooke J. Bevis, Irina B. Sears, James O’Connor, Edward K. Williamson, and Benjamin S. Glick. Golgi Structure Correlates with Transitional Endoplasmic Reticulum Organization in Pichia pastoris and Saccharomyces cerevisiae. Journal of Cell Biology, 145 (1):69–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/1/69.

Reichsman:1996:GCM

[RSC96] F. Reichsman, L. Smith, and S. Cumberledge. Glycosaminoglycans can modulate extracellular localization of the wingless protein and promote signal transduction. Journal of Cell Biology, 135(3):819–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/3/819.

Roth:1998:LPL

[RSD98] Amy F. Roth, Daniel M. Sullivan, and Nicholas G. Davis. A Large PEST-like Sequence Directs the Ubiquitination, Endocytosis, and Vacuolar Degradation of the Yeast a–Factor Receptor. Journal of Cell Biology, 142(4):949–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
4/949.

Rohrer:1995:DCT

[RSJK95] J. Rohrer, A. Schweizer, K. F. Johnson, and S. Kornfeld. A determinant in the cytoplasmic tail of the cation-dependent mannose 6-phosphate receptor prevents trafficking to lysosomes. Journal of Cell Biology, 130(6):1297–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1297.

Robinett:1996:VLD

[RSL^{+}96] C. C. Robinett, A. Straight, G. Li, C. Willhelm, G. Sudlow, A. Murray, and A. S. Belmont. In vivo localization of DNA sequences and visualization of large-scale chromatin organization using lac operator/repressor recognition. Journal of Cell Biology, 135(6):1685–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1685.

Robinson:1995:MPD

[RSP^{+}95] D. R. Robinson, T. Sherwin, A. Ploubidou, E. H. Byard, and K. Gull. Microtubule polarity and dynamics in the control of organelle positioning, segregation, and cytokinesis in the trypanosome cell cycle. Journal of Cell Biology, 128(6):1163–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/6/1163.

Rodriguez:1996:HCI

[RSR^{+}96] A. Rodríguez, E. Samoff, M. G. Rioult, A. Chung, and N. W. Andrews. Host cell invasion by trypanosomes requires lysosomes and microtubule/kinesin-mediated transport. Journal of Cell Biology, 134(2):349–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/349.

Rohrer:1996:TLL

[RSRK96] J. Rohrer, A. Schweizer, D. Russell, and S. Kornfeld. The targeting of Lamp1 to lysosomes is dependent on the spacing of its cytoplasmic tail tyrosine sorting motif relative to the membrane. Journal of Cell Biology, 132(4):565–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
4/565.

Roberts:1998:HPS

[RSS98] Thomas M. Roberts, E. D. Salmon, and Murray Stewart. Hydrostatic Pressure Shows That Lamellipodial Motility in Ascaris Sperm Requires Membrane-associated Major Sperm Protein Filament Nucleation and Elongation. Journal of Cell Biology, 140(2):367–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/2/367.

Rolls:1999:VSG

[RST^{+}99] Melissa M. Rolls, Pascal A. Stein, Stephen S. Taylor, Edward Ha, Frank McKeon, and Tom A. Rapoport. A visual screen of a Gfp-fusion library identifies a new type of nuclear envelope membrane protein. Journal of Cell Biology, 146(1):29–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
1/29.

Rodal:1999:AIC

[RTL^{+}99] Avital A. Rodal, Jonathan W. Tetreault, Pekka Lappalainen, David G. Drubin, and David C. Amberg. Aip1p Interacts with Cofilin to Disassemble Actin Filaments. Journal of Cell Biology, 145(6):1251–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/6/1251.

Rabinovitz:1999:PKC

[RTM99] Isaac Rabinovitz, Alex Toker, and Arthur M. Mercurio. Protein Kinase C–Dependent Mobilization of the α6β4 Integrin from Hemidesmosomes and Its Association with Actin–Rich Cell Protrusions Drive the Chemotactic Migration of Carcinoma Cells. Journal of Cell Biology, 146(5):1147–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
146/5/1147.

Reilein:1998:ROM

[RTP^{+}98] Amy R. Reilein, Irina S. Tint, Natalia I. Peunova, Grigori N. Enikolopov, and Vladimir I. Gelfand. Regulation of Organelle Movement in Melanophores by Protein Kinase A (PKA), Protein Kinase C (PKC), and Protein Phosphatase 2A (PP2A). Journal of Cell Biology, 142(3):803–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/3/803.

Roy:1998:PMU

[RV98] Sudipto Roy and K. VijayRaghavan. Patterning Muscles Using Organizers: Larval Muscle Templates and Adult Myoblasts Actively Interact to Pattern the Dorsal Longitudinal Flight Muscles of Drosophila. Journal of Cell Biology, 141 (5):1135–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/5/1135.

Rodier:1995:PSP

[RVD^{+}95] J. M. Rodier, A. M. Vallés, M. Denoyelle, J. P. Thiery, and B. Boyer. pp60c-src is a positive regulator of growth factor-induced cell scattering in a rat bladder carcinoma cell line. Journal of Cell Biology, 131(3):761–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/761.

Raposo:1995:MMH

[RvSL^{+}95] G. Raposo, H. M. van Santen, R. Leijendekker, H. J. Geuze, and H. L. Ploegh. Misfolded major histocompatibility complex class I molecules accumulate in an expanded ER–Golgi intermediate compartment. Journal of Cell Biology, 131(6): 1403–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1403.

Rush:1995:TMW

[RW95] J. S. Rush and C. J. Waechter. Transmembrane movement of a water-soluble analogue of mannosylphosphoryldolichol is mediated by an endoplasmic reticulum protein. Journal of Cell Biology, 130(3):529–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/3/529.

Ramanathan:1996:AIC

[RWM^{+}96] R. Ramanathan, M. F. Wilkemeyer, B. Mittal, G. Perides, and M. E. Charness. Alcohol inhibits cell–cell adhesion mediated by human L1. Journal of Cell Biology, 133(2):381–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
2/381.

Robbins:1995:MGP

[RWO95] A. R. Robbins, R. D. Ward, and C. Oliver. A mutation in glyceraldehyde 3-phosphate dehydrogenase alters endocytosis in CHO cells. Journal of Cell Biology, 130(5):1093–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/5/1093.

Rodriguez:1997:LBC

[RWOA97] Ana Rodríguez, Paul Webster, Javier Ortego, and Norma W. Andrews. Lysosomes behave as Ca^{2+}-regulated exocytic vesicles in fibroblasts and epithelial cells. Journal of Cell Biology, 137(1):93–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/1/93.

Robinson:1999:CDR

[RWS^{+}99] John T. Robinson, Edward J. Wojcik, Mark A. Sanders, Maura McGrail, and Thomas S. Hays. Cytoplasmic Dynein Is Required for the Nuclear Attachment and Migration of Centrosomes during Mitosis in Drosophila. Journal of Cell Biology, 146(3):597–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/3/597.

Relan:1999:CEI

[RYB^{+}99] Nand K. Relan, Yan Yang, Safedin Beqaj, Jeffrey H. Miner, and Lucia Schuger. Cell Elongation Induces Laminin α2 Chain Expression in Mouse Embryonic Mesenchymal Cells. Journal of Cell Biology, 147(6):1341–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/6/1341.

Reynolds:1996:RRB

[RZZ96] K. Reynolds, A. M. Zimmer, and A. Zimmer. Regulation of RAR beta 2 mRNA expression: evidence for an inhibitory peptide encoded in the 5’-untranslated region. Journal of Cell Biology, 134(4):827–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/4/827.

Scatena:1998:NMI

[SAC^{+}98] Marta Scatena, Manuela Almeida, Michelle L. Chaisson, Nelson Fausto, Roberto F. Nicosia, and Cecilia M. Giachelli. NFκB Mediates α vβ3 Integrin-induced Endothelial Cell Survival. Journal of Cell Biology, 141(4):1083–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/4/1083.

Sun:1998:HPB

[SAEV^{+}98] Xin Sun, Alla T. Alzhanova-Ericsson, Neus Visa, Youssef Aissouni, Jian Zhao, and Bertil Daneholt. The hrp23 Protein in the Balbiani Ring Pre-mRNP Particles Is Released Just before or at the Binding of the Particles to the Nuclear Pore Complex. Journal of Cell Biology, 142(5):1181–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
5/1181.

Salas:1999:ITC

[Sal99] Pedro J. I. Salas. Insoluble γ–Tubulin–Containing Structures Are Anchored to the Apical Network of Intermediate Filaments in Polarized Caco-2 Epithelial Cells. Journal of Cell Biology, 146(3):645–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/3/645.

Serrador:1997:MIC

[SALdP^{+}97] Juan M. Serrador, José L. Alonso-Lebrero, Miguel A. del Pozo, Heinz Furthmayr, Reinhard Schwartz-Albiez, Javier Calvo, Francisco Lozano, and Francisco Sánchez-Madrid. Moesin Interacts with the Cytoplasmic Region of Intercellular Adhesion Molecule-3 and Is Redistributed to the Uropod of T Lymphocytes during Cell Polarization. Journal of Cell Biology, 138(6): 1409–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1409.

Sun:1999:CAR

[SAM^{+}99] Li Sun, Olugbenga A. Adebanjo, Baljit S. Moonga, Susanne Corisdeo, Hindupur K. Anandatheerthavarada, Gopa Biswas, Toshiya Arakawa, Yoshiyuki Hakeda, Antoliy Koval, Bali Sodam, Peter J. R. Bevis, A. James Moser, F. Anthony Lai, Solomon Epstein, Bruce R. Troen, Masayoshi Kumegawa, and Mone Zaidi. Cd38/ adp-ribosyl Cyclase. Journal of Cell Biology, 146(5):1161–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/5/1161.

St-Arnaud:1995:CEC

[SAPLHD95] R. St-Arnaud, J. Prud’homme, C. Leung-Hagesteijn, and S. Dedhar. Constitutive expression of calreticulin in osteoblasts inhibits mineralization. Journal of Cell Biology, 131(5):1351–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/5/1351.

Sauve:1999:PIR

[SAR^{+}99] Debra M. Sauvé, Hilary J. Anderson, Jill M. Ray, William M. James, and Michel Roberge. Phosphorylation-induced Rearrangement of the Histone H3 NH_{2}-terminal Domain during Mitotic Chromosome Condensation. Journal of Cell Biology, 145(2):225–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/2/225.

Spagnoli:1998:IBP

[SATW98] Francesca M. Spagnoli, Laura Amicone, Marco Tripodi, and Mary C. Weiss. Identification of a Bipotential Precursor Cell in Hepatic Cell Lines Derived from Transgenic Mice Expressing Cyto–Met in the Liver. Journal of Cell Biology, 143(4): 1101–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/4/1101.

Shuster:1999:PST

[SB99a] Charles B. Shuster and David R. Burgess. Parameters That Specify the Timing of Cytokinesis. Journal of Cell Biology, 146(5):981–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/5/981.

Smith:1999:UDT

[SB99b] Christopher D. Smith and Elizabeth H. Blackburn. Uncapping and Deregulation of Telomeres Lead to Detrimental Cellular Consequences in Yeast. Journal of Cell Biology, 145(2):203–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/2/203.

Sugiura:1999:FTP

[SB99c] Tsuyoshi Sugiura and Fedor Berditchevski. Function of α3β1– Tetraspanin Protein Complexes in Tumor Cell Invasion. Evidence for the Role of the Complexes in Production of Matrix Metalloproteinase 2 (Mmp-2). Journal of Cell Biology, 146(6): 1375–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1375.

Svitkina:1999:ACA

[SB99d] Tatyana M. Svitkina and Gary G. Borisy. Arp2/3 complex and actin depolymerizing factor/cofilin in dendritic organization and treadmilling of actin filament array in Lamellipodia. Journal of Cell Biology, 145(5):1009–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/5/1009.

Sells:1999:PAK

[SBC99] Mary Ann Sells, Jonathan T. Boyd, and Jonathan Chernoff. p21-activated Kinase 1 (Pak1) Regulates Cell Motility in Mammalian Fibroblasts. Journal of Cell Biology, 145 (4):837–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/837.

Soussan:1999:EVR

[SBD^{+}99] Lior Soussan, Darya Burakov, Mathew P. Daniels, Mira Toister-Achituv, Amir Porat, Yossef Yarden, and Zvulun Elazar. Erg30, a Vap-33–Related Protein, Functions in Protein Transport Mediated by Copi Vesicles. Journal of Cell Biology, 146(2):301–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/2/301.

Schaapveld:1998:HFI

[SBG^{+}98] Roel Q. J. Schaapveld, Luca Borradori, Dirk Geerts, Manuel R. van Leusden, Ingrid Kuikman, Mirjam G. Nievers, Carien M. Niessen, Renske D. M. Steenbergen, Peter J. F. Snijders, and Arnoud Sonnenberg. Hemidesmosome formation is initiated by the β4 integrin subunit, requires complex formation of β4 and HD1/plectin, and involves a direct interaction between β4 and the bullous pemphigoid antigen 180. Journal of Cell Biology, 142(1):271–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/1/271.

Sette:1998:IPC

[SBGR98] Claudio Sette, Arturo Bevilacqua, Raffaele Geremia, and Pellegrino Rossi. Involvement of Phospholipase Cγ1 in Mouse Egg Activation Induced by a Truncated Form of the C-kit Tyrosine Kinase Present in Spermatozoa. Journal of Cell Biology, 142 (4):1063–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/4/1063.

Scott:1997:ATV

[SBOK97] Sidney V. Scott, Misuzu Baba, Yoshinori Ohsumi, and Daniel J. Klionsky. Aminopeptidase I Is Targeted to the Vacuole by a Nonclassical Vesicular Mechanism. Journal of Cell Biology, 138(1):37–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/1/37.

Serini:1998:FDE

[SBPR^{+}98] Guido Serini, Marie-Luce Bochaton-Piallat, Patricia Ropraz, Antoine Geinoz, Laura Borsi, Luciano Zardi, and Giulio Gabbiani. The Fibronectin Domain ED–A Is Crucial for Myofibroblastic Phenotype Induction by Transforming Growth Factor-β1. Journal of Cell Biology, 142(3):873–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
3/873.

Slack:1996:AMG

[SBR^{+}96] R. S. Slack, D. J. Belliveau, M. Rosenberg, J. Atwal, H. Lochmüller, R. Aloyz, A. Haghighi, B. Lach, P. Seth, E. Cooper, and F. D. Miller. Adenovirus-mediated gene transfer of the tumor suppressor, p53, induces apoptosis in postmitotic neurons. Journal of Cell Biology, 135(4):1085–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/4/1085.

Sullivan:1998:YCC

[SBR98] Donald S. Sullivan, Sue Biggins, and Mark D. Rose. The Yeast Centrin, Cdc31p, and the Interacting Protein Kinase, Kic1p, Are Required for Cell Integrity. Journal of Cell Biology, 143(3):751–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/3/751.

Soddu:1996:IPP

[SBS^{+}96] S. Soddu, G. Blandino, R. Scardigli, S. Coen, A. Marchetti, M. G. Rizzo, G. Bossi, L. Cimino, M. Crescenzi, and A. Sacchi. Interference with p53 protein inhibits hematopoietic and muscle differentiation. Journal of Cell Biology, 134(1):193–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
1/193.

Stulnig:1998:PFA

[SBS^{+}98] Thomas M. Stulnig, Markus Berger, Thomas Sigmund, Daniel Raederstorff, Hannes Stockinger, and Werner Waldhäusl. Polyunsaturated Fatty Acids Inhibit T Cell Signal Transduction by Modification of Detergent-insoluble Membrane Domains. Journal of Cell Biology, 143(3):637–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/637.

Schneiter:1999:EIT

[SBS^{+}99] Roger Schneiter, Britta Brügger, Roger Sandhoff, Günther Zellnig, Andrea Leber, Manfred Lampl, Karin Athenstaedt, Claudia Hrastnik, Sandra Eder, Günther Daum, Fritz Paltauf, Felix T. Wieland, and Sepp D. Kohlwein. Electrospray ionization tandem mass spectrometry (Esi–Ms/Ms) analysis of the lipid molecular species composition of yeast subcellular membranes reveals acyl chain–based sorting/remodeling of distinct molecular species en route to the plasma membrane. Journal of Cell Biology, 146(4):741–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/4/741.

Schroeter:1996:TDS

[SBSG96] J. P. Schroeter, J. P. Bretaudiere, R. L. Sass, and M. A. Goldstein. Three-dimensional structure of the Z band in a normal mammalian skeletal muscle. Journal of Cell Biology, 133 (3):571–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/3/571.

Seksek:1997:TDM

[SBV97] Olivier Seksek, Joachim Biwersi, and A. S. Verkman. Translational Diffusion of Macromolecule-sized Solutes in Cytoplasm and Nucleus. Journal of Cell Biology, 138(1):131–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
1/131.

Simpson:1996:NAR

[SBW^{+}96] F. Simpson, N. A. Bright, M. A. West, L. S. Newman, R. B. Darnell, and M. S. Robinson. A novel adaptor-related protein complex. Journal of Cell Biology, 133(4):749–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
4/749.

Senger:1997:RRT

[SC97] Donna L. Senger and Robert B. Campenot. Rapid Retrograde Tyrosine Phosphorylation of trkA and Other Proteins in Rat Sympathetic Neurons in Compartmented Cultures. Journal of Cell Biology, 138(2):411–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/2/411.

Sterne:1997:NEN

[SCB^{+}97] G. D. Sterne, G. R. Coulton, R. A. Brown, C. J. Green, and G. Terenghi. Neurotrophin-3–enhanced Nerve Regeneration Selectively Improves Recovery of Muscle Fibers Expressing Myosin Heavy Chains 2b. Journal of Cell Biology, 139(3): 709–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/3/709.

Shelton:1999:NMR

[SCEB99] Christopher A. Shelton, J. Clayton Carter, Gregory C. Ellis, and Bruce Bowerman. The Nonmuscle Myosin Regulatory Light Chain Gene mlc-4 Is Required for Cytokinesis, Anterior– Posterior Polarity, and Body Morphology during Caenorhabditis elegans Embryogenesis. Journal of Cell Biology, 146 (2):439–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/2/439.

Spang:1995:CBP

[SCG^{+}95] A. Spang, I. Courtney, K. Grein, M. Matzner, and E. Schiebel. The Cdc31p-binding protein Kar1p is a component of the half bridge of the yeast spindle pole body. Journal of Cell Biology, 128(5):863–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/5/863.

Schwall:1996:HID

[SCG^{+}96] R. H. Schwall, L. Y. Chang, P. J. Godowski, D. W. Kahn, K. J. Hillan, K. D. Bauer, and T. F. Zioncheck. Heparin induces dimerization and confers proliferative activity onto the hepatocyte growth factor antagonists NK1 and NK2. Journal of Cell Biology, 133(3):709–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/3/709.

Scholey:1996:KIM

[Sch96] J. M. Scholey. Kinesin-II, a membrane traffic motor in axons, axonemes, and spindles. Journal of Cell Biology, 133(1):1–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/1/1.

Schwartz:1997:IOA

[Sch97] Martin Alexander Schwartz. Integrins, Oncogenes, and Anchorage Independence. Journal of Cell Biology, 139(3):575– ??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/3/575.

Snapp:1998:DPS

[SCH^{+}98] Karen R. Snapp, Ron Craig, Michael Herron, Robert D. Nelson, Lloyd M. Stoolman, and Geoffrey S. Kansas. Dimerization of P–Selectin Glycoprotein Ligand-1 (PSGL-1) Required for Optimal Recognition of P–Selectin. Journal of Cell Biology, 142(1):263–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/1/263.

Silberstein:1995:ASS

[SCK^{+}95] S. Silberstein, P. G. Collins, D. J. Kelleher, P. J. Rapiejko, and R. Gilmore. The alpha subunit of the Saccharomyces cerevisiae oligosaccharyltransferase complex is essential for vegetative growth of yeast and is homologous to mammalian ribophorin I. Journal of Cell Biology, 128(4):525–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/525.

Silberstein:1995:EOG

[SCKG95] S. Silberstein, P. G. Collins, D. J. Kelleher, and R. Gilmore. The essential OST2 gene encodes the 16-kD subunit of the yeast oligosaccharyltransferase, a highly conserved protein expressed in diverse eukaryotic organisms. Journal of Cell Biology, 131(2):371–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/2/371.

Spertini:1996:PSG

[SCM^{+}96] O. Spertini, A. S. Cordey, N. Monai, L. Giuffrè, and M. Schapira. P-selectin glycoprotein ligand 1 is a ligand for Lselectin on neutrophils, monocytes, and CD34+ hematopoietic progenitor cells. Journal of Cell Biology, 135(2):523–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/2/523.

Schaar:1997:CFK

[SCM^{+}97] B. T. Schaar, G. K. T. Chan, P. Maddox, E. D. Salmon, and T. J. Yen. CENP-e Function at Kinetochores Is Essential for Chromosome Alignment. Journal of Cell Biology, 139(6): 1373–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/6/1373.

Shima:1998:OIS

[SCPPW98] David T. Shima, Noemí Cabrera-Poch, Rainer Pepperkok, and Graham Warren. An Ordered Inheritance Strategy for the Golgi Apparatus: Visualization of Mitotic Disassembly Reveals a Role for the Mitotic Spindle. Journal of Cell Biology, 141(4):955–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/4/955.

Segal:1998:CAK

[SCR98] Marisa Segal, Duncan J. Clarke, and Steven I. Reed. Clb5associated kinase activity is required early in the spindle pathway for correct preanaphase nuclear positioning in Saccharomyces cerevisiae. Journal of Cell Biology, 143(1):135–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/1/135.

Siomi:1995:NLD

[SD95] H. Siomi and G. Dreyfuss. A nuclear localization domain in the hnRNP A1 protein. Journal of Cell Biology, 129(3):551–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/3/551.

Stevens:1998:MLC

[SD98] Richard C. Stevens and Trisha N. Davis. Mlc1p Is a Light Chain for the Unconventional Myosin Myo2p in Saccharomyces cerevisiae. Journal of Cell Biology, 142(3):711–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
3/711.

Spencer:1997:PTP

[SDC^{+}97] Susan Spencer, Donald Dowbenko, Jill Cheng, Wenlu Li, Jennifer Brush, Suzan Utzig, Viesturs Simanis, and Laurence A. Lasky. PSTPIP: a Tyrosine Phosphorylated Cleavage Furrow– associated Protein that Is a Substrate for a PEST Tyrosine Phosphatase. Journal of Cell Biology, 138(4):845–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
4/845.

Saunders:1997:DGA

[SdCAH^{+}97] Robert D. C. Saunders, Maria do Carmo Avides, Thomas Howard, Cayetano Gonzalez, and David M. Glover. The Drosophila Gene abnormal spindle Encodes a Novel Microtubuleassociated Protein That Associates with the Polar Regions of the Mitotic Spindle. Journal of Cell Biology, 137(4):881–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/4/881.

Subtil:1997:HSC

[SDDV97] Agathe Subtil, Muriel Delepierre, and Alice Dautry-Varsat. An α-Helical Signal in the Cytosolic Domain of the Interleukin 2 Receptor β Chain Mediates Sorting Towards Degradation after Endocytosis. Journal of Cell Biology, 136(3):583–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
3/583.

Sheff:1999:RRP

[SDHM99] David R. Sheff, Elizabeth A. Daro, Michael Hull, and Ira Mellman. The Receptor Recycling Pathway Contains Two Distinct Populations of Early Endosomes with Different Sorting Functions. Journal of Cell Biology, 145(1):123–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
1/123.

Skovronsky:1998:DNI

[SDL98] Daniel M. Skovronsky, Robert W. Doms, and Virginia M.-Y. Lee. Detection of a Novel Intraneuronal Pool of Insoluble Amyloid β Protein that Accumulates with Time in Culture. Journal of Cell Biology, 141(4):1031–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/4/1031.

Stack:1995:VMP

[SDTE95] J. H. Stack, D. B. DeWald, K. Takegawa, and S. D. Emr. Vesicle-mediated protein transport: regulatory interactions between the Vps15 protein kinase and the Vps34 PtdIns 3kinase essential for protein sorting to the vacuole in yeast. Journal of Cell Biology, 129(2):321–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/2/321.

Sullivan:1999:LTL

[SEAB^{+}99] Teresa Sullivan, Diana Escalante-Alcalde, Harshida Bhatt, Miriam Anver, Narayan Bhat, Kunio Nagashima, Colin L. Stewart, and Brian Burke. Loss of a–Type Lamin Expression Compromises Nuclear Envelope Integrity Leading to Muscular Dystrophy. Journal of Cell Biology, 147(5):913–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
5/913.

Sutton:1999:CSC

[SEB99] R. Bryan Sutton, James A. Ernst, and Axel T. Brunger. Crystal Structure of the Cytosolic C2a–C2b Domains of Synaptotagmin III. Journal of Cell Biology, 147(3):589–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
3/589.

Suter:1998:ISC

[SEBF98] Daniel M. Suter, Laura D. Errante, Victoria Belotserkovsky, and Paul Forscher. The Ig Superfamily Cell Adhesion Molecule, apCAM, Mediates Growth Cone Steering by Substrate–Cytoskeletal Coupling. Journal of Cell Biology, 141 (1):227–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/227.

Slack:1998:CTR

[SEBW^{+}98] Ruth S. Slack, Hiba El-Bizri, Josée Wong, Daniel J. Belliveau, and Freda D. Miller. A Critical Temporal Requirement for the Retinoblastoma Protein Family During Neuronal Determination. Journal of Cell Biology, 140(6):1497–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1497.

Spinardi:1995:RTL

[SEC^{+}95] L. Spinardi, S. Einheber, T. Cullen, T. A. Milner, and F. G. Giancotti. A recombinant tail-less integrin beta 4 subunit disrupts hemidesmosomes, but does not suppress alpha 6 beta 4-mediated cell adhesion to laminins. Journal of Cell Biology, 129(2):473–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/473.

Sers:1997:GIA

[SEH^{+}97a] Christine Sers, Urban Emmenegger, Knut Husmann, Katharina Bucher, Ann-Catherine Andres, and Reinhold Schäfer. Growth-inhibitory Activity and Downregulation of the Class II Tumor-suppressor Gene H-rev107 in Tumor Cell Lines and Experimental Tumors. Journal of Cell Biology, 136(4):935– ??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/4/935.

Sodeik:1997:MMT

[SEH97b] Beate Sodeik, Melanie W. Ebersold, and Ari Helenius. Microtubule-mediated Transport of Incoming Herpes Simplex Virus 1 Capsids to the Nucleus. Journal of Cell Biology, 136 (5):1007–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/5/1007.

Siomi:1997:TMN

[SEK^{+}97] Mikiko C. Siomi, Paul S. Eder, Naoyuki Kataoka, Lili Wan, Qing Liu, and Gideon Dreyfuss. Transportin-mediated Nuclear Import of Heterogeneous Nuclear RNP Proteins. Journal of Cell Biology, 138(6):1181–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/6/1181.

Smith:1998:DIB

[SF98] Elizabeth A. Smith and Elaine Fuchs. Defining the Interactions Between Intermediate Filaments and Desmosomes. Journal of Cell Biology, 141(5):1229–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/5/1229.

Saoudi:1998:SRI

[SFA^{+}98] Yasmina Saoudi, Rati Fotedar, Ariane Abrieu, Marcel Dorée, Jürgen Wehland, Robert L. Margolis, and Didier Job. Stepwise Reconstitution of Interphase Microtubule Dynamics in Permeabilized Cells and Comparison to Dynamic Mechanisms in Intact Cells. Journal of Cell Biology, 142(6):1519–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1519.

Schilke:1996:CSM

[SFD^{+}96] B. Schilke, J. Forster, J. Davis, P. James, W. Walter, S. Laloraya, J. Johnson, B. Miao, and E. Craig. The cold sensitivity of a mutant of Saccharomyces cerevisiae lacking a mitochondrial heat shock protein 70 is suppressed by loss of mitochondrial DNA. Journal of Cell Biology, 134(3):603–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
3/603.

Saitou:1998:ODE

[SFD^{+}98] Mitinori Saitou, Kazushi Fujimoto, Yoshinori Doi, Masahiko Itoh, Toyoshi Fujimoto, Mikio Furuse, Hiroshi Takano, Tetsuo Noda, and Shoichiro Tsukita. Occludin-deficient Embryonic Stem Cells Can Differentiate into Polarized Epithelial Cells Bearing Tight Junctions. Journal of Cell Biology, 141 (2):397–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/2/397.

Scidmore:1996:SGD

[SFH96] M. A. Scidmore, E. R. Fischer, and T. Hackstadt. Sphingolipids and glycoproteins are differentially trafficked to the Chlamydia trachomatis inclusion. Journal of Cell Biology, 134 (2):363–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/363.

Simons:1995:BKS

[SFNRH95] J. F. Simons, S. Ferro-Novick, M. D. Rose, and A. Helenius. BiP/ kar2p serves as a molecular chaperone during carboxypeptidase Y folding in yeast. Journal of Cell Biology, 130(1):41–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/1/41.

Shi:1998:NSV

[SFR^{+}98] Gongyi Shi, Victor Faúndez, Jack Roos, Esteban C. Dell’Angelica, and Regis B. Kelly. Neuroendocrine Synaptic Vesicles Are Formed In Vitro by Both Clathrin-dependent and Clathrin-independent Pathways. Journal of Cell Biology, 143(4):947–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/4/947.

Sakakibara:1997:PIP

[SFS^{+}97] Akira Sakakibara, Mikio Furuse, Mitinori Saitou, Yuhko Ando-Akatsuka, and Shoichiro Tsukita. Possible Involvement of Phosphorylation of Occludin in Tight Junction Formation. Journal of Cell Biology, 137(6):1393–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/6/1393.

Sonoda:1999:CPE

[SFS^{+}99] Noriyuki Sonoda, Mikio Furuse, Hiroyuki Sasaki, Shigenobu Yonemura, Jun Katahira, Yasuhiko Horiguchi, and Shoichiro Tsukita. Clostridium perfringens Enterotoxin Fragment Removes Specific Claudins from Tight Junction Strands. Journal of Cell Biology, 147(1):195–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/1/195.

Shaw:1995:CSS

[SFSV95] A. Shaw, P. A. Fortes, C. D. Stout, and V. D. Vacquier. Crystal structure and subunit dynamics of the abalone sperm lysin dimer: egg envelopes dissociate dimers, the monomer is the active species. Journal of Cell Biology, 130(5):1117–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/5/1117.

Sawitzky:1995:PGA

[SG95] H. Sawitzky and F. Grolig. Phragmoplast of the green alga Spirogyra is functionally distinct from the higher plant phragmoplast. Journal of Cell Biology, 130(6):1359–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1359.

Sweet:1996:GDR

[SG96] D. J. Sweet and L. Gerace. A GTPase distinct from Ran is involved in nuclear protein import. Journal of Cell Biology, 133(5):971–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/5/971.

Shih:1997:TSP

[SG97] Theodore M. Shih and Alan L. Goldin. Topology of the Shaker Potassium Channel Probed with Hydrophilic Epitope Insertions. Journal of Cell Biology, 136(5):1037–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
5/1037.

South:1999:PSA

[SG99a] Sarah T. South and Stephen J. Gould. Peroxisome Synthesis in the Absence of Preexisting Peroxisomes. Journal of Cell Biology, 144(2):255–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/2/255.

Stinchcombe:1999:RSH

[SG99b] Jane C. Stinchcombe and Gillian M. Griffiths. Regulated Secretion from Hemopoietic Cells. Journal of Cell Biology, 147 (1):1–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/1/1.

Steinmetz:1997:CAA

[SGA97] Michel O. Steinmetz, Kenneth N. Goldie, and Ueli Aebi. A Correlative Analysis of Actin Filament Assembly, Structure, and Dynamics. Journal of Cell Biology, 138(3):559–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/559.

Sundberg:1996:RCS

[SGBD96] H. A. Sundberg, L. Goetsch, B. Byers, and T. N. Davis. Role of calmodulin and Spc110p interaction in the proper assembly of spindle pole body compenents. Journal of Cell Biology, 133 (1):111–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/1/111.

Schumacher:1998:AAI

[SGD98] Jill M. Schumacher, Andy Golden, and Peter J. Donovan. AIR-2: an Aurora/Ipl1-related protein kinase associated with chromosomes and midbody microtubules is required for polar body extrusion and cytokinesis in Caenorhabditis elegans embryos. Journal of Cell Biology, 143(6):1635–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1635.

Spang:1996:GTL

[SGGS96] A. Spang, S. Geissler, K. Grein, and E. Schiebel. gammatubulin-like Tub4p of Saccharomyces cerevisiae is associated with the spindle pole body substructures that organize microtubules and is required for mitotic spindle formation. Journal of Cell Biology, 134(2):429–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/429.

Sabourin:1999:RDP

[SGGS^{+}99] Luc A. Sabourin, Adele Girgis-Gabardo, Patrick Seale, Atsushi Asakura, and Michael A. Rudnicki. Reduced Differentiation Potential of Primary MyoD -/- Myogenic Cells Derived from Adult Skeletal Muscle. Journal of Cell Biology, 144(4): 631–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/4/631.

Shiraha:1999:IIE

[SGGW99] Hidenori Shiraha, Angela Glading, Kiran Gupta, and Alan Wells. Ip-10 Inhibits Epidermal Growth Factor–Induced Motility by Decreasing Epidermal Growth Factor Receptor– Mediated Calpain Activity. Journal of Cell Biology, 146 (1):243–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/243.

Stan:1999:ICL

[SGJP99] Radu-Virgil Stan, Lucian Ghitescu, Bruce S. Jacobson, and George E. Palade. Isolation, Cloning, and Localization of Rat PV-1, a Novel Endothelial Caveolar Protein. Journal of Cell Biology, 145(6):1189–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/6/1189.

Slot:1997:GTG

[SGM^{+}97] Jan W. Slot, Gabriella Garruti, Sally Martin, Viola Oorschot, George Posthuma, Edward W. Kraegen, Ross Laybutt, Gaétan Thibault, and David E. James. Glucose Transporter (GLUT-4) Is Targeted to Secretory Granules in Rat Atrial Cardiomyocytes. Journal of Cell Biology, 137(6):1243–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
6/1243.

Sehgal:1997:ACA

[SGR97] Ravinder N. M. Sehgal, Barry M. Gumbiner, and Louis F. Reichardt. Antagonism of Cell Adhesion by an α-Catenin Mutant, and of the Wnt-signaling Pathway by α-Catenin in Xenopus Embryos. Journal of Cell Biology, 139(4):1033–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/4/1033.

Samuels:1995:CTR

[SGS95] A. L. Samuels, T. H. Giddings, and L. A. Staehelin. Cytokinesis in tobacco BY-2 and root tip cells: a new model of cell plate formation in higher plants. Journal of Cell Biology, 130(6): 1345–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1345.

Schutz:1997:YCG

[SGSW97] Amy R. Schutz, Thomas H. Giddings, Estelle Steiner, and Mark Winey. The Yeast CDC37 Gene Interacts with MPS1 and Is Required for Proper Execution of Spindle Pole Body Duplication. Journal of Cell Biology, 136(5):969–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
5/969.

Sanders:1999:GAB

[SGTH99] Sylvia L. Sanders, Martina Gentzsch, Widmar Tanner, and Ira Herskowitz. O-glycosylation of Axl2/Bud10p by Pmt4p is required for its stability, localization, and function in daughter cells. Journal of Cell Biology, 145(6):1177–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
6/1177.

Sugiyama:1997:LIA

[SGYH97] J. E. Sugiyama, D. J. Glass, G. D. Yancopoulos, and Z. W. Hall. Laminin-induced Acetylcholine Receptor Clustering: an Alternative Pathway. Journal of Cell Biology, 139(1): 181–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/181.

Simcha:1996:STP

[SGYL^{+}96] I. Simcha, B. Geiger, S. Yehuda-Levenberg, D. Salomon, and A. Ben-Ze’ev. Suppression of tumorigenicity by plakoglobin: an augmenting effect of N-cadherin. Journal of Cell Biology, 133(1):199–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/1/199.

Shuster:1995:IAE

[SH95] C. B. Shuster and I. M. Herman. Indirect association of ezrin with F-actin: isoform specificity and calcium sensitivity. Journal of Cell Biology, 128(5):837–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/5/837.

Sanders:1996:BPY

[SH96] S. L. Sanders and I. Herskowitz. The BUD4 protein of yeast, required for axial budding, is localized to the mother/BUD neck in a cell cycle-dependent manner. Journal of Cell Biology, 134(2):413–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/413.

Small:1995:AFO

[SHA95] J. V. Small, M. Herzog, and K. Anderson. Actin filament organization in the fish keratocyte lamellipodium. Journal of Cell Biology, 129(5):1275–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/5/1275.

Sheets:1999:CRC

[SHB99] Erin D. Sheets, David Holowka, and Barbara Baird. Critical Role for Cholesterol in Lyn-mediated Tyrosine Phosphorylation of FcϵRI and Their Association with Detergent-resistant Membranes. Journal of Cell Biology, 145(4):877–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
4/877.

Schafer:1995:ICD

[SHC95] D. A. Schafer, C. Hug, and J. A. Cooper. Inhibition of CapZ during myofibrillogenesis alters assembly of actin filaments. Journal of Cell Biology, 128(1):61–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/128/1/61.

Saris:1997:HHL

[SHC^{+}97] Nina Saris, Heidi Holkeri, Rachel A. Craven, Colin J. Stirling, and Marja Makarow. The Hsp70 Homologue Lhs1p Is Involved in a Novel Function of the Yeast Endoplasmic Reticulum, Refolding and Stabilization of Heat-denatured Protein Aggregates. Journal of Cell Biology, 137(4):813–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
4/813.

Suomalainen:1996:TMM

[SHG96] M. Suomalainen, K. Hultenby, and H. Garoff. Targeting of Moloney murine leukemia virus gag precursor to the site of virus budding. Journal of Cell Biology, 135(6):1841–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/6/1841.

Schmidt:1997:SLM

[SHH97] Anne Schmidt, Matthew J. Hannah, and Wieland B. Huttner. Synaptic-like Microvesicles of Neuroendocrine Cells Originate from a Novel Compartment That Is Continuous with the Plasma Membrane and Devoid of Transferrin Receptor. Journal of Cell Biology, 137(2):445–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/445.

Slee:1999:OCC

[SHK^{+}99] Elizabeth A. Slee, Mary T. Harte, Ruth M. Kluck, Beni B. Wolf, Carlos A. Casiano, Donald D. Newmeyer, Hong-Gang Wang, John C. Reed, Donald W. Nicholson, Emad S. Alnemri, Douglas R. Green, and Seamus J. Martin. Ordering the cytochrome c-initiated caspase cascade: Hierarchical activation of caspases-2, -3, -6, -7, -8, and -10 in a caspase-9-dependent manner. Journal of Cell Biology, 144(2):281–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
2/281.

Stepp:1997:YAP

[SHL97] J. David Stepp, Kristen Huang, and Sandra K. Lemmon. The Yeast Adaptor Protein Complex, AP-3, Is Essential for the Efficient Delivery of Alkaline Phosphatase by the Alternate Pathway to the Vacuole. Journal of Cell Biology, 139(7): 1761–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1761.

Saunders:1997:SCK

[SHLD97] William Saunders, David Hornack, Valerie Lengyel, and Changchun Deng. The Saccharomyces cerevisiae Kinesinrelated Motor Kar3p Acts at Preanaphase Spindle Poles to Limit the Number and Length of Cytoplasmic Microtubules. Journal of Cell Biology, 137(2):417–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/2/417.

Schneikert:1999:NAE

[SHMC99] Jean Schneikert, Susanne Hübner, Elke Martin, and Andrew C. B. Cato. A Nuclear Action of the Eukaryotic Cochaperone Rap46 in Downregulation of Glucocorticoid Receptor Activity. Journal of Cell Biology, 146(5):929–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
5/929.

Shima:1997:PGA

[SHP^{+}97] David T. Shima, Kasturi Haldar, Rainer Pepperkok, Rose Watson, and Graham Warren. Partitioning of the Golgi Apparatus during Mitosis in Living HeLa Cells. Journal of Cell Biology, 137(6):1211–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/6/1211.

Schott:1999:CTD

[SHPB99] Daniel Schott, Jackson Ho, David Pruyne, and Anthony Bretscher. The Cooh–Terminal Domain of Myo2p, a Yeast Myosin V, Has a Direct Role in Secretory Vesicle Targeting. Journal of Cell Biology, 147(4):791–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/147/4/
791.

Soans:1996:DED

[SHPP96] C. Soans, J. A. Holash, Y. Pavlova, and E. B. Pasquale. Developmental expression and distinctive tyrosine phosphorylation of the Eph-related receptor tyrosine kinase Cek9. Journal of Cell Biology, 135(3):781–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/3/781.

Schlenstedt:1995:YDH

[SHR^{+}95] G. Schlenstedt, S. Harris, B. Risse, R. Lill, and P. A. Silver. A yeast DnaJ homologue, Scj1p, can function in the endoplasmic reticulum with BiP/Kar2p via a conserved domain that specifies interactions with Hsp70s. Journal of Cell Biology, 129(4):979–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/4/979.

Shelby:1996:DEB

[SHS96] R. D. Shelby, K. M. Hahn, and K. F. Sullivan. Dynamic elastic behavior of alpha-satellite DNA domains visualized in situ in living human cells. Journal of Cell Biology, 135(3): 545–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/545.

Shibamoto:1995:APT

[SHT^{+}95] S. Shibamoto, M. Hayakawa, K. Takeuchi, T. Hori, K. Miyazawa, N. Kitamura, K. R. Johnson, M. J. Wheelock, N. Matsuyoshi, and M. Takeichi. Association of p120, a tyrosine kinase substrate, with E-cadherin/catenin complexes. Journal of Cell Biology, 128(5):949–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/5/949.

Huang:1999:NTT

[sHWM^{+}99] Lily Jun shen Huang, Lin Wang, Yuliang Ma, Kyle Durick, Guy Perkins, Thomas J. Deerinck, Mark H. Ellisman, and Susan S. Taylor. NH_{2} -Terminal Targeting Motifs Direct Dual Specificity A–Kinase–anchoring Protein 1 (D–AKAP1) to Either Mitochondria or Endoplasmic Reticulum. Journal of Cell Biology, 145(5):951–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/5/951.

Satoh:1996:IIC

[SHY^{+}96] M. Satoh, K. Hirayoshi, S. Yokota, N. Hosokawa, and K. Nagata. Intracellular interaction of collagen-specific stress protein HSP47 with newly synthesized procollagen. Journal of Cell Biology, 133(2):469–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/2/469.

Simon:1996:PPG

[SIAS96] J. P. Simon, I. E. Ivanov, M. Adesnik, and D. D. Sabatini. The production of post-Golgi vesicles requires a protein kinase Clike molecule, but not its phosphorylating activity. Journal of Cell Biology, 135(2):355–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/2/355.

Simon:1999:ACB

[Sim99] Sanford M. Simon. An Award for Cell Biology. Journal of Cell Biology, 147(5):1–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/147/5/1.

Shimizu:1998:SEE

[SIUW98] Noriaki Shimizu, Nobuo Itoh, Hiroyasu Utiyama, and Geoffrey M. Wahl. Selective Entrapment of Extrachromosomally Amplified DNA by Nuclear Budding and Micronucleation during S Phase. Journal of Cell Biology, 140(6):1307–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
6/1307.

Shu:1995:GTC

[SJ95] H. B. Shu and H. C. Joshi. Gamma-tubulin can both nucleate microtubule assembly and self-assemble into novel tubular structures in mammalian cells. Journal of Cell Biology, 130(5): 1137–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/5/1137.

Sesaki:1999:DVF

[SJ99] Hiromi Sesaki and Robert E. Jensen. Division versus Fusion: Dnm1p and Fzo1p Antagonistically Regulate Mitochondrial Shape. Journal of Cell Biology, 147(4):699–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
4/699.

Sweet:1996:PLH

[SJA96] M. T. Sweet, K. Jones, and C. D. Allis. Phosphorylation of linker histone is associated with transcriptional activation in a normally silent nucleus. Journal of Cell Biology, 135(5): 1219–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/5/1219.

Schafer:1996:DCP

[SJC96] D. A. Schafer, P. B. Jennings, and J. A. Cooper. Dynamics of capping protein and actin assembly in vitro: uncapping barbed ends by polyphosphoinositides. Journal of Cell Biology, 135 (1):169–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/169.

Serra:1997:ETK

[SJF^{+}97] Rosa Serra, Mahlon Johnson, Ellen H. Filvaroff, James LaBorde, Daniel M. Sheehan, Rik Derynck, and Harold L. Moses. Expression of a Truncated, Kinase–Defective TGF-β Type II Receptor in Mouse Skeletal Tissue Promotes Terminal Chondrocyte Differentiation and Osteoarthritis. Journal of Cell Biology, 139(2):541–??, October 1997. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/2/541.

Shpetner:1996:PSP

[SJHC96] H. Shpetner, M. Joly, D. Hartley, and S. Corvera. Potential sites of PI-3 kinase function in the endocytic pathway revealed by the PI-3 kinase inhibitor, wortmannin. Journal of Cell Biology, 132(4):595–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/595.

Stenius:1995:SSP

[SJSJ95] K. Stenius, R. Janz, T. C. Südhof, and R. Jahn. Structure of synaptogyrin (p29) defines novel synaptic vesicle protein. Journal of Cell Biology, 131(6):1801–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1801.

Sternsdorf:1997:ECM

[SJW97] Thomas Sternsdorf, Kirsten Jensen, and Hans Will. Evidence for covalent modification of the nuclear dot-associated proteins PML and Sp100 by PIC1/SUMO-1. Journal of Cell Biology, 139(7):1621–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/7/1621.

Sternsdorf:1997:CLE

[SJZW97] Thomas Sternsdorf, Kirsten Jensen, Dirk Züchner, and Hans Will. Cellular Localization, Expression, and Structure of the Nuclear Dot Protein 52. Journal of Cell Biology, 138 (2):435–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/2/435.

Sako:1995:BLD

[SK95a] Y. Sako and A. Kusumi. Barriers for lateral diffusion of transferrin receptor in the plasma membrane as characterized by receptor dragging by laser tweezers: fence versus tether. Journal of Cell Biology, 129(6):1559–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/6/1559.

Scott:1995:VRC

[SK95b] S. V. Scott and D. J. Klionsky. In vitro reconstitution of cytoplasm to vacuole protein targeting in yeast. Journal of Cell Biology, 131(6):1727–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1727.

Santini:1996:EAR

[SK96] F. Santini and J. H. Keen. Endocytosis of activated receptors and clathrin-coated pit formation: deciphering the chicken or egg relationship. Journal of Cell Biology, 132(6):1025–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
132/6/1025.

Schadendorf:1996:TMC

[SKA^{+}96] D. Schadendorf, M. A. Kern, M. Artuc, H. L. Pahl, T. Rosenbach, I. Fichtner, W. Nürnberg, S. Stüting, E. von Stebut, M. Worm, A. Makki, K. Jurgovsky, G. Kolde, and B. M. Henz. Treatment of melanoma cells with the synthetic retinoid CD437 induces apoptosis via activation of AP-1 in vitro, and causes growth inhibition in xenografts in vivo. Journal of Cell Biology, 135(6):1889–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/6/1889.

Skov:1997:LMH

[SKC97] Søren Skov, Pia Klausen, and Mogens H. Claesson. Ligation of major histocompatability complex (MHC) class I molecules on human T cells induces cell death through PI-3 kinase-induced c-Jun NH_{2}-terminal kinase activity: A novel apoptotic pathway distinct from Fas-induced apoptosis. Journal of Cell Biology, 139(6):1523–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/6/1523.

Saunders:1995:SCK

[SKE^{+}95] W. S. Saunders, D. Koshland, D. Eshel, I. R. Gibbons, and M. A. Hoyt. Saccharomyces cerevisiae kinesin- and dyneinrelated proteins required for anaphase chromosome segregation. Journal of Cell Biology, 128(4):617–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
4/617.

Strunnikov:1995:CEC

[SKK95] A. V. Strunnikov, J. Kingsbury, and D. Koshland. CEP3 encodes a centromere protein of Saccharomyces cerevisiae. Journal of Cell Biology, 128(5):749–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/5/749.

Schweizer:1996:CCT

[SKR96] A. Schweizer, S. Kornfeld, and J. Rohrer. Cysteine34 of the cytoplasmic tail of the cation-dependent mannose 6-phosphate receptor is reversibly palmitoylated and required for normal trafficking and lysosomal enzyme sorting. Journal of Cell Biology, 132(4):577–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/4/577.

Sakai:1998:PDM

[SKR^{+}98] Yasuyoshi Sakai, Antonius Koller, Linda K. Rangell, Gilbert A. Keller, and Suresh Subramani. Peroxisome Degradation by Microautophagy in Pichia pastoris: Identification of Specific Steps and Morphological Intermediates. Journal of Cell Biology, 141(3):625–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/3/625.

Segall:1995:MKN

[SKS^{+}95] J. E. Segall, A. Kuspa, G. Shaulsky, M. Ecke, M. Maeda, C. Gaskins, R. A. Firtel, and W. F. Loomis. A MAP kinase necessary for receptor-mediated activation of adenylyl cyclase in Dictyostelium. Journal of Cell Biology, 128(3): 405–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/3/405.

Serra:1999:PHR

[SKS99] Rosa Serra, Andrew Karaplis, and Philip Sohn. Parathyroid Hormone–related Peptide (PTHrP)-dependent and - independent Effects of Transforming Growth Factor β (TGFβ) on Endochondral Bone Formation. Journal of Cell Biology, 145(4):783–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/783.

Sun:1995:ECO

[SKWY95] H. Q. Sun, K. Kwiatkowska, D. C. Wooten, and H. L. Yin. Effects of CapG overexpression on agonist-induced motility and second messenger generation. Journal of Cell Biology, 129 (1):147–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/1/147.

Shirai:1998:DEF

[SKY^{+}98] Yasuhito Shirai, Kaori Kashiwagi, Keiko Yagi, Norio Sakai, and Naoaki Saito. Distinct Effects of Fatty Acids on Translocation of γ- and ϵ-Subspecies of Protein Kinase C. Journal of Cell Biology, 143(2):511–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/2/511.

Sullivan:1995:ECB

[SL95] R. Sullivan and C. W. Lo. Expression of a connexin 43/betagalactosidase fusion protein inhibits gap junctional communication in NIH3T3 cells. Journal of Cell Biology, 130(2):419–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
130/2/419.

Smith:1996:PEP

[SL96] E. F. Smith and P. A. Lefebvre. PF16 encodes a protein with armadillo repeats and localizes to a single microtubule of the central apparatus in Chlamydomonas flagella. Journal of Cell Biology, 132(3):359–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/132/3/359.

Snapp:1997:CAI

[SL97] Erik L. Snapp and Scott M. Landfear. Cytoskeletal Association Is Important for Differential Targeting of Glucose Transporter Isoforms in Leishmania. Journal of Cell Biology, 139(7): 1775–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1775.

Sartorius:1998:MHC

[SLAS^{+}98] Carol A. Sartorius, Brian D. Lu, Leslie Acakpo-Satchivi, Renee P. Jacobsen, William C. Byrnes, and Leslie A. Leinwand. Myosin Heavy Chains IIa and IId Are Functionally Distinct in the Mouse. Journal of Cell Biology, 141(4):943–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
4/943.

Schumacher:1995:SMR

[SLEB95] J. M. Schumacher, K. Lee, S. Edelhoff, and R. E. Braun. Spnr, a murine RNA-binding protein that is localized to cytoplasmic microtubules. Journal of Cell Biology, 129(4):1023–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
4/1023.

Sadoni:1999:NOM

[SLF^{+}99] Nicolas Sadoni, Sabine Langer, Christine Fauth, Giorgio Bernardi, Thomas Cremer, Bryan M. Turner, and Daniele Zink. Nuclear Organization of Mammalian Genomes. Journal of Cell Biology, 146(6):1211–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/6/1211.

Sakaguchi:1996:ICA

[SLL96] T. Sakaguchi, G. P. Leser, and R. A. Lamb. The ion channel activity of the influenza virus M2 protein affects transport through the Golgi apparatus. Journal of Cell Biology, 133 (4):733–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/4/733.

Sonnichsen:1998:RGD

[SLL^{+}98] Birte Sönnichsen, Martin Lowe, Tim Levine, Eija Jämsä, Barbara Dirac-Svejstrup, and Graham Warren. A Role for Giantin in Docking COPI Vesicles to Golgi Membranes. Journal of Cell Biology, 140(5):1013–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/5/1013.

Sadoul:1995:SEI

[SLM^{+}95] K. Sadoul, J. Lang, C. Montecucco, U. Weller, R. Regazzi, S. Catsicas, C. B. Wollheim, and P. A. Halban. SNAP-25 is expressed in islets of Langerhans and is involved in insulin release. Journal of Cell Biology, 128(6):1019–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
6/1019.

Sakurai:1997:INO

[SLN^{+}97] Takeshi Sakurai, Marc Lustig, Moshe Nativ, John J. Hemperly, Joseph Schlessinger, Elior Peles, and Martin Grumet. Induction of Neurite Outgrowth through Contactin and Nr–CAM by Extracellular Regions of Glial Receptor Tyrosine Phosphatase β. Journal of Cell Biology, 136(4):907–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
4/907.

Starich:1996:EUR

[SLP^{+}96] T. A. Starich, R. Y. Lee, C. Panzarella, L. Avery, and J. E. Shaw. eat-5 and unc-7 represent a multigene family in Caenorhabditis elegans involved in cell–cell coupling. Journal of Cell Biology, 134(2):537–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/2/537.

Sloan-Lancaster:1998:ZAC

[SLPE^{+}98] Joanne Sloan-Lancaster, John Presley, Jan Ellenberg, Tetsuo Yamazaki, Jennifer Lippincott-Schwartz, and Lawrence E. Samelson. ZAP-70 association with T cell receptor ζ (TCRζ): Fluorescence imaging of dynamic changes upon cellular stimulation. Journal of Cell Biology, 143(3):613–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
3/613.

Schroeder:1996:LMC

[SLS^{+}96] H. Schroeder, R. Leventis, S. Shahinian, P. A. Walton, and J. R. Silvius. Lipid-modified, cysteinyl-containing peptides of diverse structures are efficiently S-acylated at the plasma membrane of mammalian cells. Journal of Cell Biology, 134 (3):647–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/3/647.

Stupack:1999:MVR

[SLS^{+}99] Dwayne G. Stupack, Erguang Li, Steve A. Silletti, Jacqueline A. Kehler, Robert L. Geahlen, Klaus Hahn, Glen R. Nemerow, and David A. Cheresh. Matrix Valency Regulates Integrin-mediated Lymphoid Adhesion via Syk Kinase. Journal of Cell Biology, 144(4):777–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/4/777.

Sapperstein:1996:AGS

[SLSW96] S. K. Sapperstein, V. V. Lupashin, H. D. Schmitt, and M. G. Waters. Assembly of the ER to Golgi SNARE complex requires Uso1p. Journal of Cell Biology, 132(5):755–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
5/755.

Sastry:1996:IAS

[SLT^{+}96] S. K. Sastry, M. Lakonishok, D. A. Thomas, J. Muschler, and A. F. Horwitz. Integrin alpha subunit ratios, cytoplasmic domains, and growth factor synergy regulate muscle proliferation and differentiation. Journal of Cell Biology, 133(1):169–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/1/169.

Scherer:1996:CNC

[SLW^{+}96] P. E. Scherer, G. Z. Lederkremer, S. Williams, M. Fogliano, G. Baldini, and H. F. Lodish. Cab45, a novel (Ca^{2+})-binding protein localized to the Golgi lumen. Journal of Cell Biology, 133(2):257–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/2/257.

Sastry:1999:QCI

[SLW^{+}99] Sarita K. Sastry, Margot Lakonishok, Stanley Wu, Tho Q. Truong, Anna Huttenlocher, Christopher E. Turner, and Alan F. Horwitz. Quantitative Changes in Integrin and Focal Adhesion Signaling Regulate Myoblast Cell Cycle Withdrawal.

Journal of Cell Biology, 144(6):1295–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/6/1295.

Shibasaki:1995:CFC

[SM95a] F. Shibasaki and F. McKeon. Calcineurin functions in Ca(2+)- activated cell death in mammalian cells. Journal of Cell Biology, 131(3):735–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/3/735.

Song:1995:AFM

[SM95b] Y. H. Song and E. Mandelkow. The anatomy of flagellar microtubules: polarity, seam, junctions, and lattice. Journal of Cell Biology, 128(1):81–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/1/81.

Stahl:1995:UTP

[SM95c] A. Stahl and B. M. Mueller. The urokinase-type plasminogen activator receptor, a GPI-linked protein, is localized in caveolae. Journal of Cell Biology, 129(2):335–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
2/335.

Stauffer:1997:CIR

[SM97] Thomas P. Stauffer and Tobias Meyer. Compartmentalized IgE Receptor–mediated Signal Transduction in Living Cells. Journal of Cell Biology, 139(6):1447–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
6/1447.

Smart:1996:CFR

[SMA96] E. J. Smart, C. Mineo, and R. G. Anderson. Clustered folate receptors deliver 5-methyltetrahydrofolate to cytoplasm of MA104 cells. Journal of Cell Biology, 134(5):1169–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/5/1169.

Shartava:1995:PMB

[SMB^{+}95] A. Shartava, C. A. Monteiro, F. A. Bencsath, K. Schneider, B. T. Chait, R. Gussio, L. A. Casoria-Scott, A. K. Shah, C. A. Heuerman, and S. R. Goodman. A posttranslational modification of beta-actin contributes to the slow dissociation of the spectrin-protein 4.1-actin complex of irreversibly sickled cells. Journal of Cell Biology, 128(5):805–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
5/805.

Sharp:1999:BKK

[SMB^{+}99] David J. Sharp, Kent L. McDonald, Heather M. Brown, Heinrich J. Matthies, Claire Walczak, Ron D. Vale, Timothy J. Mitchison, and Jonathan M. Scholey. The Bipolar Kinesin, KLP61F, Cross-links Microtubules within Interpolar Microtubule Bundles of Drosophila Embryonic Mitotic Spindles. Journal of Cell Biology, 144(1):125–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/1/125.

Song:1996:NSR

[SMC^{+}96] K. Song, K. E. Mach, C. Y. Chen, T. Reynolds, and C. F. Albright. A novel suppressor of ras1 in fission yeast, byr4, is a dosage-dependent inhibitor of cytokinesis. Journal of Cell Biology, 133(6):1307–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/6/1307.

Seaman:1997:EGR

[SMCE97] Matthew N. J. Seaman, Eric G. Marcusson, Joan Lin Cereghino, and Scott D. Emr. Endosome to Golgi Retrieval of the Vacuolar Protein Sorting Receptor, Vps10p, Requires the Function of the VPS29, VPS30, and VPS35 Gene Products. Journal of Cell Biology, 137(1):79–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/1/79.

Seaman:1998:MCC

[SME98] Matthew N. J. Seaman, J. Michael McCaffery, and Scott D. Emr. A membrane coat complex essential for endosome-to-Golgi retrograde transport in yeast. Journal of Cell Biology, 142(3):665–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/3/665.

Shimizu:1995:RPK

[SMFC95] Y. Shimizu, J. L. Mobley, L. D. Finkelstein, and A. S. Chan. A role for phosphatidylinositol 3-kinase in the regulation of beta 1 integrin activity by the CD2 antigen. Journal of Cell Biology, 131(6):1867–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1867.

Spann:1997:DNL

[SMG^{+}97] Timothy P. Spann, Robert D. Moir, Anne E. Goldman, Reimer Stick, and Robert D. Goldman. Disruption of Nuclear Lamin Organization Alters the Distribution of Replication Factors and Inhibits DNA Synthesis. Journal of Cell Biology, 136(6): 1201–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/6/1201.

Stewart:1998:LMA

[SMH98] Mairi P. Stewart, Alison McDowall, and Nancy Hogg. LFA-1-mediated adhesion is regulated by cytoskeletal restraint and by a Ca^{2+}-dependent protease, calpain. Journal of Cell Biology, 140(3):699–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/699.

Sparks:1999:SSP

[SMM99] Cynthia A. Sparks, Mary Morphew, and Dannel McCollum. Sid2p, a Spindle Pole Body Kinase That Regulates the Onset of Cytokinesis. Journal of Cell Biology, 146(4):777–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
4/777.

Scannevin:1996:ICD

[SMRT96] R. H. Scannevin, H. Murakoshi, K. J. Rhodes, and J. S. Trimmer. Identification of a cytoplasmic domain important in the polarized expression and clustering of the Kv2.1 K+ channel. Journal of Cell Biology, 135(6):1619–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
6/1619.

Siddhanta:1998:DRP

[SMS^{+}98] Uma Siddhanta, James McIlroy, Amishi Shah, Yitao Zhang, and Jonathan M. Backer. Distinct Roles for the p110α and hVPS34 Phosphatidylinositol 3’-Kinases in Vesicular Trafficking, Regulation of the Actin Cytoskeleton, and Mitogenesis. Journal of Cell Biology, 143(6):1647–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
6/1647.

Sumi:1999:CPA

[SMTN99] Tomoyuki Sumi, Kunio Matsumoto, Yoshimi Takai, and Toshikazu Nakamura. Cofilin Phosphorylation and Actin Cytoskeletal Dynamics Regulated by Rho- and Cdc42-Activated Lim–Kinase 2. Journal of Cell Biology, 147(7):1519–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/7/1519.

Smith:1999:PEP

[SMW^{+}99] Kelly P. Smith, Phillip T. Moen, Karen L. Wydner, John R. Coleman, and Jeanne B. Lawrence. Processing of Endogenous Pre-mRNAs in Association with SC-35 Domains Is Gene Specific. Journal of Cell Biology, 144(4):617–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
4/617.

Siegel:1998:DEF

[SMZ^{+}98] Richard M. Siegel, David A. Martin, Lixin Zheng, Samuel Y. Ng, John Bertin, Jeffrey Cohen, and Michael J. Lenardo. Death-effector Filaments: Novel Cytoplasmic Structures that Recruit Caspases and Trigger Apoptosis. Journal of Cell Biology, 141(5):1243–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/5/1243.

Smith:1997:MFE

[SN97] Corey Smith and Erwin Neher. Multiple Forms of Endocytosis In Bovine Adrenal Chromaffin Cells. Journal of Cell Biology, 139(4):885–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/4/885.

Sawin:1998:RCP

[SN98] Kenneth E. Sawin and Paul Nurse. Regulation of Cell Polarity by Microtubules in Fission Yeast. Journal of Cell Biology, 142 (2):457–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/457.

Sontag:1995:NPP

[SNCBM95] E. Sontag, V. Nunbhakdi-Craig, G. S. Bloom, and M. C. Mumby. A novel pool of protein phosphatase 2A is associated with microtubules and is regulated during the cell cycle. Journal of Cell Biology, 128(6):1131–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/6/1131.

Stinchcombe:1995:ART

[SNCH95] J. C. Stinchcombe, H. Nomoto, D. F. Cutler, and C. R. Hopkins. Anterograde and retrograde traffic between the rough endoplasmic reticulum and the Golgi complex. Journal of Cell Biology, 131(6):1387–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1387.

Sato:1997:MSP

[SNK^{+}97] Hiroshi Sato, Toshio Nagai, Dhandapani Kuppuswamy, Takahiro Narishige, Masaaki Koide, Donald R. Menick, and George Cooper. Microtubule Stabilization in Pressure Overload Cardiac Hypertrophy. Journal of Cell Biology, 139(4): 963–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/963.

Suomalainen:1999:MDP

[SNK^{+}99] Maarit Suomalainen, Michel Y. Nakano, Stephan Keller, Karin Boucke, Robert P. Stidwill, and Urs F. Greber. Microtubuledependent Plus- and Minus End–directed Motilities Are Competing Processes for Nuclear Targeting of Adenovirus. Journal of Cell Biology, 144(4):657–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/4/657.

Schmitt:1996:MCH

[SNL96] M. Schmitt, W. Neupert, and T. Langer. The molecular chaperone Hsp78 confers compartment-specific thermotolerance to mitochondria. Journal of Cell Biology, 134(6):1375– ??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1375.

Sako:1998:CRM

[SNT^{+}98] Yasushi Sako, Akira Nagafuchi, Shoichiro Tsukita, Masatoshi Takeichi, and Akihiro Kusumi. Cytoplasmic Regulation of the Movement of E–Cadherin on the Free Cell Surface as Studied by Optical Tweezers and Single Particle Tracking: Corralling and Tethering by the Membrane Skeleton. Journal of Cell Biology, 140(5):1227–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/5/1227.

Su:1997:CAM

[SO97] Tin Tin Su and Patrick H. O’Farrell. Chromosome Association of Minichromosome Maintenance Proteins in Drosophila Mitotic Cycles. Journal of Cell Biology, 139(1):13–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
1/13.

Su:1998:CAM

[SO98] Tin Tin Su and Patrick H. O’Farrell. Chromosome Association of Minichromosome Maintenance Proteins in Drosophila Endoreplication Cycles. Journal of Cell Biology, 140(3):451– ??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/3/451.

Shattil:1995:BEN

[SOE^{+}95] S. J. Shattil, T. O’Toole, M. Eigenthaler, V. Thon, M. Williams, B. M. Babior, and M. H. Ginsberg. Beta 3endonexin, a novel polypeptide that interacts specifically with the cytoplasmic tail of the integrin beta 3 subunit. Journal of Cell Biology, 131(3):807–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/3/807.

Stoorvogel:1996:NCC

[SOG96] W. Stoorvogel, V. Oorschot, and H. J. Geuze. A novel class of clathrin-coated vesicles budding from endosomes. Journal of Cell Biology, 132(1):21–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/1/21.

Sase:1995:RMA

[SOH^{+}95] I. Sase, T. Okinaga, M. Hoshi, G. W. Feigenson, and K. Kinosita. Regulatory mechanisms of the acrosome reaction revealed by multiview microscopy of single starfish sperm. Journal of Cell Biology, 131(4):963–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/4/963.

Schaller:1995:FAK

[SOHP95] M. D. Schaller, C. A. Otey, J. D. Hildebrand, and J. T. Parsons. Focal adhesion kinase and paxillin bind to peptides mimicking beta integrin cytoplasmic domains. Journal of Cell Biology, 130(5):1181–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/5/1181.

Stabler:1999:MCB

[SOJM99] Stacy M. Stabler, Lisa L. Ostrowski, Susan M. Janicki, and Mervyn J. Monteiro. A Myristoylated Calcium-binding Protein that Preferentially Interacts with the Alzheimer’s Disease Presenilin 2 Protein. Journal of Cell Biology, 145(6):1277–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/6/1277.

Strutz:1995:ICF

[SOL^{+}95] F. Strutz, H. Okada, C. W. Lo, T. Danoff, R. L. Carone, J. E. Tomaszewski, and E. G. Neilson. Identification and characterization of a fibroblast marker: FSP1. Journal of Cell Biology, 130(2):393–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/2/393.

Signoret:1997:PES

[SOPM^{+}97] Natalie Signoret, Joanne Oldridge, Annegret Pelchen-Matthews, Per J. Klasse, Thanh Tran, Lawrence F. Brass, Mette M.

Rosenkilde, Thue W. Schwartz, William Holmes, Walt Dallas, Michael A. Luther, Timothy N. C. Wells, James A. Hoxie, and Mark Marsh. Phorbol Esters and SDF-1 Induce Rapid Endocytosis and Down Modulation of the Chemokine Receptor CXCR4. Journal of Cell Biology, 139(3):651–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/651.

Shaywitz:1995:HSF

[SOR^{+}95] D. A. Shaywitz, L. Orci, M. Ravazzola, A. Swaroop, and C. A. Kaiser. Human SEC13Rp functions in yeast and is located on transport vesicles budding from the endoplasmic reticulum. Journal of Cell Biology, 128(5):769–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/5/769.

Sohn:1996:MTP

[SOR^{+}96] K. Sohn, L. Orci, M. Ravazzola, M. Amherdt, M. Bremser, F. Lottspeich, K. Fiedler, J. B. Helms, and F. T. Wieland. A major transmembrane protein of Golgi-derived COPI-coated vesicles involved in coatomer binding. Journal of Cell Biology, 135(5):1239–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/5/1239.

Stone:1996:AMK

[SP96] E. M. Stone and L. Pillus. Activation of an MAP kinase cascade leads to Sir3p hyperphosphorylation and strengthens transcriptional silencing. Journal of Cell Biology, 135(3): 571–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/571.

Suter:1995:BBN

[SPB^{+}95] D. M. Suter, G. E. Pollerberg, A. Buchstaller, R. J. Giger, W. J. Dreyer, and P. Sonderegger. Binding between the neural cell adhesion molecules axonin-1 and Nr-CAM/ Bravo is involved in neuron-glia interaction. Journal of Cell Biology, 131(4):1067–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/1067.

Saint-Pol:1997:TFP

[SPBCM97] Agnès Saint-Pol, Chantal Bauvy, Patrice Codogno, and Stuart E. H. Moore. Transfer of Free Polymannose-type Oligosaccharides from the Cytosol to Lysosomes in Cultured Human Hepatocellular Carcinoma HEPG2 Cells. Journal of Cell Biology, 136(1):45–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/1/45.

Simpson:1997:CAR

[SPCR97] Fiona Simpson, Andrew A. Peden, Lina Christopoulou, and Margaret S. Robinson. Characterization of the Adaptorrelated Protein Complex, AP-3. Journal of Cell Biology, 137 (4):835–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/4/835.

Sabapathy:1997:PMI

[SPK^{+}97] Kanaga T. Sabapathy, Michael S. Pepper, Friedemann Kiefer, Uta Möhle-Steinlein, Fabienne Tacchini-Cottier, Ingrid Fetka, Georg Breier, Werner Risau, Peter Carmeliet, Roberto Montesano, and Erwin F. Wagner. Polyoma middle T-induced vascular tumor formation: The role of the plasminogen activator/ plasmin system. Journal of Cell Biology, 137(4):953–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
4/953.

Scheel:1997:DCM

[SPL^{+}97] Jochen Scheel, Rainer Pepperkok, Martin Lowe, Gareth Griffiths, and Thomas E. Kreis. Dissociation of Coatomer from Membranes Is Required for Brefeldin A–induced Transfer of Golgi Enzymes to the Endoplasmic Reticulum. Journal of Cell Biology, 137(2):319–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/2/319.

Sauter:1996:ISS

[SPMB^{+}96] M. M. Sauter, A. Pelchen-Matthews, R. Bron, M. Marsh, C. C. LaBranche, P. J. Vance, J. Romano, B. S. Haggarty, T. K. Hart, W. M. Lee, and J. A. Hoxie. An internalization signal in the simian immunodeficiency virus transmembrane protein cytoplasmic domain modulates expression of envelope glycoproteins on the cell surface. Journal of Cell Biology, 132 (5):795–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/795.

Seghezzi:1998:FGF

[SPR^{+}98] Graziano Seghezzi, Sundeep Patel, Christine J. Ren, Anna Gualandris, Giuseppe Pintucci, Edith S. Robbins, Richard L. Shapiro, Aubrey C. Galloway, Daniel B. Rifkin, and Paolo Mignatti. Fibroblast Growth Factor-2 (FGF-2) Induces Vascular Endothelial Growth Factor (VEGF) Expression in the Endothelial Cells of Forming Capillaries: an Autocrine Mechanism Contributing to Angiogenesis. Journal of Cell Biology, 141(7):1659–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/7/1659.

Sciaky:1997:GTT

[SPS^{+}97a] Noah Sciaky, John Presley, Carolyn Smith, Kristien J. M. Zaal, Nelson Cole, Jorge E. Moreira, Mark Terasaki, Eric Siggia, and Jennifer Lippincott-Schwartz. Golgi Tubule Traffic and the Effects of Brefeldin A Visualized in Living Cells. Journal of Cell Biology, 139(5):1137–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/5/1137.

Screaton:1997:CAH

[SPS97b] Robert A. Screaton, Linda Z. Penn, and Clifford P. Stanners. Carcinoembryonic Antigen, a Human Tumor Marker, Cooperates with Myc and Bcl-2 in Cellular Transformation. Journal of Cell Biology, 137(4):939–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/4/939.

Sun:1995:MAM

[SPT^{+}95] X. H. Sun, F. Protasi, M. Takahashi, H. Takeshima, D. G. Ferguson, and C. Franzini-Armstrong. Molecular architecture of membranes involved in excitation-contraction coupling of cardiac muscle. Journal of Cell Biology, 129(3):659–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/659.

Sanvito:1999:IIP

[SPV^{+}99] Francesca Sanvito, Simonetta Piatti, Antonello Villa, Mario Bossi, Giovanna Lucchini, Pier Carlo Marchisio, and Stefano Biffo. The β4 integrin interactor p27^{BBP} /eIF6 is an essential nuclear matrix protein involved in 60S ribosomal subunit assembly. Journal of Cell Biology, 144(5):823–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
5/823.

Sym:1995:ZIC

[SR95] M. Sym and G. S. Roeder. Zip1-induced changes in synaptonemal complex structure and polycomplex assembly. Journal of Cell Biology, 128(4):455–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/4/455.

Sundberg:1996:SBI

[SR96] C. Sundberg and K. Rubin. Stimulation of beta1 integrins on fibroblasts induces PDGF independent tyrosine phosphorylation of PDGF beta-receptors. Journal of Cell Biology, 132(4): 741–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/4/741.

Smith:1997:YRP

[SR97] Albert V. Smith and G. Shirleen Roeder. The Yeast Red1 Protein Localizes to the Cores of Meiotic Chromosomes. Journal of Cell Biology, 136(5):957–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/5/957.

Shiraga:1999:PMR

[SRA^{+}99] Masamichi Shiraga, Alec Ritchie, Sallouha Aidoudi, Veronique Baron, David Wilcox, Gilbert White, Belen Ybarrondo, George Murphy, Andrew Leavitt, and Sanford Shattil. Primary Megakaryocytes Reveal a Role for Transcription Factor Nf–E2 in Integrin α iibβ3 Signaling. Journal of Cell Biology, 147(7):1419–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/7/1419.

Straub:1997:AMM

[SRCC97] Volker Straub, Jill A. Rafael, Jeffrey S. Chamberlain, and Kevin P. Campbell. Animal Models for Muscular Dystrophy Show Different Patterns of Sarcolemmal Disruption. Journal of Cell Biology, 139(2):375–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/2/375.

Shulga:1996:VNT

[SRG^{+}96] N. Shulga, P. Roberts, Z. Gu, L. Spitz, M. M. Tabb, M. Nomura, and D. S. Goldfarb. In vivo nuclear transport kinetics in Saccharomyces cerevisiae: a role for heat shock protein 70 during targeting and translocation. Journal of Cell Biology, 135 (2):329–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/2/329.

Schmalzing:1995:IGB

[SRH^{+}95] G. Schmalzing, H. P. Richter, A. Hansen, W. Schwarz, I. Just, and K. Aktories. Involvement of the GTP binding protein Rho in constitutive endocytosis in Xenopus laevis oocytes. Journal of Cell Biology, 130(6):1319–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/6/1319.

Sleeman:1996:RCV

[SRH^{+}96] J. Sleeman, W. Rudy, M. Hofmann, J. Moll, P. Herrlich, and H. Ponta. Regulated clustering of variant CD44 proteins increases their hyaluronate binding capacity. Journal of Cell Biology, 135(4):1139–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/4/1139.

Saada:1996:GMC

[SRR96a] A. Saada, F. Reichert, and S. Rotshenker. Granulocyte macrophage colony stimulating factor produced in lesioned peripheral nerves induces the up-regulation of cell surface expression of MAC-2 by macrophages and Schwann cells. Journal of Cell Biology, 133(1):159–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/1/159.

Sevinsky:1996:LIP

[SRR96b] J. R. Sevinsky, L. V. Rao, and W. Ruf. Ligand-induced protease receptor translocation into caveolae: a mechanism for regulating cell surface proteolysis of the tissue factordependent coagulation pathway. Journal of Cell Biology, 133 (2):293–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/2/293.

Schmitz:1997:TTD

[SRR^{+}97] Holger Schmitz, Mary C. Reedy, Michael K. Reedy, Richard T. Tregear, and Kenneth A. Taylor. Tomographic Threedimensional Reconstruction of Insect Flight Muscle Partially Relaxed by AMPPNP and Ethylene Glycol. Journal of Cell Biology, 139(3):695–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/3/695.

Stoffler:1995:NMM

[SRRB95] H. E. Stöffler, C. Ruppert, J. Reinhard, and M. Bähler. A novel mammalian myosin I from rat with an SH3 domain localizes to Con A-inducible, F-actin-rich structures at cell– cell contacts. Journal of Cell Biology, 129(3):819–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/819.

Salas:1997:ASC

[SRV^{+}97] Pedro J. I. Salas, Marcelo L. Rodriguez, Ana L. Viciana, Dora E. Vega-Salas, and Hans-Peter Hauri. The Apical Submembrane Cytoskeleton Participates in the Organization of the Apical Pole in Epithelial Cells. Journal of Cell Biology, 137(2):359–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/2/359.

Schmidt:1995:CSR

[SS95a] E. E. Schmidt and U. Schibler. Cell size regulation, a mechanism that controls cellular RNA accumulation: consequences on regulation of the ubiquitous transcription factors Oct1 and NF–Y and the liver-enriched transcription factor DBP. Journal of Cell Biology, 128(4):467–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/4/467.

Sobel:1995:HDG

[SS95b] S. G. Sobel and M. Snyder. A highly divergent gamma-tubulin gene is essential for cell growth and proper microtubule organization in Saccharomyces cerevisiae. Journal of Cell Biology, 131(6):1775–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/6/1775.

Santos:1997:TCS

[SS97] Beatriz Santos and Michael Snyder. Targeting of Chitin Synthase 3 to Polarized Growth Sites in Yeast Requires Chs5p and Myo2p. Journal of Cell Biology, 136(1):95–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/136/1/95.

Spang:1998:RRT

[SS98] Anne Spang and Randy Schekman. Reconstitution of Retrograde Transport from the Golgi to the ER In Vitro. Journal of Cell Biology, 143(3):589–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/3/589.

Shukunami:1996:CDC

[SSA^{+}96] C. Shukunami, C. Shigeno, T. Atsumi, K. Ishizeki, F. Suzuki, and Y. Hiraki. Chondrogenic differentiation of clonal mouse embryonic cell line ATDC5 in vitro: differentiation-dependent gene expression of parathyroid hormone (PTH)/PTH-related peptide receptor. Journal of Cell Biology, 133(2):457–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
2/457.

Streuli:1995:LMT

[SSB^{+}95] C. H. Streuli, C. Schmidhauser, N. Bailey, P. Yurchenco, A. P. Skubitz, C. Roskelley, and M. J. Bissell. Laminin mediates tissue-specific gene expression in mammary epithelia. Journal of Cell Biology, 129(3):591–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/3/591.

Stronach:1996:TMS

[SSB96] B. E. Stronach, S. E. Siegrist, and M. C. Beckerle. Two musclespecific LIM proteins in Drosophila. Journal of Cell Biology, 134(5):1179–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/5/1179.

Setiadi:1998:ICD

[SSEM98] Hendra Setiadi, Gerald Sedgewick, Stanley L. Erlandsen, and Rodger P. McEver. Interactions of the Cytoplasmic Domain of P–Selectin with Clathrin-coated Pits Enhance Leukocyte Adhesion under Flow. Journal of Cell Biology, 142(3):859–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
142/3/859.

Steverding:1995:TBP

[SSF^{+}95] D. Steverding, Y. D. Stierhof, H. Fuchs, R. Tauber, and P. Overath. Transferrin-binding protein complex is the receptor for transferrin uptake in Trypanosoma brucei. Journal of Cell Biology, 131(5):1173–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/5/1173.

Suzuki:1998:MNM

[SSH^{+}98] Atsushi Suzuki, Yuki Sugiyama, Yukiko Hayashi, Nobuo Nyui, Michihiko Yoshida, Ikuya Nonaka, Sho ichi Ishiura, Kiichi Arahata, and Shigeo Ohno. MKBP, a Novel Member of the Small Heat Shock Protein Family, Binds and Activates the Myotonic Dystrophy Protein Kinase. Journal of Cell Biology, 140 (5):1113–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/5/1113.

Sakai:1997:DVT

[SSI^{+}97] Norio Sakai, Keiko Sasaki, Natsu Ikegaki, Yasuhito Shirai, Yoshitaka Ono, and Naoaki Saito. Direct Visualization of the Translocation of the γ-Subspecies of Protein Kinase C in Living Cells Using Fusion Proteins with Green Fluorescent Protein. Journal of Cell Biology, 139(6):1465–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
6/1465.

Slack:1995:CDN

[SSL^{+}95] R. S. Slack, I. S. Skerjanc, B. Lach, J. Craig, K. Jardine, and M. W. McBurney. Cells differentiating into neuroectoderm undergo apoptosis in the absence of functional retinoblastoma family proteins. Journal of Cell Biology, 129(3):779–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/779.

Screpanti:1995:EGF

[SSM^{+}95] I. Screpanti, S. Scarpa, D. Meco, D. Bellavia, L. Stuppia, L. Frati, A. Modesti, and A. Gulino. Epidermal growth factor promotes a neural phenotype in thymic epithelial cells and enhances neuropoietic cytokine expression. Journal of Cell Biology, 130(1):183–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/1/183.

Straight:1998:TLM

[SSM98] Aaron F. Straight, John W. Sedat, and Andrew W. Murray. Time-lapse Microscopy Reveals Unique Roles for Kinesins during Anaphase in Budding Yeast. Journal of Cell Biology, 143(3):687–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/3/687.

Sato:1996:ERL

[SSN96] M. Sato, K. Sato, and A. Nakano. Endoplasmic reticulum localization of Sec12p is achieved by two mechanisms: Rer1pdependent retrieval that requires the transmembrane domain and Rer1p-independent retention that involves the cytoplasmic domain. Journal of Cell Biology, 134(2):279–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/279.

Simon:1995:ADM

[SSP95] V. R. Simon, T. C. Swayne, and L. A. Pon. Actin-dependent mitochondrial motility in mitotic yeast and cell-free systems: identification of a motor activity on the mitochondrial surface. Journal of Cell Biology, 130(2):345–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/2/345.

Salomon:1997:RCL

[SSR^{+}97] Daniela Salomon, Paula A. Sacco, Sujata Guha Roy, Inbal Simcha, Keith R. Johnson, Margaret J. Wheelock, and Avri Ben-Ze’ev. Regulation of β-Catenin Levels and Localization by Overexpression of Plakoglobin and Inhibition of the Ubiquitin– Proteasome System. Journal of Cell Biology, 139(5):1325– ??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1325.

StCroix:1998:CDG

[SSR^{+}98] Brad St. Croix, Capucine Sheehan, Janusz W. Rak, Vivi Ann Flørenes, Joyce M. Slingerland, and Robert S. Kerbel. Ecadherin-dependent growth suppression is mediated by the cyclin-dependent kinase inhibitor p27^{KIP1}. Journal of Cell Biology, 142(2):557–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/2/557.

Shamu:1999:PUD

[SSRP99] Caroline E. Shamu, Craig M. Story, Tom A. Rapoport, and Hidde L. Ploegh. The Pathway of Us11-Dependent Degradation of Mhc Class I Heavy Chains Involves a Ubiquitin– Conjugated Intermediate. Journal of Cell Biology, 147(1): 45–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/1/45.

Smirnova:1998:HDR

[SSRvdB98] Elena Smirnova, Dixie-Lee Shurland, Sergey N. Ryazantsev, and Alexander M. van der Bliek. A Human Dynamin-related Protein Controls the Distribution of Mitochondria. Journal of Cell Biology, 143(2):351–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/2/351.

Staddon:1995:PPR

[SSS^{+}95] J. M. Staddon, C. Smales, C. Schulze, F. S. Esch, and L. L. Rubin. p120, a p120-related protein (p100), and the cadherin/ catenin complex. Journal of Cell Biology, 130(2):369–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/369.

Simcha:1998:DNT

[SSS^{+}98] Inbal Simcha, Michael Shtutman, Daniela Salomon, Jacob Zhurinsky, Einat Sadot, Benjamin Geiger, and Avri Ben-Ze’ev. Differential Nuclear Translocation and Transactivation Potential of β-Catenin and Plakoglobin. Journal of Cell Biology, 141 (6):1433–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/6/1433.

Schleiff:1999:DMI

[SSS99a] Enrico Schleiff, John R. Silvius, and Gordon C. Shore. Direct Membrane Insertion of Voltage-dependent Anion-selective Channel Protein Catalyzed by Mitochondrial Tom20. Journal of Cell Biology, 145(5):973–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/5/973.

Shiratori:1999:DED

[SSS^{+}99b] Miwa Shiratori, Sakae Sakamoto, Noriyuki Suzuki, Yoshiki Tokutake, Yoichi Kawabe, Takemi Enomoto, Masanobu Sugimoto, Makoto Goto, Takehisa Matsumoto, and Yasuhiro Furuichi. Detection by Epitope-defined Monoclonal Antibodies of Werner DNA Helicases in the Nucleoplasm and Their Upregulation by Cell Transformation and Immortalization. Journal of Cell Biology, 144(1):1–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/1/1.

Silberstein:1998:RDH

[SSSG98] Susana Silberstein, Gabriel Schlenstedt, Pam A. Silver, and Reid Gilmore. A Role for the DnaJ Homologue Scj1p in Protein Folding in the Yeast Endoplasmic Reticulum. Journal of Cell Biology, 143(4):921–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/4/921.

Schroder:1995:GLY

[SSSKR95] S. Schröder, F. Schimmöller, B. Singer-Krüger, and H. Riezman. The Golgi-localization of yeast Emp47p depends on its di-lysine motif but is not affected by the ret1-1 mutation in alpha–COP. Journal of Cell Biology, 131(4):895–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/895.

Su:1997:THI

[SSSS97] Yuan Su, Yufang Shi, Melissa A. Stolow, and Yun-Bo Shi. Thyroid Hormone Induces Apoptosis in Primary Cell Cultures of Tadpole Intestine: Cell Type Specificity and Effects of Extracellular Matrix. Journal of Cell Biology, 139(6):1533–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/6/1533.

Sterling:1998:COP

[SSV98] Hyacinth Sterling, Charles Saginario, and Agnès Vignery. CD44 Occupancy Prevents Macrophage Multinucleation. Journal of Cell Biology, 143(3):837–??, November 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/143/3/
837.

Sydor:1996:TVP

[SSW^{+}96] A. M. Sydor, A. L. Su, F. S. Wang, A. Xu, and D. G. Jay. Talin and vinculin play distinct roles in filopodial motility in the neuronal growth cone. Journal of Cell Biology, 134(5): 1197–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/5/1197.

Saurin:1998:HPG

[SSW^{+}98] Andrew J. Saurin, Carol Shiels, Jill Williamson, David P. E. Satijn, Arie P. Otte, Denise Sheer, and Paul S. Freemont. The Human Polycomb Group Complex Associates with Pericentromeric Heterochromatin to Form a Novel Nuclear Domain. Journal of Cell Biology, 142(4):887–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/4/887.

Sakai:1996:APP

[SSY^{+}96] Y. Sakai, A. Saiganji, H. Yurimoto, K. Takabe, H. Saiki, and N. Kato. The absence of Pmp47, a putative yeast peroxisomal transporter, causes a defect in transport and folding of a specific matrix enzyme. Journal of Cell Biology, 134(1):37–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/1/37.

Schuger:1997:LCS

[SSZ^{+}97] Lucia Schuger, Amy P. N. Skubitz, Jun Zhang, Lydia Sorokin, and Li He. Laminin α1 Chain Synthesis in the Mouse Developing Lung: Requirement for Epithelial–Mesenchymal Contact and Possible Role in Bronchial Smooth muscle Development. Journal of Cell Biology, 139(2):553–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/2/553.

Scott:1996:NCP

[ST96] S. V. Scott and S. M. Theg. A new chloroplast protein import intermediate reveals distinct translocation machineries in the two envelope membranes: energetics and mechanistic implications. Journal of Cell Biology, 132(1):63–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/132/1/63.

Stearns:1997:MFK

[Ste97] Tim Stearns. Motoring to the Finish: Kinesin and Dynein Work Together to Orient the Yeast Mitotic Spindle. Journal of Cell Biology, 138(5):957–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/5/957.

Shah:1998:MBS

[STF98] Sundeep Shah, Stuart Tugendreich, and Douglass Forbes. Major Binding Sites for the Nuclear Import Receptor Are the Internal Nucleoporin Nup153 and the Adjacent Nuclear Filament Protein Tpr. Journal of Cell Biology, 141(1):31–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
1/31.

Samso:1999:TDL

[STG^{+}99] Montserrat Samsó, Ramon Trujillo, Georgina B. Gurrola, Hector H. Valdivia, and Terence Wagenknecht. Three-dimensional Location of the Imperatoxin a Binding Site on the Ryanodine Receptor. Journal of Cell Biology, 146(2):493–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
2/493.

Samejima:1998:TCD

[STK^{+}98] Kumiko Samejima, Shigenobu Toné, Timothy J. Kottke, Masato Enari, Hideki Sakahira, Carol A. Cooke, Françoise Durrieu, Luis M. Martins, Shigekazu Nagata, Scott H. Kaufmann, and William C. Earnshaw. Transition from Caspasedependent to Caspase-independent Mechanisms at the Onset of Apoptotic Execution. Journal of Cell Biology, 143(1): 225–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/1/225.

Sander:1999:RDR

[StKvD^{+}99] Eva E. Sander, Jean P. ten Klooster, Sanne van Delft, Rob A. van der Kammen, and John G. Collard. Rac Downregulates Rho Activity. Journal of Cell Biology, 147(5):1009–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
147/5/1009.

Saitoh:1996:AMF

[STN^{+}96] S. Saitoh, K. Takahashi, K. Nabeshima, Y. Yamashita, Y. Nakaseko, A. Hirata, and M. Yanagida. Aberrant mitosis in fission yeast mutants defective in fatty acid synthetase and acetyl CoA carboxylase. Journal of Cell Biology, 134(4): 949–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/949.

Smith:1996:TRD

[STP96] G. A. Smith, J. A. Theriot, and D. A. Portnoy. The tandem repeat domain in the Listeria monocytogenes ActA protein controls the rate of actin-based motility, the percentage of moving bacteria, and the localization of vasodilator-stimulated phosphoprotein and profilin. Journal of Cell Biology, 135(3): 647–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/647.

Sluder:1995:NEB

[STRM95] G. Sluder, E. A. Thompson, C. L. Rieder, and F. J. Miller. Nuclear envelope breakdown is under nuclear not cytoplasmic control in sea urchin zygotes. Journal of Cell Biology, 129 (6):1447–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1447.

Sechler:1996:ARF

[STS96] J. L. Sechler, Y. Takada, and J. E. Schwarzbauer. Altered rate of fibronectin matrix assembly by deletion of the first type III repeats. Journal of Cell Biology, 134(2):573–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/573.

Sekimata:1996:DPK

[STT^{+}96] M. Sekimata, K. Tsujimura, J. Tanaka, Y. Takeuchi, N. Inagaki, and M. Inagaki. Detection of protein kinase activity specifically activated at metaphase-anaphase transition. Journal of Cell Biology, 132(4):635–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/132/4/635.

Szilard:1995:PYY

[STVR95] R. K. Szilard, V. I. Titorenko, M. Veenhuis, and R. A. Rachubinski. Pay32p of the yeast Yarrowia lipolytica is an intraperoxisomal component of the matrix protein translocation machinery. Journal of Cell Biology, 131(6):1453–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
6/1453.

Strumpf:1998:KNC

[SV98] Dan Strumpf and Talila Volk. Kakapo, a Novel Cytoskeletalassociated Protein Is Essential for the Restricted Localization of the Neuregulin-like Factor, Vein, at the Muscle–Tendon Junction Site. Journal of Cell Biology, 143(5):1259–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
143/5/1259.

Svitkina:1996:PSM

[SVB96] T. M. Svitkina, A. B. Verkhovsky, and G. G. Borisy. Plectin sidearms mediate interaction of intermediate filaments with microtubules and other components of the cytoskeleton. Journal of Cell Biology, 135(4):991–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/4/991.

Schumacher:1997:CAL

[SVB^{+}97] Stefan Schumacher, Hansjürgen Volkmer, Fritz Buck, Albrecht Otto, Attila Tárnok, Siegfried Roth, and Fritz G. Rathjen. Chicken Acidic Leucine-rich EGF-like Domain Containing Brain Protein (CALEB), a Neural Member of the EGF Family of Differentiation Factors, Is Implicated in Neurite Formation. Journal of Cell Biology, 136(4):895–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
4/895.

Steinfeld:1996:SFG

[SVD96] R. Steinfeld, H. Van Den Berghe, and G. David. Stimulation of fibroblast growth factor receptor-1 occupancy and signaling by cell surface-associated syndecans and glypican. Journal of Cell Biology, 133(2):405–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/2/405.

Sander:1998:MDT

[SvDtK^{+}98] Eva E. Sander, Sanne van Delft, Jean P. ten Klooster, Tim Reid, Rob A. van der Kammen, Frits Michiels, and John G. Collard. Matrix-dependent Tiam1/Rac signaling in epithelial cells promotes either cell–cell adhesion or cell migration and is regulated by phosphatidylinositol 3-kinase. Journal of Cell Biology, 143(5):1385–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/5/1385.

Scheiffele:1998:CEP

[SVF^{+}98] P. Scheiffele, P. Verkade, A. M. Fra, H. Virta, K. Simons, and E. Ikonen. Caveolin-1 and -2 in the Exocytic Pathway of MDCK Cells. Journal of Cell Biology, 140(4):795–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
4/795.

Shapiro:1998:AME

[SVH^{+}98] Paul S. Shapiro, Eugeni Vaisberg, Alan J. Hunt, Nicholas S. Tolwinski, Anne M. Whalen, J. Richard McIntosh, and Natalie G. Ahn. Activation of the MKK/ERK pathway during somatic cell mitosis: Direct interactions of active ERK with kinetochores and regulation of the mitotic 3F3/2 phosphoantigen. Journal of Cell Biology, 142(6):1533–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1533.

Smyth:1999:ABM

[SVM^{+}99] Neil Smyth, H. Seda Vatansever, Patricia Murray, Michael Meyer, Christian Frie, Mats Paulsson, and David Edgar. Absence of Basement Membranes after Targeting the LAMC1 Gene Results in Embryonic Lethality Due to Failure of Endoderm Differentiation. Journal of Cell Biology, 144(1): 151–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/1/151.

Svitkina:1997:AAM

[SVMB97] Tatyana M. Svitkina, Alexander B. Verkhovsky, Kyle M. Mc-Quade, and Gary G. Borisy. Analysis of the Actin–Myosin II System in Fish Epidermal Keratocytes: Mechanism of Cell Body Translocation. Journal of Cell Biology, 139(2): 397–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/397.

Shelby:1997:ACC

[SVS97] Richard D. Shelby, Omid Vafa, and Kevin F. Sullivan. Assembly of CENP–A into Centromeric Chromatin Requires a Cooperative Array of Nucleosomal DNA Contact Sites. Journal of Cell Biology, 136(3):501–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/3/501.

Swartzman:1996:PGP

[SVT96] E. E. Swartzman, M. N. Viswanathan, and J. Thorner. The PAL1 gene product is a peroxisomal ATP-binding cassette transporter in the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 132(4):549–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/4/549.

Soullam:1995:SSF

[SW95] B. Soullam and H. J. Worman. Signals and structural features involved in integral membrane protein targeting to the inner nuclear membrane. Journal of Cell Biology, 130(1):15–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
1/15.

Spath:1998:HNF

[SW98] Gerald F. Späth and Mary C. Weiss. Hepatocyte Nuclear Factor 4 Provokes Expression of Epithelial Marker Genes, Acting As a Morphogen in Dedifferentiated Hepatoma Cells. Journal of Cell Biology, 140(4):935–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/4/935.

Shorter:1999:RVT

[SW99] James Shorter and Graham Warren. A role for the vesicle tethering protein, P115, in the post-mitotic stacking of reassembling Golgi cisternae in a cell-free system. Journal of Cell Biology, 146(1):57–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/1/57.

Sachs:1996:MMA

[SWB^{+}96] M. Sachs, K. M. Weidner, V. Brinkmann, I. Walther, A. Obermeier, A. Ullrich, and W. Birchmeier. Motogenic and morphogenic activity of epithelial receptor tyrosine kinases. Journal of Cell Biology, 133(5):1095–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/5/1095.

Sonnichsen:1996:SCC

[SWC^{+}96] B. Sönnichsen, R. Watson, H. Clausen, T. Misteli, and G. Warren. Sorting by COP I-coated vesicles under interphase and mitotic conditions. Journal of Cell Biology, 134(6):1411– ??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1411.

Shibata:1998:URD

[SWD^{+}98] A. Shibata, M. V. Wright, S. David, L. McKerracher, P. E. Braun, and S. B. Kater. Unique Responses of Differentiating Neuronal Growth Cones to Inhibitory Cues Presented by Oligodendrocytes. Journal of Cell Biology, 142(1):191–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
1/191.

Starr:1998:ZHR

[SWHG98] Daniel A. Starr, Byron C. Williams, Thomas S. Hays, and Michael L. Goldberg. ZW10 Helps Recruit Dynactin and Dynein to the Kinetochore. Journal of Cell Biology, 142(3): 763–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/3/763.

Sawamoto:1999:DRG

[SWK^{+}99] Kazunobu Sawamoto, Per Winge, Shinya Koyama, Yuki Hirota, Chiharu Yamada, Sachiyo Miyao, Shingo Yoshikawa, Ming hao Jin, Akira Kikuchi, and Hideyuki Okano. The Drosophila Ral GTPase Regulates Developmental Cell Shape Changes through the Jun NH_{2}-terminal Kinase Pathway. Journal of Cell Biology, 146(2):361–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/146/2/361.

Starr:1997:CCK

[SWL^{+}97] Daniel A. Starr, Byron C. Williams, Zexiao Li, Bijan Etemad-Moghadam, R. Kelly Dawe, and Michael L. Goldberg. Conservation of the centromere/kinetochore protein ZW10. Journal of Cell Biology, 138(6):1289–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/6/1289.

Schafer:1998:VMA

[SWM^{+}98] Dorothy A. Schafer, Matthew D. Welch, Laura M. Machesky, Paul C. Bridgman, Shelley M. Meyer, and John A. Cooper. Visualization and Molecular Analysis of Actin Assembly in Living Cells. Journal of Cell Biology, 143(7):1919–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
7/1919.

Signor:1999:RCD

[SWO^{+}99] Dawn Signor, Karen P. Wedaman, Jose T. Orozco, Noelle D. Dwyer, Cornelia I. Bargmann, Lesilee S. Rose, and Jonathan M. Scholey. Role of a Class Dhc1b Dynein in Retrograde Transport of Ift Motors and Ift Raft Particles along Cilia, but Not Dendrites, in Chemosensory Neurons of Living Caenorhabditis elegans. Journal of Cell Biology, 147(3):519–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
3/519.

Sakurai:1996:BPS

[SWR^{+}96] T. Sakurai, R. E. Westenbroek, J. Rettig, J. Hell, and W. A. Catterall. Biochemical properties and subcellular distribution of the BI and rbA isoforms of alpha 1A subunits of brain calcium channels. Journal of Cell Biology, 134(2):511–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
2/511.

Storrie:1998:RGR

[SWR^{+}98] Brian Storrie, Jamie White, Sabine Röttger, Ernst H. K. Stelzer, Tatsuo Suganuma, and Tommy Nilsson. Recycling of Golgi-resident Glycosyltransferases through the ER Reveals a Novel Pathway and Provides an Explanation for Nocodazoleinduced Golgi Scattering. Journal of Cell Biology, 143(6): 1505–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/6/1505.

Scherthan:1996:CTM

[SWS^{+}96] H. Scherthan, S. Weich, H. Schwegler, C. Heyting, M. Härle, and T. Cremer. Centromere and telomere movements during early meiotic prophase of mouse and man are associated with the onset of chromosome pairing. Journal of Cell Biology, 134(5):1109–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/5/1109.

Sechi:1997:ICT

[SWS97a] Antonio S. Sechi, Jürgen Wehland, and J. Victor Small. The Isolated Comet Tail Pseudopodium of Listeria monocytogenes: a Tail of Two Actin Filament Populations, Long and Axial and Short and Random. Journal of Cell Biology, 137(1):155–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/1/155.

Sesaki:1997:CAM

[SWS97b] Hiromi Sesaki, Estella F. S. Wong, and Chi-Hung Siu. The Cell Adhesion Molecule DdCAD-1 in Dictyostelium Is Targeted to the Cell Surface by a Nonclassical Transport Pathway Involving Contractile Vacuoles. Journal of Cell Biology, 138 (4):939–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/4/939.

Soede:1998:ZTK

[SWVR98] Ron D. M. Soede, Yvonne M. Wijnands, Ioana Van Kouteren-Cobzaru, and Ed Roos. ZAP-70 tyrosine kinase is required for LFA-1-dependent T cell migration. Journal of Cell Biology, 142(5):1371–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/5/1371.

Shutt:1995:PPR

[SWW^{+}95] D. C. Shutt, D. Wessels, K. Wagenknecht, A. Chandrasekhar, A. L. Hitt, E. J. Luna, and D. R. Soll. Ponticulin plays a role in the positional stabilization of pseudopods. Journal of Cell Biology, 131(6):1495–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1495.

Shepard:1999:YDL

[SY99] Kelly A. Shepard and Michael P. Yaffe. The Yeast Dynaminlike Protein, Mgm1p, Functions on the Mitochondrial Outer Membrane to Mediate Mitochondrial Inheritance. Journal of Cell Biology, 144(4):711–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/144/4/711.

Smart:1995:HRC

[SYA95] E. J. Smart, Y. S. Ying, and R. G. Anderson. Hormonal regulation of caveolae internalization. Journal of Cell Biology, 131(4):929–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/131/4/929.

Sharp:1995:TDM

[SYB95] D. J. Sharp, W. Yu, and P. W. Baas. Transport of dendritic microtubules establishes their nonuniform polarity orientation. Journal of Cell Biology, 130(1):93–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/130/1/93.

Sharp:1997:IMA

[SYF^{+}97] David J. Sharp, Wenqian Yu, Lotfi Ferhat, Ryoko Kuriyama, David C. Rueger, and Peter W. Baas. Identification of a Microtubule-associated Motor Protein Essential for Dendritic Differentiation. Journal of Cell Biology, 138(4):833–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
4/833.

Shaw:1997:AMD

[SYM^{+}97] Sidney L. Shaw, Elaine Yeh, Paul Maddox, E. D. Salmon, and Kerry Bloom. Astral Microtubule Dynamics in Yeast: a Microtubule-based Searching Mechanism for Spindle Orientation and Nuclear Migration into the Bud. Journal of Cell Biology, 139(4):985–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/4/985.

Suzuki:1995:MAB

[SYO95] A. Suzuki, M. Yoshida, and E. Ozawa. Mammalian alpha 1and beta 1-syntrophin bind to the alternative splice-prone region of the dystrophin COOH terminus. Journal of Cell Biology, 128(3):373–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/3/373.

Savagner:1997:ZFP

[SYT97] Pierre Savagner, Kenneth M. Yamada, and Jean Paul Thiery. The Zinc–Finger Protein Slug Causes Desmosome Dissociation, an Initial and Necessary Step for Growth Factor–induced Epithelial–Mesenchymal Transition. Journal of Cell Biology, 137(6):1403–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/6/1403.

Staudinger:1995:PPB

[SZB^{+}95] J. Staudinger, J. Zhou, R. Burgess, S. J. Elledge, and E. N. Olson. PICK1: a perinuclear binding protein and substrate for protein kinase C isolated by the yeast two-hybrid system. Journal of Cell Biology, 128(3):263–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/3/263.

Scotland:1998:NSD

[SZBB98] Paula Scotland, Daixing Zhou, Helene Benveniste, and Vann Bennett. Nervous System Defects of Ankyrin_{B} (-/-) Mice Suggest Functional Overlap between the Cell Adhesion Molecule L1 and 440-kD Ankyrin_{B} in Premyelinated Axons. Journal of Cell Biology, 143(5):1305–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/5/1305.

Sakai:1998:MIF

[SZFM98] Takao Sakai, Qinghong Zhang, Reinhard Fässler, and Deane F. Mosher. Modulation of β1A Integrin Functions by Tyrosine Residues in the β1 Cytoplasmic Domain. Journal of Cell Biology, 141(2):527–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/2/527.

Skelton:1998:GPB

[SZNS98] Timothy P. Skelton, Chunxun Zeng, Aaron Nocks, and Ivan Stamenkovic. Glycosylation Provides Both Stimulatory and Inhibitory Effects on Cell Surface and Soluble CD44 Binding to Hyaluronan. Journal of Cell Biology, 140(2):431–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
2/431.

Todorov:1995:BHM

[TAK95] I. T. Todorov, A. Attaran, and S. E. Kearsey. BM28, a human member of the MCM2-3-5 family, is displaced from chromatin during DNA replication. Journal of Cell Biology, 129 (6):1433–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1433.

Trielli:1996:DTD

[TALM96] M. O. Trielli, P. R. Andreassen, F. B. Lacroix, and R. L. Margolis. Differential Taxol-dependent arrest of transformed and nontransformed cells in the G1 phase of the cell cycle, and specific-related mortality of transformed cells. Journal of Cell Biology, 135(3):689–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/3/689.

Titov:1999:KKP

[TB99] Anton A. Titov and Günter Blobel. The karyopherin Kap122p/Pdr6p imports both subunits of the transcription factor IIA into the nucleus. Journal of Cell Biology, 147(2): 235–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/2/235.

Tuvia:1999:ABR

[TBD^{+}99] Shmuel Tuvia, Mona Buhusi, Lydia Davis, Mary Reedy, and Vann Bennett. Ankyrin-b Is Required for Intracellular Sorting of Structurally Diverse Ca^{2+} Homeostasis Proteins. Journal of Cell Biology, 147(5):995–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/147/5/995.

Turner:1999:PLM

[TBP^{+}99] Christopher E. Turner, Michael C. Brown, Joseph A. Perrotta, M. C. Riedy, Sotiris N. Nikolopoulos, A. Rosa McDonald, Shubha Bagrodia, Sheila Thomas, and Phillip S. Leventhal. Paxillin LD4 Motif Binds PAK and PIX through a Novel 95kD Ankyrin Repeat, ARF–GAP Protein: a Role in Cytoskeletal Remodeling. Journal of Cell Biology, 145(4):851–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
4/851.

Traub:1996:ACC

[TBR^{+}96] L. M. Traub, S. I. Bannykh, J. E. Rodel, M. Aridor, W. E. Balch, and S. Kornfeld. AP-2-containing clathrin coats assemble on mature lysosomes. Journal of Cell Biology, 135(6): 1801–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1801.

Tansey:1996:AHR

[TCM96] M. G. Tansey, G. C. Chu, and J. P. Merlie. ARIA/HRG regulates AChR epsilon subunit gene expression at the neuromuscular synapse via activation of phosphatidylinositol 3kinase and Ras/MAPK pathway. Journal of Cell Biology, 134 (2):465–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/465.

Tuxworth:1997:DRR

[TCM^{+}97] Richard I. Tuxworth, Janet L. Cheetham, Laura M. Machesky, George B. Spiegelmann, Gerald Weeks, and Robert H. Insall. Dictyostelium RasG Is Required for Normal Motility and Cytokinesis, But Not Growth. Journal of Cell Biology, 138 (3):605–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/3/605.

Tassin:1998:CHH

[TCMB98] Anne-Marie Tassin, Claude Celati, Mohammed Moudjou, and Michel Bornens. Characterization of the Human Homologue of the Yeast Spc98p and Its Association with γ-Tubulin. Journal of Cell Biology, 141(3):689–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/3/689.

Tilney:1996:FAB

[TCSG96] L. G. Tilney, P. Connelly, S. Smith, and G. M. Guild. F-actin bundles in Drosophila bristles are assembled from modules composed of short filaments. Journal of Cell Biology, 135(5): 1291–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/5/1291.

Tilney:1998:WTD

[TCV^{+}98] Lewis G. Tilney, Patricia S. Connelly, Kelly A. Vranich, Michael K. Shaw, and Gregory M. Guild. Why Are Two Different Cross-linkers Necessary for Actin Bundle Formation In Vivo and What Does Each Cross-link Contribute? Journal of Cell Biology, 143(1):121–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/1/121.

Taooka:1999:IMA

[TCYS99] Yasuyuki Taooka, John Chen, Ted Yednock, and Dean Sheppard. The Integrin α9β1 Mediates Adhesion to Activated Endothelial Cells and Transendothelial Neutrophil Migration through Interaction with Vascular Cell Adhesion Molecule-1. Journal of Cell Biology, 145(2):413–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/2/413.

Taverna:1998:DMM

[TDR^{+}98] Daniela Taverna, Marie-Helene Disatnik, Helen Rayburn, Roderick T. Bronson, Joy Yang, Thomas A. Rando, and Richard O. Hynes. Dystrophic Muscle in Mice Chimeric for Expression of α5 Integrin. Journal of Cell Biology, 143(3): 849–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/3/849.

Trilla:1999:CNP

[TDR99] Jose A. Trilla, Angel Durán, and Cesar Roncero. Chs7p, a New Protein Involved in the Control of Protein Export from the Endoplasmic Reticulum that Is Specifically Engaged in the Regulation of Chitin Synthesis in Saccharomyces cerevisiae. Journal of Cell Biology, 145(6):1153–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/6/1153.

Tremble:1995:CNS

[TDW95] P. Tremble, C. H. Damsky, and Z. Werb. Components of the nuclear signaling cascade that regulate collagenase gene expression in response to integrin-derived signals. Journal of Cell Biology, 129(6):1707–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/6/1707.

Tu:1995:OHN

[TEL^{+}95] P. H. Tu, G. Elder, R. A. Lazzarini, D. Nelson, J. Q. Trojanowski, and V. M. Lee. Overexpression of the human NFM subunit in transgenic mice modifies the level of endogenous NFL and the phosphorylation state of NFH subunits. Journal of Cell Biology, 129(6):1629–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/6/1629.

Tao:1996:BCA

[TET^{+}96] Y. S. Tao, R. A. Edwards, B. Tubb, S. Wang, J. Bryan, and P. D. McCrea. beta-catenin associates with the actin-bundling protein fascin in a noncadherin complex. Journal of Cell Biology, 134(5):1271–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/5/1271.

Turowski:1995:DMA

[TFF^{+}95] P. Turowski, A. Fernandez, B. Favre, N. J. Lamb, and B. A. Hemmings. Differential methylation and altered conformation of cytoplasmic and nuclear forms of protein phosphatase 2A during cell cycle progression. Journal of Cell Biology, 129 (2):397–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/2/397.

Trotter:1999:ASR

[TFS^{+}99] Kevin W. Trotter, Iain D. C. Fraser, Gregory K. Scott, M. Jackson Stutts, John D. Scott, and Sharon L. Milgram. Alternative Splicing Regulates the Subcellular Localization of a–Kinase Anchoring Protein 18 Isoforms. Journal of Cell Biology, 147(7):1481–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/7/1481.

Tuma:1999:EAE

[TFYH99] Pamela L. Tuma, Catherine M. Finnegan, Ji-Hyun Yi, and Ann L. Hubbard. Evidence for Apical Endocytosis in Polarized Hepatic Cells: Phosphoinositide 3-Kinase Inhibitors Lead to the Lysosomal Accumulation of Resident Apical Plasma Membrane Proteins. Journal of Cell Biology, 145(5):1089–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
145/5/1089.

Taniura:1995:CBS

[TGG95] H. Taniura, C. Glass, and L. Gerace. A chromatin binding site in the tail domain of nuclear lamins that interacts with core histones. Journal of Cell Biology, 131(1):33–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
1/33.

Tanaka:1997:NMR

[TGGB97] Elly M. Tanaka, Alexander A. F. Gann, Phillip B. Gates, and Jeremy P. Brockes. Newt Myotubes Reenter the Cell Cycle by Phosphorylation of the Retinoblastoma Protein. Journal of Cell Biology, 136(1):155–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/1/155.

Trombitas:1998:TES

[TGL^{+}98] Karoly Trombitás, Marion Greaser, Siegfried Labeit, Jian-Ping Jin, Miklós Kellermayer, Michiel Helmes, and Henk Granzier. Titin Extensibility In Situ: Entropic Elasticity of Permanently Folded and Permanently Unfolded Molecular Segments. Journal of Cell Biology, 140(4):853–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/4/853.

Takenaka:1997:MKR

[TGN97] Katsuya Takenaka, Yukiko Gotoh, and Eisuke Nishida. MAP Kinase Is Required for the Spindle Assembly Checkpoint but Is Dispensable for the Normal M Phase Entry and Exit in Xenopus Egg Cell Cycle Extracts. Journal of Cell Biology, 136 (5):1091–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/5/1091.

Tian:1997:GIX

[TGTL97] Jingdong Tian, Hui Gong, Gerald H. Thomsen, and William J. Lennarz. Gamete Interactions in Xenopus laevis: Identification of Sperm Binding Glycoproteins in the Egg Vitelline Envelope. Journal of Cell Biology, 136(5):1099–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
5/1099.

Tang:1999:EEC

[TGW99] Shaoqing Tang, Yunling Gao, and J. Anthony Ware. Enhancement of Endothelial Cell Migration and in Vitro Tube Formation by Tap20, a Novel β5 Integrin–Modulating, Pkcθ-Dependent Protein. Journal of Cell Biology, 147(5):1073– ??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/5/1073.

Tatu:1997:IBN

[TH97] Utpal Tatu and Ari Helenius. Interactions between Newly Synthesized Glycoproteins, Calnexin and a Network of Resident Chaperones in the Endoplasmic Reticulum. Journal of Cell Biology, 136(3):555–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/3/555.

Tachibana:1999:RTS

[TH99] Isao Tachibana and Martin E. Hemler. Role of Transmembrane 4 Superfamily (Tm4sf) Proteins Cd9 and Cd81 in Muscle Cell Fusion and Myotube Maintenance. Journal of Cell Biology, 146(4):893–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/4/893.

Theriot:1997:ATA

[The97] Julie A. Theriot. Accelerating on a treadmill: ADF/cofilin promotes rapid actin filament turnover in the dynamic cytoskeleton. Journal of Cell Biology, 136(6):1165–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
6/1165.

Tanaka:1995:RMD

[THK95] E. Tanaka, T. Ho, and M. W. Kirschner. The role of microtubule dynamics in growth cone motility and axonal growth. Journal of Cell Biology, 128(1):139–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/1/139.

Taylor:1998:HHB

[THM98] Stephen S. Taylor, Edward Ha, and Frank McKeon. The human homologue of Bub3 is required for kinetochore localization of Bub1 and a Mad3/Bub1-related protein kinase. Journal of Cell Biology, 142(1):1–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/142/1/1.

Tan:1996:SND

[THP96] P. K. Tan, J. P. Howard, and G. S. Payne. The sequence NPFXD defines a new class of endocytosis signal in Saccharomyces cerevisiae. Journal of Cell Biology, 135(6):1789– ??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/6/1789.

Tange:1998:NFY

[THS^{+}98] Yoshie Tange, Tetsuya Horio, Mizuki Shimanuki, Da-Qiao Ding, Yasushi Hiraoka, and Osami Niwa. A Novel Fission Yeast Gene, tht1^{+}, Is Required for the Fusion of Nuclear Envelopes during Karyogamy. Journal of Cell Biology, 140(2): 247–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/2/247.

Takei:1995:SDR

[THT^{+}95] Y. Takei, A. Harada, S. Takeda, K. Kobayashi, S. Terada, T. Noda, T. Takahashi, and N. Hirokawa. Synapsin I deficiency results in the structural change in the presynaptic terminals in the murine nervous system. Journal of Cell Biology, 131(6): 1789–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1789.

Theesfeld:1999:RAS

[TIBL99] Chandra L. Theesfeld, Javier E. Irazoqui, Kerry Bloom, and Daniel J. Lew. The Role of Actin in Spindle Orientation Changes during the Saccharomyces cerevisiae Cell Cycle. Journal of Cell Biology, 146(5):1019–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
5/1019.

Takebayashi:1995:HGF

[TIJ^{+}95] T. Takebayashi, M. Iwamoto, A. Jikko, T. Matsumura, M. Enomoto-Iwamoto, F. Myoukai, E. Koyama, T. Yamaai, K. Matsumoto, and T. Nakamura. Hepatocyte growth factor/scatter factor modulates cell motility, proliferation, and proteoglycan synthesis of chondrocytes. Journal of Cell Biology, 129(5):1411–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/5/1411.

Tanaka:1995:RMG

[TK95] E. Tanaka and M. W. Kirschner. The role of microtubules in growth cone turning at substrate boundaries. Journal of Cell Biology, 128(1):127–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/1/127.

Terada:1997:HDH

[TKBM97] Kazutoyo Terada, Masaki Kanazawa, Bernd Bukau, and Masataka Mori. The Human DnaJ Homologue dj2 Facilitates Mitochondrial Protein Import and Luciferase Refolding. Journal of Cell Biology, 139(5):1089–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/5/1089.

Takei:1997:DDN

[TKH^{+}97] Yosuke Takei, Satoru Kondo, Akihiro Harada, Satomi Inomata, Tetsuo Noda, and Nobutaka Hirokawa. Delayed Development of Nervous System in Mice Homozygous for Disrupted Microtubule-associated Protein 1B (MAP1B) Gene. Journal of Cell Biology, 137(7):1615–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/7/1615.

Tsumaki:1996:SCR

[TKM^{+}96] N. Tsumaki, T. Kimura, Y. Matsui, K. Nakata, and T. Ochi. Separable cis-regulatory elements that contribute to tissueand site-specific alpha 2(XI) collagen gene expression in the embryonic mouse cartilage. Journal of Cell Biology, 134(6): 1573–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1573.

Tagawa:1995:CCD

[TKVC95] T. Tagawa, T. Kuroki, P. K. Vogt, and K. Chida. The cell cycle-dependent nuclear import of v–Jun is regulated by phosphorylation of a serine adjacent to the nuclear localization signal. Journal of Cell Biology, 130(2):255–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/255.

Takahashi:1997:MEP

[TKY^{+}97] Noriko Takahashi, Takashi Kadowaki, Yoshio Yazaki, Yasushi Miyashita, and Haruo Kasai. Multiple Exocytotic Pathways in Pancreatic β Cells. Journal of Cell Biology, 138(1):55–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
1/55.

Tuvia:1998:MFM

[TLBK98] Shmuel Tuvia, Shlomo Levin, Arkady Bitler, and Rafi Korenstein. Mechanical Fluctuations of the Membrane–Skeleton Are Dependent on F–Actin ATPase in Human Erythrocytes. Journal of Cell Biology, 141(7):1551–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/7/1551.

Timmusk:1995:IBD

[TLF^{+}95] T. Timmusk, U. Lendahl, H. Funakoshi, E. Arenas, H. Persson, and M. Metsis. Identification of brain-derived neurotrophic factor promoter regions mediating tissue-specific, axotomy-, and neuronal activity-induced expression in transgenic mice. Journal of Cell Biology, 128(1):185–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/1/185.

Tian:1997:TSE

[TLF^{+}97] Guoling Tian, Sally A. Lewis, Becket Feierbach, Timothy Stearns, Heidi Rommelaere, Christophe Ampe, and Nicholas J. Cowan. Tubulin Subunits Exist in an Activated Conformational State Generated and Maintained by Protein Cofactors. Journal of Cell Biology, 138(4):821–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/138/4/821.

Takei:1996:SVC

[TMDD96] K. Takei, O. Mundigl, L. Daniell, and P. De Camilli. The synaptic vesicle cycle: a single vesicle budding step involving clathrin and dynamin. Journal of Cell Biology, 133(6):1237–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
133/6/1237.

Turner:1995:OAP

[TMM^{+}95] J. E. Turner, C. G. Minkoff, K. H. Martin, R. Misra, and K. I. Swenson. Oocyte activation and passage through the metaphase/anaphase transition of the meiotic cell cycle is blocked in clams by inhibitors of HMG–CoA reductase activity. Journal of Cell Biology, 128(6):1145–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/6/1145.

Terasaki:1997:LPM

[TMM97] Mark Terasaki, Katsuya Miyake, and Paul L. McNeil. Large Plasma Membrane Disruptions Are Rapidly Resealed by Ca^{2+}- dependent Vesicle–Vesicle Fusion Events. Journal of Cell Biology, 139(1):63–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/1/63.

Toyoshima:1997:FIC

[TMN97] Fumiko Toyoshima, Tetsuo Moriguchi, and Eisuke Nishida. Fas induces cytoplasmic apoptotic responses and activation of the MKK7-JNK/SAPK and MKK6-p38 pathways independent of CPP32-like proteases. Journal of Cell Biology, 139(4): 1005–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/1005.

Tarsounas:1999:RDF

[TMPM99] Madalena Tarsounas, Takashi Morita, Ronald E. Pearlman, and Peter B. Moens. Rad51 and Dmc1 Form Mixed Complexes Associated with Mouse Meiotic Chromosome Cores and Synaptonemal Complexes. Journal of Cell Biology, 147(2): 207–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/2/207.

Taneja:1995:FTR

[TMS^{+}95] K. L. Taneja, M. McCurrach, M. Schalling, D. Housman, and R. H. Singer. Foci of trinucleotide repeat transcripts in nuclei of myotonic dystrophy cells and tissues. Journal of Cell Biology, 128(6):995–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/6/995.

TerBush:1995:SSS

[TN95] D. R. TerBush and P. Novick. Sec6, Sec8, and Sec15 are components of a multisubunit complex which localizes to small bud tips in Saccharomyces cerevisiae. Journal of Cell Biology, 130(2):299–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/2/299.

Trapp:1997:DDP

[TNCM97] Bruce D. Trapp, Akiko Nishiyama, David Cheng, and Wendy Macklin. Differentiation and Death of Premyelinating Oligodendrocytes in Developing Rodent Brain. Journal of Cell Biology, 137(2):459–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/2/459.

Tam:1998:DRS

[TNFK^{+}98] Amy Tam, Franklin J. Nouvet, Konomi Fujimura-Kamada, Hilda Slunt, Sangram S. Sisodia, and Susan Michaelis. Dual Roles for Ste24p in Yeast a–Factor Maturation: NH_{2}-terminal Proteolysis and COOH-terminal CAAX Processing. Journal of Cell Biology, 142(3):635–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/3/635.

Takahashi:1999:NPI

[TNM^{+}99] Kenichi Takahashi, Hiroyuki Nakanishi, Masako Miyahara, Kenji Mandai, Keiko Satoh, Ayako Satoh, Hideo Nishioka, Junken Aoki, Akio Nomoto, Akira Mizoguchi, and Yoshimi Takai. Nectin/PRR: an immunoglobulin-like cell adhesion molecule recruited to cadherin-based adherens junctions through interaction with afadin, a PDZ domain–containing protein. Journal of Cell Biology, 145(3):539–??, May 1999.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/145/3/
539.

Takeda:1995:VSK

[TNY^{+}95] H. Takeda, A. Nagafuchi, S. Yonemura, S. Tsukita, J. Behrens, W. Birchmeier, and S. Tsukita. V-src kinase shifts the cadherin-based cell adhesion from the strong to the weak state and beta catenin is not required for the shift. Journal of Cell Biology, 131(6):1839–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1839.

Tirnauer:1999:YBP

[TOB^{+}99] Jennifer S. Tirnauer, Eileen O’Toole, Lisbeth Berrueta, Barbara E. Bierer, and David Pellman. Yeast Bim1p Promotes the G1-specific Dynamics of Microtubules. Journal of Cell Biology, 145(5):993–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/5/993.

Torroja:1998:SEG

[TOSF^{+}98] Laura Torroja, Daniel Ortuño-Sahagún, Alberto Ferrús, Barbara Hämmerle, and Julio A. Barbas. scully, an essential gene of Drosophila, is homologous to mammalian mitochondrial type II l-3-hydroxyacyl-CoA dehydrogenase/amyloid-β peptide-binding protein. Journal of Cell Biology, 141(4): 1009–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/4/1009.

Takai:1996:MSP

[TOT^{+}96] Y. Takai, M. Ogawara, Y. Tomono, C. Moritoh, S. Imajoh-Ohmi, O. Tsutsumi, Y. Taketani, and M. Inagaki. Mitosisspecific phosphorylation of vimentin by protein kinase C coupled with reorganization of intracellular membranes. Journal of Cell Biology, 133(1):141–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/1/141.

Tripiciano:1999:CEE

[TPM^{+}99] Antonella Tripiciano, Carmelina Peluso, Anna Rita Morena, Fioretta Palombi, Mario Stefanini, Elio Ziparo, Masashi Yanagisawa, and Antonio Filippini. Cyclic Expression of Endothelin-converting Enzyme-1 Mediates the Functional Regulation of Seminiferous Tubule Contraction. Journal of Cell Biology, 145(5):1027–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/5/1027.

Tisdale:1997:PBC

[TPMB97] Ellen J. Tisdale, Helen Plutner, Jeanne Matteson, and William E. Balch. p53/58 Binds COPI and Is Required for Selective Transport through the Early Secretory Pathway. Journal of Cell Biology, 137(3):581–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/3/581.

Tuckermann:1999:DBI

[TRA^{+}99] Jan P. Tuckermann, Holger M. Reichardt, Rosa Arribas, K. Hartmut Richter, Günther Schütz, and Peter Angel. The DNA Binding–Independent Function of the Glucocorticoid Receptor Mediates Repression of Ap-1–Dependent Genes in Skin. Journal of Cell Biology, 147(7):1365–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1365.

Thery:1999:MCD

[TRG^{+}99] Clotilde Théry, Armelle Regnault, Jérôme Garin, Joseph Wolfers, Laurence Zitvogel, Paola Ricciardi-Castagnoli, Graça Raposo, and Sebastian Amigorena. Molecular Characterization of Dendritic Cell–Derived Exosomes. Journal of Cell Biology, 147(3):599–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/3/599.

Trupp:1995:PEB

[TRJ^{+}95] M. Trupp, M. Rydén, H. Jörnvall, H. Funakoshi, T. Timmusk, E. Arenas, and C. F. Ibáñez. Peripheral expression and biological activities of GDNF, a new neurotrophic factor for avian and mammalian peripheral neurons. Journal of Cell Biology, 130(1):137–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/1/137.

Taddei:1999:DMH

[TRS^{+}99] Angela Taddei, Danièle Roche, Jean-Baptiste Sibarita, Bryan M. Turner, and Geneviève Almouzni. Duplication and Maintenance of Heterochromatin Domains. Journal of Cell Biology, 147(6):1153–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/147/6/1153.

Tumbar:1999:LSC

[TSB99] Tudorita Tumbar, Gail Sudlow, and Andrew S. Belmont. Large-scale Chromatin Unfolding and Remodeling Induced by VP16 Acidic Activation Domain. Journal of Cell Biology, 145 (7):1341–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/7/1341.

Trusolino:1998:GFD

[TSC^{+}98] Livio Trusolino, Guido Serini, Germana Cecchini, Cristina Besati, Francesco Saverio Ambesi-Impiombato, Pier Carlo Marchisio, and Rosaria De Filippi. Growth factor-dependent activation of αvβ3 integrin in normal epithelial cells: Implications for tumor invasion. Journal of Cell Biology, 142(4): 1145–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/4/1145.

Tang:1997:MAG

[TSD^{+}97] Song Tang, Ying Jing Shen, Maria Elena DeBellard, Gitali Mukhopadhyay, James L. Salzer, Paul R. Crocker, and Marie T. Filbin. Myelin-associated Glycoprotein Interacts with Neurons via a Sialic Acid Binding Site at ARG118 and a Distinct Neurite Inhibition Site. Journal of Cell Biology, 138(6): 1355–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1355.

Tortorella:1998:DTM

[TSH^{+}98] Domenico Tortorella, Craig M. Story, Johannes B. Huppa, Emmanuel J. H. J. Wiertz, Thomas R. Jones, and Hidde L. Ploegh. Dislocation of Type I Membrane Proteins from the ER to the Cytosol Is Sensitive to Changes in Redox Potential. Journal of Cell Biology, 142(2):365–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/2/365.

Takaishi:1997:RCC

[TSK^{+}97] Kenji Takaishi, Takuya Sasaki, Hirokazu Kotani, Hideo Nishioka, and Yoshimi Takai. Regulation of Cell–Cell Adhesion by Rac and Rho Small G Proteins in MDCK Cells. Journal of Cell Biology, 139(4):1047–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/4/1047.

Tomishige:1998:RML

[TSK98] Michio Tomishige, Yasushi Sako, and Akihiro Kusumi. Regulation Mechanism of the Lateral Diffusion of Band 3 in Erythrocyte Membranes by the Membrane Skeleton. Journal of Cell Biology, 142(4):989–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/4/989.

Tan:1998:RRO

[TSL^{+}98] Shirlee Tan, Yutaka Sagara, Yuanbin Liu, Pamela Maher, and David Schubert. The Regulation of Reactive Oxygen Species Production during Programmed Cell Death. Journal of Cell Biology, 141(6):1423–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/6/1423.

Taraboulos:1995:CDM

[TSS^{+}95a] A. Taraboulos, M. Scott, A. Semenov, D. Avrahami, L. Laszlo, S. B. Prusiner, and D. Avraham. Cholesterol depletion and modification of COOH-terminal targeting sequence of the prion protein inhibit formation of the scrapie isoform. Journal of Cell Biology, 129(1):121–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/1/121.

Torok:1995:OHO

[TSS^{+}95b] I. Török, D. Strand, R. Schmitt, G. Tick, T. Török, I. Kiss, and B. M. Mechler. The overgrown hematopoietic organs-31 tumor suppressor gene of Drosophila encodes an Importin-like protein accumulating in the nucleus at the onset of mitosis. Journal of Cell Biology, 129(6):1473–??, June 1995. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/6/1473.

Titorenko:1998:PYY

[TSSR98] Vladimir I. Titorenko, Jennifer J. Smith, Rachel K. Szilard, and Richard A. Rachubinski. Pex20p of the Yeast Yarrowia lipolytica Is Required for the Oligomerization of Thiolase in the Cytosol and for Its Targeting to the Peroxisome. Journal of Cell Biology, 142(2):403–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/2/403.

Tanimura:1996:CRH

[TT96] A. Tanimura and R. J. Turner. Calcium release in HSY cells conforms to a steady-state mechanism involving regulation of the inositol 1,4,5-trisphosphate receptor Ca^{2+} channel by luminal [Ca^{2+}]. Journal of Cell Biology, 132(4):607–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
4/607.

Tsumaki:1999:RCS

[TTAH^{+}99] Noriyuki Tsumaki, Kazuhiro Tanaka, Eri Arikawa-Hirasawa, Takanobu Nakase, Tomoatsu Kimura, J. Terrig Thomas, Takahiro Ochi, Frank P. Luyten, and Yoshihiko Yamada. Role of CDMP-1 in Skeletal Morphogenesis: Promotion of Mesenchymal Cell Recruitment and Chondrocyte Differentiation. Journal of Cell Biology, 144(1):161–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/1/161.

Tilney:1995:FAB

[TTG95] L. G. Tilney, M. S. Tilney, and G. M. Guild. F actin bundles in Drosophila bristles. I. Two filament cross-links are involved in bundling. Journal of Cell Biology, 130(3):629–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
3/629.

Tilney:1996:FAF

[TTG96] L. G. Tilney, M. S. Tilney, and G. M. Guild. Formation of actin filament bundles in the ring canals of developing Drosophila follicles. Journal of Cell Biology, 133(1):61–??, April 1996.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/133/1/61.

Terada:1999:IIS

[TTT^{+}99] Sumio Terada, Tetsuhiro Tsujimoto, Yosuke Takei, Tomoyuki Takahashi, and Nobutaka Hirokawa. Impairment of Inhibitory Synaptic Transmission in Mice Lacking Synapsin I. Journal of Cell Biology, 145(5):1039–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/5/1039.

Terami:1999:GOE

[TWiK^{+}99] Hiromi Terami, Benjamin D. Williams, Shin ichi Kitamura, Yasuji Sakube, Shinji Matsumoto, Shima Doi, Takashi Obinata, and Hiroaki Kagawa. Genomic Organization, Expression, and Analysis of the Troponin C Gene pat-10 of Caenorhabditis elegans. Journal of Cell Biology, 146(1): 193–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/1/193.

Togel:1998:NFL

[TWP98] Martin Tögel, Gerhard Wiche, and Friedrich Propst. Novel Features of the Light Chain of Microtubule-associated Protein MAP1B: Microtubule Stabilization, Self Interaction, Actin Filament Binding, and Regulation by the Heavy Chain. Journal of Cell Biology, 143(3):695–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/3/695.

Tran:1997:MIS

[TWS97] P. T. Tran, R. A. Walker, and E. D. Salmon. A Metastable Intermediate State of Microtubule Dynamic Instability That Differs Significantly between Plus and Minus Ends. Journal of Cell Biology, 138(1):105–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/1/105.

Tang:1998:RGC

[TWSD98] Ming-Jer Tang, Dane Worley, Michele Sanicola, and Gregory R. Dressler. The RET–Glial Cell-derived Neurotrophic Factor (GDNF) Pathway Stimulates Migration and Chemoattraction of Epithelial Cells. Journal of Cell Biology, 142(5):

1337–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/5/1337.

Tan:1995:HPP

[TWVC95] X. Tan, H. R. Waterham, M. Veenhuis, and J. M. Cregg. The Hansenula polymorpha PER8 gene encodes a novel peroxisomal integral membrane protein involved in proliferation. Journal of Cell Biology, 128(3):307–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/3/307.

Tatsumoto:1999:HEE

[TXB^{+}99] Takashi Tatsumoto, Xiaozhen Xie, Rayah Blumenthal, Isamu Okamoto, and Toru Miki. Human Ect2 is an exchange factor for rho GTPases, phosphorylated in G2/M phases, and involved in cytokinesis. Journal of Cell Biology, 147(5):921– ??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/5/921.

Taya:1998:RTA

[TYK^{+}98] Shinichiro Taya, Takaharu Yamamoto, Kyoko Kano, Yoji Kawano, Akihiro Iwamatsu, Tomoko Tsuchiya, Keiji Tanaka, Masami Kanai-Azuma, Stephen A. Wood, John S. Mattick, and Kozo Kaibuchi. The Ras Target AF-6 is a Substrate of the Fam Deubiquitinating Enzyme. Journal of Cell Biology, 142 (4):1053–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/4/1053.

Torres:1996:AWC

[TYSP^{+}96] M. A. Torres, J. A. Yang-Snyder, S. M. Purcell, A. A. De-Marais, L. L. McGrew, and R. T. Moon. Activities of the Wnt-1 class of secreted signaling factors are antagonized by the Wnt-5A class and by a dominant negative cadherin in early Xenopus development. Journal of Cell Biology, 133 (5):1123–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/5/1123.

Tokumoto:1997:ICB

[TYT^{+}97] Toshinobu Tokumoto, Masakane Yamashita, Mika Tokumoto, Yoshinao Katsu, Ryo Horiguchi, Hiroko Kajiura, and Yoshitaka Nagahama. Initiation of Cyclin B Degradation by the 26S Proteasome upon Egg Activation. Journal of Cell Biology, 138(6):1313–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/6/1313.

Takeda:1999:LRA

[TYT^{+}99] Sen Takeda, Yoshiaki Yonekawa, Yosuke Tanaka, Yasushi Okada, Shigenori Nonaka, and Nobutaka Hirokawa. Left-right asymmetry and kinesin superfamily protein KIF3A: New insights in determination of laterality and mesoderm induction by kif3A^{−}/− mice analysis. Journal of Cell Biology, 145 (4):825–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/4/825.

Totsukawa:1999:AMP

[TYY^{+}99] Go Totsukawa, Yoshihiko Yamakita, Shigeko Yamashiro, Hiroshi Hosoya, David J. Hartshorne, and Fumio Matsumura. Activation of Myosin Phosphatase Targeting Subunit by Mitosis-specific Phosphorylation. Journal of Cell Biology, 144 (4):735–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/4/735.

Tuma:1998:HKI

[TZL^{+}98] M. Carolina Tuma, Andrew Zill, Nathalie Le Bot, Isabelle Vernos, and Vladimir Gelfand. Heterotrimeric Kinesin II Is the Microtubule Motor Protein Responsible for Pigment Dispersion in Xenopus Melanophores. Journal of Cell Biology, 143(6):1547–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/6/1547.

Uchida:1996:CCA

[UHJ^{+}96] N. Uchida, Y. Honjo, K. R. Johnson, M. J. Wheelock, and M. Takeichi. The catenin/cadherin adhesion system is localized in synaptic junctions bordering transmitter release zones. Journal of Cell Biology, 135(3):767–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
3/767.

Uberall:1999:EAP

[ÜHK^{+}99] Florian Überall, Karina Hellbert, Sonja Kampfer, Karl Maly, Andreas Villunger, Martin Spitaler, James Mwanjewe, Gabriele Baier-Bitterlich, Gottfried Baier, and Hans H. Grunicke. Evidence That Atypical Protein Kinase C-λ and Atypical Protein Kinase C-ζ Participate in Ras-mediated Reorganization of the F-actin Cytoskeleton. Journal of Cell Biology, 144(3):413–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/3/413.

Ueda:1999:RON

[ULMT99] H. Ueda, J. M. Levine, R. H. Miller, and B. D. Trapp. Rat Optic Nerve Oligodendrocytes Develop in the Absence of Viable Retinal Ganglion Cell Axons. Journal of Cell Biology, 146(6): 1365–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1365.

Ungermann:1998:VVS

[UNPW98] Christian Ungermann, Benjamin J. Nichols, Hugh R. B. Pelham, and William Wickner. A Vacuolar v–t–SNARE Complex, the Predominant Form In Vivo and on Isolated Vacuoles, Is Disassembled and Activated for Docking and Fusion. Journal of Cell Biology, 140(1):61–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/140/1/61.

Urata:1995:TDS

[UPAS95] Y. Urata, S. J. Parmelee, D. A. Agard, and J. W. Sedat. A three-dimensional structural dissection of Drosophila polytene chromosomes. Journal of Cell Biology, 131(2):279–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
2/279.

Urbe:1998:HFI

[UPT98] Sylvie Urbé, Lesley J. Page, and Sharon A. Tooze. Homotypic Fusion of Immature Secretory Granules during Maturation in a Cell-free Assay. Journal of Cell Biology, 143(7): 1831–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/7/1831.

Ullrich:1996:RRR

[URU^{+}96] O. Ullrich, S. Reinsch, S. Urbé, M. Zerial, and R. G. Parton. Rab11 regulates recycling through the pericentriolar recycling endosome. Journal of Cell Biology, 135(4):913–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/913.

Ungermann:1999:TVS

[UvMJ^{+}99] Christian Ungermann, Gabriele Fischer von Mollard, Ole N. Jensen, Nathan Margolis, Tom H. Stevens, and William Wickner. Three v–SNAREs and Two t–SNAREs, Present in a Pentameric cis–SNARE Complex on Isolated Vacuoles, Are Essential for Homotypic Fusion. Journal of Cell Biology, 145 (7):1435–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/7/1435.

vonArx:1995:DNE

[vABSP95] P. von Arx, S. Bantle, T. Soldati, and J. C. Perriard. Dominant negative effect of cytoplasmic actin isoproteins on cardiomyocyte cytoarchitecture and function. Journal of Cell Biology, 131(6):1759–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/131/6/1759.

Vale:1996:SLA

[Val96] R. D. Vale. Switches, latches, and amplifiers: common themes of G proteins and molecular motors. Journal of Cell Biology, 135(2):291–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/2/291.

Varnai:1998:VPB

[VB98] Péter Várnai and Tamás Balla. Visualization of phosphoinositides that bind pleckstrin homology domains: Calcium- and agonist-induced dynamic changes and relationship to Myo-[^{3} H]inositol-labeled phosphoinositide pools. Journal of Cell Biology, 143(2):501–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/2/501.

VanHaastert:1996:CDF

[VBG96] P. J. Van Haastert, J. D. Bishop, and R. H. Gomer. The cell density factor CMF regulates the chemoattractant receptor cAR1 in Dictyostelium. Journal of Cell Biology, 134(6): 1543–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1543.

Vogel:1996:PDA

[VBZ96] S. S. Vogel, P. S. Blank, and J. Zimmerberg. Poissondistributed active fusion complexes underlie the control of the rate and extent of exocytosis by calcium. Journal of Cell Biology, 134(2):329–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/2/329.

Vecchi:1997:CPE

[VC97] Manuela Vecchi and Graham Carpenter. Constitutive Proteolysis of the ErbB-4 Receptor Tyrosine Kinase by a Unique, Sequential Mechanism. Journal of Cell Biology, 139(4):995– ??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/995.

Varkey:1999:ACC

[VCJA99] Johnson Varkey, Po Chen, Ronald Jemmerson, and John M. Abrams. Altered Cytochrome c Display Precedes Apoptotic Cell Death in Drosophila. Journal of Cell Biology, 144(4): 701–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/4/701.

VanRheenen:1998:SNP

[VCL^{+}98] Susan M. VanRheenen, Xiaochun Cao, Vladimir V. Lupashin, Charles Barlowe, and M. Gerard Waters. Sec35p, a Novel Peripheral Membrane Protein, Is Required for ER to Golgi Vesicle Docking. Journal of Cell Biology, 141(5):1107–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/141/
5/1107.

VanRheenen:1999:SPR

[VCS^{+}99] Susan M. VanRheenen, Xiaochun Cao, Stephanie K. Sapperstein, Elbert C. Chiang, Vladimir V. Lupashin, Charles Barlowe, and M. Gerard Waters. Sec34p, a Protein Required for Vesicle Tethering to the Yeast Golgi Apparatus, Is in a Complex with Sec35p. Journal of Cell Biology, 147(4):729– ??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/4/729.

Vainio:1996:HAM

[VDA^{+}96] O. Vainio, D. Dunon, F. Aïssi, J. P. Dangy, K. M. McNagny, and B. A. Imhof. HEMCAM, an adhesion molecule expressed by c-kit+ hemopoietic progenitors. Journal of Cell Biology, 135(6):1655–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1655.

vandenBerg:1995:EGP

[vdBCH^{+}95] C. W. van den Berg, T. Cinek, M. B. Hallett, V. Horejsi, and B. P. Morgan. Exogenous glycosyl phosphatidylinositolanchored CD59 associates with kinases in membrane clusters on U937 cells and becomes Ca(2+)-signaling competent. Journal of Cell Biology, 131(3):669–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/131/3/669.

vanderBijl:1996:SNS

[vdBLCvM96] P. van der Bijl, M. Lopes-Cardozo, and G. van Meer. Sorting of newly synthesized galactosphingolipids to the two surface domains of epithelial cells. Journal of Cell Biology, 132 (5):813–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/5/813.

vanderLoop:1996:SNC

[vdLST^{+}96] F. T. van der Loop, G. Schaart, E. D. Timmer, F. C. Ramaekers, and G. J. van Eys. Smoothelin, a novel cytoskeletal protein specific for smooth muscle cells. Journal of Cell Biology, 134(2):401–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/2/401.

vanDelft:1997:ACE

[vDSH^{+}97] Sanne van Delft, Christopher Schumacher, Willem Hage, Arie J. Verkleij, and Paul M. P. van Bergen en Henegouwen. Association and Colocalization of Eps15 with Adaptor Protein-2 and Clathrin. Journal of Cell Biology, 136(4): 811–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/4/811.

Vida:1995:NVS

[VE95] T. A. Vida and S. D. Emr. A new vital stain for visualizing vacuolar membrane dynamics and endocytosis in yeast. Journal of Cell Biology, 128(5):779–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/128/5/779.

Vida:1999:CFA

[VG99] Thomas Vida and Brenda Gerhardt. A cell-free assay allows reconstitution of Vps33p-dependent transport to the yeast vacuole/lysosome. Journal of Cell Biology, 146(1):85–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
1/85.

Vaisberg:1996:MCE

[VGM96] E. A. Vaisberg, P. M. Grissom, and J. R. McIntosh. Mammalian cells express three distinct dynein heavy chains that are localized to different cytoplasmic organelles. Journal of Cell Biology, 133(4):831–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/4/831.

Vartanian:1997:ANS

[VGVF97] Timothy Vartanian, Andrew Goodearl, Andrea Viehöver, and Gerald Fischbach. Axonal Neuregulin Signals Cells of the Oligodendrocyte Lineage through Activation of HER4 and Schwann Cells through HER2 and HER3. Journal of Cell Biology, 137(1):211–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/1/211.

vanGenderen:1995:DTG

[vGvM95] I. van Genderen and G. van Meer. Differential targeting of glucosylceramide and galactosylceramide analogues after synthesis but not during transcytosis in Madin–Darby canine kidney cells. Journal of Cell Biology, 131(3):645–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/645.

Valtz:1996:PPY

[VH96] N. Valtz and I. Herskowitz. Pea2 protein of yeast is localized to sites of polarized growth and is required for efficient mating and bipolar budding. Journal of Cell Biology, 135(3): 725–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/3/725.

Vandevoorde:1997:IET

[VHF97] Veronique Vandevoorde, Guy Haegeman, and Walter Fiers. Induced Expression of Trimerized Intracellular Domains of the Human Tumor Necrosis Factor (TNF) p55 Receptor Elicits TNF Effects. Journal of Cell Biology, 137(7):1627–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
7/1627.

vanHengel:1997:PKC

[vHGB^{+}97] Jolanda van Hengel, Lionel Gohon, Erik Bruyneel, Stefan Vermeulen, Maria Cornelissen, Marc Mareel, and Frans van Roy. Protein Kinase C Activation Upregulates Intercellular Adhesion of α-Catenin–negative Human Colon Cancer Cell Variants via Induction of Desmosomes. Journal of Cell Biology, 137 (5):1103–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/5/1103.

vantHof:1997:EAB

[vHMV^{+}97] Wouter van’t Hof, Abha Malik, S. Vijayakumar, Jizeng Qiao, Janet van Adelsberg, and Qais Al-Awqati. The Effect of Apical and Basolateral Lipids on the Function of the Band 3 Anion Exchange Protein. Journal of Cell Biology, 139(4): 941–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/941.

Visa:1996:NCB

[VIF^{+}96] N. Visa, E. Izaurralde, J. Ferreira, B. Daneholt, and I. W. Mattaj. A nuclear cap-binding complex binds Balbiani ring pre-mRNA cotranscriptionally and accompanies the ribonucleoprotein particle during nuclear export. Journal of Cell Biology, 133(1):5–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/1/5.

vanIJzendoorn:1998:GSS

[vIH98] Sven C. D. van IJzendoorn and Dick Hoekstra. (glyco)sphingolipids are sorted in sub–apical compartments in HepG2 cells: a role for non–Golgi-related intracellular sites in the polarized distribution of (glyco)sphingolipids. Journal of Cell Biology, 142 (3):683–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/3/683.

vanIJzendoorn:1997:SGS

[vIZKH97] Sven C. D. van IJzendoorn, Mirjam M. P. Zegers, Jan Willem Kok, and Dick Hoekstra. Segregation of glucosylceramide and sphingomyelin occurs in the apical to basolateral transcytotic route in HepG2 cells. Journal of Cell Biology, 137(2):347–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/2/347.

Voigt:1996:SSD

[VJHR96] S. Voigt, B. Jungnickel, E. Hartmann, and T. A. Rapoport. Signal sequence-dependent function of the TRAM protein during early phases of protein transport across the endoplasmic reticulum membrane. Journal of Cell Biology, 134(1):25–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
134/1/25.

Vollenweider:1998:MLE

[VKIH98] Florence Vollenweider, Felix Kappeler, Christian Itin, and Hans-Peter Hauri. Mistargeting of the Lectin ERGIC-53 to the Endoplasmic Reticulum of HeLa Cells Impairs the Secretion of a Lysosomal Enzyme. Journal of Cell Biology, 142 (2):377–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/377.

Vastrik:1995:EMG

[VKP^{+}95] I. Västrik, A. Kaipainen, T. L. Penttilä, A. Lymboussakis, R. Alitalo, M. Parvinen, and K. Alitalo. Expression of the mad gene during cell differentiation in vivo and its inhibition of cell growth in vitro. Journal of Cell Biology, 128(6):1197–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/6/1197.

Vilquin:1996:PLA

[VKR^{+}96] J. T. Vilquin, I. Kinoshita, B. Roy, M. Goulet, E. Engvall, F. Tomé, M. Fardeau, and J. P. Tremblay. Partial laminin alpha2 chain restoration in alpha2 chain-deficient dy/dy mouse by primary muscle cell culture transplantation. Journal of Cell Biology, 133(1):185–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/1/185.

Verschoor:1999:FGS

[VL99] Adriana Verschoor and Trevor Lithgow. A First Glimpse at the Structure of the Tom Translocase from the Mitochondrial Outer Membrane. Journal of Cell Biology, 147(5):905– ??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/5/905.

Verhey:1998:LCD

[VLA^{+}98] Kristen J. Verhey, Donna L. Lizotte, Tatiana Abramson, Linda Barenboim, Bruce J. Schnapp, and Tom A. Rapoport. Light chain-dependent regulation of kinesin’s interaction with microtubules. Journal of Cell Biology, 143(4):1053–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
4/1053.

Verheggen:1998:PPR

[VLAHV98] Céline Verheggen, Sophie Le Panse, Geneviève Almouzni, and Danièle Hernandez-Verdun. Presence of Pre-rRNAs before Activation of Polymerase I Transcription in the Building Process of Nucleoli during Early Development of Xenopus laevis. Journal of Cell Biology, 142(5):1167–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/5/1167.

vanLeeuwen:1997:GNE

[vLKvdK^{+}97] Frank N. van Leeuwen, Hendrie E. T. Kain, Rob A. van der Kammen, Frits Michiels, Onno W. Kranenburg, and John G. Collard. The Guanine Nucleotide Exchange Factor Tiam1 Affects Neuronal Morphology; Opposing Roles for the Small GTPases Rac and Rho. Journal of Cell Biology, 139(3):797– ??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/3/797.

VanBrocklyn:1998:DAS

[VLM^{+}98] James R. Van Brocklyn, Menq-Jer Lee, Ramil Menzeleev, Ana Olivera, Lisa Edsall, Olivier Cuvillier, Dianne M. Thomas, Peter J. P. Coopman, Shobha Thangada, Catherine H. Liu, Timothy Hla, and Sarah Spiegel. Dual Actions of Sphingosine-1-Phosphate: Extracellular through the G_{i}-coupled Receptor Edg-1 and Intracellular to Regulate Proliferation and Survival. Journal of Cell Biology, 142(1):229–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/1/229.

Vachon:1996:MLM

[VLX^{+}96] P. H. Vachon, F. Loechel, H. Xu, U. M. Wewer, and E. Engvall. Merosin and laminin in myogenesis; specific requirement for merosin in myotube stability and survival. Journal of Cell Biology, 134(6):1483–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/6/1483.

Volkmer:1996:NIN

[VLZR96] H. Volkmer, R. Leuschner, U. Zacharias, and F. G. Rathjen. Neurofascin induces neurites by heterophilic interactions with axonal NrCAM while NrCAM requires F11 on the axonal surface to extend neurites. Journal of Cell Biology, 135(4): 1059–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/4/1059.

Vaduva:1997:ABV

[VMH97] Gabriela Vaduva, Nancy C. Martin, and Anita K. Hopper. Actin-binding Verprolin Is a Polarity Development Protein Required for the Morphogenesis and Function of the Yeast Actin Cytoskeleton. Journal of Cell Biology, 139(7):1821–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
139/7/1821.

Verde:1995:FYC

[VMN95] F. Verde, J. Mata, and P. Nurse. Fission yeast cell morphogenesis: identification of new genes and analysis of their role during the cell cycle. Journal of Cell Biology, 131(6): 1529–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1529.

vonMollard:1997:YVS

[vMNS97] Gabriele Fischer von Mollard, Steven F. Nothwehr, and Tom H. Stevens. The Yeast v–SNARE Vti1p Mediates Two Vesicle Transport Pathways through Interactions with the t– SNAREs Sed5p and Pep12p. Journal of Cell Biology, 137 (7):1511–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/7/1511.

Vaillant:1999:DNC

[VMT^{+}99] A. R. Vaillant, I. Mazzoni, C. Tudan, M. Boudreau, D. R. Kaplan, and F. D. Miller. Depolarization and Neurotrophins Converge on the Phosphatidylinositol 3–Kinase–Akt Pathway to Synergistically Regulate Neuronal Survival. Journal of Cell Biology, 146(5):955–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/146/5/955.

Venuti:1995:MRL

[VMV^{+}95] J. M. Venuti, J. H. Morris, J. L. Vivian, E. N. Olson, and W. H. Klein. Myogenin is required for late but not early aspects of myogenesis during mouse development. Journal of Cell Biology, 128(4):563–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/4/563.

Vivinus-Nebot:1999:LHT

[VNTM^{+}99] Mylène Vivinus-Nebot, Michel Ticchioni, Florence Mary, Paul Hofman, Vito Quaranta, Patricia Rousselle, and Alain Bernard. Laminin 5 in the human thymus: Control of T cell proliferation via α_{6}β_{4} integrins. Journal of Cell Biology, 144 (3):563–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/3/563.

Volz:1995:SRR

[VOP^{+}95] B. Volz, G. Orberger, S. Porwoll, H. P. Hauri, and R. Tauber. Selective reentry of recycling cell surface glycoproteins to the biosynthetic pathway in human hepatocarcinoma HepG2 cells. Journal of Cell Biology, 130(3):537–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/3/537.

Valtz:1995:FRO

[VPH95] N. Valtz, M. Peter, and I. Herskowitz. FAR1 is required for oriented polarization of yeast cells in response to mating pheromones. Journal of Cell Biology, 131(4):863–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/4/863.

Vashishtha:1998:SPM

[VPM^{+}98] Malini Vashishtha, Thomas Phalen, Marianne T. Marquardt, Jae S. Ryu, Alice C. Ng, and Margaret Kielian. A Single Point Mutation Controls the Cholesterol Dependence of Semliki Forest Virus Entry and Exit. Journal of Cell Biology, 140 (1):91–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/91.

Vincent:1996:MMA

[VRD96] I. Vincent, M. Rosado, and P. Davies. Mitotic mechanisms in Alzheimer’s disease? Journal of Cell Biology, 132(3): 413–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/413.

Voos:1998:RRL

[VS98] Wolfgang Voos and Tom H. Stevens. Retrieval of resident late-Golgi membrane proteins from the prevacuolar compartment of Saccharomyces cerevisiae is dependent on the function of Grd19p. Journal of Cell Biology, 140(3):577–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
3/577.

Verkhovsky:1995:MIF

[VSB95] A. B. Verkhovsky, T. M. Svitkina, and G. G. Borisy. Myosin II filament assemblies in the active lamella of fibroblasts: their morphogenesis and role in the formation of actin filament bundles. Journal of Cell Biology, 131(4):989–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
4/989.

Verschoor:1996:NSE

[VSGF96] A. Verschoor, S. Srivastava, R. Grassucci, and J. Frank. Native 3D structure of eukaryotic 80s ribosome: morphological homology with E. coli 70S ribosome. Journal of Cell Biology, 133(3):495–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/3/495.

Vogel:1997:CIS

[VSRP97] Jacalyn M. Vogel, Tim Stearns, Conly L. Rieder, and Robert E. Palazzo. Centrosomes Isolated from Spisula solidissima Oocytes Contain Rings and an Unusual Stoichiometric Ratio of α /β Tubulin. Journal of Cell Biology, 137(1):193–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
137/1/193.

Vagner:1996:TCA

[VTG^{+}96] S. Vagner, C. Touriol, B. Galy, S. Audigier, M. C. Gensac, F. Amalric, F. Bayard, H. Prats, and A. C. Prats. Translation of CUG- but not AUG-initiated forms of human fibroblast growth factor 2 is activated in transformed and stressed cells. Journal of Cell Biology, 135(5):1391–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
5/1391.

Vijayakumar:1999:HRA

[VTHAA99] S. Vijayakumar, Jiro Takito, Chinami Hikita, and Qais Al-Awqati. Hensin Remodels the Apical Cytoskeleton and Induces Columnarization of Intercalated Epithelial Cells: Processes that Resemble Terminal Differentiation. Journal of Cell Biology, 144(5):1057–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/5/1057.

Hof:1997:RPM

[vtHR97] Wouter van ’t Hof and Marilyn D. Resh. Rapid plasma membrane anchoring of newly synthesized p59^{fyn}: Selective requirement for NH_{2}-terminal myristoylation and palmitoylation at Cysteine-3. Journal of Cell Biology, 136(5):1023–??, March 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
5/1023.

Vozzi:1995:ACM

[VUC^{+}95] C. Vozzi, S. Ullrich, A. Charollais, J. Philippe, L. Orci, and P. Meda. Adequate connexin-mediated coupling is required for proper insulin production. Journal of Cell Biology, 131(6): 1561–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1561.

Vaughan:1995:CDB

[VV95] K. T. Vaughan and R. B. Vallee. Cytoplasmic dynein binds dynactin through a direct interaction between the intermediate chains and p150Glued. Journal of Cell Biology, 131(6): 1507–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1507.

Verschure:1999:SRB

[VvdKMvD99] Pernette J. Verschure, Ineke van der Kraan, Erik M. M. Manders, and Roel van Driel. Spatial Relationship between Transcription Sites and Chromosome Territories. Journal of Cell Biology, 147(1):13–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/1/13.

Vickery:1999:DDD

[VvZ99] Ross G. Vickery and Mark von Zastrow. Distinct Dynamindependent and -independent Mechanisms Target Structurally Homologous Dopamine Receptors to Different Endocytic Membranes. Journal of Cell Biology, 144(1):31–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
1/31.

vanWeert:1995:TLE

[vWDG^{+}95] A. W. van Weert, K. W. Dunn, H. J. Gueze, F. R. Maxfield, and W. Stoorvogel. Transport from late endosomes to lysosomes, but not sorting of integral membrane proteins in endosomes, depends on the vacuolar proton pump. Journal of Cell Biology, 130(4):821–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/4/821.

Vleminckx:1997:APC

[VWG^{+}97] Kris Vleminckx, Ellen Wong, Kathy Guger, Bonnee Rubinfeld, Paul Polakis, and Barry M. Gumbiner. Adenomatous Polyposis Coli Tumor Suppressor Protein Has Signaling Activity in Xenopus laevis Embryos Resulting in the Induction of an Ectopic Dorsoanterior Axis. Journal of Cell Biology, 136 (2):411–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/2/411.

Vilquin:1995:SHM

[VWK^{+}95] J. T. Vilquin, E. Wagner, I. Kinoshita, R. Roy, and J. P. Tremblay. Successful histocompatible myoblast transplantation in dystrophin-deficient mdx mouse despite the production of antibodies against dystrophin. Journal of Cell Biology, 131(4): 975–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/4/975.

Vidwans:1999:MRG

[VWO99] Smruti J. Vidwans, Mei Lie Wong, and Patrick H. O’Farrell. Mitotic Regulators Govern Progress through Steps in the Centrosome Duplication Cycle. Journal of Cell Biology, 147(7): 1371–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/7/1371.

Vijayasaradhi:1995:IST

[VXBH95] S. Vijayasaradhi, Y. Xu, B. Bouchard, and A. N. Houghton. Intracellular sorting and targeting of melanosomal membrane proteins: identification of signals for sorting of the human brown locus protein, gp75. Journal of Cell Biology, 130(4): 807–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/4/807.

Verhey:1995:DSG

[VYB95] K. J. Verhey, J. I. Yeh, and M. J. Birnbaum. Distinct signals in the GLUT4 glucose transporter for internalization and for targeting to an insulin-responsive compartment. Journal of Cell Biology, 130(5):1071–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/5/1071.

Volkmer:1998:DCM

[VZNR98] Hansjürgen Volkmer, Ute Zacharias, Ursel Nörenberg, and Fritz G. Rathjen. Dissection of Complex Molecular Interactions of Neurofascin with Axonin-1, F11, and Tenascin–R, Which Promote Attachment and Neurite Formation of Tectal Cells. Journal of Cell Biology, 142(4):1083–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
4/1083.

Wang:1995:CDB

[WA95] S. Z. Wang and R. Adler. Chromokinesin: a DNA-binding, kinesin-like nuclear protein. Journal of Cell Biology, 128(5): 761–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/5/761.

Wallace:1995:RIN

[Wal95] B. G. Wallace. Regulation of the interaction of nicotinic acetylcholine receptors with the cytoskeleton by agrin-activated protein tyrosine kinase. Journal of Cell Biology, 128(6):1121–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
128/6/1121.

Weber:1996:SRA

[WAMS96] C. Weber, R. Alon, B. Moser, and T. A. Springer. Sequential regulation of alpha 4 beta 1 and alpha 5 beta 1 integrin avidity by CC chemokines in monocytes: implications for transendothelial chemotaxis. Journal of Cell Biology, 134 (4):1063–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/1063.

White:1998:SAA

[WAV98] Phoebe White, Hermann Aberle, and Jean-Paul Vincent. Signaling and Adhesion Activities of Mammalian β-Catenin and Plakoglobin in Drosophila. Journal of Cell Biology, 140(1): 183–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/183.

Wilde:1996:VPA

[WB96] A. Wilde and F. M. Brodsky. In vivo phosphorylation of adaptors regulates their interaction with clathrin. Journal of Cell Biology, 135(3):635–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/3/635.

Williams:1999:ECR

[WB99] McRae W. Williams and Robert J. Bloch. Extensive but coordinated reorganization of the membrane skeleton in myofibers of dystrophic (mdx) mice. Journal of Cell Biology, 144(6): 1259–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/6/1259.

Wittmann:1998:LKL

[WBA^{+}98] Torsten Wittmann, Haralabia Boleti, Claude Antony, Eric Karsenti, and Isabelle Vernos. Localization of the Kinesin-like Protein Xklp2 to Spindle Poles Requires a Leucine Zipper, a Microtubule-associated Protein, and Dynein. Journal of Cell Biology, 143(3):673–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/3/673.

Whitehouse:1997:TSE

[WBG^{+}97] Caroline Whitehouse, Joy Burchell, Stephen Gschmeissner, Inka Brockhausen, Kenneth O. Lloyd, and Joyce Taylor-Papadimitriou. A Transfected Sialyltransferase That Is Elevated in Breast Cancer and Localizes to the medial/trans -Golgi Apparatus Inhibits the Development of core-2–based O–Glycans. Journal of Cell Biology, 137(6):1229–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
6/1229.

Wolins:1997:ADP

[WBKB97] Nathan Wolins, Herbert Bosshart, Helmut Küster, and Juan S. Bonifacino. Aggregation as a determinant of protein fate in post-Golgi compartments: Role of the luminal domain of furin in lysosomal targeting. Journal of Cell Biology, 139(7): 1735–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/7/1735.

West:1997:RAR

[WBR97] Michele A. West, Nicholas A. Bright, and Margaret S. Robinson. The Role of ADP-ribosylation Factor and Phospholipase D in Adaptor Recruitment. Journal of Cell Biology, 138(6): 1239–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1239.

Wu:1998:VMD

[WBR^{+}98] Xufeng Wu, Blair Bowers, Kang Rao, Qin Wei, and John A. Hammer. Visualization of Melanosome Dynamics within Wild–Type and Dilute Melanocytes Suggests a Paradigm for Myosin V Function In Vivo. Journal of Cell Biology, 143(7): 1899–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/7/1899.

Wei:1998:GTR

[WBS^{+}98] Maria L. Wei, Frank Bonzelius, Rebecca M. Scully, Regis B. Kelly, and Gary A. Herman. GLUT4 and Transferrin Receptor Are Differentially Sorted Along the Endocytic Pathway in CHO Cells. Journal of Cell Biology, 140(3):565–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
3/565.

Wilsch-Brauninger:1997:SLS

[WBSNV97] Michaela Wilsch-Bräuninger, Heinz Schwarz, and Christiane Nüsslein-Volhard. A sponge-like structure involved in the association and transport of maternal products during Drosophila oogenesis. Journal of Cell Biology, 139(3):817–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/817.

Wilson:1997:TIA

[WC97a] Jean M. Wilson and Tamara L. Colton. Targeting of an Intestinal Apical Endosomal Protein to Endosomes in Nonpolarized Cells. Journal of Cell Biology, 136(2):319–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
2/319.

Wu:1997:CCB

[WC97b] Rong-Rong Wu and John R. Couchman. cDNA cloning of the basement membrane chondroitin sulfate proteoglycan core protein, bamacan: a five domain structure including coiled– coil motifs. Journal of Cell Biology, 136(2):433–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/136/
2/433.

Waters:1998:LMK

[WCMS98] Jennifer C. Waters, Rey-Huei Chen, Andrew W. Murray, and E. D. Salmon. Localization of Mad2 to Kinetochores Depends on Microtubule Attachment, Not Tension. Journal of Cell Biology, 141(5):1181–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/5/1181.

Webb:1997:GPC

[WCV^{+}97] Helena Webb, Nicola Carnall, Luc Vanhamme, Sylvie Rolin, Jakke Van Den Abbeele, Sue Welburn, Etienne Pays, and Mark Carrington. The GPI–Phospholipase C of Trypanosoma brucei Is Nonessential But Influences Parasitemia in Mice. Journal of Cell Biology, 139(1):103–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/1/103.

Wang:1998:VVP

[WCW98] Yong-Xu Wang, Natalie L. Catlett, and Lois S. Weisman. Vac8p, a Vacuolar Protein with Armadillo Repeats, Functions in both Vacuole Inheritance and Protein Targeting from the Cytoplasm to Vacuole. Journal of Cell Biology, 140(5): 1063–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/5/1063.

Wyatt:1995:RNG

[WD95] S. Wyatt and A. M. Davies. Regulation of nerve growth factor receptor gene expression in sympathetic neurons during development. Journal of Cell Biology, 130(6):1435–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1435.

Wessely:1998:MGM

[WDL^{+}98] Oliver Wessely, Eva-Maria Deiner, Kim Chew Lim, Georg Mellitzer, Peter Steinlein, and Hartmut Beug. Mammalian Granulocyte–Macrophage Colony-stimulating Factor Receptor Expressed in Primary Avian Hematopoietic Progenitors: Lineage-specific Regulation of Proliferation and Differentiation. Journal of Cell Biology, 141(4):1041–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/4/
1041.

Welch:1997:HAC

[WDV^{+}97] Matthew D. Welch, Angela H. DePace, Suzie Verma, Akihiro Iwamatsu, and Timothy J. Mitchison. The human Arp2/ 3 complex is composed of evolutionarily conserved subunits and is localized to cellular regions of dynamic actin filament assembly. Journal of Cell Biology, 138(2):375–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
2/375.

Wendland:1998:PYE

[WE98] Beverly Wendland and Scott D. Emr. Pan1p, Yeast eps15, Functions as a Multivalent Adaptor That Coordinates Protein–Protein Interactions Essential for Endocytosis. Journal of Cell Biology, 141(1):71–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/1/71.

Wong:1996:DNE

[WF96] M. H. Wong and M. T. Filbin. Dominant-negative effect on adhesion by myelin Po protein truncated in its cytoplasmic domain. Journal of Cell Biology, 134(6):1531–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/134/
6/1531.

Wein:1996:DNK

[WFBC96] H. Wein, M. Foss, B. Brady, and W. Z. Cande. DSK1, a novel kinesin-related protein from the diatom Cylindrotheca fusiformis that is involved in anaphase spindle elongation. Journal of Cell Biology, 133(3):595–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/133/3/595.

Wubbolts:1996:DVT

[WFBO^{+}96] R. Wubbolts, M. Fernandez-Borja, L. Oomen, D. Verwoerd, H. Janssen, J. Calafat, A. Tulp, S. Dusseljee, and J. Neefjes. Direct vesicular transport of MHC class II molecules from lysosomal structures to the cell surface. Journal of Cell Biology, 135(3):611–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/3/611.

Wigley:1999:DAP

[WFL^{+}99] W. Christian Wigley, Rosalind P. Fabunmi, Min Goo Lee, Christopher R. Marino, Shmuel Muallem, George N. De-Martino, and Philip J. Thomas. Dynamic Association of Proteasomal Machinery with the Centrosome. Journal of Cell Biology, 145(3):481–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/3/481.

Wilson:1997:CMT

[WFS97] Nedra F. Wilson, Mary J. Foglesong, and William J. Snell. The Chlamydomonas Mating Type Plus Fertilization Tubule, a Prototypic Cell Fusion Organelle: Isolation, Characterization, and In Vitro Adhesion to Mating Type Minus Gametes. Journal of Cell Biology, 137(7):1537–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/137/7/1537.

Wong:1997:SPC

[WG97] Vivian Wong and Barry M. Gumbiner. A Synthetic Peptide Corresponding to the Extracellular Domain of Occludin Perturbs the Tight Junction Permeability Barrier. Journal of Cell Biology, 136(2):399–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/2/399.

Wen:1999:PPQ

[WG99] Chenhui Wen and Iva Greenwald. p24 Proteins and Quality Control of LIN-12 and GLP-1 Trafficking in Caenorhabditis elegans. Journal of Cell Biology, 145(6):1165–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
6/1165.

Walensky:1998:KFB

[WGF^{+}98] Loren D. Walensky, Philippe Gascard, Michael E. Field, Seth Blackshaw, John G. Conboy, Narla Mohandas, and Solomon H. Snyder. The 13-kD FK506 Binding Protein, FKBP13, Interacts with a Novel Homologue of the Erythrocyte Membrane Cytoskeletal Protein 4.1. Journal of Cell Biology, 141(1):143–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/1/143.

Williams:1996:BSA

[WGG96] B. C. Williams, M. Gatti, and M. L. Goldberg. Bipolar spindle attachments affect redistributions of ZW10, a Drosophila centromere/kinetochore component required for accurate chromosome segregation. Journal of Cell Biology, 134(5):1127– ??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/5/1127.

White:1998:TAC

[WGP98] Thomas W. White, Daniel A. Goodenough, and David L. Paul. Targeted Ablation of Connexin50 in Mice Results in Microphthalmia and Zonular Pulverulent Cataracts. Journal of Cell Biology, 143(3):815–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/3/815.

Wahlberg:1995:TGR

[WGR^{+}95] J. M. Wahlberg, I. Geffen, F. Reymond, T. Simmen, and M. Spiess. trans-Golgi retention of a plasma membrane protein: mutations in the cytoplasmic domain of the asialoglycoprotein receptor subunit H1 result in trans-Golgi retention. Journal of Cell Biology, 130(2):285–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/2/285.

Wang:1997:SMB

[WH97] Peijing Jeremy Wang and Tim C. Huffaker. Stu2p: a microtubule-binding protein that is an essential component of the yeast spindle pole body. Journal of Cell Biology, 139(5): 1271–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1271.

Woods:1999:CIM

[WHB^{+}99] Linda M. Woods, Craig A. Hodges, Esther Baart, Sean M. Baker, Michael Liskay, and Patricia A. Hunt. Chromosomal Influence on Meiotic Spindle Assembly: Abnormal Meiosis I in Female Mlh1 Mutant Mice. Journal of Cell Biology, 145 (7):1395–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/7/1395.

Wheatley:1997:CIR

[WHG^{+}97] Sally P. Wheatley, Edward H. Hinchcliffe, Michael Glotzer, Anthony A. Hyman, Greenfield Sluder, and Yu li Wang. CDK1 Inactivation Regulates Anaphase Spindle Dynamics and Cytokinesis In Vivo. Journal of Cell Biology, 138(2):385–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
2/385.

Wang:1996:MTE

[WHH96] Z. Z. Wang, S. F. Hardy, and Z. W. Hall. Membrane tethering enables an extracellular domain of the acetylcholine receptor alpha subunit to form a heterodimeric ligand-binding site. Journal of Cell Biology, 135(3):809–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/135/3/
809.

Wang:1996:IUC

[WHM^{+}96] W. Wang, P. A. Hansen, B. A. Marshall, J. O. Holloszy, and M. Mueckler. Insulin unmasks a COOH-terminal Glut4 epitope and increases glucose transport across T-tubules in skeletal muscle. Journal of Cell Biology, 135(2):415–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
2/415.

Woods:1996:DPR

[WHP^{+}96] D. F. Woods, C. Hough, D. Peel, G. Callaini, and P. J. Bryant. Dlg protein is required for junction structure, cell polarity, and proliferation control in Drosophila epithelia. Journal of Cell Biology, 134(6):1469–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/134/6/1469.

Wolter:1997:MBC

[WHS^{+}97] Keith G. Wolter, Yi-Te Hsu, Carolyn L. Smith, Amotz Nechushtan, Xu-Guang Xi, and Richard J. Youle. Movement of Bax from the Cytosol to Mitochondria during Apoptosis. Journal of Cell Biology, 139(5):1281–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/5/1281.

Weil:1996:CEM

[WJC^{+}96] M. Weil, M. D. Jacobson, H. S. Coles, T. J. Davies, R. L. Gardner, K. D. Raff, and M. C. Raff. Constitutive expression of the machinery for programmed cell death. Journal of Cell Biology, 133(5):1053–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/5/1053.

Wigge:1998:ASS

[WJH^{+}98] Philip A. Wigge, Ole N. Jensen, Simon Holmes, Sylvie Souès, Matthias Mann, and John V. Kilmartin. Analysis of the Saccharomyces spindle pole by matrix-assisted laser desorption/ ionization (MALDI) mass spectrometry. Journal of Cell Biology, 141(4):967–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/4/967.

White:1999:RCN

[WJM^{+}99] Jamie White, Ludger Johannes, Frédéric Mallard, Andreas Girod, Stephan Grill, Sigrid Reinsch, Patrick Keller, Barbara Tzschaschel, Arnaud Echard, Bruno Goud, and Ernst H. K. Stelzer. Rab6 Coordinates a Novel Golgi to ER Retrograde Transport Pathway in Live Cells. Journal of Cell Biology, 147(4):743–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/4/743.

Jin:1998:YND

[wJTSSL98] Quan wen Jin, Edgar Trelles-Sticken, Harry Scherthan, and Josef Loidl. Yeast Nuclei Display Prominent Centromere Clustering That Is Reduced in Nondividing Cells and in Meiotic Prophase. Journal of Cell Biology, 141(1):21–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/141/1/21.

Wolf:1998:GSD

[WJWM^{+}98] Anne A. Wolf, Michael G. Jobling, Susan Wimer-Mackin, Margaret Ferguson-Maltzman, James L. Madara, Randall K. Holmes, and Wayne I. Lencer. Ganglioside Structure Dictates Signal Transduction by Cholera Toxin and Association with Caveolae-like Membrane Domains in Polarized Epithelia. Journal of Cell Biology, 141(4):917–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/4/917.

Woolf:1995:RHG

[WKJH^{+}95] A. S. Woolf, M. Kolatsi-Joannou, P. Hardman, E. Andermarcher, C. Moorby, L. G. Fine, P. S. Jat, M. D. Noble, and E. Gherardi. Roles of hepatocyte growth factor/scatter factor and the met receptor in the early development of the metanephros. Journal of Cell Biology, 128(1):171–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
1/171.

Wilkerson:1995:MRI

[WKK^{+}95] C. G. Wilkerson, S. M. King, A. Koutoulis, G. J. Pazour, and G. B. Witman. The 78,000 M(r) intermediate chain of Chlamydomonas outer arm dynein isa WD-repeat protein required for arm assembly. Journal of Cell Biology, 129(1):169–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
1/169.

Wada:1996:CMI

[WKL^{+}96] J. Wada, A. Kumar, Z. Liu, E. Ruoslahti, L. Reichardt, J. Marvaldi, and Y. S. Kanwar. Cloning of mouse integrin alphaV cDNA and role of the alphaV-related matrix receptors in metanephric development. Journal of Cell Biology, 132(6): 1161–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/6/1161.

Wang:1999:AKP

[WKL^{+}99] Xiaolu Wang, Mark Kibschull, Michael M. Laue, Beate Lichte, Elisabeth Petrasch-Parwez, and Manfred W. Kilimann. Aczonin, a 550-Kd Putative Scaffolding Protein of Presynaptic Active Zones, Shares Homology Regions with Rim and Bassoon and Binds Profilin. Journal of Cell Biology, 147(1): 151–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/1/151.

Weiss:1998:VPC

[WKR^{+}98] Elisabeth E. Weiss, Martina Kroemker, Angelika-H. Rüdiger, Brigitte M. Jockusch, and Manfred Rüdiger. Vinculin Is Part of the Cadherin–Catenin Junctional Complex: Complex Formation between α-Catenin and Vinculin. Journal of Cell Biology, 141(3):755–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/141/3/755.

Weskamp:1996:MWE

[WKRB96] G. Weskamp, J. Krätzschmar, M. S. Reid, and C. P. Blobel. MDC9, a widely expressed cellular disintegrin containing cytoplasmic SH3 ligand domains. Journal of Cell Biology, 132(4): 717–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/4/717.

Waddle:1996:MCA

[WKWC96] J. A. Waddle, T. S. Karpova, R. H. Waterston, and J. A. Cooper. Movement of cortical actin patches in yeast. Journal of Cell Biology, 132(5):861–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/5/861.

Wang:1998:SDE

[WLD^{+}98] Huan-You Wang, Wen Lin, Jacqueline A. Dyck, Joanne M. Yeakley, Zhou Songyang, Lewis C. Cantley, and Xiang-Dong Fu. SRPK2: A differentially expressed sr protein-specific kinase involved in mediating the interaction and localization of pre-mRNA splicing factors in mammalian cells. Journal of Cell Biology, 140(4):737–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/4/737.

Ward:1998:CFD

[WLF98] Robert E. Ward, Rebecca S. Lamb, and Richard G. Fehon. A conserved functional domain of Drosophila coracle is required for localization at the septate junction and has membraneorganizing activity. Journal of Cell Biology, 140(6):1463–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
140/6/1463.

Wang:1999:IAP

[WLF99] Xue-Qing Wang, Frederik P. Lindberg, and William A. Frazier. Integrin-associated Protein Stimulates α2β1-Dependent Chemotaxis via GI–Mediated Inhibition of Adenylate Cyclase and Extracellular–Regulated Kinases. Journal of Cell Biology, 147(2):389–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/2/389.

Wennerberg:1996:BID

[WLG^{+}96] K. Wennerberg, L. Lohikangas, D. Gullberg, M. Pfaff, S. Johansson, and R. Fässler. Beta 1 integrin-dependent and - independent polymerization of fibronectin. Journal of Cell Biology, 132(1):227–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/1/227.

Ward:1997:HLF

[WLK97] Diane M. Ward, Jonathan D. Leslie, and Jerry Kaplan. Homotypic Lysosome Fusion in Macrophages: Analysis Using an In Vitro Assay. Journal of Cell Biology, 139(3):665–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
3/665.

Warren:1995:FEC

[WLML95] K. S. Warren, J. L. Lin, J. P. McDermott, and J. J. Lin. Forced expression of chimeric human fibroblast tropomyosin mutants affects cytokinesis. Journal of Cell Biology, 129(3):697–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
129/3/697.

Warmerdam:1996:IIC

[WLR96] P. A. Warmerdam, E. O. Long, and P. A. Roche. Isoforms of the invariant chain regulate transport of MHC class II molecules to antigen processing compartments. Journal of Cell Biology, 133(2):281–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/2/281.

Wordeman:1995:IPC

[WM95] L. Wordeman and T. J. Mitchison. Identification and partial characterization of mitotic centromere-associated kinesin, a kinesin-related protein that associates with centromeres during mitosis. Journal of Cell Biology, 128(1):95–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/128/
1/95.

Wells:1996:ASC

[WM96] W. A. Wells and A. W. Murray. Aberrantly segregating centromeres activate the spindle assembly checkpoint in budding yeast. Journal of Cell Biology, 133(1):75–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/133/1/75.

Wang:1995:RRP

[WMC^{+}95a] H. G. Wang, J. A. Millan, A. D. Cox, C. J. Der, U. R. Rapp, T. Beck, H. Zha, and J. C. Reed. R-Ras promotes apoptosis caused by growth factor deprivation via a Bcl-2 suppressible mechanism. Journal of Cell Biology, 129(4):1103–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
4/1103.

Wong:1995:INS

[WMC^{+}95b] P. C. Wong, J. Marszalek, T. O. Crawford, Z. Xu, S. T. Hsieh, J. W. Griffin, and D. W. Cleveland. Increasing neurofilament subunit NF–M expression reduces axonal NF–H, inhibits radial growth, and results in neurofilamentous accumulation in motor neurons. Journal of Cell Biology, 130(6): 1413–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/6/1413.

Winey:1995:TDU

[WMO^{+}95] M. Winey, C. L. Mamay, E. T. O’Toole, D. N. Mastronarde, T. H. Giddings, K. L. McDonald, and J. R. McIntosh. Threedimensional ultrastructural analysis of the Saccharomyces cerevisiae mitotic spindle. Journal of Cell Biology, 129(6): 1601–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1601.

Wu:1998:MCC

[WMP^{+}98] Xin Wu, Jon E. Mogford, Steven H. Platts, George E. Davis, Gerald A. Meininger, and Michael J. Davis. Modulation of calcium current in arteriolar smooth muscle by α_{v}β_{3} and α_{5}β_{1} integrin ligands. Journal of Cell Biology, 143(1):241–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
1/241.

Wedaman:1996:SSL

[WMR^{+}96] K. P. Wedaman, D. W. Meyer, D. J. Rashid, D. G. Cole, and J. M. Scholey. Sequence and submolecular localization of the 115-kD accessory subunit of the heterotrimeric kinesin-II (KRP85/95) complex. Journal of Cell Biology, 132(3): 371–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/3/371.

Weerasinghe:1998:RIT

[WMR^{+}98] Dheepika Weerasinghe, Kevin P. McHugh, Frederick P. Ross, Eric J. Brown, Roland H. Gisler, and Beat A. Imhof. A Role for the α vβ3 Integrin in the Transmigration of Monocytes. Journal of Cell Biology, 142(2):595–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/2/595.

Waser:1997:RCG

[WMSM97] Mathilde Waser, Nasrin Mesaeli, Charlotte Spencer, and Marek Michalak. Regulation of Calreticulin Gene Expression by Calcium. Journal of Cell Biology, 138(3):547–??, August 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/138/
3/547.

Wendland:1996:NFA

[WMXE96] B. Wendland, J. M. McCaffery, Q. Xiao, and S. D. Emr. A novel fluorescence-activated cell sorter-based screen for yeast endocytosis mutants identifies a yeast homologue of mammalian eps15. Journal of Cell Biology, 135(6):1485–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/6/1485.

Wiemer:1995:HPT

[WNB^{+}95] E. A. Wiemer, W. M. Nuttley, B. L. Bertolaet, X. Li, U. Francke, M. J. Wheelock, U. K. Anné, K. R. Johnson, and S. Subramani. Human peroxisomal targeting signal-1 receptor restores peroxisomal protein import in cells from patients with fatal peroxisomal disorders. Journal of Cell Biology, 130 (1):51–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/1/51.

West:1997:TSV

[WNB97] Anne E. West, Rachael L. Neve, and Kathleen M. Buckley. Targeting of the Synaptic Vesicle Protein Synaptobrevin in the Axon of Cultured Hippocampal Neurons: Evidence for Two Distinct Sorting Steps. Journal of Cell Biology, 139(4): 917–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/917.

Wu:1998:RNN

[WOM98] L. Wu, S. A. Osmani, and P. M. Mirabito. A Role for NIMA in the Nuclear Localization of Cyclin B in Aspergillus nidulans. Journal of Cell Biology, 141(7):1575–??, June 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/7/1575.

Ward:1995:SEI

[WPLK95] D. M. Ward, C. M. Perou, M. Lloyd, and J. Kaplan. ”synchronized” endocytosis and intracellular sorting in alveolar macrophages: the early sorting endosome is a transient organelle. Journal of Cell Biology, 129(5):1229–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/129/5/
1229.

Wolfsberg:1995:ANF

[WPMW95] T. G. Wolfsberg, P. Primakoff, D. G. Myles, and J. M. White. ADAM, a novel family of membrane proteins containing A Disintegrin And Metalloprotease domain: multipotential functions in cell–cell and cell-matrix interactions. Journal of Cell Biology, 131(2):275–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/131/2/275.

Wang:1996:FMD

[WPS^{+}96] C. Wang, W. F. Pralong, M. F. Schulz, G. Rougon, J. M. Aubry, S. Pagliusi, A. Robert, and J. Z. Kiss. Functional N– methyl–D–aspartate receptors in O-2A glial precursor cells: a critical role in regulating polysialic acid-neural cell adhesion molecule expression and cell migration. Journal of Cell Biology, 135(6):1565–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/135/6/1565.

Wess:1998:CDS

[WPS^{+}98] T. J. Wess, P. P. Purslow, M. J. Sherratt, J. Ashworth, C. A. Shuttleworth, and C. M. Kielty. Calcium Determines the Supramolecular Organization of Fibrillin-rich Microfibrils. Journal of Cell Biology, 141(3):829–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/3/829.

Weaver:1997:RMP

[WPW^{+}97] V. M. Weaver, O. W. Petersen, F. Wang, C. A. Larabell, P. Briand, C. Damsky, and M. J. Bissell. Reversion of the Malignant Phenotype of Human Breast Cells in Three–Dimensional Culture and In Vivo by Integrin Blocking Antibodies. Journal of Cell Biology, 137(1):231–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
1/231.

Wang:1996:RMM

[WPZ^{+}96] X. M. Wang, J. G. Peloquin, Y. Zhai, J. C. Bulinski, and G. G. Borisy. Removal of MAP4 from microtubules in vivo produces no observable phenotype at the cellular level. Journal of Cell Biology, 132(3):345–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/132/3/345.

Walker:1999:MTG

[WRC^{+}99] Macie B. Walker, Laura M. Roy, Eric Coleman, Rodger Voelker, and Alice Barkan. The Maize tha4 Gene Functions in Sec–Independent Protein Transport in Chloroplasts and Is Related to hcf106, tatA, and tatB. Journal of Cell Biology, 147 (2):267–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/147/2/267.

Waterham:1997:PTI

[WRdVC97] Hans R. Waterham, Kimberly A. Russell, Yne de Vries, and James M. Cregg. Peroxisomal Targeting, Import, and Assembly of Alcohol Oxidase in Pichia pastoris. Journal of Cell Biology, 139(6):1419–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/6/1419.

Wong:1998:EFE

[WRG98] Melissa H. Wong, Bonnee Rubinfeld, and Jeffrey I. Gordon. Effects of Forced Expression of an NH_{2}-terminal Truncated β-Catenin on Mouse Intestinal Epithelial Homeostasis. Journal of Cell Biology, 141(3):765–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/3/765.

Williams:1995:DKL

[WRW^{+}95] B. C. Williams, M. F. Riedy, E. V. Williams, M. Gatti, and M. L. Goldberg. The Drosophila kinesin-like protein KLP3A is a midbody component required for central spindle assembly and initiation of cytokinesis. Journal of Cell Biology, 129 (3):709–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/3/709.

Walensky:1995:ITR

[WS95a] L. D. Walensky and S. H. Snyder. Inositol 1,4,5-trisphosphate receptors selectively localized to the acrosomes of mammalian sperm. Journal of Cell Biology, 130(4):857–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/130/4/
857.

Weisenberger:1995:PMI

[WS95b] D. Weisenberger and U. Scheer. A possible mechanism for the inhibition of ribosomal RNA gene transcription during mitosis. Journal of Cell Biology, 129(3):561–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/3/561.

Wahle:1996:MTH

[WS96] S. Wahle and W. Stoffel. Membrane topology of the highaffinity L-glutamate transporter (GLAST-1) of the central nervous system. Journal of Cell Biology, 135(6):1867–??, December 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
135/6/1867.

Wahlberg:1997:MDD

[WS97] Johanna M. Wahlberg and Martin Spiess. Multiple Determinants Direct the Orientation of Signal–Anchor Proteins: The Topogenic Role of the Hydrophobic Signal Domain. Journal of Cell Biology, 137(3):555–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/3/555.

Wood:1998:SCB

[WS98] S. J. Wood and C. R. Slater. β-spectrin Is Colocalized with Both Voltage-gated Sodium Channels and Ankyrin_{G} at the Adult Rat Neuromuscular Junction. Journal of Cell Biology, 140(3):675–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/140/3/675.

Weipoltshammer:1999:IAR

[WSA^{+}99] Klara Weipoltshammer, Christian Schöfer, Marlene Almeder, Vlada V. Philimonenko, Klemens Frei, Franz Wachtler, and Pavel Hozák. Intranuclear Anchoring of Repetitive DNA Sequences. Journal of Cell Biology, 147(7):1409–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1409.

Walch-Solimena:1995:SSS

[WSBE^{+}95] C. Walch-Solimena, J. Blasi, L. Edelmann, E. R. Chapman, G. F. von Mollard, and R. Jahn. The t–SNAREs syntaxin 1 and SNAP-25 are present on organelles that participate in synaptic vesicle recycling. Journal of Cell Biology, 128(4): 637–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/4/637.

Walch-Solimena:1997:SMN

[WSCN97] Christiane Walch-Solimena, Ruth N. Collins, and Peter J. Novick. Sec2p mediates nucleotide exchange on Sec4p and is involved in polarized delivery of post-Golgi vesicles. Journal of Cell Biology, 137(7):1495–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/137/7/1495.

Way:1995:SDO

[WSG^{+}95] M. Way, M. Sanders, C. Garcia, J. Sakai, and P. Matsudaira. Sequence and domain organization of scruin, an actin-cross-linking protein in the acrosomal process of Limulus sperm.

Journal of Cell Biology, 128(1):51–??, January 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/128/1/51.

Wunsch:1998:DSH

[WSG^{+}98] Stefan Wünsch, Cecilia P. Sanchez, Michael Gekle, Lars Große-Wortmann, Jochen Wiesner, and Michael Lanzer. Differential stimulation of the Na^{+} /H^{+} exchanger determines chloroquine uptake in Plasmodium falciparum. Journal of Cell Biology, 140(2):335–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/2/335.

Waterman-Storer:1995:MMT

[WSGPS95] C. M. Waterman-Storer, J. Gregory, S. F. Parsons, and E. D. Salmon. Membrane/ microtubule tip attachment complexes (TACs) allow the assembly dynamics of plus ends to push and pull membranes into tubulovesicular networks in interphase Xenopus egg extracts. Journal of Cell Biology, 130(5): 1161–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/5/1161.

Wendel:1998:BMP

[WSH98] Mikael Wendel, Yngve Sommarin, and Dick Heinegård. Bone Matrix Proteins: Isolation and Characterization of a Novel Cell-binding Keratan Sulfate Proteoglycan (Osteoadherin) from Bovine Bone. Journal of Cell Biology, 141(3):839–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/3/839.

Weiss:1997:ERC

[WSR^{+}97] Johannes M. Weiss, Jonathan Sleeman, Andreas C. Renkl, Henning Dittmar, Christian C. Termeer, Sabine Taxis, Norma Howells, Martin Hofmann, Gabriele Köhler, Erwin Schöpf, Helmut Ponta, Peter Herrlich, and Jan C. Simon. An Essential Role for CD44 Variant Isoforms in Epidermal Langerhans Cell and Blood Dendritic Cell Function. Journal of Cell Biology, 137(5):1137–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/5/1137.

Waterman-Storer:1997:ABR

[WSS97] Clare M. Waterman-Storer and E. D. Salmon. Actomyosinbased Retrograde Flow of Microtubules in the Lamella of Migrating Epithelial Cells Influences Microtubule Dynamic Instability and Turnover and Is Associated with Microtubule Breakage and Treadmilling. Journal of Cell Biology, 139(2): 417–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/417.

Wei:1999:TDV

[WSSB99] Xiangyun Wei, Suryanarayan Somanathan, Jagath Samarabandu, and Ronald Berezney. Three-dimensional Visualization of Transcription Sites and Their Association with Splicing Factor–Rich Nuclear Speckles. Journal of Cell Biology, 146 (3):543–??, August 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/3/543.

Wanker:1995:FCK

[WSSM95] E. E. Wanker, Y. Sun, A. J. Savitz, and D. I. Meyer. Functional characterization of the 180-kD ribosome receptor in vivo. Journal of Cell Biology, 130(1):29–??, July 1995. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/130/1/29.

Wang:1995:NCR

[WSW95] S. Wang, H. Sakai, and M. Wiedmann. NAC covers ribosomeassociated nascent chains thereby forming a protective environment for regions of nascent chains just emerging from the peptidyl transferase center. Journal of Cell Biology, 130(3): 519–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/3/519.

Wojciak-Stothard:1999:MAS

[WSWR99] Beata Wójciak-Stothard, Lynn Williams, and Anne J. Ridley. Monocyte Adhesion and Spreading on Human Endothelial Cells Is Dependent on Rho-regulated Receptor Clustering. Journal of Cell Biology, 145(6):1293–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/6/1293.

Whalley:1995:DMR

[WTCV95] T. Whalley, M. Terasaki, M. S. Cho, and S. S. Vogel. Direct membrane retrieval into large vesicles after exocytosis in sea urchin eggs. Journal of Cell Biology, 131(5):1183–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
131/5/1183.

Wendt:1999:VHH

[WTM^{+}99] Thomas Wendt, Dianne Taylor, Terri Messier, Kathleen M. Trybus, and Kenneth A. Taylor. Visualization of Head– Head Interactions in the Inhibited State of Smooth Muscle Myosin. Journal of Cell Biology, 147(7):1385–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
7/1385.

Watabe-Uchida:1998:CVI

[WUUI^{+}98] Mitsuko Watabe-Uchida, Naoshige Uchida, Yuzo Imamura, Akira Nagafuchi, Kazushi Fujimoto, Tadashi Uemura, Stefan Vermeulen, Frans van Roy, Eileen D. Adamson, and Masatoshi Takeichi. α-catenin-vinculin Interaction Functions to Organize the Apical Junctional Complex in Epithelial Cells. Journal of Cell Biology, 142(3):847–??, August 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/142/3/847.

Wesseling:1995:EMO

[WvdVV^{+}95] J. Wesseling, S. W. van der Valk, H. L. Vos, A. Sonnenberg, and J. Hilkens. Episialin (MUC1) overexpression inhibits integrin-mediated cell adhesion to extracellular matrix components. Journal of Cell Biology, 129(1):255–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
1/255.

Walczak:1997:XKR

[WVM97] Claire E. Walczak, Suzie Verma, and Timothy J. Mitchison. XCTK2: a Kinesin-related Protein That Promotes Mitotic Spindle Assembly in Xenopus laevis Egg Extracts. Journal of Cell Biology, 136(4):859–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/4/859.

Wu:1995:GPB

[WVVD95] L. Wu, R. Valkema, P. J. Van Haastert, and P. N. Devreotes. The G protein beta subunit is essential for multiple responses to chemoattractants in Dictyostelium. Journal of Cell Biology, 129(6):1667–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1667.

Weiss:1996:SCS

[WW96a] E. Weiss and M. Winey. The Saccharomyces cerevisiae spindle pole body duplication gene MPS1 is part of a mitotic checkpoint. Journal of Cell Biology, 132(1):111–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
1/111.

Wheatley:1996:MMB

[WW96b] S. P. Wheatley and Y. Wang. Midzone microtubule bundles are continuously required for cytokinesis in cultured epithelial cells. Journal of Cell Biology, 135(4):981–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/135/
4/981.

Wiemer:1997:VPC

[WWD^{+}97] Erik A. C. Wiemer, Thibaut Wenzel, Thomas J. Deerinck, Mark H. Ellisman, and Suresh Subramani. Visualization of the Peroxisomal Compartment in Living Mammalian Cells: Dynamic Behavior and Association with Microtubules. Journal of Cell Biology, 136(1):71–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/1/71.

Wiesner:1998:CNG

[WWM^{+}98] Burkhard Wiesner, Jocelyn Weiner, Ralf Middendorff, Volker Hagen, U. Benjamin Kaupp, and Ingo Weyand. Cyclic Nucleotide-gated Channels on the Flagellum Control Ca^{2+} Entry into Sperm. Journal of Cell Biology, 142(2):473–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
2/473.

Wiley:1998:RMA

[WWO^{+}98] H. Steven Wiley, Margaret F. Woolf, Lee K. Opresko, Patrick M. Burke, Birgit Will, Jeffrey R. Morgan, and Douglas A. Lauffenburger. Removal of the Membrane-anchoring Domain of Epidermal Growth Factor Leads to Intracrine Signaling and Disruption of Mammary Epithelial Cell Organization. Journal of Cell Biology, 143(5):1317–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/143/
5/1317.

Wilding:1996:LPC

[WWP^{+}96] M. Wilding, E. M. Wright, R. Patel, G. Ellis-Davies, and M. Whitaker. Local perinuclear calcium signals associated with mitosis-entry in early sea urchin embryos. Journal of Cell Biology, 135(1):191–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/1/191.

Wang:1996:DMB

[WXS96] H. Wang, Z. Xie, and R. E. Scott. Differentiation modulates the balance of positive and negative Jun/AP-1 DNA binding activities to regulate cellular proliferative potential: different effects in nontransformed and transformed cells. Journal of Cell Biology, 135(4):1151–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/4/1151.

Wei:1999:RCU

[WYL^{+}99] Ying Wei, Xiuwei Yang, Qiumei Liu, John A. Wilkins, and Harold A. Chapman. A Role for Caveolin and the Urokinase Receptor in Integrin-mediated Adhesion and Signaling. Journal of Cell Biology, 144(6):1285–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/6/1285.

Watanabe:1995:KGN

[WYY95] K. Watanabe, H. Yamada, and Y. Yamaguchi. K-glypican: a novel GPI-anchored heparan sulfate proteoglycan that is highly expressed in developing brain and kidney. Journal of Cell Biology, 130(5):1207–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/130/5/1207.

Wang:1997:RMA

[WZF97] Xiao Min Wang, Ye Zhai, and James E. Ferrell. A Role for Mitogen-activated Protein Kinase in the Spindle Assembly Checkpoint in XTC Cells. Journal of Cell Biology, 137 (2):433–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/2/433.

Xu:1996:EMA

[XC96] J. Xu and R. A. Clark. Extracellular matrix alters PDGF regulation of fibroblast integrins. Journal of Cell Biology, 132 (1):239–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/1/239.

Xu:1997:TDC

[XC97] Jiahua Xu and Richard A. F. Clark. A three-dimensional collagen lattice induces protein kinase C-ζ activity: Role in α_{2} integrin and collagenase mRNA expression. Journal of Cell Biology, 136(2):473–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/2/473.

Xu:1998:CAS

[XCLM98] Xian-Zhong Shawn Xu, Atish Choudhury, Xiaoling Li, and Craig Montell. Coordination of an Array of Signaling Proteins through Homo- and Heteromeric Interactions Between PDZ Domains and Target Proteins. Journal of Cell Biology, 142 (2):545–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/545.

Xia:1996:AMI

[XGC96] Y. Xia, S. G. Gil, and W. G. Carter. Anchorage mediated by integrin alpha6beta4 to laminin 5 (epiligrin) regulates tyrosine phosphorylation of a membrane-associated 80kD protein. Journal of Cell Biology, 132(4):727–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/132/
4/727.

Xiao:1995:RMS

[XGR95] Y. Xiao, U. Grieshammer, and N. Rosenthal. Regulation of a muscle-specific transgene by retinoic acid. Journal of Cell Biology, 129(5):1345–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/5/1345.

Xu:1996:MFS

[XHUC96] J. Q. Xu, B. A. Harder, P. Uman, and R. Craig. Myosin filament structure in vertebrate smooth muscle. Journal of Cell Biology, 134(1):53–??, July 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/134/1/53.

Xing:1995:NGO

[XJM^{+}95] Y. Xing, C. V. Johnson, P. T. Moen, J. A. McNeil, and J. Lawrence. Nonrandom gene organization: structural arrangements of specific pre-mRNA transcription and splicing with SC-35 domains. Journal of Cell Biology, 131(6):1635– ??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/6/1635.

Xu:1996:SCN

[XML^{+}96] Z. Xu, J. R. Marszalek, M. K. Lee, P. C. Wong, J. Folmer, T. O. Crawford, S. T. Hsieh, J. W. Griffin, and D. W. Cleveland. Subunit composition of neurofilaments specifies axonal diameter. Journal of Cell Biology, 133(5):1061–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
5/1061.

Xu:1997:HTC

[XMMW97] Zuoyu Xu, Andreas Mayer, Eric Muller, and William Wickner. A heterodimer of thioredoxin and I_{2}^{B} cooperates with Sec18p (NSF) to promote yeast vacuole inheritance. Journal of Cell Biology, 136(2):299–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/136/2/299.

Xu:1996:TRV

[XW96] Z. Xu and W. Wickner. Thioredoxin is required for vacuole inheritance in Saccharomyces cerevisiae. Journal of Cell Biology, 132(5):787–??, March 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/5/787.

Xia:1997:AAL

[XWK^{+}97] Houhui Xia, Sara T. Winokur, Wen-Lin Kuo, Michael R. Altherr, and David S. Bredt. Actinin-associated LIM Protein: Identification of a Domain Interaction between PDZ and Spectrin-like Repeat Motifs. Journal of Cell Biology, 139(2): 507–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/2/507.

Xiao:1997:DDC

[XZMD97] Zhan Xiao, Ning Zhang, Douglas B. Murphy, and Peter N. Devreotes. Dynamic Distribution of Chemoattractant Receptors in Living Cells During Chemotaxis and Persistent Stimulation. Journal of Cell Biology, 139(2):365–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/2/365.

Xu:1996:PIA

[XZSC96] J. Xu, M. M. Zutter, S. A. Santoro, and R. A. Clark. PDGF induction of alpha 2 integrin gene expression is mediated by protein kinase C-zeta. Journal of Cell Biology, 134(5):1301– ??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/5/1301.

Xu:1998:TDC

[XZSC98] Jiahua Xu, Mary M. Zutter, Samuel A. Santoro, and Richard A. F. Clark. A Three-dimensional Collagen Lattice Activates NF-κB in Human Fibroblasts: Role in Integrin α_{2} Gene Expression and Tissue Remodeling. Journal of Cell Biology, 140 (3):709–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/3/709.

Yao:1997:MDM

[YAC97] Xuebiao Yao, Karen L. Anderson, and Don W. Cleveland. The Microtubule-dependent Motor Centromere–associated Protein E (CENP–E) Is an Integral Component of Kinetochore Corona Fibers That Link Centromeres to Spindle Microtubules. Journal of Cell Biology, 139(2):435–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/2/435.

Yang:1997:RAC

[YAD97] Shirley Yang, Kathryn R. Ayscough, and David G. Drubin. A Role for the Actin Cytoskeleton of Saccharomyces cerevisiae in Bipolar Bud–Site Selection. Journal of Cell Biology, 136 (1):111–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/1/111.

Yasui:1998:RRA

[YAiN^{+}98] Yoshihiro Yasui, Mutsuki Amano, Koh ichi Nagata, Naoyuki Inagaki, Hideo Nakamura, Hideyuki Saya, Kozo Kaibuchi, and Masaki Inagaki. Roles of Rho-associated Kinase in Cytokinesis; Mutations in Rho-associated Kinase Phosphorylation Sites Impair Cytokinetic Segregation of Glial Filaments. Journal of Cell Biology, 143(5):1249–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/5/1249.

Yohn:1998:TCP

[YCR^{+}98] Christopher B. Yohn, Amybeth Cohen, Cristen Rosch, Michael R. Kuchka, and Stephen P. Mayfield. Translation of the Chloroplast psbA mRNA Requires the Nuclear-encoded Poly(A)-binding Protein, RB47. Journal of Cell Biology, 142 (2):435–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/2/435.

Yang:1998:KAP

[YDFB98] Jiacheng Yang, Judith A. Drazba, Donald G. Ferguson, and Meredith Bond. A-kinase Anchoring Protein 100 (AKAP100) is Localized in Multiple Subcellular Compartments in the Adult Rat Heart. Journal of Cell Biology, 142(2):511–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
2/511.

Ye:1999:AAW

[YF99] Yihong Ye and Mark E. Fortini. Apoptotic activities of wildtype and Alzheimer’s disease-related mutant presenilins in Drosophila melanogaster. Journal of Cell Biology, 146(6): 1351–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/6/1351.

Yoshida:1996:LAB

[YFST^{+}96] S. Yoshida, A. Fujisawa-Sehara, T. Taki, K. Arai, and Y. Nabeshima. Lysophosphatidic acid and bFGF control different modes in proliferating myoblasts. Journal of Cell Biology, 132(1):181–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/1/181.

Yang:1997:IMP

[YGG97a] Li Yang, Tinglu Guan, and Larry Gerace. Integral Membrane Proteins of the Nuclear Envelope Are Dispersed throughout the Endoplasmic Reticulum during Mitosis. Journal of Cell Biology, 137(6):1199–??, June 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/6/1199.

Yang:1997:LBF

[YGG97b] Li Yang, Tinglu Guan, and Larry Gerace. Lamin-binding Fragment of LAP2 Inhibits Increase in Nuclear Volume during the Cell Cycle and Progression into S Phase. Journal of Cell Biology, 139(5):1077–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/139/5/1077.

Yamamoto:1996:PIA

[YGK96a] A. Yamamoto, V. Guacci, and D. Koshland. Pds1p, an inhibitor of anaphase in budding yeast, plays a critical role in the APC and checkpoint pathway(s). Journal of Cell Biology, 133(1):99–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/1/99.

Yamamoto:1996:PRF

[YGK96b] A. Yamamoto, V. Guacci, and D. Koshland. Pds1p is required for faithful execution of anaphase in the yeast, Saccharomyces cerevisiae. Journal of Cell Biology, 133(1):85–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/133/1/85.

Yu:1997:NMC

[YHC^{+}97] Hong-Guo Yu, Evelyn N. Hiatt, Annette Chan, Mary Sweeney, and R. Kelly Dawe. Neocentromere-mediated Chromosome Movement in Maize. Journal of Cell Biology, 139(4):831– ??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/4/831.

Yonemura:1998:ERM

[YHD^{+}98] Shigenobu Yonemura, Motohiro Hirao, Yoshinori Doi, Nobuyuki Takahashi, Takahisa Kondo, Sachiko Tsukita, and Shoichiro Tsukita. Ezrin/ Radixin/ moesin (ERM) Proteins Bind to a Positively Charged Amino Acid Cluster in the Juxta– Membrane Cytoplasmic Domain of CD44, CD43, and ICAM-2. Journal of Cell Biology, 140(4):885–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/140/4/
885.

Yamamoto:1997:RTA

[YHK^{+}97] Takaharu Yamamoto, Naozumi Harada, Kyoko Kano, Shin ichiro Taya, Eli Canaani, Yoshiharu Matsuura, Akira Mizoguchi, Chizuka Ide, and Kozo Kaibuchi. The Ras Target AF-6 Interacts with ZO-1 and Serves as a Peripheral Component of Tight Junctions in Epithelial Cells. Journal of Cell Biology, 139(3):785–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/139/3/785.

Yamamoto-Hino:1998:AVB

[YHMS^{+}98] Miki Yamamoto-Hino, Atsushi Miyawaki, Akihisa Segawa, Eijiro Adachi, Shohei Yamashina, Toyoshi Fujimoto, Tomoyasu Sugiyama, Teiichi Furuichi, Mamoru Hasegawa, and Katsuhiko Mikoshiba. Apical Vesicles Bearing Inositol 1,4,5trisphosphate Receptors in the Ca^{2+} Initiation Site of Ductal Epithelium of Submandibular Gland. Journal of Cell Biology, 141(1):135–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/1/135.

Yonekawa:1998:DSV

[YHO^{+}98] Yoshiaki Yonekawa, Akihiro Harada, Yasushi Okada, Takeshi Funakoshi, Yoshimitsu Kanai, Yosuke Takei, Sumio Terada, Tetsuo Noda, and Nobutaka Hirokawa. Defect in Synaptic Vesicle Precursor Transport and Neuronal Cell Death in KIF1A Motor Protein–deficient Mice. Journal of Cell Biology, 141(2):431–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/141/2/431.

Yamaguchi:1999:SAE

[YHS^{+}99] Atsushi Yamaguchi, Osamu Hori, David M. Stern, Enno Hartmann, Satoshi Ogawa, and Masaya Tohyama. Stressassociated endoplasmic reticulum protein 1 (Serp1)/ribosomeassociated membrane protein 4 (Ramp4) stabilizes membrane proteins during stress and facilitates subsequent glycosylation. Journal of Cell Biology, 147(6):1195–??, December 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
6/1195.

Yuen:1995:DSF

[YJB^{+}95] I. S. Yuen, R. Jain, J. D. Bishop, D. F. Lindsey, W. J. Deery, P. J. Van Haastert, and R. H. Gomer. A density-sensing factor regulates signal transduction in Dictyostelium. Journal of Cell Biology, 129(5):1251–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/5/1251.

Yang:1995:IDP

[YKC95] T. Yang, K. M. Kozopas, and R. W. Craig. The intracellular distribution and pattern of expression of Mcl-1 overlap with, but are not identical to, those of Bcl-2. Journal of Cell Biology, 128(6):1173–??, March 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/128/6/1173.

Yu:1997:RSU

[YKO97] W. Yu, J. Kim, and L. Ossowski. Reduction in Surface Urokinase Receptor Forces Malignant Cells into a Protracted State of Dormancy. Journal of Cell Biology, 137(3):767–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/137/
3/767.

Yoshimori:1996:DBT

[YKRS96] T. Yoshimori, P. Keller, M. G. Roth, and K. Simons. Different biosynthetic transport routes to the plasma membrane in BHK and CHO cells. Journal of Cell Biology, 133(2):247–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
2/247.

Yeaman:1997:GSD

[YLB^{+}97] Charles Yeaman, Annick H. Le Gall, Anne N. Baldwin, Laure Monlauzeur, Andre Le Bivic, and Enrique Rodriguez-Boulan. The O-glycosylated stalk domain is required for apical sorting of neurotrophin receptors in polarized MDCK cells. Journal of Cell Biology, 139(4):929–??, November 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/139/4/929.

Yang:1997:STG

[YLD97] Jun Yang, Jimin Liu, and Donald B. DeFranco. Subnuclear Trafficking of Glucocorticoid Receptors In Vitro: Chromatin Recycling and Nuclear Export. Journal of Cell Biology, 137 (3):523–??, May 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/137/3/523.

Yanez-Mo:1998:REC

[YMAC^{+}98] María Yáñez-Mó, Arántzazu Alfranca, Carlos Cabañas, Mónica Marazuela, Reyes Tejedor, M. Angeles Ursa, Leonie K. Ashman, Manuel O. de Landázuri, and Francisco Sánchez-Madrid. Regulation of endothelial cell motility by complexes of tetraspan molecules CD81/TAPA-1 and CD151/PETA-3 with α3β1 integrin localized at endothelial lateral junctions. Journal of Cell Biology, 141(3):791–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/141/3/791.

Yu:1999:MHC

[YMD99] Hong-Guo Yu, Michael G. Muszynski, and R. Kelly Dawe. The Maize Homologue of the Cell Cycle Checkpoint Protein MAD2 Reveals Kinetochore Substructure and Contrasting Mitotic and Meiotic Localization Patterns. Journal of Cell Biology, 145(3):425–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/145/3/425.

Yoon:1998:MPV

[YMPG98] Miri Yoon, Robert D. Moir, Veena Prahlad, and Robert D. Goldman. Motile Properties of Vimentin Intermediate Filament Networks in Living Cells. Journal of Cell Biology, 143 (1):147–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/143/1/147.

Yan:1995:ARD

[YN95] H. Yan and J. Newport. An analysis of the regulation of DNA synthesis by cdk2, Cip1, and licensing factor. Journal of Cell Biology, 129(1):1–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/1/1.

Yap:1998:JRC

[YNG98] Alpha S. Yap, Carien M. Niessen, and Barry M. Gumbiner. The Juxtamembrane Region of the Cadherin Cytoplasmic Tail Supports Lateral Clustering, Adhesive Strengthening, and Interaction with p120^{ctn}. Journal of Cell Biology, 141(3):779–??, May 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
141/3/779.

Yamaguchi:1999:LDT

[YNM^{+}99] Rumi Yamaguchi, Masanobu Nakamura, Nobuyoshi Mochizuki, Steve A. Kay, and Akira Nagatani. Light-dependent Translocation of a Phytochrome B–GFP Fusion Protein to the Nucleus in Transgenic Arabidopsis. Journal of Cell Biology, 145 (3):437–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/3/437.

Yamazaki:1995:KBH

[YNOH95] H. Yamazaki, T. Nakata, Y. Okada, and N. Hirokawa. KIF3A/ b: a heterodimeric kinesin superfamily protein that works as a microtubule plus end-directed motor for membrane organelle transport. Journal of Cell Biology, 130(6):1387–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
6/1387.

Yuan:1998:SCP

[YPB^{+}98] Li Yuan, Jeanette Pelttari, Eva Brundell, Birgitta Björkroth, Jian Zhao, Jian-Guo Liu, Hjalmar Brismar, Bertil Daneholt, and Christer Höög. The Synaptonemal Complex Protein SCP3 Can Form Multistranded, Cross-striated Fibers In Vivo. Journal of Cell Biology, 142(2):331–??, July 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/142/2/331.

Yoon:1998:NDL

[YPDM98] Yisang Yoon, Kelly R. Pitts, Sophie Dahan, and Mark A. Mc-Niven. A Novel Dynamin-like Protein Associates with Cytoplasmic Vesicles and Tubules of the Endoplasmic Reticulum in Mammalian Cells. Journal of Cell Biology, 140(4): 779–??, February 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/4/779.

Yuan:1997:RDD

[YPM97] Ruiyong Yuan, Paul Primakoff, and Diana G. Myles. A Role for the Disintegrin Domain of Cyritestin, a Sperm Surface Protein Belonging to the ADAM Family, in Mouse Sperm–Egg Plasma Membrane Adhesion and Fusion. Journal of Cell Biology, 137(1):105–??, April 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/137/1/105.

Yang:1996:GAA

[YRM^{+}96] J. T. Yang, T. A. Rando, W. A. Mohler, H. Rayburn, H. M. Blau, and R. O. Hynes. Genetic analysis of alpha 4 integrin functions in the development of mouse skeletal muscle. Journal of Cell Biology, 135(3):829–??, November 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/135/3/829.

Yu:1996:MTA

[YSB96] W. Yu, M. J. Schwei, and P. W. Baas. Microtubule transport and assembly during axon growth. Journal of Cell Biology, 133 (1):151–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/133/1/151.

Yeh:1995:SDC

[YSC^{+}95] E. Yeh, R. V. Skibbens, J. W. Cheng, E. D. Salmon, and K. Bloom. Spindle dynamics and cell cycle regulation of dynein in the budding yeast, Saccharomyces cerevisiae. Journal of Cell Biology, 130(3):687–??, August 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/130/3/687.

Yagami:1999:MGP

[YSEI^{+}99] Kimitoshi Yagami, Jo-Young Suh, Motomi Enomoto-Iwamoto, Eiki Koyama, William R. Abrams, Irving M. Shapiro, Maurizio Pacifici, and Masahiro Iwamoto. Matrix Gla Protein Is a Developmental Regulator of Chondrocyte Mineralization And, When Constitutively Expressed, Blocks Endochondral and Intramembranous Ossification in the Limb. Journal of Cell Biology, 147(5):1097–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/147/5/1097.

Yamazaki:1995:TCS

[YSK95] T. Yamazaki, D. J. Selkoe, and E. H. Koo. Trafficking of cell surface beta-amyloid precursor protein: retrograde and transcytotic transport in cultured neurons. Journal of Cell Biology, 129(2):431–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/2/431.

Yu:1997:IMM

[YSK^{+}97] Wenqian Yu, David J. Sharp, Ryoko Kuriyama, Prabhat Mallik, and Peter W. Baas. Inhibition of a Mitotic Motor Compromises the Formation of Dendrite-like Processes from Neuroblastoma Cells. Journal of Cell Biology, 136(3): 659–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/136/3/659.

Yang:1995:SRH

[YSM^{+}95] Y. Yang, E. Spitzer, D. Meyer, M. Sachs, C. Niemann, G. Hartmann, K. M. Weidner, C. Birchmeier, and W. Birchmeier. Sequential requirement of hepatocyte growth factor and neuregulin in the morphogenesis and differentiation of the mammary gland. Journal of Cell Biology, 131(1):215–??, October 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
1/215.

Yamashita:1995:OPB

[YtDH^{+}95] H. Yamashita, P. ten Dijke, D. Huylebroeck, T. K. Sampath, M. Andries, J. C. Smith, C. H. Heldin, and K. Miyazono. Osteogenic protein-1 binds to activin type II receptors and induces certain activin-like effects. Journal of Cell Biology, 130(1):217–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/130/1/217.

Yeakley:1999:PRV

[YTO^{+}99] Joanne M. Yeakley, Hélène Tronchère, James Olesen, Jacqueline A. Dyck, Huan-You Wang, and Xiang-Dong Fu. Phosphorylation regulates in vivo interaction and molecular targeting of serine/arginine-rich pre-mRNA splicing factors. Journal of Cell Biology, 145(3):447–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/3/447.

Yonemura:1999:DIE

[YTT99] Shigenobu Yonemura, Sachiko Tsukita, and Shoichiro Tsukita. Direct involvement of ezrin/radixin/moesin (ERM)-binding membrane proteins in the organization of microvilli in collaboration with activated ERM proteins. Journal of Cell Biology, 145(7):1497–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/145/7/1497.

Yamakita:1999:DFP

[YTY^{+}99] Yoshihiko Yamakita, Go Totsukawa, Shigeko Yamashiro, David Fry, Xiaoe Zhang, Steven K. Hanks, and Fumio Matsumura. Dissociation of FAK/p130^{CAS} /c–Src complex during mitosis: Role of mitosis-specific serine phosphorylation of FAK. Journal of Cell Biology, 144(2):315–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
2/315.

Yokota:1995:EOC

[YvdEH^{+}95] H. Yokota, G. van den Engh, J. E. Hearst, R. K. Sachs, and B. J. Trask. Evidence for the organization of chromatin in megabase pair-sized loops arranged along a random walk path in the human G0/G1 interphase nucleus. Journal of Cell Biology, 130(6):1239–??, September 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/130/6/1239.

Yamamoto:1999:CDH

[YWMH99] Ayumu Yamamoto, Robert R. West, J. Richard McIntosh, and Yasushi Hiraoka. A Cytoplasmic Dynein Heavy Chain Is Required for Oscillatory Nuclear Movement of Meiotic Prophase and Efficient Meiotic Recombination in Fission Yeast. Journal of Cell Biology, 145(6):1233–??, June 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/145/6/1233.

Yoshida:1996:LAV

[YWW^{+}96] M. Yoshida, W. F. Westlin, N. Wang, D. E. Ingber, A. Rosenzweig, N. Resnick, and M. A. Gimbrone. Leukocyte adhesion to vascular endothelium induces E-selectin linkage to the actin cytoskeleton. Journal of Cell Biology, 133(2):445–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/133/
2/445.

Yasunaga:1996:IFT

[YYH^{+}96] M. Yasunaga, T. Yagi, N. Hanzawa, M. Yasuda, Y. Yamanashi, T. Yamamoto, S. Aizawa, Y. Miyauchi, and S. Nishikawa. Involvement of Fyn tyrosine kinase in progression of cytokinesis of B lymphocyte progenitor. Journal of Cell Biology, 132 (1):91–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/132/1/91.

Yang:1997:IMA

[YYSB97] Sam S. Yang, Elaine Yeh, E. D. Salmon, and Kerry Bloom. Identification of a Mid-anaphase Checkpoint in Budding Yeast. Journal of Cell Biology, 136(2):345–??, January 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/136/2/345.

Zhang:1998:RKA

[ZB98] Xu Zhang and Vann Bennett. Restriction of 480/270kD Ankyrin_{G} to axon proximal segments requires multiple Ankyrin_{G}-specific domains. Journal of Cell Biology, 142(6): 1571–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/142/6/1571.

Zhong:1999:ACC

[ZBG99] Yun Zhong, William M. Brieher, and Barry M. Gumbiner. Analysis of C-cadherin Regulation during Tissue Morphogenesis with an Activating Antibody. Journal of Cell Biology, 144 (2):351–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/144/2/351.

Zhou:1997:SMB

[ZBW^{+}97] Daixing Zhou, Connie S. Birkenmeier, McRae W. Williams, John J. Sharp, Jane E. Barker, and Robert J. Bloch. Small, Membrane-bound, Alternatively Spliced Forms of Ankyrin 1 Associated with the Sarcoplasmic Reticulum of Mammalian Skeletal Muscle. Journal of Cell Biology, 136(3):621–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
136/3/621.

Zeng:1999:RAC

[ZC99] Guisheng Zeng and Mingjie Cai. Regulation of the actin cytoskeleton organization in yeast by a novel serine/threonine kinase Prk1p. Journal of Cell Biology, 144(1):71–??, January 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
1/71.

Zecevic:1998:AMK

[ZCE^{+}98] Maja Zecevic, Andrew D. Catling, Scott T. Eblen, Luigina Renzi, James C. Hittle, Tim J. Yen, Gary J. Gorbsky, and Michael J. Weber. Active MAP Kinase in Mitosis: Localization at Kinetochores and Association with the Motor Protein CENP–E. Journal of Cell Biology, 142(6):1547–??, September 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/142/
6/1547.

Zakharenko:1999:NSA

[ZCOP99] Stanislav Zakharenko, Sunghoe Chang, Michael O’Donoghue, and Sergey V. Popov. Neurotransmitter Secretion along Growing Nerve Processes: Comparison with Synaptic Vesicle Exocytosis. Journal of Cell Biology, 144(3):507–??, February 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/144/
3/507.

Zhong:1998:RMC

[ZCWB^{+}98] Cuiling Zhong, Magdalena Chrzanowska-Wodnicka, James Brown, Amy Shaub, Alexey M. Belkin, and Keith Burridge. Rho-mediated Contractility Exposes a Cryptic Site in Fibronectin and Induces Fibronectin Matrix Assembly. Journal of Cell Biology, 141(2):539–??, April 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/141/2/539.

Zeng:1999:PTP

[ZDK^{+}99] Li Zeng, Luca D’Alessandri, Markus B. Kalousek, Lloyd Vaughan, and Catherine J. Pallen. Protein Tyrosine Phosphatase α (Ptpα) and Contactin Form a Novel Neuronal Receptor Complex Linked to the Intracellular Tyrosine Kinase Fyn. Journal of Cell Biology, 147(4):707–??, November 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
4/707.

Zabel:1996:NRN

[ZDT^{+}96] U. Zabel, V. Doye, H. Tekotte, R. Wepf, P. Grandi, and E. C. Hurt. Nic96p is required for nuclear pore formation and functionally interacts with a novel nucleoporin, Nup188p. Journal of Cell Biology, 133(6):1141–??, June 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/6/1141.

Zhou:1999:DDM

[ZGYS99] Heather Zhou, David J. Glass, George D. Yancopoulos, and Joshua R. Sanes. Distinct Domains of Musk Mediate Its Abilities to Induce and to Associate with Postsynaptic Specializations. Journal of Cell Biology, 146(5):1133–??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/146/
5/1133.

Zegers:1997:STA

[ZH97] Mirjam M. P. Zegers and Dick Hoekstra. Sphingolipid Transport to the Apical Plasma Membrane Domain in Human Hepatoma Cells Is Controlled by PKC and PKA Activity: a Correlation with Cell Polarity in HepG2 Cells. Journal of Cell Biology, 138(2):307–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/138/2/307.

Zhang:1995:DEF

[ZHR95] H. Zhang, W. Hu, and F. Ramirez. Developmental expression of fibrillin genes suggests heterogeneity of extracellular microfibrils. Journal of Cell Biology, 129(4):1165–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
4/1165.

Zhang:1997:NOT

[ZhTA^{+}97] Bin Zhang, Pang hsien Tu, Farhad Abtahian, John Q. Trojanowski, and Virginia M.-Y. Lee. Neurofilaments and Orthograde Transport Are Reduced in Ventral Root Axons of Transgenic Mice that Express Human SOD1 with a G93A Mutation. Journal of Cell Biology, 139(5):1307–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/139/
5/1307.

Zigmond:1997:RAP

[ZJB^{+}97] Sally H. Zigmond, Michael Joyce, Jane Borleis, Gary M. Bokoch, and Peter N. Devreotes. Regulation of Actin Polymerization in Cell-free Systems by GTPγS and Cdc42. Journal of Cell Biology, 138(2):363–??, July 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/138/2/363.

Zigmond:1998:MCI

[ZJY^{+}98] Sally H. Zigmond, Michael Joyce, Changsong Yang, Kevin Brown, Minzhou Huang, and Martin Pring. Mechanism of Cdc42-induced Actin Polymerization in Neutrophil Extracts.

Journal of Cell Biology, 142(4):1001–??, August 1998. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/142/4/
1001.

Zhai:1995:KMD

[ZKB95] Y. Zhai, P. J. Kronebusch, and G. G. Borisy. Kinetochore microtubule dynamics and the metaphase-anaphase transition. Journal of Cell Biology, 131(3):721–??, November 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/131/
3/721.

Zheng:1995:MAC

[ZKH^{+}95] Z. Zheng, S. Katoh, Q. He, K. Oritani, K. Miyake, J. Lesley, R. Hyman, A. Hamik, R. M. Parkhouse, and A. G. Farr. Monoclonal antibodies to CD44 and their influence on hyaluronan recognition. Journal of Cell Biology, 130(2):485–??, July 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/130/
2/485.

Zhou:1999:DMB

[ZKN99] Lan Zhou, Eileen M. Kasperek, and Bruce J. Nicholson. Dissection of the Molecular Basis of pp60^{v−src} Induced Gating of Connexin 43 Gap Junction Channels. Journal of Cell Biology, 144(5):1033–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/144/5/1033.

Zeng:1999:SCP

[ZKS^{+}99] Xuemei Zeng, Jason A. Kahana, Pamela A. Silver, Mary K. Morphew, J. Richard McIntosh, Ian T. Fitch, John Carbon, and William S. Saunders. Slk19p Is a Centromere Protein That Functions to Stabilize Mitotic Spindles. Journal of Cell Biology, 146(2):415–??, July 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/146/2/415.

Zhai:1996:MDG

[ZKSB96] Y. Zhai, P. J. Kronebusch, P. M. Simon, and G. G. Borisy. Microtubule dynamics at the G2/M transition: abrupt breakdown of cytoplasmic microtubules at nuclear envelope breakdown and implications for spindle morphogenesis. Journal of Cell Biology, 135(1):201–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/135/1/201.

Zhang:1995:LTP

[ZL95a] J. S. Zhang and F. M. Longo. LAR tyrosine phosphatase receptor: alternative splicing is preferential to the nervous system, coordinated with cell growth and generates novel isoforms containing extensive CAG repeats. Journal of Cell Biology, 128 (3):415–??, February 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/128/3/415.

Zhang:1995:PPS

[ZL95b] J. W. Zhang and P. B. Lazarow. PEB1 (PAS7) in Saccharomyces cerevisiae encodes a hydrophilic, intra-peroxisomal protein that is a member of the WD repeat family and is essential for the import of thiolase into peroxisomes. Journal of Cell Biology, 129(1):65–??, April 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/129/1/65.

Zhu:1995:FMP

[ZL95c] X. Zhu and I. Lindberg. 7b2 facilitates the maturation of proPC2 in neuroendocrine cells and is required for the expression of enzymatic activity. Journal of Cell Biology, 129 (6):1641–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/129/6/1641.

Zhang:1996:PPI

[ZL96] J. W. Zhang and P. B. Lazarow. Peb1p (Pas7p) is an intraperoxisomal receptor for the NH2-terminal, type 2, peroxisomal targeting sequence of thiolase: Peb1p itself is targeted to peroxisomes by an NH2-terminal peptide. Journal of Cell Biology, 132(3):325–??, February 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/3/325.

Zollner:1997:SHM

[ZLB^{+}97] Olaf Zöllner, Martin C. Lenter, James E. Blanks, Eric Borges, Martin Steegmaier, Hans-Günther Zerwes, and Dietmar Vestweber. L-selectin from Human, but Not from Mouse Neutrophils Binds Directly to E–Selectin. Journal of Cell Biology, 136(3):707–??, February 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/136/3/707.

Zhang:1996:ICD

[ZLC^{+}96] C. X. Zhang, M. P. Lee, A. D. Chen, S. D. Brown, and T. Hsieh. Isolation and characterization of a Drosophila gene essential for early embryonic development and formation of cortical cleavage furrows. Journal of Cell Biology, 134(4): 923–??, August 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/4/923.

Zhu:1998:DNH

[ZLL^{+}98] Qinzhang Zhu, Michael Lindenbaum, Françoise Levavasseur, Hélène Jacomy, and Jean-Pierre Julien. Disruption of the NF– H Gene Increases Axonal Microtubule Content and Velocity of Neurofilament Transport: Relief of Axonopathy Resulting from the Toxin β,β’-Iminodipropionitrile. Journal of Cell Biology, 143(1):183–??, October 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/1/183.

Zhou:1998:ARC

[ZLM^{+}98] Daixing Zhou, Stephen Lambert, Peter L. Malen, Scott Carpenter, Linda M. Boland, and Vann Bennett. Ankyrin_{G} Is Required for Clustering of Voltage-gated Na Channels at Axon Initial Segments and for Normal Action Potential Firing. Journal of Cell Biology, 143(5):1295–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/143/5/1295.

Zeidman:1999:PRD

[ZLPL99] Ruth Zeidman, Bjarne Löfgren, Sven Påhlman, and Christer Larsson. PKCϵ, via its regulatory domain and independently of its catalytic domain, induces neurite-like processes in neuroblastoma cells. Journal of Cell Biology, 145(4):713–??, May 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/145/
4/713.

Zarnegar:1995:MFH

[ZM95] R. Zarnegar and G. K. Michalopoulos. The many faces of hepatocyte growth factor: from hepatopoiesis to hematopoiesis. Journal of Cell Biology, 129(5):1177–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/5/1177.

Zimmerli:1996:PLF

[ZMG^{+}96a] S. Zimmerli, M. Majeed, M. Gustavsson, O. Stendahl, D. A. Sanan, and J. D. Ernst. Phagosome-lysosome fusion is a calcium-independent event in macrophages. Journal of Cell Biology, 132(1):49–??, January 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/132/1/49.

Zondag:1996:LAB

[ZMG96b] G. C. Zondag, W. H. Moolenaar, and M. F. Gebbink. Lack of association between receptor protein tyrosine phosphatase RPTP mu and cadherins. Journal of Cell Biology, 134(6): 1513–??, September 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/134/6/1513.

Zambruno:1995:TGF

[ZMM^{+}95] G. Zambruno, P. C. Marchisio, A. Marconi, C. Vaschieri, A. Melchiori, A. Giannetti, and M. De Luca. Transforming growth factor-beta 1 modulates beta 1 and beta 5 integrin receptors and induces the de novo expression of the alpha v beta 6 heterodimer in normal human keratinocytes: implications for wound healing. Journal of Cell Biology, 129(3):853–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/129/
3/853.

Zhang:1995:CIS

[ZN95a] D. Zhang and R. B. Nicklas. Chromosomes initiate spindle assembly upon experimental dissolution of the nuclear envelope in grasshopper spermatocytes. Journal of Cell Biology, 131(5): 1125–??, December 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/131/5/1125.

Zhang:1995:ICC

[ZN95b] D. Zhang and R. B. Nicklas. The impact of chromosomes and centrosomes on spindle assembly as observed in living cells. Journal of Cell Biology, 129(5):1287–??, June 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/129/5/1287.

Zhu:1996:ADC

[ZOB^{+}96] X. Zhu, M. Ohtsubo, R. M. Böhmer, J. M. Roberts, and R. K. Assoian. Adhesion-dependent cell cycle progression linked to the expression of cyclin D1, activation of cyclin E-cdk2, and phosphorylation of the retinoblastoma protein. Journal of Cell Biology, 133(2):391–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/2/391.

Zhu:1999:TIP

[ZOKS99] Yong Zhu, Anush Oganesian, Douglas R. Keene, and Linda J. Sandell. Type IIA Procollagen Containing the Cysteine-rich Amino Propeptide Is Deposited in the Extracellular Matrix of Prechondrogenic Tissue and Binds to TGF-β1 and BMP-2. Journal of Cell Biology, 144(5):1069–??, March 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/144/5/1069.

Zitt:1997:ETC

[ZOS^{+}97] Christof Zitt, Alexander G. Obukhov, Carsten Strübing, Andrea Zobel, Frank Kalkbrenner, Andreas Lückhoff, and Günter Schultz. Expression of TRPC3 in Chinese Hamster Ovary Cells Results in Calcium-activated Cation Currents Not Related to Store Depletion. Journal of Cell Biology, 138(6): 1333–??, September 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/138/6/1333.

Zakharenko:1998:DAM

[ZP98] Stanislav Zakharenko and Sergey Popov. Dynamics of Axonal Microtubules Regulate the Topology of New Membrane Insertion into the Growing Neurites. Journal of Cell Biology, 143(4):1077–??, November 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/143/4/1077.

Zha:1998:STI

[ZPL^{+}98] Xiaohui Zha, Lynda M. Pierini, Philip L. Leopold, Paul J. Skiba, Ira Tabas, and Frederick R. Maxfield. Sphingomyelinase Treatment Induces ATP-independent Endocytosis. Journal of Cell Biology, 140(1):39–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/140/1/39.

Zeile:1996:RTC

[ZPS96] W. L. Zeile, D. L. Purich, and F. S. Southwick. Recognition of two classes of oligoproline sequences in profilin-mediated acceleration of actin-based Shigella motility. Journal of Cell Biology, 133(1):49–??, April 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/133/1/49.

Zerges:1998:LDM

[ZR98] William Zerges and Jean-David Rochaix. Low Density Membranes Are Associated with RNA-binding Proteins and Thylakoids in the Chloroplast of Chlamydomonas reinhardtii. Journal of Cell Biology, 140(1):101–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/140/1/101.

Zhao:1998:RCC

[ZRG98] Ji-He Zhao, Heinz Reiske, and Jun-Lin Guan. Regulation of the Cell Cycle by Focal Adhesion Kinase. Journal of Cell Biology, 143(7):1997–??, December 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/143/7/1997.

Zhang:1999:ICI

[ZRHO99] Jiandi Zhang, Mary C. Reedy, Yusuf A. Hannun, and Lina M. Obeid. Inhibition of Caspases Inhibits the Release of Apoptotic Bodies: Bcl-2 Inhibits the Initiation of Formation of Apoptotic Bodies in Chemotherapeutic Agent-induced Apoptosis. Journal of Cell Biology, 145(1):99–??, April 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/145/1/99.

Ziv:1998:IGC

[ZS98] Noam E. Ziv and Micha E. Spira. Induction of Growth Cone Formation by Transient and Localized Increases of Intracellular Proteolytic Activity. Journal of Cell Biology, 140(1): 223–??, January 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/140/1/223.

Zhang:1999:NRA

[ZSB99] H. L. Zhang, R. H. Singer, and G. J. Bassell. Neurotrophin Regulation of β-Actin mRNA and Protein Localization within Growth Cones. Journal of Cell Biology, 147(1):59–??, October 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/147/
1/59.

Zheng:1995:KGF

[ZSMV95] J. Zheng, O. Saksela, S. Matikainen, and A. Vaheri. Keratinocyte growth factor is a bifunctional regulator of HPV16 DNA-immortalized cervical epithelial cells. Journal of Cell Biology, 129(3):843–??, May 1995. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/129/3/843.

Zacchi:1998:RRM

[ZSP^{+}98] Paola Zacchi, Harald Stenmark, Robert G. Parton, Donata Orioli, Filip Lim, Angelika Giner, Ira Mellman, Marino Zerial, and Carol Murphy. Rab17 Regulates Membrane Trafficking through Apical Recycling Endosomes in Polarized Epithelial Cells. Journal of Cell Biology, 140(5):1039–??, March 1998. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/140/
5/1039.

Zarnescu:1999:ASE

[ZT99] Daniela C. Zarnescu and Graham H. Thomas. Apical spectrin is essential for epithelial morphogenesis but not apicobasal polarity in Drosophila. Journal of Cell Biology, 146(5):1075– ??, September 1999. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/146/5/1075.

Zechner:1997:RPM

[ZTH^{+}97] Dietmar Zechner, Donna J. Thuerauf, Deanna S. Hanford, Patrick M. McDonough, and Christopher C. Glembotski. A Role for the p38 Mitogen-activated Protein Kinase Pathway in Myocardial Cell Growth, Sarcomeric Organization, and Cardiac-specific Gene Expression. Journal of Cell Biology, 139 (1):115–??, October 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/1/115.

Zhang:1997:MPR

[ZWT^{+}97] Tao Zhang, Siew Heng Wong, Bor Luen Tang, Yue Xu, Frank Peter, V. Nathan Subramaniam, and Wanjin Hong. The mammalian protein (rbet1) homologous to yeast bet1p is primarily associated with the pre-Golgi intermediate compartment and is involved in vesicular transport from the endoplasmic reticulum to the Golgi apparatus. Journal of Cell Biology, 139(5): 1157–??, December 1997. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/139/5/1157.

Zinkl:1996:AAI

[ZZvdB^{+}96] G. M. Zinkl, A. Zuk, P. van der Bijl, G. van Meer, and K. S. Matlin. An antiglycolipid antibody inhibits Madin–Darby canine kidney cell adhesion to laminin and interferes with basolateral polarization and tight junction formation. Journal of Cell Biology, 133(3):695–??, May 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:/

/jcb.rupress.org/content/133/3/695.

Zhang:1996:LHR

[ZZYG96] G. Zhang, M. L. Zapp, G. Yan, and M. R. Green. Localization of HIV-1 RNA in mammalian nuclei. Journal of Cell Biology, 135(1):9–??, October 1996. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/135/1/9.